A Complete Bibliography of Publications in the Journal of Cell Biology: 2005–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

24 May 2021 Version 1.02

Title word cross-reference

1 [BRY^{+}09, DPK^{+}05, GSiK^{+}09, SWH^{+}06, Wil09a]. 2 [Wil09a]. 23 [MRD^{+}08]. 3 [AWM^{+}08, CCL^{+}08, FKML^{+}06, GdTAR07, HSTM06, LFL^{+}05, MCF^{+}05, MES^{+}08, MAK^{+}06, Wil09a, XSVGR05]. 4 [BSA^{+}08, CHCB05, GMM^{+}08, LPA^{+}07, PWW^{+}09, SHDT05, WSB^{+}09]. 5 [CYC^{+}06, LPA^{+}07]. ^{+} [CULS^{+}07, FKN^{+}07, HFP^{+}05, HMR^{+}07, KWS^{+}08, Ras08a, Rob08o, SSD^{+}09, pTyLZ^{+}09]. ^{−/−} [GHDS^{+}07, THN^{+}06]. ^{2+} [BSMT05, CZD^{+}09, CLK^{+}07, CSB^{+}06, FM05, GRC^{+}05, HJ06, HSK^{+}05, HPH^{+}05, ITG^{+}08, JBL^{+}07, LeB06-57, LeB06-75, LWG^{+}05, LWBL06, LGG^{+}05, Mel08a, Mua05, NJS^{+}06, OFY^{+}05, RDY^{+}05, RMP06, SFEK07, SBV^{+}06, pTyLZ^{+}09, WBK^{+}06, Wel05-44, WNF^{+}05b, WBLL06, XLH^{+}07, ZDMD06]. ^{307} [NEF^{+}06]. ^{ARF} [NCB^{+}08, NCB^{+}09, MZH^{+}08]. ^{βTrCP} [FBJB^{+}09]. ^{Cdc55} [CRLP07]. ^{Cdh1} [FBJB^{+}09, KST^{+}07, MVC^{+}07]. ^{fyn} [BLS^{+}05, SSLS05]. ^{Glued} [JRZ^{+}07, LSC^{+}06a, RKvdK^{+}09]. ^{Ink4a} [YTH^{+}09]. ^{S} [ASK^{+}09, CWM^{+}09, Wil09k]. ^{Shc} [MMW^{+}07]. ^{Slimb} [RRR^{+}09]. _{1} [TDP^{+}06]. 2 [BWJ05, BHG^{+}05, DTW06, GC05, IPAR05, KNKH06, MPS^{+}07, MM05, SMH^{+}09, SCK^{+}07, VBSM06, Wil07-47, vRSvR^{+}07]. _{3} [FLA^{+}08, HHFC06, MHM^{+}05, SKO^{+}06, SFEK07, TDT^{+}08]. _{6} [SMWdV09]. CS [UWiY^{+}07]. _{EL} [XJF05]. _{iα} [CK07]. _{L} [BbCQ^{+}09, KPM05]. _{o} [BKE^{+}07]. S [WPH^{+}07, FNH^{+}05]. _{V} [TGRD09, CZD^{+}09, LPB^{+}08, XLH^{+}07]. α [AXM^{+}05, BZHA^{+}08, BBM^{+}08, BLS^{+}05, BSL^{+}08, BBG^{+}05, CPAH08, CNM^{+}05, CCL^{+}08, CZD^{+}09, CSM^{+}06, CC09, CSL^{+}05, DNG^{+}06, DVF^{+}06, ERM^{+}09b, EVH^{+}08, FHNG05, FHO^{+}05, FCDM^{+}07, GHX^{+}07, GPN^{+}06, GGMBF08, GPC^{+}06, GCS^{+}07, GBNS07, HOM^{+}05, HLL^{+}08, HCRT08, JTN^{+}05, KBY^{+}08, KKW^{+}09, KBC^{+}07b, LLH^{+}09, LMC^{+}09, LGV08, LLC^{+}08, LCL^{+}05, MWMS07, MKH^{+}07, MYZ^{+}09, MBL^{+}07, NEF^{+}06, NSM^{+}09, NRJ^{+}09, OMF^{+}05, PDM^{+}08, PHB^{+}07, PMK^{+}05, PZG^{+}07, RSL06, RM05b, SSD^{+}09, SOR06, SCM^{+}07, SLZ^{+}06, SGKP^{+}07a, SGKP^{+}07b, SFS^{+}07, TKB^{+}06, TPZG06, Tum05-71, UFY^{+}07, VCC^{+}08, WBC^{+}06, WWX^{+}06, WBP^{+}08, WCR^{+}05, WCN07, XMA^{+}09, YM09, YLM^{+}06b, ZKR^{+}07, vKST^{+}07]. α_{1} [LYV^{+}08, NRW^{+}08]. α_{12/13} [LSC^{+}05]. α_{12/13} [LCS^{+}09]. α_{4} [AFM^{+}05]. α_{4}β_{1} [AFM^{+}05, EVH^{+}07]. α_{5}β_{1} [BMR^{+}08]. α_{e} [EFG^{+}05]. α_{i} [GRC^{+}05]. α_{IIb} [SZX08]. β [AXM^{+}05, AJR^{+}09, AM05b, ADC^{+}07, ABY^{+}08, BRKR06, BFG^{+}06b, BCT06, BMRP07, CTGVGFF08, CNM^{+}05, CCL^{+}08, CCM^{+}07, CTP^{+}09, CSL^{+}05, CML^{+}06, DRD^{+}05, EVH^{+}08, FHNG05, FFW^{+}07, qGyWX^{+}08, GCSG^{+}05, GG06, GDA^{+}08, HWN^{+}07, HOM^{+}05, HJB07, HZL^{+}07, HFvdV^{+}05, JGAS08, JRZ^{+}07, JKW^{+}09, JTN^{+}05, JLG^{+}06, KBY^{+}08, KKW^{+}09, KKB^{+}06, KTI^{+}05, KKES05, KKRP05, KLV^{+}08, KLG^{+}09, LD07, LeB05-75, LeB05-69, LZL^{+}05, LPN^{+}08, Les07-38, Les09s, LMHT08, LPF^{+}06, LCB^{+}08, LLC^{+}08, MFS^{+}09, MCA^{+}06, MSM^{+}07, Miz05, MM05, NLC^{+}05, NPV^{+}07, NZQ^{+}07, NWR^{+}08, PCR^{+}07, PCG^{+}05, PAV^{+}06, dJPAW^{+}09, PST^{+}06, RHMW^{+}05, RKN^{+}07, RM05b, SBC^{+}05, SMH^{+}09, SCD^{+}08, SHX^{+}07, Sho09z, Sho09-52, SHjC^{+}07, SNG^{+}05, SLZ^{+}06, SLK^{+}08, SOC^{+}07, TTO^{+}08, TKB^{+}06, TVP^{+}06, Tum05d, WBP^{+}08, WCR^{+}05, WCN07, WRMH07, WWS05b, WWC^{+}07, XMA^{+}09, XYC^{+}07, YM06, YMD^{+}07, YOHR07, YCK^{+}05, YLM^{+}06b, ZKR^{+}07, tKJCH06]. β_{1} [MFBG^{+}08, OKN06, SDN^{+}05b]. β_{2} [WPL^{+}06, XLH^{+}07]. β_{3} [FMM^{+}08, MQWP08, SZX08]. β_{4} [XLH^{+}07]. β_{e} [EFG^{+}05]. δ [BTD^{+}07, BLT09b, CTP^{+}09, MSM^{+}07, Sho09-52]. ϵ [KBW^{+}06, NYI^{+}05b]. η [AM05b]. γ [AM05b, BDS^{+}08, BVS^{+}09, DMDO08, GLS^{+}09, HRB^{+}06, IDTT^{+}09, JSPT05, KKC^{+}07, KKD^{+}07, KTA^{+}06, LVL^{+}05, LBS^{+}08, LHRP05, LBC^{+}07, MYZ^{+}09, MDET08, NEF^{+}06, PGO^{+}07, SDDV05, STG05, Sho09a, SRD^{+}07, SZX08, SLWA07, SFEK07, TMAL^{+}06, Tum06-27, VCD^{+}06, ZCJ^{+}08, vKST^{+}07]. κ [BWL^{+}08, BAR^{+}05, CWC^{+}07, GBNS07, KBW^{+}06, LDVV06, LHM^{+}09, OSB^{+}05, PGP^{+}07, RRA^{+}09, SLZ^{+}06, STL^{+}07, TCW^{+}06, Tum05t, ZTKK05]. μ [GSW^{+}08, HGJ06, LBC^{+}07]. ζ [BFP^{+}07, EMMN^{+}05, KNT^{+}07, MWY^{+}09, MDW05].

-3 [VKR^{+}09]. -4 [CYO^{+}09]. -8 [YBT^{+}05]. -AAA [ERM^{+}09a]. -actin [MCA^{+}06, NRJ^{+}09]. -adrenoceptor [SHX^{+}07]. -amyloid [KKB^{+}06, YCK^{+}05]. -antitrypsin [NRW^{+}08]. -arrestins [PCR^{+}07]. -barrel [HWN^{+}07]. -binding [GG06]. -Catenin [RHMW^{+}05, BRKR06, qGyWX^{+}08, GCSG^{+}05, HFvdV^{+}05, KKW^{+}09, LeB05-75, LPN^{+}08, Les09s, LMHT08, LCB^{+}08, MFS^{+}09, Tum05d, dJPAW^{+}09]. -conjugated [MAK^{+}06]. -cytoplasmic [AFM^{+}05]. -dependent [AFM^{+}05, GSiK^{+}09, GLS^{+}09, MWY^{+}09, BSMT05, BRY^{+}09, GC05, ITG^{+}08, LGG^{+}05, WWX^{+}06]. -dystroglycan [RKN^{+}07]. -EF1 [BLT09b]. -H2AX [KKD^{+}07, LBS^{+}08]. -helical [SGKP^{+}07a, SGKP^{+}07b]. -Herpesvirus [DNG^{+}06]. -independent [FHNG05, LRS05, SSHW09]. -induced [MDET08, NEF^{+}06, HLL^{+}08, LD07, PDM^{+}08, WPH^{+}07]. -inducible [LLH^{+}09]. -integrin [MFBG^{+}08]. -kinase [BSA^{+}08, CHCB05, LPA^{+}07, LFL^{+}05, MCF^{+}05, PWW^{+}09, SHDT05, WSB^{+}09]. -lamin [PHB^{+}07]. -mediated [FHO^{+}05, BSL^{+}08, LMC^{+}09, FBJB^{+}09, OFY^{+}05]. -MKK4-JNK [SLZ^{+}06]. -p53 [MZH^{+}08, NCB^{+}08, NCB^{+}09]. -phosphate [AWM^{+}08, LPA^{+}07, MES^{+}08, SWH^{+}06]. -Pix [tKJCH06]. -related [PRB^{+}09]. -rich [MPS^{+}07]. -secretase [KKC^{+}07, SRD^{+}07]. -secretase-mediated [IDTT^{+}09]. -site [PCG^{+}05]. -smooth [GPC^{+}06]. -spectrin [HJB07]. -Syntrophin [LCL^{+}05]. -synuclein [FCDM^{+}07]. -terminal [BWJ05, KNKH06, MM05, VBSM06]. -TrCP-dependent [SCD^{+}08]. -triggered [HJ06]. -Tubulin [BVS^{+}09, HRB^{+}06, JSPT05, STG05, Tum06-27, VCD^{+}06, ZCJ^{+}08]. -tubulin-mediated [DMDO08]. -TuRC [Sho09a, VCD^{+}06].

/BMP [XYC^{+}07]. /BMP-activated [XYC^{+}07]. /Ca [pTyLZ^{+}09]. /calmodulin [FM05]. /g [RSS^{+}09]. /H [SSD^{+}09]. /NF [GBNS07]. /NF- [GBNS07]. /NSF [JWSS05]. /NSF-independent [JWSS05].

1 [AHM^{+}05, BCJ^{+}07, BAR^{+}08, BMB^{+}07, CHSV07, CPAH08, CNM^{+}05, CCM^{+}07, CMF08, CTP^{+}09, CBIO05, DPD^{+}09, DXC^{+}09, DP06, DHT^{+}09, DPMB^{+}07, DGP^{+}05, EPV^{+}05, FHNG05, FVSC08, GG06, GEH^{+}07, GNPC^{+}08, GGEd^{+}07, HAAHD05, HCB^{+}08, IUT^{+}07, KZC^{+}08, KHW^{+}07, KHW^{+}09, KKRP05, KDD^{+}09, LCO07, LeB06-49, LDGK08, LCvD^{+}08, LLH^{+}09, LCT^{+}09, LSdWL07, LFL^{+}05, MLV09, MFBG^{+}08, MDH^{+}06, NEF^{+}06, NSM^{+}09, NKK^{+}06, OOS07, OOS09, OBY^{+}06, OLC^{+}07, PXD^{+}07, PMK^{+}05, PVM^{+}05, PST^{+}06, PZG^{+}07, SSS^{+}08a, Sed09g, Sho09-50, SCS^{+}05, SRD^{+}07, SiTT^{+}08, Tum05-52, WNL^{+}09, Wel06t, Wil07-41, Wil09n, WRMH07, WSFD07, XJF05, YWEK08, yYMM05, YMD^{+}07, vKST^{+}07, TSN^{+}06]. 1-dependent [KKW^{+}09]. 1-integrin [BMRP07, SHjC^{+}07]. 1-integrins [PAV^{+}06]. 1-mediated [HOM^{+}05]. 1-null [YMD^{+}07]. 1/ [GHSG07, KHW^{+}07, KHW^{+}09]. 1/clathrin [LRCH^{+}05]. 1/Mfn1 [RLDC09]. 10 [GEP08, WOWB08]. 10-SNARE [GEP08]. 108/ [SHL^{+}09]. 12 [LeB07-60, MSLS07, ZRH^{+}09]. 12R [EFM^{+}07]. 12R-lipoxygenase [EFM^{+}07]. 13s [TAM^{+}06]. 14 [DOVD08]. 14-3-3 [FHO^{+}05, KRZ^{+}08, LMHT08, STMG05]. 17/20 [YWW^{+}08]. 170 [BKLT^{+}08, DvCvH^{+}08, LHL^{+}08]. 19 [SHL^{+}09]. 1a [MT07]. 1b [WAW08, ASHD^{+}05, API^{+}07, FSP^{+}07, LBC^{+}07]. 1C [HK06, Pow06j]. 1f [RKN^{+}07]. 1R [XKYW07].

2 [BRY^{+}09, DXC^{+}08, DSBK09, FBY^{+}06, GGO^{+}09, GNPC^{+}08, LOG^{+}06, LPN^{+}08, LLHH07, MQWP08, MKB^{+}09, MDET08, MSSF05, MGH^{+}09, MKV^{+}09, NBB^{+}09, OOY^{+}06, OLLG05, OBY^{+}06, RLDC09, SBF^{+}06, SHL^{+}09, THN^{+}06, UDC^{+}07, WPWF08, WZC^{+}08, WPH^{+}07, WKC^{+}09, WF05, WF07, ZDMD06, WBC^{+}06]. 2- [AM05b, HGJ06, RLDC09]. 2/ [AM05b]. 2/ErbB [BHJ^{+}05]. 2/HspA2 [SMM^{+}08a]. 20/Sema2 [WZC^{+}08]. 2005 [LK06]. 2006 [Wel06o, WL07]. 2007 [LW08]. 2008 [dSW09]. 206 [KLS^{+}06, RGA^{+}06]. 20h [BGJ^{+}07]. 25K/ [SLK^{+}08]. 2A [PPK08]. 2C [KTA^{+}06]. 2J [HLJ^{+}07, JLB^{+}05].

3 [AXM^{+}05, CM06, FIK^{+}20, IEGV06, LeB06-53, LBI^{+}08, MXR^{+}05, OOY^{+}06, OFfL^{+}07, POCS^{+}06, RXA^{+}08, RP05, SBF^{+}06, SBMP05, dGSWW^{+}06, SWH^{+}06, WBP^{+}08, WLK^{+}05, XMA^{+}09, YLM^{+}06b, BFG^{+}06b, BHJ^{+}05]. 3-integrin [SMH^{+}09]. 3/Opa1 [RLDC09]. 35 [SBJ^{+}07]. 367a [Wil06d]. 3F3 [WF05, WF07]. 3F3/2 [WF05, WF07]. 3K [LCT^{+}09]. 3R [DRD^{+}05]. 3T3 [LLS^{+}09]. 3T3-L1 [LLS^{+}09]. 3’UTR [CTS^{+}05].

4 [BRM^{+}07, DMDO08, DGZ^{+}05, ERM^{+}09b]. 4.1 [KHW^{+}07, KHW^{+}09]. 40S [ZWO^{+}09]. 43 [LH05, RBC05]. 45 [LLH^{+}07]. 48th [dSW09]. 4D [OKN06]. 4D/Plexin [OKN06]. 4E [FBD^{+}05b]. 4GalNAcTB [JKW^{+}09]. 4K [PST^{+}06].

5 [FLA^{+}08, FIK^{+}05, LeB07-28, LCT^{+}09, MB07, RYD08, TSN^{+}07, Wil07v]. 5- [SDB^{+}05]. 5’-6’-epoxyeicosatrienoic [FLA^{+}08]. 5-bisphosphate [JKKM08, VTRB06, vZPG^{+}07]. 5-MCDE [LCT^{+}09]. 5-triphosphate [LMC^{+}09]. 5-trisphosphate [HPH^{+}05, MuMI^{+}06, dCD^{+}07]. 5/ [HOH^{+}07]. 53 [NRW^{+}08]. 53BP1 [BJLM^{+}05, MDWH^{+}08, TNS07]. 5P [DTW06].

6 [FLA^{+}08, DXC^{+}08, MKH^{+}07, MWSK06, SSM^{+}08, VSG09]. 6-Å [MWSK06]. 60S [KOVP09, KRPH08, LSD^{+}06].

7 [WGL^{+}06]. 70 [IKZ^{+}07, KFK^{+}05]. 70-kD [KFK^{+}05].

8 [FTB^{+}09, Sco08]. 84 [BMSJ07, JMG^{+}06].

9 [RLDC09, SD07]. 9-1-1 [YM09]. 9-Å [SD07]. 95 [NBS^{+}08, NFI^{+}09, Wil08r]. 98 [MQL^{+}06].

= [LeB05-107, Rob07g, Rob08-36].

A-type [AG09]. A/C [GNPC^{+}08, MDL^{+}09, Sed09f]. A/C-mediated [MDL^{+}09]. A2 [PWM^{+}06a, WGKA^{+}08, vKGK07]. A2-dependent [WGKA^{+}08]. A30P [FCDM^{+}07]. A7 [JGL^{+}06]. AAA [ADD^{+}06, ERM^{+}09a, GD05, HMR^{+}07, KRPH08]. Aberrant [DNK^{+}05, SS07]. ability [GSB^{+}06a]. Abl [HWJ^{+}08, NNS^{+}08]. Abl- [NNS^{+}08]. Ablation [NLC^{+}05, BvRJ^{+}05, HJK^{+}06]. Abnormal [LHS^{+}05]. abrogate [BvRJ^{+}05]. abrogated [HOH^{+}07]. Absence [YMD^{+}07, DG05, HYB08, KYK^{+}06, MKB^{+}09, OLH^{+}08, SIL^{+}07, SRXvdK06, WSS^{+}06]. Abundance [Rob08a, MBF^{+}07, MBF^{+}08]. abundant [KOKM07]. Abuzz [Sed08b]. ACA [DKR^{+}06]. ACAP1 [LPB^{+}07]. ACAP1-containing [LPB^{+}07]. accelerate [KR07]. accelerates [SMM^{+}08a]. access [BZA^{+}06, Ros05, Wil08k]. accessible [CDW^{+}05]. accessories [Mat06]. accompanies [DNK^{+}05]. accomplices [Sed07g]. according [SLP^{+}06, ZDMD06]. accumulate [WBLL06]. accumulates [MAG^{+}08]. Accumulation [NRD^{+}06, JMM^{+}06, KP08, NNK^{+}08, RMKS08, WZCB^{+}06, ZKG^{+}08]. Accuracy [Wat09]. acentrosomal [CBIO05, MFC^{+}09, RYN05]. Acetylate [LeB06a]. Acetylation [LWC09, iHBD08, KCD^{+}06a]. acetylcholine [HCRT08, KBC^{+}08]. acetyltransferase [HCK^{+}07]. achieve [Les07-30]. Achlorhydria [TKH^{+}05]. acid [BPL^{+}05, BW08, FLA^{+}08, KGL07, ML05, PRZ^{+}08, RCM^{+}09, SKM^{+}08, SB09, ZJH08]. acids [WHR^{+}09, WEL^{+}07n]. acinar [HOH^{+}08]. acini [PH07]. acquisition [Nor06]. Acrobatic [LeB07a]. across [AACS^{+}06, BYMB06, Les08-29, Ros08c, WC07a]. act [BYK08, GNL^{+}06, JNHB08, KGKV07, Les07-79, Sho09f, Ter06, WPWF08, WHB^{+}07]. Actin [Bon05a, Bon05b, CCL^{+}05, DOS^{+}07, KCMB08, LeB05a, LeB07b, LSM^{+}05, Les08a, MTH^{+}06, Pow05a, Pow05b, Rob08b, SSK^{+}09, THL^{+}06, Tum05-77, Tum06a, Wil07a, YWJ^{+}07, vBDF^{+}09, ACsHS08, BRH05, BMS^{+}08, Bon05g, BKBM06, CM07, CAR^{+}08, CHCB05, CHCM06, FBD^{+}08, GSJ^{+}08, GMG^{+}07, GPC^{+}06, GFD^{+}05, HJK^{+}06, HdL06, HAN^{+}09, HL06, HLS^{+}09, HLP06, HWS^{+}07, KRZ^{+}08, KOM07, KOKM07, KCOS06, LeB05b, LeB06-68, LeB07-38, LeB08-30, LeB08-36, LBI^{+}08, LSNY08, Les07d, Les09g, LHL^{+}08, LCS^{+}09, MT07, MCA^{+}06, MMS07, NAS08, NRJ^{+}09, OD07, POGL08, Ped08, PV06, QHB^{+}07, RKH^{+}09, RXA^{+}08, RKK09, SHS^{+}09, SRN^{+}07, SDB^{+}05, SOR06, SOR^{+}05, STU^{+}08, SBMP05, SS09, SKO^{+}06, SCK^{+}07, SMG07, TMI07, TSALS05, TSP08, TSS^{+}06, Tum05-31, Tum05-41, VLFA^{+}07, Wel06-39, Wel06-41, Wil08a, WOWB08, YHJ^{+}05, ZHI^{+}08, Wil07b, NHK^{+}08, SJL^{+}07]. Actin- [LSM^{+}05, KCOS06]. actin-based [VLFA^{+}07]. Actin-dependent [DOS^{+}07, SCK^{+}07]. actin-depolymerizing [POGL08]. actin-mediated [OD07]. Actin-myosin [CCL^{+}05, Tum05-77]. acting [MB07]. Action [vDVM^{+}09, BPL^{+}08, CFBM05, HKIY09, HGK^{+}06, Kat06d, Les06a, LWL08a, LWL08b, POCS^{+}06, RvVM^{+}08, Wel06-69, Wil07-57]. actions [HWJ^{+}08, Wil08r]. activate [ALB^{+}09, BCJ^{+}07, IH05, JTN^{+}05, LNG^{+}05, QU08, SKE^{+}05, SSLS05, SMM08b]. activated [CCC^{+}05, FBY^{+}06, KCOS06, LeB05-31, LLL^{+}06, MDET08, MBF^{+}07, MBF^{+}08, OSS^{+}07, RG06a, WMT07, XYC^{+}07, YCK^{+}05]. activates [BKHR05, BCT06, CGLC^{+}08, CC09, EYB^{+}06, FBY^{+}06, KPM05, KKRP05, NZQ^{+}07, SS07, WRMH07, YMCT08, vRSvR^{+}07]. Activating [Tum06b, CKG^{+}08, GSH^{+}08, LKL^{+}05, MAH^{+}05, NMO^{+}07, ZDD^{+}06, FHS^{+}07a, KBN^{+}06, TGH^{+}07]. Activation [GGMBF08, JRZ^{+}07, MGBH^{+}06, RP05, TKB^{+}06, TSN^{+}07, Tom08, WGKA^{+}08, AWR^{+}07, BAR^{+}05, BLS^{+}05, CLK^{+}07, CML^{+}06, DMBF^{+}07, DPL^{+}08, DRD^{+}05, DP07, ERM^{+}09b, FHNG05, FvD08, FPP^{+}08, GGO^{+}09, GCF^{+}05, GG06, GPJ^{+}06, HOH^{+}08, HSB^{+}07, KSK^{+}06, LDVV06, LD07, LeB07-41, LC08b, LWC09, LLJ^{+}05, LSC^{+}07, LYWK08, LTDH05, LWBL06, LLC^{+}08, MKB^{+}09, MRO^{+}05, MLHS06, NRD^{+}06, NNS^{+}08, OYKY^{+}05, OSB^{+}05, OSS^{+}07, OTM^{+}07, PGP^{+}07, PG06, PhXD^{+}06, PRK^{+}07, PRK^{+}09, PCPW08, PFGSV06, ROKG08, RMKS08, SBM^{+}07, SYKM08, SSD06, SLV^{+}07, SFS^{+}07, TTYH09, pTyLZ^{+}09, VLK^{+}08, VDA^{+}08, VMKW^{+}08, Wel06-56, XJF05, XNM^{+}09, YKH^{+}08, YMD^{+}07, YSB^{+}07, Yu06, YLM^{+}06b, ZDMD06, dCD^{+}07, tKJCH06]. Activation-dependent [MGBH^{+}06]. activator [FNH^{+}05, HOM^{+}05, KRP^{+}09, KNT^{+}06, MQWP08, MRO^{+}05, YOSM06]. Active [Les06b, PNO^{+}09, BBvD05, BB05, BLR^{+}06, BGdN^{+}08, CSSvM08, DOS^{+}07, FOW^{+}09, HFvdV^{+}05, LeB05-75, LeB08x, MCD^{+}07, PFD06, SPS^{+}08, SWB^{+}07, XC08]. activities [FMD^{+}09, OYM^{+}09, YN07]. Activity [CTS^{+}06a, LRS05, MFS^{+}09, AFV^{+}05, ASK^{+}09, BMS^{+}08, BFP^{+}07, BSL^{+}08, CBK^{+}06, CZD^{+}09, CSMP^{+}09, CG06, CLK^{+}07, CWM^{+}09, Dal09, DGSN06, DS07, ERM^{+}09b, EFG^{+}05, FKN^{+}07, GNPC^{+}08, GDA^{+}08, GHK^{+}09, HBK^{+}06, HOH^{+}07, HCK^{+}07, HF06, IDTT^{+}09, JLL^{+}07, KBY^{+}08, KTI^{+}05, LPWP^{+}07, LeB06z, Les06f, LMHT08, LWHC08, LYY08, LMC^{+}09, LSdWL07, LLT^{+}08, LCS^{+}09, MYO^{+}05, MS09, MM08, MGR^{+}05, NLL^{+}06, NSM^{+}09, NTY^{+}05, NFI^{+}09, OKN06, PPBF08, PEW^{+}05, RKT^{+}07, Sed08e, SOR06, SDD^{+}07, SIL^{+}07, SKR^{+}07, SSK^{+}09, SLNS06, TST08, TMI07, ToMY^{+}08, VWK^{+}06, WMM^{+}09, WMS^{+}06, WWO^{+}08, WL09a, WSB^{+}06, WBG^{+}08, WKC^{+}09, WVB07, XOV^{+}09, YMC05, ZPGB^{+}07, ZKG^{+}08]. Activity-dependent [LRS05, SLNS06]. activity-independent [MM08]. activity-sensitive [NFI^{+}09]. acto [IGFG^{+}08]. acto/myosin [IGFG^{+}08]. actomyosin [HYB08, STK^{+}05, dRKD^{+}05]. acts [BIH^{+}05, BIH^{+}07, CSM^{+}06, DTW06, HEW07, JMB^{+}07, LR05, PKK09, Rou08, RRA^{+}09]. acutely [RCM^{+}09]. Acyl [LeB05c, FHFB^{+}06]. acylated [CNMM^{+}06]. acylglycerol [BPL^{+}05]. acyltransferase [SB09, SSSL05]. ADAM12 [ADC^{+}07]. ADAM17 [WPC^{+}08]. Adaptation [BXN^{+}06]. adapted [DSH^{+}07]. adapter [SSSB07, yYMM05]. adaptin [GSW^{+}08, LBC^{+}07]. adaptive [MFBG^{+}08]. Adaptor [Tum05a, FN06, GHX^{+}07, JWG^{+}07, MDW05, Wel06j]. ADAR2 [VBL^{+}05].

ADAR2-mediated [VBL^{+}05]. addition [DMDO08]. adds [Sho08]. Adenovirus [CC09, HAT^{+}05b, SCM^{+}07]. adenylation [HOZ^{+}09]. adenylyl [TDT^{+}08]. Aderem [LeB08b]. ADF [WHB^{+}07, vBDF^{+}09]. ADF/cofilin [WHB^{+}07, vBDF^{+}09]. Adherens [Gra07, DP08, HT08, IUT^{+}07, LBC^{+}07, MMH09, SHS^{+}09, WL09a, WL09b, Wel06-68]. adherent [BZYM^{+}08, Wil07-64]. Adhering [Sho09b]. adhesin [OHH^{+}06]. adhesion [AS06, AFM^{+}05, BZYM^{+}08, BMR^{+}08, BMK^{+}07, CCH09, CG06, EBMG09, GLBS06, GFH^{+}06, GJL^{+}08, GPC^{+}06, GPJ^{+}06, HWS^{+}07, JS07, JLM^{+}07, JLL^{+}07, KTN07, LMM^{+}09, LeB05a, LeB05-77, LeB05-108, LeB08a, LSdWL07, LLT^{+}08, LYK^{+}06, LLC^{+}08, MFA^{+}07, MFS^{+}09, MRKH05, MKV^{+}09, NWB^{+}06, NN07a, OFY^{+}05, OPT^{+}07, PR05, PAV^{+}06, PT07, PLR^{+}06, QCGM05, RHMW^{+}05, SMSPK08, SRN^{+}07, SPA^{+}05, SWI06, TSP08, Tum05-74, VMZW^{+}07a, VMZW^{+}07b, Wel07j, Wil08p, WKIKG09, YN07, ZHI^{+}08]. adhesion-dependent [GPJ^{+}06, LMM^{+}09]. adhesions [GSJ^{+}08, Les09o, TSN^{+}06, Tum06-36, WKWL06, YHJ^{+}05]. adhesive [LZ08]. adhesome [ZB09]. adipocyte [SQB^{+}08]. adipocytes [WP08]. adipogenesis [Mel08a, Sho09-46]. adipose [LMZ^{+}05]. Adjacent [TGH^{+}07]. ADP [COB^{+}08, LDGK08, WDWC08]. ADP/ATP [COB^{+}08]. adrenoceptor [SHX^{+}07]. adrenoleukodystrophy [GDF^{+}05]. adrenoleukodystrophy-related [GDF^{+}05]. adult [BP05, DCAC05, JAV^{+}07, KCS^{+}06, LRS05, MNF^{+}07, NSS^{+}08, RMZ^{+}06, SDL^{+}05, VMY^{+}09, VFRJ^{+}07, WLLF^{+}08, WPDZ^{+}06]. advance [Wel06-51]. advances [YKH^{+}08]. advantage [LeB07p]. aeruginosa [KGDR^{+}07]. afadin [SHS^{+}09]. affect [DDH09, MLS^{+}07, SLCB05]. affecting [BMRP07, FMA^{+}09, PTF^{+}06]. affects [CFH^{+}08, EHHK09, GTB^{+}07, MBS^{+}08, NBM^{+}08, OMGW^{+}06, PLH^{+}08, VTT^{+}09b, Wel05n]. affinity [FHNG05, Kat06b, MFBG^{+}08, TMH^{+}06, WWS05b]. affinity-dependent [FHNG05]. after [DBK^{+}07, EMPK^{+}07, HFY^{+}06, LeB08x, Rob07n, ULN^{+}07, Wel06-46]. again [Pow09a]. against [SBC^{+}07, Wil07l]. age [Les07-79, MWY^{+}09, Wel05-61]. age-dependent [MWY^{+}09]. Aged [Tum06c]. agent [OF08]. agent-based [OF08]. agents [DDO^{+}08, MCD^{+}07]. aggregate [PDM^{+}08]. aggregates [BLB^{+}05, CSSvM08, FSR^{+}07, LeB05t, LeB06-32, MSH^{+}05, NSS^{+}08, WZM^{+}08, YCR06]. aggregation [KL05, LSC^{+}06a, Wel05w]. aggresomes [OLC^{+}07]. Aging [Kat06a, BJM^{+}06, EVF^{+}08, MG08b, Tum06d, WBK^{+}06]. agonist [WBP^{+}08, YLM^{+}06b]. agonist-induced [WBP^{+}08, YLM^{+}06b]. Agrin [MAG05, Tum05b]. ahead [Les07v, Rob07l, Wel05y]. aid [Wel06x]. aided [MZ08]. AIDS [Wel06o, AGL^{+}08]. ain’t [Wil08c]. Aip1 [BKBM06, KCMB08]. airway [VTT^{+}09b]. AIS [Les07a]. Ajuba [PEW^{+}05]. Aki1 [Les09a, NAF09]. Akt [BBL^{+}08, CZD^{+}09, DDO^{+}08, CTS^{+}06a, CR06, CTO^{+}08, KRZ^{+}08, LCT^{+}09, NBM^{+}08, PhXD^{+}06, STMG05, YGW06]. AKT-mediated [CTS^{+}06a, STMG05]. Akt1 [IPG^{+}05, SL05]. Akt2 [IPG^{+}05, YOS^{+}05]. Alan [LeB08b]. alarm [LeB08f]. Albatross [SIS^{+}08]. Alberts [Wil08h]. Alexander [Sho09c]. Alien [LeB06b]. alignment [MFC^{+}09, VMSV08, ZAM07a, ZAM07b]. Alison [Sed08a]. alive [Tum05-72, Wel05-34]. alleles [AMERT^{+}07]. alleviates [SPT^{+}09, SPT^{+}21]. allosteric [KVS^{+}09]. allow [GSHD08, LCO07]. allowed [Les09z]. allows [KKM^{+}08]. along [CGS05, ÇB09, LSM^{+}05, MEG^{+}07, MT07, Sch08]. alphaviruses [ZMGC05]. ALT [DMBF^{+}07]. alteration [FWC^{+}07, KCOS06]. alterations [SNA^{+}07]. altered [CMK06]. Alternative [SGKP^{+}07a, SGKP^{+}07b, GHY^{+}05, JZN^{+}09, Les07-36, PAM^{+}07, SDI^{+}06, Wil08b]. alters [CCH09, GFC^{+}08, LZL^{+}05, LLH^{+}09]. alveolar [BIH^{+}05, BIH^{+}07]. always [Wil08c]. Alzheimer [HZL^{+}07, LeB07-55, Les07u, Les08b, PCG^{+}05, Rob07g, Sed07a, YCK^{+}05]. ameliorate [MBL^{+}07]. American [LP05, LK06, LW08, WL07, dSW09]. amino [KGL07, WHR^{+}09, WEL^{+}07n]. amiodarone [PKM^{+}05]. amiodarone-induced [PKM^{+}05]. AML [BDS^{+}08]. AML-associated [BDS^{+}08]. ammonia [VP05]. amoeboid [SSHW09]. AMP [BAdSG05, OFY^{+}05, VKR^{+}09]. AMPA [FTT^{+}05]. AMPAR [Tum05u]. amphibians [AS09]. amphiphysin [LWL08a, LWL08b]. AMPK [MSK^{+}07]. amplification [AAG^{+}09, SSAD08, WCR^{+}08, ZCJ^{+}08]. amplifies [HPT^{+}07]. Amy [Sed08b]. amyloid [GDA^{+}08, HZL^{+}07, KKB^{+}06, MM05, SKO^{+}08, YCK^{+}05]. amyloid- [MM05]. amyotrophic [KR07, MSH^{+}05]. Ana3 [SDW^{+}09]. Analogue [Wel06a]. analyses [STG05]. Analysis [TS06, YMTC05, BFWH07, BMBD^{+}09, CCSBM^{+}07, CML^{+}06, DZW^{+}08, EMB^{+}08, EMD^{+}08, GBNK08a, GBNK08b, HQF^{+}07, LJS^{+}09b, OIFN^{+}07, PALW05, PGL^{+}09, RM05b, SKM^{+}08, SRBL06, Wil07-45, WKIKG09, WNF^{+}05b, YCX^{+}08]. anaphase [CNW^{+}05, CCSBM^{+}07, CRLP07, CGGT09, GHM^{+}08, KLRS07, LeB07f, MTRJA05, MPL^{+}07a, MOS^{+}09, MHFH06, QU08, RW05, TST08, YKH^{+}08]. anaphase-promoting [CNW^{+}05]. anchor [CYO^{+}09, DDB^{+}05, LeB07-64, NS07, Tum05-29]. anchorage [BCT06, CTY^{+}06, DRD^{+}05, DBR09]. anchorage-dependent [DRD^{+}05]. anchorage-independent [BCT06, DBR09]. Anchored [Pow06a, BLC^{+}09, CTY^{+}06, HIK^{+}08, Nic09a, PPCZ06, Ras08a, SFS^{+}07, CZM^{+}07, NS06, SS06a, SFEK07]. Anchoring [Tum05c, BMSJ07, MKSJ06, ZHYE08]. anchors [NS06]. Andrea [Wil08b]. Androgen [EIK^{+}05, vRCB^{+}07]. Androgen-dependent [EIK^{+}05]. anemia [MLH^{+}06]. Aneuploidy [Tum06d, TC08, WC09]. Ang [CNM^{+}05]. Ang-1 [CNM^{+}05]. angiogenesis [FWC^{+}07, Les08-59, dJPAW^{+}09, SPA^{+}05, WLLF^{+}08]. angiogenic [FMM^{+}08, SSS^{+}08a]. Angiogenin [YIfHA09]. angiopoietin [SKE^{+}05]. angle [PFGSV06]. anillin [CWRG^{+}08, HO08]. anillin-related [CWRG^{+}08]. animals [HJB07, Mat05]. ankyrin [BFB^{+}08, DBDD06, DZG^{+}07, MPD09, LPB^{+}08]. AnkyrinG [HOR08, YOHR07]. annealing [ÇB09]. Announcing [Hil08]. Annual [dSW09]. anoikis [MMW^{+}07]. answers [Sho09e, VW08]. antagonises [MBdR^{+}08]. Antagonism [YBT^{+}09, RSS^{+}09]. antagonist [FKS^{+}06, YTY^{+}07]. Antagonistic [YR09, GHN^{+}09]. antagonistically [TEIS09]. antagonize [KZH^{+}07]. antagonizes [MAK^{+}06, STMG05, WL09a]. antagonizing [NLL^{+}06, SLV^{+}07]. antenna [Sch08]. anterior [TA07]. anterior-posterior [TA07]. anterograde [CGS05]. Anthony [Wil07c]. anthrax [ALvdG06]. anti [BLA^{+}09, FMD^{+}09, WPH^{+}07]. anti-apoptotic [WPH^{+}07]. anti-recombinase [BLA^{+}09, FMD^{+}09]. antiangiogenic [PWM^{+}06a]. antiapoptotic [GGO^{+}09]. antibiotic [Wil06o]. antibody [WZM^{+}08]. antigen [CBG^{+}06, KNKH06, LGH^{+}05, LeB07-30, MSM^{+}06, NS07, VLFA^{+}07]. antigen-stimulated [LGH^{+}05]. antigens [LeB05-79, LeB07-33]. antitrypsin [NRW^{+}08]. Antiviral [Wel05a]. Antonio [Wil09b]. any [Wel05-63]. AP [API^{+}07, DHT^{+}09, FSP^{+}07, GNPC^{+}08, LDGK08, LCT^{+}09, LRCH^{+}05, WPWF08, WBC^{+}06]. AP-1 [DHT^{+}09, GNPC^{+}08, LDGK08, LCT^{+}09, LRCH^{+}05]. AP-1/clathrin [LRCH^{+}05]. AP-1B-dependent [API^{+}07, FSP^{+}07]. AP-2 [WPWF08, WBC^{+}06]. AP1 [GBNS07]. AP2 [SSM^{+}08]. ApAF [LLL^{+}06, MDH^{+}06, PVM^{+}05, WSFD07]. Apaf-1 [MDH^{+}06, PVM^{+}05, WSFD07]. ApAF-ApC [LLL^{+}06]. ApAF-ApC/EBP [LLL^{+}06]. apart [Kak08, LeB07o, LeB07-40, Sho09o, DP08]. APC [CGK07, DP07, DSN^{+}07b, EMMN^{+}05, FBJB^{+}09, KST^{+}07, LeB05d, LeB06c, MVC^{+}07, ROKG08, RP08, RAP08, SRE^{+}05, Wil08c, YKH^{+}08, ZAM07a, ZAM07b]. APC/C [DP07, ROKG08, YKH^{+}08]. APC/EB1 [ZAM07a, ZAM07b]. ApC/EBP [LLL^{+}06]. ApCREB [LLL^{+}06]. APH [SRD^{+}07]. APH-1 [SRD^{+}07]. Apical [SLM^{+}06, BLP^{+}05, DGZ^{+}05, HP05, HT08, HSTM06, IMY^{+}09, JKDH08, JLC08, KGDR^{+}07, LSR07, LWHC08, LBG^{+}06, MT07, PPCZ06, SHS^{+}09, VVMS05, WL09a, WL09b, vBAS^{+}07]. Apoptosis [LeB07c, Les07b, TdVM^{+}07, AAG^{+}09, BP05, BFL^{+}05, BWJ05, BEH^{+}06, DSLY^{+}05, DSN^{+}07b, EIK^{+}05, GEH^{+}07, KNKH06, KHSM07, LCvD^{+}08, LMC^{+}09, MLC^{+}06, MHM^{+}05, MDC^{+}08, MRD^{+}08, NMV^{+}07, NSM05, PUG^{+}06, PPS^{+}06, PVM^{+}05, Pow07a, RHMS05, SWB^{+}07, TCHL05, Tum05-44, UFY^{+}07, WRKC^{+}09, WSS^{+}06, WPH^{+}07, WWS^{+}06, XKYW07, YHC^{+}06, YTB07, YZO^{+}07]. Apoptotic [LeB07d, Les09b, CCL^{+}05, JGL^{+}06, Les09-30, MYMZ08, Sco08, SMR09, SLZ^{+}06, WZM07, WPH^{+}07, WSFD07]. APP [LeB05e, LZL^{+}05, RP05, Wil08q]. apparatus [ARKS^{+}08, AM05b]. appendage [GSL^{+}07]. application [Hil08]. applies [INB06]. apposed [SCR^{+}09]. Apq12 [SHC07]. Apratoxin [Sho09d]. apraxia [SBC^{+}07]. April [lWHMH06b]. Aquaporin [Rob08c, NHK^{+}08]. aquaporin-2 [NHK^{+}08]. Arabidopsis [PST^{+}06, SSK^{+}09, VCN^{+}05]. architecture [BSM^{+}08, HGJ06, JBF^{+}08, MFH05, NPBP07, Sho09-36, Tak06, ZRH^{+}09]. ARE-containing [HAT^{+}05b]. ARE-mediated [vRG08]. Arf [HAAHD05, KCE^{+}06]. Arf-1 [HAAHD05]. Arf6 [EYB^{+}06]. ArfGAP [KBI^{+}08]. ARFGAP1 [LYH^{+}05, LDPLS05, WBE^{+}08]. ArfGAP2 [WBE^{+}08]. ArfGAP3 [WBE^{+}08]. Arg [LMM^{+}09]. arginine [CMD^{+}08, GTO^{+}07]. Ari [Wil08d]. ARK [MDH^{+}06]. ARL [KCE^{+}06]. arm [PKK09]. arms [BSM^{+}08, BSM^{+}09, HQF^{+}07, Wel05i, LCL^{+}05]. ARMS-dependent [LCL^{+}05]. Arp2 [LBI^{+}08, Les08c, MXR^{+}05, RXA^{+}08, SBMP05]. Arp2/ [Les08c]. Arp2/3 [LBI^{+}08, MXR^{+}05, RXA^{+}08, SBMP05]. Arp3 [MXR^{+}05]. arrangement [DL08, KOM07]. arrangements [Sho09p]. array [MB07, SSK^{+}09]. arrays [HRLH08, RMV06]. arrest [DUR^{+}05, FNH^{+}05, KPS^{+}07, LeB05-64, LTDH05, MYO^{+}05, MVC^{+}07, NTB^{+}06, NPV^{+}07, PWD^{+}06, WF07]. arrested [MMU^{+}09]. arrestins [PCR^{+}07]. arrhythmogenic [PRB^{+}09]. arsenite [SLZ^{+}06]. asb11 [DBvdW^{+}06]. Ase1 [CGGT09]. Ase1/ [CGGT09]. Asi1 [BZA^{+}06]. ASK1 [OTM^{+}07, SL05]. ASK1-dependent [OTM^{+}07]. Asp816Val [ZDD^{+}06]. asparagine-rich [DTP07]. aspartyl [VDA^{+}08]. aspects [YMC05]. Aspergillus [PUG^{+}06]. assay [PGWH08]. assemble [DZG^{+}07, LZ08, MMH09]. assembles [DPMB^{+}07, HXO^{+}07]. assemblies [CZM^{+}07, GC05]. Assembling [CSTT^{+}06, Tum06e]. Assembly [BSA^{+}08, DLH^{+}09, HYB08, TMH^{+}05, WSK^{+}06, AACS^{+}06, AGL^{+}08, ASH^{+}09, AHD^{+}09, BZHA^{+}08, BJLM^{+}05, BBM^{+}08, BFCS^{+}07, BSDA07, BYR08, BMP08, CM06, CSSE06, CMK06, DKR^{+}06, DBDD06, DTP07, Den05, DPP^{+}07, ELJ^{+}08, FZT^{+}07, FOW^{+}09, GS07, GHA^{+}05, GSV07, GL06, HRB^{+}06, HWPH^{+}08, HL06, HCL^{+}07, JS07, JRZ^{+}07, KCAS05, KLRS07, KNP^{+}05, KCH^{+}06, KTN07, LEH^{+}05, LLH^{+}07, LGZ^{+}05, LFH07, LLM^{+}05, LFW^{+}09a, LFW^{+}09b, LDPLS05, LRJY06, LLW^{+}09, MHM^{+}07, MMCW06, MSL^{+}09, MS09, MSV^{+}05, MD09, MFC^{+}09, MDPG07, NYM^{+}08, NdCF^{+}05, OW06, OLH^{+}08, OLKK09, OSM^{+}09, OYM^{+}09, ORM^{+}08, RBJL06, Rex09, RGCWG09, Rob08-29, Rob08-42, SHC07, SOR06, STK^{+}05, SAM^{+}07, Sho09-44, SBMP05, SSE05, SRD^{+}07, SGDM06, SNPCG06, SLNS06, TLM^{+}07, TWL07, TPNS07a, TPNS07b, TWG^{+}09, UTK^{+}09, ULN^{+}07, UPBM06, UZ08, VS07, Wil07-61, YDL^{+}06, YGR^{+}08, ZTM^{+}08]. assembly [vBAS^{+}07]. assists [KGL07, Rob08v]. associate [BJZ^{+}09, IGFG^{+}08]. associated [BSW^{+}05, BJM^{+}06, BvOM^{+}07, BXN^{+}06, BBG^{+}05, CFU^{+}09, CBG^{+}06, DVF^{+}06, DS07, FKJ^{+}09, FSR^{+}07, GMG^{+}07, HlQY^{+}06, JGL^{+}06, JZN^{+}09, LSC^{+}06a, LJY^{+}07, MS09, MSH^{+}05, MQL^{+}06, NRL^{+}05, NdCF^{+}05, OOS07, OOS09, OFfL^{+}07, PHM^{+}09, QIH^{+}06, SCFKB07, TMZ^{+}09, VCSG07, WLY06, WBK^{+}06, WBLL06, ZS08, vdBKB^{+}07, BDS^{+}08, OOY^{+}06, SS08a, SS08b, SN05, Tum05-63, WHR^{+}09]. associates [GFC^{+}08, HKIY09, MRvD^{+}05, WLK^{+}05]. Association [BGdN^{+}08, AFM^{+}05, CWRG^{+}08, EMMN^{+}05, JTH09, KLG^{+}09, LDGK08, LPM^{+}07, NS06, PGB^{+}06, PCN^{+}05, SSHP^{+}07, SDD^{+}07, Tum05u, WZM07, ZOH^{+}08]. associations [LC08a]. assume [SRXvdK06]. astralogy [Can09]. Astrin [TKSNG07]. Astrocyte [Wil08e]. astrocytes [BSF^{+}08]. Asymmetric [DPD^{+}09, YF08, CSW08, LS09, NVC05, RP07, TSK^{+}06]. asymmetrically [GAS06, KMK^{+}05b, RW05]. Asymmetry [BSM^{+}09, Tum05e, Lin08, RMAA08, Wel07c]. asynapsis [MBdR^{+}08]. Asynchronous [GHP06, GLH^{+}05]. ATAD3 [HMR^{+}07]. ataxia [SBC^{+}07]. ataxin [JMM^{+}06, WLY06]. ataxin-3 [WLY06]. ataxin-7 [JMM^{+}06]. ATF4 [YAV^{+}05]. ATF4-mediated [YAV^{+}05]. ATF6 [NSM05]. Atg11 [HSY^{+}06]. Atg7 [KWU^{+}05]. Atg7-deficient [KWU^{+}05]. Atg9 [HSY^{+}06]. atherogenesis [OSS^{+}07]. Atherosclerosis [Rob08d]. ATM [DPL^{+}08, DCA^{+}06, LBS^{+}08, SDC09, YHC^{+}06]. ATM/ATR [YHC^{+}06]. ATP [CCC^{+}06, JGL^{+}06, MSK05, COB^{+}08]. ATP-binding [JGL^{+}06]. ATPase [ADD^{+}06, CULS^{+}07, KRPH08, LBG^{+}06, CWC^{+}08]. ATR [DCA^{+}06, LLC^{+}07, SFB09, YLM06a, YHC^{+}06]. ATR-mediated [LLC^{+}07, YLM06a]. atrial [RKB^{+}07]. AtRMR1 [PLLH05]. atrophy [CBG^{+}09, JBL^{+}07, Les09x, RGW^{+}07]. ATRX [RSM^{+}08]. attach [Les05e]. attachment [HHZ^{+}08, JBF^{+}08, KLG^{+}09, PCN^{+}05, Sho09-28, VMSV08]. attachments [HFF^{+}07a, Nyb05f, YKBR09]. attenuates [KZT^{+}08, KZT^{+}09, LNG^{+}05, RKK09, vKST^{+}07]. attenuation [HGK^{+}06, SGK^{+}07]. atToc159 [WAKS08]. attract [LeB07s, Tum06-35]. attraction [Les08-45, Les09-31, MFC06]. attractive [Wil07n]. attractors [CFBM05]. atypical [HlQY^{+}06]. augment [VWK^{+}06]. Augmin [GMZ^{+}08, MFC^{+}09]. August [Wel06o]. aureus [BDR^{+}06, SDS^{+}05]. AURKB [HBPL09]. AURKB-mediated [HBPL09]. aurora [XOV^{+}09, BKHR05, BH08b, ELJ^{+}08, KNP^{+}05, MOS^{+}09, Sho09e, VML06, YK07, ZSMW09]. Aurora/Ipl1p [ZSMW09]. Australin [GGB^{+}08]. auto [RYD08]. auto-regulation [RYD08]. autocrine [NVJ^{+}08]. Autoimmune [LeB05f]. autoimmunity [Rob08a]. autoinduction [DLL^{+}08, YWRA07]. Autoinhibition [IEGV06, SOR06, NSM^{+}09]. autointegration [MNF^{+}07]. Automatic [Wel06b]. automatically [Hil07]. autonomous [CC09, IPC09]. autonomously [LNG^{+}05]. Autonomy [DGG^{+}08, GSHD08]. autophagic [FSR^{+}07, Miz05]. Autophagosome [AWM^{+}08, HTK^{+}08, McB08, Wel06-32]. Autophagosomes [Les09c]. Autophagy [KDD^{+}09, Rob08e, SG09, Sho09f, Wel05b, YCR06, BLB^{+}05, CCSK08, DDO^{+}08, GBNK08a, GBNK08b, HOH^{+}08, HSY^{+}06, HCB^{+}08, JFM^{+}09, JHY^{+}08, KWU^{+}05, LeB09a, NTSY08, PCMSB09, RVBZ^{+}08, RCT09, SFB^{+}05, Sed08m, SZL^{+}09, ST09, Tum05-49, WLE09, iYOW^{+}06]. Autophagy-mediated [YCR06]. autophagy-related [GBNK08a, GBNK08b]. Autophosphorylation [SDC09, UWiY^{+}07, DPL^{+}08]. auxilin [HBK^{+}06]. availability [KVC^{+}09]. avenues [LeB07z]. average [Sho09-46]. avoid [LeB05-99]. awareness [MJG07]. away [RW05, Wel05-33]. axis [DLL^{+}08, NWR^{+}08, PWM^{+}06a, PRZ^{+}08, TA07]. axodendritic [TÖL^{+}06b]. axoglial [CWNF^{+}09]. Axon [Nyb05a, DZG^{+}07, HXO^{+}07, HOR08, JBL^{+}07, LeB05-30, LeB07-28, Les08w, Les09-35, MB07, PT07, RAP08, SSJ^{+}07, STU^{+}08, WZC^{+}05, WZC^{+}08, YBT^{+}05, YOHR07, LeB06d, MEG^{+}07]. Axonal [GMSS06, BFB^{+}08, CKKG05, EGR^{+}05, FDB^{+}09, GKN^{+}07, HFY^{+}06, HFM^{+}05, KHB^{+}05, MS06, NZQ^{+}07, Tum06f, YWK^{+}08]. axonemal [AGL^{+}08, GWS09]. Axons [HJB07, HLJ^{+}07, Les08d, Bea06, CWM^{+}09, DG08, HCRT08, LeB08-35, MMT^{+}05, PBB^{+}09, Rob08-40, Wil08c, Wil08r]. axotomy [EMPK^{+}07].

B [GBNS07, Kat06b, KBW^{+}06, CLA^{+}05, SMWdV09, BWL^{+}08, BAR^{+}05, CCSBM^{+}07, CLK^{+}07, CWC^{+}07, CAL^{+}07, DTL^{+}07, ELJ^{+}08, GKCM^{+}08, HPT^{+}07, KBW^{+}06, KLRS07, LDVV06, LeB07f, LHM^{+}09, MOS^{+}09, OSB^{+}05, PWM^{+}06a, PGP^{+}07, QWD^{+}08, RMKS08, RRA^{+}09, Sed09g, Sho09e, STP08, SLZ^{+}06, STL^{+}07, SWB^{+}07, TCW^{+}06, Tum05t, VML06, WSS^{+}06, WM05, XOV^{+}09, ZTKK05, ZBKJ07, GBNS07]. B- [GBNS07]. B-mediated [ZTKK05]. B/thromboxane [PWM^{+}06a]. B1 [LKL^{+}05, MHL^{+}06, MLS^{+}07, MLV09, OKN06]. B1-mediated [OKN06]. B23 [YWM^{+}08]. B23/nucleophosmin [YWM^{+}08]. BACE [LZL^{+}05]. back [LeB06-45, LeB08l, LeB08s, Les08-60, Les09-35, Pow06b, Tum06g, VWK^{+}06, WPDZ^{+}06, WVB07]. Backtracking [Les08e, Les06t]. backups [Les06s]. Bacteria [Wel05c, LeB05-42, MRKH05, NBB^{+}09]. bacteria-induced [MRKH05]. bacterial [ABSK06, CJW07, CM07, LeB05a, SSE05, Wel05-34]. Bad [Les08f, Wil08q]. BAF [LCO07, Wil07-41]. BAF-1 [Wil07-41]. Bak [KPM05, PUG^{+}06, UDC^{+}07, WZM07]. balance [ELM09, MWMS07, Sho09-41, Wil07m]. balance-point [ELM09]. balances [SGK^{+}07, Sho09f]. balancing [WHB^{+}07, WKC^{+}09]. band [ERS08, WSK^{+}06]. BAR [MPS^{+}07]. barbed [FBD^{+}08, KCMB08, SMG07]. barbed-end [KCMB08]. Barr [FFS06]. barrel [HWN^{+}07]. Barrier [MNF^{+}07, Bon05c, EFM^{+}07, FSK^{+}07, IFF^{+}05, LCR^{+}05, LCB^{+}08, Pol08, SNG^{+}07, dGSWW^{+}06]. bars [CFV07]. Barth [CMK06]. Basal [POG^{+}09, Pow05c, Tum05f, KPR^{+}07a, KPR^{+}07b, KKY^{+}05, LeB05r, RYN05, STG05, SDW^{+}09, Tum05-54]. based [Bas07d, CCL^{+}05, EMPK^{+}07, GHM^{+}08, Hil08, LeB05-71, MFS^{+}09, OF08, PGWH08, RKT^{+}07, SSZR07, SWI06, VLFA^{+}07]. basement [LLM^{+}05]. basic [SJL^{+}07]. Basigin [BMK^{+}07]. basis [JWG^{+}07, LBH^{+}05, Mar08, MWSK06, NOH^{+}07, RXA^{+}08, SZX08]. basket [LFH07]. basolateral [API^{+}07, FHMC09, FSP^{+}07, HSTM06, KGDR^{+}07, LeB07-35]. Bassoon [FDB^{+}09]. Bat3 [SMM^{+}08a]. Bax [GGO^{+}09, GEH^{+}07, Let09, WZM07]. Bax/ [WZM07]. Bazooka [MMH09]. BBB [Rob08f]. BBS [Sho09g]. BBSome [LJS^{+}09a]. BC1 [MIBT06, WIL^{+}05]. Bcl [BbCQ^{+}09, KPM05, BRY^{+}09, RLDC09, GGO^{+}09, MGH^{+}09, PAM^{+}07, UDC^{+}07, WPH^{+}07, ZDMD06]. Bcl-2 [GGO^{+}09, UDC^{+}07, WPH^{+}07, ZDMD06]. BCL-2-dependent [BRY^{+}09]. BCL-2-like [RLDC09, MGH^{+}09]. Bcl-X [KPM05, BbCQ^{+}09, PAM^{+}07]. BCR [VLFA^{+}07]. BCR-driven [VLFA^{+}07]. BDNF [BRKR06, BSF^{+}08, Sho09h, YJJ^{+}09]. be [BYR08, LeB06-37, VD05, WPDZ^{+}06, WGL^{+}06]. bead [TMBL^{+}08]. bear [BMK^{+}06, Sho09-38]. beat [LeB06q, Les08-65]. beaten [LDVV06]. Beating [Tum05g]. because [BMK^{+}06, Wil08j]. become [Wel05-30]. becomes [Pow05z]. beefs [Les08g]. before [LeB07-62, Les09c, NMO^{+}07, Sho09l, STvdB^{+}09, Wel05-52, Wel06-61]. BEG4 [GQSHO05]. beginner [Nor06]. beginning [SD07]. behave [DCA^{+}06]. behavior [DvCvH^{+}08, HO08, SWM^{+}07, YF08]. behaviors [LSC^{+}05]. behind [LeB05-50, Les06a]. Being [BH08a]. believing [Nor06]. belt [IUT^{+}07]. belt-like [IUT^{+}07]. benchwarmer [DNK^{+}05]. bend [Nyb05g]. Bending [Ros07a, Wil07d, DLHW09, Wel05i]. bends [Les09r]. bent [LeB05g]. Bethesda [lWHMH06b]. betrayal [Wel06-57]. Better [RW08a, Nyb05o, Pow06f, Rob08g]. between [AMERT^{+}07, BJC^{+}06, BH08b, BSH^{+}07, BGdN^{+}08, COC^{+}05, CNM^{+}05, CBMCSS08, CRV^{+}06, DC07, Di 08, ERS08, FKB^{+}05, FKN^{+}07, JCMG08, JBF^{+}08, KGM^{+}09, KSpLL06, LGH^{+}05, LVL^{+}05, LR07, Mat05, MCA^{+}06, MBL^{+}07, MOCC05, NFCF08, OIFN^{+}07, OYTO07, OSI^{+}07, PLR^{+}06, QHB^{+}07, RSS^{+}09, RKK09, SCR^{+}09, SSZR07, TC08, TD08, TSS^{+}06, YBT^{+}09]. beyond [AKD06, IPC09, MGH^{+}09]. BFA [Pow06b]. Bfa1 [FDV^{+}06, GSdBS09]. bFGF [ML05]. bFGF-mediated [ML05]. BH3 [GGO^{+}09, LCvD^{+}08, MDET08, MGH^{+}09, NMV^{+}07, UDC^{+}07]. BH3-only [GGO^{+}09, MGH^{+}09, NMV^{+}07, MDET08]. bi [BJKF05]. bi-orient [BJKF05]. Biased [Wel06c]. Bid [TdVM^{+}07]. BiFC [SRBL06]. bifurcation [SSJ^{+}07]. Big [LeB06e, Wel07a, LeB05-103, Pow05d, Rob08-41, Rob08-48, Sed09i]. Bik [MDET08]. Bik-induced [GEH^{+}07]. Bik/Blk/Nbk [MDET08]. bilayer [Heu05, LeB05-47]. bilayers [BYMB06]. bilobe [dGHW08]. Bim [Pow07a, XJF05, MGH^{+}09, WPH^{+}07]. Bim1 [GHM^{+}08]. Bim1p [ZSMW09]. binary [DPD^{+}09]. bind [LDL^{+}09a, MKGL^{+}05, Sho09t]. Binding [LeB05g, LDGK08, YNL^{+}08, ABSK06, AHD^{+}09, BW08, BFP^{+}07, BHG^{+}05, BCJ^{+}07, CJMG05, DXC^{+}09, FHO^{+}05, FCB^{+}09, FCB^{+}19, FBD^{+}08, GRGP07, GHM^{+}08, GTB^{+}07, HBPL09, HSLI09, HM08, HMR^{+}07, JGL^{+}06, KLR^{+}07, KVS^{+}09, KDG^{+}09, KZS07, LHL^{+}08, LLM^{+}05, LPF^{+}06, MBR^{+}06, MXR^{+}05, MRO^{+}05, MSM^{+}07, MGBM^{+}09, MSK05, NRD^{+}06, OLC^{+}07, RIHP06, RYD08, RW08a, RRA^{+}09, SMH^{+}09, SCR^{+}09, STG05, SOW^{+}09, SGKM08, SRD^{+}07, SIS^{+}08, TCW^{+}06, TKB^{+}06, WAKS08, WCR^{+}05, Wil07r, ZNP^{+}09, BKLT^{+}08, BEN^{+}05, CHCB05, FBD^{+}05b, GG06, GFA^{+}09, GTX^{+}06, KCE^{+}06, KII^{+}08, LDGO^{+}09, MKSJ06, PAM^{+}07, VSR^{+}08]. binds [ABvBHH06, CCS^{+}06, FN06, KHBV05, LGV08, NdCF^{+}05, SSS^{+}08b, TCW^{+}06, YOY08]. bioactivity [HGJ^{+}08]. bioavailability [CCM^{+}07, LJN^{+}05]. Biogenesis [KKC^{+}07, SMWdV09, ASSST06, BWL^{+}08, BMGPB^{+}08, BDW^{+}09, CMF^{+}09, DHT^{+}09, GRGP07, GG09, GTO^{+}07, HWN^{+}07, HRS^{+}05, JhKC08, KOVP09, KRPH08, KZS07, LNL^{+}08, MRO^{+}05, MZH^{+}08, RCM^{+}09, RCKD06, SKM^{+}08, STM^{+}07, STG05, SBW^{+}07, SGKP^{+}07a, SGKP^{+}07b, TM06, dGHW08]. biological [EW08, Mar08, VD09]. biologically [MCD^{+}07]. biologist [Pow05z, Wil07-68]. biologists [Sho09k]. Biology [LK06, LW08, Mel08b, WL07, dSW09, BTP^{+}07, CFV07, HS09a, JKCW07, Lin08, Mel05, Mis09a, Sed08b, Wel07i, Wil07-55, LP05, Mis09b]. biomass [BbCQ^{+}09]. biorientation [HH06b, PMK^{+}07]. biosynthesis [KRG^{+}08, KDD^{+}09, VKR^{+}09]. BiP [ASB^{+}05b, LeB05h, Pow05d]. BiP-mediated [ASB^{+}05b]. biphasic [WWS^{+}06]. biphasically [GSJ^{+}08]. bipolar [JSPT05, TWG^{+}09]. Birth [Les09d]. bis [SDB^{+}05]. bisphosphate [JKKM08, VTRB06, vZPG^{+}07]. Bissell [Sho09-40]. bivalent [NVC05]. bizarre [Pow05b]. black [LeB06-81]. Blackburn [Wil08j]. blacklist [Wil08l]. bladder [Sho09-45]. blastoderm [BRH05, FML^{+}06]. blastula [MSO09]. ble [Les08-41]. blebbing [FG08]. blebs [CHCM06]. Blinded [Les07c]. blistering [WSB^{+}06]. Blk [MDET08]. BLM [OSI^{+}07, TNS07]. block [RRR^{+}09]. blockade [LeB06-63, Wel07b]. blocked [SBJ^{+}07]. blocking [KWY^{+}06]. blocks [MNF^{+}07, TCKV05, Wel06-54, ZHS06, ZL05]. Blood [LeB05i, Wil09c, Bon05c, CBG^{+}06, JKCW07, KZC^{+}08, KSE^{+}07, LeB07-37, LCB^{+}08, Pol08, Rob07m, WKC^{+}09]. Bloom [MGW^{+}08]. blow [Rob08e, Wil06c]. blueprints [Sho09-36]. BMP [HGK^{+}06, KLW^{+}05, LeB05-107, PRK^{+}07, PRK^{+}09, WHB^{+}07, KRP^{+}09, Sho09i]. BMP-activated [XYC^{+}07]. BMP1 [GG06]. BMP4 [FKS^{+}06]. BMP6 [PRK^{+}07, PRK^{+}09]. BMP6-dependent [PRK^{+}07, PRK^{+}09]. Bmper [Sho09i, KRP^{+}09]. Bnip [KBY^{+}08]. Bod1 [PMK^{+}07]. bodies [BJC^{+}06, BF09, BMP08, CBML06, DCA^{+}06, FSR^{+}07, JZN^{+}09, KSA^{+}05, KKY^{+}05, LTK09, LRCH^{+}05, Mat06, Rob08j, SNC^{+}06, SDW^{+}09, Wil07-40, DCMB^{+}07, HCC^{+}07, Wil07k, Wil07-31]. body [BAR^{+}08, CHA05, DTP07, GAS06, HCN06, JMG^{+}06, KPR^{+}07a, KPR^{+}07b, LJA09, LSC^{+}06b, LMB^{+}06, LNL^{+}08, NTS^{+}05, NdCF^{+}05, NAG^{+}08, POG^{+}09, PGO^{+}07, QIH^{+}06, STG05, STK^{+}05, TTYH09, Tum05f, Tum05-54, FBD^{+}05b, ZEC^{+}08]. bomb2 [NVEF07]. bonds [KGM^{+}09, LeB05j, Tho06]. Bone [dJPAW^{+}09, BYK08, FMM^{+}08, GHDS^{+}07, KR07, PXD^{+}07, Pow05i, SDL^{+}05, WKH^{+}06, YAV^{+}05, ZHS06, ZKR^{+}07]. bones [CJW07, LeB05-37, LeB06e, LeB07-52]. boost [Wil07-35]. Bop1 [HRS^{+}05]. Bora [SCD^{+}08]. born [LeB05-57, LeB05-67]. Both [GT08, Les09e, BJML09, Di 08, LBS^{+}08, Ped08, Sho09-29, Tum05-38, WSS^{+}06]. Botulinum [BFL^{+}05, DTL^{+}07]. bound [CBMCSS08, CFL^{+}05, DTR^{+}05, FDV^{+}06, JSPT05, SAM^{+}07, SKO^{+}06, YLM06a]. boundaries [GDW^{+}07, Pow07i, Wil08v]. Boundary [Tum05h, Mat05, RHMW^{+}05, TDD^{+}05]. bounded [PMW05]. bouquet [CYO^{+}09, Pow07b]. box [EHWL06, RMC^{+}05]. Bqt2p [TJC06]. Bqt3 [CYO^{+}09]. Brain [Les09f, Wil09d, BP05, Bon05c, GTX^{+}06, HTKR06, LS09, LCB^{+}08, NSS^{+}08, Pol08, SLCB05, WZW^{+}09, Wel05-49]. Brain-repairing [Les09f]. brain-specific [GTX^{+}06]. brains [LeB06-28, Wel06a, Wil08s]. brake [MB07, Wel06-62]. brakes [Les07-50, Sed09a]. branch [CB08]. branches [CFBM05]. branching [KZC^{+}08, LH05, LeB07-28, RXA^{+}08]. branchy [LeB05-103]. BRCA1 [KDD08, LeB06f, PL06]. breadcrumb [Sed08c]. break [BFCS^{+}07, HJB07, KG05, LeB05v, LeB08u, PGW^{+}08, TFC^{+}06b, YMTC05, YEL^{+}08]. breakage [DBK^{+}07]. breakdown [HMG^{+}08, Les07-47, Les08-57, MK07]. Breaking [LeB05j, Les07d, Nyb05b, Les07y, PvdGS06, Sed08a]. Breaks [Les08f, BJLK^{+}06, DCA^{+}06, DG05, KCD^{+}06b, Les07-51, MG09, Tum05k, UWiY^{+}07, YWJ^{+}07]. breakup [Les09u]. breast [CCH09, CCP^{+}08, LJM^{+}07, LCM^{+}08, MGK^{+}09, VFRJ^{+}07, YWW^{+}08]. breed [Wil09q]. breeding [TKK^{+}05]. Brett [Wil08f]. Brian [Wil08g]. bridge [COC^{+}05, TRS^{+}08]. Brief [Les06c]. bring [Les06j, Pow05f, Rob08y]. bringeth [LeB06-62]. Bringing [LeB05k, RH09, Sho09j, Wil07-55]. brings [Wel06k]. Brink [Les08r]. broad [WSK^{+}06]. broaden [HS09a]. browser [Hil08]. browser-based [Hil08]. Bruce [HCB^{+}08, Wil08h]. brucei [LN07, dGHW08]. Bruchpilot [FOW^{+}09]. brush [Sho09-39]. brush-off [Sho09-39]. Bryan [Sav09a]. BtubA [SSE05]. BtubB [SSE05]. BUB1 [NDS^{+}07, BSDA07, JMB^{+}07, KKM09a]. Bub2 [FDV^{+}06]. Bub3 [BJM^{+}06]. Bub3/Rae1 [BJM^{+}06]. BubR1 [HHZ^{+}08, MDC05, RGL^{+}09, WF07, ZAM07a, ZAM07b]. BubR1-dependent [MDC05]. bud [BFF^{+}06, LeB06m, LeB07-62, PZG^{+}07]. budding [BMSJ07, CRLP07, CN06, FDV^{+}06, GRGP07, HSY^{+}06, HLP06, LSC^{+}06b, MHFH06, Nyb05g, SAM^{+}07, TS06, ZKN^{+}06, ZSMW09]. Bug [LeB07e, LeB06-31, Wel05-63]. Bugs [Wel05d]. build [Les07-71, Rob08r]. Building [BH08b, JKCW07, LeB05l, Les09g, NAS08, Pow05e, Rob07a, Rob08f, Wel06-54, Wil09d]. built [BF09, Wel06-28]. bulk [HCN06, KKM^{+}08]. bull [Rob07a]. bundles [JSPT05, MT07, NAS08, STK^{+}05]. bundling [Tom08]. bursts [KCMB08]. business [LeB06c, Pow05u, Sho09g]. Buszczak [Sed09i].

c [CBM^{+}07, DFA05, NTB^{+}06, PBAF^{+}08, SDDV05, BTD^{+}07, BGZ^{+}07, BFL^{+}05, BFP^{+}07, BXN^{+}06, CG06, CC09, CFBM05, DP07, GNPC^{+}08, HLK^{+}09, KGKV07, KNT^{+}07, ROKG08, Sed09f, SZX08, SWB^{+}07, SFEK07, VBL^{+}05, Wel05-48, WEL^{+}07n, Wil08f, YKH^{+}08, BWJ05, BDS^{+}08, qGyWX^{+}08, GBNS07, KNKH06, KP08, LMS05, MKV^{+}09, MM05, YMDL09, ZKR^{+}07]. C-cadherin [CG06]. c-Jun [PBAF^{+}08]. C-mediated [HLK^{+}09, MDL^{+}09]. c-Met [CBM^{+}07]. c-myc [NTB^{+}06]. C-terminal [WEL^{+}07n]. C. [CLG^{+}07, GCS^{+}07, HK06, IEGV06, KHW^{+}07, KHW^{+}09, KO07, LDGO^{+}09, MNF^{+}07, MS07, OV09, POCS^{+}06, RLDC09, TA07, WZC^{+}08, ZRH^{+}09]. C/ [BTD^{+}07]. C/D [VBL^{+}05]. C2 [BJML09]. C3G [NNS^{+}08]. C3G-mediated [NNS^{+}08]. CA [LK06, WL07, dSW09, SFEK07, pTyLZ^{+}09, BSMT05, CZD^{+}09, CLK^{+}07, CSB^{+}06, FM05, GRC^{+}05, HJ06, HSK^{+}05, HPH^{+}05, ITG^{+}08, JBL^{+}07, LeB06-57, LeB06-75, LWG^{+}05, LWBL06, LGG^{+}05, Mel08a, Mua05, NJS^{+}06, OFY^{+}05, RDY^{+}05, RMP06, SBV^{+}06, pTyLZ^{+}09, TGRD09, WBK^{+}06, Wel05-44, WNF^{+}05b, WBLL06, XLH^{+}07, ZDMD06]. cable [Wil07b]. Caching [Les07e]. cacophony [LYV^{+}08]. Cad99C [DTY^{+}05b]. Cadherin [LeB05m, LeB06g, Les06d, BRKR06, DTY^{+}05b, HHY^{+}08, HLA^{+}06, LOG^{+}06, LNPC06, LR05, MCSY07, MFS^{+}09, MMH09, RHMW^{+}05, SMSPK08, TMH^{+}06, Tum05l, YA06, CG06, KSR^{+}08, LCS^{+}09, LBC^{+}07, LR05, NN07a, PR05, QCGM05, SFC^{+}08, SYM^{+}08b]. cadherin-based [MFS^{+}09]. cadherin-mediated [RHMW^{+}05, YA06]. cadherins [DCEM09, IMY^{+}09, Pow05u]. Caenorhabditis [ASSST06, AHM^{+}05, BAR^{+}08, DPD^{+}09, ECR^{+}09, ESG^{+}06c, GSW^{+}08, HKLM06, HKLM07, JNHB08, KR06, LLH^{+}07, LBG^{+}06, MYMZ08, MQL^{+}06, NAG^{+}08, OOS07, OOS09, OBY^{+}06, SHL^{+}09, YWEK08]. Caf4p [GGC05]. Cage [LeB06h]. Cajal [CBML06, LMB^{+}06, Mat06, PGO^{+}07]. calcineurin [LPWP^{+}07, LGG^{+}05]. calcitonin [PCR^{+}07]. Calcium [LeB05n, Les09h, Tum05i, Wil07e, BAdSG05, BCC^{+}09, HPT^{+}07, LeB05s, LeB05-54, LeB05-33, LeB06-63, Les08b, LWL08a, LWL08b, LYV^{+}08, MuMI^{+}06, NYI^{+}05b, PMW05, Put05, RSL06, Rob08w, Tum05-65, TTE^{+}07, VBG^{+}06, WWS^{+}06, Les09i]. calcium-independent [NYI^{+}05b]. call [Les08-30]. calls [Les08z]. calmodulin [FM05]. Calpain [CCP^{+}08, DBC^{+}06]. Calreticulin [LGG^{+}05, SQB^{+}08]. CAM [STU^{+}08, DS07, HLL^{+}08, ROKG08]. CaM- [DS07]. CaMKII [BSL^{+}08, HSK^{+}05, WBG^{+}08]. cAMP [TDP^{+}06, GRC^{+}05, HHT^{+}07, RMP06, SBM^{+}07, TDT^{+}08]. can [BSK07, BMK^{+}06, BEH^{+}06, GHY^{+}05, HK07, HMM^{+}05, LCvD^{+}08, MTRD^{+}06, NMO^{+}07, NKK^{+}06, PH07, Pow05l, SN05, WHC^{+}07, WPDZ^{+}06, WGL^{+}06]. Canada [Wel06o]. Cancer [Les07f, BPL^{+}05, CCH09, CCP^{+}08, DLBSD^{+}05, Les07n, Les07-37, Les07-65, Les08t, Les08-61, Les08-62, Les09-37, LJM^{+}07, LCM^{+}08, MYO^{+}05, MGK^{+}09, OMLP^{+}08, PhXD^{+}06, SSHW09, Sed08k, Sed09j, Sed09k, SWM^{+}07, Wel06s, Wil06h, YWW^{+}08]. cancer-free [Wil06h]. cancers [GCSG^{+}05]. Cande [Wil07-68]. cannot [BYR08]. Canoe [SHS^{+}09]. Canonical [LVG^{+}09, DXC^{+}08]. can’t [Les07l, Les09-37]. CAP [PTF^{+}06]. CAP1 [LCR^{+}05]. CAP1/ [LCR^{+}05]. Capacitative [Put05]. Capacity [WLPB^{+}09, DL08, SSS^{+}08a]. Capital [LW08]. cappers [KCMB08]. capping [MTH^{+}06, SOR^{+}05]. Cappuccino [QHB^{+}07]. CAPRI [LWG^{+}05]. Caprin [DXC^{+}08]. Caprin-2 [DXC^{+}08]. CAPS [HWSJ08]. Capture [SM09, GC05, HH06b, MDC05, WF06, ZMA^{+}07]. CAR [AHM^{+}05, DP06]. CAR-1 [AHM^{+}05, DP06]. carbohydrate [DNK^{+}05, GDW^{+}07]. carbs [Les07e]. carcinoembryonic [NS07]. carcinogenesis [GBNS07]. carcinoma [KNKH06, SWD^{+}08, vRSvR^{+}07]. Cardiac [ONW^{+}07, TMW^{+}05, AS09, GPN^{+}06, MAZ^{+}08, NFC^{+}08, NRJ^{+}09, SHX^{+}07, SAD^{+}09, WBL^{+}06]. Cardio [Les08g]. Cardiolipin [COB^{+}08, GSH^{+}08, DDRL^{+}09, KRG^{+}08, OHP^{+}09, Sco08, TEIS09]. cardiomyocytes [CGLC^{+}08, CZSG08, ONW^{+}07, PVM^{+}05]. cardiomyopathies [ZS08]. cardiomyopathy [PRB^{+}09]. cardiovascular [ZMW^{+}08]. Career [Les08h, Wil07f, Wil09l]. cargo [ALB^{+}09, API^{+}07, FHO^{+}05, FSP^{+}07, HR09, JWG^{+}07, KLR^{+}07, LJS^{+}09a, LYH^{+}05, LDGK08, Les07-46, MSM^{+}07, NTHN^{+}08, PLLH05, PCN^{+}05, SOW^{+}09, SSSB07, VVMS05, yYMM05, vBDF^{+}09]. cargo-binding [MSM^{+}07]. cargo-specific [SSSB07]. cargos [Les07-76]. carrier [COB^{+}08, HHS^{+}09, SMM^{+}08c]. carriers [BBPM07, Wel05-39]. cascade [LGG^{+}05, SHC^{+}05, SLZ^{+}06]. cascades [NBM^{+}08]. case [Pow05e, Ros07b, Wil06k]. casein [PPK08]. CASK [CHH^{+}08]. Caspase [Les07g, MDC^{+}08, AAG^{+}09, GSH^{+}08, HCB^{+}08, KKM09b, LDVV06, LYY08, NDS^{+}07, Pow09d, RHMS05, RKT^{+}07, Rob08-27, SLK^{+}08, YSB^{+}07, YHC^{+}06, DOVD08, MKB^{+}09, Sco08]. caspase-1 [Rob08-27]. caspase-12 [SLK^{+}08]. Caspase-14 [DOVD08]. Caspase-2 [MKB^{+}09]. caspase-3 [YHC^{+}06]. caspase-8 [GSH^{+}08, Sco08]. caspase-dependent [AAG^{+}09]. Caspase-free [Les07g]. caspase-independent [NDS^{+}07]. Caspase-mediated [MDC^{+}08, LDVV06, RHMS05]. Caspases [LDVV06, Les08u, LYY08, LVDR05, VDVV07]. cassette [JGL^{+}06]. cassettes [KCD^{+}06a]. Casting [Les07h]. casts [Les07-41]. Catalytic [Wel05e, BSL^{+}08, ToMY^{+}08]. Catalytically [LMMHP^{+}05, WDWC08]. catalyzed [BSMT05]. Catastrophe [LeB07f, CCSBM^{+}07]. Catch [HK07, KGM^{+}09, Pow05w, Sho09w, Tho06, Tum05-75, Tum06-49]. Catching [Wel05f]. Catenin [RHMW^{+}05, BRKR06, DBR09, qGyWX^{+}08, GCSG^{+}05, HFvdV^{+}05, KKW^{+}09, LeB05-75, LPN^{+}08, Les09s, LMHT08, LCB^{+}08, LGV08, MFS^{+}09, MMH09, SYM^{+}08b, Tum05d, YA06, dJPAW^{+}09]. cathepsin [LMMHP^{+}05, SWB^{+}07]. Caught [Les07i, Les07-58]. cause [EVF^{+}08, MTRD^{+}06, PhXD^{+}06, WRKC^{+}09]. caused [DC07]. causes [DMBF^{+}07, HJK^{+}06, JSFP08, JLG^{+}06, RMR^{+}09, WSB^{+}06, WBL^{+}06, WBLL06]. caveolae [Les08-39]. caveolar [BLH^{+}09, SHjC^{+}07, TMH^{+}05]. caveolin [CWC^{+}08, CTY^{+}06, DPK^{+}05, GJL^{+}08, KFC^{+}05, Wel06-28, GGEd^{+}07]. caveolin- [DPK^{+}05]. caveolin-1 [CWC^{+}08, GJL^{+}08, GGEd^{+}07]. caveolin-independent [KFC^{+}05]. Cavin [BLH^{+}09]. Cavin-4 [BLH^{+}09]. CBF3 [GTHK05]. Cbk1 [JBYW06]. Cbl [EIK^{+}05]. CBP80-mRNPs [dPMF^{+}09]. CCCH [HOZ^{+}09]. CCCH-type [HOZ^{+}09]. CCM1 [GHSG07]. CCN1 [TCHL05]. CCN3 [FKML^{+}06]. CD151 [SKvdBW^{+}06]. CD2 [KDV^{+}09]. CD2AP [JSC08]. CD2AP/CIN85 [JSC08]. CD3 [MDW05]. CD317 [RKH^{+}09]. CD317/tetherin [RKH^{+}09]. CD36 [SDS^{+}05]. CD36-mediated [SDS^{+}05]. CD43 [IKZ^{+}07]. CD44 [KZH^{+}07, LWC09, Les09j, THN^{+}06]. CD44-mediated [THN^{+}06]. CD53 [DPMB^{+}07]. CD81 [DPMB^{+}07]. CD9 [DPMB^{+}07, EMD^{+}08]. CD95 [GHM^{+}09]. CD95-induced [GHM^{+}09]. CD98hc [FZT^{+}07]. Cdc14 [CWRG^{+}08, JhKC08, KLRS07, MHC^{+}09, QU08]. Cdc14-regulated [KLRS07]. Cdc14B [WCR^{+}08]. Cdc20 [LFW^{+}09a, LFW^{+}09b, NRD^{+}06]. Cdc25 [PPK08, RMAA08]. CDC25A [ZNP^{+}09, KBK^{+}05, LKL^{+}05]. Cdc25B [LKL^{+}05]. Cdc28 [EHHK09]. Cdc28/ [EHHK09]. Cdc37 [MLC^{+}07]. cdc42 [BTP^{+}07, BB05, CFL^{+}05, DP08, EMMN^{+}05, FKN^{+}07, HT08, JKDH08, KBY^{+}08, Lec08, Les08i, OYKY^{+}05, OIAT06, RFBB07, SYRL^{+}08, SMH07, TGH^{+}07, VWK^{+}06, WL09a, ZOH^{+}08]. Cdc45 [AH05]. Cdc6 [Tum06k, YHC^{+}06]. CDK [TWL07, CSMP^{+}09]. CDK-related [TWL07]. Cdk1 [EHHK09, RMKS08, HDS^{+}07, LKL^{+}05, SIL^{+}07, ToMY^{+}08, WF07, YGY^{+}05].

Cdk11 [YGR^{+}08]. Cdk2 [AH05, JAV^{+}07, SLMK08]. CDK4 [ZTKK05]. Cdk5 [LeB05o, LVDR05]. CDK6 [ZNP^{+}09]. cdks [Les07w]. Cdo [KBY^{+}08, TKB^{+}06]. CEACAM [MRKH05]. CEACAM1 [KVS^{+}09, MKV^{+}09]. CEACAM1-L [MKV^{+}09]. CEACAM1-S [MKV^{+}09]. CED [RLDC09]. CED-9 [RLDC09]. Cell [FG08, FKB^{+}05, GAK^{+}05, HCFM08, Kat06c, KR06, LD07, LP05, LeB06i, LK06, LeB07g, LW08, Les08j, Lin08, Mel08b, MFBG^{+}08, OFY^{+}05, Pow07c, SWT^{+}05, Sho09k, Sho09l, Sho09m, Tum06g, ULN^{+}07, WL07, dSW09, AKD06, AM05b, ASSST06, BFG^{+}06b, BTD^{+}07, BDP^{+}07, BP05, BGZ^{+}07, BB05, BTP^{+}07, BBM^{+}08, BAdSG05, BMK^{+}07, BFP^{+}07, BLB^{+}05, BvOM^{+}07, BLMS06, Bon05f, BFF^{+}06, BMG^{+}07, BHK^{+}06, CJW07, CZM^{+}07, CMD^{+}08, CGLC^{+}08, CLL07, CF07, CNM^{+}05, CCL^{+}08, CHCM06, CG06, CSW08, CMF08, CK07, CULS^{+}07, CCP^{+}08, CWC^{+}07, DL08, DLT^{+}07, DBR09, DVF^{+}06, DWMY09, DMC^{+}07, DCEM09, EPV^{+}05, EMB^{+}08, EVH^{+}07, EVH^{+}08, EVF^{+}08, EMMN^{+}05, FFW^{+}07, FHMC09, FSG^{+}07, FAG^{+}07, FIK^{+}05, FIK^{+}20, GdTAR07, GGO^{+}09, GMW^{+}05, GH05, GMvdKC07, GHM^{+}09, GGMBF08, GAS06, GHSG07, GPJ^{+}06, GGEd^{+}07, GEE07, GGH^{+}08, HOM^{+}05]. cell [HKN09, HXZ^{+}07, HEW07, HCN06, HIK^{+}08, HLK^{+}07, HRS^{+}05, HHFC06, HLA^{+}06, HSTM06, HWJ^{+}08, IUT^{+}07, IPC09, IPG^{+}05, IH05, JGAS08, JLB^{+}05, JKCW07, JMB^{+}07, JLM^{+}07, JLL^{+}07, JTN^{+}05, KRZ^{+}08, KNKH06, Kat06j, KBK^{+}05, hKDH^{+}08, KKW^{+}09, KLW^{+}05, KCD^{+}06a, KTN07, KBR^{+}08, KSR^{+}08, KWY^{+}06, LNG^{+}05, LS09, LMM^{+}09, LeB05j, LeB05-107, LeB06-44, LeB08-34, LLJ^{+}05, LSC^{+}07, Les06a, Les07b, Les07-68, Les07-62, Les08y, Les08-50, Les09e, LPF^{+}06, LLT^{+}08, LSC^{+}05, LCL^{+}07, LHM^{+}09, LTDH05, LG07, LG08, MCSY07, MFA^{+}07, MYO^{+}05, MPL^{+}07a, MCF^{+}05, MFS^{+}09, MWMS07, ML05, MNF^{+}07, Mar08, MCRVC^{+}08, MEG^{+}07, MSO09, MDET08, MAZ^{+}08, Mel05, MAG^{+}08, MDH^{+}06, MKGL^{+}05, MSK^{+}07, Mis09a, MMU^{+}09, MSC^{+}07, MLHS06, MSM^{+}06, MRKH05, MDW05, NPMZ06, NPHT08, NN07a, NDS^{+}07, NYI^{+}05b, NTY^{+}05, NYN^{+}06, NNS^{+}08, Nyb05h, OW06, OKN06, OYTO07, OBY^{+}06]. cell [OSI^{+}07, OV09, PR05, PMS^{+}07, PPA^{+}08, PLH^{+}08, PEER^{+}05, PTL^{+}05, PMW05, PB05, PG06, PCMSB09, PYR08, PGWH08, PAV^{+}06, PBB^{+}09, PT07, PWD^{+}06, PCPW08, Pow05l, Pow05w, Pow05z, PJMR06, PEW^{+}05, PdBR^{+}07, PVJJ07, QCGM05, QWD^{+}08, QP05, RVBZ^{+}08, RHMW^{+}05, RYD08, RMAA08, Rob07d, Rob08-42, RRA^{+}09, RP07, SSHW09, SFP^{+}05, SGMPV07, SON06, SYRL^{+}08, SRL05, SMSPK08, SWD^{+}08, SJL^{+}07, SBF^{+}06, SMH07, Sch08, Sed08b, Sed09i, SYA^{+}08, SLM^{+}06, SRR^{+}08, SDN^{+}05b, SCFKB07, SKR^{+}07, ShCM^{+}06, SRXvdK06, STP08, SYM^{+}08b, SGDM06, SAD^{+}09, SWG^{+}05, SLV^{+}07, SLWA07, SWB^{+}07, SRQL09, TMH^{+}05, TKB^{+}06, TTG^{+}09, THL^{+}06, TP09, TMI07, TMW^{+}05, pTyLZ^{+}09, TLK^{+}06, TMAL^{+}06, TCKV05, TWC^{+}07, TGRD09, Tum05-35, Tum05-64, Tum06z, Tum06-30, Tum06-34, VP05, VBG^{+}08, VDSV06, VCSG07, VFRJ^{+}07, WZO^{+}05, WMM^{+}09, WC07a, WLC08, WLE09, WL09a, WZM07]. cell [WLLF^{+}08, WM05, Wel05g, Wel05-29, Wel06b, WCN07, Wil06j, Wil07-52, Wil07-55, Wil07-56, Wil07-68, Wil08p, WZCB^{+}06, WPDZ^{+}06, WXM^{+}05, XSVGR05, YWJ^{+}07, YN07, YF08, YMCT08, YA06, YR09, YTB07, YKP^{+}08, YWW^{+}08, YMDL09, YCX^{+}08, ZXM^{+}06, ZDMD06, dRKD^{+}05, vdBKB^{+}07, LeB06-52, Mis09b, lWHMH06b]. cell-based [PGWH08]. cell-free [OW06]. cell-mediated [DLT^{+}07]. cell-penetrating [BGZ^{+}07]. Cells [GNE^{+}07, LP05, LeB05p, LK06, LW08, Les08f, Les08x, PGP^{+}07, Rob07b, WL07, YKBR09, AOT06, ABA^{+}06, ALB^{+}09, AOP^{+}05, AMS^{+}05, ANI^{+}07, API^{+}07, BTD^{+}07, BAR^{+}05, Bas07a, BPL^{+}05, BW08, BYK08, BvRJ^{+}05, BLP^{+}05, BH08b, BMB^{+}07, Bon05b, BDR^{+}06, BMK^{+}06, BRY^{+}09, BBG^{+}05, BEH^{+}06, CHSV07, CYC^{+}06, CS09, CKG^{+}08, CSMP^{+}09, CCC^{+}05, CSL^{+}05, CBG^{+}06, DMBF^{+}07, DKR^{+}06, DLT^{+}07, DGZ^{+}05, DBC^{+}06, DLBSD^{+}05, DDLC^{+}06, DHML05, DZW^{+}08, EIK^{+}05, EBMG09, FHMC09, FMM^{+}08, FSP^{+}07, FBY^{+}06, FPP^{+}08, FKN^{+}07, Fuc08, GEP08, GCSG^{+}05, GEN^{+}06, GHP06, GT08, GKN^{+}07, GSY^{+}09, HTK^{+}08, HOH^{+}08, HTYNK09, HGA^{+}09, HMG^{+}08, HWPH^{+}08, HMM^{+}05, HDS^{+}07, HLL^{+}09, Hor09, HL06, HLJ^{+}07, HSTM06, IFF^{+}05, INB06, JAV^{+}07, JCVB08, JGL^{+}06, JLL^{+}07, KDV^{+}09, KR07, KMK^{+}05b, Kat06b, Kat06e, Kat06g, KWS^{+}08, KG05, KKK^{+}05, KAD^{+}08, KZH^{+}07]. cells [KCD^{+}06b, KCOS06, KS08, LEH^{+}05, LPG^{+}07, LHS^{+}05, LSDFS08, LGH^{+}05, LeB05z, LeB05-29, LeB05-61, LeB06k, LeB06-77, LeB07-37, LeB07-56, LeB07-63, LeB08j, LeB08q, LeB08-33, LW07b, LSM^{+}05, Les07f, Les07-30, Les07-34, Les07-37, Les07-65, Les07-78, Les07-79, Les08c, Les08n, Les08t, Les08z, Les08-51, Les08-47, Les08-60, Les08-61, Les08-62, Les08-64, Les09b, Les09f, Les09-30, Les09-37, LJM^{+}07, LLM^{+}05, LCT^{+}09, Lin08, LDPLS05, LRJY06, LLC^{+}07, LVG^{+}09, LMZ^{+}05, LJS^{+}09b, LCM^{+}08, LRCH^{+}05, LLC^{+}08, MCSY07, MYO^{+}05, MYMZ08, MuMI^{+}06, MSSF05, MG08b, MMT^{+}05, MOM^{+}06, MZH^{+}08, MRKH05, NBB^{+}09, NCB^{+}08, NCB^{+}09, NPMZ06, NTB^{+}06, NZQ^{+}07, NFCF08, NFZ^{+}09, OLS08, OIT08, OMLP^{+}08, OIAT06, ONW^{+}07, PPCZ06, PMB^{+}09, PH07, PhXD^{+}06, Pow05n, PST^{+}06, PKB^{+}06, PVJJ07, RYN05, RMZ^{+}06, RKT^{+}07, Rob07m, Rob08-47, Rob08m, Rob08-38, Rob08-54, RSSC06, RKH^{+}09]. cells [RLH^{+}09, SBC^{+}05, SWD^{+}08, SSS^{+}08a, Sed08c, STM^{+}07, SPA^{+}05, SFY^{+}05, SSM05, Sho09y, Sho09-27, Sho09-38, SWM^{+}07, SMR09, SRXvdK06, SHK^{+}07, SNPCG06, SIS^{+}08, TDP^{+}06, TC08, TMW^{+}05, TD08, Tum06-51, ULN^{+}07, VSG09, VMY^{+}09, VTRB06, VMZW^{+}07a, VMZW^{+}07b, VMKW^{+}08, VVMS05, VCC^{+}08, VS07, Wel06-39, Wel06-59, Wel07a, WPC^{+}08, WMSB^{+}06, Wil07-33, Wil07-66, Wil07-64, Wil09j, Wil09o, Wil09p, WWS05b, WWL07, WPDZ^{+}06, WVB07, WLPB^{+}09, YHC^{+}06, YSP^{+}06, YZO^{+}07, ZHI^{+}08, ZXM^{+}06, ZNP^{+}09, ZEC^{+}08, ZL05, vRSvR^{+}07, vRCB^{+}07]. cellubrevin [FSP^{+}07]. Cellular [GHM^{+}09, HK05, Les08k, LBH^{+}05, Wil07f, Wil07g, BH08a, CTS^{+}06c, FCB^{+}09, FCB^{+}19, FSTI06, GZBM^{+}06, KDD08, LSR07, LPB^{+}08, MFC06, MSECB05, OMGW^{+}06, PvdGS06, Pow05x, RSL06, SOR06, SLR^{+}06, TTK^{+}06, TGH^{+}07, Wel05-69, Wil09l, ZJH08]. cellularization [NFC^{+}08]. cellulose [Wel06-63]. CENP [HGJ06, ASH^{+}09, KHWC08, Les07-32, Les08l, MHM^{+}07, MDC05, MMC08, PNO^{+}09]. CENP-E [KHWC08, Les08l]. CENs [Mal06]. center [COC^{+}05]. Centering [Les07j]. centers [CSSvM08, KG05]. central [ACsHS08, CWNF^{+}09, CTS^{+}06c, DHR^{+}08, GKCM^{+}08, LW07a, LYH^{+}05, RW08c, YBK^{+}06, ZMW^{+}08, ZTM^{+}08, Hil07]. centralspindlin [YPG05]. centration [GCS^{+}07, KO07]. centrin [AHD^{+}09, LeB05-43, LSC^{+}06b]. centrin-binding [AHD^{+}09]. Centriolar [LeB07h, DMDO08]. centriole [BSW^{+}05, DPP^{+}07, GSL^{+}07, HRB^{+}06, LEH^{+}05, Les07j, NAF09, RRR^{+}09, ULN^{+}07, VS07, WCR^{+}08]. centriole- [DPP^{+}07]. centriole-associated [BSW^{+}05]. centrioles [AHD^{+}09, DMDO08, Les09a, Les09d, LR07, SDW^{+}09]. Centrobin [ZLB^{+}05]. Centromere [JCVB08, Wil09e, BSDA07, CS09, EMB^{+}08, HGJ06, HFF^{+}07a, Les07o, Les07-32, MSC^{+}07]. centromere-specific [HGJ06]. Centromeres [Rob08h, HCK^{+}07, JBF^{+}08, KYK^{+}06, PL06]. Centromeric [CS07, LeB05q, HW09, JBFC07, MMC08, PNO^{+}09, YK07]. Centrosomal [NAF09, BKHR05, KO07, ULN^{+}07]. Centrosome [LeB05r, Les06e, RW05, BSW^{+}05, BWW^{+}08, BBL^{+}08, DCEM09, FDM^{+}08, GMZ^{+}08, HRB^{+}06, KR06, KNP^{+}05, LeB07n, LKL^{+}05, MPL^{+}07a, Rob07l, RP07, SSHP^{+}07, SGDM06, TKSNG07, TMZ^{+}09, VCD^{+}06, Wel05-51, Wel06-31, YF08]. centrosome-associated [TMZ^{+}09]. centrosome-dependent [KNP^{+}05]. Centrosome-free [LeB05r]. centrosome-independent [GMZ^{+}08]. Centrosomes [Wil06c, BKHR05, CK07, LeB06i, LeB08a, Les07p, RBJL06, TWG^{+}09, Tum05-41]. Centrosomin [LR07]. Cep164 [GSL^{+}07]. ceramide [BVGA^{+}05, FHS^{+}07b, VTT^{+}09a, WRKC^{+}09, ZJH08]. ceramide-rich [BVGA^{+}05]. cerebral [ABY^{+}08, IMY^{+}09]. cerevisiae [EHHK09, FMK^{+}09, HH06b, MSB06, NSO^{+}09, AFN^{+}08, BMP08, CCSK08, DTP07, ESZK06, FFT^{+}05, GAS06, KKGD05, NRD^{+}06, TKK^{+}05, TP09]. CERT [Sed09a]. CFTR [GOM^{+}09]. CGH [AHM^{+}05, BAR^{+}08]. CGH-1 [AHM^{+}05, BAR^{+}08]. chain [BMT^{+}07, FHNG05, FDB^{+}09, GMSS06, KKM^{+}08, KMS07, KHSM07, LB08, PKK09, RG06b, WPL^{+}06, WvdLH^{+}07]. chains [LJA09, MLC^{+}07]. Chait [Wil08g]. challenge [Ros09]. chance [LC08a]. change [ABvBHH06, Les08h]. Changes [KCD^{+}06b, AFV^{+}05, CHSV07, HlQY^{+}06, LeB08c, MS09, oSCW07, SMRS^{+}07, VTRB06]. changing [Mis09a, LeB07z]. Channel [LeB08c, AFV^{+}05, BCC^{+}09, CZD^{+}09, CLK^{+}07, FLA^{+}08, IBS^{+}05, JBL^{+}07, KBGG^{+}08, LeB05-96, Les09h, LPB^{+}08, LYV^{+}08, RDY^{+}05, VBG^{+}06, WJM05, WBG^{+}08, Wil08i, XLH^{+}07]. channels [BH08a, BFB^{+}08, COTAE^{+}05, HFP^{+}05, HSK^{+}05, Les08-40, LWBL06, Ras08a, SBV^{+}06, pTyLZ^{+}09, Wil07e]. Chaperone [SBV^{+}06, BMGPB^{+}08, JASC08, KGL07, LLH^{+}07, Rob08p, SZL^{+}09, Wil07m, ZKG^{+}08]. Chaperone-mediated [SBV^{+}06, SZL^{+}09]. chaperones [ERS08, KL05, LeB05h, LeB06-69, LeB08s, Sho09d, Wel06-43]. characterization [TDD^{+}05, VS07]. Charcot [TT09]. charge [BXHD08, YHD^{+}09]. Charged [Wel06d]. check [Les06w, Sho09l, Wel07k]. check-up [Wel07k]. Checkpoint [CN06, ESZK06, HKLM06, HKLM07, Kat07a, BJKF05, BLT^{+}09a, BYR08, CCS^{+}08, DBK^{+}07, FVSC08, GSV07, KKW^{+}05, LeB06w, LeB06-42, LYWK08, Les07-64, LFW^{+}09a, LFW^{+}09b, LLC^{+}07, MDC05, MS09, MD09, MTV^{+}05, MDPG07, MBF^{+}07, MBF^{+}08, NRD^{+}06, OLH^{+}08, ORM^{+}08, Pow06m, RGL^{+}09, Rob08-29, SS07, Tum05-45, UTK^{+}09, VDSV06, WH06, WF05, WF07, YKH^{+}08, YLM06a, YKBR09, Yu06]. checkpoint-mediated [LFW^{+}09a, LFW^{+}09b]. checkpoints [LeB05-99]. Cheeseman [Sho09-28]. chemical [DCMB^{+}07, HDS^{+}07, MAG05]. chemoattractant [KBR^{+}08]. Chemoattractants [KGKV07]. Chemokine [Wel07b, GMvdKC07, LeB08-29]. chemokine-induced [GMvdKC07]. chemokines [Wel06-38]. chemorepellents [KGKV07]. chemosensing [XMSYJ07]. Chemosensitivity [Sho09n]. chemotactic [KNT^{+}07]. chemotaxing [JLL^{+}07]. Chemotaxis [Pow07d, Wel05h, KNT^{+}06, PB05, Tum06-34, VKGV08, WFJ06, vKGK07]. chemotherapy [BRY^{+}09]. chewing [Rob07c]. Chibby [LMHT08, VTT^{+}09b]. chill [Les08q]. CHIP [LHZ^{+}08]. Chironomus [KHBV05]. chitin [Nyb05m]. CHK [KHW^{+}07, KHW^{+}09]. CHK-1 [KHW^{+}07, KHW^{+}09]. Chk1 [RMKS08, SFB09, TMZ^{+}09, YLM06a]. CHK2 [BLT^{+}09a, DCA^{+}06]. Chlamydia [PCV^{+}08]. chlamydial [PCV^{+}08]. Chlamydomonas [LW07a, Les09k, BSM^{+}08, BSM^{+}09, Den05, GWS09, HDM^{+}07, LJS^{+}09a, TWL07, UZ08, iWK06]. chloroplast [UZ08]. chloroplasts [CCC^{+}06, LS06, WAKS08]. CHMP5 [SXH^{+}06]. choice [GLH^{+}05, Les07k]. Cholesterol [DSLY^{+}05, RKvdK^{+}09, Sed07a, CTY^{+}06, CC09, LeB07x]. Cholesterol-induced [DSLY^{+}05]. chondrocyte [ZBKJ07]. chondroitin [OBY^{+}06]. choosy [Tum05-81]. CHOP [Sho09-27]. choreography [Les06e]. Christian [Sho09o]. chromatid [KPS^{+}07, TKSNG07, WWK^{+}06]. chromatids [JGY09]. Chromatin [AH05, Tum06h, WSFD07, AVHH09a, AVHH09b, AG09, BLMS06, BEN^{+}05, BGdN^{+}08, CS07, CDW^{+}05, DGSN06, DTR^{+}05, GLX^{+}09a, GLX^{+}09b, GTB^{+}07, HGJ06, HBPL09, HM08, HKP^{+}09, JBFC07, KCD^{+}06b, LeB06-27, Les07k, Les09-36, LJS^{+}09b, MHM^{+}07, MBdR^{+}08, MMC08, McN09, MBJF^{+}08, MRK^{+}07, MJF^{+}07, NWR^{+}08, OLKK09, PNO^{+}09, Sed08g, Sed09d, SMT^{+}08, TDD^{+}05, TMAL^{+}06, WLS^{+}09, Wil07-45, Wil09o, XNM^{+}09, vdBKB^{+}07]. chromatin-associated [vdBKB^{+}07]. chromatin-bound [DTR^{+}05]. chromokinesin [TTO^{+}08]. chromosomal [BSO^{+}08, CYO^{+}09, GGB^{+}08, TC08, VML06, YDL^{+}06, YEL^{+}08]. Chromosome [BLvH^{+}05, Com07, DBK^{+}07, Pow06c, DSK^{+}06, GLH^{+}05, HBPL09, iHBD08, HH06b, JWC^{+}08, JMB^{+}07, KKM09a, KCH^{+}06, MBdR^{+}08, MLS^{+}07, MFH05, MPLF^{+}09, MFC^{+}09, OOS07, OOS09, PMK^{+}07, Pow05e, RSM^{+}08, RP08, SLP^{+}06, TTO^{+}08, WWK^{+}06, WWvD07, XOV^{+}09, YRSM^{+}09, ZAM07a, ZAM07b]. Chromosomes [LeB07i, Les07l, BJKF05, Bon05j, CM09, HH06a, KSL^{+}05, LeB08n, Les09p, Sed09m, WH06, ZRBS07]. chronic [Wel06s]. Chs2p [ZKN^{+}06]. Chs3 [LDS^{+}06]. cIAP1 [VCC^{+}08]. CIB1 [LLJ^{+}05, YLM^{+}06b]. Cilia [QP05, Wel05i, Wil07h, Cho09a, HRK^{+}09, HDR09, KAD^{+}08, Les07x, Les08-65, POCS^{+}06, Sch08, SBW^{+}07, SR07, VTT^{+}09b]. ciliary [AFV^{+}05, LDR^{+}08, Mar08, Smi07]. ciliated [ESG^{+}06c, VS07]. ciliogenesis [FFW^{+}07, POG^{+}09, SGB^{+}06]. cilium [DMC^{+}07, GSL^{+}07, JNHB08, MSR09, PALW05, SSD^{+}09, STM^{+}07, iYEF^{+}07]. CIN [RP08, YD08]. CIN85 [JSC08]. Cin8p [CGJ09]. Cindr [JSC08]. circle [HGJ06]. circuit [SJL^{+}07]. circuitry [HR09]. circular [HRLH08]. Cirque [Ban09]. cis [KVS^{+}09, LLC^{+}08]. cis- [KVS^{+}09]. citron [GNL^{+}06]. CK1 [GWS09]. CK2 [BFB^{+}08, TSP^{+}07, Wil08i]. Cl41 [LCT^{+}09]. CLASP [CBZY08, HH06a, KLG^{+}09, Sho09z, WWS05b]. CLASP1 [MKGL^{+}05]. CLASP2 [MKGL^{+}05]. Claspin [FBJB^{+}09, YLM06a]. Class [MCF^{+}05, ZMGC05, AFN^{+}08, CMF^{+}09, JHY^{+}08, LK05, ST09, STvdB^{+}09, VLFA^{+}07, GTO^{+}07]. Classical [DCEM09]. clastosomes [JMM^{+}06]. Clathrin [DPK^{+}05, EBMG09, BHH^{+}06, DBW^{+}06, FN06, KKM^{+}08, LPB^{+}07, LTAM^{+}09, SDDV05, SSM^{+}08, SCK^{+}07, WMT07, LRCH^{+}05]. Clathrin- [DPK^{+}05, SCK^{+}07]. clathrin-coated [BHH^{+}06, SSM^{+}08]. clathrin-dependent [DBW^{+}06, LTAM^{+}09, WMT07]. clathrin-independent [SDDV05]. Claudin [Sed08d]. Clean [Wil07i, Kat06e, Les08o, Wel05-69, Wil08e]. Cleaning [LeB08d]. cleans [Sed08l]. clear [LeB08-27, Sed08f, Sho09x, Wel06w]. clearance [FSR^{+}07, YCR06]. clears [LeB06-32, Rob08c]. Cleavage [VDA^{+}08, YHC^{+}06, ACsHS08, AAG^{+}09, BBvD05, CVACF07, GG06, GT08, GKvdB07, HOH^{+}07, IDTT^{+}09, LeB05-74, Les08a, LVDR05, MDC^{+}08, NSO^{+}09, OF08, Rob08v, TSN^{+}07]. cleavage-independent [CVACF07]. cleaves [TP09, YIfHA09]. cleft [CBG^{+}06]. cleft-associated [CBG^{+}06]. client [SZL^{+}09]. climb [LeB09c]. CLIP [DvCvH^{+}08, Wit08, BKLT^{+}08, LHL^{+}08]. CLIP-170 [BKLT^{+}08, LHL^{+}08]. Clipping [Tum05j]. CLIPs [Wit08]. clock [Kat06a, LeB05-28, Les07-59, Les08-60]. Clonal [Wel07c]. cloning [Wil08u]. close [Les06v, Les07-35, RW08b, RW08e, Sed07e]. closed [ABvBHH06, NRD^{+}06, Wil06k]. closely [SCR^{+}09, WBLL06]. closes [LWL08a, LWL08b]. closet [ZHS06]. Closing [Les07m]. closure [STK^{+}05, TPNS07a, TPNS07b]. clothing [Wil07-68]. Clp1 [CWRG^{+}08]. Clp1/Cdc14 [CWRG^{+}08]. ClpX [Wel05j]. cluster [CMK06, KIKI^{+}07, SFS^{+}07, XC08]. Clustering [Wil07j, CSL^{+}05, JBL^{+}07, LeB07y, SMH^{+}09, TJC06, TSALS05]. clusters [HBD^{+}05, Kat06c, MMH09, Nyb05i, SFEK07, SFS^{+}07]. CMT2A [DC07]. CNP [LGZ^{+}05]. CNS [DG08, WTE^{+}08]. co [MQWP08]. co-activator [MQWP08]. Coassembly [DDRL^{+}09]. coat [GL06, Heu05, LPB^{+}07, LDPLS05, LRCH^{+}05, WHOS06]. Coated [Pow05f, BHH^{+}06, SSM^{+}08]. Coatomer [CFL^{+}05, MSM^{+}07, RCT09]. Coatomer-bound [CFL^{+}05]. cochaperone [CCC^{+}06]. Coddling [Les07n]. code [RAS^{+}07, RAS^{+}08b, Sho09c, Wil08k]. codes [MIBT06]. cofactor [MRL^{+}08, STL^{+}07]. Cofilin [FBD^{+}08, KOKM07, OD07, Sed08e, SWM^{+}07, BKBM06, CM06, KCMB08, LeB06-53, NTY^{+}05, OYM^{+}09, WMS^{+}06, vRSvR^{+}07, vBDF^{+}09, DvRF^{+}05b, WHB^{+}07]. Cog3p [ZL05]. cognate [CCS^{+}06, KFK^{+}05]. coherent [MS06]. Cohesin [AJR^{+}09, LBVN09, LeB06j, NWR^{+}08, Sho09p, HCK^{+}07, KKGD05, NAF09, TSALS05, YK07, ZKS^{+}08]. cohesin-dependent [KKGD05]. Cohesinopathy [GLX^{+}09a, GLX^{+}09b]. Cohesins [Tum05k, DSK^{+}06]. cohesion [BSW^{+}05, CS09, HS09b, RSM^{+}08, TKSNG07]. coil [BBF^{+}07, BLC05, HW09, KHWC08, SGKM08]. coiled [BBF^{+}07, BLC05, KHWC08, SGKM08]. coiled-coil [BBF^{+}07, BLC05, SGKM08]. Coilin [Les07o]. coincidence [Sed08e]. coli [Kat06d, BM07]. collaborate [LeB08r]. collaboration [BJC^{+}06]. collaborator [Les09-33]. Collagen [SFC^{+}08, CEY^{+}05, LeB05-102, RG06a]. collapse [RW08d]. collateral [EW08]. collecting [NIC^{+}09b]. Collective [KBR^{+}08, Pow07e, LCL^{+}07]. colon [GCSG^{+}05, OMLP^{+}08]. colonies [Tum05-53, VP05]. color [Wil07-55]. combination [DSN^{+}07b]. combinations [BJZ^{+}09]. combinatorial [KLW^{+}05, ZMA^{+}07]. combined [HWJ^{+}08]. combines [KHWC08]. come [Les09o, Rob08r]. comes [Les07q, Ped08, Sho09m]. comet [BKBM06]. commitment [SQB^{+}08, VWM^{+}07, dPMF^{+}09]. committed [GNE^{+}07]. committee [Ros07b]. common [RMV06, Wil09s]. communicate [HdM09]. communication [BW08, vZPG^{+}07]. compaction [DSK^{+}06, GTB^{+}07, LJS^{+}09b, MJF^{+}07]. Comparative [BHH^{+}06]. compartment [DBvdW^{+}06]. Compartmentalization [vRCB^{+}07, BMK^{+}07, LFS^{+}05]. compartmentalized [FML^{+}06, TDP^{+}06]. compartments [AWM^{+}08, HR05, LOG^{+}06, SHX^{+}07, Tum05-80]. COMPASS [NLG^{+}09]. Compensatory [WLLF^{+}08]. compete [Wil07t]. competence [KS08, YTB07]. competent [YTY^{+}07]. competes [SRD^{+}07]. competing [KII^{+}08]. competition [TTNA^{+}06]. Competitive [WOP^{+}06]. Complementary [Rob08i]. Complementation [DC07, GSHD08]. Complete [GSB^{+}06a, IBS^{+}05, GHY^{+}05, XMA^{+}09]. completion [DHR^{+}08]. complex [AXM^{+}05, ARKS^{+}08, ASH^{+}09, AHM^{+}05, BSA^{+}08, BBF^{+}07, BKE^{+}07, BFCS^{+}07, CNW^{+}05, CSSE06, CLR^{+}06, DGG^{+}08, DXC^{+}09, DTW06, DS07, ERM^{+}09b, FKJ^{+}09, FN06, FDV^{+}06, GMW^{+}05, qGyWX^{+}08, GEP08, GFH^{+}06, GMvdKC07, GMM^{+}08, GTHK05, GDR^{+}05, GMZ^{+}08, HSLI09, HJ06, HRB^{+}06, HRK^{+}09, HRS^{+}05, HZR05, HPH^{+}05, KCAS05, KAL^{+}06, KCH^{+}06, KBGG^{+}08, KCOS06, LU05, LH05, LBI^{+}08, Les05f, LPB^{+}07, LDL^{+}09a, LFL^{+}05, MMCW06, MFS^{+}09, MSL^{+}09, MMC08, MXR^{+}05, MRO^{+}05, MQL^{+}06, NWB^{+}06, NRD^{+}06, NdCF^{+}05, NNS^{+}08, OHK07, OMF^{+}05, OSM^{+}09, POGL08, PSL^{+}06, PNO^{+}09, PMC08, PGO^{+}07, Rap05, Rex09, Rob08-30, Rob08-44, RKH^{+}09, RXA^{+}08, RWM^{+}05, SRA^{+}05, SYM^{+}08a, SYRL^{+}08, SB09, SWT08, SSP^{+}07, SBMP05, SMM08b, SRD^{+}07, SMT^{+}08, SNPCG06, SGKP^{+}07a, SGKP^{+}07b, SMM^{+}08c, SKO^{+}06, SLNS06, THY^{+}06, TW07, VML06, VBG^{+}08, VBSM06, VCSG07, VCC^{+}08, WZO^{+}05, WHOS06]. complex [WKCY^{+}08, XOV^{+}09, YM09, YDL^{+}06, YPG05, ZWG^{+}05, ZKS^{+}08, ZKG^{+}08, ZCJ^{+}08, ZMH^{+}09, deC09]. Complexes [LeB05s, BKHR05, BLH^{+}09, BBM^{+}08, BVS^{+}09, DLH^{+}09, ELD^{+}08, JSPT05, hKDH^{+}08, KGM^{+}07, LeB05-82, LPM^{+}07, MWSK06, OFfL^{+}07, PHB^{+}07, SHC07, SRBL06, SM09, ST09, SHS^{+}06, TDD^{+}05, Wel05l]. complexins [RWB^{+}05]. complexity [ZB09]. component [CSSE06, EFCP07, FKJ^{+}09, FTM^{+}06, GGC05, POG^{+}09, RDY^{+}05, SNC^{+}06, SORM05]. components [CBG^{+}09, KPR^{+}07a, KPR^{+}07b, PPA^{+}08, SWT^{+}05, XGX^{+}09]. composite [BKLT^{+}08]. composition [CCH09]. comprehensive [YCX^{+}08]. compressive [BMK^{+}06]. compromised [WBK^{+}06]. Computing [Wel06e]. Con [Bea06]. Con-nectin [Bea06]. Conboy [Pow09a]. concentration [EMPK^{+}07, KRP^{+}09, YM06]. concentration-dependent [KRP^{+}09]. concentrations [YKBR09]. Concentric [BB05]. concept [HK08, KSpLL06]. concert [TMH^{+}06, WPWF08, ZKR^{+}07]. Concerted [GJL^{+}08, LWL08a, LWL08b, POCS^{+}06]. Condensation [GCF^{+}05, BLMS06, LeB05-40, MTRJA05, McN09]. condensation-mediated [MTRJA05]. condense [Kat07e, Kat07f]. Condensed [CDW^{+}05, HKP^{+}09, Tum05y]. Condensin [BLMS06, NNK^{+}08]. Conditional [JLG^{+}06, SPA^{+}05, YCX^{+}08]. conditions [WGL^{+}06]. cone [FSN^{+}06, HHT^{+}07, LeB06-66, Tum06q]. cones [Bon05d, LeB07-44, OFY^{+}05, WHB^{+}07]. Conference [Wel06o]. confers [DLBSD^{+}05]. configuration [OIAT06]. confined [FFS06]. conformation [WCR^{+}05]. Conformational [oSCW07, ABvBHH06, CHSV07, HM08, PKK09, SWS^{+}08]. conformers [NRD^{+}06]. congression [MSB06, RSM^{+}08]. conjugated [MAK^{+}06]. conjugating [AZP^{+}05]. conjugation [HDR09]. conjunction [HEW07]. connect [JMG^{+}06]. connected [AWM^{+}08, BMT^{+}07, Les09t]. connection [BP05, Hor09, LeB05-66, Les07m, Les09n, LR07, Rob08x]. connections [Les08-31, LLT^{+}08, RH09]. connects [HCL^{+}07, RG06b]. connexin [Les07m]. connexin43 [vZPG^{+}07]. Consensus [KBI^{+}08, SBC^{+}06]. consequences [WBP^{+}08]. consequently [MBdR^{+}08]. Conservation [Mat05]. conserved [ASH06, ADD^{+}06, BFWH07, BMGPB^{+}08, CBIO05, DS07, GD05, HOZ^{+}09, IH05, JBF^{+}08, LFMO^{+}09, OHP^{+}09, POG^{+}09, RDY^{+}05, SDI^{+}06, SSP^{+}07, SSS^{+}08b, SDW^{+}09]. consolidation [LLL^{+}06, RCS^{+}09]. conspiracy [Les06-29]. constituents [FTT^{+}05]. constitutes [IFF^{+}05]. Constitutive [SMT^{+}08, BXN^{+}06, FNH^{+}05, KWU^{+}05]. constriction [SHS^{+}09]. Constructing [Wil07k]. construction [BW06, LeB06l, NAS08]. consumption [Wel06-34]. Contact [CCL^{+}07, NPHT08, BvRJ^{+}05, FIK^{+}05, FIK^{+}20, OLLG05, WZCB^{+}06]. Contact-dependent [CCL^{+}07, NPHT08]. contact-induced [FIK^{+}05, FIK^{+}20]. Contacting [Pow05g]. contacts [BBM^{+}08, IFF^{+}05, KKW^{+}09, MCSY07, MFS^{+}09, RKvdK^{+}09]. contain [PH07, SGKM08]. containing [AHM^{+}05, CSSvM08, DPMB^{+}07, HFF07b, JFM^{+}09, KBI^{+}08, LBG^{+}06, MHM^{+}07, CHA05, HAT^{+}05b, hKDH^{+}08, KCOS06, LPB^{+}07, YOS^{+}05]. content [Hil07, LeB06b]. context [SN05, Sho09-40]. context-specific [SN05]. contiguous [SLP^{+}06, TDP^{+}06]. continued [KYK^{+}06, SFB09]. continuous [PVJJ07, SWT08]. Contractile [Tum06i, BLL^{+}06, CHCM06, JTH09, KOM07, NAS08, RGCWG09, SWI06, WSK^{+}06]. contractility [BRH05, SGMPV07]. contraction [CCL^{+}05, RG06a, WRMH07, dRKD^{+}05]. contractions [Pow05x]. contracts [Rob07j]. contrasting [LWG^{+}05]. Contrasts [LeB05t, OF08]. contribute [BJZ^{+}09, NAS08, PPA^{+}08, TMW^{+}05, ZRBS07]. contributes [ADC^{+}07, BBF^{+}07, BFB^{+}08, BWFdL09, CWRG^{+}08, FDV^{+}06, FMD^{+}09, GSdBS09, MDC^{+}08, MSO09, RKT^{+}07, SFY^{+}05]. Contribution [YHD^{+}09, SDL^{+}05]. contributions [OLKK09]. Control [GCS^{+}07, SRL05, AS06, BFG^{+}06b, BHG^{+}05, CFU^{+}09, CN06, DvRF^{+}05b, DTR^{+}05, DLH^{+}09, DCEM09, ELM09, ESZK06, EMMN^{+}05, GMW^{+}05, GC05, HBK^{+}06, HK06, HO08, HAT^{+}05b, HDS^{+}07, HG07, JBYW06, KBW^{+}06, KR06, KKW^{+}07, KDG^{+}09, LeB06-72, LW07b, Les08-54, Les08-58, Les09d, LFMO^{+}09, LLC^{+}07, MLC^{+}07, MAZ^{+}08, NCB^{+}08, NCB^{+}09, NN07b, NMZ^{+}09, OLS08, OYM^{+}09, OHP^{+}09, PGWH08, PCPW08, Rob08i, Rob08-37, RKvdK^{+}09, SMH^{+}09, SRN^{+}07, TKK^{+}05, Tum05-34, Tum06-32, TTE^{+}07, WIL^{+}05, WH06, WRP^{+}06, Wel05g, Wel06t, Wil08b, YBT^{+}05, YNL^{+}08, YWW^{+}08, YWRA07, ZTKK05, dPMF^{+}09]. controlled [BJLM^{+}05, MVC^{+}07, MFBG^{+}08, OMGW^{+}06, SYM^{+}08a, XMSYJ07]. controller [Vau05, Wil08i]. Controlling [Tum05l, Tum05m, EMB^{+}08, FA08, GH05, LR05, SWM^{+}06, TM05]. controls [BMK^{+}07, BMGPB^{+}08, CLL07, DXC^{+}09, EMPK^{+}07, FSG^{+}07, FKML^{+}06, GG06, GPC^{+}06, GFD^{+}05, HFF07b, HWS^{+}07, JKDH08, JWC^{+}08, JLM^{+}07, JLL^{+}07, KKGD05, KP08, KLRS07, LPA^{+}07, LOG^{+}06, LeB05a, Les06f, Les08b, LCB^{+}08, MLV09, MCRVC^{+}08, MAOM06, MRO^{+}05, MKSJ06, NIC^{+}09b, PAV^{+}06, PBD^{+}06, RHMW^{+}05, SGB^{+}06, SCD^{+}08, SRS^{+}09, SHB^{+}07, SCF^{+}07, Tum05t, Tum05-59, VTT^{+}09a, Wel06-45, WKC^{+}09, WAW08, WZCB^{+}06, WF06, WF07, ZHI^{+}08]. Controversy [SS06a]. convergence [BMR^{+}08, LSC^{+}05, RG06a, TMI07, YKP^{+}08]. convergent [KHH08]. converter [Wil06e]. converting [PCR^{+}07]. converts [KRP^{+}09]. COOH [IPAR05, SDS^{+}05]. COOH-terminal [IPAR05, SDS^{+}05]. cooperate [BCJ^{+}07, FSS^{+}09, MCSY07, RGCWG09, Rob08z, SMH07, ZAM07a, ZAM07b]. cooperates [KHH08, LMHT08, LKL^{+}05, QU08]. Cooperation [KGM^{+}07, YKP^{+}08, BHJ^{+}05, IPAR05, LeB06-38, NIC^{+}09b]. Cooperative [LeB05u, Tum06j, MSK05, SFC^{+}08]. cooperatively [SOW^{+}09]. Coordinate [YBT^{+}05, DHT^{+}09, HSLI09, LeB05s, LeB08-31, LSNY08, LHM^{+}09, MPG^{+}09, SWD^{+}08]. Coordinated [MM05, NCB^{+}08, NCB^{+}09, SSM^{+}08, BPL^{+}08, OHP^{+}09]. coordinately [DRD^{+}05, FvD08, HF06]. coordinates [AWR^{+}07, CCL^{+}08, HCN06, HPH^{+}05, LHL^{+}08, LSIT^{+}08, NWO06b, THS^{+}09]. Coordination [ST09, TSS^{+}06, FKB^{+}05]. cop [Les09q]. COPI [ARKS^{+}08, CFL^{+}05, LYH^{+}05, LDPLS05, WBE^{+}08]. COPII [BHG^{+}05, GL06, YSP^{+}06]. Coping [LeB06k]. copying [Les09z]. cord [HFY^{+}06, Les08m, SSJ^{+}07]. core [DR07, ECR^{+}09, HWSJ08, HZR05, LPF^{+}06, LWL08a, LWL08b, SHL^{+}09]. core-binding [LPF^{+}06]. coreceptor [KDV^{+}09, WZC^{+}08]. coreceptors [Tum05j]. Coregulated [BLR^{+}06]. Coregulation [SBF^{+}06]. corepressor [SLMK08]. coronin [BKBM06, KCMB08]. correct [BH08b, HFF^{+}07a, LeB06o]. Correction [ASB^{+}05a, AM05a, BFG^{+}06a, BC05, CRS^{+}06, DTY^{+}05a, DvRF^{+}05a, DSN^{+}07a, ESG^{+}06a, ESG^{+}06b, FBD^{+}05a, FCB^{+}19, FIK^{+}20, GSB^{+}06c, GSB^{+}06b, KMK^{+}05a, MSW^{+}17, NWO06a, NYI^{+}05a, PWM^{+}06b, RMZ^{+}07, SPT^{+}21, SFY^{+}06, SDN^{+}05a, TÖL^{+}06a, TFC^{+}06a, Tum05n, WIL^{+}06a, lWHMH06a, Wel06f, Wil06d, WWS05a, WNF^{+}05a, WTH06, YWW^{+}05]. corrects [CGGT09]. correlate [PWD^{+}06]. correlates [GSJ^{+}08]. corridors [LeB06-48]. Cortactin [OYM^{+}09, SDDV05, LMM^{+}09]. Cortex [LeB06l, CHCM06, IMY^{+}09, MKGL^{+}05, RYN05, SON06, VSG09, yYMM05]. cortical [BSF^{+}08, CKG^{+}08, HYB08, JTH09, JWSS05, KO07, LYTL07, SSK^{+}09, UFY^{+}07, WMSB^{+}06]. Cosmc [Wil06e, JASC08]. costameres [KFR^{+}08]. costume [LeB08v]. cotranslation [CSTT^{+}06]. cotranslational [CJMG05, JCMG08]. cotransport [MTL^{+}06]. Could [Wil07l]. count [Rob08-47]. counter [SWT^{+}05]. counteracting [GFD^{+}05]. counteracts [FBJB^{+}09, MRKH05, WM05]. counterdiffusion [DPD^{+}09]. counterparts [BRY^{+}09]. Counting [Tum05o]. couple [DP07, PRZ^{+}08]. Coupled [SS09, NVC05, TSP^{+}07, YOY08]. couples [DGSN06, DTW06]. Coupling [CLK^{+}07, CLR^{+}06, AFV^{+}05, CCL^{+}08, DFA05, Joh09, LYH^{+}05, MAG05, SHH^{+}08, SBV^{+}06, TDT^{+}08, WBG^{+}08, WvdLH^{+}07]. course [LeB06-53]. court [Sed08m]. CPG [OBY^{+}06, BEN^{+}05, SRL^{+}06]. CPG-1 [OBY^{+}06]. CPG-2 [OBY^{+}06]. CpG-induced [SRL^{+}06]. CR3 [LHL^{+}08]. CR3-mediated [LHL^{+}08]. CRAC [LWBL06]. Cracking [PvdGS06]. CRAG [QIH^{+}06]. crazy [Wil07i]. CRB [OFfL^{+}07]. CRD [SSSL05]. create [Pow07i]. creates [SFB09, VMKW^{+}08, Wil07j]. CREB [WBG^{+}08]. crest [KLW^{+}05, TMW^{+}05, WPDZ^{+}06]. crew [Les06z]. Cripto [DLL^{+}08, WNL^{+}09]. Cripto-1 [WNL^{+}09]. crista [RSS^{+}09]. cristae [RSS^{+}09]. Critical [GXJF07, PMS^{+}07, ASSST06, BZHA^{+}08, BHJ^{+}05, JAV^{+}07, JBYW06, JCMG08, LGG^{+}05, NLG^{+}09, NTY^{+}05, PUG^{+}06, RKH^{+}09, Wil07m]. critically [BGZ^{+}07, SCFKB07]. Crk [SMM08b]. CrkII [HWJ^{+}08]. CrkL [NNS^{+}08]. CRM1 [HFvdV^{+}05, HAT^{+}05b]. CRM1-dependent [HAT^{+}05b]. CRMP2 [RP05]. cross [BPL^{+}05, CBMCSS08, CTY^{+}06, KKW^{+}08, LC09, Rob08h, YTH^{+}09]. cross-linked [CTY^{+}06, LC09]. cross-linkers [CBMCSS08]. cross-linking [KKW^{+}08]. crosses [GDW^{+}07]. Crossing [NVC05]. Crossovers [LeB05v]. crowd [LeB05-38, Les08-54, Nic09a, Rob08-28]. crucial [BSH^{+}07, iHBD08, LW07b, SHS^{+}06, ZXM^{+}06]. Crumblin’ [DP08]. crumbs [LWHC08, LWHC08]. Crumbs3 [FFW^{+}07]. cryo [MWSK06]. cryoelectron [NPBP07, OHK07]. Cryopyrin [LMV^{+}09]. Cryopyrin/Nalp3 [LMV^{+}09]. cryptic [GHM^{+}09]. Crystal [XMA^{+}09]. Cse4p [HGJ06]. Csk [VSR^{+}08]. Csk-binding [VSR^{+}08]. CSP [OVL^{+}07]. CTD [CVACF07]. cten [LSdWL07]. Cu [MSH^{+}05]. Cu/Zn [MSH^{+}05]. Cubic [AKD06]. cue [Les06h]. cues [HP05, LeB05-54, ZMA^{+}07]. Cul3 [MOS^{+}09]. culture [KAD^{+}08]. cultured [WWL07, XLH^{+}07]. cultures [ABA^{+}06, UFY^{+}07]. cup [BMG^{+}07, SS08b, vBAS^{+}07]. Curbside [Pow05h]. Curing [Wel06g]. curved [MGBM^{+}09]. Curves [Wil07n]. cushion [NBM^{+}08, NFC^{+}08]. customers [LeB08c]. cut [LeB06t, Les08-66, Wil07p]. Cut12 [TTYH09]. cuts [LeB07-27, Sho09-29]. Cutting [Les08m]. Cx43 [SMRS^{+}07]. cyclase [SSJ^{+}07, TDT^{+}08]. cycle [BTD^{+}07, BP05, BAdSG05, BvOM^{+}07, BMB^{+}07, CGLC^{+}08, CSW08, CWC^{+}07, DVF^{+}06, EMB^{+}08, GAS06, GGH^{+}08, HXZ^{+}07, JGAS08, KR06, LeB06i, LeB06-44, LZM^{+}06, Les07-62, LHM^{+}09, LTDH05, LG07, LG08, MYO^{+}05, MPL^{+}07a, MSO09, MAG^{+}08, MMU^{+}09, NPMZ06, PMS^{+}07, PMW05, PWD^{+}06, Pow05x, QP05, RMAA08, RP07, Sho09l, SD07, SGDM06, SWM^{+}06, TWC^{+}07, Tum05-35, Tum06z, Tum06-30, ULN^{+}07, VDSV06, WZO^{+}05, Wil07-56, WWK^{+}06, YR09, ZTZ^{+}09]. cycle-dependent [EMB^{+}08, GGH^{+}08]. Cycles [LeB05w, GHP06, LeB05c, Mis05, Sed08n, Wil09o]. Cyclic [BAdSG05, OFY^{+}05, PBD^{+}06]. Cyclin [KST^{+}07, Les06f, CTS^{+}05, KZH^{+}07, LKL^{+}05, MHL^{+}06, MFO09, PLH^{+}08, RMKS08, TWC^{+}07, YWW^{+}08]. cyclin-dependent [PLH^{+}08]. Cycling [Les07p, CHA05, KBK^{+}05, TM05, Tum05-60]. Cyclodepsipeptide [SZL^{+}09]. cycloheximide [Wel07l]. Cyk3 [NSO^{+}09]. CYR61 [TCHL05]. cysteine [TRS^{+}08]. cysteines [SPK^{+}09]. cystic [JSFP08]. cytoarchitecture [OMGW^{+}06]. cytochrome [SMWdV09, CLA^{+}05]. Cytokine [KBK^{+}05, TPA^{+}08, LeB07-45, LHRP05, SRL^{+}06, SDDV05]. Cytokine-driven [KBK^{+}05]. cytokinesis [AHM^{+}05, BBFN07, CWRG^{+}08, DHR^{+}08, GCM09, GTHK05, GT08, GNL^{+}06, HO08, HCFM08, JD06, LeB05-86, LeB07-65, LB08, MOS^{+}09, PWW^{+}09, Pow05q, RBJL06, Tum05e, VSG09, vDVM^{+}09]. Cytokinetic [Can09, CGK07, CAR^{+}08, FvD08, WSK^{+}06]. Cytomegalovirus [Les09l, PPS^{+}06]. cytomegalovirus-encoded [PPS^{+}06]. Cytopathic [PPS^{+}06]. Cytopathicity [PCV^{+}08]. cytoplasm [CLR^{+}06, MDC^{+}08, MHC^{+}09]. Cytoplasmic [BKMW08, PFG^{+}05, AFM^{+}05, CHA05, CJMG05, HLM^{+}08, HFM^{+}05, HCC^{+}07, LSDFS08, LPN^{+}08, LGBD^{+}07, MSB06, Rob07h, SNC^{+}06, SDS^{+}05, VDSV06, WHC^{+}07, WZM^{+}08, WBT^{+}08, WPL^{+}06, ZWO^{+}09]. cytoskeletal [GQSHO05, RH09, SSZR07, WLK^{+}05, YHJ^{+}05, ZRH^{+}09]. Cytoskeleton [BBM^{+}08, CJW07, DBDD06, JSC08, LeB07-61, LCS^{+}09, MPD09, NAS08, Pow05y, RKH^{+}09, SHS^{+}09, TSP08, TSS^{+}06, WWL07, ZJH08]. cytosol [vKST^{+}07]. Cytosolic [MuMI^{+}06, RMP06]. cytostatic [DUR^{+}05].

Cytotoxic [WSS^{+}06]. Czech [LeB06-52].

D [BKHR05, CCL^{+}08, DNG^{+}06, FHS^{+}07b, FKML^{+}06, GdTAR07, GMM^{+}08, HSTM06, LJR^{+}05, LMMHP^{+}05, Les07-69, MRD^{+}08, OMLP^{+}08, VBL^{+}05, Wil09a, XSVGR05]. D-loop [Les07-69]. D-mediated [FHS^{+}07b]. D.C [LW08]. D.C. [LP05]. D1 [CTS^{+}05, KZH^{+}07, YWW^{+}08]. D1/microRNA [YWW^{+}08]. DA-Raf1 [YTY^{+}07]. Dab2 [TTG^{+}09]. Dab2/integrin [TTG^{+}09]. Dachsous [IMY^{+}09]. Damage [LeB06m, BJLM^{+}05, BLMS06, CCS^{+}08, CKKG05, CSMP^{+}09, DCA^{+}06, EW08, HKLM06, HKLM07, KKK^{+}05, KHW^{+}07, KHW^{+}09, LeB05z, LeB05-65, LBS^{+}08, LDL^{+}09b, MBJM^{+}09, MBJF^{+}08, MDWH^{+}08, MRL^{+}08, MJF^{+}07, PGP^{+}07, Rob07n, SS07, SRS^{+}09, SDC09, SBC^{+}07, TRC^{+}08, WH06, WCB^{+}08, Wel06-46]. damaged [SMT^{+}08, SiTT^{+}08]. dance [Wil08m, HOH^{+}07]. Dances [LZ08]. danger [PGP^{+}07]. Dap160 [KKW^{+}07]. DAPK [KWY^{+}06]. dark [EFG^{+}05]. Darkness [Mar06]. data [RVH07, Sed08i]. DataViewer [Hil08]. Daughter [Les08n, GT08, ZKN^{+}06]. David [Sho09q]. DBA [HLJ^{+}07, JLB^{+}05]. DBA/2J [HLJ^{+}07, JLB^{+}05]. Dbf2 [MHC^{+}09]. Dcp [HCB^{+}08]. Dcp-1 [HCB^{+}08]. DDA3 [JWC^{+}08]. DDR1 [FKML^{+}06]. DE-cadherin [PR05]. Deacetylate [LeB05x]. deactivation [TD08]. Deadenylation [ZEC^{+}08]. deadly [Les06-29]. deal [Pol08]. dear [Les06l]. Death [LeB07j, Les07q, Tum06k, Wil07o, ASSST06, Ash08, BLB^{+}05, BLMS06, BEH^{+}06, CMD^{+}08, CLL07, FAG^{+}07, GH05, GHM^{+}09, HZL^{+}07, HlQY^{+}06, IH05, JMB^{+}07, LeB05-56, LeB05-55, LeB07-54, Les06b, Les07g, Les07u, Les09e, LCL^{+}07, MDET08, MDH^{+}06, NDS^{+}07, Nyb05c, PKM^{+}05, SFP^{+}05, THL^{+}06, TP09, TGRD09, Tum05-53, Tum06-42, VP05, WLE09, WZM07, Wel05-55, Wil06c, Wil09h, Wil09l, XJF05, YTB07]. Death-defying [Wil07o]. decay [DCMB^{+}07, FBD^{+}05b, HLL^{+}09, LGBD^{+}07, ZEC^{+}08, vRG08]. December [LK06, LW08, WL07, dSW09]. decide [Rob08-47, Sho09y, SMR09]. decision [LRBM09, Wel06-42]. decision-making [Wel06-42]. decisions [Wil07f]. decodes [LeB06-57]. Decoding [Mua05, vRG08, RSL06, Wil08m]. Decondensation [Tum05p, AH05, WLS^{+}09]. Deconstructing [YCX^{+}08, Wil08f]. Decorin [GHN^{+}09, Sho09r]. decrease [PCG^{+}05]. Decreased [GHK^{+}09, TRC^{+}08]. decrepit [Wel05d]. dedicated [HSK^{+}05]. Deep [Heu05, LeB07-61, Sho09u]. Deep-etch [Heu05]. deepest [HB08]. deeply [CB08]. defasciculation [YBT^{+}05]. default [BHY^{+}08, LeB07j]. Defaulty [LeB06n]. defect [dGSWW^{+}06]. Defective [GHDS^{+}07, JBL^{+}07, MLH^{+}06, HFF^{+}07a, SNG^{+}07, SBC^{+}07, THN^{+}06]. Defects [MLS^{+}07, CRLP07, CC09, DNK^{+}05, DC07, DSN^{+}07b, EVF^{+}08, KHB^{+}05, MDL^{+}09, OFfL^{+}07, PhXD^{+}06, RMR^{+}09, RSM^{+}08, SCFKB07]. defense [SBC^{+}07]. Deficiency [WHM^{+}07, YZO^{+}07, BGJ^{+}07, EFM^{+}07, HTKR06, KPS^{+}07, SMM^{+}08a, Sed09f, SVS^{+}07a, SVS^{+}07b]. deficient [BLL^{+}06, EYB^{+}06, KR07, KWU^{+}05, LHS^{+}05, MBL^{+}07, SNG^{+}07, SNPCG06, WBL^{+}06, FPP^{+}08]. deficits [HZL^{+}07, HJK^{+}06]. define [FHS^{+}07a]. defined [KZS07]. defines [COB^{+}08, MCF^{+}05, PFD06, SRQL09, ZRH^{+}09]. Defining [HLS^{+}09, SBJ^{+}07, Wel05k, Wel05l]. definition [ELD^{+}08]. deform [MPS^{+}07]. deformation [BHG^{+}05, HRLH08, TSS^{+}06]. defusing [Les07u]. defying [Wil07o]. degeneration [ASK^{+}09, BFL^{+}05, EGR^{+}05, JLB^{+}05, LJR^{+}05, WZC^{+}05, WRKC^{+}09]. deglycanation [MBR^{+}06]. degradation [CHO^{+}09b, CEY^{+}05, FCDM^{+}07, GGH^{+}08, LLH^{+}07, LCvD^{+}08, LHZ^{+}08, MLC^{+}07, MG08a, MDW05, OOY^{+}06, Pow06d, QIH^{+}06, RRP^{+}08, SGMPV07, SMM^{+}08a, SS08a, SCD^{+}08, SZL^{+}09, SN05, TdVM^{+}07, TAK^{+}09, TGRD09, Tum05-63, VCC^{+}08, WHR^{+}09, YGW06]. degraded [BLB^{+}05, CBG^{+}09, Les05a]. degranulation [NYI^{+}05b, Tum05q]. dehydration [LeB07-50]. delaying [RMKS08]. delays [Wil06q]. Deletion [HZL^{+}07, JSFP08, JLG^{+}06, dGSWW^{+}06]. deliver [Rob08e, Wil06c, Wil07-37]. deliveries [Les08-44]. delivery [CSB^{+}06, KP08, KBR^{+}08]. Delta [HBK^{+}06]. Deltas [Les07-61]. DeMali [Wil08p]. demonstrates [BFWH07, PGWH08]. Demonstration [KGM^{+}09]. Dendrite [Ehl05, HZR05, HFF07b, LeB06-71]. dendrites [Bea06, LeB07r]. Dendritic [WIL^{+}05, Wil07p, dPMF^{+}09, BMRP07, CHH^{+}08, GTX^{+}06, HLS^{+}09, IDTT^{+}09, LeB07-63, LSC^{+}07, LLHH07, MTH^{+}06, MIBT06, Pow05m, SSD06, TSN^{+}07, Wel05-60, Wel06q, WPC^{+}08, YGW06]. dense [DR07, ECR^{+}09, HWSJ08, LWL08a, LWL08b, SHL^{+}09]. dense-core [LWL08a, LWL08b]. density [GHDS^{+}07, LeB08k, MFBG^{+}08, PKJS^{+}08]. deoxyhypusine [PCMSB09]. Depalmitoylated [GDR^{+}05]. depend [YM06]. dependence [DLBSD^{+}05, GGO^{+}09]. dependency [GEH^{+}07, Sho09n]. dependent [AFM^{+}05, AAG^{+}09, ANI^{+}07, API^{+}07, BSMT05, BM07, BRM^{+}07, BBvD05, BAdSG05, BvOM^{+}07, BRY^{+}09, CBG^{+}09, CNMM^{+}06, CGGT09, CCL^{+}07, CSB^{+}06, DMBF^{+}07, DRD^{+}05, DBW^{+}06, DOS^{+}07, DZG^{+}07, EIK^{+}05, EMB^{+}08, FHNG05, FHO^{+}05, FKB^{+}05, FSP^{+}07, FD09, FPP^{+}08, GdTAR07, GSiK^{+}09, GHDS^{+}07, GLS^{+}09, GPC^{+}06, GPJ^{+}06, GC05, GQSHO05, GC07, GGH^{+}08, GBNS07, HRB^{+}06, HO08, HAT^{+}05b, HLL^{+}09, HCC^{+}07, ITG^{+}08, JKKM08, KRZ^{+}08, KNY07, KKGD05, KRP^{+}09, KGDR^{+}07, KMSLS06, KKW^{+}09, KKB^{+}06, KNP^{+}05, KZH^{+}07, KBC^{+}08, LMM^{+}09, LSNY08, LC08b, LZM^{+}06, LCR^{+}05, LRS05, LCM^{+}08, LPB^{+}08, LG07, LG08, LTAM^{+}09, LFS^{+}05, LCL^{+}05, LGG^{+}05, MWY^{+}09, MDC05, MVC^{+}07, MGBH^{+}06, MBJM^{+}09, MPLF^{+}09, MF08, MG08a, MFBG^{+}08, MTH^{+}06, MK07, NPHT08, NYI^{+}05b, OTM^{+}07, ORM^{+}08, PLH^{+}08, PWL^{+}09, PRK^{+}07, PRK^{+}09, RCS^{+}09, RLDC09, RCFB08, SSHW09, SSD06, SCD^{+}08, SCS^{+}05]. dependent [SGDM06, SSSB07, SCK^{+}07, SWB^{+}07, SLNS06, TKKT07, TMH^{+}07, TNS07, WMM^{+}09, WZC^{+}05, WKH^{+}06, WWX^{+}06, WSS^{+}06, WPC^{+}08, WC07b, WGKA^{+}08, WMT07, WWL07, YGR^{+}08, YRSM^{+}09, YZO^{+}07, ZBKJ07, dRKD^{+}05, MWY^{+}09]. depending [SYM^{+}08b]. depends [CTY^{+}06, KFR^{+}08, LeB07g, OBY^{+}06, RSS^{+}09, Sho09n]. Dephosphorylation [PGO^{+}07, KTA^{+}06, TLDB08, ToMY^{+}08]. depletion [MFC06, WBLL06, WCR^{+}08, ZL05]. depolymerase [JWC^{+}08]. depolymerization [BKBM06, HLS^{+}09, NLL^{+}06, ROKG08, THL^{+}06]. depolymerizes [WDWC08]. depolymerizing [POGL08]. deposited [Hil07]. deposition [HOM^{+}05, LZL^{+}05]. depression [XLH^{+}07, YJJ^{+}09]. deprivation [XJF05]. derepression [BGJ^{+}07]. derived [BSK07, FMM^{+}08, WPDZ^{+}06]. Derlin [OOY^{+}06]. Derlin-2 [OOY^{+}06]. Derlin-3 [OOY^{+}06]. dermal [BFG^{+}06b]. dermomyotome [BYK08]. descends [Mar06]. desensitize [vKST^{+}07]. Desmin [PDM^{+}08, KFR^{+}08]. Desmoglein [GSiK^{+}09, Sho09s]. Desmoplakin [GHA^{+}05, LF07]. desmosomes [Tum05-46]. despite [MPD09]. destabilizing [ZHYE08]. destination [BHY^{+}08]. Destroying [Wel07d]. destruction [Sho09d, Wel05-45, Wel07i, Wil06q]. detachment [MRKH05]. detail [Rob07k]. details [Sed07b]. detection [LeB06-55]. determinant [BXHD08, NLG^{+}09]. determinants [GOM^{+}09, SRE^{+}05, WSG06]. determination [LS09, OYTO07, Wel05-31]. determine [JGAS08, KDD08, MNH05, SPS^{+}08]. determined [CZM^{+}07, NPBP07, SSZR07, XLH^{+}07]. determines [BRM^{+}07, CLA^{+}05, GEH^{+}07, KKM09b, Les06i, NDS^{+}07, NWR^{+}08, SWM^{+}07, TÖL^{+}06b, YBT^{+}09]. deubiquitinating [WLY06]. deubiquitylating [ASH06, MW06]. developers [Wil09f]. developing [BLT09b, LSR07, SLCB05, YJJ^{+}09]. Development [Wel05m, ABY^{+}08, BTP^{+}07, Ehl05, EVH^{+}07, EVH^{+}08, GKCM^{+}08, GPN^{+}06, GXJF07, HZR05, KKW^{+}07, KKM09b, LDKT06, LSG^{+}06, LCB^{+}08, NSM05, NAG^{+}08, PZG^{+}07, SNC^{+}06, SPA^{+}05, SG09, Sho09f, TLM^{+}07, TSN^{+}07, VMY^{+}09, WZW^{+}09, WKC^{+}09, WFJ06, WWC^{+}07, ZHS06, ZMW^{+}08]. developmental [CLL07, LN07, LW07b, Wil08m]. developmentally [PB05]. Dhh1 [BAR^{+}08]. Dhh1/RCK [BAR^{+}08]. DHHC [JKW^{+}09, LDS^{+}06, NFI^{+}09, SSSL05]. Dia1 [BMG^{+}07, Les07-54]. DIAP2 [RKT^{+}07]. diaphanous [SOR06]. diaphanous-related [SOR06]. diastrophic [GS07]. Dicer [TRC^{+}08]. dictate [BLC^{+}09, WCN07]. dictated [FMM^{+}08, OD07]. Dictyostelium [JTH09, MBS^{+}08, PB05, Sed08c, VKGV08, vKGK07]. Did2 [NWO06b]. Diego [WL07]. Diet [Wel05n]. Dieting [Les08o]. Difference [Wel05o, Les07-66]. Differences [AMERT^{+}07, OIFN^{+}07]. Different [RCS^{+}09, BFL^{+}05, BJZ^{+}09, DCAC05, LRS05, LWG^{+}05, Nyb05k, Sho09t, Sho09-30, TDP^{+}06, YMC05]. Differential [CS09, MWMS07, PG06, WBE^{+}08, AMS^{+}05, MKH^{+}07]. differentially [BYK08, KCD^{+}06a, SMG07, WMT07, ZDMD06]. differentiate [LeB05x, ONW^{+}07]. differentiating [LLS^{+}09]. differentiation [AS06, BWL^{+}08, BDP^{+}07, BEN^{+}05, CTGVGFF08, DFB^{+}07, DVF^{+}06, GXJF07, GFH^{+}06, GEN^{+}06, GSiK^{+}09, GLS^{+}09, HPT^{+}07, IPAR05, KBY^{+}08, KLS^{+}06, Las09, LSG^{+}06, LHZ^{+}08, LTDH05, MAH^{+}05, ML05, NVC05, NCB^{+}08, NCB^{+}09, NLC^{+}05, NRL^{+}05, PZG^{+}07, RYD08, SSH^{+}05, SYA^{+}08, SLCB05, Sho09s, SMW^{+}07, SQB^{+}08, SWM^{+}06, Tum06x, WPWF08, WLPB^{+}09, YTY^{+}07, ZHS06, ZXM^{+}06, vdBKB^{+}07]. diffusing [CBMCSS08, MCA^{+}06]. diffusion [BKMW08, HSLI09, Kat07b, LeB06-71]. dimensional [AXM^{+}05, DWMY09, KOM07, MMC08, MFM^{+}06, NPBP07, OHK07, RLH^{+}09, SSHW09]. dimensions [GHA^{+}05]. dimer [CZM^{+}07]. Dimeric [BBPM07, YOY08]. dimerization [MKV^{+}09]. diminishes [BvRJ^{+}05]. dipeptide [BAR^{+}05]. Direct [LeB06o, MS06, PJMR06, RMP06, UPBM06, VMSV08, YLM06a, GOM^{+}09, HWS^{+}07, LSR07, LBC^{+}07, MKH^{+}07, NTHN^{+}08, NFC^{+}08, RMAA08, TDT^{+}08, Tum05i, ZNP^{+}09]. directed [BWFdL09, GNV05, LVL^{+}05]. directing [DP08]. direction [Rob08-28, SMR09, Wel06e]. directional [BRM^{+}07, GGEd^{+}07, KKW^{+}08, LPA^{+}07, NTY^{+}05, SSD^{+}09, SLWA07]. directionally [PEER^{+}05]. directly [FN06, MF08, NN07a, PPCZ06, VW08, WMS^{+}06, YMCT08, YM09]. directs [KWS^{+}08, SDB^{+}05, Sho09s, Sho09-49, Wel06q]. dis [DZW^{+}08]. dis- [DZW^{+}08]. Disappearance [FDV^{+}06]. disassembly [Den05, EBMG09, HDR09, KKB^{+}06, KCMB08, MK07, NVC05, PWL^{+}09, SWI06, WPC^{+}08]. disc [FSN^{+}06]. discoideum [PB05, VKGV08, vKGK07]. discoidin [SFC^{+}08]. discovering [Pow05u]. discovery [Pow05m, Wel05p, Wel05q]. Discussion [Wel06-67]. Disease [MGK^{+}09, BW08, CC09, DC07, FSR^{+}07, GDW^{+}07, JSFP08, JFM^{+}09, KO06, LDKT06, Les07-45, Les08-28, LSC^{+}06a, Mar08, MBL^{+}07, PSL^{+}06, Sho09-29, Tum05x, WWO^{+}08, WLE09, YCK^{+}05, Mis09b]. disease- [KO06]. Disease-specific [MGK^{+}09]. disengagement [NAF09]. Dishevelled [AACS^{+}06]. Dishing [Pow05i]. disintegration [CCL^{+}05]. disintegrin [ADC^{+}07]. disjunction [WWK^{+}06]. disk [PDM^{+}08]. disks [KFR^{+}08]. dislocation [MLP06, STvdB^{+}09]. dismutase [MSH^{+}05]. Disorder [LeB06p]. disorders [IPC09]. dispensable [DPL^{+}08, TLM^{+}07, VCD^{+}06]. displacement [HMR^{+}07, KO07]. displaces [FOB^{+}09]. display [TMY^{+}09]. disrupt [CZP^{+}05, GLX^{+}09a, GLX^{+}09b]. disrupted [JFM^{+}09, LLC^{+}07]. disrupting [BRKR06]. disruption [MPD09]. disrupts [EFM^{+}07, MBdR^{+}08]. Dissecting [WJM05]. Dissection [NNK^{+}08, WJM07, VCSG07, iYEF^{+}07, tDAK^{+}05]. dissociates [NMO^{+}07]. Dissociation [SKR^{+}07, SL05, MTH^{+}06, NWO06b]. distal [YBT^{+}09]. distance [Les08-30, vDVM^{+}09]. Distinct [AZP^{+}05, IGFG^{+}08, IPG^{+}05, KCD^{+}06a, KCS^{+}06, MDWH^{+}08, SRX^{+}07, WL09b, WFJ06, ZWO^{+}09, AVHH09a, AVHH09b, ABSK06, BTP^{+}07, CCS^{+}06, DZG^{+}07, EVH^{+}07, EVH^{+}08, FSS^{+}09, GEP08, GSC^{+}07, GTO^{+}07, HL06, HCC^{+}07, JYZL06, hKDH^{+}08, KSL^{+}05, KFR^{+}08, KKRP05, KDD08, LSC^{+}05, MLC^{+}07, RMZ^{+}06, SH05, TPNS07a, TPNS07b, Tum05-80, TMY^{+}09, WTE^{+}08, WSS^{+}06, WPL^{+}06, WP08, ZTM^{+}08]. distinctive [SZX08]. distinguishes [GEP08]. distress [BJKF05, LeB05-83]. distribution [CZM^{+}07, CCC^{+}05, CBZY08, HKV^{+}05, LMHT08, LLH^{+}09, SLP^{+}06]. disturbs [JBL^{+}07]. Disulfide [COC^{+}05, BMT^{+}07, FSnP^{+}06, SMM^{+}08c].

Divergent [GSC^{+}07, CAL^{+}07, KBS^{+}06, LNG^{+}05]. diverse [NAG^{+}08]. diversity [LeB07-48, Wel07c]. diverts [Pow06m]. Divided [Wel05r, Les07-55]. Dividing [LeB05y, LeB06q]. Division [Les06g, Wel05s, CK07, FFW^{+}07, GHP06, GGBV^{+}07, KNY07, LeB07-29, LeB07-31, Les07b, MAG^{+}08, MH07, Nyb05h, OBY^{+}06, PVMH05, PVJJ07, RW05, RP07, TCKV05, Tum05m, Wel05o, YF08]. divisions [CSW08, LS09]. DIX [Wil07q]. DJ [OLC^{+}07]. DJ-1 [OLC^{+}07]. DLC [LSdWL07]. DLC-1 [LSdWL07]. DLG [OFfL^{+}07]. Dlg1 [EMMN^{+}05]. dLKR [CMD^{+}08]. dLKR/ [CMD^{+}08]. DLL1 [GSC^{+}07]. DLL3 [GSC^{+}07, LNG^{+}05]. Dmoesin [COTAE^{+}05]. DNA [BY09, BJLM^{+}05, BJLK^{+}06, BLMS06, CCS^{+}08, CSMP^{+}09, DMBF^{+}07, DCA^{+}06, DP07, DJV^{+}05, ESZK06, GTB^{+}07, HKLM06, HKLM07, KMK^{+}05b, KG05, KKK^{+}05, KHW^{+}07, KHW^{+}09, KCD^{+}06b, LeB05z, LeB05-59, LeB06j, LeB06r, LeB06-40, LeB08e, Les07-31, Les07-53, Les08p, Les08q, Les08-33, LBS^{+}08, LCM^{+}08, LDL^{+}09b, MTRJA05, MBJM^{+}09, MSO09, MBJF^{+}08, MG09, MDWH^{+}08, MRL^{+}08, MBF^{+}07, MBF^{+}08, MJF^{+}07, PGW^{+}08, Rob08z, Rob08-51, SS07, SSS^{+}08b, SRS^{+}09, SFB09, SDC09, SHK^{+}07, SiTT^{+}08, SBC^{+}07, TRC^{+}08, TFC^{+}06b, Tum05-33, Tum06l, UWiY^{+}07, UPBM06, WH06, Wel05n, Wel05-32, Wel05-33, Wel05-64, Wel06-37, WC07b, YMTC05, YM09, YEL^{+}08]. DNA-activated [MBF^{+}07, MBF^{+}08]. DNA-dependent [WC07b]. Dnm1 [IPM^{+}05]. Dnm1p [GGC05]. DNMT1 [SHK^{+}07]. Do [LeB06s, BYR08, CN06, LeB05p, Pow05l, Rob08t, Wil09s, YKBR09]. Do-it-yourself [LeB06s]. Doa4 [ASH06]. dock [HWSJ08]. DOCK180 [SMM08b]. DOCK2 [KNT^{+}06]. Dock7 [YMCT08]. Docked [LeB05-27]. docking [DDLC^{+}06, YOS^{+}05]. docks [GMK^{+}05, SPK^{+}09]. dodge [Les07-33]. Does [LeB05-28, DP07, LNG^{+}05, LeB05-84, Les09k, Rap05, Tum06d, Tum06-44]. doesn’t [Les08-44]. Dolichol [KBC^{+}07a]. Dolichol-linked [KBC^{+}07a]. Domain [LK05, AFM^{+}05, API^{+}07, BLC05, BMSJ07, DTP07, DPMB^{+}07, GD05, HWN^{+}07, HFF07b, JMG^{+}06, KBI^{+}08, KPR^{+}07a, KPR^{+}07b, KVS^{+}09, LPG^{+}07, LGDN09, LSdWL07, MHM^{+}07, MPS^{+}07, MSSF05, MTB^{+}09, NLS^{+}05, NSM^{+}09, POGL08, RCFB08, RRA^{+}09, RP05, STG05, SFC^{+}08, SAM^{+}07, SDD^{+}07, ShCM^{+}06, SDS^{+}05, TSS^{+}06, VBSM06, VCN^{+}05]. domain-containing [KBI^{+}08]. domain-dependent [RCFB08]. domain-like [MPS^{+}07]. domain-mediated [ShCM^{+}06]. domains [BJML09, BFF^{+}06, BVGA^{+}05, BYMB06, DZG^{+}07, FSS^{+}09, HBD^{+}05, LeB05-32, LeB05-33, MNH05, RGCWG09, SRX^{+}07, SRBL06, SOW^{+}09, SBJ^{+}07, TMH^{+}05, TDP^{+}06]. dominant [LK05, YTY^{+}07]. dominant-negative [LK05, YTY^{+}07]. dominated [SSK^{+}09]. Don [Wil09l]. Don’T [Les08f, LeB06t, Hat05a, Les07-65, RW08d, Wel05-62]. Doom [Les08f]. dormancy [SBC^{+}05]. dormant [TMW^{+}05, WGL^{+}06]. dose [MFO09]. Dot1 [NLG^{+}09]. Double [Rob08-31, BFCS^{+}07, DCA^{+}06, DG05, GTX^{+}06, KG05, KCD^{+}06b, Les07-56, MG09, PGW^{+}08, TFC^{+}06b, UWiY^{+}07, YMTC05, YEL^{+}08]. double-strand [DCA^{+}06, DG05, KCD^{+}06b, MG09, UWiY^{+}07, YMTC05, YEL^{+}08]. double-stranded [GTX^{+}06]. doublet [BSM^{+}09]. Down [YTB07, CLS^{+}06, Les08-48, Les08-66, Les09b, Les09-32, LFL^{+}05, MYMH09, MDW05, Nyb05b, RMC^{+}05, SXH^{+}06, SMRS^{+}07]. down-regulate [CLS^{+}06]. down-regulates [LFL^{+}05, MYMH09, MDW05]. down-regulating [RMC^{+}05]. Down-regulation [YTB07, SXH^{+}06, SMRS^{+}07]. downside [Les09m]. downstream [GFD^{+}05, LLL^{+}06, SYRL^{+}08, SMG07, WCN07, vBAS^{+}07]. DP [MYO^{+}05]. Dpm1p [FKB^{+}05]. DPP [LeB05-29, WLC08]. Drainage [Wil09f]. Dreifuss [MDL^{+}09]. DRhoGEF2 [BRH05]. drICE [RKT^{+}07]. dRip11 [LSR07]. drive [PWW^{+}09, RLH^{+}09, oSCW07, Tum05-35, YN07]. Driven [Les08r, CHCB05, FMK^{+}09, GWS09, KBK^{+}05, LSM^{+}05, MSB06, PWL^{+}09, Wel05-37, TM05, VLFA^{+}07]. Driver [CKKG05]. drives [CMF08, Cho09a, LeB07f, MHC^{+}09, Sho09-45, Tum05b, Wel05o, Wil07r, YKP^{+}08, ZMGC05]. driving [DP06, LSC^{+}05, MFC06, Pow05x]. droplet [EYB09]. droplets [BDW^{+}09, FSG^{+}08, Les08-53, Sho09-31]. Drosophila [AEJ^{+}06, CSW08, DTY^{+}05b, GEE07, HBK^{+}06, HKN09, Mat06, NVEF07, VCD^{+}06, BRH05, BLP^{+}05, BMK^{+}07, BBL^{+}08, CBIO05, DL08, DNK^{+}05, EFG^{+}05, FOW^{+}09, FML^{+}06, GGB^{+}08, Gra07, HP05, HT08, HCN06, HCB^{+}08, JLC08, JSC08, JKW^{+}09, JHY^{+}08, KKH^{+}06, KHBV05, KKM09b, KSR^{+}08, LCO07, LC07, LSR07, LTK09, LMB^{+}06, LR07, MSO09, MFO09, MAZ^{+}08, MTB^{+}09, MOCC05, MPG^{+}09, MDH^{+}06, MKSJ06, NSS^{+}08, OLS08, PMW05, PCMSB09, PFD06, PWW^{+}09, RCM^{+}09, RGW^{+}07, SRX^{+}07, SS07, SRL05, SMSPK08, SHS^{+}09, SNC^{+}06, SB06, SG09, SORM05, SWG^{+}05, SRQL09, TRS^{+}08, VLK^{+}08, VSG09, WJM05, WJM07, WLC08, WLE09, WFJ06]. DRP1 [WZM07, HLL^{+}08, KNY07, WZW^{+}09]. dry [Wel05e]. dsg2 [PRB^{+}09]. DSL [LNG^{+}05, NMO^{+}07]. dTACC [TLDB08]. DTNBP1 [CFH^{+}08]. Dual [MMS07, PWW^{+}09, SYM^{+}08a, CLA^{+}05, LeB07k, LLW^{+}09, MMNH09]. dual-specificity [LLW^{+}09]. DUBway [MW06]. ducts [SBC^{+}05]. due [LWG^{+}05]. Dumbfounded [MOCC05]. dunk [Sed08m]. duplication [HRB^{+}06, LSC^{+}06b, WCR^{+}08]. During [CBG^{+}09, BJKF05, BFG^{+}06b, BBvD05, BTP^{+}07, BAR^{+}08, BEN^{+}05, CPAH08, CTGVGFF08, CCC^{+}06, CGGT09, CEY^{+}05, DHT^{+}09, Den05, DSK^{+}06, EW08, ESZK06, FOB^{+}09, FvD08, FSnP^{+}06, GKCM^{+}08, GAS06, GT08, GER^{+}07, HJ06, HP05, HTYNK09, HEW07, HO08, HHY^{+}08, HKLM06, HKLM07, HLS^{+}09, HCB^{+}08, HCFM08, IUT^{+}07, JKDH08, JCMG08, JGY09, JSC08, KMS07, KR06, KES^{+}09, KSL^{+}05, KNT^{+}06, LN07, LC07, LSNY08, LSC^{+}07, LHL^{+}08, LFS^{+}05, MCSY07, MHL^{+}06, ML05, MAG05, MuMI^{+}06, MM08, MSSF05, MAZ^{+}08, MSW^{+}07, MSW^{+}17, MAG^{+}08, MHC^{+}09, NYM^{+}08, NSM05, NNK^{+}08, NRD^{+}06, NAG^{+}08, NZQ^{+}07, OLH^{+}08, OHH^{+}06, OD07, PB05, PG06, PYR08, PSS^{+}05, PWW^{+}09, PEW^{+}05, PKB^{+}06, PVJJ07, Rex09, SFP^{+}05, SMSPK08, SMM^{+}08a, SHS^{+}09, SDB^{+}05, SS06b, SOW^{+}09, SXH^{+}06, SPK^{+}09, ST09, SBMP05, SL05, SLR^{+}06, SWG^{+}05, SCK^{+}07, TPNS07a, TPNS07b]. during [TTK^{+}06, TP09, TST08, TWG^{+}09, TNS07, VSG09, VMY^{+}09, WC07a, WAKS08, WZM07, WLLF^{+}08, WBT^{+}08, WPL^{+}06, WFJ06, WXM^{+}05, YGY^{+}05, YLM06a, YDL^{+}06, YEL^{+}08, YHD^{+}09, YKP^{+}08, YK07, vDVM^{+}09]. duties [Les08n]. duty [Les07-56]. Dvl [qGyWX^{+}08]. Dying [Bas07a, Les06h]. Dynactin [Wil07r, KLR^{+}07, LSC^{+}06a, LGV08, UHM^{+}08]. dynactin-mediated [LGV08]. Dynamic [BJLM^{+}05, BVGA^{+}05, CHA05, CBMCSS08, DvCvH^{+}08, HH06b, KG05, PYR08, SFEK07, TT09, VBNR06, Wel06h, CSTT^{+}06, DMDO08, DGSN06, DCMB^{+}07, FvD08, GMG^{+}07, HH06a, HT08, MCA^{+}06, OF08, TM05, Wel06-35]. dynamically [AWM^{+}08, BKLT^{+}08, KSA^{+}05]. Dynamics [HWPH^{+}08, LZM^{+}06, VSG09, Wil09g, Ban09, BVS^{+}09, CZM^{+}07, CCH09, CSL^{+}05, CCP^{+}08, EMD^{+}08, GTHK05, GHA^{+}05, GJL^{+}08, GFD^{+}05, HTYNK09, HDM^{+}07, JWC^{+}08, JSC08, KBW^{+}06, LOWTL^{+}08, LSNY08, LDPLS05, MuMI^{+}06, MKC^{+}05, MKGL^{+}05, NWB^{+}06, PFGSV06, RMP06, SHC07, SRN^{+}07, SBW^{+}07, SDD^{+}07, SSK^{+}09, TPNS07a, TPNS07b, TÖL^{+}06b, UWiY^{+}07, VMZW^{+}07a, VMZW^{+}07b, Wil07q, WF06]. dynamin [GKvdB07, KBC^{+}08, NLS^{+}05, SDDV05, TT09, WZW^{+}09]. dynamin-independent [KBC^{+}08]. dynamin-like [GKvdB07]. dynamin-related [WZW^{+}09]. dynamitin [LGV08]. Dynein [FDB^{+}09, LWHC08, Sed08f, Tum05r, AGL^{+}08, BSM^{+}08, BSM^{+}09, CFU^{+}09, CFL^{+}05, CLG^{+}07, DS07, GWS09, GSV07, HLM^{+}08, HFM^{+}05, HRK^{+}09, HQF^{+}07, JRZ^{+}07, KHB^{+}05, LKC05, OHK07, PKK09, PFG^{+}05, SON06, SCMV07, UHM^{+}08, VDSV06, VMSV08, WBT^{+}08, Wil07r, deC09, LeB07l, ZRH^{+}09]. dynein-driven [GWS09]. Dynein-mediated [LWHC08]. dysfunction [EVF^{+}08, hKDH^{+}08, MHM^{+}05, SFB09]. dysfunctional [HS09b]. dysplasia [GS07]. dystroglycan [MBL^{+}07, RKN^{+}07]. dystrophic [GSB^{+}06a, PMC08, RKN^{+}07]. Dystrophin [Les09n, PHM^{+}09, RMR^{+}09]. dystrophy [MBL^{+}07, MDL^{+}09, PRB^{+}09, SBJ^{+}07, SVS^{+}07a, SVS^{+}07b].

E-cadherin [LNPC06, KSR^{+}08, LCS^{+}09, LBC^{+}07, NN07a, QCGM05, SYM^{+}08b]. E-cadherin-mediated [SMSPK08]. E-catenin [LGV08]. E2 [SLK^{+}08]. E2-25K [SLK^{+}08]. E2-25K/ [SLK^{+}08]. E2F [DVF^{+}06, MYO^{+}05]. E2F/DP [MYO^{+}05]. E3 [GBNS07, KNY07, LJY^{+}07, MOS^{+}09, ORM^{+}08, STvdB^{+}09, TTE^{+}07, WHC^{+}07]. E4orf6 [HAT^{+}05b]. each [LeB06-80]. Early [BJM^{+}06, GMG^{+}07, HZOL05, LeB07m, LeB08f, RCT09, Tum06m, Wil07s, BBL^{+}08, CZP^{+}05, CCS^{+}08, HXZ^{+}07, Heu05, HAAHD05, HLJ^{+}07, JLG^{+}06, KKC^{+}07, KHW^{+}07, KHW^{+}09, KFC^{+}05, KZS07, KPS^{+}07, LeB05-30, LeB05-74, MFO09, MM08, NAG^{+}08, PSL^{+}06, QU08, SNC^{+}06, SYA^{+}08, SBJ^{+}07, SRD^{+}07, UPBM06, Wil07f, dPMF^{+}09]. easy [Les07z, Wel05m]. Eat [Wil07t, LeB08j, Les09c, RLDC09]. EAT-3 [RLDC09]. EAT-3/Opa1-dependent [RLDC09]. Eating [Tum05s, Bas07b, Les08k, SS08b, Wil07-58]. eats [Wel05j]. EB1 [BKLT^{+}08, MKGL^{+}05, SRE^{+}05, Vau05, WTH05, ZSMW09, ZAM07a, ZAM07b]. EB1-mediated [SRE^{+}05]. EB1/tubulin [BKLT^{+}08]. EB1/tubulin-binding [BKLT^{+}08]. EBP [BTD^{+}07, LLL^{+}06]. EBs [Wit08]. ecdysone [SSAD08]. ECM [Les06i, Pow06d, Tum05t]. ECT2 [YPG05, OYKY^{+}05]. ectodomain [AXM^{+}05, XMA^{+}09]. Edc3p [DTP07]. Edg [SWH^{+}06]. Edg-3 [SWH^{+}06]. edge [BWW^{+}08, GC05, KBR^{+}08, LMM^{+}09, LeB07-51, LeB08-36]. editing [VBL^{+}05]. education [Wil08h, Wil08h]. Eed [AMERT^{+}07]. EF1 [BLT09b]. effect [BLB^{+}05, DLL^{+}08]. Effector [HCB^{+}08, ASP06, AWR^{+}07, CAR^{+}08, GHSG07, GPJ^{+}06, GAG^{+}06, LLL^{+}06, PSL^{+}06, PST^{+}06, SHL^{+}09]. effectors [SHC^{+}05]. Effects [MXR^{+}05, HBPL09, JGAS08, Kat07d, OF08, OMLP^{+}08, PPS^{+}06, SYM^{+}08b, XSVGR05]. efficacy [PFD06]. efficiency [DLT^{+}07, YM06]. Efficient [BPL^{+}08, JSPT05, GNL^{+}06, HH06b, NdCF^{+}05, TFC^{+}06b, WF06]. efflux [FKN^{+}07, RBC05]. effort [Wel06r]. Efr3 [BSA^{+}08]. Eg5 [UHM^{+}08]. EGF [HK06, HIK^{+}08, HPH^{+}05, Les05a, SLWA07, WBC^{+}06, vRSvR^{+}07]. EGF-induced [HPH^{+}05, vRSvR^{+}07]. EGF-stimulated [SLWA07]. EGFR [BPL^{+}05, CCL^{+}07, GSiK^{+}09, LPG^{+}07, Rob08-46, Sho09d, SWH^{+}06]. EGFR1 [CCL^{+}08]. Egg [Rob08j, BBM^{+}07, LM05, MFH05, OW06, Pow05l, UHM^{+}08]. EGO [LeB05-64]. EGP [Wil07-40]. eIF [LeB08g]. eIF2 [CHA05, LeB05-31]. eIF2B [CHA05]. eIF2B-containing [CHA05]. eIF4E [CTS^{+}05, CTS^{+}06c, CTO^{+}08, FBD^{+}05b, LeB05-87, LeB06u, LDGO^{+}09]. eIF4E-binding [FBD^{+}05b, LDGO^{+}09]. eIF4G [KII^{+}08]. eIF4GI [RVBZ^{+}08]. eIF5A [PCMSB09]. eisosomes [FMA^{+}09]. Elaine [Sho09u]. elasticity [LeB06-58]. elastogenic [HOH^{+}07]. Electric [Wel06i]. electrical [BM07, MAG05]. Electron [PGL^{+}09, HKV^{+}05, MFM^{+}06]. Electron-tomographic [PGL^{+}09]. electrostatic [FOB^{+}09]. elegans [ASSST06, AHM^{+}05, BAR^{+}08, CLG^{+}07, DPD^{+}09, ECR^{+}09, ESG^{+}06c, GCS^{+}07, GSW^{+}08, HK06, HKLM06, HKLM07, IEGV06, JNHB08, KR06, KHW^{+}07, KHW^{+}09, KO07, LLH^{+}07, LDGO^{+}09, LBG^{+}06, MYMZ08, MNF^{+}07, MQL^{+}06, MS07, NAG^{+}08, OOS07, OOS09, OBY^{+}06, OV09, POCS^{+}06, RLDC09, SHL^{+}09, TA07, WZC^{+}08, YWEK08, ZRH^{+}09]. element [CTS^{+}05]. elementary [LWBL06]. Elevated [WM05]. elicit [TVP^{+}06]. elicited [CMF08]. elicits [TRC^{+}08]. elimination [KR06, SRL05]. elixir [Wel06z]. Elizabeth [Sed07b, Wil08j]. elongate [ÇB09]. elongation [BGJ^{+}07, BLS^{+}05, CGGT09, MHFH06, Wel06-48]. elucidation [SH05]. embryo [BRH05, BBL^{+}08, FML^{+}06, GEN^{+}06, Les06k, MFO09, YBT^{+}09]. embryogenesis [SPA^{+}05, SXH^{+}06]. embryology [Wil07-57]. Embryonic [Les07s, SRXvdK06, WMSB^{+}06, AHM^{+}05, IMY^{+}09, KAD^{+}08, LW07b, Les06a, NTB^{+}06, NPB^{+}05, OBY^{+}06, TPZG06, VMY^{+}09, WZW^{+}09, WRKC^{+}09, Wil09s, ZNP^{+}09]. Embryos [Rob08k, Rob08l, FTB^{+}09, HKLM06, HKLM07, KHW^{+}07, KHW^{+}09, LDGO^{+}09, MSO09, PMW05, SWG^{+}05, TA07, Wil07-58]. Emerin [LeB07n, LHS^{+}05, SSHP^{+}07]. emerin-deficient [LHS^{+}05]. Emery [MDL^{+}09]. Emi1 [DP07, Les07t, MVC^{+}07]. Emi1-dependent [MVC^{+}07]. Emi2 [MHL^{+}06]. employs [TVP^{+}06]. EMT [NBM^{+}08]. Ena [FSK^{+}07, TSP08]. Ena/ [FSK^{+}07]. Ena/VASP [LLHH07, TSP08]. enabled [TGH^{+}07]. enables [BJKF05, LLM^{+}05]. encoded [PPS^{+}06, YOSM06]. encodes [WBG^{+}08]. encoding [Wel05-66]. encounter [Les06c]. encounters [HZOL05]. end [BVS^{+}09, ÇB09, DG05, FBD^{+}08, GHM^{+}08, GNV05, KDG^{+}09, KCMB08, LeB06-42, MKGL^{+}05, MSB06, SRE^{+}05, SMG07, WC07b, WTH05]. end-directed [GNV05]. end-to-end [ÇB09]. Endo180 [CEY^{+}05, SWI06]. endocardial [NBM^{+}08]. endochondral [PXD^{+}07]. endocycle [SSAD08]. endocycle/gene [SSAD08]. endocytic [AMS^{+}05, BLC^{+}09, DNK^{+}05, GOM^{+}09, IGFG^{+}08, JWG^{+}07, Kak08, KRP^{+}09, KFC^{+}05, KBC^{+}08, LPB^{+}07, Pow05w, SHC^{+}05, SORM05, SSSB07, SCK^{+}07]. Endocytosis [LeB07o, Nic09a, ALvdG06, BMBD^{+}09, DPK^{+}05, EBMG09, HT08, HCN06, ITG^{+}08, JHY^{+}08, KHH08, KBC^{+}07b, LeB06g, MS08, NMO^{+}07, OD07, OWM05, SDDV05, Sed09h, TSS^{+}06, Wel06-66, Wil07-47]. endoderm [YBT^{+}09]. endogenous [BYF^{+}05, DTR^{+}05, PTL^{+}05, PVM^{+}05, YLM^{+}06b]. endomembrane [TM06]. Endoplasmic [LHRP05, NSM05, PVJJ07, ASB^{+}05b, AH08, ADO07a, ADO07b, AWM^{+}08, CHO^{+}09b, JASC08, KPM05, LC07, LMC^{+}09, LFS^{+}05, NTHN^{+}08, PMW05, RCFB08, SPT^{+}09, SPT^{+}21, SBV^{+}06, WWS^{+}06, ZKN^{+}06]. Endosomal [VLK^{+}08, CSSE06, DHT^{+}09, JRZ^{+}07, MYMH09, PCR^{+}07, PAV^{+}06, RCT09, TTK^{+}06, TDG^{+}06, XGX^{+}09, YWK^{+}08]. endosomally [RRP^{+}08]. Endosome [Tum06n, ASH06, BFF^{+}06, CZP^{+}05, CHCB05, FN06, MKH^{+}07, PSL^{+}06, Rob08-43, RKvdK^{+}09, SCM^{+}07, SXH^{+}06]. endosome-like [BFF^{+}06]. Endosomes [Les06j, SWI06, GEP08, GMG^{+}07, HLM^{+}08, HSB^{+}07, LeB06-54, LeB07-42, LTAM^{+}09, MZ08, NWO06b, NFCF08, RCT09, RvVM^{+}08, Wel07h]. Endothelial [BZYM^{+}08, Bon05c, BHK^{+}06, BSK07, BBM^{+}08, BDR^{+}06, CNM^{+}05, DDLC^{+}06, FHFB^{+}06, FSK^{+}07, FWC^{+}07, GHSG07, KWS^{+}08, LOG^{+}06, Les05h, LZ08, LPF^{+}06, LRCH^{+}05, LLC^{+}08, MHM^{+}05, MMM^{+}05, PdBR^{+}07, SBF^{+}06, SSS^{+}08a, SPA^{+}05, SDN^{+}05b, WLLF^{+}08, WXM^{+}05, ZXM^{+}06, vBAS^{+}07]. Endothelin [PCR^{+}07]. Endothelin-converting [PCR^{+}07]. endothelium [BYK08]. ends [BKLT^{+}08, BGP^{+}06, DvCvH^{+}08, LeB06-35, LeB06-44, LeB08y, PTF^{+}06, Tum05-38, Tum05-66, Vau05, Wit08]. energetic [MSK^{+}07]. energy [Pow05x, Wel06-58]. engage [PVM^{+}05]. engagement [DSLY^{+}05, LNPC06, MRKH05, vBAS^{+}07]. engaging [UDC^{+}07]. engulfment [TD08]. enhance [BSL^{+}08, CFBM05, MPD09, SSLS05, VVL^{+}08]. enhanced [BMK^{+}06, GSB^{+}06a, LYK^{+}06, MFS^{+}09, PS08]. Enhancement [WNL^{+}09]. enhanceosome [STL^{+}07]. enhancers [KHBV05]. enhances [BRKR06, BFP^{+}07, DXC^{+}08, HFvdV^{+}05, JGL^{+}06, LPWP^{+}07, LDGK08, LCL^{+}05, MRKH05, SDL^{+}05, SSP^{+}07, WGKA^{+}08]. enhancing [Ros05]. enigma [Lin08]. enlargement [MOCC05]. enough [LeB05-73, Les07w]. enriched [AWM^{+}08, DMC^{+}07, LZ08, NKK^{+}06]. enrichment [DPD^{+}09]. ensure [CYO^{+}09, TLDB08]. ensures [Les07t, SYM^{+}08a, SiTT^{+}08, TMH^{+}07]. enter [LeB07-49, LRBM09, MHL^{+}06, Sho09-37]. enterica [SHB^{+}07]. enterocyte [DGZ^{+}05, MHS^{+}09]. enterocyte-like [DGZ^{+}05]. ential [Sho09e]. Entropic [CM09]. Entropy [LeB06v, Wel06-50]. Entry [NPMZ06, CTS^{+}06b, DTL^{+}07, GLS^{+}09, LeB05h, LSM^{+}05, LRBM09, LWBL06, MFO09, Put05, SBJ^{+}07, TMZ^{+}09, VLK^{+}08, WZO^{+}05, Wel06-45]. envelope [AH08, AVHH09a, AVHH09b, CYO^{+}09, EVF^{+}08, GNPC^{+}08, HMG^{+}08, Heu05, JMG^{+}06, LeB06r, LeB07i, LeB08-34, Les09-29, LFH07, LS06, MK07, RW08b, TTYH09, Tum05-68, UPBM06, ZRH^{+}09]. environment [BGdN^{+}08]. envoplakin [SNG^{+}07]. enzymatic [ASK^{+}09, SHC^{+}05]. Enzyme [LeB08h, Rob07d, ASH06, AZP^{+}05, NFI^{+}09, PCR^{+}07, WLY06]. enzyme-1 [PCR^{+}07]. enzymes [Les08-48, MW06, WRP^{+}06, ZRBS07]. EPB41L5 [HHY^{+}08]. EphA4 [BMRP07, IDTT^{+}09, LCL^{+}05]. EphB2 [PCPW08]. ephrin [QWD^{+}08]. EPI64 [HGR^{+}06]. epiblast [GEN^{+}06, GNE^{+}07]. epiboly [LCS^{+}09]. Epidermal [VMY^{+}09, BFG^{+}06b, DSBK09, EFM^{+}07, FHS^{+}07a, GSiK^{+}09, LCR^{+}05, LCT^{+}09, LGT^{+}07, NTB^{+}06, RRP^{+}08, SNG^{+}07, SCFKB07, dGSWW^{+}06, ZTKK05]. epidermal-specific [dGSWW^{+}06]. epidermis [LF07, Sho09-38, WBC^{+}06, WPWF08]. Epigenetic [JDK^{+}05]. Epigenetics [Hor09]. epilepsy [Les08-31]. epileptogenesis [WKW^{+}08]. epithelia [AS06, WL09b]. Epithelial [SHH^{+}08, AMS^{+}05, AFV^{+}05, BAR^{+}05, BLP^{+}05, CTP^{+}09, FHMC09, FSP^{+}07, HP05, HHY^{+}08, HSTM06, IFF^{+}05, IUT^{+}07, IPG^{+}05, JKDH08, KKW^{+}09, LD07, LDKT06, LJM^{+}07, LWHC08, MBR^{+}06, MCSY07, MSK^{+}07, MMS07, MRKH05, NN07a, NPV^{+}07, NFCF08, OIAT06, PH07, QCGM05, RYN05, RKH^{+}09, STM^{+}07, SLM^{+}06, SSM05, SIS^{+}08, TPNS07a, TPNS07b, TVP^{+}06, VS07, WWX^{+}06, WL09a, WWS05b, WWC^{+}07, XSVGR05, YN07, dRKD^{+}05]. epithelial-restricted [MBR^{+}06]. epithelium [BDP^{+}07, JLG^{+}06, NLC^{+}05, SMR09, WC07a]. epitope [WF07]. epoxyeicosatrienoic [FLA^{+}08]. Eps15 [KKW^{+}07, RRP^{+}08]. epsin [LZ09]. epsin1 [DBW^{+}06]. Epstein [FFS06]. Epstein-Barr [FFS06]. equal [CGJ09, HGJ06, LeB05-45, Tum06d]. equality [Les09p]. equalization [MPLF^{+}09]. equation [vRG08]. equilibrium [DGSN06, MCA^{+}06]. ER-associated [OOY^{+}06, SS08a, SS08b, SN05, Tum05-63, WHR^{+}09]. ER-stress [KIKI^{+}07]. era [EGL^{+}09]. ERAD [KGL07, LDL^{+}09a, Tum06-50, WHC^{+}07]. Eran [Sed08g]. erasin [LDL^{+}09a]. ErbB [BHJ^{+}05, HCRT08]. ErbB-2 [BHJ^{+}05]. ErbB-2/ErbB-3 [BHJ^{+}05]. ErbB-3 [BHJ^{+}05]. ErbB2 [YMCT08]. ERGIC [NRW^{+}08]. ERGIC-53 [NRW^{+}08]. Ergosterol [JMG08, BVGA^{+}05]. ergosterol- [BVGA^{+}05].

ERK [BWW^{+}08, HSB^{+}07, JGL^{+}06, MWMS07, YTY^{+}07]. ERK1 [GNPC^{+}08, MDET08, THN^{+}06, DSBK09]. ERK1/2 [GNPC^{+}08, MDET08, DSBK09]. ERK1c [SS06b]. ERK5 [CWC^{+}07, SYKM08, Wil07-28]. ERK5-NF [CWC^{+}07]. ERKs [Wel05-66]. ERM [HLK^{+}09, HF06]. ERO1 [LMC^{+}09]. ERO1- [LMC^{+}09]. Ero1L [TRS^{+}08]. Error [CFV07, Wel05x]. errors [MGW^{+}08]. ERs [Bas07b]. erythrocyte [MBR^{+}05, OHH^{+}06]. erythropoiesis [EVH^{+}07, EVH^{+}08]. erythropoietin [EVH^{+}07, EVH^{+}08]. escape [LeB05-64, Les08-27, Wel06-52]. Escaping [LeB06w]. Escherichia [BM07]. ESCRT [CSSE06, HRLH08, HR09, LeB06-59, LeB08i, NWO06b, SOW^{+}09]. ESCRTs [ADD^{+}06]. EspF [AWR^{+}07]. Essential [BTP^{+}07, BLP^{+}05, BCC^{+}09, GKCM^{+}08, KNT^{+}07, LLJ^{+}05, LSC^{+}05, MYZ^{+}09, PdBR^{+}07, Wel06j, vKGK07, ASH^{+}09, ANI^{+}07, ASHD^{+}05, BAR^{+}05, BLS^{+}05, BSDA07, BHK^{+}06, CBM^{+}07, CTP^{+}09, CBG^{+}06, CBIO05, DBR09, FVSC08, GSH^{+}08, GSV07, GSW^{+}08, HGJ06, HSY^{+}06, HDS^{+}07, HH06b, KHWC08, KHH08, KVS^{+}09, LRCH^{+}05, MSL^{+}09, MFH05, MEG^{+}07, NYM^{+}08, NNK^{+}08, NHK^{+}08, PXD^{+}07, PPBF08, RDY^{+}05, RAP08, SSJ^{+}07, SWT^{+}05, SLM^{+}06, SXH^{+}06, SOC^{+}07, SHDT05, STL^{+}07, TLM^{+}07, TJC06, TFC^{+}06b, TMAL^{+}06, WOWB08, XNM^{+}09, YA06, YEL^{+}08]. establishing [HSK^{+}05]. establishment [HP05, PNO^{+}09, SHC09]. estradiol [LCM^{+}08]. Estrogen [SWH^{+}06, PKJS^{+}08, SKR^{+}07]. etch [Heu05]. Etd1 [GCM09]. ethylmaleimide [MMM^{+}05]. eukaryotes [HG07]. Eukaryotic [HS09a, AWR^{+}07, BW08, MGBH^{+}06, PALW05, XMSYJ07]. even [Hat05a]. event [SMRS^{+}07]. events [ESZK06, ST09]. everything [Sho09-40]. Evidence [Tum05u, VFRJ^{+}07, MS06, MTH^{+}06, ZHS06]. evoked [pTyLZ^{+}09]. Evolution [DFA05, YBK^{+}06, ZB09, LeB05-35, Tum05-61]. evolutionary [GDW^{+}07, Wil07-68]. Evolving [Wel07e]. exacerbates [SVS^{+}07a, SVS^{+}07b, Wel06s]. Examining [TC08]. Excess [EFG^{+}05, Pow05j, WGL^{+}06, LeB05-41, LeB06d, Les07-53]. exchange [ÇB09, GSdBS09, HM08, KTA^{+}06, LeB06-30, LLT^{+}08, LLW^{+}09, MBS^{+}08, Wil07u, YMCT08, tKJCH06]. exchanger [FKN^{+}07, PB05, SSD^{+}09]. exchangers [pTyLZ^{+}09]. excision [MBJM^{+}09]. excitability [JBL^{+}07]. excitation [WBG^{+}08]. excitement [Les08-32]. excitotoxic [CYY^{+}08, TGRD09]. excitotoxicity [KZT^{+}08, KZT^{+}09, SFY^{+}05]. excluding [Nic09a]. Exclusion [Wel05t]. exclusively [BKHR05]. exhibit [BLL^{+}06, CAL^{+}07]. existence [Sho09-31]. exists [FML^{+}06, MCA^{+}06]. Exit [Tum06o, ZKN^{+}06, BvOM^{+}07, CN06, DNG^{+}06, FDV^{+}06, HHZ^{+}08, JBFC07, LDS^{+}06, MHC^{+}09, NYM^{+}08, Poo09, Rob08-27, SGDM06, Wel05-44]. Exo70p [HXZ^{+}07]. Exocyst [LeB05-36, LH05, RWM^{+}05, SYRL^{+}08, SORM05, ZWG^{+}05]. exocytic [HXZ^{+}07]. exocytose [GT08]. Exocytosis [Tum05v, BSMT05, BLP^{+}05, Gle07, HJ06, LOWTL^{+}08, LSNY08, MMM^{+}05, OIFN^{+}07, OVL^{+}07, RBC05, RWM^{+}05, SRA^{+}05, TDG^{+}06, Tum05w, ZWG^{+}05]. Exomer [WHOS06]. exon [ELD^{+}08]. exonic [KHBV05]. exosome [Wil07u]. Exosomes [BFF^{+}06, Tum06p, BW08, LeB07-30, LBG^{+}06]. expand [Sho09k]. expanded [QIH^{+}06]. expanding [PYR08, Ska07]. expands [GSY^{+}09]. expansion [PST^{+}06, SPT^{+}09, SPT^{+}21, YN07]. experience [Rob08k]. experimental [CFV07]. exploits [KGDR^{+}07]. exploring [Tak06]. export [CTS^{+}05, FKJ^{+}09, FTM^{+}06, HFvdV^{+}05, HOZ^{+}09, KFK^{+}05, LJS^{+}09a, LCM^{+}08, MBR^{+}05, SRBL06, SCVW09, SWT^{+}05, Sho09g, TW07, Tum05-50]. Exportin [GGE^{+}09]. exports [LeB06u]. exposed [Les09-33]. exposes [KBC^{+}07b]. express [GNE^{+}07]. expressed [GCSG^{+}05, HMM^{+}05, NBB^{+}09]. expressing [RP05, ZDD^{+}06]. Expression [JGAS08, SFY^{+}05, BXN^{+}06, CBG^{+}06, DGP^{+}05, GFC^{+}08, JLG^{+}06, JDK^{+}05, KBW^{+}06, KHBV05, KCD^{+}06a, KZH^{+}07, LBVN09, MLS^{+}07, MDW05, OMLP^{+}08, PCV^{+}08, RMC^{+}05, RGA^{+}06, RLH^{+}09, SKM^{+}08, SKR^{+}07, SYM^{+}08b, TRC^{+}08, WM05]. extends [ESG^{+}06c, KHB^{+}05, Les08-28, MGH^{+}09]. extension [BJKF05, KHH08, LSC^{+}05, Mat05, PRK^{+}07, PRK^{+}09, RFBB07, YKP^{+}08]. Extensive [MBdR^{+}08]. External [SWL^{+}05]. Extra [Tum05w]. Extracellular [WvNS^{+}07, BRM^{+}07, CEY^{+}05, DMC^{+}07, FAG^{+}07, GXJF07, GRC^{+}05, HXO^{+}07, HBD^{+}05, MFBG^{+}08, OSB^{+}05, WLS^{+}09, WRMH07, WOP^{+}06]. extract [UHM^{+}08]. extracted [CMF08]. extracts [BBM^{+}07, LM05, MFH05, OW06, Pow05l]. extrinsic [SRXvdK06]. extrude [SMR09]. eye [Rob07a, Sed09e, Tum06q]. eyelids [STK^{+}05]. ezrin [IKZ^{+}07, TKH^{+}05].

F [EHWL06, GMG^{+}07, KOM07, RMC^{+}05, SS09, ZMA^{+}07]. F-actin [SJL^{+}07, GMG^{+}07, KOM07, SS09]. F-box [EHWL06, RMC^{+}05]. F-spondin [ZMA^{+}07]. faces [RH09]. facilitate [HEW07, SS09]. facilitates [BLT^{+}09a, GSW^{+}08, KFK^{+}05, WWvD07, YWK^{+}08]. facilitating [GSY^{+}09]. Facilitation [XLH^{+}07, HSK^{+}05, LLL^{+}06]. factor [AOP^{+}05, ANI^{+}07, BSK07, BCR^{+}08, BBG^{+}05, CSM^{+}06, DLBSD^{+}05, DUR^{+}05, ELD^{+}08, FBY^{+}06, FHS^{+}07a, HM08, HZL^{+}07, HIK^{+}08, KTA^{+}06, LeB08-35, LDGK08, LPF^{+}06, LLT^{+}08, MGBH^{+}06, MYMH09, MNF^{+}07, MMM^{+}05, MBS^{+}08, MOM^{+}06, NJS^{+}06, POGL08, PUG^{+}06, PTF^{+}06, RRP^{+}08, SSH^{+}05, SCVW09, TVP^{+}06, VVL^{+}08, Wil06f, Wil07-67, YMCT08, YGR^{+}08, tKJCH06, BAR^{+}05]. factor- [TVP^{+}06]. factories [XC08]. factors [BLC^{+}09, BZA^{+}06, BDR^{+}06, CDW^{+}05, DCMB^{+}07, GGE^{+}09, KBN^{+}06, KKK^{+}05, LeB07s, LSD^{+}06, MMS07, OYTO07, WPWF08, Wil07-30]. fails [BvRJ^{+}05]. failure [PDM^{+}08, Rob08v, SKvdBW^{+}06]. failures [CGK07]. FAK [BHJ^{+}05, BHK^{+}06, CCH09, GKN^{+}07, Les08e, LLT^{+}08, PLR^{+}06, Tum06s, WLLF^{+}08]. falciparum [BAdSG05]. fall [LeB05-50, LeB07-66]. false [Ros08a, Ros08b]. FAM29A [ZCJ^{+}08]. familial [ZS08]. family [BLH^{+}09, CTY^{+}06, DBD^{+}09, DDB^{+}05, GGO^{+}09, GRGP07, GGE^{+}09, GAG^{+}06, GFD^{+}05, HHFC06, JKW^{+}09, KCE^{+}06, Las09, MHM^{+}07, PSS^{+}05, RGCWG09, SWD^{+}08, SGKM08, VWM^{+}07, LeB06-52]. Fanconi [MLH^{+}06].

FAPP2 [VVMS05]. Fas [AAG^{+}09, HBD^{+}05, PMK^{+}05, WWS^{+}06]. Fas-mediated [WWS^{+}06]. Fasciclin2 [KTN07]. fascin [ViKA^{+}06]. FasL [HBD^{+}05]. Fast [GNPC^{+}08, DvCvH^{+}08, MTH^{+}06, Rob08-37]. faster [YKBR09]. fat [HCN06, Wel05u, Wel05-39, Wil09h, IMY^{+}09]. Fat1 [HLA^{+}06]. fatal [Les08-32]. Fate [Wel05v, AWE^{+}05, HAT^{+}05b, KCD^{+}06a, LS09, LeB07g, Les06i, NDS^{+}07, OYTO07, Rob07d, SRXvdK06, SAD^{+}09, VWM^{+}07]. fatty [SKM^{+}08]. Fbh1 [FMD^{+}09]. Fbw7 [BDS^{+}08, GGH^{+}08]. Fc [MYZ^{+}09, NYI^{+}05b]. Fcj1 [RSS^{+}09]. FcRn [TMY^{+}09]. features [CRV^{+}06, WPDZ^{+}06]. feed [LeB05-102]. feedback [ANI^{+}07, FKN^{+}07, KBW^{+}06, LRBM09, MOCC05, PCPW08, XKYW07, YWW^{+}08, YMDL09]. feel [Pow09a]. Feeling [Les08s]. Fellow [Wil07c]. Female [LeB07p, DL08, DLT^{+}07, MFC^{+}09]. females [KPS^{+}07]. fences [LeB06-68]. feng [Les08t]. Fer [LPN^{+}08]. fertility [LeB05i]. fetal [FKS^{+}06, NRJ^{+}09]. Fever [Wel06k]. FG [SRX^{+}07, TW07]. FGF [KKH^{+}06, LeB05-37, MAH^{+}05, MRD^{+}08]. FGF- [MRD^{+}08]. FGFR1 [FCB^{+}09, FCB^{+}19, XSVGR05]. FGFR2 [MAH^{+}05]. FHA [ShCM^{+}06]. FHL1 [CMN^{+}08, Sho08]. Fhl2 [WHM^{+}07]. FIAT [YAV^{+}05]. Fiber [Nyb05d, AWE^{+}05, NFZ^{+}09, TMI07, Wel06-36]. fibers [GPC^{+}06, HL06, KRZ^{+}08, RKN^{+}07, Tum06-48]. fibril [TLM^{+}07]. fibrillar [DWMY09]. Fibrillin [CCM^{+}07]. Fibrillin-1 [CCM^{+}07]. fibrinogen [SZX08]. fibroblast [LMMHP^{+}05, MYMH09]. Fibroblasts [Sho09v, LeB05-90, LLM^{+}05, NMV^{+}07, PHB^{+}07, RG06a, SH05, TCHL05, THN^{+}06]. Fibronectin [Sho09w, AXM^{+}05, FZT^{+}07, GS07, TLM^{+}07]. fibrosis [SBM^{+}07]. Fibulin [HK06, HOH^{+}07, Wil07v]. Fibulin-1C [HK06]. Fibulin-5 [HOH^{+}07, Wil07v]. Fibulin-5/ [HOH^{+}07]. field [Les06p]. Fifty [Mel05]. Fig4p [DTW06]. Fight [LeB06x, LeB06-76, LeB08l]. Fighting [Les07u]. filament [BLL^{+}06, CSTT^{+}06, CBG^{+}09, MTH^{+}06, RXA^{+}08, SOR^{+}05, SSK^{+}09, WKWL06, WAW08]. filamentous [LSNY08, Ped08]. filaments [BSW^{+}05, CM07, ÇB09, HRLH08, KG05, LBI^{+}08, LC09, LSC^{+}06b, MSV^{+}05, MQL^{+}06, Wel06-41, WOWB08, YHJ^{+}05]. Filamin [ZBKJ07, NOH^{+}07]. files [Hil08]. Filling [Pow06e]. Filopodia [Les07v, NYN^{+}06, HHT^{+}07, LeB07-67, LSM^{+}05, LLHH07, NAS08, TMI07, Wil07d]. filopodial [PRK^{+}07, PRK^{+}09, ViKA^{+}06]. filter [LeB08d]. filtering [LWG^{+}05]. final [BXHD08]. finally [Mal06]. find [CF07, Tum05z]. finding [Smi07]. finds [LeB05-36, LeB08-32]. fine [Pow09d, Rob08n, WSG06]. fine-tunes [Rob08n]. fine-tuning [WSG06]. finger [HOZ^{+}09]. fingers [LeB07-58]. FIP200 [GMW^{+}05, GPN^{+}06, HTK^{+}08]. fire [Tum06-55]. first [HB08, Les07-45, OIFN^{+}07, Pow05k, Rob08-49]. first- [OIFN^{+}07]. Fis1p [KCAS05]. fish [AS09, Wil09s]. fishing [Les08l, Wil09m]. Fission [Tum05x, BbCQ^{+}09, BHG^{+}05, BBPM07, CGGT09, DSK^{+}06, GGC05, HYB08, KOM07, KCAS05, NNK^{+}08, NN07b, SON06, SBMP05, TJC06, Wel05-37, WSK^{+}06]. fit [IPM^{+}05]. Fixed [Wil08p]. Fixing [Wel06l]. flagella [BSM^{+}08, BSM^{+}09, Den05, GWS09, HRK^{+}09, HQF^{+}07, HDM^{+}07, HDR09, LeB07v, LeB08-31, LJS^{+}09a, SR07, TWL07, YOY08]. flagellar [DS07, ELM09, LW07a, OHK07, PKK09, iWK06, WNF^{+}05b]. flagellum [DBD^{+}09]. Flatland* [AKD06]. Flavivirus [Wil08k, FSH08, Har08]. Fld1p [FSG^{+}08]. fleeting [Wil07-43]. flex [Nyb05d]. flexibility [LeB07a, Wil07-43]. Flexible [Wel06m, KKH^{+}06, KHWC08]. flick [RW08a]. flies [CTGVGFF08]. Flightless [LYY08]. Flightless-I [LYY08]. FLIM [LJS^{+}09b]. flip [BXHD08, Rob08-36]. flips [LeB07v]. Flk [KZC^{+}08]. Flk-1 [KZC^{+}08]. floated [Pow06i]. flooding [LeB07-53]. Flow [LYK^{+}06, Wel06n, CMF08, GSJ^{+}08, HLP06, LSIT^{+}08, Nyb05n, OSB^{+}05, STU^{+}08, Tum05t, Wil07-66, Wil07-67, Wil08a]. flow-elicited [CMF08]. Flow-enhanced [LYK^{+}06]. flow-induced [LSIT^{+}08, Tum05t]. flowing [Wel05c]. Flt [KZC^{+}08]. Flt-1 [KZC^{+}08]. Fluid [Pow07f, AFV^{+}05]. fluidlike [SAM^{+}07]. fluorescence [GBNK08a, GBNK08b, HKV^{+}05, Wat09]. fluorescent [DDH09]. Flux [LeB06y, CYC^{+}06, HR09, MPLF^{+}09, SGK^{+}07, YCM^{+}08, ZRBS07]. flux-dependent [MPLF^{+}09]. fluxes [LRS05]. fly [Wil09m]. FM1 [RBC05]. FM1-43 [RBC05]. fMLP [KNT^{+}07, SMG07]. FN14 [VCC^{+}08]. Focal [GPC^{+}06, Les09o, SRN^{+}07, WKWL06, CCH09, EBMG09, GSJ^{+}08, GJL^{+}08, HWS^{+}07, LSdWL07, LLT^{+}08, OPT^{+}07, SPA^{+}05, SCS^{+}05, TSN^{+}06, WKIKG09, YHJ^{+}05]. foci [PL06]. focus [LeB07z]. focusing [GNV05]. fold [Rob08-49, SSS^{+}08b]. Folding [LeB05-38, SLP^{+}06, GOM^{+}09, KGL07, LeB06h, Sav09a, SNA^{+}07, SPK^{+}09, Ska07, Wel07e]. follicle [AOP^{+}05, LGT^{+}07, OLS08, SCFKB07]. Follistatin [PBD^{+}06]. follow [FHMC09, Kat06c, Les09f, Wel05-62]. Following [Wil09h]. force [Bas07e, KO07, MFC06, MPLF^{+}09, PJMR06, SD07, YHJ^{+}05]. force-generating [SD07]. forces [Rob08-44, TWG^{+}09]. forget [Les07-34]. Fork [CLL07, FPP^{+}08, LM05, SiTT^{+}08, Tum05p, Tum05-67, Thu07]. forkhead [SNG^{+}05]. forks [CCS^{+}08, YM09]. form [BLH^{+}09, Bon05c, BBPM07, qGyWX^{+}08, KBGG^{+}08, LZ08, LCL^{+}07, TAM^{+}06, Wil07d]. Formation [EMPK^{+}07, KCOS06, Les08-63, Pol08, RSS^{+}09, ARKS^{+}08, AH08, AVHH09a, AVHH09b, AWM^{+}08, BRKR06, BMG^{+}07, BHK^{+}06, COC^{+}05, CSM^{+}06, CYO^{+}09, DLHW09, DNG^{+}06, DDLC^{+}06, FBD^{+}05b, FvD08, FBD^{+}08, GFH^{+}06, GSL^{+}07, HTK^{+}08, HBD^{+}05, HWS^{+}07, IUT^{+}07, IKZ^{+}07, IDTT^{+}09, JSPT05, KRZ^{+}08, KG05, KKY^{+}05, KIKI^{+}07, KOKM07, KES^{+}09, LCO07, LYH^{+}05, LLT^{+}08, LLHH07, LRJY06, LFL^{+}05, MES^{+}08, MKB^{+}09, MCRVC^{+}08, MSSF05, MAZ^{+}08, MAK^{+}06, MMNH09, NTS^{+}05, NBM^{+}08, NSS^{+}08, NBS^{+}08, NSO^{+}09, NYN^{+}06, NIC^{+}09b, OIT08, PDM^{+}08, PXD^{+}07, Pow05i, QIH^{+}06, SYM^{+}08a, SMSPK08, SYKM08, STBL09, SSM05, SAM^{+}07, SMM08b, STP08, SMG07, TRS^{+}08, TMI07, TSALS05, TA07, Tum05-59, Tum06i, WTE^{+}08, WLS^{+}09, Wel06c, Wel06-49, WKWL06, WKIKG09, WF07, YLK^{+}05, YBT^{+}09, iYEF^{+}07, ZHI^{+}08, ZEC^{+}08]. formed [SCS^{+}05, VBG^{+}06]. Formin [LeB05-39, Tum06r, BMS^{+}08, LeB08l, SWD^{+}08, SOR^{+}05]. forming [DMDO08, LDL^{+}09a, LRCH^{+}05, SWL^{+}05]. formins [SOR06]. forms [BSW^{+}05, BBF^{+}07, BLB^{+}05, IPM^{+}05]. forth [Les07-27]. Forward [Rob07e].

Fos [GNPC^{+}08]. Four [Rob08m, VKGV08, GHA^{+}05, Rou08, PWW^{+}09]. FOXC2 [NIC^{+}09b]. FoxK [CTGVGFF08]. FOXO [HMS^{+}08, LPWP^{+}07]. FOXO-regulated [HMS^{+}08]. FOXO1a [BIH^{+}05, BIH^{+}07]. fracture [Wel05-31]. Fragile [Tum05y, PPBF08]. fragility [Wil06f]. fragment [TdVM^{+}07, XMA^{+}09]. fragmentation [CBML06, LU05, SS06b]. fragments [CYY^{+}08, Les08-32, RW05]. Francisco [LK06, dSW09]. Frand [Sed08a]. Fraser [Wil07-57]. Frauke [Sed09b]. fray [LeB07-49]. free [FBD^{+}08, HSLI09, LeB05r, Les07g, OW06, SMG07, Wil06h]. Freedom [LeB05-40, Wel05w, Les09i]. freely [CBMCSS08]. Freeman [Wil09m]. frees [LeB07-42]. freeze [Wel05-31]. frequency [SWM^{+}07]. frequent [MTH^{+}06]. frequently [BLR^{+}06]. FRET [CHSV07, LJS^{+}09b]. Friedl [Sed09j]. Frizzled [KHH08, WZCB^{+}06]. FRL [SOR06]. Frog [Pow05l]. front [GCSG^{+}05, JLL^{+}07, Ste08, Tum06g, VWK^{+}06]. Frontiers [lWHMH06b]. Fstl1 [RGA^{+}06]. Fuchs [Sho09u]. fuel [FA08]. fuels [SS06a]. Full [Wel06o, AHD^{+}09]. full-length [AHD^{+}09]. fully [PPCZ06]. fumigatus [PUG^{+}06]. Function [HDM^{+}07, PSS^{+}05, ASH06, ASSST06, BLL^{+}06, BLP^{+}05, BBFN07, BLA^{+}09, CMF^{+}09, CCP^{+}08, CFBM05, DP06, DTP07, DTR^{+}05, DLH^{+}09, EFM^{+}07, FSN^{+}06, FMA^{+}09, FSK^{+}07, GHX^{+}07, HWN^{+}07, HMM^{+}05, KO06, LeB06-50, LKC05, LSC^{+}07, LSC^{+}06a, LCR^{+}05, LVG^{+}09, MMCW06, MXR^{+}05, MBS^{+}08, MMNH09, NAF09, NS06, Nyb05j, NKK^{+}06, OYKY^{+}05, OD07, PR05, PLH^{+}08, PG06, PT09, RGW^{+}07, RBJL06, RDY^{+}05, SSS^{+}08a, SB09, Sed07e, STG05, SOW^{+}09, SXH^{+}06, Smi07, SGKP^{+}07a, SGKP^{+}07b, TS06, VDSV06, VTT^{+}09b, WCR^{+}05, WPL^{+}06, WP08, Wil09m, WOWB08, XNM^{+}09, YBK^{+}06, YPG05, YCX^{+}08, ZS08, ZMW^{+}08, dCD^{+}07]. Functional [ESG^{+}06c, FSR^{+}07, HQF^{+}07, LPM^{+}07, MHM^{+}07, OSI^{+}07, VBG^{+}06, iYEF^{+}07, BSDA07, CCC^{+}05, CRV^{+}06, FKJ^{+}09, FSG^{+}08, GSHD08, HGJ06, JS07, JYZL06, Joh09, LSD^{+}06, MYO^{+}05, NS07, PCN^{+}05, Ped08, ZOH^{+}08, LeB06-52]. functionally [AVHH09a, AVHH09b, FML^{+}06, GFA^{+}09, KBS^{+}06, RWB^{+}05, SYA^{+}08]. functions [AZP^{+}05, BSW^{+}05, BJZ^{+}09, BDW^{+}09, CLG^{+}07, DBDD06, DBD^{+}09, GSC^{+}07, HHY^{+}08, JTN^{+}05, JhKC08, KSL^{+}05, KKM09b, LK05, MLC^{+}06, MYMZ08, MS08, MWSK06, MDWH^{+}08, MBS^{+}08, MBF^{+}07, MBF^{+}08, NOH^{+}07, OOS07, OOS09, PLLH05, PV06, RGL^{+}09, RMZ^{+}06, RKT^{+}07, SRX^{+}07, STM^{+}07, SSSB07, THS^{+}09, VVL^{+}08, VBSM06, VTRB06, WKH^{+}06, WZC^{+}08, WTE^{+}08, WL09b, YWEK08]. funded [Ros05]. fungal [KBC^{+}07a]. Funky [Tum06s]. furrow [ACsHS08, CAR^{+}08, FvD08, GT08, Les08a, NSO^{+}09, OF08, Tum05-59]. furrows [LeB08w]. fury [Wel06o]. Fus2p [PYR08, Sho09l, YR09]. fuse [Pow05j]. fused [HS09b]. Fusing [Les07y]. Fusion [Les07z, BbCQ^{+}09, BVGA^{+}05, DDRL^{+}09, DC07, DHML05, ERM^{+}09a, EW08, EHWL06, FSH08, GHY^{+}05, GMK^{+}05, Har08, HEW07, HHS^{+}09, JKKM08, JMG08, LeB05w, Les07-39, LK05, LMZ^{+}05, MNF^{+}07, MSW^{+}07, MSW^{+}17, PPA^{+}08, PYR08, PBD^{+}06, RM05a, RLDC09, SDL^{+}05, Sho09o, Sho09-33, SCF^{+}07, Tum06-33, YOS^{+}05, ZMGC05]. fusions [LeB05-101]. Fyn [CMF08, WGKA^{+}08]. FZO [RLDC09]. FZO-1 [RLDC09]. FZO-1/Mfn1 [RLDC09]. Fzo1 [EHWL06].

G [KTI^{+}05, LeB05-41, Tum05z, BFB^{+}08, CK07, CLG^{+}07, DTL^{+}07, DZG^{+}07, EFG^{+}05, FNH^{+}05, GRC^{+}05, GGMBF08, Gle07, GCS^{+}07, Les08r, LSC^{+}05, LCS^{+}09, LPB^{+}08, NHK^{+}08, SJL^{+}07, SGKM08, SFS^{+}07, TSP^{+}07, RSS^{+}09]. G-actin [NHK^{+}08]. G-dependent [DZG^{+}07, LPB^{+}08]. G1 [RMC^{+}05, ZNP^{+}09]. G2 [CWC^{+}07, DBK^{+}07, DSBK09, MAG^{+}08]. G2/M [DSBK09]. Gab2 [ML05]. GABA [Les09f, Rob08n, pTyLZ^{+}09]. GABA-induced [pTyLZ^{+}09]. GABAergic [DCAC05]. GADD45 [SLZ^{+}06]. Gag [BFF^{+}06]. Gaga [Sav09b]. gain [Rob07n]. GAKIN [HHFC06]. galectin [GJL^{+}08, SWT^{+}05, DGZ^{+}05]. galectin-1 [SWT^{+}05]. galectin-3 [GJL^{+}08]. Galectin-4 [DGZ^{+}05]. game [Sed09i]. gang [Les08-38]. ganglion [HLJ^{+}07, JLB^{+}05]. Ganglioside [NRL^{+}05]. Ganglioside-induced [NRL^{+}05]. gap [LeB07-57, Pow06e, Wel05k, vZPG^{+}07, BPL^{+}08, CN06, LH05, NLS^{+}05, OKN06, RFBB07]. GAP-43 [LH05]. garbologist [Wil08t]. Gardner [Wil08t]. gas [Les07-74, Tum05-27]. gastrulation [LSC^{+}05, LHM^{+}09, YKP^{+}08]. Gata [dGSWW^{+}06]. Gata-3 [dGSWW^{+}06]. gate [Hat05a, LK06]. gated [BH08a, LPB^{+}08, LYV^{+}08]. gatekeeper [SMRS^{+}07]. gateways [NKK^{+}06]. gating [ASB^{+}05b]. Gawky [SNC^{+}06]. GC [LeB06b]. GEECs [Nic09a]. GEF [OIAT06]. GEF1 [HSB^{+}07]. gel [Les05g]. Gellular [Wel06p]. Gene [LC08a, CFH^{+}08, DBvdW^{+}06, DOS^{+}07, GLS^{+}09, GDF^{+}05, GQSHO05, GFC^{+}08, JDK^{+}05, KBW^{+}06, KCD^{+}06a, Les08-45, Les09q, MLS^{+}07, McN09, MM08, MGK^{+}09, OMLP^{+}08, PTL^{+}05, SKM^{+}08, SLP^{+}06, SSAD08, WKIKG09]. gene-expression [KCD^{+}06a]. generate [BYK08, NZQ^{+}07, SNA^{+}07, SWI06]. generated [DCAC05, HL06, TWG^{+}09, LJR^{+}05]. generates [CHO^{+}09b, DHML05, TDP^{+}06, ZMA^{+}07]. generating [KOKM07, MHS^{+}09, SD07, YOY08, YCK^{+}05]. Generation [HR05, GMZ^{+}08, HZL^{+}07, Miz05, PCG^{+}05, Pow09c, WJM07]. Generous [LeB05-42]. Genes [LeB06z, Sho09x, AMERT^{+}07, BLR^{+}06, BGdN^{+}08, CMD^{+}08, CF07, FMA^{+}09, GFC^{+}08, HKP^{+}09, LW07b, MTRD^{+}06, Rob08y, SPS^{+}08, Tum06-31, Wel06-53, XC08]. Genetic [CML^{+}06, FWC^{+}07, HJK^{+}06, Les08t, MAZ^{+}08, GSHD08, KS08, OHP^{+}09, YCX^{+}08, ZHS06]. genetically [IH05]. genetics [HDS^{+}07]. Genome [BFWH07, EMB^{+}08, Kat06f, SKM^{+}08, BJLK^{+}06, DLH^{+}09, EHHK09, FMA^{+}09, FMD^{+}09, TMH^{+}07, YSB^{+}07]. Genome-wide [BFWH07, EMB^{+}08, SKM^{+}08, FMA^{+}09, YSB^{+}07]. genomes [OLH^{+}08]. genomic [LYWK08, MGW^{+}08, MSM^{+}06, SLP^{+}06, TDD^{+}05, Wil07-45]. genomics [MHM^{+}07]. genotoxic [MCD^{+}07]. genotypes [CGK07]. geodetic [Heu05]. geometric [CFBM05]. geometry [LeB06-66, NFZ^{+}09]. George [Mel08b]. gephyrin [MTL^{+}06]. Germ [Sho09y, EIK^{+}05, KSR^{+}08, PMB^{+}09, SRL05, WC07a]. germline [DL08, HKN09, LTK09, SRQL09, WLC08]. Gero [Wel07f]. get [LeB07-56, Les07-31, Les08u, Les09t, Mel06, Pow06a, Ras08a, RM07, Wel05d, Wel05i, Wel05-41, Wel06-35]. get-together [Les08u]. gets [Bon05j, CS07, LeB05-30, LeB05-37, Les09r, Rob08q, RW08e, Tum05-81]. Getting [Gle07, Les06k, Wel05y, Wel05z, Les09-29, Sed08g, Sed09c, Ros08c, Sho09-53, Wil08b]. GFP [DvCvH^{+}08]. Gia [Sed08h]. giant [LeB06-28, ZWG^{+}05]. Gibson [Sho09-36]. Gilbert [Wil07w]. Giraldez [Wil09b]. GIT1 [NWB^{+}06]. GIV [GGMBF08]. give [LeB06-56, Sho09-39, VMY^{+}09]. gives [Sed08f, Sho09-27, Sho09a]. gland [Les07q, SKR^{+}07]. glandular [NLC^{+}05]. glaucoma [HLJ^{+}07]. glia [PT07, TDG^{+}06, Wil09m]. Glial [ZTM^{+}08, LeB05l, SFY^{+}05]. gliomedin [EFCP07]. gliotoxin [PUG^{+}06]. gliotoxin-induced [PUG^{+}06]. Global [MJF^{+}07, JNHB08, WSG06]. globally [Ter06]. globin [BLR^{+}06]. globular [Ped08]. glucocorticoid [KTI^{+}05]. glucose [YZO^{+}07]. glucosylceramide [HNvD^{+}07]. glue [KWA^{+}05, PT07, Wel06-64]. GLUT1 [VKR^{+}09]. GLUT4 [LMZ^{+}05, Tum05-28, WP08, YOS^{+}05]. GLUT4-containing [YOS^{+}05]. glutamate [KVC^{+}09, OWM05, SSD06]. glutamine [DTP07, YZO^{+}07]. glutamine/asparagine [DTP07]. glutamine/asparagine-rich [DTP07]. glutaredoxin [AAG^{+}09]. glutathionylation [AAG^{+}09]. Gly [LeB06-58, PTF^{+}06]. Glyburide [LMV^{+}09]. Glycan [LeB07q, CHO^{+}09b]. glycans [GOM^{+}09, SN05]. glycine [MTL^{+}06, Rob08n]. Glycogen [BYF^{+}05]. Glycolipids [Tum05-29]. glycoprotein [CHO^{+}09b, DNG^{+}06, KAL^{+}06, PGWH08, PMC08, SGK^{+}07, VCSG07]. glycoproteins [NTHN^{+}08, SN05]. glycosaminoglycan [RRA^{+}09]. glycosaminoglycan-binding [RRA^{+}09]. glycosphingolipid [HNvD^{+}07, SHjC^{+}07]. glycosyl [CTY^{+}06]. glycosyl-phosphatidylinositol [CTY^{+}06]. Glycosylation [Wil07x, FKB^{+}05, JASC08]. glycosylphosphatidylinositol [BLC^{+}09]. glycosylphosphatidylinositol-anchored [BLC^{+}09]. glypican [DLBSD^{+}05]. glypican-1 [DLBSD^{+}05]. glypicans [HKN09]. GM1 [VSR^{+}08]. GMP [PBD^{+}06]. GNL3L [ZMH^{+}09]. Go [Les07-27, GG09, Les07l, Les07s, Les08w, Les09-37, Pow05n, Rob08-52, Tum05-28, Wel05-36, Wil07i, Wil07-66]. Godfather [Wil08u]. goes [Les08l, Les09-28, Sco08, Tum05-43, Wel05-28, Wel05-51, Wil08o]. Going [Les07-28, Mis05]. Golden [LK06]. Golgi [ARKS^{+}08, AM05b, Ban09, BE05, BBF^{+}07, BWW^{+}08, BCR^{+}08, CZP^{+}05, EGL^{+}09, FHFB^{+}06, GEP08, GDR^{+}05, GL06, HNvD^{+}07, HEW07, HAAHD05, JKW^{+}09, KES^{+}09, LeB05-43, LeB07g, LeB07r, LeB08v, LDPLS05, NYM^{+}08, Pow05v, SRA^{+}05, Sav09b, SB09, STBL09, SS06b, SSP^{+}07, Sho09-32, Sho09-42, SGKM08, SHDT05, WHOS06, WRP^{+}06, WS09, Wel06q, Wil07y, WSB^{+}09, ZL05, dGHW08, vBDF^{+}09]. Golgi-localized [FHFB^{+}06]. Golgi-resident [HEW07]. Goliath [Ros09]. GOLPH3 [WSB^{+}09]. gonad [ZRH^{+}09]. Gonadotropin [PKJS^{+}08, GMM^{+}08].

Gonadotropin-releasing [PKJS^{+}08]. Gone [MG08b, LeB06-37, LeB07-62]. good [Les08-37, Sho09w]. goods [Wil07-37]. got [Les08-46]. govern [TTNA^{+}06]. governs [CTS^{+}06c, PGO^{+}07, PWD^{+}06]. Gp135 [MSSF05]. Gp135/ [MSSF05]. gp210 [SNPCG06]. gp210-deficient [SNPCG06]. GPCR [KSR^{+}08, XMSYJ07]. GPCR-mediated [XMSYJ07]. GPI [CZM^{+}07, NS06, NS07, PPCZ06, SS06a, SFEK07, SFS^{+}07]. GPI-anchored [CZM^{+}07, NS06, SS06a, SFEK07, PPCZ06, SFS^{+}07]. GPR3 [FNH^{+}05]. Gq [EFG^{+}05]. gRab [Rob08-39]. Gradient [Nyb05e, CCSBM^{+}07, LeB07f, SH05]. gradient-sensing [SH05]. gradients [WHB^{+}07]. granule [BMP08, DP06]. granules [AOT06, AK06, DR07, GMK^{+}05, HCC^{+}07, KSA^{+}05, NYI^{+}05b, NAG^{+}08, UZ08, WGKA^{+}08]. Granuphilin [GMK^{+}05]. granzyme [CAL^{+}07, SWB^{+}07]. granzymes [KBS^{+}06, WSS^{+}06]. Grapes [RMKS08]. GRASP65 [BBF^{+}07, BWW^{+}08]. Grd19 [SSSB07]. Grd19/ [SSSB07]. great [Wil07a]. Gremlin [FKS^{+}06]. gross [BSO^{+}08]. Group [Wel06r, SSH^{+}05, AMERT^{+}07]. grow [Rob08-33]. growing [BKLT^{+}08, Wit08]. Growth [Bon05d, DLBSD^{+}05, HK06, LeB07s, Tum05-78, AOP^{+}05, ANI^{+}07, BSK07, BLT09b, BCT06, BCR^{+}08, DBR09, FKB^{+}05, FBY^{+}06, FHS^{+}07a, GEE07, GFD^{+}05, HHT^{+}07, HFY^{+}06, HXZ^{+}07, HCN06, HIK^{+}08, HLA^{+}06, HFF07b, JP06, KDG^{+}09, LMMHP^{+}05, LeB06m, LeB06-72, LeB07q, LeB07-29, LeB07-44, LeB07-59, LeB08o, LSNY08, Les09w, LSdWL07, LJY^{+}07, LFS^{+}05, MYMH09, MPG^{+}09, MH07, MB07, NTB^{+}06, NPV^{+}07, OLLG05, OFY^{+}05, OTM^{+}07, PTL^{+}05, PCMSB09, PSS^{+}05, RBJL06, RRP^{+}08, SGB^{+}06, Sed09a, ShCM^{+}06, SYM^{+}08b, SSK^{+}09, TVP^{+}06, VCN^{+}05, WHB^{+}07, Wil07-35, YMDL09, ZTKK05]. growth-dependent [FKB^{+}05]. gruberi [KKY^{+}05]. GSK [RP05]. GSK-3 [RP05]. GSK3 [KLG^{+}09, Sho09z, WWS05b]. GT1 [PGWH08]. GTP [BHG^{+}05, KCE^{+}06, LDGK08, Mis05, RWM^{+}05, SCM^{+}07, SSE05, TM05, Tum05-79, WAKS08]. GTP-binding [KCE^{+}06]. GTP-driven [TM05]. GTPase [BPL^{+}08, FHS^{+}07a, GAG^{+}06, GFD^{+}05, KBN^{+}06, LSNY08, MYMZ08, PAV^{+}06, QIH^{+}06, RCS^{+}09, RWM^{+}05, oSCW07, TGH^{+}07, WZW^{+}09]. GTPase-activating [FHS^{+}07a, KBN^{+}06, TGH^{+}07]. GTPase-dependent [RCS^{+}09]. GTPases [GGEd^{+}07, HFF07b, LH05, PG06, SHB^{+}07, SWG^{+}05, iYEF^{+}07]. guanine [LLT^{+}08]. guanosine [TTO^{+}08]. guanylyl [SSJ^{+}07]. guard [Les09b, Wil09q]. guards [LeB06v]. guidance [HHT^{+}07, HKIY09, Pow07e, WZC^{+}08, ZMA^{+}07]. guide [Mal06, Nor06]. guiding [VML06]. gulp [Rob08-48]. Gurdon [Wil08u]. gut [LeB06-34, Sho09-39]. gyroscope [Pow05n].

H [Rob08o, SSD^{+}09, DKR^{+}06, FKN^{+}07, GMG^{+}07, PB05]. H/ACA [DKR^{+}06]. H1 [AH05, FSTI06, LeB06-27, MFH05]. H2A [KTA^{+}06, MBJM^{+}09]. H2AX [KKD^{+}07, LBS^{+}08]. H2B [FD09, KTA^{+}06]. H2BK123 [NLG^{+}09]. H2BK123Rgument [LD09]. H3 [BvOM^{+}07, FD09, GSY^{+}09, HTYNK09, iHBD08, XGX^{+}09]. H3K4 [NLG^{+}09]. H3K79 [NLG^{+}09]. H4 [SS07]. hair [AOP^{+}05, LGT^{+}07, OTM^{+}07, PST^{+}06, SCFKB07, Wil07-35]. hairs [VCN^{+}05]. hairy [LeB05-44]. half [Pow05o]. half-spindle [Pow05o]. halt [Tum06-30]. halting [LMZ^{+}05]. Hand [Rob07f, GG09]. handcuff [ZKS^{+}08]. handle [Sed09c]. Handy [Wil06g]. Hang [LeB07u]. HAP40 [PSL^{+}06]. haploinsufficient [BJM^{+}06]. hare [LeB05-89, Wel06-62]. Hauling [Les08v]. HB [HIK^{+}08]. HDAC3 [ZXM^{+}06]. HDAC4 [LRS05]. HDAC6 [OLC^{+}07]. head [CLL07, SMH^{+}09, Thu07]. headpiece [PFGSV06]. Heald [Sed08n]. healer [Les08p]. Healing [Les07-29, BFG^{+}06b, CBM^{+}07, CTP^{+}09, Wel06i, WHM^{+}07, WFJ06]. heals [Sho09-52]. health [BW08]. healthy [Wil09l]. hearing [Wel05-48]. Heart [Wil07z, Wil09i, BH08a, Gle07, HB08, LeB05-61, LPB^{+}08, PDM^{+}08, SMSPK08, TMW^{+}05, Wil08b]. Heat [MG08a, KFK^{+}05, YOY08]. Heather [Ros08c]. heavy [GMSS06, KKM^{+}08, RG06b]. HECT [MSECB05, ORM^{+}08]. Hedgehog [LeB06-28, GQSHO05, KAD^{+}08, JLG^{+}06, LYWK08, LBG^{+}06, WMM^{+}09]. Hedgehog-related [LBG^{+}06]. hedgehog-responsive [GQSHO05]. hedgehog/Hes1 [WMM^{+}09]. HEK293 [TDP^{+}06]. HeLa [LEH^{+}05, YHC^{+}06, ZL05]. held [LeB06-45]. Helenius [Wil08d]. helical [SGKP^{+}07a, SGKP^{+}07b]. helicase [AHM^{+}05, FMD^{+}09, LeB05-91, PPBF08, TNS07]. helicase-dependent [TNS07]. help [Ras08a]. Helping [Les07-30, Les08w, Les08x]. helps [LeB07r, Les09a, MWMS07]. hemagglutinin [HKV^{+}05]. hemifusion [GHY^{+}05, RM05a]. hemocyte [WFJ06]. heparan [FWC^{+}07, KKH^{+}06, KSpLL06]. Heparanase [MBR^{+}06]. hepatitis [Wil08f]. hepatocyte [CSM^{+}06]. here [Les07-31, Nyb05c]. Herp [TTE^{+}07]. herpes [CTS^{+}06b]. Herpesvirus [DNG^{+}06]. Hes1 [MOM^{+}06, WMM^{+}09]. Heterochromatin [KKD^{+}07, LDL^{+}09b, Wel05-27, BY09, GER^{+}07, KWA^{+}05, LG07, LG08, PL06, WSG06]. heterodimer [LLL^{+}06]. heterodimers [NMO^{+}07]. heterogeneities [STP08]. Heterogeneity [LLS^{+}09]. heterophilic [TMH^{+}06]. Heterotrimeric [CLG^{+}07]. hexagonal [NFZ^{+}09]. Hey [Les07-31, Wil07-27]. HGF [BCT06, PTL^{+}05]. HGF-induced [BCT06]. HGF/Met [PTL^{+}05]. HGF/Met-induced [PTL^{+}05]. HHMI [CDM^{+}07]. Hidde [LeB07t]. Hidden [LeB06-29, Les09-33]. Hiding [Tum05-30]. hierarchy [VFRJ^{+}07]. HIF [LLH^{+}09, XJF05, PZG^{+}07]. HIF-1 [LLH^{+}09, XJF05, PZG^{+}07]. Hif1 [Tum05-71]. High [CLS^{+}06, HBD^{+}05, LeB08-27, SSH^{+}05, Wil06g]. high-resolution [Wil06g]. Higher [RCKD06, DLT^{+}07, MMC08]. Highly [Wel05-28, MAK^{+}06, SSS^{+}08b]. Highwire [Sho09f]. hinders [LeB07b]. hinge [LYK^{+}06, PFGSV06]. hinges [Tho06]. HINT1 [LBS^{+}08]. Hip [SLK^{+}08, VD05]. Hip-2 [SLK^{+}08]. Hippel [SGB^{+}06]. hippocampal [DTL^{+}07, FM05, PKJS^{+}08, RBC05, XLH^{+}07]. histidines [FSH08]. Histone [iHBD08, LeB06-30, MFH05, NLG^{+}09, Rob08p, SLCB05, WLS^{+}09, AH05, BvOM^{+}07, CMD^{+}08, FSTI06, GTB^{+}07, GER^{+}07, GSY^{+}09, HGJ06, HTYNK09, JET^{+}07, KTA^{+}06, MBF^{+}07, MRK^{+}07, MBF^{+}08, SS07, XGX^{+}09]. histones [LeB06-74]. History [Tum05-31]. hit [Rob08w]. hitch [DP06]. hitches [Les08a]. hitchhiker [Mal06]. hits [Les07-50]. HIV [BMB^{+}07, BFF^{+}06, CPAH08, DPMB^{+}07, KDD^{+}09, LeB08t, LeB09a, Les08-38, NKK^{+}06, Wel06t, Wil07-33, Wil07-34]. HIV-1 [BMB^{+}07, CPAH08, DPMB^{+}07, KDD^{+}09, NKK^{+}06, Wel06t]. HMGA [Wil06k]. HMGA2 [TVP^{+}06]. HMGB1 [PGP^{+}07]. hnRNP [KHBV05, WGKA^{+}08]. Hoarding [LeB06-31]. Hochedlinger [Wil07-32]. Hof1 [NSO^{+}09]. Hold [Les06l, Wel05-35]. Holding [Tum05-32]. holds [Tum05r]. hole [LeB06-81, Wel06l]. hollow [HB08]. home [Les08-33, Wil07-34]. homeodomain [LCL^{+}07]. homeostasis [CTP^{+}09, TTK^{+}06, TTE^{+}07, VTT^{+}09a, VMY^{+}09]. Homing [Tum06t, NS06]. homologous [AMERT^{+}07, BSO^{+}08, BFCS^{+}07, SiTT^{+}08, TNS07]. homologue [FSG^{+}08, JGY09, MWY^{+}09, MDH^{+}06, MLP06, NSS^{+}08, PCMSB09, SHS^{+}09, WZC^{+}08, WOP^{+}06, ZSMW09]. homologues [CS09, SCR^{+}09]. homology [BWFdL09, JMG^{+}06, POGL08]. homology-directed [BWFdL09]. Homophilic [MKV^{+}09]. homotypic [STBL09, YSP^{+}06]. honeycomb [Heu05, SWL^{+}05]. hooks [LeB07n]. Hoon [Wil08v]. hops [LeB08y, LU05]. Hormone [LCM^{+}08, CMD^{+}08, GMM^{+}08, LeB07-45, MGR^{+}05, PKJS^{+}08, Wil07z]. hormone-1 [GMM^{+}08]. Hormone-dependent [LCM^{+}08]. hormone-mediated [CMD^{+}08]. hormone-releasing [GMM^{+}08]. Horsley [Sho09-53]. host [INB06, MBR^{+}05, PUG^{+}06, PG06]. hotspots [WKWL06]. house [Kat06e, Les08o, Sed08l]. housekeeping [BBFN07]. Howard [LeB07u]. HoxD [Kat06h]. HP1 [ABY^{+}08, JZN^{+}09]. HP1- [ABY^{+}08]. hPOC5 [AHD^{+}09]. HR [KKK^{+}05]. HRas [FM05]. Hrd3p [MLP06]. hRio2 [ZWO^{+}09]. hRME [SSM^{+}08]. hRME-6 [SSM^{+}08]. Hrp59 [KHBV05]. Hrs [RRP^{+}08]. hsc70 [KFK^{+}05]. hSgo2 [HFF^{+}07a]. Hsp70 [SMM^{+}08a]. Hsp70-2 [SMM^{+}08a]. Hsp70-2/HspA2 [SMM^{+}08a]. Hsp90 [BMGPB^{+}08, SZL^{+}09, ZKG^{+}08]. Hsp90a [ERS08]. HspA2 [SMM^{+}08a]. hSSB1 [SRS^{+}09]. hSSB1-mediated [SRS^{+}09]. Htm1 [CHO^{+}09b]. Htm1p [Sed09c]. hub [VD05]. hubs [LeB05-35]. Human [CAL^{+}07, FMD^{+}09, KAD^{+}08, MSM^{+}06, BTD^{+}07, BH08b, BLR^{+}06, CS09, CFU^{+}09, CMN^{+}08, DTY^{+}05b, DKR^{+}06, DDLC^{+}06, DHML05, FSG^{+}08, FFS06, FKS^{+}06, FKML^{+}06, GEP08, KCE^{+}06, KBI^{+}08, KBS^{+}06, KCH^{+}06, LMMHP^{+}05, LeB07-45, LRJY06, LVG^{+}09, MRKH05, PHB^{+}07, RG06a, SHK^{+}07, TTO^{+}08, TC08, TD08, ULN^{+}07, VFRJ^{+}07, YZO^{+}07, ZDD^{+}06, ZNP^{+}09]. HUMMR [LLH^{+}09]. Hungry [LeB08j, Les07-34, Pow06f]. Hunter [Wil08w]. Hunting [Les06m]. Huntingtin [MGR^{+}05, OVL^{+}07, BLB^{+}05, LeB05o, LVDR05, SFY^{+}05, WZM^{+}08, YCR06, PSL^{+}06, TAK^{+}09, Tum06y]. Huntingtin-HAP40 [PSL^{+}06]. huntingtin-induced [BLB^{+}05]. Huntingtin-interacting [OVL^{+}07]. Huntington [Les08-28, PSL^{+}06, Sho09-29, WWO^{+}08]. HuR [MDC^{+}08]. HURP [Wil06b, WF06]. Hwang [Ros07b]. Hyaluronan [KZH^{+}07]. Hydin [LeB07v, Smi07, LW07a, LDR^{+}08]. Hydrogen [GH05, MMM^{+}05]. hydrolysis [BHG^{+}05, CCC^{+}06, HLK^{+}09, RWM^{+}05, SDB^{+}05, SSE05, WAKS08, vRSvR^{+}07]. hydrophilic [KMS07]. hydrophobic [BLC05]. hydroxylase [PCMSB09]. hydroxylated [WHR^{+}09]. Hyman [Wil07c]. hypercitrullination [WLS^{+}09]. Hyperdynamic [Kat06g]. hyperosmotic [ZPGB^{+}07]. hyperplasia [JLG^{+}06]. hypersensitivity [BvRJ^{+}05]. hypertrophy [ZBKJ07]. hypothesis [KMK^{+}05b]. Hypoxia [BBG^{+}05, LLH^{+}09]. hypoxia- [LLH^{+}09]. Hypoxia-inducible [BBG^{+}05]. Hypoxic [LeB07w, BKE^{+}07, LJM^{+}07]. Hyung [Wil09l]. hZW10 [FVSC08]. hZwint [FVSC08]. hZwint-1 [FVSC08].

I-mediated [SFC^{+}08]. i3 [GGMBF08]. IA [GEE07]. Iain [Sho09-28]. IAP [HCB^{+}08]. IAP1 [KKM09b]. IAPs [GHM^{+}09]. ICAM [TSN^{+}07]. ICAM-5 [TSN^{+}07]. ICAM1 [vBAS^{+}07]. ICAP [LeB08k, MFBG^{+}08]. ICAP-1-dependent [MFBG^{+}08]. iceberg [Wil09b]. Icrac [CLK^{+}07]. idea [Pow06i]. ideas [Sed08a, Sed08h]. Identification [CJMG05, CMN^{+}08, FHFB^{+}06, FSH08, GMW^{+}05, GFA^{+}09, MMU^{+}09, NS07, NRW^{+}08, OBY^{+}06, THY^{+}06, UFY^{+}07, KFC^{+}05]. identified [BMBD^{+}09, MCD^{+}07]. identifies [MHM^{+}07, TTG^{+}09]. identify [KPR^{+}07a, KPR^{+}07b]. Identifying [TMBL^{+}08]. identities [Wel05r]. identity [Les08d, Sed09i]. If [Rob08r, HK07, LeB05-66]. IFN [MDET08]. IFT [Den05, ESG^{+}06c, HQF^{+}07, LJS^{+}09a, POCS^{+}06]. IFT20 [JSFP08]. IFT46 [HQF^{+}07]. Ig [BMK^{+}07, KVS^{+}09, SSS^{+}08b]. Ig-fold [SSS^{+}08b]. IgE [LGH^{+}05, Too05]. IGF [RYD08, SDL^{+}05, XKYW07, DGP^{+}05]. Igf-1 [DGP^{+}05]. IGF-1R [XKYW07]. IGF-I [SDL^{+}05]. IGFBP [RYD08]. IGFBP-5 [RYD08]. IgG [TMY^{+}09]. Ihh [BLT09b]. II [BBM^{+}08, ADC^{+}07, CM07, CHCB05, CVACF07, GCS^{+}07, HHT^{+}07, HLP06, JLM^{+}07, KPS^{+}07, LeB06-46, Les07-51, LK05, LGBD^{+}07, MHL^{+}06, MCF^{+}05, MBS^{+}08, MRL^{+}08, POCS^{+}06, PNB07, PWD^{+}06, ROKG08, RYD08, TD08, VLFA^{+}07, WTE^{+}08, WVB07, YKH^{+}08, YMC05, ZMGC05]. IIA [KCOS06, VMZW^{+}07a, VMZW^{+}07b]. IIb [WBP^{+}08, YLM^{+}06b, VMZW^{+}07a, VMZW^{+}07b, VMKW^{+}08]. III [AWR^{+}07, BKE^{+}07, HCRT08, HRLH08, JHY^{+}08, LK05, NWO06b, ST09]. IKK [BWL^{+}08, Sho09-29, SLZ^{+}06, TAK^{+}09]. IKK- [NEF^{+}06]. IKK/ [BWL^{+}08]. IL-1 [CTP^{+}09]. IL-3R [DRD^{+}05]. IL-6 [MKH^{+}07]. I’m [Wil07-27]. image [Hil08, Nor06]. Imaging [OIFN^{+}07, Sed07c, Wil07-30, HSTM06, PH07, SWG^{+}05, lWHMH06b, YTH^{+}09, lWHMH06b]. immature [CZSG08, SWL^{+}05]. immobilization [CNMM^{+}06, SFS^{+}07]. immobilized [SFEK07]. Immortal [Tum05-33, KMK^{+}05b]. Immune [IKZ^{+}07, Wel06u, LeB07-32, LHRP05, Sed07c, SX08, STP08]. Immunity [LeB07x, Bas07d, Wel05-42]. immunoglobulin [TLK^{+}06]. immunoglobulin-like [TLK^{+}06]. immunological [BUJ^{+}08, Rob07a, TLK^{+}06]. Immunologist [LeB07t]. immunology [Sed09j]. immunoreceptors [ZKR^{+}07]. IMP [MKSJ06]. IMP-binding [MKSJ06]. impair [LDR^{+}08, SNA^{+}07]. Impaired [ECR^{+}09, WWO^{+}08, LHS^{+}05, NTSY08, TT09, WBL^{+}06]. Impairment [KWU^{+}05, dGSWW^{+}06]. impairs [HTKR06, KST^{+}07, LYWK08, RRA^{+}09, WH06, WHM^{+}07]. implicates [WZM^{+}08]. implication [BSMT05, CK07]. implications [BW08, CMN^{+}08, VBG^{+}06]. import [CCC^{+}06, FHO^{+}05, GGE^{+}09, KFK^{+}05, KGM^{+}07, LeB05-98, LZM^{+}06, Les06n, MF08, PES^{+}07, RM05b, SMM^{+}08c, THS^{+}09, TTNA^{+}06, Tum06-46, WAKS08, XYC^{+}07, YM06]. Important [WKW^{+}08, GFA^{+}09, SYA^{+}08, WEL^{+}07n]. importin [FFW^{+}07, FHO^{+}05, RM05b, YM06, TTO^{+}08]. importin- [RM05b, TTO^{+}08]. imports [Les08g]. impose [LWG^{+}05]. improper [OFfL^{+}07]. In-and-out [Tum06u]. inaccessibility [LeB07d]. inactivates [UTK^{+}09]. Inactivation [KKM^{+}08, VTT^{+}09b, HGA^{+}09, LW07b, MBdR^{+}08, RMR^{+}09, ZTZ^{+}09]. inactive [HBPL09, LMMHP^{+}05]. inactivity [Wil07-51]. INCENP [XOV^{+}09, JhKC08]. INCENP-aurora [XOV^{+}09]. included [Mat06]. inclusion [BZYM^{+}08]. incompatibility [Tum06a]. incompatible [THL^{+}06]. increased [BXN^{+}06, GHDS^{+}07, HJK^{+}06]. increases [BbCQ^{+}09, SDL^{+}05]. Independent [KKK^{+}05, LeB07y, MMH09, BCT06, CYC^{+}06, CPAH08, CVACF07, DPK^{+}05, DBW^{+}06, DBR09, EHWL06, FHNG05, GMSS06, GMZ^{+}08, GCS^{+}07, HCK^{+}07, KFC^{+}05, KKM09a, KBC^{+}08, LSdWL07, LRS05, MTRJA05, MM08, NDS^{+}07, NYI^{+}05b, PNB07, RWM^{+}05, SSHW09, SJL^{+}07, SDDV05, TNS07, WJM05, DPP^{+}07, JWSS05, SRL^{+}06, WMM^{+}09]. independently [BMS^{+}08, DLL^{+}08, DTW06, GLH^{+}05, HFvdV^{+}05, MKSJ06, ROKG08, RHMW^{+}05, SPT^{+}09, SPT^{+}21]. India [VD09]. indicates [KSL^{+}05]. indispensible [SFB09]. individual [FML^{+}06, HQF^{+}07, RBC05, Wil07-30, WvNS^{+}07]. induce [BEN^{+}05, CSM^{+}06, KPM05, LKL^{+}05, OTM^{+}07, RHMS05, THL^{+}06, WHC^{+}07]. induced [BSF^{+}08, BLB^{+}05, CLL07, CBML06, DSLY^{+}05, DLBSD^{+}05, FIK^{+}05, FIK^{+}20, GMvdKC07, HCB^{+}08, KOKM07, KWU^{+}05, KNT^{+}07, LNG^{+}05, LYWK08, LMC^{+}09, LSIT^{+}08, MFA^{+}07, ML05, MBJM^{+}09, MDET08, MCD^{+}07, MRKH05, NEF^{+}06, NRL^{+}05, PUG^{+}06, PGW^{+}08, PKM^{+}05, SKM^{+}08, SRL^{+}06, SKO^{+}06, pTyLZ^{+}09, Tum05t, WCB^{+}08, WBP^{+}08, WSS^{+}06, WPH^{+}07, WPC^{+}08, XJF05, YIfHA09, YJJ^{+}09, YLM^{+}06b, ZTZ^{+}09, vRSvR^{+}07, BCT06, GHM^{+}09, GEH^{+}07, HLL^{+}08, HPH^{+}05, KNKH06, LD07, MRD^{+}08, NYN^{+}06, PDM^{+}08, PTL^{+}05, SYKM08, VSR^{+}08, ZXM^{+}06]. inducer [MLC^{+}06]. induces [BSL^{+}08, DSN^{+}07b, FCB^{+}09, FCB^{+}19, HlQY^{+}06, IPAR05, JLC08, JYZL06, KLV^{+}08, LYWK08, LJM^{+}07, LSC^{+}06a, LLHH07, MYO^{+}05, MZH^{+}08, dJPAW^{+}09, SFB^{+}05, SYM^{+}08b, SFEK07, TCHL05, Tum06-33, UZ08, VCC^{+}08, YGY^{+}05, YZO^{+}07]. Inducible [BvRJ^{+}05, HGA^{+}09, BBG^{+}05, HKP^{+}09, PTL^{+}05, VTRB06, LLH^{+}09]. inducing [KGKV07, RGA^{+}06, STK^{+}05]. Induction [BDR^{+}06, HWJ^{+}08, JZN^{+}09, KTN07, LCT^{+}09, MAH^{+}05, MMS07, PBD^{+}06, TdVM^{+}07]. industry [Wel07h]. Inequality [LeB05-45, LeB07h, Les08-36]. infect [LeB07-67]. infected [CBG^{+}06]. Infection [LeB05-46, BXN^{+}06, DDLC^{+}06, LeB06x, LeB06-76, SHjC^{+}07]. inflammasome [LMV^{+}09]. inflammation [LSIT^{+}08, Rob07d, SWG^{+}05]. inflammatory [KBW^{+}06]. inflate [Sho09-32]. Influence [MFO09, HMG^{+}08, KCD^{+}06a, Tum05-31, VTRB06]. Influences [PVMH05, PEW^{+}05, SRXvdK06]. influenza [HKV^{+}05]. influx [LeB06-75]. information [BUJ^{+}08, HSLI09, LWG^{+}05, MPL^{+}07a]. ingest [LeB06-65]. inherent [TMY^{+}09]. Inheritance [LYTL07, CMF^{+}09, FFT^{+}05, MMNH09, Sed09e, Sho09-48, WS09]. inhibit [BWL^{+}08, GHM^{+}09, LB08]. inhibited [CTS^{+}06a, KCMB08, VBL^{+}05]. inhibiting [KST^{+}07, KZT^{+}08, KZT^{+}09, LCS^{+}09, MGH^{+}09, SFB^{+}05]. Inhibition [CNW^{+}05, DCMB^{+}07, FIK^{+}05, Mel08a, SDN^{+}05b, SHjC^{+}07, SGDM06, TD08, WSB^{+}06, XJF05, BSO^{+}08, BDR^{+}06, CM06, CCL^{+}07, DDO^{+}08, HFY^{+}06, HGJ^{+}08, LFW^{+}09a, LFW^{+}09b, MMM^{+}05, NSM^{+}09, NS06, OLLG05, OYTO07, RVBZ^{+}08, SBM^{+}07, SGMPV07, SYM^{+}08a, Ste08, TGH^{+}07, VCSG07, FIK^{+}20]. Inhibitions [Les08x, Sho09-50]. inhibitor [BYF^{+}05, BGZ^{+}07, KRP^{+}09, LK05, PTL^{+}05, RRA^{+}09, TCW^{+}06, YTB07, YLM^{+}06b]. inhibitors [PMS^{+}07, Rob08e]. inhibitory [KCOS06, PLR^{+}06, PPS^{+}06, TLK^{+}06, XMSYJ07]. inhibits [FWC^{+}07, HZL^{+}07, JGY09, KKM^{+}08, KNKH06, KII^{+}08, KWY^{+}06, LMV^{+}09, LPWP^{+}07, LZL^{+}05, MAH^{+}05, MDET08, NSM^{+}09, OKN06, RBJL06, RMC^{+}05, RGA^{+}06, SQB^{+}08, TCW^{+}06, Wel05-63]. initial [DZG^{+}07, HXO^{+}07, HOR08, WNB08, YOHR07]. initiate [CWNF^{+}09, RCS^{+}09, TMI07, YWJ^{+}07]. initiated [DBW^{+}06, KKB^{+}06]. initiates [BFL^{+}05, GGBV^{+}07, KSR^{+}08, LLM^{+}05, ROKG08]. initiating [TJC06]. initiation [KII^{+}08, MGBH^{+}06, NN07a, NFC^{+}08, oSCW07, YN07]. injured [ITG^{+}08, NJS^{+}06]. injury [HFY^{+}06, LeB05e]. INK4A [LTDH05]. Inn1 [NSO^{+}09]. Innate [Wel06v]. inner [BZA^{+}06, BSDA07, BSM^{+}08, BSM^{+}09, COB^{+}08, CNMM^{+}06, DDRL^{+}09, DJV^{+}05, GFC^{+}08, HWPH^{+}08, HIK^{+}08, HHS^{+}09, HFF^{+}07a, HCL^{+}07, LS06, MNH05, Sed09h, VBNR06]. innervation [Pow05c, SHX^{+}07]. inositol [HSLI09, HPH^{+}05, LMC^{+}09, MuMI^{+}06, SFB^{+}05, dCD^{+}07]. Inp1p [FFT^{+}05]. ins [Les08b]. insensitive [BKBM06]. insertion [HSTM06, LeB06s, LeB06-82, OTH^{+}07, Rap05]. insight [BBM^{+}07, LeB06-51]. insights [LDKT06, PFGSV06]. InsP [MHM^{+}05]. instability [LYWK08, MGW^{+}08, MSC^{+}07, MSM^{+}06, RP08, TT09, TC08]. instincts [LeB06-34]. instructing [Wel07f]. instructs [Pow05c]. insulative [DG08]. Insulin [LeB06-32, LMZ^{+}05, GMK^{+}05, HGJ^{+}08, OIFN^{+}07, YCR06]. insulin-signaling [YCR06]. insulted [HLJ^{+}07]. intact [SM09, VTRB06]. integral [MMCW06, SHC07, SHS^{+}06]. integrates [HLA^{+}06]. Integration [BCR^{+}08, TTYH09]. integrator [HSK^{+}05]. Integrin [CWNF^{+}09, FDM^{+}08, GLBS06, KKW^{+}09, Les06o, LGT^{+}07, PBB^{+}09, WCB^{+}08, dRKD^{+}05, AXM^{+}05, AS06, ASHD^{+}05, BMR^{+}08, BCT06, BMRP07, CNM^{+}05, CCL^{+}08, CSL^{+}05, CML^{+}06, DRD^{+}05, EVH^{+}07, EVH^{+}08, EBMG09, FHNG05, FMM^{+}08, FZT^{+}07, HHY^{+}08, JS07, KGM^{+}09, KWY^{+}06, LeB05g, LeB06-34, LLC^{+}08, MQL^{+}06, MFBG^{+}08, NLC^{+}05, NZQ^{+}07, OKN06, PFGSV06, RBJL06, SMH^{+}09, Sho09w, SHjC^{+}07, SZX08, WBP^{+}08, WCR^{+}05, WCN07, WPL^{+}06, XMA^{+}09, YMD^{+}07, YLM^{+}06b, ZB09, ZKR^{+}07, TTG^{+}09]. integrin-associated [MQL^{+}06]. Integrin-dependent [dRKD^{+}05]. Integrin-linked [FDM^{+}08, LGT^{+}07, PBB^{+}09, WCB^{+}08]. Integrin-mediated [CWNF^{+}09, GLBS06, AS06, KWY^{+}06, YMD^{+}07]. Integrins [DvRF^{+}05b, LeB06-33, OLS08, Tum05-34, EVH^{+}07, EVH^{+}08, JLG^{+}06, LSIT^{+}08, MQWP08, NNS^{+}08, PAV^{+}06, SFC^{+}08, SDN^{+}05b, Tum06-37]. integrity [CAR^{+}08, KFR^{+}08, KSE^{+}07, MNF^{+}07, NVEF07, SDW^{+}09, THY^{+}06, TKSNG07]. inter [BSM^{+}09]. inter-doublet [BSM^{+}09]. interact [BMG^{+}07, MMH09, Pow05a, ZWG^{+}05]. interacting [LCL^{+}07, MGR^{+}05, MM05, OVL^{+}07, HTK^{+}08, TWC^{+}07]. Interaction [SSP^{+}07, AGL^{+}08, ADC^{+}07, BH08b, BSH^{+}07, CNM^{+}05, COTAE^{+}05, DNG^{+}06, EMD^{+}08, FVSC08, FIK^{+}05, FIK^{+}20, GMW^{+}05, qGyWX^{+}08, GNPC^{+}08, JCMG08, KIKI^{+}07, LVL^{+}05, LSdWL07, LBC^{+}07, MYMH09, NPHT08, NHK^{+}08, OSM^{+}09, RKK09, SKE^{+}05, SYRL^{+}08, STBL09, ShCM^{+}06, SHK^{+}07, TMBL^{+}08, WPH^{+}07, ZTZ^{+}09, ZRH^{+}09]. Interactions [KSpLL06, SBMP05, ASK^{+}09, BRKR06, BXHD08, BFB^{+}08, CWNF^{+}09, FFW^{+}07, FKB^{+}05, GL06, HWS^{+}07, KCAS05, LGH^{+}05, MEG^{+}07, MSB06, NSO^{+}09, OSI^{+}07, PG06, QHB^{+}07, RKN^{+}07, TMH^{+}06, Tum05-69, UPBM06, XOV^{+}09, YNL^{+}08, vRCB^{+}07]. interactome [COB^{+}08, OHP^{+}09]. interacts [FTM^{+}06, GD05, HSB^{+}07, JMG^{+}06, KTI^{+}05, LMM^{+}09, MPG^{+}09, RRP^{+}08, STU^{+}08]. intercalated [PDM^{+}08]. intercalations [YKP^{+}08]. intercellular [BZHA^{+}08, BW08, LSIT^{+}08]. interdomain [LYK^{+}06, Tho06]. interface [CVD^{+}05, MFM^{+}06]. interference [TCKV05]. interferes [DDLC^{+}06]. interferon [KBC^{+}07b, LHRP05]. interferon- [KBC^{+}07b, LHRP05]. interkinesis [MHL^{+}06]. interkinetochore [OLH^{+}08, WF06]. Interleukins [Tum05-35]. Intermediate [ÇB09, Bon05g, CPAH08, CSTT^{+}06, LC09, WAW08]. intermembrane [HK07, SDI^{+}06, SPK^{+}09]. internal [NLS^{+}05]. Internalization [SB06, BGZ^{+}07, HBK^{+}06, LOG^{+}06, SCK^{+}07, WMT07]. internalized [DBW^{+}06, KBC^{+}08, Les05a]. International [Wel06o]. Interneurite [TMH^{+}06]. interneurons [DCAC05]. internode [MEG^{+}07]. interphase [CM09, HTYNK09, KYK^{+}06, LeB06-40, MLS^{+}07, MSC^{+}07, PL06, Rex09, VDSV06]. Interplay [SAD^{+}09, BAdSG05]. intersects [KDD^{+}09]. intestinal [BAR^{+}05, BDP^{+}07, JLG^{+}06]. intimate [BJC^{+}06]. intra [SSP^{+}07, WOP^{+}06, MTV^{+}05]. intra- [WOP^{+}06]. intra-Golgi [SSP^{+}07]. intrabodies [PCG^{+}05]. Intracellular [CEY^{+}05, DPMB^{+}07, FM05, LOG^{+}06, LYY08, LGV08, MuMI^{+}06, NRW^{+}08, RP05, WZM^{+}08, WBK^{+}06, WP08, YSSH^{+}07]. Intraflagellar [Den05, ELM09, Sch08, AGL^{+}08, IEGV06, PGL^{+}09]. intrafusal [AWE^{+}05]. Intrakinetochore [MS09]. Intramembrane [OHH^{+}06]. intramolecular [VBSM06]. intranuclear [DOS^{+}07]. Intranucleolar [KZS07]. intravasation [WMS^{+}06]. intrinsic [YTY^{+}07]. Introduction [Wel06-67]. intron [ELD^{+}08]. INTS3 [SRS^{+}09]. invade [INB06, Rob08-34, Sho09v]. invadopodia [CCH09, CCP^{+}08]. invadopodium [OYM^{+}09, YLK^{+}05]. invasion [BHJ^{+}05, CCH09, CCP^{+}08, GdTAR07, LD07, LeB07-34, OHH^{+}06, RLH^{+}09, SSHW09, SYRL^{+}08, Sed07g, WMS^{+}06, Wil07-33]. invasive [GCSG^{+}05, LMMHP^{+}05, Sho09q]. invasiveness [YA06]. invention [Wel05-31]. inverse [MPS^{+}07]. inversely [ELM09]. involucrin [SNG^{+}07]. involved [AFV^{+}05, CJMG05, DUR^{+}05, ELD^{+}08, HMR^{+}07, HDR09, LSD^{+}06, LDL^{+}09a, MLP06, OMGW^{+}06, SRA^{+}05, SCMV07, SHL^{+}09, SBC^{+}07, TSK^{+}06, TA07, TSS^{+}06, VW08, VVMS05, XMSYJ07, iYEF^{+}07, ZWG^{+}05, vBAS^{+}07, vdBKB^{+}07]. Involvement [CULS^{+}07, HOH^{+}08]. involves [AH05, ABSK06, EHWL06, HGK^{+}06, KMSLS06, MLHS06, WvdLH^{+}07, XJF05]. involving [KIKI^{+}07, TSN^{+}06]. ionizing [LYWK08]. IP [FLA^{+}08, SFEK07, TDT^{+}08, Wil08l]. Ipl1 [YK07]. Ipl1p [ZSMW09]. IQGAP1 [BMG^{+}07, SYRL^{+}08]. Ira [Wel07h]. Ire1 [HLL^{+}09, KIKI^{+}07, Wil07-29]. Ire1-dependent [HLL^{+}09]. Irina [Pow09a]. irradiation [KDD08]. Irreproducible [RVH08]. irreversible [BvOM^{+}07]. IRS [LFL^{+}05, NEF^{+}06]. IRS-1 [LFL^{+}05, NEF^{+}06]. IRSp53 [MPS^{+}07, RKK09, SKO^{+}06]. Irving [LeB09b]. isn’t [Wil09a]. isoenzymes [ERM^{+}09a]. Isoform [GGH^{+}08, DGP^{+}05, EYB^{+}06, FFW^{+}07, HLM^{+}08, HK06, HLP06, MCD^{+}07, RRP^{+}08, WAW08]. Isoform- [GGH^{+}08]. isoforms [DDRL^{+}09, HlQY^{+}06, KFR^{+}08, ZTM^{+}08]. Isolating [Les05b]. Isolation [ABA^{+}06, Les05c]. isomerase [FSnP^{+}06]. isomerase-like [FSnP^{+}06]. ISWI [YRSM^{+}09]. ITAM [Rob08a, ZKR^{+}07]. itself [Les08y]. IV [YOHR07, AOT06].

Jagunal [LC07]. JAK [WLC08]. JAK/STAT [WLC08]. JAK1 [SMW^{+}07]. JAM [LLC^{+}08]. JAM-L-mediated [LLC^{+}08]. JCB [Hil07, Hil08, Mel08b, Mis09b]. Jennifer [Wil08m]. Jernej [Wil08n]. Jidong [Wil07-31]. JIP1 [SL05]. JLP [TKB^{+}06]. JNK [SL05, SLZ^{+}06, STMG05, ZTKK05]. JNK1 [TÖL^{+}06b]. job [Mel08a, Tum05d, Wil08e]. Jodi [Sed09e]. John [Wil08u]. Join [Rob08-28]. joining [DG05, WC07b]. joint [Bas07a, PZG^{+}07, SOC^{+}07]. Joseph [Ros08c]. journeyman [LeB07t]. Journeys [EGL^{+}09]. Jump [Hac07, Les07-76]. Jump-starting [Hac07]. Jun [PBAF^{+}08, KNKH06, BWJ05, qGyWX^{+}08, GBNS07, MM05]. Jun/AP1 [GBNS07]. JunB [LPF^{+}06]. Junction [Wil08o, BZHA^{+}08, CM06, CULS^{+}07, Gra07, LBC^{+}07, LJY^{+}07, LCL^{+}05, MDL^{+}09, MPG^{+}09, MRO^{+}05, NVJ^{+}08, PYR08, PFD06, SWT08, SSM05, Tak06]. Junctional [HPH^{+}05, OIAT06, Wel05l, vZPG^{+}07]. Junctions [Tum05-36, ABY^{+}08, Bon05c, DG08, DP08, GHSG07, HT08, IUT^{+}07, JSC08, LeB07-39, LeB07-57, Les08-56, LSIT^{+}08, LWBL06, MMH09, MMT^{+}05, Pow05t, RSS^{+}09, SHS^{+}09, WCB^{+}08, WL09a, WL09b, Wel05k, Wel06-68]. Junying [LeB07z]. Jussi [Sed08i]. Just [Tum05-37, DDH09, Les07-72, Les07-63, Les08q, Tum06-45]. Just-right [Tum05-37].

K123 [FD09]. K4 [FD09, GSY^{+}09]. K63 [LJA09, OLC^{+}07]. K63-linked [LJA09, OLC^{+}07]. K79 [FD09]. KA1 [CYY^{+}08]. Kai [LeB05-28, Sed08j]. kainate [CYY^{+}08]. Kaksonen [Sed09h]. Kank [KRZ^{+}08, RKK09]. Kapoor [Sed09m]. Kar3 [GHM^{+}08]. Kar’ted [LeB06-35]. Karyopherins [Les06p, MPL^{+}07b]. karyosome [LCO07, LeB07i]. KASH [RH09]. Katanin [LeB07-27, MAOM06, SBW^{+}07, LDGO^{+}09]. KCA [yYMM05]. KCA-1 [yYMM05]. KChIP1 [HFP^{+}05]. kD [KFK^{+}05]. KDEL [RAS^{+}07, RAS^{+}08b]. keep [Kat06g, LeB06-77, Wel05c, Wil07-38]. Keeping [Les07-35, Wel05-33, Wel06w, Sed09e]. Keeps [Rob08-45, LeB06-64, Rob08-40, Sho09l]. Kept [Wel05-34]. Keratin [NTS^{+}05, KO06, LC09, OMGW^{+}06, SIS^{+}08, WKWL06, WWL07]. keratin-binding [SIS^{+}08]. keratinocyte [GFH^{+}06]. keratinocytes [MRO^{+}05, NMV^{+}07]. Keratins [KC09, VKR^{+}09, BLC05, LeB06-67, SBC^{+}06]. Kex2 [HEW07]. key [ASP06, JD06, LLL^{+}06, MFA^{+}07, NLC^{+}05, OHH^{+}06, RSL06, SWG^{+}05, SMW^{+}07, VKGV08]. keys [WC07b]. kick [LeB06i, RW08b]. kickstarts [Les08c]. kid [Wil07-32, TTO^{+}08]. Kidney [SKvdBW^{+}06, HMM^{+}05, JSFP08, LeB08d]. kidneys [Wil09s]. KIF13A [DHT^{+}09]. KIF14 [GNL^{+}06]. Kif2a [JWC^{+}08]. Kiger [Sed08b]. kill [LeB06a, LeB06-70]. Killer [Pow06g, Les07e, RHMS05, Rob07m, TLK^{+}06]. Killers [Wel06x]. killing [WSS^{+}06, Les09s]. Kim [Wil08v]. kinase [BYF^{+}05, BSA^{+}08, BPL^{+}05, BFP^{+}07, BFB^{+}08, BPL^{+}08, CKG^{+}08, CNM^{+}05, CHCB05, CM06, CFBM05, CBIO05, ELJ^{+}08, EYB^{+}06, FBO^{+}07, FDM^{+}08, FHS^{+}07b, GXJF07, GHM^{+}09, GWS09, GGEd^{+}07, GNL^{+}06, HHT^{+}07, HLL^{+}08, HF06, KNKH06, KLV^{+}08, KNT^{+}07, LU05, LPA^{+}07, LGH^{+}05, LeB05-44, LeB08a, LPN^{+}08, LBC^{+}07, LCL^{+}07, LGT^{+}07, LFL^{+}05, MCF^{+}05, MLC^{+}07, MGBH^{+}06, MPG^{+}09, MBF^{+}07, MBF^{+}08, MM05, NYM^{+}08, NPB^{+}05, NSM^{+}09, NTY^{+}05, NYN^{+}06, OSS^{+}07, OPT^{+}07, PPK08, PBB^{+}09, PWW^{+}09, ROKG08, RG06a, RMAA08, SKE^{+}05, SRN^{+}07, SPA^{+}05, SDD^{+}07, STP08, SHDT05, SWH^{+}06, SLWA07, TWL07, TST08, TSP^{+}07, VKR^{+}09, WCB^{+}08, WHB^{+}07, WC07b, WCN07, WGKA^{+}08, WF05, WSB^{+}09, XOV^{+}09, YHC^{+}06, YK07, ZPGB^{+}07, dGHW08, Wil08w, KGKV07, ST09, vKGK07]. kinase-1 [SWH^{+}06]. kinase-3 [BYF^{+}05]. kinase-deficient [EYB^{+}06]. kinase-MAPK [GXJF07]. kinase-mediated [YHC^{+}06]. kinases [CTY^{+}06, CLK^{+}07, DCA^{+}06, DSBK09, FOB^{+}09, LVL^{+}05, Les07-33, LRS05, OMGW^{+}06, PLH^{+}08, YMC05]. KIND [HFF07b]. Kindlin [MQWP08].

Kindlin-2 [MQWP08]. Kinesin [CYC^{+}06, DBD^{+}09, LeB06-36, Tum05-38, CBMCSS08, GMSS06, Hac07, HHFC06, KKW^{+}08, LeB05u, LeB07-41, LeB08r, MTB^{+}09, MGBM^{+}09, MSK05, RG06b, SD07, Tom08, TS06, WDWC08, ZHYE08, BCJ^{+}07, CHSV07, LeB07-28, MB07, Rob08s, TAM^{+}06, VSG09, yYMM05, POCS^{+}06]. Kinesin-1 [BCJ^{+}07, CHSV07, yYMM05]. kinesin-13 [WDWC08]. Kinesin-13s [TAM^{+}06]. kinesin-3 [HHFC06]. kinesin-5 [CBMCSS08, KKW^{+}08, LeB07-28, MB07]. Kinesin-6 [VSG09]. kinesins [ESG^{+}06c]. kinetic [CLA^{+}05, DZW^{+}08, TTNA^{+}06]. kinetically [HCC^{+}07]. kinetics [CFH^{+}08]. Kinetochore [TWG^{+}09, UTK^{+}09, Wel06y, ASH^{+}09, CYC^{+}06, CFU^{+}09, ELJ^{+}08, FVSC08, GTHK05, GSV07, HWPH^{+}08, HFF^{+}07a, HHZ^{+}08, JBF^{+}08, JhKC08, KHWC08, KKM09a, KCH^{+}06, KKW^{+}05, LRJY06, MS09, MHFH06, PMK^{+}07, Pow06h, Pow06l, Sed07e, TKKT07, VDSV06, VMSV08, Wel06-36, Wil06b, WF06, WWvD07, YWEK08]. kinetochore-associated [CFU^{+}09]. Kinetochore-generated [TWG^{+}09]. kinetochore-independent [KKM09a]. kinetochore-microtubule [JBF^{+}08, VMSV08]. kinetochore/cell [VDSV06]. Kinetochores [Les07-36, Wel05-35, MPLF^{+}09, MDPG07, NNK^{+}08, OLKK09, SCVW09, Sho09-28, SCMV07, TST08, TS06, WBT^{+}08, WF05]. king [Wil08w]. Kip1p [CGJ09]. KIR [KCOS06]. kit [Les08j, ZDD^{+}06]. KLHL21 [MOS^{+}09]. Klotho [MRD^{+}08]. KLP10A [MTB^{+}09]. knock [Rob08e]. knock-out [Rob08e]. knockdown [TKH^{+}05, WKIKG09]. Knocking [Tum06w]. knockout [OFfL^{+}07, SPA^{+}05, YCX^{+}08]. Knows [Wil08v]. Konrad [Wil07-32]. Krantz [Sav09a]. KRas [FM05, LTAM^{+}09, Sho09-30]. Kris [Wil08p]. KRIT [GHSG07]. KRIT-1 [GHSG07]. KRIT-1/ [GHSG07]. Ku80 [PGW^{+}08]. Kv4 [HFP^{+}05].

L [CZD^{+}09, HSK^{+}05, LLC^{+}08, MKV^{+}09, WBG^{+}08]. L-type [CZD^{+}09, HSK^{+}05, WBG^{+}08]. L1 [GCSG^{+}05, LLS^{+}09, STU^{+}08, YWK^{+}08]. L1-CAM [STU^{+}08]. L1/NgCAM [YWK^{+}08]. L1CAM [WZC^{+}08]. L7Ae [BMGPB^{+}08]. labor [Les06g]. Lack [DMBF^{+}07]. lacking [ECR^{+}09, HJB07, MYO^{+}05, SKvdBW^{+}06, WLLF^{+}08]. lacks [Tum06c]. lacritin [MBR^{+}06, WWX^{+}06]. lactational [AS06]. LAD [WZC^{+}08]. LAD-2 [WZC^{+}08]. Ladd [Wil08b]. laevis [KHH08, LM05, MFH05, OW06, UHM^{+}08]. laforin [GDW^{+}07]. lamella [KLG^{+}09, NAS08]. Lamellae [Tum05-39, NZQ^{+}07]. Lamellipodia [LeB08l]. lamellipodial [BHK^{+}06, OPT^{+}07, PJMR06]. lamellipodium [GAK^{+}05, KOKM07]. Lamin [MLV09, MDL^{+}09, Rob08-45, Sed09g, Sed09f, GNPC^{+}08, MLS^{+}07, SSS^{+}08b, ZS08, PHB^{+}07]. Lamina [DVF^{+}06, Les05c, Pow05c]. Lamina-associated [DVF^{+}06]. Laminin [Les08-29, LLM^{+}05, CYY^{+}08, HOM^{+}05, MBL^{+}07, Tum05-29, YBT^{+}05]. laminin- [MBL^{+}07]. laminin-2 [YBT^{+}05]. laminin-5 [HOM^{+}05]. Laminins [NVJ^{+}08, MBL^{+}07]. laminopathies [MG08b]. Lamins [LeB08m, Tum06x, Wel06z, AG09]. lands [Wil07y]. landscape [MG09]. Langerhans [Wil07-33, BvRJ^{+}05]. LANP [HAT^{+}05b]. LAP1 [GD05]. LAP2 [PHB^{+}07]. LAPTM5 [PGB^{+}06]. LARG [HKIY09]. Large [DR05, HKP^{+}09, BYMB06, BEN^{+}05, CHA05]. Large-scale [DR05, HKP^{+}09, BEN^{+}05]. largely [PCV^{+}08]. larvae [ZWG^{+}05]. last [Tum05-76]. late [ASH06, CZP^{+}05, FN06, GMG^{+}07, HSB^{+}07, KOVP09, LTAM^{+}09, MTRJA05, MYMH09, NVC05, RKvdK^{+}09, SPA^{+}05, SXH^{+}06, Wel05-27, Wil07s]. latent [BZA^{+}06, GG06, WRMH07]. Lateral [MSR09, BMK^{+}06, KR07, MSH^{+}05, Pow06c, SRL05]. Lattice [LeB05-47, Pow05r, SWL^{+}05, WKCY^{+}08]. Lattices [LGDN09]. lay [Sed08c]. layer [Kat07b]. lead [CGK07, Sho09-31, Tum05-39]. leaders [Kat06c]. leading [BWW^{+}08, qGyWX^{+}08, GC05, HSB^{+}07, LeB08-36, RMR^{+}09, SGMPV07, THN^{+}06]. leads [FAG^{+}07, GHDS^{+}07, LTAM^{+}09, MOCC05, PS08, Pow05m, RSM^{+}08, SGDM06, WCR^{+}08]. leaks [Sed08d]. leaky [Wel05-30]. lean [Les07-37]. Learn [Rob08t]. learning [HZL^{+}07, LeB05-48, Sho09c]. leave [LDVV06, LeB07i, Les08-46]. Lechler [LeB07-61]. legend [AKD06]. length [AHD^{+}09, BH08b, ÇB09, DTY^{+}05b, ELM09, MAOM06, TWL07, TÖL^{+}06b, ZHYE08]. Lengthening [Wel05-36, JZN^{+}09]. lengths [BLL^{+}06, MSV^{+}05]. lens [NFZ^{+}09, Sho09m]. Leptin [NBM^{+}08, HGJ^{+}08, Les08-30]. Less [Tum05-40, Hat05a, LeB05-50, Wel07g]. let [Les07l, Les09b, Les09-37]. Lethal [ZWG^{+}05]. lethality [JLG^{+}06, WRKC^{+}09, WBL^{+}06]. letting [Wel05-36]. leucine [NPB^{+}05]. leucoprotease [TCW^{+}06]. leukemia [BRY^{+}09, DCA^{+}06, QIH^{+}06]. leukocyte [DDLC^{+}06, LLC^{+}08, vBAS^{+}07]. leukocytes [BZYM^{+}08, Les07-57, LZ08, Rob08r]. level [MJG07]. levels [CLS^{+}06, FDB^{+}09, LDGO^{+}09, PVM^{+}05, RCM^{+}09, RRR^{+}09, SPS^{+}08, VTRB06, WvNS^{+}07, YTB07]. LFA [FHNG05, SCS^{+}05]. LFA-1 [FHNG05, SCS^{+}05]. lgl [GAG^{+}06, BFP^{+}07]. Liberating [Les07-38]. license [Tum06-55, WGL^{+}06]. licensing [LLC^{+}07]. life [KHB^{+}05, Les06l, Les08-34, SWM^{+}06, Tak06, Tum05-40]. lifespan [LeB07-46]. ligand [GHN^{+}09, KGM^{+}09, LNG^{+}05, LeB05-88, LCvD^{+}08, NMO^{+}07, WPC^{+}08]. ligands [GSC^{+}07, LNG^{+}05, UDC^{+}07]. ligase [CNW^{+}05, GGH^{+}08, GBNS07, KNY07, LJY^{+}07, MOS^{+}09, MKC^{+}05, ORM^{+}08, PGB^{+}06, RRR^{+}09, STvdB^{+}09, TTE^{+}07, WHC^{+}07]. ligases [EYB09, MSECB05]. Light [COTAE^{+}05, FDB^{+}09, GMSS06, LB08, PKK09]. Light-regulated [COTAE^{+}05]. like [BFF^{+}06, CTS^{+}06b, DGZ^{+}05, FSnP^{+}06, GKvdB07, HWN^{+}07, IUT^{+}07, MLC^{+}06, MYO^{+}05, MPS^{+}07, MEG^{+}07, MGH^{+}09, MDW05, PCR^{+}07, Pow05y, RMAA08, RLDC09, RKB^{+}07, SRR^{+}08, TLK^{+}06, Tum05-51, Wil09a, WVB07, dGHW08]. LIM [CM06, NTY^{+}05, PEW^{+}05, WHB^{+}07, WHM^{+}07, ZTZ^{+}09]. LIM-only [WHM^{+}07]. limb [MLHS06, PZG^{+}07]. limbal [BTD^{+}07]. limbs [Wil07-38]. limit [WL09a, Wil07-33]. limited [BWJ05, PLR^{+}06]. Limiting [Kat07b, Lec08, GHM^{+}09, RM05a, SYKM08]. limits [AH08, MJF^{+}07]. Lin12 [TRS^{+}08]. Lin12-Notch [TRS^{+}08]. LINC [CLR^{+}06]. Lindau [SGB^{+}06]. Lindenbach [Wil08f]. line [ERS08, Pow09d, WBL^{+}06, Wel05g, YCX^{+}08]. lineage [GNE^{+}07, PVMH05]. lineages [WPDZ^{+}06]. linear [HKP^{+}09, KBC^{+}07b]. Lining [Les06q]. link [HRK^{+}09, MSECB05, PLR^{+}06, Rob07d, TC08, Wel05-45, Wel06-39]. linkage [CRV^{+}06, KYK^{+}06, SHS^{+}09, WAW08]. linkages [KKGD05]. linked [CTY^{+}06, CC09, FDM^{+}08, KSA^{+}05, KBC^{+}07a, LC09, LGT^{+}07, PL06, PBB^{+}09, SGK^{+}07, WCB^{+}08, LJA09, OLC^{+}07]. Linker [MBL^{+}07, FSTI06, GTB^{+}07, WLK^{+}05]. linkers [CBMCSS08]. Linking [EGR^{+}05, LeB07-29, MBS^{+}08, Tum05-41, Tum05-42, Tum06y, Tum06z, GHX^{+}07, HFF07b, KKW^{+}08, LFH07, NN07a]. Links [LeB05-49, AS06, AOP^{+}05, BSM^{+}09, CKKG05, GDW^{+}07, HH06a, KKW^{+}09, KKW^{+}05, LeB05q, LeB05-58, LTDH05, MQL^{+}06, MBF^{+}07, MBF^{+}08, RVBZ^{+}08, SRA^{+}05, SHC07, SMT^{+}08, WSB^{+}09]. Lipid [Les09t, Pow06i, Rob07g, Rob08u, SCR^{+}09, Sho09-31, Sho09-32, dGSWW^{+}06, Wel05-37, Wel06-27, BJC^{+}06, BCR^{+}08, BXHD08, BDW^{+}09, DBW^{+}06, EYB09, FKB^{+}05, FSG^{+}08, LJR^{+}05, LeB05-79, Les08-53, Lev07, Nic09a, SRL05, SSLS05, SGK^{+}07, SDD^{+}07, VTRB06, Wel06-44]. lipid-anchored [Nic09a]. Lipid-driven [Wel05-37]. lipid-linked [SGK^{+}07]. Lipid-regulated [SCR^{+}09]. Lipids [LeB06-37, HSLI09]. lipin [Les08g]. lipoxygenase [EFM^{+}07]. liprin [OMF^{+}05]. LIS [Hat05a, LeB05-50]. LIS-less [Hat05a, LeB05-50]. LIS1 [TCKV05, HMG^{+}08]. Listeria [BKBM06]. Lithium [SFB^{+}05]. little [Ras08a, Rob08-41, Tum05-43]. Liu [Wil07-31]. Live [SWG^{+}05, Wil06h, ALB^{+}09, BMB^{+}07, CHSV07, HSTM06, KG05, LeB05-41, Wil07-57, lWHMH06b]. lived [SS08a]. liver [GPN^{+}06, Wel06-28]. living [BMK^{+}06, CCC^{+}05, CSL^{+}05, DZW^{+}08, HTYNK09, HWPH^{+}08, KCD^{+}06b, KS08, LGH^{+}05, LDPLS05, LJS^{+}09b, MuMI^{+}06, RKT^{+}07, RSSC06, VTRB06, vRCB^{+}07]. LKB1 [FA08, JLC08, MSK^{+}07]. LL5 [KKES05]. lll [JRZ^{+}07]. Lloyd [Sed08k]. LMP4 [CBK^{+}06]. loaded [CS07, Ski05]. loading [TTO^{+}08]. loads [BMK^{+}06]. Local [KO07, LeB05-51, CAR^{+}08, DGP^{+}05, HSK^{+}05, Les08z, LWBL06, WZC^{+}05, Wel05h, WZCB^{+}06]. Localization [BLA^{+}09, CK07, KSL^{+}05, ASP06, BFWH07, CBK^{+}06, CKX^{+}08, CSW08, CGJ09, FKML^{+}06, GSC^{+}07, GSdBS09, JGAS08, KZS07, LSDFS08, LZM^{+}06, Les06f, LWHC08, LYY08, LSdWL07, LTK09, LCL^{+}05, MMC08, MDPG07, MHFH06, PYR08, PGO^{+}07, PEW^{+}05, SRBL06, SOR06, Tum06r, VKR^{+}09, WPH^{+}07, XOV^{+}09, YR09, YK07, YPG05, ZPGB^{+}07]. localize [WVB07]. Localized [LeB05-52, SSD06, Sed07d, YN07, EMMN^{+}05, JASC08, KL05, RRP^{+}08, SGMPV07, SWI06, WvNS^{+}07, FHFB^{+}06]. localizes [COC^{+}05, FDM^{+}08, GRGP07, NWB^{+}06]. localizing [PWW^{+}09]. Locally [XMSYJ07, Ter06, WBG^{+}08, vRSvR^{+}07]. location [DDH09]. lock [WC07b]. locks [Sed09g]. locomotion [Bon05a]. Locus [MM08, DOS^{+}07, HGJ06, KS08]. Locus-specific [MM08]. loitering [Les07-58]. Long [MSK05, BF09, LeB09c, LLL^{+}06, Les06x, Les07s, Les07x, Les07-39, Les08w, Les08-30, RCS^{+}09, Rob08-33, Rob08-50, Sho09-33, SCFKB07, Tum05-43, Wil06h]. long-distance [Les08-30]. Long-range [MSK05]. long-term [LLL^{+}06, RCS^{+}09, SCFKB07]. longer [Pow06a]. longest [LeB08-33]. Longevity [MGW^{+}08]. look [Sed07e]. looks [Pow05y]. loop [ANI^{+}07, HMR^{+}07, MOCC05, NWR^{+}08, RYD08, Rob08z, Wil07-49, XKYW07, YWW^{+}08, YMDL09, YWRA07, Les07-69]. looping [BLvH^{+}05]. loops [LeB06j, WEL^{+}07n]. Loopy [LeB05-53]. loose [Rob08-33, Wil07p]. loosen [Les08-56]. loosens [Sho09z]. lopsided [Les09k]. Lose [Les08x, Les08d, Les08-65]. Losing [LeB08n, LeB07z]. Loss [DSN^{+}07b, FSTI06, GdTAR07, LFW^{+}09a, LFW^{+}09b, MZH^{+}08, RSM^{+}08, DMBF^{+}07, JLG^{+}06, LeB06-44, Sed09a, Sho09-49, SCFKB07, Wel06u]. Lost [Les05d, RG06b, Wel06t, ZS08]. lot [DR07, Rob08h]. Louis [LeB09c]. love [BEH^{+}06, Sed08i, Sho09u]. low [MCA^{+}06, MS06, Wel06-58, XJF05]. low-mobility [MCA^{+}06]. loyalties [Les07-55]. LPA [RG06a]. LPA-stimulated [RG06a]. LRP5 [DXC^{+}08]. LRP5/6 [DXC^{+}08]. LSm1 [dPMF^{+}09]. LSm1/CBP80 [dPMF^{+}09]. LSm1/CBP80-mRNPs [dPMF^{+}09]. Lsm4p [DTP07]. Lte1 [GSdBS09]. Lucas [Sho09-34]. Lumen [LeB05-54, Pow07f, SMSPK08]. lumens [Lec08]. Luminal [GZBM^{+}06]. lunch [Wel06-30]. lures [LeB05z]. lush [Les07x]. lymphatic [NIC^{+}09b]. Lymphocyte [Wel06-29, GCF^{+}05, HPT^{+}07, HLK^{+}09, WSS^{+}06, dCD^{+}07]. lymphocytes [CLK^{+}07, LeB05-80, SCS^{+}05, SWB^{+}07, Wel06k]. Lyn [CLK^{+}07, LGH^{+}05, STP08, SFS^{+}07]. lysine [BvOM^{+}07, iHBD08, SS07, TMH^{+}07, WHC^{+}07, XGX^{+}09]. lysis [EW08, WSS^{+}06]. Lysophosphatidic [SB09, BPL^{+}05, PRZ^{+}08]. Lysosomal [ASSST06, CSB^{+}06, DNK^{+}05, VCC^{+}08]. lysosome [Bon05e, TAK^{+}09]. Lysosomes [Wel06-30, LTAM^{+}09, PGB^{+}06, Pow05j, Sho09i, Wel05f]. lysosomotropic [DDO^{+}08].

M [WBL^{+}06, ERM^{+}09a, HGJ06, CWC^{+}07, DSBK09, KHBV05]. MAB [WZC^{+}08]. MAB-20 [WZC^{+}08]. MAB-20/Sema2-mediated [WZC^{+}08]. machine [Wel05-57]. machinery [AGL^{+}08, BM07, BJLK^{+}06, BMGPB^{+}08, FKJ^{+}09, FTM^{+}06, GMK^{+}05, GGC05, Kak08, KAD^{+}08, SWT^{+}05, SHK^{+}07, TMH^{+}06, ZRBS07]. machines [EW08, Joh09]. macro [Kat07d]. macroapertures [BDR^{+}06]. Macroautophagy [YCK^{+}05, DBC^{+}06]. macrophage [DSLY^{+}05, Les08-27, MKH^{+}07, Pow05m, Wil07-34]. Macrophages [Wil07-35, DPMB^{+}07, JGL^{+}06, KDD^{+}09, Les07c, OTM^{+}07, OPT^{+}07, PPA^{+}08, vKST^{+}07]. Mad [MPG^{+}09]. MAD1 [MDPG07]. Mad1p [SCVW09]. MAD2 [MDPG07, NRD^{+}06, ORM^{+}08, SWS^{+}08, TST08, Yu06]. Mad2-dependent [ORM^{+}08]. made [LeB05-45, SWL^{+}05, Wil07-65]. MAGUK [CHH^{+}08]. mainly [MG08a, NMV^{+}07]. maintain [GSHD08, MSK^{+}07, PHB^{+}07, ZAM07a, ZAM07b, ZKG^{+}08]. Maintaining [LR07, NVEF07, PPBF08, THY^{+}06, WL09b]. maintains [CAR^{+}08, KS08, LU05, MNF^{+}07, MOM^{+}06, SBC^{+}05, SSS^{+}08a, YK07]. maintenance [AOP^{+}05, COTAE^{+}05, CULS^{+}07, DJV^{+}05, FMD^{+}09, HOR08, HWPH^{+}08, KMSLS06, KLW^{+}05, KWA^{+}05, KRG^{+}08, OSM^{+}09, SHC09, SHK^{+}07, SNPCG06, TKSNG07, XNM^{+}09]. major [AFN^{+}08, COB^{+}08, CBG^{+}06, CEY^{+}05, HRK^{+}09, KBS^{+}06, LGBD^{+}07, PTF^{+}06, Rob07h]. Make [Wel05-38, Bas07a, Bon05d, Bon05e, Hat05a, Les07-72, Les08z, Les08-44, Les08-64, Sho09x, Wel07g]. maker [Vau05]. makes [LeB06-28, Les07-51, Les07-79, Les09n, Mal06, Sho09p, Wel06n]. Making [Kat06h, Les06r, Les07-40, Les08-31, LNL^{+}08, Nyb05f, RMV06, SR07, Tum05-44, Tum06-27, Wel06-32, Wel06-31, Wel06-42, Pow09a]. Malaria [Wel06-33, CBG^{+}06, MBR^{+}05, OHH^{+}06, Tum06w]. male [EIK^{+}05, GGB^{+}08, PMB^{+}09, SRQL09, YDL^{+}06, YEL^{+}08]. males [KPS^{+}07]. Malhotra [Sav09b]. Malignancy [LeB06-38, CEY^{+}05]. Mallory [NTS^{+}05]. MALS [OMF^{+}05, OFfL^{+}07]. MALS-3 [OFfL^{+}07]. MALS-liprin- [OMF^{+}05]. Mammalian [BWFdL09, HRS^{+}05, IMY^{+}09, KDG^{+}09, BJLK^{+}06, BFP^{+}07, CSMP^{+}09, CLA^{+}05, DBC^{+}06, GT08, HTK^{+}08, HGA^{+}09, HLL^{+}09, JCVB08, KKK^{+}05, KMSLS06, KYK^{+}06, LLC^{+}07, MCSY07, MLP06, NSS^{+}08, OOY^{+}06, PVJJ07, QCGM05, RAS^{+}07, RAS^{+}08b, SBC^{+}06, SSM05, SHC09, TMW^{+}05, WH06, WLPB^{+}09, YOSM06, YSP^{+}06, ZEC^{+}08]. mammals [AS09, KKD^{+}07, SPS^{+}08, Wel07a]. mammary [AS06, GSY^{+}09, NLC^{+}05, PH07, SKR^{+}07, WMS^{+}06, XSVGR05, XNM^{+}09]. mammary-specific [XNM^{+}09]. Man [TMAL^{+}06, LeB06-62]. manager [Les09q]. Managing [Les09u]. manipulation [JDK^{+}05]. manner [HCK^{+}07, KKM09a, LCL^{+}05, SSD06, ZBKJ07]. Mannose [JP06, Tum06-28]. many [RH09]. map [LeB06-61, Sed09k, DSBK09, LeB06-39, OMGW^{+}06, Pow06j, YRSM^{+}09]. MAP2 [HFF07b]. MAPK [GXJF07, JTN^{+}05, KBY^{+}08, Las09, PCPW08, TKB^{+}06, VDA^{+}08, WWL07]. MAPK-dependent [WWL07]. mapmaker [Sho09-35]. Mapping [LRJY06, NKK^{+}06, Tum05-45, WP08, ELD^{+}08]. MAPs [Pow05a]. Marc [Wil09m]. MARCH [Les07-44]. MARCH5 [KNY07]. marches [Les06a]. Marian [Sho09-35]. Marie [TT09]. mark [Ros07b, dPMF^{+}09]. marker [DMC^{+}07, Vau05]. Marko [Sed09h]. Marks [LeB05-55]. marrow [FMM^{+}08, KR07, SDL^{+}05]. Mars [TLDB08]. maskin [KNP^{+}05, PSS^{+}05]. mass [CMN^{+}08, TMBL^{+}08, Wil08g, YAV^{+}05]. mast [NYI^{+}05b]. master [Sed08j, Vau05, Wil08g]. Mastocytosis [ZDD^{+}06]. Matching [LeB08o]. mate [Sho09l]. Maternal [NPB^{+}05, NAG^{+}08, BAR^{+}08, Rob08j]. math [Wil08x]. mating [HEW07, JMG08, MSW^{+}07, MSW^{+}17, PYR08, PKB^{+}06]. matrices [GdTAR07]. Matrix [LeB05-56, Les08-32, OSS^{+}07, BRM^{+}07, CEY^{+}05, EFCP07, FZT^{+}07, GS07, HXO^{+}07, LeB05-48, LeB08k, LJN^{+}05, Les07a, LFMO^{+}09, LR07, MFBG^{+}08, OSB^{+}05, Pow05g, RG06a, SAM^{+}07, TCHL05, WKW^{+}08, WRMH07].

Matrix-specific [OSS^{+}07]. Matt [Sho09-36]. matter [Tum06-44]. matters [Nyb05a]. Maturation [FOW^{+}09, AOT06, ECR^{+}09, FBO^{+}07, GKCM^{+}08, IBS^{+}05, KOVP09, LeB05-58, LJY^{+}07, MLC^{+}07, NVJ^{+}08, NIC^{+}09b, OYM^{+}09, SCM^{+}07, SHB^{+}07, SHL^{+}09, WNL^{+}09, WBL^{+}06, YHD^{+}09, ZWO^{+}09]. mature [Nyb05f, Wil07p, WSFD07]. maturing [PGWH08]. maturity [LeB07d]. Maurer [CBG^{+}06]. Max [WZO^{+}05, WZO^{+}05]. may [Rob07d, LeB06-52]. mBet3p [YSP^{+}06]. MCAK [HFF^{+}07a, MRvD^{+}05, WWvD07]. MCDE [LCT^{+}09]. MCL [BRY^{+}09, GEH^{+}07, LCvD^{+}08]. MCL- [BRY^{+}09]. Mcl-1 [GEH^{+}07, LCvD^{+}08]. Mcm2 [WGL^{+}06]. Mcp5 [SON06]. MD [lWHMH06b]. MDC1 [MBJM^{+}09, MBJF^{+}08, MDWH^{+}08, SMT^{+}08, BJLM^{+}05]. Mdc1/NFBD1 [BJLM^{+}05]. MDCK [API^{+}07, MSSF05, PPCZ06, VVMS05]. MDF [YWEK08]. MDF-1 [YWEK08]. mDia1 [LHL^{+}08, SOR06, ZHI^{+}08]. mDia2 [BMS^{+}08]. Mdm2 [LSG^{+}06, XKYW07]. Mdm30 [EHWL06]. Mdm31 [DJV^{+}05]. Mdm32 [DJV^{+}05]. Mdm38 [FTM^{+}06]. Mdv1p [KCAS05]. mdx [RKN^{+}07, SVS^{+}07a, SVS^{+}07b]. me [HK07, HW09, RVH07]. meal [Les08-64]. mean [Les07-37]. means [LeB09a, Les09p, Rob08-55, Wel05-48]. measurement [PJMR06]. measuring [Wel06-40]. Mechanical [YHJ^{+}05, AFM^{+}05]. Mechanism [DTL^{+}07, POCS^{+}06, BLC^{+}09, CC09, CLA^{+}05, DBW^{+}06, EMB^{+}08, HAT^{+}05b, KRP^{+}09, LeB06-39, MWY^{+}09, MPS^{+}07, NN07a, NVJ^{+}08, PhXD^{+}06, RKT^{+}07, SKO^{+}08, SRL^{+}06, SH05, SOR^{+}05, TM05, WZC^{+}05]. mechanism-based [RKT^{+}07]. Mechanisms [GNV05, MS08, WBP^{+}08, ASB^{+}05b, BW08, CSL^{+}05, DZG^{+}07, FHNG05, HL06, HLS^{+}09, JYZL06, LGBD^{+}07, MPD09, Mat05, PR05, TKKT07, VCSG07, WIL^{+}05, WFJ06, XMSYJ07, YLK^{+}05, YF08]. mechanistic [OIFN^{+}07]. mechano [FLA^{+}08]. mechano- [FLA^{+}08]. Mechanotransduction [CMF08, LHS^{+}05]. mediate [EVH^{+}07, EVH^{+}08, HKIY09, LTK09, MEG^{+}07, MDET08, OMLP^{+}08, PT07, Rap05, RRP^{+}08, TTO^{+}08]. mediated [AS06, CGS05, CMD^{+}08, CWNF^{+}09, CEY^{+}05, DMDO08, DPD^{+}09, DLT^{+}07, DCMB^{+}07, ERM^{+}09b, FVSC08, FHO^{+}05, FKB^{+}05, FBJB^{+}09, FCB^{+}09, FCB^{+}19, GLBS06, HBPL09, HOM^{+}05, HFP^{+}05, HGA^{+}09, IDTT^{+}09, KBK^{+}05, KSK^{+}06, KHW^{+}07, KHW^{+}09, KWY^{+}06, LDVV06, LGZ^{+}05, LPN^{+}08, LWC09, LWHC08, LFW^{+}09a, LFW^{+}09b, LGV08, LRS05, LLC^{+}08, MTRJA05, MHM^{+}05, MYZ^{+}09, MBR^{+}05, MDC^{+}08, MDL^{+}09, NYN^{+}06, OD07, OLC^{+}07, OFY^{+}05, PhXD^{+}06, RHMW^{+}05, RHMS05, RM05b, SON06, SMSPK08, SZL^{+}09, SDN^{+}05b, ShCM^{+}06, SL05, SLK^{+}08, SBV^{+}06, THN^{+}06, TGH^{+}07, TNS07, VBL^{+}05, WLE09, XKYW07, YCR06, YA06, YMD^{+}07, YHC^{+}06, ZL05, tKJCH06, ASB^{+}05b, AM05b, BLS^{+}05, BSL^{+}08, BCC^{+}09, CTS^{+}06a, DTR^{+}05, DBR09, EGR^{+}05, FHS^{+}07b, HLK^{+}09, KHH08, LHL^{+}08, LMC^{+}09, LLC^{+}07, ML05, NWO06b, NYI^{+}05b, NNS^{+}08, OKN06, SFC^{+}08, SRS^{+}09, SRE^{+}05, SDS^{+}05, STMG05, TSN^{+}07, WZC^{+}08, WKIKG09, WF07, WWS^{+}06]. mediated [XMSYJ07, YGY^{+}05, YLM06a, YAV^{+}05, ZTKK05, vRG08]. mediates [AVHH09a, AVHH09b, BLC05, CTGVGFF08, CM06, CG06, DDRL^{+}09, DFB^{+}07, DS07, EBMG09, GPJ^{+}06, GGE^{+}09, HHT^{+}07, HXZ^{+}07, HGJ^{+}08, JMB^{+}07, KTA^{+}06, KVS^{+}09, LLH^{+}07, LBG^{+}06, MMW^{+}07, MAG05, NSM05, NFI^{+}09, OHH^{+}06, OLLG05, PDM^{+}08, QIH^{+}06, SNG^{+}05, TDT^{+}08, WZC^{+}05, WLS^{+}09, WS09, WAW08, yYMM05]. mediating [FZT^{+}07, VKGV08]. mediation [NDS^{+}07]. mediator [VBG^{+}08]. mediators [LCT^{+}09]. medicine [LeB09b]. meet [Les09c]. Meeting [dSW09, LC08a, Mis09b]. meetings [LeB06-87]. MEF2C [LGG^{+}05]. megabytes [LeB08b]. Megator [LFMO^{+}09]. Megf10 [HLK^{+}07]. meiocytes [AJR^{+}09]. meiosis [DPP^{+}07, GGB^{+}08, JGY09, KKGD05, LeB05q, MHL^{+}06, MFC^{+}09, YKH^{+}08, YDL^{+}06, YEL^{+}08, YK07]. meiosis-specific [YKH^{+}08]. Meiotic [DSK^{+}06, Les08f, TSALS05, BGP^{+}06, CBIO05, FNH^{+}05, HH06a, KSL^{+}05, MBdR^{+}08, MAOM06, NVC05, NWR^{+}08, Pow07b, ROKG08, SON06, Sed07f, WH06, yYMM05, YCM^{+}08]. Meissner [Sho09c]. MEK [SL05]. MEK1 [TTK^{+}06]. Melanin [Les06s]. melanoblasts [MOM^{+}06]. melanocyte [MOM^{+}06, WM05]. melanocytes [FKML^{+}06, LTDH05]. melanogaster [AEJ^{+}06, DTY^{+}05b, HBK^{+}06, VCD^{+}06, BMK^{+}07, BBL^{+}08, CBIO05, DL08, GGB^{+}08, Gra07, HCN06, HCB^{+}08, JLC08, JKW^{+}09, KSR^{+}08, LMB^{+}06, MAZ^{+}08, MDH^{+}06, MKSJ06, NSS^{+}08, RGW^{+}07, SS07, SRL05, SB06, SORM05, SRQL09, TRS^{+}08, VLK^{+}08, WFJ06]. melanogenic [WRP^{+}06]. melanoma [BBG^{+}05, WM05]. Melanophilin [WTH05]. melanosome [DHT^{+}09, LCGRB09]. melatonin [BAdSG05]. melatonin-dependent [BAdSG05]. Melchior [Sed09b]. MELK [NPB^{+}05]. Mellman [Wel07h]. melting [Wil08x]. member [GAG^{+}06, HRS^{+}05, VWM^{+}07]. members [BLH^{+}09, DBD^{+}09, GGO^{+}09, RGCWG09, SWD^{+}08]. Membrane [ABSK06, BAR^{+}05, CYO^{+}09, EW08, FHMC09, HIK^{+}08, Kat07c, KGL07, Les06t, Les09v, SKO^{+}08, SPT^{+}09, STP08, ZOH^{+}08, AKD06, AWR^{+}07, AVHH09a, AVHH09b, AWM^{+}08, BSA^{+}08, BHY^{+}08, Ban09, BHG^{+}05, BZA^{+}06, BXHD08, BFF^{+}06, BBPM07, BVGA^{+}05, BSH^{+}07, CTS^{+}06a, CWC^{+}08, COB^{+}08, CNMM^{+}06, CSB^{+}06, DLHW09, DDRL^{+}09, DGZ^{+}05, DPMB^{+}07, DJV^{+}05, DDB^{+}05, DMC^{+}07, FG08, FFT^{+}05, FOB^{+}09, FML^{+}06, FSH08, GCF^{+}05, GLBS06, GOM^{+}09, GG09, GT08, GMS^{+}08, GFC^{+}08, GGBV^{+}07, HWSJ08, HRLH08, HLK^{+}09, Har08, HKV^{+}05, Heu05, HHS^{+}09, HZOL05, HSTM06, HPH^{+}05, HCL^{+}07, ITG^{+}08, IGFG^{+}08, IMY^{+}09, JKKM08, JMG08, JWSS05, KDV^{+}09, KKM^{+}08, KMS07, KGDR^{+}07, KKW^{+}07, KL05, LPG^{+}07, LGDN09, LeB07e, LeB07y, LeB07-43, LeB08i, LeB08v, Les08-52, Les09r, LLM^{+}05, LS06, LK05, LZ09, LWBL06, MMCW06, MF08, MT07, MNH05, MSR09, MFM^{+}06, NYI^{+}05b, NFCF08]. membrane [NFZ^{+}09, OLLG05, OTH^{+}07, Poo09, Pow07g, Pow07h, PKB^{+}06, Rap05, RIHP06, RM05a, RCFB08, SSHP^{+}07, SMH^{+}09, SHC07, Sed09l, STBL09, SNA^{+}07, SAM^{+}07, ST09, Ska07, SSZR07, SCF^{+}07, SHS^{+}06, SKO^{+}06, TDG^{+}06, TSS^{+}06, Tum06p, Tum06-41, UPBM06, VCN^{+}05, VBNR06, WHOS06, Wel05-31, Wel06-39, WLK^{+}05, Wil07d, WZCB^{+}06, WBLL06, iYOW^{+}06, YHD^{+}09, ZMGC05, ZMA^{+}07, vKST^{+}07, SPT^{+}21, Sed08j]. Membrane-anchored [HIK^{+}08]. membrane-bound [SAM^{+}07, SKO^{+}06]. membrane-integral [SHS^{+}06]. membrane-localized [KL05]. Membranes [CZK06, Sho09-37, AKD06, BFB^{+}08, CLA^{+}05, JTH09, LDPLS05, MPS^{+}07, SMWdV09, SCR^{+}09]. membrin [HAAHD05]. Memo [ZHI^{+}08]. Memo-RhoA-mDia1 [ZHI^{+}08]. memories [Les09g]. memorize [Wil07-36]. memory [HZL^{+}07, LeB08-30, Pow06f, Tum05-64]. men [Sed08k]. Mena [HJK^{+}06]. Mena/VASP [HJK^{+}06]. meningitidis [DDLC^{+}06]. menu [Les08y]. Merkel [Sho09-38, VMY^{+}09]. Merlin [CCL^{+}07, HF06, Les07-41, OLLG05]. Merlin/ [OLLG05]. merotely [CGGT09]. mesenchymal [CTP^{+}09, HHY^{+}08, IPG^{+}05, KZH^{+}07, LD07, LDKT06, LJM^{+}07, LVG^{+}09, MMS07, TVP^{+}06, WWC^{+}07, YA06, RLH^{+}09]. mesenchyme [PZG^{+}07]. mesh [LeB07b]. Meshorer [Sed08g]. mesoangioblasts [GSB^{+}06a]. mesoderm [YKP^{+}08]. message [Ros08c, Sho09b]. messenger [BJZ^{+}09, BFWH07, BAR^{+}08, FKJ^{+}09, FLA^{+}08, HLL^{+}09, Les09s, Pow07k, Wel07d, YSSH^{+}07]. messengers [BAdSG05]. Met-induced [PTL^{+}05]. Metabolic [CR06, Wel06-34, RMR^{+}09, SL05]. metabolism [GDW^{+}07, TEIS09, Wel06m]. metacaspases [VDVV07]. metalloproteinase [ADC^{+}07, WKW^{+}08]. metalloproteinases [LJN^{+}05]. metaphase [KPS^{+}07, MVC^{+}07, YCM^{+}08, ZAM07a, ZAM07b]. metaplasia [LBH^{+}05]. metastases [Wel06-51]. metastasis [MPS^{+}07, WMS^{+}06]. metastatic [SWM^{+}07]. metazoan [SHS^{+}06]. method [BGP^{+}06]. Methyl [BEN^{+}05]. Methylate [Wil07-36]. Methylation [FD09, BvOM^{+}07, CMD^{+}08, GTB^{+}07, GSY^{+}09, SHK^{+}07, TMH^{+}07]. methyltransferase [HBN^{+}09, JET^{+}07, XGX^{+}09]. methyltransferases [GTO^{+}07]. Meting [Les07-42]. Mfn1 [DC07, RLDC09]. Mfn2 [DC07]. MgcRacGAP [KBN^{+}06, OYKY^{+}05]. Mgm1 [DDRL^{+}09]. Mgm1p [SDI^{+}06]. MHC [LeB07-30, Les09l, STvdB^{+}09, VLFA^{+}07, WHC^{+}07, Wil07-42]. Mia40 [SPK^{+}09]. MICA [TRC^{+}08]. MICB [TRC^{+}08]. Mice [SNG^{+}07, Wel05-39, BJM^{+}06, BLL^{+}06, CGK07, CWM^{+}09, CML^{+}06, EVF^{+}08, GHDS^{+}07, HZL^{+}07, JLB^{+}05, KHB^{+}05, KSL^{+}05, KWU^{+}05, KPS^{+}07, LeB06-79, LDR^{+}08, LFW^{+}09a, LFW^{+}09b, MBdR^{+}08, NRJ^{+}09, OFfL^{+}07, PUG^{+}06, PXD^{+}07, RP05, SKvdBW^{+}06, Sed08k, SVS^{+}07a, SVS^{+}07b, WZW^{+}09, WWO^{+}08, WRKC^{+}09, WLLF^{+}08, YMD^{+}07, YAV^{+}05, ZDD^{+}06]. Michael [Sed09i, Wel07i]. microbes [Les07h]. Microbots [Wil07-37]. Microcephalin [TMZ^{+}09]. Microclustered [Tum06-29]. Microclusters [TLK^{+}06, KVS^{+}09, STP08]. Microdomain [Wil08q, GC05, SKO^{+}08]. Microdomains [PMW05, DBW^{+}06, GMS^{+}08, KDV^{+}09, LZ08, NKK^{+}06, Sco08]. microenvironments [CCL^{+}08]. Microglia [Les07-43]. microphthalmia [BBG^{+}05]. microphthalmia-associated [BBG^{+}05]. microRNA [KLS^{+}06, PS08, Wil09b, vRG08, YWW^{+}08]. MicroRNAs [Kat07d, FTB^{+}09, KVC^{+}09]. microscopy [GBNK08a, GBNK08b, HKV^{+}05, OHK07, Wat09, Wel06-27, Wil06g]. Microsomes [Wel05-40]. microtrabecular [Pow05r]. Microtubule [KKW^{+}08, KLR^{+}07, MDC05, Rob08v, Wil06i, WNB08, BBvD05, BKLT^{+}08, BVS^{+}09, BGP^{+}06, CKG^{+}08, CCSBM^{+}07, CBMCSS08, DvCvH^{+}08, DHR^{+}08, EMPK^{+}07, EGR^{+}05, GHM^{+}08, GHDS^{+}07, GWS09, GMZ^{+}08, JWC^{+}08, JSPT05, JBF^{+}08, KKB^{+}06, KNP^{+}05, KDG^{+}09, KLG^{+}09, LF07, LGZ^{+}05, Les09m, LHL^{+}08, MPD09, MFC^{+}09, MKGL^{+}05, MSB06, MK07, MGBM^{+}09, MSK05, MB07, NYI^{+}05b, OHK07, OOS07, OOS09, PTF^{+}06, PWL^{+}09, PSS^{+}05, PHM^{+}09, RYN05, RBJL06, RMV06, SGB^{+}06, SRE^{+}05, SSZR07, TKKT07, TÖL^{+}06b, Tom08, ToMY^{+}08, VMSV08, Vau05, VCD^{+}06, WEL^{+}07n, Wil07r, Wil07-60, WKCY^{+}08, Wit08, WWvD07, WTH05, YCM^{+}08, YGR^{+}08, ZRBS07, ZCJ^{+}08, ZHYE08, KLG^{+}09]. microtubule-associated [OOS07, OOS09, PHM^{+}09]. microtubule-based [EMPK^{+}07, GHM^{+}08, SSZR07]. microtubule-binding [LHL^{+}08]. microtubule-bound [CBMCSS08, JSPT05]. microtubule-dependent [BBvD05, GHDS^{+}07, MK07, NYI^{+}05b, TKKT07]. microtubule-destabilizing [ZHYE08]. Microtubules [Bon05f, BMK^{+}06, BYR08, LeB05-57, Pow05n, Pow05o, Pow05p, TA07, Wel05-41, Wel06-35, ACsHS08, BKHR05, BMS^{+}08, BFWH07, CYC^{+}06, DMDO08, FvD08, GZBM^{+}06, GC05, HH06a, HFM^{+}05, HH06b, LeB05l, LeB05-100, LeB07b, LeB07-36, Les06q, Les08l, MOS^{+}09, MRvD^{+}05, MS07, NLL^{+}06, OF08, Pow05a, Pow06c, Rob08-55, RMV06, RW05, SON06, SBW^{+}07, Sho09j, Sho09z, Sho09a, SMR09, SHH^{+}08, TAM^{+}06, TT09, Tum05-62, Tum06-38, Tum06-47, VSG09, WDWC08, WWS05b, ZHI^{+}08]. microvillar [HGR^{+}06, MT07, Pow07g, Pow07h]. Microvilli [Sho09-39, CSM^{+}06, DTY^{+}05b, LeB05m]. microvillus [MHS^{+}09]. mid [MSO09]. mid-blastula [MSO09]. Mid1 [CWRG^{+}08]. Mid1p [HYB08]. Midbody [DMC^{+}07, CK07]. middle [LeB06-43, Les07i, Wil07-50]. middleman [Pow05s]. midgut [CTGVGFF08]. midline [Tum05-73]. midzone [KLRS07, MOS^{+}09]. Miesenböck [Wel07f]. Mig [MQWP08]. Mig-2 [MQWP08]. migrate [ONW^{+}07, WMSB^{+}06]. Migrating [LeB06-40, Wel07j, Wil07-38, EBMG09, FKN^{+}07, PGP^{+}07, PJMR06, SCS^{+}05, VMZW^{+}07a, VMZW^{+}07b, VMKW^{+}08, WWS05b, ZHI^{+}08]. migration [BRM^{+}07, BMG^{+}07, BBL^{+}08, CCL^{+}08, DWMY09, EPV^{+}05, GdTAR07, GSB^{+}06a, GGMBF08, GMM^{+}08, GPJ^{+}06, GGEd^{+}07, GAK^{+}05, HOM^{+}05, HTKR06, HLA^{+}06, IPG^{+}05, KRZ^{+}08, KBR^{+}08, KSR^{+}08, LeB07-58, LLJ^{+}05, MFA^{+}07, MCF^{+}05, MWMS07, McN09, NPHT08, NBM^{+}08, NTY^{+}05, NYN^{+}06, OKN06, OMGW^{+}06, OV09, PR05, PEER^{+}05, PTL^{+}05, PRK^{+}07, PRK^{+}09, PEW^{+}05, PdBR^{+}07, QCGM05, SRL05, SSD^{+}09, SDN^{+}05b, SWM^{+}07, SWG^{+}05, SLWA07, TTG^{+}09, pTyLZ^{+}09, Tum05-34, Tum05-52, WC07a, Wel05s, WCN07, WXM^{+}05, XSVGR05, YMCT08, vBAS^{+}07]. Mih1 [PPK08]. Mih1/Cdc25 [PPK08]. Miller [Sed07b]. Milton [LeB06-41, GMSS06, RG06b]. MIM [GQSHO05]. MIM/BEG4 [GQSHO05]. mimicking [SCM^{+}07]. mimicry [INB06]. Min [CGK07]. Mina [Sho09-40]. mind [CN06, NVEF07]. Minor [Rob07h]. minus [BGP^{+}06, GNV05, LeB06-42, Tum05-66]. minute [Tum05-76]. miR [FTB^{+}09, KLS^{+}06, RGA^{+}06]. miR-206 [KLS^{+}06, RGA^{+}06]. miR-8 [FTB^{+}09]. miRNA [LeB05y, MZH^{+}08]. miRNAs [Tum06-30]. miro [RG06b]. Mirror [LeB07-31]. Mis12 [KCH^{+}06]. Mischel [Wil08s]. misfires [Rob07i]. misfolded [BXN^{+}06, OLC^{+}07, Sed09c, SN05]. Misfolding [Rob07i]. mislocalization [CMK06, HJK^{+}06]. misorientation [JSFP08]. misregulated [EW08]. missegregation [JMB^{+}07]. misses [Ros07b]. Missing [MPS^{+}07]. Missing-in-metastasis [MPS^{+}07]. Misu [HBN^{+}09]. MITF [BBG^{+}05, LeB05-58, LTDH05, WM05]. Mito [Wel05-42]. Mitochondria [LeB07-32, Les07-44, Rob08w, BMT^{+}07, CGS05, CRV^{+}06, GRGP07, GMSS06, GSH^{+}08, GGC05, HK07, IPM^{+}05, LeB06-41, LeB06-50, LeB06-73, LeB08h, LeB08z, Les07-42, Les08-59, McB08, MNH05, MS06, NTSY08, OHP^{+}09, PKM^{+}05, RSS^{+}09, RG06b, Sco08, THY^{+}06, TEIS09, VW08, WZM07]. Mitochondrial [CMK06, GSHD08, GG09, HHS^{+}09, Les08-33, Les09w, PhXD^{+}06, Rob08-35, Sho09-41, SMM^{+}08c, UDC^{+}07, WRKC^{+}09, AFN^{+}08, BWL^{+}08, BKE^{+}07, BbCQ^{+}09, CBZY08, COB^{+}08, CLA^{+}05, DDRL^{+}09, DC07, DJV^{+}05, ERM^{+}09a, EHWL06, FTM^{+}06, GRGP07, GGC05, GKvdB07, HWN^{+}07, HLL^{+}08, HMR^{+}07, KNY07, KCAS05, KGM^{+}07, KRG^{+}08, KHSM07, Les07-39, Les08g, LLH^{+}09, MHM^{+}05, MLH^{+}06, NRL^{+}05, OTH^{+}07, PUG^{+}06, Pow05x, Rap05, Rob08x, RLDC09, SSS^{+}08a, Sed09e, SDI^{+}06, Sho09-33, SPK^{+}09, SCF^{+}07, SGKP^{+}07a, SGKP^{+}07b, SBV^{+}06, THS^{+}09, VBNR06, WPH^{+}07, WvdLH^{+}07]. mitochondrion [HCL^{+}07, WAW08]. mitofusins [LeB07-48]. Mitogen [PTL^{+}05, KZH^{+}07, MBR^{+}06]. mitogen-dependent [KZH^{+}07]. Mitogen-inducible [PTL^{+}05]. mitogenic [DLBSD^{+}05, VSR^{+}08]. Mitosis [SIL^{+}07, DP07, DSN^{+}07b, JBFC07, Kat07c, KST^{+}07, KNP^{+}05, KYK^{+}06, LeB06-33, LeB06-60, Les07t, LFMO^{+}09, LKL^{+}05, LRBM09, MHC^{+}09, NNK^{+}08, OLH^{+}08, OYKY^{+}05, PSS^{+}05, SS06b, SWS^{+}08, TST08, VML06, Wel06-52, WBT^{+}08, YGY^{+}05, ZKN^{+}06]. Mitotic [CFU^{+}09, Rob08b, Sho09-42, BLMS06, BFWH07, BYR08, BBM^{+}07, CKG^{+}08, CBMCSS08, CDW^{+}05, CN06, CBIO05, DP07, DZW^{+}08, FVSC08, FDM^{+}08, FDV^{+}06, GNV05, HCK^{+}07, HHZ^{+}08, HBN^{+}09, JCVB08, JWC^{+}08, JSFP08, KKW^{+}05, Les07g, Les08h, Les08-36, LRBM09, LZ09, MDC05, MFH05, MFO09, MAOM06, NYM^{+}08, NDS^{+}07, OLKK09, SCD^{+}08, SCMV07, TTYH09, TLDB08, TMZ^{+}09, Tom08, WS09, WOWB08, YKBR09, Yu06, ZCJ^{+}08, ZMH^{+}09]. mix [BKMW08]. Mixed [Wel06-37, Wel06-36, Sho09h]. Mixed-up [Wel06-37]. Miz1 [GFH^{+}06]. mK3 [WHC^{+}07]. MKK4 [SLZ^{+}06]. MLC2 [SWI06]. MLC2-based [SWI06].

Mlp2p [NdCF^{+}05]. MMP [TSN^{+}07]. MMP-mediated [TSN^{+}07]. MMPs [LeB05-59]. MMR [KSL^{+}05]. Mms2 [AZP^{+}05]. Mnt [WZO^{+}05]. Mob1 [MHC^{+}09]. Mobile [NFI^{+}09, Tum06-31, DTR^{+}05, LMZ^{+}05]. mobility [KCD^{+}06b, MCA^{+}06, SSH^{+}05]. mobilizes [BRKR06, Wil07-42, YHJ^{+}05]. mode [PVMH05, Tum05-48]. model [CMF08, DGP^{+}05, ELM09, KR07, LKC05, MGW^{+}08, MGBM^{+}09, OF08, RGW^{+}07, RAP08, SWS^{+}08, XSVGR05, Yu06, ZKS^{+}08]. Modeling [Tum05-80]. models [JKCW07]. modern [Mis09a]. modes [CCS^{+}06, EMD^{+}08, LWG^{+}05, RBC05, TPNS07a, TPNS07b]. modification [JYZL06, KKD^{+}07, LM05, Les07-48, WSFD07]. modifications [GTB^{+}07, HTYNK09, KKH^{+}06, MRK^{+}07, SLCB05]. modified [HHS^{+}09, MBJF^{+}08, Wel05s]. modifiers [JNHB08]. modulate [AEJ^{+}06, DSK^{+}06, XOV^{+}09]. modulated [BGdN^{+}08, SHB^{+}07]. modulates [BPL^{+}05, CSMP^{+}09, KZC^{+}08, KLG^{+}09, LHRP05, LBC^{+}07, MCRVC^{+}08, MSLS07, MGR^{+}05, OYTO07, OSB^{+}05, PAM^{+}07, SRN^{+}07]. modulating [CZD^{+}09, ELJ^{+}08, HGJ^{+}08, LYY08, LCS^{+}09, NBM^{+}08, PMK^{+}05]. Modulation [iWK06, ESG^{+}06c, LOWTL^{+}08, TSN^{+}06, XKYW07]. modulator [KHSM07, TPZG06]. module [TTG^{+}09, WvdLH^{+}07]. modules [oSCW07]. Moesin [CKG^{+}08, IKZ^{+}07, HF06]. mold [Les07-60]. Molecular [BSM^{+}08, FTT^{+}05, JBF^{+}08, JWG^{+}07, Les06u, MPD09, OV09, PV06, TKKT07, VCSG07, VS07, YLK^{+}05, ZKG^{+}08, tDAK^{+}05, ASB^{+}05b, BCJ^{+}07, BKMW08, FSG^{+}07, INB06, JYZL06, Joh09, JTN^{+}05, JASC08, KTN07, Les06-28, MJG07, NN07a, PPA^{+}08, RAS^{+}07, RAS^{+}08b, SWT08, SAD^{+}09, SWM^{+}06, Tak06]. molecularly [GMK^{+}05]. molecule [BMK^{+}07, DGG^{+}08, DCMB^{+}07, EMD^{+}08, HKIY09, KTN07, SFEK07, SFS^{+}07, UFY^{+}07, ZMA^{+}07]. molecules [KGH^{+}05, LeB08b, MBL^{+}07, OFY^{+}05, STvdB^{+}09]. molting [Sed08a]. Mom [Rob08l]. monitoring [RMP06]. monoallelic [LBVN09]. monocarboxylate [FA08, JLC08]. monocytes [TCW^{+}06, vKST^{+}07]. monocytes/macrophages [vKST^{+}07]. monolayer [CNMM^{+}06]. Monomer [CZM^{+}07, BKBM06]. monomeric [DSH^{+}07]. monomers [KOKM07]. monomethylation [SS07]. Monooriented [LeB06-43]. monophosphatase [SFB^{+}05]. monopolar [HCFM08]. monoubiquitination [NLG^{+}09]. Mopping [Wel06-38]. morphogen [CSM^{+}06]. morphogenesis [AWE^{+}05, CJW07, CG06, FSN^{+}06, GSiK^{+}09, GTX^{+}06, HLS^{+}09, JKDH08, JBYW06, LSR07, LPF^{+}06, LGT^{+}07, LR05, MCSY07, MAZ^{+}08, NLC^{+}05, SOC^{+}07, TCKV05]. morphogenetic [BYK08, CRLP07, ESZK06, dJPAW^{+}09]. morphogens [Wil08x]. morphology [HLL^{+}08, LeB05-82, SSD06, VBG^{+}08]. most [MDH^{+}06]. mothballs [Les06u]. Mother [Wil07-39, FDV^{+}06, ZKN^{+}06]. mother-bound [FDV^{+}06]. Motherly [LeB06-44]. motif [GFA^{+}09, KBC^{+}07b, MKSJ06, TLM^{+}07]. motifs [CML^{+}06, MSECB05, SMH^{+}09, Wel05-59]. motile [DvRF^{+}05b, HL06, KHWC08, LeB06-41, PH07, SBW^{+}07, VTT^{+}09b].

Motility [LeB07-34, LeB07-33, AFN^{+}08, BFG^{+}06b, CHCB05, DSH^{+}07, FG08, GRGP07, GC05, HJK^{+}06, IEGV06, JP06, JLM^{+}07, JTH09, KKW^{+}08, KNT^{+}06, KWY^{+}06, KNT^{+}07, LW07a, LDR^{+}08, LLT^{+}08, OPT^{+}07, PSL^{+}06, PLH^{+}08, PKK09, SGMPV07, SJL^{+}07, Smi07, TCKV05, Tum06y, Tum06-28, iWK06, WNF^{+}05b, YWJ^{+}07, YA06]. motion [NTHN^{+}08]. motogenic [THN^{+}06]. motoneurons [JBL^{+}07]. Motor [LeB08p, MTB^{+}09, PWL^{+}09, AFN^{+}08, BCJ^{+}07, CGJ09, DGP^{+}05, GHM^{+}08, GNV05, IEGV06, KHWC08, LeB07-64, Les07v, LSC^{+}06a, NEF^{+}06, PGO^{+}07, Pow06j, THS^{+}09, TMI07, TS06, VMSV08, VLFA^{+}07, WvdLH^{+}07, ZHYE08]. Motor-dependent [PWL^{+}09]. Motoring [Pow07g, Pow07h, Wel07k, Les09-28]. Motors [LeB05-60, Rob08y, ALB^{+}09, BKMW08, JRZ^{+}07, POCS^{+}06, Sch08, Sho09-47, SS09, Tum06j]. Mouse [MHL^{+}06, YDL^{+}06, YEL^{+}08, BvRJ^{+}05, DC07, DGP^{+}05, DUR^{+}05, FNH^{+}05, GER^{+}07, JSFP08, KBS^{+}06, KR07, LS09, LCT^{+}09, LG07, LG08, MVC^{+}07, NTS^{+}05, NFZ^{+}09, PMB^{+}09, PRB^{+}09, RMP06, SXH^{+}06, TPZG06, WSG06, YBT^{+}09]. mouth [LWL08a, LWL08b]. move [LeB05c, LeB07-32, Les07-46, Tum06-37, Wil08q]. movement [FIK^{+}05, FIK^{+}20, HSLI09, JWC^{+}08, LPA^{+}07, LeB05a, LSM^{+}05, RIHP06, Rob08-55, SSHW09, SON06, SS09, TSP08, WNF^{+}05b, WWvD07]. movements [KHH08, LSC^{+}05, Sho09-45]. moves [Dal09, Rob08-53, ZRBS07]. Moving [Tum05-46, Wel06-39, LeB07-55, Rob08-40, Sch08, Tum05-48, VW08, Wel05z]. MP1 [TTK^{+}06]. MP1-MEK1 [TTK^{+}06]. Mph1p [BSO^{+}08]. MPS1 [MDPG07, TST08]. Mps2 [JMG^{+}06]. Mps3 [BMSJ07, JMG^{+}06]. MRE11 [SMT^{+}08, DG05]. mRNA [CVACF07, DSH^{+}07, DCMB^{+}07, FBD^{+}05b, GNE^{+}07, HOZ^{+}09, LeB05-68, LeB05-92, Les07-48, Les07-51, LFH07, LGBD^{+}07, MBF^{+}07, MBF^{+}08, MKSJ06, Rob08j, RSSC06, SRBL06, SNC^{+}06, Sed07d, SBJ^{+}07, SCF^{+}07, SLR^{+}06, TW07, YOSM06, ZEC^{+}08]. mRNAs [CTS^{+}05, HAT^{+}05b, KHBV05, LeB06u, LeB06-45, LeB07s, NAG^{+}08, PMB^{+}09, WGKA^{+}08, Wil07-40, WvNS^{+}07]. mRNP [DP06, KSA^{+}05, NAG^{+}08, Sho09-43]. mRNPs [HCC^{+}07, dPMF^{+}09]. Mrt4 [LLW^{+}09]. Msb2 [VDA^{+}08]. Msk [XYC^{+}07]. MSK1 [OMLP^{+}08]. Msps [BKHR05]. MT [LeB05r]. MTOC [MCRVC^{+}08]. mTOR [GPN^{+}06, LD07, LeB08q, RVBZ^{+}08, RMR^{+}09, WWX^{+}06]. MTs [Pow06k]. mucin [VDA^{+}08]. MUG [Sho09-44]. MukBs [Rob08z]. multiadaptor [RFBB07]. multicellular [VP05]. Multifunctional [Wil07-41]. multiinnervated [NBS^{+}08]. multinucleated [GHP06]. Multiparametric [WKIKG09]. Multiple [DPL^{+}08, SKE^{+}05, Wil08r, EMD^{+}08, FHMC09, OTH^{+}07, RKN^{+}07, ST09, SGKM08, SCK^{+}07, TDP^{+}06, TCKV05, UDC^{+}07, VTRB06, WC07b, WPDZ^{+}06, YSB^{+}07]. multiply [Les07-27, MH07]. multipotent [MLHS06, NPB^{+}05, TMW^{+}05]. multiprotein [DLH^{+}09, KCOS06]. multispanning [KMS07]. multistep [TM05]. multivesicular [FSR^{+}07, LJA09]. Munc18 [CCS^{+}06, DXC^{+}09]. Munc18-1 [DXC^{+}09]. munching [Wil06i]. muramyl [BAR^{+}05]. MURC [BLH^{+}09].

MURC/ [BLH^{+}09]. MuRF1 [CBG^{+}09]. MuRF1-dependent [CBG^{+}09]. murine [CAL^{+}07, DPL^{+}08, NPBP07]. MUS81 [FPP^{+}08]. MUS81-dependent [FPP^{+}08]. Musashi1 [KII^{+}08]. Muscle [DGP^{+}05, KLS^{+}06, LeB05-61, Les07-45, Les09x, RMR^{+}09, WBK^{+}06, AWE^{+}05, BLL^{+}06, BYK08, CBG^{+}09, CMN^{+}08, DLT^{+}07, DFB^{+}07, FBY^{+}06, FKS^{+}06, GSB^{+}06a, GEN^{+}06, GNE^{+}07, GPC^{+}06, HLA^{+}06, JP06, LPWP^{+}07, Las09, LeB05w, LeB05-34, Les07-40, LRS05, MNF^{+}07, MSV^{+}05, MQL^{+}06, MLHS06, NPMZ06, NSM05, NVEF07, NRJ^{+}09, OYTO07, RGW^{+}07, RMZ^{+}06, RYD08, RKN^{+}07, Rob07i, Rob07j, RMP06, SDL^{+}05, SYA^{+}08, Sho08, Tum05-61, Tum06c, WCB^{+}08, Wil09i, ZMW^{+}08]. Muscle-specific [KLS^{+}06]. Muscular [Wel06-40, JBL^{+}07, MBL^{+}07, MDL^{+}09, RGW^{+}07, SVS^{+}07a, SVS^{+}07b]. Musculin [HMM^{+}05]. Musculin/ [HMM^{+}05]. muskelin [VBG^{+}08]. must [Wel06-53]. mutant [BDS^{+}08, CMK06, HMS^{+}08, JMM^{+}06, SFY^{+}05, SBJ^{+}07, TT09, WDWC08, WZM^{+}08, WRKC^{+}09]. Mutants [CZP^{+}05, ECR^{+}09, LeB05-53, LeB06e, ZS08]. Mutation [Rob08-31, KHB^{+}05, LSC^{+}06a, MGW^{+}08, ZDD^{+}06]. Mutational [STG05]. Mutations [LDR^{+}08, CGK07, DC07, GLX^{+}09a, GLX^{+}09b, MWY^{+}09, MAH^{+}05, MXR^{+}05]. Muthuswamy [Sed09k]. MVB [SOW^{+}09]. Myb [MHM^{+}07]. Myc [GFH^{+}06, HPT^{+}07, YMDL09, NTB^{+}06, YZO^{+}07, BDS^{+}08, NCB^{+}08, NCB^{+}09, WZO^{+}05]. MYC-dependent [YZO^{+}07]. MyD88 [KR07]. MyD88-deficient [KR07]. Myelin [BW06, GHX^{+}07, LGZ^{+}05, TDG^{+}06, Tum05-36, VCSG07, WTE^{+}08, YBK^{+}06]. myelin-associated [VCSG07]. myelinated [DG08, MMT^{+}05]. myelinating [LHRP05]. Myelination [Cha07, CWNF^{+}09, GKCM^{+}08, MEG^{+}07, NZQ^{+}07, PBAF^{+}08, Sal08, YBT^{+}05]. myeloid [NBB^{+}09]. myelomonocytic [SDL^{+}05]. Myo1c [NEF^{+}06]. Myo2 [AFN^{+}08]. Myo4p [DSH^{+}07]. Myo6 [Tum05-81]. myoblast [DFB^{+}07, JP06, PBD^{+}06, Tum06-33]. Myoblasts [PPA^{+}08, SMW^{+}07]. myocardial [CULS^{+}07, PRB^{+}09]. myocardin [MMS07]. myocardin-related [MMS07]. Myocyte [PRB^{+}09]. myocytes [SHX^{+}07]. MyoD [GEN^{+}06, GNE^{+}07, IPAR05, RGA^{+}06]. MyoD-positive [GEN^{+}06]. Myofiber [KFR^{+}08]. myofibril [ERS08]. Myofibroblast [WRMH07]. myogenesis [BWL^{+}08, KCS^{+}06, TKB^{+}06]. myogenic [FKS^{+}06, GLS^{+}09, HLK^{+}07, IPAR05, KBY^{+}08, OYTO07, YTY^{+}07]. MyoJ [JTH09]. myopathy [CMN^{+}08, RMR^{+}09]. MyoR [HMM^{+}05]. Myosin [API^{+}07, FMK^{+}09, LeB05-62, LeB07-36, LeB08s, LeB08r, LSR07, MCSY07, NEF^{+}06, PCN^{+}05, Pow05q, Rob07j, Rob08b, Sho09-45, WTE^{+}08, Wil07-42, AFN^{+}08, Bon05g, CMF^{+}09, CCL^{+}05, DSH^{+}07, FvD08, GCS^{+}07, HLP06, JLM^{+}07, JTH09, KCOS06, LLH^{+}07, LeB06-46, LeB07k, LeB07-35, LSM^{+}05, LB08, MBS^{+}08, OWM05, SRA^{+}05, SS09, TPNS07a, TPNS07b, Tum05-77, VSG09, VW08, VLFA^{+}07, VMKW^{+}08, WVB07, WTH05, YWJ^{+}07, YMC05, LeB06-47, MT07, WOWB08, IGFG^{+}08, PRK^{+}07, PRK^{+}09, SBMP05, TMI07, TD08]. Myosin-10 [WOWB08]. Myosin-1a [MT07]. Myosin-driven [FMK^{+}09, LSM^{+}05]. Myosins [Les07-46, Les09y, VMZW^{+}07a, VMZW^{+}07b]. myotendinous [WCB^{+}08]. myotonic [SBJ^{+}07]. myotube [AWE^{+}05, MOCC05]. Myristoylation [LeB05-63, CLA^{+}05]. mystery [Pow05m]. Myt1 [NYM^{+}08].

N [CLA^{+}05, GOM^{+}09, HTKR06, LR05, Cha07, CHO^{+}09b, CWM^{+}09, HWN^{+}07, KVS^{+}09, MMM^{+}05, MBJF^{+}08, OYM^{+}09, SFC^{+}08, TdVM^{+}07]. N-cadherin [LR05, SFC^{+}08]. N-ethylmaleimide [MMM^{+}05]. N-glycan [CHO^{+}09b]. N-glycans [GOM^{+}09]. N-myristoylation [CLA^{+}05]. N-syndecan [HTKR06]. N-terminal [CWM^{+}09, HWN^{+}07, KVS^{+}09, TdVM^{+}07]. N16 [ASK^{+}09]. Na/K [CWC^{+}08]. NAADP [BCC^{+}09]. NAADP-mediated [BCC^{+}09]. NAD [Tum05-47, WZC^{+}05]. NAD-dependent [WZC^{+}05]. NADH [CLA^{+}05]. NADH-cytochrome [CLA^{+}05]. NADPH [CKX^{+}08]. Naegleria [KKY^{+}05]. Nalp3 [LMV^{+}09]. name [Wel05-41]. NANOG [ZNP^{+}09]. nanoplatforms [BZYM^{+}08]. Nanotubes [BW08, LeB08t]. Nap1 [NPHT08]. narcissism [LeB06-59]. nascent [CSB^{+}06, HZOL05, KMS07, MLC^{+}07, PNB07, iYOW^{+}06]. natriuretic [RKB^{+}07]. natural [TLK^{+}06]. naturally [LPWP^{+}07]. Nbk [GEH^{+}07, MDET08]. Nbk/Bik [GEH^{+}07]. Nbk/Bik-induced [GEH^{+}07]. NBR1 [MYMH09]. NBS1 [DCA^{+}06, MBJF^{+}08, SMT^{+}08]. NCAM [BLS^{+}05, BSL^{+}08, FCB^{+}09, FCB^{+}19, SSLS05, SLNS06]. NCAM-mediated [BLS^{+}05]. Nce102 [FMA^{+}09]. NCS [LeB06-49]. NCS-1 [LeB06-49]. NDC1 [SHS^{+}06, OSM^{+}09]. Ndc10 [MHFH06]. Ndc1p [MMCW06]. Ndc80 [WKCY^{+}08]. NDEL1 [ToMY^{+}08, HMG^{+}08]. Nebulin [BLL^{+}06, MSV^{+}05]. Nebulin-deficient [BLL^{+}06]. Necdin [DFB^{+}07]. necessary [AOT06, DLHW09, HRB^{+}06, LBI^{+}08, LLL^{+}06, Les09e, NPV^{+}07, PMB^{+}09, PB05, Sho09p, Wil08c]. neck [ZKN^{+}06]. necks [Wel05-60]. Necl [FIK^{+}20, FIK^{+}05]. Necl-5 [FIK^{+}20, FIK^{+}05]. Necls [PT07]. necrosis [HZL^{+}07, PRB^{+}09, UFY^{+}07]. necrotic [ASSST06, SFP^{+}05]. Nectin [MEG^{+}07, Bea06, FIK^{+}05, TMH^{+}06, FIK^{+}20]. nectin-3 [FIK^{+}05, FIK^{+}20]. Nectin-like [MEG^{+}07]. NEDD1 [HRB^{+}06, ZCJ^{+}08]. NEDD1- [ZCJ^{+}08]. NEDD1-dependent [HRB^{+}06]. Nedd4 [PGB^{+}06]. Need [LeB05-64, LeB06-47, LeB07k, Rob08l]. needed [DP07]. Needs [Rob08d]. NEEP21 [YWK^{+}08]. negative [ANI^{+}07, BLT09b, KBW^{+}06, LK05, MLC^{+}06, PBAF^{+}08, SSH^{+}05, SL05, WKH^{+}06, YTY^{+}07]. negatively [EHHK09, KLV^{+}08, LHZ^{+}08, NLS^{+}05, SLMK08, YOSM06]. negotiations [Sho09-37]. neighborhood [LC08a]. Neighbors [Tum06-32, Rob08t]. Neisseria [DDLC^{+}06]. Nemo [MPG^{+}09, NEF^{+}06]. NEMO/IKK [NEF^{+}06]. NEMO/IKK- [NEF^{+}06]. neocortex [ABY^{+}08, DCAC05]. nephrocystins [JNHB08]. nero [PCMSB09]. Nerve [LeB05-65, ANI^{+}07, HLJ^{+}07, WHB^{+}07]. nerves [EFCP07]. Nervous [LeB07-37, Les07-47, BTP^{+}07, CWNF^{+}09, GKCM^{+}08, PBB^{+}09, Wel07f, YBK^{+}06, ZTM^{+}08]. Nesprin [WLK^{+}05, LeB05-66]. Nesprin-3 [WLK^{+}05]. nets [Les09-36, Les07h]. network [BHK^{+}06, CSTT^{+}06, JBYW06, KES^{+}09, KFR^{+}08, LSD^{+}06, LRBM09, MTH^{+}06, NRL^{+}05, OSM^{+}09, RYN05, STBL09, SS09, SHB^{+}07, ShCM^{+}06, WHOS06, Wel05-38, WAW08, YWJ^{+}07, ZHI^{+}08, vBDF^{+}09]. networks [LC09, SKM^{+}08, Sho09k]. Neural [KII^{+}08, KLW^{+}05, WPDZ^{+}06, DBvdW^{+}06, DMC^{+}07, HMG^{+}08, HTKR06, JAV^{+}07, KMK^{+}05b, NCB^{+}08, NCB^{+}09, NPB^{+}05, PVMH05, QWD^{+}08, SRXvdK06, TMW^{+}05, TCKV05, WMSB^{+}06]. Neuregulin1 [HCRT08]. Neuregulin1-ErbB [HCRT08]. neurite [BFL^{+}05, BLS^{+}05, BSL^{+}08, CFBM05, HSB^{+}07, LH05, LeB05-30, MLC^{+}06, PTL^{+}05, RFBB07, SSLS05, VVL^{+}08, VCSG07]. neurites [LeB05-62, LeB06-36]. neuritogenesis [ANI^{+}07]. neuroblast [CSW08]. neuroblasts [OV09]. neurodegeneration [CYY^{+}08, DNK^{+}05]. neurodegenerative [FSR^{+}07, IPC09, Les07-45, WLE09]. neuroectoderm [HT08]. Neurofascin [HXO^{+}07, ZTM^{+}08]. neurofascin’d [Les07a]. neurofibromin [LLHH07]. neurofibromin-PKA-Ena [LLHH07]. neurofibromin-PKA-Ena/VASP [LLHH07]. neurofilaments [HFM^{+}05]. neurokinin [PWM^{+}06a]. neurological [WZM^{+}08]. neuromuscular [ABY^{+}08, LJY^{+}07, LCL^{+}05, MDL^{+}09, MPG^{+}09, NVJ^{+}08, PFD06, YJJ^{+}09]. Neuron [TDG^{+}06, HLM^{+}08, LSC^{+}06a, Pow09b, Rob07i, Wil07-43]. neuron-specific [HLM^{+}08]. Neuronal [Dal09, LeB06-48, MTL^{+}06, BFL^{+}05, BWJ05, CGS05, DXC^{+}09, FTT^{+}05, GDW^{+}07, GMM^{+}08, HOR08, HHFC06, HlQY^{+}06, LeB05t, LeB05-82, LeB08-28, Les06b, MLC^{+}06, ML05, MSH^{+}05, NJS^{+}06, RAP08, SSLS05, SFY^{+}05, Ste08, TSK^{+}06, UFY^{+}07, VVL^{+}08, VWM^{+}07, WvNS^{+}07, WNB08, YGW06, ZTM^{+}08]. Neurons [LeB05-67, LeB06-49, Les08-34, BbCQ^{+}09, DNG^{+}06, DGP^{+}05, DTL^{+}07, EMPK^{+}07, ESG^{+}06c, FM05, HWSJ08, Hat05a, HZR05, Kat06i, KKRP05, LeB05-50, LeB05-104, LeB06n, Les06-27, NJS^{+}06, PGO^{+}07, Wil07l, Wil07-38, WSFD07, XLH^{+}07]. neuropil [WZM^{+}08]. neuroprotective [YBK^{+}06]. neuroprotects [Wil09k]. neurotoxicity [SLK^{+}08]. neurotoxin [BFL^{+}05, DTL^{+}07]. neurotransmission [MAG05]. Neurotransmitter [OMF^{+}05, AACS^{+}06, BSMT05, MSLS07, SYM^{+}08a]. neurotrophic [VVL^{+}08]. Neurotrophin [JYZL06, HSB^{+}07, SBM^{+}07]. neutrophil [FAG^{+}07, KNT^{+}06, LPA^{+}07, VWK^{+}06, WLS^{+}09, WVB07]. neutrophil-like [WVB07]. Newly [MCD^{+}07, DMDO08, MG08a, MBF^{+}07, MBF^{+}08]. Newt [LeB06-51]. nexin [HRK^{+}09]. next [Pow09c]. NF [BWL^{+}08, CWC^{+}07, KBW^{+}06, LDVV06, LHM^{+}09, OLLG05, OSB^{+}05, PGP^{+}07, RRA^{+}09, SLZ^{+}06, STL^{+}07, TCW^{+}06, Tum05t, ZTKK05]. NF- [BWL^{+}08, GBNS07, KBW^{+}06, LDVV06, LHM^{+}09, OSB^{+}05, PGP^{+}07, RRA^{+}09, SLZ^{+}06, STL^{+}07, TCW^{+}06, Tum05t, ZTKK05]. NF-2 [OLLG05]. Nf2 [CCL^{+}07]. Nf2/Merlin [CCL^{+}07]. NFAT [WWX^{+}06]. NFATc1 [NIC^{+}09b]. NFBD1 [BJLM^{+}05]. NG2 [MWMS07, pTyLZ^{+}09]. NgCAM [AMS^{+}05, YWK^{+}08]. NGF [XJF05]. NH [BWJ05, BHG^{+}05, IPAR05, KNKH06, MM05, VBSM06]. NHE1 [FKN^{+}07, SSD^{+}09]. NHEJ [HS09b, KKK^{+}05]. NHERF [MSSF05].

NHERF-2 [MSSF05]. NHK [CBIO05, LCO07]. NHK-1 [CBIO05, LCO07]. nicastrin [SRD^{+}07]. niche [DL08, HKN09, WLC08, Wel05r]. Nicotinic [KBC^{+}08, HCRT08]. Niemann [CC09]. NIH [OIT08, Ros05]. NIH-funded [Ros05]. NIH-src [OIT08]. nitric [FHFB^{+}06, NBS^{+}08, PBD^{+}06]. NK [KCOS06]. NMD [DCMB^{+}07, DCMB^{+}07]. NMDA [KZT^{+}08, KZT^{+}09, Ste08, TSN^{+}07]. Nmnat [CWM^{+}09]. Nmnat1 [ASK^{+}09]. nNOS [CVD^{+}05]. No [LeB05-70, Les07-49, Les09z, Rob08-28, Rob08-29, Rob08-30, Wel05-44, LeB07-62, Les08-46, Wel05-39, Wel05-48, Wil06l, Wil07-54, Tum06-33, Wel05-45]. NOD2 [BAR^{+}05]. Nodal [DLL^{+}08]. node [CTS^{+}06c, HRK^{+}09, Les07-71]. Nodes [DZG^{+}07, HYB08, WSK^{+}06, YOHR07, ZTM^{+}08]. nodulin [VCN^{+}05]. NOMA [RFBB07]. Nomenclature [KCE^{+}06, KBI^{+}08, PFG^{+}05, SBC^{+}06]. Non [IPC09]. Non-cell [IPC09]. Nonadhesion [LeB05-71]. Nonadhesion-based [LeB05-71]. noncanonical [SMH07]. noncentrosomal [RMV06]. nonclassical [GDF^{+}05]. noncoding [BF09, FFS06]. nondisjunction [MTRD^{+}06]. nonhomologous [DG05, WC07b]. nonidentical [HR05]. noninvasive [PH07]. nonmechanical [KO06]. nonmuscle [Bon05b]. nonnatural [SHjC^{+}07]. nonproliferative [PMS^{+}07]. nonquiescent [ABA^{+}06]. nonribosomal [PT09]. nonsense [DCMB^{+}07]. nonsense-mediated [DCMB^{+}07]. nonspecific [TTNA^{+}06]. Nonubiquitous [Tum05-49]. Nonvesicular [RIHP06, GDR^{+}05]. Nop17 [ZKG^{+}08]. norm [Sed07d]. normal [FSN^{+}06, HYB08, LYTL07, PL06]. Notch [BYK08, CGLC^{+}08, LeB07-39, MOM^{+}06, NPV^{+}07, CLS^{+}06, GSC^{+}07, Gra07, HBK^{+}06, LNG^{+}05, LeB05-72, LeB07o, MG08b, NFC^{+}08, SLV^{+}07, SSAD08, TRS^{+}08, TRB09, VLK^{+}08, WMM^{+}09, WPWF08, WNL^{+}09]. Notch-independent [WMM^{+}09]. Notch1 [CZSG08, NMO^{+}07]. notochord [RHMW^{+}05]. Novel [FA08, FAG^{+}07, MSM^{+}07, NJS^{+}06, VBG^{+}08, AS09, BPL^{+}05, BW08, BFP^{+}07, BBFN07, CTS^{+}06b, CCP^{+}08, DBvdW^{+}06, DDB^{+}05, FFW^{+}07, GCSG^{+}05, GTHK05, GFA^{+}09, GHN^{+}09, GGE^{+}09, GD05, GSL^{+}07, GSV07, HFP^{+}05, HRS^{+}05, IFF^{+}05, KTA^{+}06, LGZ^{+}05, LCvD^{+}08, LRCH^{+}05, MCF^{+}05, ML05, MF08, MSC^{+}07, NS07, OOS07, OOS09, OBY^{+}06, OTH^{+}07, PSL^{+}06, PMK^{+}07, QIH^{+}06, RP07, RKB^{+}07, SSHP^{+}07, STG05, SOR^{+}05, SPK^{+}09, STL^{+}07, SRQL09, TWC^{+}07, TSS^{+}06, Tum05-50, WLK^{+}05, XKYW07, YSSH^{+}07, yYMM05, YCK^{+}05, ZTZ^{+}09]. novo [DLHW09, KKY^{+}05, KMSLS06, LEH^{+}05, ULN^{+}07]. Noxa [NMV^{+}07]. NPC [DTR^{+}05, Les07-28, Rob07k]. NPC-mediated [DTR^{+}05]. Npr2 [SSJ^{+}07]. Nsa1 [KRPH08]. NSF-independent [JWSS05]. NSun2 [HBN^{+}09]. Nuclear [CSSvM08, EVF^{+}08, qGyWX^{+}08, HdL06, KGH^{+}05, LN07, Les09-27, MSW^{+}07, MSB06, NN07b, SNPCG06, TW07, Wel06-41, YM06, AH08, AVHH09a, AVHH09b, BAR^{+}05, BZA^{+}06, BF09, BGdN^{+}08, BMSJ07, CSM^{+}06, CYO^{+}09, CCL^{+}05, CGJ09, CTS^{+}05, DGG^{+}08, DLHW09, DCA^{+}06, DDB^{+}05, DZW^{+}08, FHO^{+}05, FKJ^{+}09, FSS^{+}09, FBY^{+}06, GTB^{+}07, GHP06, GGE^{+}09, GNPC^{+}08, GCS^{+}07, GFC^{+}08, HMG^{+}08, HFvdV^{+}05, HIK^{+}08, HAN^{+}09, HG07, HOZ^{+}09, JMM^{+}06, JMG^{+}06, KBN^{+}06, KP08, KFK^{+}05, KS08, LHS^{+}05, LC08a, LeB06r, LeB08-34, LWC09, Les05c, Les05f, Les06n, Les07-72, Les09j, Les09-29, LFH07, LRS05, LPM^{+}07, LCM^{+}08, MMCW06, MSL^{+}09, MDET08, MGR^{+}05, MSM^{+}06, MK07, NdCF^{+}05, OW06, OSM^{+}09, PNB07, Ped08, PV06, QIH^{+}06, Rex09, RM05b, Rob08u, Rob08-30, RW08c, RMKS08, SRX^{+}07, SON06, SSHP^{+}07, SHC07, SCVW09, SSS^{+}08b, SHS^{+}06, TTYH09, TFC^{+}06b, TTNA^{+}06]. nuclear [Tum05-68, Tum06-39, Tum06-46, UPBM06, Wel06-45, WBG^{+}08, WLK^{+}05, Wil07-32, XYC^{+}07, YR09, ZRH^{+}09, LeB06-52, MSW^{+}17, Wil08t]. nucleate [LBI^{+}08]. nucleation [BMS^{+}08, MTH^{+}06, VCD^{+}06]. Nucleator [Wil07-44, LeB07-38]. nucleators [QHB^{+}07]. nuclei [FML^{+}06, LeB05-55, PL06, Wel05z]. nucleic [BW08]. nucleo [RH09]. nucleo-cytoskeletal [RH09]. nucleocytoplasmic [FFS06, LSD^{+}06, VBG^{+}08]. nucleoid [HMR^{+}07]. nucleoids [DJV^{+}05, GSHD08]. Nucleolar [YWM^{+}08, HBN^{+}09, PT09, VBL^{+}05, Wel05-46]. nucleolus [MKC^{+}05, Mis05, Rob07f, Rob08l, Tum05-79, Tum05-30, VBL^{+}05]. Nucleophosmin [BDS^{+}08, Rob08-31, Di 08, YWM^{+}08]. Nucleoplasmic [MCA^{+}06, PHB^{+}07]. nucleoporin [DTR^{+}05, FSS^{+}09, SHS^{+}06]. nucleoporins [FSS^{+}09, MMCW06, MSL^{+}09, TW07]. nucleosomes [HW09]. nucleostemin [TM05]. Nucleotide [MBJM^{+}09, GSdBS09, LLT^{+}08, MXR^{+}05, OD07, STG05]. nucleotide-binding [MXR^{+}05, STG05]. Nucleotides [Wil06j]. Nucleus [Rob07l, CF07, CLR^{+}06, DCEM09, FFS06, HG07, KTI^{+}05, LCO07, LeB06-80, LeB07n, Les07-63, MAG^{+}08, Sho09-43, SHDT05, Tum05a, Wil06p, LeB06-52]. NudE [SCMV07]. NudEL [SCMV07]. Nuf [CAR^{+}08]. null [NRJ^{+}09, DDO^{+}08, SWB^{+}07, YMD^{+}07]. NuMA [SHC09]. Numb [CLS^{+}06]. number [SPS^{+}08, TKK^{+}05]. numbers [LeB05-94]. Numblike [CLS^{+}06]. Nunnari [Sed09e]. Nup153 [SRX^{+}07]. Nup157p [MSL^{+}09]. Nup170 [FSS^{+}09]. Nup170p [MSL^{+}09]. Nup214 [SRX^{+}07]. Nup2p [DTR^{+}05]. nutrient [RVBZ^{+}08, WOP^{+}06]. NXF1 [SRBL06].

O [Tum06-52, JASC08]. O-glycosylation [JASC08]. observed [MBF^{+}07, MBF^{+}08]. occur [NMO^{+}07]. occurring [LPWP^{+}07]. occurs [BGdN^{+}08, FSN^{+}06, GLH^{+}05, KCMB08, MSW^{+}07, MSW^{+}17, SRL^{+}06]. Oct [MLV09, Sed09g]. Oct-1 [MLV09, Sed09g]. Oct4 [SAD^{+}09]. oculomotor [SBC^{+}07]. ODA16 [AGL^{+}08]. odd [LeB08n]. oenocytes [Wel05-47]. off [LeB07-43, LeB08y, Les07-67, Nyb05n, Sal08, Sho09-39, Wil06b]. offense [Pow07a]. offloading [LKC05]. OL [NPHT08]. old [Les07c, Mel06, Pow09a, Wel05-43]. olfactory [AMERT^{+}07]. oligodendrocyte [GKCM^{+}08, SLCB05]. oligodendrocytes [LGZ^{+}05, LHRP05]. oligodendroglial [WGKA^{+}08]. oligomeric [SSP^{+}07]. oligosaccharide [KBC^{+}07a, SGK^{+}07]. oligosaccharyltransferase [KBC^{+}07a]. Oliver [Wil07-45]. OMA1 [ERM^{+}09a, HGA^{+}09]. Oncogene [MWY^{+}09]. oncogenesis [Di 08]. oncogenic [BHJ^{+}05, Rob08-54]. oncoprotein [EIK^{+}05]. One [DWMY09, LeB05-73, Tum05-51, Tum06-34, Wel05-48, ALB^{+}09, Les05e, Les07w, LNL^{+}08, Rob08-36, Rob08-55, Tum06w]. One-dimensional [DWMY09]. one-way [Rob08-36, Rob08-55]. oneness [Les07-30]. only [MDET08, GGO^{+}09, MGH^{+}09, NMV^{+}07, WHM^{+}07]. onset [CRLP07, KR07, YKH^{+}08]. onto [Pow05v, SPK^{+}09]. oocyte [LCO07, SORM05, yYMM05]. oocytes [DUR^{+}05, DPP^{+}07, FNH^{+}05, MVC^{+}07, WH06, ZHYE08]. oogenesis [BAR^{+}08, Gra07, HCB^{+}08, KR06, LC07]. OPA1 [ERM^{+}09a, HGA^{+}09, SCF^{+}07, GKvdB07]. Opa1-dependent [RLDC09]. open [ABvBHH06, LeB07-47, NRD^{+}06, Wil06k]. open-to-closed [ABvBHH06]. opened [PFGSV06]. opening [RM05a]. opens [LWL08a, LWL08b, Wil09o]. operated [LWBL06, RDY^{+}05]. Opposing [Sho09-47, ELJ^{+}08, FSnP^{+}06, SYM^{+}08b]. Opposite [ALB^{+}09, KGKV07, OF08]. Opposite-polarity [ALB^{+}09]. Opposites [Tum06-35]. optic [HLJ^{+}07]. optical [Wel05-49]. Optimization [SKO^{+}06]. optimize [AS09]. Optineurin [SRA^{+}05]. Orb2p [LB08]. orchestrate [WPWF08]. orchestrating [BFG^{+}06b]. Orchestration [CCS^{+}08]. orchestrator [Wil07a]. order [GLBS06, LeB06c, LeB06-64, MMC08, RCKD06, Wel06y]. Ordered [Tum06-36]. orders [Tum05-56]. organ [KKM09b, LeB07-59, WWC^{+}07]. Organelle [STBL09, LU05, Les05i, MHS^{+}09, SKM^{+}08, SRQL09, WAW08]. Organelles [Rob08-32, LeB05-105, LNL^{+}08, SM09, SSZR07, Wel05-64, Wil07-39]. organism [IH05]. organisms [KBC^{+}07a]. Organization [MRK^{+}07, SHX^{+}07, AKD06, AG09, AM05b, BRH05, BJLK^{+}06, BFB^{+}08, CHSV07, CKG^{+}08, CM09, DHR^{+}08, FDM^{+}08, GLX^{+}09a, GLX^{+}09b, GHDS^{+}07, GTB^{+}07, HMR^{+}07, HK05, IMY^{+}09, KLR^{+}07, LPG^{+}07, LeB05r, LF07, LC09, LZ09, LYTL07, MFC06, NWR^{+}08, NFZ^{+}09, RYN05, RKH^{+}09, SRA^{+}05, SLP^{+}06, TSK^{+}06, ToMY^{+}08, YCM^{+}08, LeB06-52]. organizer [HOH^{+}07]. organizing [BVGA^{+}05]. organs [HMS^{+}08]. orient [BJKF05]. Orientation [WKCY^{+}08, BWW^{+}08, BBL^{+}08, CFU^{+}09, Com07, JKDH08, VMSV08]. oriented [Bon05j]. Orienting [LeB05-74, SGB^{+}06]. origin [GHK^{+}09, HZR05, KMSLS06, Wel05-46]. Original [RP08, Hil08]. Origins [Les08-36, CCS^{+}08, LeB08f, WGL^{+}06]. ORP1L [JRZ^{+}07, RKvdK^{+}09]. orthologue [BBF^{+}07, DTY^{+}05b, JSC08]. oscillations [KBW^{+}06, MuMI^{+}06]. oskar [MKSJ06]. OSM [IEGV06, POCS^{+}06]. OSM-3 [IEGV06, POCS^{+}06]. osmotic [FTB^{+}09]. osmotransducing [FLA^{+}08]. Osteoblast [LSG^{+}06, GXJF07, LHZ^{+}08, MAH^{+}05, SSH^{+}05, WKH^{+}06]. osteoblast-dependent [WKH^{+}06]. osteoblastogenesis [WKH^{+}06]. osteoblasts [PRZ^{+}08]. osteocalcin [HGJ^{+}08]. osteoclast [MMU^{+}09]. osteoclastic [ZKR^{+}07]. osteoclastogenesis [RRA^{+}09, WKH^{+}06]. osteoclasts [GHDS^{+}07]. osteogenesis [LVG^{+}09, PRZ^{+}08]. osteoporosis [LeB07c]. other [JMM^{+}06, LNG^{+}05, PR05, Sed09b]. outcomes [GHY^{+}05]. Outdoor [Wel06-42]. outer [AGL^{+}08, ASH^{+}09, BHY^{+}08, CLA^{+}05, HHS^{+}09, HQF^{+}07, NPBP07, OTH^{+}07, PKK09, Rap05, SSHP^{+}07, Tum06q, WLK^{+}05]. outgrowth [BSL^{+}08, HSB^{+}07, LGZ^{+}05, MLC^{+}06, RFBB07, RAP08, SSLS05, STU^{+}08, VVL^{+}08, VCSG07]. outlier [LeB05-72]. outs [Les08b]. ovary [OLS08]. Overall [CSMP^{+}09]. overcome [FPP^{+}08]. overdone [LeB05d]. overexpression [KST^{+}07, LZL^{+}05, NTS^{+}05, WCR^{+}08]. overflow [LeB05-92]. overgrowth [HMS^{+}08]. overlapping [MDWH^{+}08, RMZ^{+}06]. Ovol1 [NTB^{+}06]. own [HR08, LeB06q, LeB06-80, LeB09a, Les05h, Les07-28, Rob08c, Wel05-38]. oxidant [Ter06]. oxidants [WXM^{+}05]. oxidase [CKX^{+}08, SMM^{+}08c, TRS^{+}08]. Oxidation [PWD^{+}06, LeB05-52, LeB06x, SMM^{+}08c]. oxidative [MLV09, MLH^{+}06, SPK^{+}09, SL05, SBC^{+}07, TP09, HK07]. oxide [FHFB^{+}06, NBS^{+}08, PBD^{+}06]. oxygen [BKE^{+}07, LeB07w, MG08a, Wel05c, XJF05]. oxysterol [RIHP06, SCR^{+}09]. oxysterol-binding [RIHP06, SCR^{+}09].

P [BMP08, GC05, MPS^{+}07, SMH^{+}09, SCK^{+}07, CRS^{+}06, DSN^{+}07a, ESG^{+}06b, GSB^{+}06c, GSB^{+}06b, NWO06a, PWM^{+}06b, RMZ^{+}07, TÖL^{+}06a, lWHMH06a, Wel06f, Wil06d, BAR^{+}08, DLL^{+}08, DCMB^{+}07, FBD^{+}05b, HCC^{+}07, NSS^{+}08, NAG^{+}08, Rob08j, Wil07k, Wil07-31, WSB^{+}09, ZEC^{+}08]. P-bodies [DCMB^{+}07, HCC^{+}07, Wil07k, Wil07-31]. P-body [FBD^{+}05b, ZEC^{+}08]. P0 [GHX^{+}07, LLW^{+}09]. p107 [LeB05-76, RMC^{+}05, VWM^{+}07]. P115 [SMR09]. p120 [DBR09, SYM^{+}08b, YA06]. p120-catenin [DBR09]. p120catenin [LPN^{+}08]. p130Cas [PEW^{+}05, SMM08b]. p14 [TTK^{+}06]. p150 [JRZ^{+}07, RKvdK^{+}09, LSC^{+}06a]. p16 [YTH^{+}09]. p19 [MZH^{+}08, NCB^{+}08, NCB^{+}09]. p190Rho [LLT^{+}08]. p190RhoGAP [BMR^{+}08]. p21 [JZN^{+}09, OSS^{+}07, RG06a]. p21-activated [OSS^{+}07, RG06a]. p24 [ARKS^{+}08]. P2X7 [PRZ^{+}08]. p32 [BFP^{+}07]. p38 [GLS^{+}09, JTN^{+}05, KBY^{+}08, Las09, TKB^{+}06, WWL07]. p38- [GLS^{+}09]. p38MAPK [OMLP^{+}08]. p50 [SLZ^{+}06]. p53 [DMBF^{+}07, GdTAR07, JZN^{+}09, KCD^{+}06a, LeB06e, Les07-50, MYO^{+}05, MZH^{+}08, NCB^{+}08, NCB^{+}09, SGDM06, Tum05-40, WKH^{+}06, Wil07-63, XKYW07, YTH^{+}09, ZHS06, LSG^{+}06, XKYW07]. p53-dependent [SGDM06]. p53-Mdm2 [XKYW07]. p53-mediated [XKYW07]. p53/p21 [JZN^{+}09]. p59 [BLS^{+}05, SSLS05]. p62 [BLB^{+}05, NSS^{+}08]. p62/SQSTM1 [BLB^{+}05]. p65 [TCW^{+}06]. p66 [MMW^{+}07]. p66shc [KSK^{+}06, SNG^{+}05]. p73 [HlQY^{+}06]. p75 [SBM^{+}07]. p85 [LFL^{+}05]. p90Rsk [DUR^{+}05]. p97 [WLY06, LDL^{+}09a]. p97-associated [WLY06]. p97/VCP [LDL^{+}09a]. PABP [KII^{+}08]. pace [Sho09v]. paces [LeB07-36]. Pack [Rob08-33]. packing [Sho09m]. Pacman [ZRBS07]. Pacman-flux [ZRBS07]. pact [Wil07g]. PAG [VSR^{+}08]. Painting [Sed09j]. pair [LW07a, THS^{+}09]. pairing [JGY09]. PAK [NWB^{+}06, LB08]. Pak-related [LB08]. PAK1 [LLJ^{+}05]. Pak1p [LB08]. Pak1p/Orb2p [LB08]. Palade [LRCH^{+}05, Mel08b, Wel05-56]. palmitoyl [OVL^{+}07]. palmitoylate [FHFB^{+}06]. palmitoylating [NFI^{+}09]. Palmitoylation [DDB^{+}05, GC07, LDS^{+}06, ALvdG06, Dal09]. Palmitoylation-dependent [GC07]. pancreas [Les08-30]. pancreatic [HOH^{+}08]. pancreatitis [Les08k]. Paneth [BDP^{+}07]. pannexin [VBG^{+}06]. Par-3 [CM06, LeB06-53]. Par3 [Les07x, STM^{+}07]. Par6 [EMMN^{+}05]. Par6-PKC [EMMN^{+}05]. Paradigm [RG06b]. paradox [Wil08a]. parallel [Pow05o, Rob08-32]. paralogues [Les07-73]. parasite [MBR^{+}05, OHH^{+}06]. Paraspeckles [BF09]. Paraxial [CG06]. park [Pow05o]. Parkin [McB08, NTSY08, OLC^{+}07, Sed08l]. Parkin-mediated [OLC^{+}07]. Parkinson [Wil07l]. ParM [LeB07-40]. PARP [SiTT^{+}08]. PARP-1 [SiTT^{+}08]. part [vRG08]. partial [BSO^{+}08, SWS^{+}08]. partially [SM09]. participate [MSSF05]. participates [BUJ^{+}08, FZT^{+}07]. particle [ABSK06, ELM09, PGL^{+}09, oSCW07]. particles [DMC^{+}07, GZBM^{+}06, KRPH08, LeB06-29, POCS^{+}06, SWL^{+}05, Wel05-56]. partition [KMK^{+}05b]. Partitioning [LeB06-54, CTS^{+}06a, CBMCSS08, CLA^{+}05, EMD^{+}08, FMK^{+}09, HGJ06, RCFB08]. Partner [Rob08-34, FFS06]. Partners [LeB07-41, BCJ^{+}07, Rob08g, TMBL^{+}08, Tum05z]. parts [Rob08s]. party [Sed09d, Wel06-51]. Pass [Wel07l]. Passenger [Les06w, GGB^{+}08, VML06, XOV^{+}09]. Passionate [Sed08n]. past [Wil06n]. patch [SBMP05]. path [Rob08c]. pathogen [LeB07-49]. pathogenesis [LeB07u]. pathogens [MRKH05]. pathology [PMC08]. pathway [BCT06, CCP^{+}08, CWC^{+}07, CTO^{+}08, CEY^{+}05, DvRF^{+}05b, DBDD06, DVF^{+}06, DPP^{+}07, ERM^{+}09b, FA08, FPP^{+}08, GXJF07, GGE^{+}09, GDR^{+}05, Gra07, GEE07, HBK^{+}06, HFP^{+}05, HR09, IH05, JWSS05, KBY^{+}08, KOVP09, KGDR^{+}07, KMSLS06, KKC^{+}07, KBC^{+}08, KWY^{+}06, KDD^{+}09, LEH^{+}05, LD07, LJA09, LeB06-32, LPN^{+}08, LLHH07, LPB^{+}08, LTAM^{+}09, MCF^{+}05, MSECB05, MF08, MSW^{+}07, MSW^{+}17, NCB^{+}08, NCB^{+}09, OMLP^{+}08, OTH^{+}07, PhXD^{+}06, PBD^{+}06, PCPW08, SBM^{+}07, Sed07b, SHC^{+}05, Sho09p, SFB09, SL05, SRD^{+}07, SMW^{+}07, Tum05-50, WJM07, WLC08, Wel05-50, WSFD07, YCR06, iYOW^{+}06, YTY^{+}07, YCK^{+}05, YWM^{+}08, ZMGC05, ZJH08, ZKN^{+}06]. pathways [AWR^{+}07, AMS^{+}05, BMRP07, CCS^{+}08, DSLY^{+}05, FHMC09, FHS^{+}07a, GPN^{+}06, GFD^{+}05, HCL^{+}07, KDD08, LRJY06, dJPAW^{+}09, RCS^{+}09, SMH07, TMY^{+}09, VKGV08, WWX^{+}06, WCN07]. PATJ [SSM05]. patrol [Wil07-52]. pattern [Pow07c, SLP^{+}06]. Patterned [LeB05-77]. patterning [JSC08, LJN^{+}05]. patterns [KCD^{+}06a]. Paul [Wil08s]. pause [Sho09a]. Pax3 [KCS^{+}06, RMZ^{+}06]. Pax7 [KCS^{+}06, OYTO07, RMZ^{+}06]. Paxillin [NWB^{+}06, AFM^{+}05]. PC12 [AOT06, ANI^{+}07]. PC4 [MRL^{+}08]. PCDH15 [DTY^{+}05b]. PCH [RGCWG09, TSS^{+}06]. PCNA [GFA^{+}09, LeB08m, LM05, SSS^{+}08b]. PCNA-binding [GFA^{+}09]. PDE4 [SBM^{+}07]. PDE4/cAMP/ [SBM^{+}07]. PDGF [RG06a, VSR^{+}08]. PDGF- [RG06a]. PDGF-induced [VSR^{+}08]. PDGFR [SSD^{+}09]. PDGFR- [SSD^{+}09]. PDK1 [PdBR^{+}07]. Pds5 [JGY09]. PDZ [HSB^{+}07, STBL09]. PDZRhoGEF [WVB07]. peace [BEH^{+}06]. pebbles [MOCC05]. PECAM [CMF08]. PECAM-1 [CMF08]. pedal [Tum05-27]. Peering [CB08]. Pegging [Rob08-35]. Pelkmans [Sho09-34]. pemphigus [WSB^{+}06]. penetrating [BGZ^{+}07]. Peptidase [LeB07-42, IBS^{+}05]. peptide [CCC^{+}06, KNT^{+}07, PCG^{+}05, RKB^{+}07, SZX08, YCK^{+}05]. peptide-generating [YCK^{+}05]. Peptides [Wil07-46, BW08]. perfect [LeB07-52]. perforin [SWB^{+}07, WSS^{+}06]. perforin-dependent [SWB^{+}07, WSS^{+}06]. perform [DBD^{+}09]. Pericentric [GER^{+}07, LG07, LG08, PL06, WSG06]. pericentrin [TMZ^{+}09]. pericentriolar [LR07]. pericyte [SBF^{+}06]. perinodal [EFCP07]. peripheral [BTP^{+}07, Bon05i, EFCP07, MMT^{+}05, PBB^{+}09]. periphery [BMSJ07, KS08, RW08c, TFC^{+}06b, TMI07, Tum05-72, Tum06-39]. periplakin [SNG^{+}07]. PERK [MGBH^{+}06, SGK^{+}07]. PERKs [Wil07x]. perlecan [ZMW^{+}08]. permanent [Wil07-65]. permeability [OSS^{+}07, VBG^{+}06, Wil07-67]. permeabilization [UDC^{+}07]. permeases [KGL07]. Peroxide [LeB06-55, GH05, MMM^{+}05]. Peroxiredoxin [KSE^{+}07, MLH^{+}06, PWD^{+}06]. peroxiredoxin-3 [MLH^{+}06]. peroxisomal [BVGA^{+}05, FFT^{+}05, GGBV^{+}07, LZM^{+}06, LNL^{+}08, MF08, PES^{+}07, TM06]. Peroxisome [Sho09-48, TM06, BVGA^{+}05, CMF^{+}09, FFT^{+}05, FMK^{+}09, GDF^{+}05, GGBV^{+}07, MMNH09, Rou08]. peroxisomes [BJC^{+}06, Bon05h, KMSLS06, Les07-66, MF08, MH07, Pow09c, Tum05-32]. persist [WMSB^{+}06]. persistence [OPT^{+}07, WZCB^{+}06]. persistent [HS09b, KDG^{+}09, MPD09, PEER^{+}05, WCN07]. persists [SFP^{+}05, SIL^{+}07]. perspective [Sed08g]. Perturbing [RBJL06, HAT^{+}05b]. perturbs [LSC^{+}06a]. Pes1 [HRS^{+}05]. Peter [Sed09j]. PEX16 [KMSLS06]. PEX16-dependent [KMSLS06]. Pex16p [MF08]. Pex16p-dependent [MF08]. Pex19p [MF08]. Pex19p- [MF08]. Pex3 [CMF^{+}09]. Pex3p [MF08, MMNH09]. Pex5p [PES^{+}07]. Pez [WWC^{+}07]. Pfa3p [SSSL05]. Pfa4 [LDS^{+}06]. Pfeffer [Sed09l]. PGE [TDP^{+}06]. pH [FBD^{+}08, FSH08, Har08]. phagocytic [BMG^{+}07, Les09-31, NBB^{+}09, TD08]. phagocytosis [CTS^{+}06b, FOB^{+}09, JGL^{+}06, LHL^{+}08, MYZ^{+}09, SDB^{+}05, SDS^{+}05, WPL^{+}06]. phagocytosis-like [CTS^{+}06b]. phagolysosome [RCKD06]. phagosome [FBO^{+}07, SHB^{+}07, YHD^{+}09]. phagosomes [CSB^{+}06]. PHAP [MDC^{+}08]. pharmacological [LLS^{+}09]. phase [ABA^{+}06, BJKF05, HDS^{+}07, LeB05-76, LeB08f, LLC^{+}07, MAG^{+}08, MBF^{+}07, MBF^{+}08, OIFN^{+}07, RMC^{+}05, TMH^{+}07, AH05, CCS^{+}08, JET^{+}07, PL06]. phenotype [OMLP^{+}08, YMD^{+}07]. phenotypes [BJM^{+}06]. Phenylbutyrate [GDF^{+}05]. pheromone [JMG08, PKM^{+}05, TJC06, YR09]. pheromone- [PKM^{+}05]. phosphatase [BCR^{+}08, CRLP07, CWRG^{+}08, ELJ^{+}08, GDW^{+}07, GQSHO05, JhKC08, KTA^{+}06, LLW^{+}09, MHC^{+}09, PPK08, QU08, SIL^{+}07, SWM^{+}06, ToMY^{+}08, WHB^{+}07, WWC^{+}07]. phosphatase-dependent [GQSHO05]. phosphatases [SRL05]. Phosphate [LeB06-56, AWM^{+}08, FOB^{+}09, LPA^{+}07, LBC^{+}07, MES^{+}08, SRL05, SDB^{+}05, SWH^{+}06, SLWA07]. phosphatidic [RCM^{+}09]. phosphatidylcholine [TPA^{+}08]. Phosphatidylethanolamine [BGZ^{+}07, OHP^{+}09]. Phosphatidylinositol [JKKM08, SDB^{+}05, AWM^{+}08, CHCB05, CTY^{+}06, FOB^{+}09, LPA^{+}07, LBC^{+}07, MES^{+}08, SHDT05, SLWA07, VTRB06, VCN^{+}05, vZPG^{+}07]. Phosphatidylinositol-4 [SDB^{+}05]. phosphatidylserine [YHD^{+}09]. phosphodestruction [MFS^{+}09]. phosphodiesterases [TDP^{+}06]. Phosphoinositide [JD06, FBD^{+}08, LFL^{+}05, MCF^{+}05, SHC^{+}05]. phosphoinositides [CTY^{+}06, RIHP06]. Phospholipase [HLK^{+}09, KGKV07, LJR^{+}05, SFEK07, vKGK07]. phospholipid [BYMB06, LeB05-97]. phospholipids [GG09, ZOH^{+}08]. phosphoprotein [MSLS07]. Phosphoregulation [JBYW06, ZSMW09]. phosphorylated [GJL^{+}08, MM05]. phosphorylates [LCO07, LB08, TAK^{+}09, YOS^{+}05]. phosphorylating [STMG05]. Phosphorylation [HHZ^{+}08, HF06, LeB08u, LBI^{+}08, MBJF^{+}08, SNG^{+}05, SMRS^{+}07, TSP^{+}07, WBT^{+}08, YGY^{+}05, AH05, BSMT05, BLT^{+}09a, BWJ05, BSL^{+}08, CTS^{+}06a, CM06, CMF08, DXC^{+}08, FHO^{+}05, FHS^{+}07b, GHX^{+}07, HLL^{+}08, HTYNK09, KKW^{+}09, KNP^{+}05, KO06, KLG^{+}09, Les08c, LVDR05, MWMS07, MS09, MTB^{+}09, NEF^{+}06, NWB^{+}06, RP05, SRL^{+}06, SMT^{+}08, TÖL^{+}06b, WF07, YLM06a, dCD^{+}07]. phosphorylation-dependent [FHO^{+}05]. phosphorylation-related [KO06]. phosphorylations [FHNG05]. phosphotyrosine [LSdWL07]. phosphotyrosine-independent [LSdWL07]. photoreceptor [BLP^{+}05, tDAK^{+}05]. photoreceptors [BHY^{+}08, COTAE^{+}05, EFG^{+}05, FSN^{+}06, LSR07, RCM^{+}09]. photosynthetic [SMWdV09]. phototransduction [LJR^{+}05, WJM07]. physical [CRV^{+}06]. physically [NMO^{+}07, SMT^{+}08]. physics [Wel06-43]. Physiological [VP05, LLS^{+}09]. physiologically [GDA^{+}08]. pi [LTK09, GC05, JHY^{+}08, LCT^{+}09, MPS^{+}07, PWW^{+}09, PST^{+}06, SMH^{+}09]. PI-3K [LCT^{+}09]. PI-4K [PST^{+}06]. pi-bodies [LTK09]. PI3 [DTW06, KGKV07, ST09, vKGK07]. PI3-kinase [KGKV07, ST09, vKGK07]. PI3K [SJL^{+}07, KRZ^{+}08]. PI4P [iYOW^{+}06]. PI4P-signaling [iYOW^{+}06]. Piccolo [LOWTL^{+}08]. Pick [CC09]. picture [Sed09j]. PIDDosome [MKB^{+}09]. PIE [DPD^{+}09, Wil09n]. PIE-1 [DPD^{+}09, Wil09n]. piece [Les07-51]. Piecing [Rex09]. pigments [LeB07-56]. Pih1 [ZKG^{+}08]. Pih1/Nop17 [ZKG^{+}08]. Pik1 [SHDT05]. PIKfyve [Pow09b, TGRD09]. Pilings [Wel06-44]. pilot [JKW^{+}09]. Pin1 [SLMK08]. pinch [Nyb05h]. Pinched [LeB07-43]. PIP [HHFC06, SKO^{+}06, Wil07-47, vRSvR^{+}07]. PIP2 [HLK^{+}09]. PIP3 [KGDR^{+}07, VWK^{+}06]. PIP3-dependent [KGDR^{+}07]. PIP5K [MYZ^{+}09]. PIP5K- [MYZ^{+}09]. PIPKI [LVL^{+}05]. PIPs [LeB05k]. piRNAs [LTK09]. pitfalls [Nor06]. pits [Pow05f]. Pitx2 [SYA^{+}08]. pivotal [PUG^{+}06]. PIX [NWB^{+}06, LeB06-84, tKJCH06]. PK [UWiY^{+}07]. PKA [LLHH07, BSMT05, LeB07-44, LLL^{+}06, SBM^{+}07]. PKA-activated [LLL^{+}06]. PKA-catalyzed [BSMT05]. PKC [AM05b, BZHA^{+}08, ERM^{+}09b, EMMN^{+}05, GHX^{+}07, LeB06-57, MWMS07, SMRS^{+}07, WWX^{+}06, vKST^{+}07, RSL06]. PKD [AM05b, BBPM07]. PKD-mediated [AM05b]. PKD2 [HDM^{+}07]. Place [Rob08-45, Les07-54, Les08-40, Sho09-48, Wil07-39, Wil07-53]. places [LeB05-67, Pow05b]. Plakophilin [BZHA^{+}08]. plane [BBvD05, RW05]. planes [Tum05m]. Plant [LeB07-45, GH05, LeB07-49]. plants [Mat05]. plaque [Rob07g]. Plasma [GMS^{+}08, HRLH08, LPG^{+}07, LeB08v, PKB^{+}06, BSA^{+}08, BFF^{+}06, CNMM^{+}06, DPMB^{+}07, DDB^{+}05, FG08, GCF^{+}05, HWSJ08, ITG^{+}08, IGFG^{+}08, JMG08, JWSS05, KDV^{+}09, LGDN09, MSR09, MFM^{+}06, NYI^{+}05b, OLLG05, RIHP06, Tum06p, WHOS06, WZCB^{+}06, WBLL06, LWBL06]. Plasmid [MJG07, CM07, CGJ09, HGJ06, Les09-28]. plasmids [CM07, LeB07-40]. plasminogen [SBM^{+}07]. Plasmodium [BAdSG05]. plasticity [BJML09, SGMPV07]. plastid [IBS^{+}05]. plate [BLT09b]. Platelet [LeB07-46, ASHD^{+}05, BSK07, SZX08, WBP^{+}08]. platelet-derived [BSK07]. plates [LRJY06]. platform [BVGA^{+}05, GSH^{+}08, Sco08]. platforms [LZ08]. play [FSnP^{+}06, Les06p]. Playing [Di 08]. plays [AFN^{+}08, JD06, LYH^{+}05, LGBD^{+}07, LRCH^{+}05, Pow07a, RKH^{+}09]. Plectin [OMGW^{+}06, RKN^{+}07, WAW08, KFR^{+}08, WLK^{+}05]. Plectin-controlled [OMGW^{+}06]. Pleiotropic [XSVGR05]. Plexin [OKN06]. PlexinB1 [GMM^{+}08]. Plk1 [SCD^{+}08, WF07, YGY^{+}05]. Plk1- [SCD^{+}08]. Plk1-mediated [WF07, YGY^{+}05]. Plk4 [RRR^{+}09]. Plk4/Sak [RRR^{+}09]. Ploegh [LeB07t]. plot [KC09]. pluripotency [SAD^{+}09]. plus [BVS^{+}09, DvCvH^{+}08, GHM^{+}08, LeB06-35, LeB08y, MKGL^{+}05, MSB06, PTF^{+}06, SRE^{+}05, Vau05, WTH05, XMA^{+}09]. plus-end [MKGL^{+}05]. Plx1 [WF05]. PMC [Hil07]. PML [DMBF^{+}07, JMM^{+}06, JZN^{+}09, Wel05-51]. PML-dependent [DMBF^{+}07]. Png1 [KAL^{+}06]. PNS [WTE^{+}08]. Poc1 [POG^{+}09]. pocket [MXR^{+}05]. podocalyxin [MSSF05]. podosome [GHDS^{+}07, OIT08, SYKM08, WPC^{+}08]. Podosomes [Wil07-48]. Point [Wil06l, BMR^{+}08, ELM09, JBF^{+}08, RG06a, YWRA07]. poise [iWK06]. Pol [LeB06-46, Les07-51, MWY^{+}09]. polarity [ALB^{+}09, BFP^{+}07, DP08, DCEM09, FKN^{+}07, GMvdKC07, HP05, HOR08, HHFC06, KGKV07, KNT^{+}06, KWY^{+}06, LPA^{+}07, LeB06-85, LWHC08, MRO^{+}05, MSK^{+}07, MS07, NN07a, OFfL^{+}07, PB05, QCGM05, RAP08, SMH07, SLM^{+}06, SSM05, SHH^{+}08, Tum05w, TMY^{+}09, VWK^{+}06, VMZW^{+}07a, VMZW^{+}07b, Wil07-49, YGW06]. polarity/proton [Wil07-49]. polarization [EMMN^{+}05, GMvdKC07, GGEd^{+}07, HCFM08, IUT^{+}07, MSSF05, OFfL^{+}07, PKB^{+}06, RWM^{+}05, SIS^{+}08, TSK^{+}06, WNB08]. polarized [API^{+}07, FHMC09, FSP^{+}07, HXZ^{+}07, HSTM06, LFS^{+}05, NFCF08, PPCZ06, PST^{+}06, RYN05, RKH^{+}09, STM^{+}07, ShCM^{+}06, VVMS05, YWJ^{+}07, YKP^{+}08, ZWG^{+}05]. polarizes [VCN^{+}05]. Polarizing [LeB08w, Les07-52, WVB07]. pole [FDV^{+}06, JMG^{+}06, LSC^{+}06b, NdCF^{+}05, TTYH09, Tum05-42]. poles [GNV05, JWC^{+}08, SHC09, TKKT07]. Poleward [UHM^{+}08, CYC^{+}06]. policy [Ros05]. poliovirus [Wel06r]. pollen [GFD^{+}05, THL^{+}06]. Polo [dGHW08, RMAA08]. Polo-like [dGHW08, RMAA08]. Polycomb [AMERT^{+}07]. polyglutamine [CSSvM08, JMM^{+}06, QIH^{+}06]. polyglutamine-containing [CSSvM08]. polyglutamylated [SHH^{+}08]. Polymerase [LeB05-78, CVACF07, LeB05-68, Les07-59, Les07-58, Les08-44, LGBD^{+}07, MRL^{+}08, PNB07, YM09]. polymerase- [YM09]. polymeric [MCA^{+}06]. polymerization [CAR^{+}08, CHCB05, HLS^{+}09, SKO^{+}06]. polymerization-driven [CHCB05]. polymerize [HdL06]. polymerizing [MRvD^{+}05]. polymodal [KBGG^{+}08]. polypeptide [BYMB06, COC^{+}05, DVF^{+}06]. polyploidy [DSN^{+}07b]. polytopic [KL05]. polyubiquitination [MSM^{+}06, OLC^{+}07]. POM121 [SNPCG06]. POM121- [SNPCG06]. Pom152p [MMCW06]. Pom1p [Wil07-50]. pombe [TTYH09, CBZY08, GCM09, LB08, RGCWG09]. pool [Mel08a]. pools [LeB05-104]. population [HMM^{+}05, MLHS06, ONW^{+}07]. populations [MCA^{+}06]. Pore [Wel05-52, BCC^{+}09, COC^{+}05, DGG^{+}08, DLHW09, DZW^{+}08, FKJ^{+}09, LeB06v, LeB07-47, Les05f, Les07-72, Les09-27, LFH07, LPM^{+}07, MMCW06, MSL^{+}09, NdCF^{+}05, OSM^{+}09, RM05a, Rex09, Rob08-30, SHC07, SHS^{+}06, SNPCG06, TW07, Tum05-42, Wel06-65, WEL^{+}07n]. pores [FSS^{+}09, Tum06e]. Porterplasm [Pow05r]. Portioning [Wil09n]. portrait [Les05f]. POSH [TTE^{+}07]. Position [Rob08-35, DCEM09, JKDH08, LeB08m, MLS^{+}07, Sed09m, Sho09x]. positional [MPL^{+}07a]. Positioning [Wil06m, Wil07-51, BE05, CLG^{+}07, DHT^{+}09, HP05, iHBD08, OLS08, OF08, RKvdK^{+}09, TGH^{+}07]. Positive [FKN^{+}07, ASHD^{+}05, GEN^{+}06, MOCC05]. positively [EHHK09, WLC08]. possess [KAD^{+}08]. post [HNvD^{+}07, HHY^{+}08, WRP^{+}06, HFP^{+}05]. post-Golgi [HNvD^{+}07, WRP^{+}06]. post-transcriptionally [HHY^{+}08]. posterior [KO07, TA07]. postmitotic [HZR05, OW06]. postnatal [GKCM^{+}08, JLG^{+}06, WMSB^{+}06]. Postsynaptic [BJML09, KTN07, NVJ^{+}08, OWM05, PFD06, SLNS06]. posttranscriptional [LTK09]. posttranslational [MLC^{+}07, LeB07u]. pot [Wil08x]. potent [LVG^{+}09]. potential [BM07, ONW^{+}07, SCF^{+}07, WJM05]. potential-dependent [BM07]. potentially [SHC07]. potentiate [DG08]. potentiation [RCS^{+}09]. power [LeB05-51, Les07-52, YOY08]. Powered [Les05g]. powers [KRPH08, MT07, Pow05q]. pox [Rob08-48]. poxvirus [Heu05, SWL^{+}05]. PP1 [TMAL^{+}06]. PP2A [CRLP07, YK07]. PP2C [PGO^{+}07]. pp32 [HAT^{+}05b, MDC^{+}08]. pp32/LANP [HAT^{+}05b]. pp32/PHAP [MDC^{+}08]. PPAR [CTP^{+}09, Sho09-52, vKST^{+}07]. PPH [KHW^{+}07, KHW^{+}09]. PPH-4.1-mediated [KHW^{+}07, KHW^{+}09]. PPM1G [PGO^{+}07]. PPM1G/PP2C [PGO^{+}07]. PpPex20p [LZM^{+}06]. PPXY [MSECB05]. PR [TMH^{+}07]. PR-Set7-dependent [TMH^{+}07]. practical [Nor06]. Prague [LeB06-52]. pRB [MYO^{+}05]. Prc1 [CGGT09, LW07b]. Prc1-dependent [CGGT09]. Pre [HNvD^{+}07, BJZ^{+}09, KRPH08, LSD^{+}06, YOSM06]. Pre- [HNvD^{+}07]. pre-60S [KRPH08, LSD^{+}06]. pre-messenger [BJZ^{+}09]. pre-mRNA [YOSM06]. preadipocytes [LLS^{+}09]. preapical [MSSF05]. preautophagosomal [HSY^{+}06]. precede [BLMS06]. precedes [LSM^{+}05]. precise [Pow07c]. precision [Wat09]. precocious [KGL07]. precursor [MM05, Rap05, SKO^{+}08, SMM^{+}08c, WKWL06]. precursors [DCAC05, MMU^{+}09, SDL^{+}05, ZWO^{+}09]. predicted [CCSBM^{+}07]. preferences [CAL^{+}07]. preferential [SRBL06]. preferentially [HLM^{+}08]. prefibrillar [KKB^{+}06]. preformed [BZYM^{+}08]. pregnancy [Wel06x]. PRELP [RRA^{+}09]. premature [EVF^{+}08, MGW^{+}08, MNF^{+}07, SMW^{+}07]. premature-aging [EVF^{+}08]. preprotein [WAKS08]. prerequisite [ZEC^{+}08]. presence [MSK05]. presenilin [GDA^{+}08]. present [GD05, SORM05]. presentation [VLFA^{+}07]. Presenting [LeB05-79]. presequence [WvdLH^{+}07]. preservation [Pow09b]. preset [LeB07-46]. presomitic [YKP^{+}08]. pressure [LeB05i, LeB05-95, Tum05g]. Presynaptic [HCRT08, AACS^{+}06, NLS^{+}05, OMF^{+}05]. prevent [JMM^{+}06, KL05, LeB05-27, PCG^{+}05]. Preventing [Les09-29, DDLC^{+}06, KGL07, RKK09, SMW^{+}07]. prevents [CRLP07, DP08, EFG^{+}05, HZL^{+}07, LeB07c, LLC^{+}07, MGW^{+}08, MRD^{+}08, Rob08a, RMKS08, SMSPK08, SMRS^{+}07, TGRD09, Wil07-50, WCR^{+}08]. price [Les07-53]. Primary [PS08, DMC^{+}07, GSL^{+}07, KAD^{+}08, MZH^{+}08, SSD^{+}09, STM^{+}07, SLP^{+}06, WRKC^{+}09, iYEF^{+}07]. primate [DHML05]. prime [Les08-66]. primed [HCC^{+}07]. priming [DXC^{+}09]. primitive [LeB05-105, LeB06-88, SRXvdK06]. Prion [KZT^{+}08, KZT^{+}09, SSLS05, Ste08]. Prions [Les08-37, Wil07-46]. PRISM [Ros07a]. Prm1 [HEW07]. pro [FMD^{+}09, LeB06-58, LeB07-48, YJJ^{+}09, BSF^{+}08]. pro- [FMD^{+}09]. Pro-BDNF-induced [YJJ^{+}09]. Pro-diversity [LeB07-48]. Probing [Wil07-31]. problem [Rob07g, Rob08-41]. procedures [Sho09q]. proceed [LCvD^{+}08]. Process [LGZ^{+}05]. processes [CGS05]. Processing [LJN^{+}05, DTP07, ERM^{+}09a, GAS06, KSA^{+}05, LZL^{+}05, LWG^{+}05, MLS^{+}07, SKO^{+}08, SNC^{+}06, SCF^{+}07]. processive [KHWC08, SOR^{+}05]. produce [KHWC08]. produces [BPL^{+}05]. production [BKE^{+}07, GDA^{+}08, PRZ^{+}08, PS08, Rou08]. productive [LeB08p]. products [Pow05j]. profiles [IGFG^{+}08]. Progenitor [WMM^{+}09, AOP^{+}05, DBvdW^{+}06, FKS^{+}06, GSY^{+}09, JAV^{+}07, KWS^{+}08, NTB^{+}06, NPB^{+}05, QWD^{+}08, RMZ^{+}06, SRR^{+}08, SLCB05, VWM^{+}07]. progenitors [DMC^{+}07, VMY^{+}09]. progeria [MG08b]. program [FAG^{+}07, GLS^{+}09, HLK^{+}07, LeB06-30, Sho09-49]. programmed [GH05, MDH^{+}06, PKM^{+}05, THL^{+}06]. programming [RCKD06]. programs [BFL^{+}05, RLH^{+}09, SSHW09, SLZ^{+}06]. Progranulin [VVL^{+}08]. progression [CGLC^{+}08, CWC^{+}07, CBIO05, DVF^{+}06, DSBK09, HLK^{+}07, JET^{+}07, KST^{+}07, KRPH08, LM05, MVC^{+}07, MFO09, RMC^{+}05, SCD^{+}08, SCMV07, SiTT^{+}08, TWC^{+}07, ULN^{+}07, WWK^{+}06]. progressive [HlQY^{+}06, PRB^{+}09]. Prohibitin [SSS^{+}08a]. Prohibitin-1 [SSS^{+}08a]. prohibitins [OHP^{+}09]. proinflammatory [LYY08]. proliferating [MSM^{+}06]. Proliferation [LG07, LG08, BDP^{+}07, CZSG08, CTS^{+}06c, DRD^{+}05, FKS^{+}06, FIK^{+}05, FIK^{+}20, HRS^{+}05, JAV^{+}07, LeB06u, LeB08g, Les07-41, LNPC06, MWMS07, MRD^{+}08, NPB^{+}05, OLS08, PLR^{+}06, RVBZ^{+}08, SRR^{+}08, SMW^{+}07, TTK^{+}06, Tum06s, WMM^{+}09, WBC^{+}06, WWX^{+}06, WM05, Wil07-28, XSVGR05, YWW^{+}08]. Proliferation-dependent [LG07, LG08]. proliferative [PHB^{+}07]. proliferator [GDF^{+}05]. Proline [MNH05, LVL^{+}05]. proline-directed [LVL^{+}05]. prominin [DMC^{+}07]. prominin-1 [DMC^{+}07]. promote [BWL^{+}08, BYR08, BMP08, CCL^{+}08, CYY^{+}08, CLG^{+}07, FBY^{+}06, GHY^{+}05, GC05, GNL^{+}06, KHH08, KBC^{+}07b, LMM^{+}09, MLC^{+}07, NEF^{+}06, NVJ^{+}08, SKE^{+}05, SMH07, STU^{+}08, SWI06, WZM07, WF06, YMCT08]. promoter [BZA^{+}06, LN07]. Promoters [LeB07-49, Les07b]. promotes [BSO^{+}08, BHG^{+}05, CTS^{+}05, CTO^{+}08, DLL^{+}08, DDO^{+}08, GdTAR07, GGO^{+}09, GSiK^{+}09, HGJ06, HFY^{+}06, JMG08, JHY^{+}08, KLS^{+}06, KOKM07, LM05, Les08k, LHZ^{+}08, LFW^{+}09a, LFW^{+}09b, MWY^{+}09, MES^{+}08, MAG^{+}08, NTS^{+}05, NTSY08, NdCF^{+}05, NBS^{+}08, Pow09b, RLDC09, SYKM08, SG09, SZL^{+}09, SMM08b, SLNS06, TTYH09, TLDB08, TP09, TSN^{+}07, TMI07, YIfHA09, YHC^{+}06, ZCJ^{+}08, ZMH^{+}09]. Promoting [LeB06-59, BLT^{+}09a, CAR^{+}08, CNW^{+}05, DFB^{+}07, GGBV^{+}07, NSO^{+}09, PRZ^{+}08, SMW^{+}07, WC09, WWvD07]. Promotion [FHO^{+}05, NPHT08]. prompt [FPP^{+}08, Wil07-46]. prompts [IDTT^{+}09, Wil07-28]. Promyelocytic [DCA^{+}06, QIH^{+}06]. Propagation [JBFC07, EMB^{+}08]. Proper [GCM09, KCH^{+}06, LR07, SB06, WF06]. properties [DG08, HMR^{+}07, MSH^{+}05]. Prophase [MVC^{+}07, DSK^{+}06, FNH^{+}05, KSL^{+}05, KPS^{+}07, NVC05]. prosecretory [MBR^{+}06]. prostate [BPL^{+}05, Tum05-55]. prostatic [SBC^{+}05]. prosurvival [MGH^{+}09, UDC^{+}07]. Protease [SSHW09, ERM^{+}09a, HGA^{+}09, HEW07, LFH07, LCR^{+}05, PCV^{+}08, PXD^{+}07, VDA^{+}08, WMT07]. protease-activated [WMT07]. Protease-dependent [SSHW09]. proteases [Sho09-41, YWM^{+}08]. proteasomal [DGSN06]. proteasome [CPAH08, EHWL06, LSDFS08, Rob07c, TAK^{+}09, WWO^{+}08, YGW06]. proteasome-independent [CPAH08, EHWL06]. Proteasomes [Wil07-52, Wil06m]. protect [CWM^{+}09, MLC^{+}07, Rob08j, Wil07l, YK07]. Protection [BAR^{+}08, WZC^{+}05]. protective [BLB^{+}05]. protects [AJR^{+}09, DC07, DGP^{+}05, LeB05o, WCB^{+}08]. Protein [BFB^{+}08, FSnP^{+}06, LeB07-50, Les08r, SFP^{+}05, SWM^{+}06, ToMY^{+}08, AMERT^{+}07, AFN^{+}08, AHM^{+}05, AHD^{+}09, BY09, BGZ^{+}07, BBF^{+}07, BYK08, BFP^{+}07, BBFN07, BLB^{+}05, BvOM^{+}07, BAR^{+}08, BZA^{+}06, BXHD08, BPL^{+}08, BSH^{+}07, BMSJ07, CZM^{+}07, CBK^{+}06, CCS^{+}06, CCL^{+}08, CZD^{+}09, CHH^{+}08, CSSvM08, CSW08, CRLP07, CCC^{+}06, COB^{+}08, CHO^{+}09b, CWM^{+}09, CBG^{+}06, CLG^{+}07, CFBM05, Dal09, DvCvH^{+}08, ELJ^{+}08, EHWL06, EYB^{+}06, FFT^{+}05, FHO^{+}05, FBD^{+}05b, FSR^{+}07, FFS06, FTM^{+}06, FNH^{+}05, FHS^{+}07b, GHX^{+}07, GGO^{+}09, GGB^{+}08, GRGP07, GHM^{+}08, GG06, GBNK08a, GBNK08b, GTX^{+}06, GWS09, GD05, GMZ^{+}08, GSL^{+}07, GC07, GGC05, GSV07, GKvdB07, GFC^{+}08, GL06, HHT^{+}07, HRLH08, HTK^{+}08, HMR^{+}07, HK07, HHS^{+}09, HHFC06, HQF^{+}07, HCB^{+}08, HZOL05, HF06, HOZ^{+}09, IBS^{+}05, JCMG08, JKW^{+}09, JASC08, JFM^{+}09, JWSS05, JhKC08, KKGD05, KII^{+}08, KBN^{+}06, KC09, KZT^{+}08, KZT^{+}09]. protein [KMS07, KHBV05, KPR^{+}07a, KPR^{+}07b, KTA^{+}06, KTI^{+}05, KFK^{+}05, KNT^{+}07, KGM^{+}07, KRG^{+}08, LDS^{+}06, LJA09, LeB05-78, LeB07y, LW07a, LGZ^{+}05, LVL^{+}05, Les06x, Les07-41, Les08b, Les08h, Les08-28, Lev07, LSC^{+}06a, LFH07, LHL^{+}08, LS06, LDGO^{+}09, LLH^{+}09, LFMO^{+}09, LCL^{+}07, LNL^{+}08, LB08, LDL^{+}09b, MTL^{+}06, MHM^{+}07, MLC^{+}07, MYMH09, MSECB05, MF08, MDET08, MGH^{+}09, MW06, MGR^{+}05, MTH^{+}06, MHC^{+}09, MDPG07, MHFH06, MLP06, MBF^{+}07, MBF^{+}08, MM05, MDW05, NMV^{+}07, NYM^{+}08, NSS^{+}08, NS06, NRL^{+}05, NdCF^{+}05, OOY^{+}06, OVL^{+}07, OOS07, OOS09, PPK08, PUG^{+}06, PLLH05, PAM^{+}07, PT09, dJPAW^{+}09, PSS^{+}05, PPS^{+}06, PMK^{+}07, PEW^{+}05, PST^{+}06, PHM^{+}09, QCGM05, QIH^{+}06, RIHP06, RM05b, Rob08-41, RKvdK^{+}09, RMC^{+}05, RLDC09, RKB^{+}07, SRX^{+}07, SON06, SKO^{+}08, SSHP^{+}07, SSLS05, SJL^{+}07, SHC07, SGB^{+}06, SPT^{+}09, SPT^{+}21, SCR^{+}09, SDI^{+}06, oSCW07, SWT08, SNA^{+}07]. protein [SDD^{+}07, Ska07, SIL^{+}07, ShCM^{+}06, SSSL05, SGDM06, Ste08, SDW^{+}09, SMM^{+}08c, SIS^{+}08, SWL^{+}05, SRQL09, TKB^{+}06, THY^{+}06, THS^{+}09, TCHL05, TGH^{+}07, TSK^{+}06, TSP^{+}07, TMBL^{+}08, TWC^{+}07, VTT^{+}09a, VDSV06, VLFA^{+}07, VSR^{+}08, VKR^{+}09, VCN^{+}05, WAKS08, WRKC^{+}09, Wel05-32, WC07b, WLK^{+}05, Wil08o, WHM^{+}07, WWC^{+}07, YMTC05, YOY08, YHD^{+}09, YOSM06, YWM^{+}08, ZPGB^{+}07, ZMGC05, vRCB^{+}07, vdBKB^{+}07, UPBM06]. protein-1 [MM05]. protein-dependent [BvOM^{+}07]. protein-independent [SJL^{+}07]. protein-mediated [TGH^{+}07]. protein-sorting [JWSS05, MSECB05]. Proteins [LeB07-51, Les08-38, AVHH09a, AVHH09b, BHY^{+}08, BLC^{+}09, BJZ^{+}09, Bon05d, BEN^{+}05, BLA^{+}09, CCS^{+}06, CMF^{+}09, CDW^{+}05, CTY^{+}06, CYO^{+}09, CK07, CNMM^{+}06, DLHW09, DR07, DJV^{+}05, ESZK06, FHMC09, FSnP^{+}06, FHS^{+}07a, GG09, GNV05, GMS^{+}08, HWN^{+}07, HSLI09, HLK^{+}09, HT08, HSTM06, JMM^{+}06, JZN^{+}09, KCE^{+}06, KBI^{+}08, Kat06g, KIKI^{+}07, KSL^{+}05, KDG^{+}09, KL05, KSpLL06, KZS07, LeB05-41, LJS^{+}09a, LW07b, Les06w, Les08-46, Les08-66, Les09v, LK05, LBG^{+}06, MBR^{+}05, MEG^{+}07, MG08a, MNH05, MGH^{+}09, Nic09a, OTH^{+}07, PPCZ06, PTF^{+}06, Pow05v, Pow06b, Rap05, RIHP06, RMAA08, Rob08m, RCFB08, RAP08, SKE^{+}05, SWD^{+}08, Sav09a, SS06a, SS08a, Sed09b, Sed09c, SZL^{+}09, SAM^{+}07, Sho09g, SGKM08, SNG^{+}05, SGKP^{+}07a, SGKP^{+}07b, THS^{+}09, TSS^{+}06, Tum05z, Tum05-30, Tum06-37, WHOS06, WPH^{+}07, WF05, ZMGC05, ZWG^{+}05]. proteoglycan [MWMS07]. proteoglycans [OBY^{+}06]. Proteolysis [ZMA^{+}07, FBJB^{+}09, LeB05d, MCD^{+}07, NMO^{+}07, OHH^{+}06, SIL^{+}07]. proteolysis-resistant [MCD^{+}07]. Proteolytic [CYY^{+}08, CSSvM08, GKvdB07, HGA^{+}09, HOH^{+}07, HCL^{+}07]. proteomes [TMBL^{+}08]. Proteomic [PALW05, TDD^{+}05]. proteomics [BHH^{+}06, TTG^{+}09, Tum05h]. prothymosin [UFY^{+}07]. prothymosin- [UFY^{+}07]. protist [KBC^{+}07a]. proto [EIK^{+}05]. proto-oncoprotein [EIK^{+}05]. protocadherin [CG06, NPHT08]. protofilaments [MGBM^{+}09]. proton [Rob08-36, Wil07-49]. protrusion [LMM^{+}09, NWB^{+}06, VMZW^{+}07a, VMZW^{+}07b, ViKA^{+}06]. protrusions [SWD^{+}08]. protrusive [PJMR06]. provide [BW08, BBM^{+}07, GHM^{+}08, Sco08]. provides [GSH^{+}08, KBW^{+}06, MFA^{+}07, YTB07]. proximal [SBC^{+}05, SMH^{+}09]. proximity [BLR^{+}06]. Prp20p [DTR^{+}05]. PRP4 [MDPG07]. Prss8 [LCR^{+}05]. PSD [NBS^{+}08, NFI^{+}09, Wil08r]. PSD-95 [NBS^{+}08, NFI^{+}09, Wil08r]. PSD95 [CVD^{+}05]. Pseudokinase [LeB08x]. Pseudomonas [KGDR^{+}07]. Pseudopodia [DDH09]. PtdIns [BSA^{+}08, SCK^{+}07, WSB^{+}09]. PtdIns4 [WSB^{+}09]. PTEN [DDO^{+}08, Sed08k]. PTEN-null [DDO^{+}08]. PtK1 [CYC^{+}06]. PTP [ASHD^{+}05, WKC^{+}09]. PTP-1B [ASHD^{+}05]. PTP1 [SWM^{+}06]. PTP1B [CCP^{+}08]. public [Ros05]. publications [Ros05]. PubMed [Hil07]. Puf3p [GRGP07]. pull [Les07-46, Les08s, Les09y, Rob08s, Wel05-34]. pulling [KO07]. pulling-force [KO07]. pulls [LeB07o, Wil07-59]. pullyu [Las09]. pulmonary [dJPAW^{+}09]. pulse [LeB06z]. Puma [Let09, GGO^{+}09]. Pumilio [GRGP07]. pump [GDA^{+}08]. Pumping [Les08-39]. punch [Tum06w]. punctuation [Kat06f]. puncture [Les08j]. Pursuing [Pow05s]. push [Wil07-48]. pushes [LeB07-40, Pow05b]. Pushing [VW08, CM07, TWG^{+}09]. pushmi [Las09]. pushmi-pullyu [Las09]. put [Tum06l, Wil09g]. puts [Les08y, Sed09a, Wel06-62, Wil07-39]. Putting [LeB06-60, Les07-54, Les08-40, Les09-30, Nyb05h, PT07, RAP08, Wil07-53]. pXBP1 [YOSM06]. Pygo2 [GSY^{+}09, Hor09, Wil09o]. PyK2 [LLT^{+}08, GHDS^{+}07, LeB07-52, WLLF^{+}08]. pylori [GMG^{+}07].

Q [BKE^{+}07, MBS^{+}08, OV09]. Qing [Sed08m]. quality [MLC^{+}07, PGWH08]. Quantitative [CCSBM^{+}07, GBNK08a, GBNK08b, HKV^{+}05, LJS^{+}09b, SH05, TTG^{+}09, BMBD^{+}09, HSTM06, TMBL^{+}08, Wat09]. question [Sho09e]. Quick [LeB05-80, LeB08y, Rob08-36, Tum05-65]. quiescence [LeB05-58]. quiescent [ABA^{+}06, CGLC^{+}08, JTN^{+}05, LSDFS08, MMU^{+}09]. quiet [Ste08].

R [GPJ^{+}06, OKN06, WP08]. R2TP [ZKG^{+}08]. R7BP [DDB^{+}05]. Rab [CCL^{+}08, Les08-41, Tum06-37, SHL^{+}09, BPL^{+}08, FHS^{+}07a, GAG^{+}06, RCKD06, SGKM08, SHB^{+}07, iYEF^{+}07]. RAB-2 [SHL^{+}09]. Rab-coupling [CCL^{+}08]. Rab11 [CAR^{+}08, LSR07, PWW^{+}09]. Rab13 [NFCF08]. Rab2 [ECR^{+}09, MYMZ08, Sho09-49]. Rab21 [PAV^{+}06]. Rab32 [WRP^{+}06]. Rab38 [WRP^{+}06]. rab5 [SSM^{+}08, ADO07a, ADO07b, MES^{+}08, PSL^{+}06, RvVM^{+}08, SHC^{+}05]. Rab7 [JRZ^{+}07, RKvdK^{+}09, RvVM^{+}08, SCM^{+}07]. Rab9 [ASP06]. RabA4b [PST^{+}06]. Rabs [Mar06]. Rac [Wel06-45, Cha07, GPJ^{+}06, HOM^{+}05, KBC^{+}08, KNT^{+}06, MRO^{+}05, OLLG05, PEER^{+}05, SMM08b, SKO^{+}06, Tum05-43, Tum06b, YN07]. Rac-dependent [KBC^{+}08]. rac1 [BTP^{+}07, KSK^{+}06, AS06, BRM^{+}07, DBR09, ERM^{+}09b, EYB^{+}06, KBN^{+}06, LeB06-84, Les08-42, LNPC06, MAG^{+}08, NZQ^{+}07, dJPAW^{+}09, PEW^{+}05, RKK09, SMG07, WWS05b, tKJCH06]. Rac1-mediated [DBR09]. Rac2 [SMG07]. RACK1 [GHX^{+}07]. Rad23 [KAL^{+}06]. RAD50 [SMT^{+}08]. RAD51C [BLT^{+}09a, KPS^{+}07]. Rad53 [ShCM^{+}06]. radar [Pow05d]. radial [DS07, NZQ^{+}07, PBB^{+}09, YOY08]. radiation [LYWK08, NMV^{+}07]. radicals [MG08a]. Rae1 [BJM^{+}06]. RAF [WM05, EPV^{+}05, GKCM^{+}08, NSM^{+}09, PMK^{+}05, Sho09-50, Tum05-52]. Raf-1 [EPV^{+}05, NSM^{+}09, PMK^{+}05, Sho09-50, Tum05-52]. Raf1 [YTY^{+}07]. raft [Pow06i]. rafts [LGDN09, Les09t, SSLS05, Sho09-32]. raises [LeB07x]. Ral [LH05]. RalBP1 [GPJ^{+}06]. RAM [JBYW06]. RAMP4 [Poo09]. Ran [HM08, TTO^{+}08]. RanBP3 [HFvdV^{+}05]. RanBP9 [VBG^{+}08]. RanBPM [DL08]. Rando [Wil07-45]. Random [Rob08-37, PEER^{+}05, Rob07e]. randomly [Wel05j]. range [MSK05]. RanGTP [DPP^{+}07, YGR^{+}08, YRSM^{+}09]. RanGTP-dependent [YGR^{+}08, YRSM^{+}09]. RanGTP-independent [DPP^{+}07]. RanGTPase [MK07]. Ranvier [DZG^{+}07, YOHR07, ZTM^{+}08]. Rap1 [GMvdKC07, GHSG07, HSB^{+}07, JLM^{+}07, JLL^{+}07, NNS^{+}08]. Rap1- [GMvdKC07]. Rap1-PDZ-GEF1 [HSB^{+}07]. Rape [Wel07i]. RapGAP1 [JLL^{+}07]. Rapid [BKBM06, HSLI09, ITG^{+}08, MAG05, SWT08, SSK^{+}09]. Rapidly [VTRB06, MCA^{+}06, Pow05p, SCS^{+}05]. Raptor [VKR^{+}09]. Ras [GPJ^{+}06, GDR^{+}05, HFF07b, KWS^{+}08, LeB05c, LWG^{+}05, MBS^{+}08, OKN06, Pow06g, RKB^{+}07, SJL^{+}07, YTY^{+}07]. Ras-like [RKB^{+}07]. Ras/PI3K/F [SJL^{+}07]. Ras/PI3K/F-actin [SJL^{+}07]. RASAL [LWG^{+}05]. RasB [MBS^{+}08]. RasGEF [HFF07b, MBS^{+}08]. rat [LMZ^{+}05, SLCB05]. rate [AH08, FN06, RM05a, RBC05, VWM^{+}07, YGR^{+}08]. rates [TTNA^{+}06]. rather [GSdBS09]. Rb [LeB08z]. RCC1 [HM08]. RCK [BAR^{+}08]. rDNA [Rob08-38]. rDNAs [NNK^{+}08]. Reach [Rob08-39, LeB05-93, Les08-28]. reaches [Les07-32, Les08-29]. Reaching [LLR^{+}05]. reaction [LeB07w]. reactivate [RLH^{+}09]. reactive [BKE^{+}07, LeB07w, Ter06]. readout [Pow07c]. Ready [Wel05-53, Tum05-28]. Real [PH07, WNF^{+}05b, YTH^{+}09, BMB^{+}07]. Real-time [PH07, WNF^{+}05b, YTH^{+}09, BMB^{+}07]. Really [Wel05-54, Pow05y]. rear [Les06j, VMKW^{+}08, YWJ^{+}07]. rearrange [LeB05-59]. rearrangement [BSO^{+}08]. rearrangements [KVS^{+}09]. reason [Les08-46]. Reassembly [CHCM06, DZW^{+}08]. Rebecca [LeB07-53, Sed08n]. rebuffed [Les08-43]. rebuild [Les07d]. RecA [KG05]. recapitulates [YMD^{+}07]. Rec’d [LeB05-81]. reception [oSCW07]. Receptor [ALvdG06, GQSHO05, KP08, AMERT^{+}07, ABSK06, ADC^{+}07, BHJ^{+}05, BMRP07, CNM^{+}05, CYY^{+}08, CSM^{+}06, DSLY^{+}05, EYB^{+}06, FCB^{+}09, FCB^{+}19, GHN^{+}09, GCS^{+}07, HWN^{+}07, HZL^{+}07, HSB^{+}07, HPH^{+}05, JP06, JCMG08, JRZ^{+}07, KZC^{+}08, KVC^{+}09, KTI^{+}05, KLV^{+}08, KBC^{+}07b, KBC^{+}08, LGDN09, LeB05-88, LeB06g, LeB07-54, Les07u, Les07-55, LLM^{+}05, LMC^{+}09, LCM^{+}08, MLC^{+}06, MTL^{+}06, MHM^{+}05, MYZ^{+}09, MYMH09, MF08, MDW05, NBB^{+}09, NEF^{+}06, NYN^{+}06, NNS^{+}08, NRW^{+}08, OFY^{+}05, OWM05, PCR^{+}07, PLLH05, PES^{+}07, Rob08n, RRP^{+}08, SBM^{+}07, SKE^{+}05, SSLS05, SDDV05, SSJ^{+}07, SSD06, oSCW07, SRR^{+}08, SXH^{+}06, SFC^{+}08, Sho09r, SKR^{+}07, Ste08, SWH^{+}06, SMG07, SFEK07, SFS^{+}07, Too05, TSP^{+}07, TLK^{+}06, Tum06-28, VLK^{+}08, VCSG07, WJM05, WBC^{+}06, WAKS08, WNL^{+}09, WCR^{+}05, Wel06-56, WMT07, YWEK08, ZDD^{+}06, ZDMD06, dCD^{+}07, vRCB^{+}07, STP08]. receptor-1 [WMT07]. receptor-dependent [SSD06]. receptor-independent [GCS^{+}07]. receptor-like [HWN^{+}07, PCR^{+}07]. receptor-mediated [MHM^{+}05, MYZ^{+}09]. receptors [BSK07, BZYM^{+}08, CMF^{+}09, FTT^{+}05, HCRT08, KZT^{+}08, KZT^{+}09, KKRP05, KFK^{+}05, LGH^{+}05, LeB07-42, LeB08-28, MGR^{+}05, PRZ^{+}08, PRK^{+}07, PRK^{+}09, RAS^{+}07, RAS^{+}08b, Rob08i, Rob08-54, SWT^{+}05, TSN^{+}07, TDT^{+}08]. Reciprocal [NHK^{+}08, OYTO07]. Recognition [WEL^{+}07n, AMS^{+}05, ABSK06, BAR^{+}05, HZR05, JWG^{+}07, SZX08, WSB^{+}09]. recognizing [BKLT^{+}08]. recombinant [SKE^{+}05]. recombinase [BLA^{+}09, FMD^{+}09]. Recombination [KKGD05, BSO^{+}08, BGJ^{+}07, BFCS^{+}07, BLA^{+}09, JCVB08, LeB05-81, MGW^{+}08, MTV^{+}05, SiTT^{+}08, TNS07, Tum05-67, WH06, Wil07s]. Reconstitution [LS06, CCSK08]. reconstructing [Wil08f]. reconstruction [MFM^{+}06]. recovery [HFY^{+}06, PWD^{+}06, YOY08]. RECQ1 [Les07-56]. recruit [GMSS06, SFS^{+}07, YM09]. recruited [BZYM^{+}08, DMDO08, LDL^{+}09b, NTSY08, YOHR07]. recruiting [MES^{+}08]. Recruitment [AVHH09a, AVHH09b, HSY^{+}06, MRL^{+}08, Wil07-54, AH05, BAR^{+}05, BLMS06, CFL^{+}05, GHM^{+}09, GPC^{+}06, HRB^{+}06, IKZ^{+}07, KKK^{+}05, MGBH^{+}06, MKV^{+}09, OLLG05, OD07, OTM^{+}07, PNB07, PTF^{+}06, Poo09, RvVM^{+}08, SSD06, SRE^{+}05, SCMV07, SFEK07, WBP^{+}08, ZCJ^{+}08]. recruits [GGC05, GSV07, HFF^{+}07a, JWC^{+}08, LeB06r, SSLS05, TMAL^{+}06]. Recycling [ADD^{+}06, BSF^{+}08, CCL^{+}08, GSW^{+}08, KKM^{+}08, Kat07c, LSD^{+}06, LPB^{+}07, MKH^{+}07, NFCF08, PES^{+}07, Pow05h, SSSB07, TMY^{+}09, WCN07]. Red [Rob07m, CBG^{+}06, EVH^{+}07, EVH^{+}08]. redecorates [LeB07e]. Redox [AAG^{+}09, PhXD^{+}06, iWK06]. redox-mediated [PhXD^{+}06]. reduce [FN06, KR07]. Reduced [PVM^{+}05, BLL^{+}06]. reduces [LVDR05, PMC08]. reductase [CLA^{+}05]. Reduction [MYO^{+}05]. redundancy [LRBM09]. reduplication [RRR^{+}09]. reentry [CGLC^{+}08, Les07-62]. reestablishing [MHL^{+}06]. Ref [NSS^{+}08]. reflects [AMS^{+}05]. refractory [KKD^{+}07]. regenerating [EMPK^{+}07, LeB07p]. regeneration [AS09, DLT^{+}07, DFB^{+}07, LPWP^{+}07, LeB06-51, MLHS06, Wil07w]. regenerative [KCS^{+}06, LeB09b]. region [ADD^{+}06, SBC^{+}05, SLP^{+}06, TFC^{+}06b, Tho06]. Regional [YCM^{+}08, JBF^{+}08]. regions [GHK^{+}09, IPAR05, WPL^{+}06, WBLL06]. reglucosylation [PGWH08]. regulate [ALvdG06, BDS^{+}08, BVS^{+}09, CCH09, CLS^{+}06, CM06, CCP^{+}08, DRD^{+}05, DP07, EMMN^{+}05, FHNG05, FHFB^{+}06, FTB^{+}09, GEN^{+}06, GMS^{+}08, GL06, HBPL09, HKN09, HCN06, HHY^{+}08, HMM^{+}05, HCB^{+}08, HHZ^{+}08, IMY^{+}09, KLV^{+}08, LH05, LMHT08, LLT^{+}08, LCS^{+}09, LSIT^{+}08, MKGL^{+}05, MKV^{+}09, OYKY^{+}05, OFY^{+}05, SDD^{+}07, SLMK08, SMG07, TEIS09, TMZ^{+}09, Tum06v, VVL^{+}08, VKR^{+}09, WCN07, WPL^{+}06, WvNS^{+}07, WFJ06, YOS^{+}05, YMC05, YAV^{+}05, ZKR^{+}07]. Regulated [HLL^{+}09, NTHN^{+}08, Tum05-53, AFM^{+}05, ADD^{+}06, BBM^{+}08, COTAE^{+}05, DCA^{+}06, EIK^{+}05, FM05, GCM09, GXJF07, GAS06, Gle07, GDA^{+}08, GEE07, GFC^{+}08, HMS^{+}08, HF06, LD07, LPN^{+}08, LSG^{+}06, LWL08a, LWL08b, LYK^{+}06, LG07, LG08, LLC^{+}08, MPL^{+}07b, NJS^{+}06, NAF09, NAG^{+}08, OOY^{+}06, OMF^{+}05, PSL^{+}06, PB05, PBD^{+}06, SCR^{+}09, SCF^{+}07, TMH^{+}06, KLRS07]. regulates [BWL^{+}08, BTD^{+}07, BRH05, BZHA^{+}08, BDP^{+}07, BFP^{+}07, BWW^{+}08, BBPM07, BMRP07, BBL^{+}08, CMD^{+}08, CBK^{+}06, CNM^{+}05, CZD^{+}09, CHH^{+}08, CCM^{+}07, CFL^{+}05, CBZY08, CSSE06, DTY^{+}05b, DL08, DVF^{+}06, DDB^{+}05, EPV^{+}05, FFW^{+}07, FBO^{+}07, FSG^{+}08, FDB^{+}09, FDM^{+}08, GRGP07, GLBS06, GFH^{+}06, GMvdKC07, GMM^{+}08, GHSG07, GDF^{+}05, GGEd^{+}07, GDA^{+}08, GBNS07, HBK^{+}06, HLL^{+}08, HGR^{+}06, HLK^{+}07, HHFC06, HZR05, HOZ^{+}09, IEGV06, JKKM08, JP06, JSC08, KRZ^{+}08, KBY^{+}08, KAL^{+}06, KKM09a, KNT^{+}06, KDD^{+}09, LPG^{+}07, LS09, LDS^{+}06, LBVN09, LOWTL^{+}08, LJN^{+}05, LHZ^{+}08, LYY08, LBS^{+}08, LDGO^{+}09, LSdWL07, LGV08, LFL^{+}05, LCL^{+}05, MYMH09, MSV^{+}05, MDW05, MB07, NTB^{+}06, NPB^{+}05, NLS^{+}05, NWB^{+}06, NLL^{+}06, NFCF08, NNS^{+}08, OSS^{+}07, OYM^{+}09, OIAT06, PCR^{+}07, PSL^{+}06, PEER^{+}05, PMB^{+}09, PAV^{+}06, PKJS^{+}08, PZG^{+}07, QCGM05, RYD08, Rob08p, RRR^{+}09, RFBB07, SBM^{+}07, SGMPV07, SJL^{+}07, SHS^{+}09]. regulates [SB09, SDI^{+}06, SOR06, SCM^{+}07, SBW^{+}07, SS06b, STK^{+}05, SHC^{+}05, SSM05, SLK^{+}08, SRD^{+}07, SLR^{+}06, SIS^{+}08, SLV^{+}07, SRQL09, TWL07, TTK^{+}06, ToMY^{+}08, TWC^{+}07, TGRD09, Tum05-52, Tum06-28, TTE^{+}07, UWiY^{+}07, VLK^{+}08, VWM^{+}07, VLFA^{+}07, VSR^{+}08, WZO^{+}05, WLC08, WLE09, WBT^{+}08, WMT07, WWC^{+}07, YGR^{+}08, YOSM06, YHJ^{+}05, YMDL09, YPG05, YWM^{+}08, ZS08, ZKG^{+}08, ZDMD06, ZHYE08, dRKD^{+}05, vBAS^{+}07]. Regulating [DP08, Tum05-54, Tum05-55, BFB^{+}08, CG06, DXC^{+}08, GTHK05, IPG^{+}05, KSR^{+}08, LU05, LLJ^{+}05, LPF^{+}06, OKN06, PLR^{+}06, PdBR^{+}07, RMC^{+}05, SKM^{+}08, SSS^{+}08a, SWS^{+}08, TTG^{+}09]. Regulation [BHG^{+}05, CWC^{+}08, CKX^{+}08, CTP^{+}09, CWC^{+}07, ERM^{+}09a, EHWL06, FKS^{+}06, FNH^{+}05, FHS^{+}07b, GWS09, GKvdB07, HM08, JLL^{+}07, JASC08, JhKC08, KVC^{+}09, LJR^{+}05, LVL^{+}05, LC08b, LSC^{+}07, LJY^{+}07, MKC^{+}05, OPT^{+}07, PPK08, PCPW08, QWD^{+}08, RvVM^{+}08, RKB^{+}07, SSAD08, VMZW^{+}07a, VMZW^{+}07b, WLY06, WCR^{+}05, ZPGB^{+}07, vZPG^{+}07, AWE^{+}05, AZP^{+}05, BJML09, BRM^{+}07, BTP^{+}07, BAdSG05, CS09, CYY^{+}08, CR06, DS07, GMW^{+}05, GPN^{+}06, GGMBF08, GJL^{+}08, GNPC^{+}08, HHT^{+}07, HT08, HK08, JLM^{+}07, KKM09b, LGDN09, LW07b, LZM^{+}06, LFH07, MVC^{+}07, MMM^{+}05, MDC^{+}08, MTB^{+}09, MW06, MHFH06, NTY^{+}05, PLH^{+}08, PCMSB09, PT09, PSS^{+}05, PVM^{+}05, RYD08, Rob08u, RWM^{+}05, SWD^{+}08, SSM^{+}08, STG05, SXH^{+}06, SFC^{+}08, SSZR07, SMRS^{+}07, SiTT^{+}08, TRC^{+}08, TSP^{+}07, VP05, VBG^{+}08, WWS05b, YKH^{+}08, YA06, YR09, YTB07, ZTKK05, ZOH^{+}08]. regulator [ASHD^{+}05, BLT09b, CMN^{+}08, DBvdW^{+}06, LF07, MLC^{+}06, NRL^{+}05, PBAF^{+}08, RKT^{+}07, SSH^{+}05, SL05, TTYH09, Tum06-54, WBC^{+}06, WKH^{+}06, YWK^{+}08]. Regulators [BMBD^{+}09, CSW08, KR06, LeB06i, LVG^{+}09, OHP^{+}09, RMAA08, YSB^{+}07]. Regulatory [PR05, QHB^{+}07, CC09, HRK^{+}09, KIKI^{+}07, LB08, LFL^{+}05, NSM^{+}09, OYTO07, PWM^{+}06a, PKJS^{+}08, SKO^{+}08, VBSM06, VTRB06, YWW^{+}08, deC09]. regulon [CTS^{+}06c, CTO^{+}08]. Reichardt [LeB09c]. Reinforced [Tum06-38]. reinforcement [BMK^{+}06, YHJ^{+}05]. reinhardtii [BSM^{+}08, HDM^{+}07, LW07a, LJS^{+}09a, UZ08, iWK06]. reins [LeB05-62, LeB07-28]. rejuvenating [Wil07l]. related [CWRG^{+}08, GBNK08a, GBNK08b, GDF^{+}05, JKW^{+}09, KO06, LB08, MMS07, NAG^{+}08, PRB^{+}09, RIHP06, SOR06, VTT^{+}09a, WZW^{+}09, WMS^{+}06, LBG^{+}06, TWL07]. relationships [TTNA^{+}06]. Relative [OLKK09]. relatives [UDC^{+}07]. relax [LeB06-27, Les07-72]. Relaxing [TSP08, Wel06-46]. relay [BMT^{+}07, LeB06-85]. release [AACS^{+}06, BMK^{+}07, CFH^{+}08, CCC^{+}06, CLK^{+}07, CVACF07, DBK^{+}07, DMC^{+}07, HBPL09, LeB05-54, LeB07-60, MSLS07, OMF^{+}05, SYM^{+}08a, SM09, WBK^{+}06, Wil07z, WWS^{+}06]. releases [HLK^{+}09, vRSvR^{+}07]. Releasing [MD09, GMM^{+}08, PKJS^{+}08]. reliance [SFB09]. relies [LeB05-43, UDC^{+}07]. relive [LeB05p]. relocalization [HCC^{+}07, MHC^{+}09]. remains [PVJJ07, Rob07j]. Remodeling [Wel06-47, ZJH08, BMRP07, DGSN06, EGR^{+}05, GQSHO05, Gra07, GGBV^{+}07, HJK^{+}06, KSA^{+}05, MMS07, RKK09, SDB^{+}05, SWT08, SLNS06, WKH^{+}06, WL09b, XNM^{+}09, vBDF^{+}09]. remodels [LeB07-39]. removal [IKZ^{+}07, MYMZ08, PGW^{+}08, ULN^{+}07]. remyelination [PBB^{+}09]. Renal [OFfL^{+}07]. rendered [HS09b]. renewal [BTD^{+}07, JAV^{+}07, Lin08, NCB^{+}08, NCB^{+}09, SRQL09, WLPB^{+}09]. reorganization [BEN^{+}05, LHL^{+}08, PDM^{+}08, YWJ^{+}07]. reorganizing [LC07]. Repair [DG05, ITG^{+}08, LeB06-61, Tum05-56, Tum06-39, BY09, BFCS^{+}07, BWFdL09, GSB^{+}06a, KG05, LeB05e, LeB08u, Les08j, MBJM^{+}09, TFC^{+}06b, THN^{+}06, Wil07-54, YMTC05, YEL^{+}08]. Repairing [Wel05-55, Les09f]. repeated [MKSJ06]. repeats [HK06, MBJF^{+}08, SDN^{+}05b, TRS^{+}08]. repertoire [ESG^{+}06c]. replicating [BJKF05, Wil07-27]. Replication [FPP^{+}08, LeB05-82, BWFdL09, CCS^{+}08, DP07, ESZK06, GLH^{+}05, GHK^{+}09, LeB06f, LeB07m, LM05, MSO09, MTV^{+}05, MBF^{+}07, MBF^{+}08, PL06, Rob08p, SSS^{+}08b, SHK^{+}07, SiTT^{+}08, TMH^{+}07, TNS07, Tum05-27, Tum05-54, Wel06-47, WGL^{+}06, WSG06, YM09]. replicative [WGL^{+}06]. reply [CDM^{+}07]. Repo [TMAL^{+}06, LeB06-62]. Repo-Man [TMAL^{+}06]. repositioning [DOS^{+}07, LN07, MM08, MGK^{+}09]. represses [NTB^{+}06, YAV^{+}05, ZBKJ07]. repression [HlQY^{+}06, KO07, YIfHA09]. reproduced [PCV^{+}08]. reprogramming [GER^{+}07, Sho09c, Wil07-32]. Republic [LeB06-52]. Repulsion [SMSPK08, PCPW08]. repulsive [HKIY09]. require [AMERT^{+}07, GHX^{+}07, PGP^{+}07, ZWO^{+}09]. required [ARKS^{+}08, ASH06, AM05b, API^{+}07, ABY^{+}08, AHM^{+}05, AHD^{+}09, BM07, BDP^{+}07, BKE^{+}07, BW06, BFCS^{+}07, CKG^{+}08, CHCB05, COTAE^{+}05, DBC^{+}06, DJV^{+}05, DSBK09, DHR^{+}08, FFT^{+}05, FSP^{+}07, FSR^{+}07, FSK^{+}07, GGB^{+}08, GMG^{+}07, GTX^{+}06, GTO^{+}07, GMZ^{+}08, GSL^{+}07, GEE07, GKN^{+}07, HOM^{+}05, HTK^{+}08, HOR08, HRS^{+}05, HQF^{+}07, HBN^{+}09, JS07, JGY09, JET^{+}07, JFM^{+}09, KNY07, KBN^{+}06, KOVP09, KHBV05, KLR^{+}07, KNP^{+}05, KCH^{+}06, KSE^{+}07, KRG^{+}08, LW07a, LJS^{+}09a, LC07, LSC^{+}07, LWHC08, LPF^{+}06, LYTL07, LGT^{+}07, LYV^{+}08, MBR^{+}06, MHM^{+}07, MHL^{+}06, MOS^{+}09, MFC^{+}09, MDH^{+}06, MDPG07, MHFH06, NSS^{+}08, NVEF07, NFC^{+}08, NYN^{+}06, NFZ^{+}09, OW06, OOY^{+}06, OVL^{+}07, ORM^{+}08, PR05, PCMSB09, PHB^{+}07, PBB^{+}09, PES^{+}07, PMK^{+}07, RCT09, SON06, SYRL^{+}08, SRL^{+}06, SDDV05, SNC^{+}06, SSD^{+}09, SB06, SHK^{+}07, SCMV07, SDW^{+}09, SCK^{+}07, SLWA07, TKSNG07, TRS^{+}08, pTyLZ^{+}09, UPBM06]. required [VDA^{+}08, WZW^{+}09, WJM07, XYC^{+}07, YDL^{+}06, YTY^{+}07, YRSM^{+}09, YSP^{+}06, ZTKK05, dGHW08, vBDF^{+}09]. Requirement [IUT^{+}07, WWS^{+}06, YGW06, BCC^{+}09, CVACF07, LZ09, PMS^{+}07, YLM06a]. Requirements [GS07, SHC09]. requires [ASK^{+}09, BSA^{+}08, BJML09, BH08b, BPL^{+}08, BMSJ07, CLK^{+}07, CWM^{+}09, DDRL^{+}09, DSLY^{+}05, GMSS06, HHS^{+}09, IKZ^{+}07, IBS^{+}05, JBFC07, JSPT05, JZN^{+}09, JKW^{+}09, KBR^{+}08, LCL^{+}07, LLW^{+}09, LPB^{+}08, LR07, MTV^{+}05, NPMZ06, NRD^{+}06, NYI^{+}05b, PGB^{+}06, PCN^{+}05, POG^{+}09, RIHP06, SFC^{+}08, SHH^{+}08, SDS^{+}05, TdVM^{+}07, TT09, TW07, TPA^{+}08, TSALS05, WPH^{+}07]. Rer1p [SRD^{+}07]. rereplication [LLC^{+}07]. Rescue [NRJ^{+}09, Les07o, Les07-43, Tum05-47]. rescues [CC09, KHB^{+}05, Tum05w]. research [Les06a, Ros05]. researcher [Wil07w]. Reshaping [AH08]. residency [FVSC08]. resident [HEW07]. Residual [SWB^{+}07]. residues [CJMG05, MNH05, WHC^{+}07]. resilences [Rob08-51]. resistant [MCD^{+}07]. resolution [KKGD05, MWSK06, SD07, TSALS05, Wil06g]. resolved [LGH^{+}05]. resorption [ZKR^{+}07]. respectively [CS09]. respiration [PhXD^{+}06]. respiratory [BMT^{+}07, KHSM07, WvdLH^{+}07]. Response [SDS^{+}05, BRM^{+}07, BJLK^{+}06, BBFN07, BLMS06, CNM^{+}05, CSMP^{+}09, CRLP07, DCA^{+}06, FMM^{+}08, FTB^{+}09, FCB^{+}09, FCB^{+}19, HKLM06, HKLM07, JMB^{+}07, KHW^{+}07, KHW^{+}09, KDD08, LBS^{+}08, LHRP05, MBJM^{+}09, MFBG^{+}08, MDWH^{+}08, MSC^{+}07, MJF^{+}07, OOY^{+}06, OMGW^{+}06, PGP^{+}07, RVH08, SPT^{+}09, Sed07c, Sho09-51, SRS^{+}09, SLZ^{+}06, Tum05-65, YOSM06, SPT^{+}21]. responses [DLBSD^{+}05, LLS^{+}09, MLV09]. responsible [CCL^{+}05, JTH09, WF05]. responsive [GQSHO05]. responsiveness [LJR^{+}05]. restrain [BJKF05]. Restraining [LeB05-83]. restrains [TST08]. Restricted [WWX^{+}06, EMPK^{+}07, MBR^{+}06]. restriction [CPAH08, YWRA07]. restricts [BZA^{+}06, GLS^{+}09, HMS^{+}08, SRR^{+}08, WSFD07]. restructuring [KPM05]. result [DSN^{+}07b]. resulting [Ros05]. results [BGJ^{+}07, KPS^{+}07, MLH^{+}06, RVH08]. Retention [FSN^{+}06, BJLM^{+}05, MBJF^{+}08, PS08]. reticulons [RW08b]. reticulum [ASB^{+}05b, AH08, ADO07a, ADO07b, AWM^{+}08, CHO^{+}09b, JASC08, KPM05, LC07, LMC^{+}09, LHRP05, LFS^{+}05, NTHN^{+}08, NSM05, PMW05, PVJJ07, RCFB08, RMP06, SPT^{+}09, SPT^{+}21, SBV^{+}06, WWS^{+}06, ZKN^{+}06]. retina [PVMH05, WMM^{+}09]. Retinal [JLB^{+}05, HLJ^{+}07, LJR^{+}05, RWB^{+}05, SRR^{+}08]. Retinoblastoma [BvOM^{+}07, VWM^{+}07, DVF^{+}06]. retinoic [ML05]. retinoid [HGK^{+}06]. Retraction [KTYT05, FSG^{+}07, HWJ^{+}08, YJJ^{+}09]. retrieval [KKW^{+}07]. retroelement [LTK09]. Retrograde [Tum05-57, ARKS^{+}08, GSJ^{+}08, HLP06, STU^{+}08, WSB^{+}09, ZL05]. retromer [RvVM^{+}08, SSSB07]. retromer-dependent [SSSB07]. retrotranslocation [FSnP^{+}06, WLY06]. return [Wil06l]. returning [Wel07m]. reunion [LeB05-70]. Rev1 [MWY^{+}09]. Rev1/Pol [MWY^{+}09]. reveal [STG05]. revealed [CHSV07, HDS^{+}07, HSTM06, MWSK06, NTHN^{+}08, RBC05]. reveals [BBFN07, BGP^{+}06, BLA^{+}09, DOVD08, DvCvH^{+}08, DCMB^{+}07, EMB^{+}08, EMD^{+}08, FMA^{+}09, Heu05, LCvD^{+}08, MBJM^{+}09, MK07, NLC^{+}05, OIFN^{+}07, PH07, RGW^{+}07, SPA^{+}05, ShCM^{+}06, SWG^{+}05, VCSG07, Wil08k, YTH^{+}09, YCM^{+}08, YSB^{+}07, ZTZ^{+}09]. reversal [DR05, Tum05-58]. reverse [LeB06-60, Pow07d]. reversed [CR06, ERM^{+}09b]. reverses [SCFKB07]. Reversible [FM05, HSLI09, LSDFS08]. Reversion [ZTZ^{+}09]. Reversion-induced [ZTZ^{+}09]. review [Ros07b]. reviewing [RP08]. revisited [BP05, SSHW09]. Revisiting [BY09, MS07, ELM09]. RGD [TLM^{+}07]. RGMa [HFY^{+}06]. RGS2 [NMZ^{+}09]. RGS7 [DDB^{+}05]. Rhabdomere [RCM^{+}09]. rhabdomyosarcoma [BIH^{+}05, BIH^{+}07]. Rhamm [THN^{+}06]. rhesus [CPAH08]. Rho [HO08, RWM^{+}05, DvRF^{+}05b, EPV^{+}05, GGEd^{+}07, GFD^{+}05, HWJ^{+}08, LeB05-36, LSNY08, PG06, RCS^{+}09, SYKM08, SWI06, WSB^{+}06, WCN07, YN07, YMC05]. Rho-dependent [HO08]. Rho-GTPase-dependent [LSNY08]. Rho1 [WL09a, WL09b]. RhoA [dJPAW^{+}09, ZHI^{+}08, BBvD05, BB05, BDR^{+}06, GdTAR07, KRZ^{+}08, LC08b, LLT^{+}08, MMW^{+}07, OMLP^{+}08, PLR^{+}06, Rob08i, SGMPV07, SYRL^{+}08, SSD06, Tum05-59, VWK^{+}06, WPL^{+}06, WVB07, YPG05]. RhoA-dependent [KRZ^{+}08, LC08b]. RhoG [ERM^{+}09b, Les07-57, vBAS^{+}07]. RhoGAP [RFBB07]. RhoGDI1 [ERM^{+}09b]. RhoGEF [SMR09]. RI [NYI^{+}05b]. RI-mediated [NYI^{+}05b]. ribbon [RWB^{+}05, Sho09-42, WS09, tDAK^{+}05]. ribonucleoproteins [GTO^{+}07]. ribosomal [KZS07, Les07-73, MTRJA05, MTRD^{+}06, SPS^{+}08, ZWO^{+}09]. Ribosome [LeB05-84, LLW^{+}09, Tum05-60, CJMG05, GEE07, HRS^{+}05, Joh09, JhKC08, KRPH08, KZS07, Poo09]. Ribosomes [Les08-43, Wel05-56, FTM^{+}06, LeB08j]. ribosylation [LDGK08]. RIC [SHL^{+}09]. RIC-19 [SHL^{+}09]. rich [BVGA^{+}05, DTP07, MPS^{+}07, RM07]. RICH2 [RKH^{+}09]. Richard [Wil08t]. richer [RM07]. RID [CC09, SCM^{+}07]. RID- [CC09]. ride [Les08a, Pow05w, Wel05u]. Riding [MW06]. Rif1 [BWFdL09]. right [Les07-63, PRB^{+}09, Sed09m, Sho09-41, Sho09-48, Tum05-37, Tum06-45]. RILP [JRZ^{+}07, RKvdK^{+}09]. Ring [LeB05-85, Les06y, BVS^{+}09, HRB^{+}06, KOM07, PYR08, RGCWG09, Tum06i, Tum06-27, Wel05-35, WSK^{+}06]. Ring1B [LW07b]. rings [HYB08, LeB08i, TAM^{+}06]. RIP1 [GHM^{+}09]. rise [VMY^{+}09]. rising [Fuc08]. Rix7 [KRPH08]. RLIP76 [GPJ^{+}06]. RMD [OOS07, OOS09]. RMD-1 [OOS07, OOS09]. RNA [AK06, AHM^{+}05, BF09, CTS^{+}06c, CTO^{+}08, CVACF07, DKR^{+}06, FKJ^{+}09, GRGP07, GTX^{+}06, HBPL09, HBN^{+}09, KII^{+}08, LeB06-46, Les06u, Les07-59, Les07-58, Les08-44, LGBD^{+}07, MRL^{+}08, MIBT06, PAM^{+}07, PNB07, SPS^{+}08, TCKV05, UZ08, VBL^{+}05, WIL^{+}05, Wel05-33, Wel05-57, WGKA^{+}08, Wil08n]. RNA-binding [GTX^{+}06, KII^{+}08, PAM^{+}07]. RNAi [WKIKG09, YSB^{+}07]. RNAi-mediated [WKIKG09]. RNAs [BJZ^{+}09, BFWH07, BAR^{+}08, FFS06, HLL^{+}09, VBL^{+}05]. RNase [TP09]. RNF8 [MBJM^{+}09]. RNP [DKR^{+}06]. RNPs [BMGPB^{+}08, PNB07]. Rny1p [TP09]. road [LeB06-61, Les09j, Sed09k, Wil08o]. roadblocks [Wil07s]. roar [RW08d]. Robo [Rob08-40]. robustness [CWRG^{+}08, DGG^{+}08]. ROC2 [GBNS07]. ROCK [SWI06, GdTAR07, OMLP^{+}08, PLR^{+}06, STK^{+}05, Tum05-51]. ROCK-dependent [GdTAR07]. ROCK-I [STK^{+}05]. ROCK1 [HWJ^{+}08]. ROCKII [LC08b]. rod [FSN^{+}06, NPBP07]. Roger [Wil07-55]. Rok [NSM^{+}09, PMK^{+}05]. Rok- [NSM^{+}09, PMK^{+}05]. Role [AG09, GPN^{+}06, HFM^{+}05, LMC^{+}09, NSO^{+}09, OSM^{+}09, PKM^{+}05, SS08a, TPZG06, VDSV06, ViKA^{+}06, AACS^{+}06, AFN^{+}08, ANI^{+}07, ADO07a, ADO07b, BY09, BFG^{+}06b, BWJ05, BBFN07, CCSBM^{+}07, CK07, CTS^{+}06b, DLT^{+}07, DBR09, EYB09, FBD^{+}05b, FA08, FBD^{+}08, GKCM^{+}08, GMW^{+}05, GXJF07, GTHK05, HDS^{+}07, JD06, KCAS05, KNT^{+}07, Lec08, LYH^{+}05, LLJ^{+}05, LSC^{+}06b, LKL^{+}05, LDPLS05, LRCH^{+}05, MHM^{+}05, MMCW06, MPL^{+}07b, ML05, MGH^{+}09, MS07, MK07, NJS^{+}06, NLC^{+}05, OWM05, PGWH08, PLR^{+}06, PKJS^{+}08, PST^{+}06, PdBR^{+}07, RKH^{+}09, RP07, SSHP^{+}07, SPA^{+}05, SSS^{+}08b, Ska07, ShCM^{+}06, SWH^{+}06, TDP^{+}06, Thu07, TM06, VBG^{+}08, VMSV08, WAKS08, WC09, WLLF^{+}08, WKW^{+}08, Wil07y, WNF^{+}05b, XKYW07, XGX^{+}09, ZNP^{+}09, vKGK07, CLR^{+}06]. Roles [HLP06, BTP^{+}07, CR06, ELJ^{+}08, FSnP^{+}06, IPG^{+}05, KKH^{+}06, hKDH^{+}08, KCS^{+}06, LSC^{+}05, LGBD^{+}07, MLC^{+}07, MYZ^{+}09, MMS07, PRK^{+}07, PRK^{+}09, PWW^{+}09, RP08, SHDT05, SWG^{+}05, TNS07, WJM05, WBE^{+}08, ZTM^{+}08]. Roll [Les05h, Cha07, Les09-27, Les09-29]. Roll-your-own [Les05h]. rolling [LeB06-77, MOCC05]. romance [LC08a]. room [Les07-40]. root [LeB05k, PST^{+}06, VCN^{+}05]. Rootletin [BSW^{+}05]. Ror2 [NYN^{+}06]. ROS [CKX^{+}08, Tum05-44]. Rosettes [Wel06-48]. rotate [LeB05-28]. Roundabout [SMSPK08]. roundup [Les09h]. rouser [Les08-41]. route [LeB05-69, Les09i, LTAM^{+}09, Sho09-30]. routes [GEP08, Les09e]. row [AGL^{+}08]. Royal [Wil07c]. Rpb7p [LGBD^{+}07]. RPTP [BLS^{+}05, BSL^{+}08]. rRNA [Wel05-46]. RRP [Gle07]. RSF [PNO^{+}09]. Rtt109 [iHBD08]. ruler [LeB05b]. Rules [Les08-45, Les09-31, LeB05m]. run [Sho09g, Wel05e]. runaway [Wil08n]. runs [Rob07l]. Runx2 [LHZ^{+}08, ZBKJ07]. Runx2/Smad3 [ZBKJ07]. Runx2/Smad3-dependent [ZBKJ07]. ryanodine [OFY^{+}05]. Ryk [KHH08]. Ryoo [Wil09l].

S [YOSM06, AH05, ASH^{+}09, AAG^{+}09, BJKF05, CCS^{+}08, EHHK09, FMK^{+}09, HDS^{+}07, HH06b, JET^{+}07, LeB05-76, LeB08f, LLC^{+}07, MTV^{+}05, MSB06, MBF^{+}07, MBF^{+}08, MKV^{+}09, NSO^{+}09, PL06, RMC^{+}05, TTYH09, TMH^{+}07, ZNP^{+}09]. S-glutathionylation [AAG^{+}09]. S-phase [AH05, CCS^{+}08, JET^{+}07, PL06]. S2 [VSG09]. S365 [SMRS^{+}07]. SA1 [CS09]. SA2 [CS09]. SAC1 [BCR^{+}08]. Sac1p [FKB^{+}05]. Saccharomyces [AFN^{+}08, BMP08, CCSK08, DTP07, ESZK06, FFT^{+}05, GAS06, KKGD05, NRD^{+}06, TKK^{+}05, TP09]. Sad1 [BMSJ07, JMG^{+}06]. Sad1-UNC-84 [BMSJ07, JMG^{+}06]. SAG [GBNS07]. SAG/ [GBNS07]. Sak [RRR^{+}09]. Saka [Wil08x]. Salamander [MLHS06]. salivary [Les07q]. Salmonella [Wil06n, MES^{+}08, PG06, SHB^{+}07]. Salmonella-host [PG06]. SAM [SGKP^{+}07a, SGKP^{+}07b]. Sam68 [PAM^{+}07, PMB^{+}09]. same [ZMGC05]. San [LK06, WL07, dSW09, HCK^{+}07]. SAR [KCE^{+}06]. Sar1 [BHG^{+}05]. sarcolemma [RKN^{+}07]. sarcomere [LLH^{+}07, WBL^{+}06]. sarcoplasmic [RMP06]. Sarcospan [PMC08, Rob08-41]. SAS [DMDO08]. SAS-4 [DMDO08]. satellite [HLK^{+}07, JTN^{+}05, MLHS06, NPMZ06, SLV^{+}07, Wel05-29]. satisfied [BYR08, OLH^{+}08]. satisfy [YKBR09]. save [LeB06a, Les08-61, Les08-62]. saves [Rob08-38, Tum05-62, Wil09i]. Saving [Wel05-58]. saviors [Wil06j]. says [LeB05n, Rob08-52]. SC-35 [SBJ^{+}07]. SC35 [SRBL06]. scaffold [BZHA^{+}08, EYB^{+}06, GTHK05, MTL^{+}06, PFD06, Pow05e, SWL^{+}05, TKB^{+}06, Wil07-60]. scaffolding [RKN^{+}07, WJM05]. scaffolds [LGDN09]. scale [BEN^{+}05, DR05, HKP^{+}09, Wil07-45]. scales [ELM09]. scaling [BBM^{+}07]. SCAPER [TWC^{+}07, Wil07-56]. Scarred [LeB07-54]. Scattering [Tum06-40, dRKD^{+}05]. scavenger [DSLY^{+}05]. Scc3 [CS09]. SCF [FBJB^{+}09, RRR^{+}09]. SCG10 [TÖL^{+}06b]. SCH9 [MGW^{+}08, MWY^{+}09]. schizophrenia [CFH^{+}08, Les08-49]. Schizosaccharomyces [CBZY08, GCM09, LB08, RGCWG09]. Schwann [BTP^{+}07, GKN^{+}07, Les08-60, LLM^{+}05, MEG^{+}07, MMT^{+}05, NZQ^{+}07, PBB^{+}09, PT07, YMCT08]. science [HdM09, LeB07-53, LeB09c, Sho09u, Wil08j]. sciences [VD09]. Scientific [RVH08]. scientist [Wil07c]. scientists [Mel09]. sclerosis [KR07, MSH^{+}05]. sclerosis-associated [MSH^{+}05]. scope [HS09a]. Scoring [Sed08m]. Scott [Wil07-57]. Scramblases [AEJ^{+}06]. screams [Pow06l]. screen [FMA^{+}09, Pow05d, YSB^{+}07]. Scribble [QCGM05]. Sculpting [deC09]. SDH [CMD^{+}08]. SDT [MBJF^{+}08]. sea [Wel06-44, WNF^{+}05b]. seal [Wel05-52]. seals [Pol08]. seamless [WC07a]. search [AS09, PT09]. Sec14p [VCN^{+}05]. Sec14p-nodulin [VCN^{+}05]. SEC18 [JWSS05]. Sec1p [CCS^{+}06]. Sec1p/Munc18 [CCS^{+}06]. Sec3 [ZOH^{+}08]. Sec4p [GAG^{+}06]. Sec5 [SORM05]. Sec6 [BLP^{+}05]. Sec61 [ASB^{+}05b, LeB05-53]. Sec61p [CJMG05, Poo09, SS08a]. second [BAdSG05, OIFN^{+}07, YSSH^{+}07]. second-phase [OIFN^{+}07]. secret [HB08]. secretase [IDTT^{+}09, KKC^{+}07, SRD^{+}07, Wil07-61]. Secreted [EFCP07]. Secretion [LeB05-86, BSF^{+}08, HXZ^{+}07, HGJ^{+}08, LeB05-27, LSR07, LBG^{+}06, MKH^{+}07, Rob08-27, RKB^{+}07, SRL^{+}06, TPA^{+}08, Tum06z, Tum06-53, Wel05-50]. Secretory [TCW^{+}06, AOT06, AMS^{+}05, DR07, FML^{+}06, FHS^{+}07b, GOM^{+}09, GAG^{+}06, KKC^{+}07, KES^{+}09, Pow05j, Sed07b, SRD^{+}07, ZKN^{+}06]. secrets [DOVD08, Sed09h, Wil07-31]. Securing [Pow09c]. security [Ros08a, Ros08b, Wil06n, Wil09q]. SecY [COC^{+}05, LeB05-73, SNA^{+}07]. Sed5 [SSP^{+}07]. see [Rob08g]. Seeing [Bon05h, Nor06, Tum05-61, Wel06-49]. seek [Smi07]. segment [BHY^{+}08, HXO^{+}07, HOR08, Poo09]. segments [DZG^{+}07, KMS07, NPBP07, OTH^{+}07, Tum06q, YOHR07]. segregate [PMW05]. segregated [WBK^{+}06]. Segregation [KES^{+}09, VMKW^{+}08, Wel06-50, CM07, CGJ09, HGJ06, HP05, KKM09a, KCH^{+}06, Les09p, MTRJA05, MFH05, MPLF^{+}09, MJG07, OOS07, OOS09, YRSM^{+}09]. seipin [FSG^{+}08]. SEL1L [MLP06]. select [WHOS06]. selectin [LYK^{+}06]. selection [KBC^{+}07a]. Selective [LeB06-63, MHM^{+}05, Nyb05i, TDT^{+}08, BIH^{+}05, BIH^{+}07, HSY^{+}06, KYK^{+}06, dGSWW^{+}06, iYOW^{+}06]. selectively [FWC^{+}07, LCvD^{+}08, LYY08, NTSY08]. Selectivity [LeB05-87, CLL07]. Self [Bas07b, LC09, Les07-60, RYN05, Rob08-42, SS08b, Wil07-58, BTD^{+}07, HK05, JAV^{+}07, LeB07-33, Les08k, Lin08, NCB^{+}08, NCB^{+}09, NS06, Pow09b, SAM^{+}07, SRQL09, THL^{+}06, TD08, Tum06a, WLPB^{+}09]. self-antigens [LeB07-33]. Self-assembly [Rob08-42, SAM^{+}07]. self-association [NS06]. Self-eating [Bas07b, SS08b, Wil07-58, Les08k]. self-incompatibility [Tum06a]. self-incompatible [THL^{+}06]. Self-organization [LC09, RYN05, HK05]. self-preservation [Pow09b]. self-renewal [BTD^{+}07, JAV^{+}07, Lin08, NCB^{+}08, NCB^{+}09, SRQL09, WLPB^{+}09]. Self-starting [Les07-60]. selfish [CGJ09]. Sem1 [FKJ^{+}09]. Sema2-mediated [WZC^{+}08]. Semaphorin [GMM^{+}08, OKN06]. seminiferous [WC07a]. Senataxin [SBC^{+}07]. Sending [DR07, Wel06-51]. sends [LeB08q, Pow06b, Sho09h, Sho09i]. senescence [FSTI06, MYO^{+}05, MZH^{+}08, SSS^{+}08a, Wil06l]. senescence-like [MYO^{+}05]. SENP3 [YWM^{+}08]. SENP5 [YWM^{+}08]. sense [Ros08a, Ros08b, SR07]. senses [Tum06a, Wil07-29]. Sensing [Nyb05j, LeB05-97, LeB05-90, LeB06b, LeB08k, RVBZ^{+}08, SH05, TJC06, WOP^{+}06]. sensitive [BRY^{+}09, MMM^{+}05, NFI^{+}09, RCM^{+}09]. sensitivity [MLH^{+}06, PMK^{+}05, WNL^{+}09]. sensitization [HPH^{+}05, TDT^{+}08]. sensitize [VCC^{+}08]. sensitizes [DDO^{+}08, FLA^{+}08]. sensor [FBD^{+}08, Har08, KIKI^{+}07, NJS^{+}06, RKvdK^{+}09]. sensor-1 [NJS^{+}06]. sensors [Bas07e, DCA^{+}06, FSH08, GRC^{+}05]. sensory [DNG^{+}06, ESG^{+}06c, KKM09b, KBGG^{+}08, SSJ^{+}07]. Senthil [Sed09k]. Separase [QU08, WWK^{+}06, KYK^{+}06, NAF09, Tum06o]. separase-regulated [NAF09]. Separate [Les07-61, DBD^{+}09, LeB05-49, TWG^{+}09]. separately [Les09y]. Separating [LeB05-88, RGL^{+}09]. separation [KPS^{+}07, MTV^{+}05]. Septin [LFS^{+}05, GTHK05, LYTL07, SHH^{+}08]. Septin-dependent [LFS^{+}05]. Septins [LeB08-27, MPL^{+}07b, Wel06-55]. septum [NSO^{+}09]. sequence [CWM^{+}09, NS07, SLP^{+}06]. Sequential [OIT08, PRK^{+}07, PRK^{+}09, DBW^{+}06, KKK^{+}05, RCS^{+}09, RvVM^{+}08]. sequestration [LFL^{+}05]. Ser273 [NWB^{+}06]. SERCA [GDA^{+}08]. serine [LCR^{+}05, NEF^{+}06, SDC09, WHC^{+}07]. serovar [SHB^{+}07]. Serpin [KNKH06]. Serum [LeB06-64]. serve [BVGA^{+}05, DNG^{+}06]. serves [BHY^{+}08, HSLI09]. set [Wel05-53]. Set7 [TMH^{+}07]. SET8 [JET^{+}07]. sets [PMK^{+}05]. severe [RMR^{+}09]. Severing [Tum05-62, EGR^{+}05, MTH^{+}06, RMV06, SSK^{+}09, WEL^{+}07n, ZRBS07]. sex [DLT^{+}07, MBdR^{+}08]. Sfi1p [LSC^{+}06b]. Sgk3 [AOP^{+}05]. SH2 [LSdWL07, SDD^{+}07]. SH3 [RGCWG09]. shape [Bon05f, DL08, DJV^{+}05, FvD08, LHS^{+}05, LeB06-65, LeB07r, Les06k, Rob08q, Sed08h, SDI^{+}06, Tum05-64, Wil07-59, Wil09j, WAW08]. shapes [LeB05-29, LeB05-34, LeB08h, Rob08-44]. shaping [WWL07, Sed08h]. share [Les08n, PPA^{+}08, Rob08-54]. Shc [LSIT^{+}08, Rob08d]. shear [Tum05c, ZXM^{+}06]. shear- [ZXM^{+}06]. sheath [WTE^{+}08]. sheaths [Tum05-36]. shedding [DLBSD^{+}05, OHH^{+}06]. sheds [Sho09-50]. sheet [PVJJ07]. sheet-to-tubule [PVJJ07]. Sherwood [Sho09q]. shields [LeB07-50]. Shiga [FHS^{+}07a]. Shigella [INB06]. SHIP2 [ANI^{+}07]. SHIP2-dependent [ANI^{+}07]. shock [KFK^{+}05, MG08a, YOY08]. Shoichiro [Tak06]. Shootin1 [STU^{+}08, TSK^{+}06]. Short [Les08-46, CWM^{+}09, LeB07m, Les07-39, SS08a, SCFKB07]. short-lived [SS08a]. shortage [Les07-74]. shortening [Cho09a]. shorter [BLL^{+}06]. shot [Les06x, Sho09-44]. Shotgun [SMSPK08]. Shotgun/E [SMSPK08]. Shotgun/E-cadherin-mediated [SMSPK08]. Show [RVH07, Les07-68, Les08-37]. shows [Rob08-53]. SHP [FBY^{+}06, LPN^{+}08, MKV^{+}09]. SHP-2 [FBY^{+}06, LPN^{+}08, MKV^{+}09]. SHP2 [RFBB07, BCT06, LC08b]. SHPRH [MSM^{+}06]. Shr3 [KGL07]. Shrinking [Bas07c]. Shuffling [LeB08-28]. shugoshin [HCK^{+}07]. shugoshin-independent [HCK^{+}07]. shui [Les08t]. shut [LeB06-83]. Shuttling [ERS08, DDB^{+}05, FFS06]. side [HMM^{+}05, Les07-68, Les07-74, Les08-37, ONW^{+}07]. sides [Di 08]. sight [Wel05f]. Sights [LeB06-66]. Signal [MBR^{+}05, Tum05-63, ASH06, AMS^{+}05, ABSK06, BSK07, CKX^{+}08, CHO^{+}09b, FN06, GXJF07, GH05, GGBV^{+}07, IBS^{+}05, KP08, KNT^{+}07, LJA09, Les06o, MPL^{+}07a, NS07, PGP^{+}07, SMH07, oSCW07, SPK^{+}09, SL05, TSK^{+}06, Tum05-79, WBG^{+}08]. Signal-mediated [MBR^{+}05]. signal-regulated [GXJF07]. signaler [RW08e]. Signaling [AACS^{+}06, CNMM^{+}06, LeB05-89, LeB06-67, Rob08-43, Tum05-64, AOP^{+}05, AWR^{+}07, AG09, ASHD^{+}05, ADC^{+}07, BLT^{+}09a, BWL^{+}08, BKE^{+}07, BHJ^{+}05, BCT06, BCR^{+}08, BMRP07, BCC^{+}09, BPL^{+}08, BSH^{+}07, CKKG05, CLS^{+}06, CTP^{+}09, CZSG08, CLG^{+}07, CWC^{+}07, CTO^{+}08, CCL^{+}07, DXC^{+}08, EPV^{+}05, EYB^{+}06, FKB^{+}05, FZT^{+}07, FBY^{+}06, FMA^{+}09, GPN^{+}06, GCSG^{+}05, GSiK^{+}09, GCS^{+}07, HLM^{+}08, HPT^{+}07, HBK^{+}06, HCRT08, HGK^{+}06, Hor09, HWJ^{+}08, HPH^{+}05, JGAS08, JD06, JGL^{+}06, JMG08, JHY^{+}08, KRZ^{+}08, KKH^{+}06, KBY^{+}08, KZC^{+}08, KWS^{+}08, KBW^{+}06, KGKV07, KRP^{+}09, KKW^{+}09, KAD^{+}08, KLW^{+}05, KKW^{+}05, KLV^{+}08, KCOS06, LJR^{+}05, LNG^{+}05, LPG^{+}07, LGDN09, LOG^{+}06, LJS^{+}09a, LDKT06, LSG^{+}06, LYWK08, Les07-75, LLM^{+}05, LMHT08, LCB^{+}08, LSC^{+}05, LTAM^{+}09, LFL^{+}05, LCL^{+}05, LGG^{+}05, MCF^{+}05, MAH^{+}05, MDC05, MYMH09, MCRVC^{+}08, MZ08, MRD^{+}08, MOM^{+}06, MZH^{+}08, NPMZ06, NSM05, NBM^{+}08]. signaling [NPV^{+}07, NBS^{+}08, NNS^{+}08, PRZ^{+}08, PMK^{+}05, PLR^{+}06, PBD^{+}06, Pol08, Pow07j, RCS^{+}09, SKM^{+}08, SRN^{+}07, SB06, SHX^{+}07, SXH^{+}06, SHjC^{+}07, SOC^{+}07, SLV^{+}07, SSAD08, SFEK07, SLNS06, TTK^{+}06, Ter06, TRS^{+}08, THN^{+}06, pTyLZ^{+}09, Too05, TDG^{+}06, TLK^{+}06, Tum06n, Tum06-29, Tum06u, VDA^{+}08, VKGV08, VSR^{+}08, VCC^{+}08, WBC^{+}06, WLC08, WNL^{+}09, WBK^{+}06, Wel05-54, Wel06d, Wel07k, WCN07, Wil07h, WSB^{+}09, YCR06, YBT^{+}09, iYOW^{+}06, YLM06a, YR09, ZHI^{+}08]. Signaling-dependent [CNMM^{+}06]. signalling [CTGVGFF08]. signals [BMR^{+}08, DDH09, GRC^{+}05, HLA^{+}06, HSK^{+}05, KDV^{+}09, LDGK08, LLR^{+}05, LSIT^{+}08, LGG^{+}05, Mua05, OIT08, Pow06d, Sco08, SFC^{+}08, Sho09h, SWI06, STMG05, TG05, VWK^{+}06, Wil09j, Wil09r, ZDMD06]. signatures [OV09]. significance [CCC^{+}05, CRV^{+}06]. signifying [Wel06o]. SIK [KLV^{+}08]. silences [MDC05]. silencing [GLS^{+}09, GSV07, HKLM06, HKLM07, KHW^{+}07, KHW^{+}09, LN07, Les07k, LTK09, MBdR^{+}08, Rob08-38, Wil07-31, Wil07-65]. silent [HPH^{+}05, Rob08-51]. Silnoon [JLC08]. Similar [XC08]. Simons [Sed08j]. Simple [LeB05-90, Rob08-44, TTNA^{+}06, Les07-62, OIAT06]. simplex [CTS^{+}06b]. SINEs [Pow07i]. Single [EMD^{+}08, SN05, BB05, BJZ^{+}09, DGG^{+}08, Heu05, KGH^{+}05, LeB05-47, NTHN^{+}08, PCV^{+}08, Pow07f, SFEK07, SFS^{+}07, Wel06-31, WNF^{+}05b, ZMA^{+}07]. Single-molecule [EMD^{+}08, DGG^{+}08, SFEK07, SFS^{+}07]. sink [KRP^{+}09]. Sir [Wil08u]. SIRT1 [Les09-32, YMDL09]. SIRT2 [Les08-47, PLH^{+}08]. Sister [HS09b, JGY09, KPS^{+}07, Les09-34, TKSNG07, WWK^{+}06]. sisters [Les07-35]. Sisyphean [LeB05-91]. sit [Wil07-40]. Site [KBC^{+}07b, BKE^{+}07, CVACF07, DKR^{+}06, GCF^{+}05, JBF^{+}08, MSM^{+}07, MRK^{+}07, PCG^{+}05, PPBF08, YBT^{+}09, ZHI^{+}08, PXD^{+}07]. Site-1 [PXD^{+}07]. Site-specific [KBC^{+}07b]. sites [ABSK06, BJLM^{+}05, BKLT^{+}08, DPL^{+}08, DNG^{+}06, EMPK^{+}07, HHZ^{+}08, JS07, KSA^{+}05, KKK^{+}05, KZS07, KCD^{+}06b, MRL^{+}08, PGP^{+}07, PS08, SGKM08, SDC09, TCW^{+}06, TDG^{+}06]. Sitting [Tum06-41]. situ [BSM^{+}08, PGL^{+}09]. Size [Les08-35, DBvdW^{+}06, ELM09, FSG^{+}08, GMW^{+}05, GPC^{+}06, GL06, HG07, LD07, NN07b, Nyb05a, PFD06, Tum06-44]. size-wise [HG07]. Sizing [LeB06-68, Les07-63, Sho09-51, HW09]. Skating [Pow06k]. skeletal [BWL^{+}08, BLL^{+}06, CMN^{+}08, DLT^{+}07, DFB^{+}07, DFA05, FBY^{+}06, FKS^{+}06, GSB^{+}06a, GXJF07, GEN^{+}06, GNE^{+}07, LPWP^{+}07, Las09, LSG^{+}06, LRS05, MLHS06, NRJ^{+}09, RMP06, SDL^{+}05, ZMW^{+}08]. skeletogenesis [HGK^{+}06]. Skeleton [Les06z, MFM^{+}06, NFZ^{+}09]. Skeletons [ZHS06]. Ski [JGAS08]. Skin [Bas07d, CJW07, Fuc08, CBM^{+}07, CTP^{+}09, FKML^{+}06, GBNS07, KSE^{+}07, LeB07-56, LeB07-61, Les07-29, NMV^{+}07, OTM^{+}07, Sho09u, Sho09-53, dGSWW^{+}06, THN^{+}06, WBC^{+}06, WPWF08, WSB^{+}06, Wil07v, Wil09s, WHM^{+}07, WPDZ^{+}06]. Skin-based [Bas07d]. skinny [Wel05-39]. Skp2 [RMC^{+}05, YWRA07]. Skps [LeB05-76]. slam [Sed08m]. slate [Wel05-69]. SLC3A2 [FZT^{+}07]. Sleuthing [Sed07b]. Sli15 [JhKC08]. Sli15/INCENP [JhKC08]. slides [LeB07-47]. sliding [GWS09, LeB06y, MT07]. Slik [CKG^{+}08, HF06]. slime [Les07-60]. Slingshot [NTY^{+}05, WHB^{+}07]. slippage [BYR08]. slips [LeB05-39]. Slit [HB08, SMSPK08]. sloppy [Les07-64]. Slow [LeB07-55, Les08-48, Les08-49, Wel06-52, KHWC08, LeB06-71, Tum06-42, Wel07a]. Slow-moving [LeB07-55]. slowdown [Les08-50]. slowly [HlQY^{+}06, Wel05-28]. slows [Wil07-56]. Slug [NFC^{+}08, MMS07]. sluggish [Wil09p]. Sm [AHM^{+}05, GTO^{+}07]. Sm-class [GTO^{+}07]. SMA [Wil07e]. SMAC [LMS05]. Smad [KKW^{+}09]. Smad1 [YBT^{+}09]. Smad2 [YBT^{+}09]. Smad3-dependent [ZBKJ07]. Smads [XYC^{+}07]. Small [Les07-65, PAV^{+}06, BLC05, DG08, HFF07b, MYMZ08, PNB07, SWG^{+}05, TTO^{+}08, Tum05v, VBL^{+}05]. smart [Wel05-61]. SMC1 [AJR^{+}09, NWR^{+}08]. Smell [Wel06-53, Nyb05k]. SMK [KHW^{+}07, KHW^{+}09]. SMK-1 [KHW^{+}07, KHW^{+}09]. SMK-1/ [KHW^{+}07, KHW^{+}09]. SMN [PGO^{+}07, PGO^{+}07, RGW^{+}07]. smooth [BYK08, GPC^{+}06, HLA^{+}06, SYA^{+}08, Xia06]. Smoothened [MSR09]. smoothes [Wel06-34]. SMRT [SLMK08]. SMSr [VTT^{+}09a]. Smurf1 [SGMPV07]. Smurf2 [ORM^{+}08]. snail [Sho09v, BYF^{+}05]. Snail1 [RLH^{+}09]. Snail1a [LHM^{+}09]. SNARE [CCS^{+}06, DXC^{+}09, FSP^{+}07, GEP08, HJ06, HAAHD05, JKKM08, SYM^{+}08a, SM09, SSP^{+}07, VBSM06, WP08, YNL^{+}08]. SNARE-dependent [JKKM08]. SNAREs [GHY^{+}05]. snatching [Pow06c]. sneaks [Wil06n]. snoRNA [ZKG^{+}08]. snRNP [Les06m]. Snx3p [SSSB07]. Soaking [Tum06-43]. Society [LP05, LK06, LW08, WL07, dSW09, Wil07c]. sodium [BH08a, BFB^{+}08]. Soleil [Ban09]. somatodendritic [YWK^{+}08]. some [LeB07-27, Sho08]. something [Rob08-39, Tum05-73, Wel06o]. Sometimes [Tum06-44]. somite [RHMW^{+}05]. Sonic [LYWK08, WMM^{+}09]. soon [LeB06t]. SopB [MES^{+}08]. Sopping [LeB08-29]. sort [DR07]. Sorting [LeB05-92, WvdLH^{+}07, API^{+}07, CG06, CSSE06, DHT^{+}09, EMPK^{+}07, FSP^{+}07, FN06, GC07, GKN^{+}07, JWG^{+}07, JWSS05, LJA09, LeB05-71, LeB07-35, LYH^{+}05, LDGK08, LNL^{+}08, MSECB05, MNH05, NZQ^{+}07, PCR^{+}07, PBB^{+}09, RHMW^{+}05, SOW^{+}09, Wel05-37, vBDF^{+}09, Sed09l]. Sos [KSK^{+}06]. Sos-mediated [KSK^{+}06]. sound [LeB08f, Wel06o]. source [Tum05i]. sources [Tum06-48]. Sox [GGE^{+}09]. Sox17 [SAD^{+}09]. Sox2 [MAH^{+}05, SAD^{+}09]. Sox8 [SSH^{+}05]. Sox9 [BDP^{+}07]. SP [BXN^{+}06]. space [HK07, SDI^{+}06, SPK^{+}09]. Spaced [Tum06-45]. spacing [PFD06]. span [KMS07, KHB^{+}05]. spares [LeB07-27]. spark [WBK^{+}06]. sparks [Mel06, Tum06c]. Spartin [EYB09, Sho09-31]. Spartin/SPG20 [EYB09]. Spastin [Wil07-59, EGR^{+}05, WEL^{+}07n]. Spastin-mediated [EGR^{+}05]. Spatial [BJLK^{+}06, CCC^{+}05, ELD^{+}08, HHT^{+}07, MIBT06, NTY^{+}05, WWS05b, BLR^{+}06, FKML^{+}06, HSLI09, MJG07, SH05, TTE^{+}07]. spatially [EMMN^{+}05, HCC^{+}07, KZC^{+}08]. Spatiotemporal [LFMO^{+}09, AWR^{+}07, HTYNK09]. SPB [Wel06-54]. SPDL [YWEK08]. SPDL-1 [YWEK08]. Special [LeB07-56]. Specialized [KL05, LeB06-69, GAS06, HXO^{+}07, XC08]. species [BKE^{+}07, MCA^{+}06, MAK^{+}06]. Specific [BJZ^{+}09, FHNG05, FHS^{+}07a, KKH^{+}06, NS06, BLH^{+}09, BLC^{+}09, BAR^{+}08, FCB^{+}09, FCB^{+}19, FSH08, GTX^{+}06, HLM^{+}08, HGJ06, HXZ^{+}07, JGAS08, JLB^{+}05, KLS^{+}06, KBC^{+}07b, LJA09, LJS^{+}09a, MM08, MGK^{+}09, OSS^{+}07, PCG^{+}05, RRA^{+}09, SN05, SSSB07, dGSWW^{+}06, TW07, TMBL^{+}08, VCSG07, XNM^{+}09, YKH^{+}08]. specifically [GGB^{+}08, KPM05, Rob08-43, SLZ^{+}06, WvNS^{+}07]. specification [AWE^{+}05, BBvD05, DLL^{+}08, Kat06j, KWS^{+}08]. Specificity [Ter06, KSpLL06, LLW^{+}09, NS07, RAS^{+}07, RAS^{+}08b, YNL^{+}08]. specifies [WNB08]. specify [LRJY06]. specifying [SAD^{+}09]. speckles [BGdN^{+}08]. spectraplakins [SRE^{+}05]. Spectrin [Bon05i, DBDD06, BBM^{+}08, HJB07, JRZ^{+}07, MPD09, YOHR07, PFD06]. spectrin/ankyrin/microtubule [MPD09]. spectrometry [TMBL^{+}08, Wil08g]. spectroscopy [HKV^{+}05]. speed [GSJ^{+}08, Kat06j, LeB06-47, LeB08-27]. Speeding [Tum06-46]. speeds [Wil07-56]. Speedy [Wil09p]. spell [Les07-41]. sperm [WNF^{+}05b, Tum05-65]. spermatogenesis [PMB^{+}09, SMM^{+}08a, WC07a]. spermatogonial [WLPB^{+}09]. spermiogenesis [GER^{+}07]. SPG20 [EYB09]. spherical [SWL^{+}05]. sphingolipid [FMA^{+}09, NPMZ06]. sphingolipids [KES^{+}09]. Sphingomyelin [VTT^{+}09a]. sphingomyelinase [ZJH08]. sphingomyelinase/ceramide [ZJH08]. sphingosine [SWH^{+}06]. Spill [LeB06-70]. Spin [Les08-35]. spinal [HFY^{+}06, JBL^{+}07, RGW^{+}07, SSJ^{+}07]. Spindle [Les08-35, MTRJA05, Rob08b, YKH^{+}08, ACsHS08, BJKF05, BH08b, BYR08, BBM^{+}07, BGP^{+}06, BBL^{+}08, CFU^{+}09, CGGT09, CLG^{+}07, DPP^{+}07, DHR^{+}08, FDM^{+}08, FDV^{+}06, GHM^{+}08, GNV05, GMZ^{+}08, GSV07, HH06a, HRB^{+}06, HBN^{+}09, JKDH08, JWC^{+}08, JMG^{+}06, JSFP08, LeB06-86, LeB08e, Les08i, LSC^{+}06b, LFW^{+}09a, LFW^{+}09b, LFMO^{+}09, LZ09, MPL^{+}07a, MS09, MD09, MAOM06, MDPG07, MHFH06, NRD^{+}06, NdCF^{+}05, OLH^{+}08, OLKK09, ORM^{+}08, Pow05o, Pow07b, RGL^{+}09, ROKG08, RBJL06, Rob08-29, Sed07f, Sho09-44, Sho09-42, SHC09, TTYH09, TLDB08, TKKT07, TWG^{+}09, UTK^{+}09, WS09, Wel06c, Wil07j, WF05, WF06, WF07, WOWB08, yYMM05, YCM^{+}08, YGR^{+}08, Yu06, ZCJ^{+}08, ZHYE08, Les09-33]. Spindle-independent [MTRJA05]. spindles [CBIO05, LeB05-83, TLDB08, UHM^{+}08]. Spindly [GSV07, LeB07l, CFU^{+}09]. Spine [LeB08-30, BMRP07, GTX^{+}06, HLS^{+}09, HZR05, IDTT^{+}09, LLHH07, NBS^{+}08, PKJS^{+}08, SSD06, TSN^{+}07]. Spines [CFBM05, LeB05-93, LeB06-71, Bas07c, Wil07p]. Spinning [Wel06-55]. spinogenesis [CHH^{+}08]. Spinophilin [BUJ^{+}08, SX08]. spirals [IPM^{+}05, Tum05-37]. Spire [QHB^{+}07]. Spitz [LeB06-72]. spliceosomal [PNB07]. spliceosome [Rob07h]. Splicing [CVACF07, PNB07, ELD^{+}08, KHBV05, PAM^{+}07, Rob08g, SCF^{+}07, Wil08b]. Splicing- [CVACF07]. Splicing-independent [PNB07]. Split [Wel05-59, Les07-70, Rob08w]. Splitting [Les07-66]. spoke [DS07]. Spokes [LeB08-31, YOY08]. spondin [ZMA^{+}07]. spontaneous [EFG^{+}05, JMB^{+}07, LeB07-60, MSLS07]. spontaneously [YWJ^{+}07]. Spore [TKK^{+}05]. spores [Tum05o]. sport [LeB08-33]. spot [Wel05-29]. spots [LeB05-31]. spread [Kat06b, LeB08t]. Spreading [LeB06-73, FSG^{+}07, HOM^{+}05, Wel05t, Wil07-50]. spreads [Sho09-50]. Spred2 [MYMH09]. spring [LeB08e]. springs [Wel06-69]. sprinkled [Pow05v]. spur [Les07b]. spurt [Les09w]. Spying [Too05]. SQSTM1 [BLB^{+}05]. squamous [KNKH06, LBH^{+}05]. squeeze [Les09-30]. squeezed [Kat06d, LeB06-83]. Squeezing [Wel05-60]. SR [BJZ^{+}09, LeB08-32]. Src [BCT06, CTY^{+}06, CCP^{+}08, DBR09, GGEd^{+}07, MKV^{+}09, MDW05, SYKM08, SDD^{+}07, VSR^{+}08, ZTZ^{+}09, ZKR^{+}07, OIT08]. Src- [DBR09]. Src-induced [SYKM08]. Src-like [MDW05]. SRC1 [RW08c, GFC^{+}08]. Sro7p [GAG^{+}06]. SRP [JCMG08, LeB05-63, Les05b]. Srs2 [BLA^{+}09]. Ssy1p [WOP^{+}06]. Stability [LeB05-94, BLC05, BHK^{+}06, CKG^{+}08, CZD^{+}09, DJV^{+}05, EHHK09, GOM^{+}09, HBD^{+}05, HBN^{+}09, LeB07-51, Les09m, MPD09, POG^{+}09, PGO^{+}07, PPBF08, SSP^{+}07, Sho09j, SLR^{+}06, TLDB08, TMH^{+}07]. stabilization [LYV^{+}08, PMC08, SBF^{+}06, WNB08, YGR^{+}08, qGyWX^{+}08]. stabilize [BKHR05, HT08, VWK^{+}06, YKBR09]. stabilized [DMDO08, Heu05, Rob07c]. stabilizer [Wil09e]. stabilizes [BMS^{+}08, HCK^{+}07, SDC09, Tum06-47, WCB^{+}08, ZKG^{+}08, ZMH^{+}09]. Stabilizing [Tum05-66, Wil07-60, LeB06-39]. Stable [CNM^{+}05, FVSC08, FvD08, ASH^{+}09, OF08, VMSV08, WZM07, Wil07b]. stably [GNE^{+}07]. stage [GBNS07]. stage-dependent [GBNS07]. stages [HXZ^{+}07, MBL^{+}07, OYM^{+}09, SYA^{+}08, SCK^{+}07, TCKV05]. stalk [LLW^{+}09]. stalled [Tum05-67, YM09]. Stalling [Wil07-61, FPP^{+}08, Pow07j]. stalls [Tum05-67]. stand [Les08-51]. Staph [Tum05s]. Staphylococcus [BDR^{+}06, SDS^{+}05]. Star [Rob08u]. start [Wil07f]. Starting [Les07-67, Hac07, Hat05a, Les07-60]. starts [Sho09-43]. starvation [HCB^{+}08, KWU^{+}05]. starvation-induced [HCB^{+}08, KWU^{+}05]. starving [Rob08-38]. STAT [KBN^{+}06, WLC08]. STAT1 [SMW^{+}07]. STAT3 [KP08, LWC09, SMW^{+}07, NLL^{+}06, Tum06-47]. STAT5 [DRD^{+}05, XNM^{+}09]. state [OD07, PHB^{+}07, PWD^{+}06, QWD^{+}08, SWT08, Yu06]. states [CBMCSS08, LLS^{+}09, PMS^{+}07]. stathmin [NLL^{+}06]. stationary [ABA^{+}06]. stationary-phase [ABA^{+}06]. status [WMS^{+}06]. Staufen2 [GTX^{+}06]. stay [Les09a]. Staying [Wel05-61, Wel05y, Wil09g]. stays [Les08-33, Tum06l]. Ste13p [FN06]. steady [SWT08]. Stealing [LeB06-74]. steer [WHB^{+}07]. steering [OFY^{+}05, Wel05h]. steers [LeB07-44]. Stefan [Sed07e]. Stem [LeB08-33, Les07-68, Les08-50, Les08-51, BTD^{+}07, DL08, DLT^{+}07, DMC^{+}07, EVF^{+}08, Fuc08, HKN09, HMG^{+}08, Hor09, JKCW07, KMK^{+}05b, Kat06g, KAD^{+}08, KLW^{+}05, LeB05-107, Les06a, Les07-79, Lin08, LVG^{+}09, MG08b, MOM^{+}06, NCB^{+}08, NCB^{+}09, OLS08, SBC^{+}05, Sed09i, SCFKB07, SKR^{+}07, SRXvdK06, SAD^{+}09, SRQL09, TMW^{+}05, TCKV05, VFRJ^{+}07, WLC08, Wel06-59, WMSB^{+}06, WPDZ^{+}06, WLPB^{+}09, YF08, ZXM^{+}06, ZNP^{+}09]. step [KOVP09, Les07-75, MSW^{+}07, MSW^{+}17, Tum05q, Tum05-74]. Steps [Wil07-62, BM07, EVH^{+}07, EVH^{+}08, KIKI^{+}07, LeB05-39, Les07z, MK07, NNK^{+}08, RCS^{+}09, SBJ^{+}07, Tum05v, WP08, ZWO^{+}09, dPMF^{+}09]. Stepwise [DKR^{+}06, Wel06-56, JRZ^{+}07]. stereoisomer [SHjC^{+}07]. Steric [BLC^{+}09]. steroid [CLL07, YTB07]. steroid-induced [CLL07]. steroid-triggered [YTB07]. sterol [RIHP06, SCR^{+}09]. sterols [KES^{+}09]. Stick [LeB07-57, Les09v, Wel06-40]. sticker [Sho09i]. Sticking [Pow05t, Tum05-68]. sticks [Les09l]. Sticky [LeB07-58, Tum05-69, Pow05u]. Stiffening [LeB05-95]. still [DR07, Tum05r, Wel05y, Wil07-27]. STIM1 [LWBL06, RDY^{+}05, WBLL06]. stimulate [MG08a]. stimulated [ANI^{+}07, LGH^{+}05, SSD^{+}09, RG06a, SLWA07, WM05]. stimulates [CZSG08, HPT^{+}07, LNPC06, LMZ^{+}05, MRD^{+}08]. stimulating [GSdBS09, LeB06-75]. Stimulation [CCC^{+}06, LMC^{+}09, TDP^{+}06, Wil07-62]. stimulatory [DLL^{+}08]. stimuli [WvNS^{+}07]. stimulus [KOKM07]. stimulus-induced [KOKM07]. Stirring [Wil08x]. stochastic [WLPB^{+}09]. stock [EGL^{+}09]. stoichiometry [CHSV07, GBNK08a, GBNK08b]. Stomata [LeB06-76]. stonin [JWG^{+}07]. stop [LeB05n, LeB05-80, Sho09-49]. stop-loss [Sho09-49]. Stopping [Les09-34]. stops [LeB06g, LeB06m, Les08-47, Nyb05c, Sed08d]. storage [DNK^{+}05, PLLH05, TDG^{+}06]. store [CLK^{+}07, LWBL06, RDY^{+}05, WBLL06]. store-operated [LWBL06, RDY^{+}05]. Stra13 [SLV^{+}07]. Strabismus [LS09]. straighten [Les07f]. straightjacket [LYV^{+}08]. strain [AFM^{+}05, Sho09-47]. strand [BJLK^{+}06, BFCS^{+}07, DCA^{+}06, DG05, KMK^{+}05b, KG05, KCD^{+}06b, MG09, PGW^{+}08, TFC^{+}06b, UWiY^{+}07, YMTC05, YEL^{+}08]. stranded [GTX^{+}06]. strategies [AS09, LeB06k]. strategy [DvRF^{+}05b, KP08]. stray [Les07-65]. strays [Les07-77]. strength [LeB05-106, MJF^{+}07, ZDMD06]. strengthening [AFM^{+}05]. Stress [HL06, HCC^{+}07, KSA^{+}05, LeB07-59, Tum06-48, UZ08, BMP08, BWFdL09, DSLY^{+}05, ESZK06, FTB^{+}09, GSJ^{+}08, GPC^{+}06, KRZ^{+}08, LeB05-34, LMC^{+}09, LHRP05, MLV09, MSK^{+}07, MLH^{+}06, NSM05, OMGW^{+}06, SPT^{+}09, SPT^{+}21, SNA^{+}07, Sho09-27, Sho09-51, SL05, SLK^{+}08, SLR^{+}06, TP09, TNS07, Tum05c, Tum06-43, WCB^{+}08, Wel06s, Wil07-29, Wil07-40, Wil09c, WGL^{+}06, YIfHA09, YOSM06, ZPGB^{+}07, KIKI^{+}07]. Stress-dependent [HCC^{+}07]. stress-induced [WCB^{+}08, YIfHA09]. stress-mediated [SLK^{+}08, TNS07]. Stressful [Wel06-57]. Stretch [Wel05-62, CMF08, MS09, Rob08-29]. stretch- [CMF08]. Stretchable [Bas07e]. Stretching [Tum05-70, UTK^{+}09]. striatal [WMSB^{+}06]. striated [BYK08, MSV^{+}05, RGW^{+}07]. striatum [DCAC05]. strict [PVM^{+}05]. strictly [FD09]. strike [Sho09-41]. strikes [Let09]. stripe [BLC05]. stripping [KRPH08]. strokes [YOY08]. strong [Rob07j, Sho09-28]. Structural [CRV^{+}06, HJ06, LSC^{+}06b, MSH^{+}05, MWSK06, NOH^{+}07, SRE^{+}05, SZX08, WH06, Yu06, CHSV07, CDW^{+}05, GHM^{+}08, JYZL06, Joh09, KKW^{+}05, MSC^{+}07, PWD^{+}06, RXA^{+}08, SDW^{+}09]. Structurally [RWB^{+}05, FSS^{+}09, IPM^{+}05, KBS^{+}06]. Structure [POGL08, Sed07f, AXM^{+}05, ASH^{+}09, BLC^{+}09, DMDO08, HGR^{+}06, HSY^{+}06, HKP^{+}09, JNHB08, KPR^{+}07a, KPR^{+}07b, KCD^{+}06b, LeB05f, MMC08, MWSK06, OWM05, SB09, SMRS^{+}07, WS09, Wel05-31, Wel06-60, WKCY^{+}08, XMA^{+}09, iYOW^{+}06]. structure-specific [BLC^{+}09]. structures [DDLC^{+}06, GMG^{+}07, IGFG^{+}08, OHK07, TGH^{+}07]. structuring [ADO07a, ADO07b]. Stt4 [BSA^{+}08]. Stu2p [ABvBHH06]. Stuck [LeB05-96, Les08-52]. study [DGG^{+}08, Sed09m, SFEK07, SFS^{+}07]. stunts [LeB07-59]. style [Les06-28, Tum05-34]. subapical [RKH^{+}09]. Subcellular [WXM^{+}05, CBK^{+}06, LZL^{+}05, LMHT08]. subcompartmentalization [VBNR06]. Subcompartments [MKH^{+}07]. subdomains [HGR^{+}06, PCN^{+}05]. subendothelial [OSB^{+}05]. subnuclear [GLX^{+}09a, GLX^{+}09b]. subset [FSS^{+}09, KHBV05]. substrate [CAL^{+}07, GGH^{+}08, MGBH^{+}06]. substrates [VBL^{+}05, WHR^{+}09]. subtelomeric [GFC^{+}08, TFC^{+}06b]. subunit [CYY^{+}08, KVC^{+}09, LGBD^{+}07, LFL^{+}05, LYV^{+}08, MFC^{+}09, ToMY^{+}08, ZWO^{+}09]. subunits [ÇB09, MSM^{+}07, XLH^{+}07]. subventricular [JAV^{+}07]. subvert [INB06]. success [Rob08v, Wil08n]. sufficient [HBD^{+}05, LLL^{+}06, SWB^{+}07, THL^{+}06]. sugar [Sed09c]. sugar-handle [Sed09c]. Sugars [Pow05v]. suggests [MMC08]. suicide [Wel05-58, Wil07g]. sulfate [FWC^{+}07, GS07, KKH^{+}06, KSpLL06]. sulfatide [LLM^{+}05]. sulfatides [DGZ^{+}05]. sulfhydryl [SMM^{+}08c]. summits [LeB09c]. Sumo [Rob08-45, LM05, LFH07, Sed09b, YWM^{+}08]. SUMO2 [MAK^{+}06]. SUMO2/ [MAK^{+}06]. Sumoylation [MHFH06, ZS08, MPL^{+}07b, WZM07, CHH^{+}08, HAN^{+}09]. SUN [RH09, Tum05-71, WL07]. Sun1 [LPM^{+}07]. Sunday [CKKG05]. sunlight [Wel06g]. Super [HW09]. Super-coil [HW09]. superoxide [KNT^{+}07, MHM^{+}05, MWY^{+}09, MSH^{+}05]. Supersensitive [LeB05-97]. Supersized [Les08-53]. Supervillin [TSN^{+}06]. supper [Pow05k]. supply [KOKM07]. Support [KMK^{+}05b, GHM^{+}08, Les08-34]. supports [BGZ^{+}07, LeB05-47, Tum06g]. suppress [ASK^{+}09, JMB^{+}07]. suppressed [SIL^{+}07]. suppresses [KTI^{+}05, MSM^{+}06]. suppressing [OLLG05, WC09]. Suppression [ERM^{+}09b, WZM^{+}08, Di 08, GSiK^{+}09, KNKH06, LSdWL07, XJF05]. suppressor [BvOM^{+}07, BIH^{+}05, BIH^{+}07, HF06, KWY^{+}06, Les06d, LBS^{+}08, SGB^{+}06, ZHS06]. supramolecular [HBD^{+}05]. surface [AM05b, BLC05, CZM^{+}07, CBG^{+}06, FBO^{+}07, FHMC09, Fuc08, Heu05, HIK^{+}08, JKDH08, LeB08-34, NN07a, PPCZ06, SWT^{+}05, TKB^{+}06, YHD^{+}09]. surprise [Rob07h]. surprises [Wel06-67]. surveillance [BJLK^{+}06, LeB05-65, LFH07, Wil08d]. Survival [MZ08, BHK^{+}06, BWFdL09, DFB^{+}07, KR07, hKDH^{+}08, LSC^{+}07, Mal06, ML05, MOM^{+}06, NJS^{+}06, NVEF07, PhXD^{+}06, PGO^{+}07, SCFKB07, STMG05, Tum05-71, VVL^{+}08, XSVGR05]. survive [LeB06-49]. Survivin [YCX^{+}08]. Surviving [Wel06-58]. susceptibility [BXN^{+}06, CFH^{+}08]. SUSP1 [MAK^{+}06]. Sustained [XNM^{+}09, BJLM^{+}05, HSB^{+}07, MCRVC^{+}08]. sustaining [PMS^{+}07]. sustains [BCT06]. Suv4 [BGJ^{+}07]. Suv4-20h [BGJ^{+}07]. Suzanne [Sed09l]. SV40 [SHjC^{+}07]. Sweeter [Rob08-46]. swelling [KPM05]. switch [BWL^{+}08, CCSBM^{+}07, FOB^{+}09, FSG^{+}07, HM08, JTN^{+}05, LeB06-36, LeB07q, MAG05, PEER^{+}05, PKK09, SAD^{+}09, SSAD08, SWM^{+}06, UFY^{+}07, WvdLH^{+}07]. switches [KO07]. Switching [Sal08, SSZR07, RYD08, SKO^{+}08, SWS^{+}08, WZO^{+}05, Wil08q]. SYCE2 [BFCS^{+}07]. SYCP2 [YDL^{+}06]. Syk [ZKR^{+}07, CLK^{+}07]. symmetry [LeB05v, LeB07-31, Nyb05b, PvdGS06, YWJ^{+}07]. sympathetic [HGJ^{+}08, SHX^{+}07]. Sympathy [Wel06-59]. symptoms [WZM^{+}08]. Synapse [Les08-54, AACS^{+}06, BRKR06, Dal09, IKZ^{+}07, KKW^{+}07, LeB07-32, Les08-29, LJY^{+}07, MPD09, Pow05h, SX08, SG09, Sho09f, STP08, TD08, Wel06-61, tDAK^{+}05]. synapse-associated [LJY^{+}07]. Synapses [LPN^{+}08, BMK^{+}07, BUJ^{+}08, JYZL06, LeB05-51, Les08-49, Les09-34, RWB^{+}05, RBC05, TLK^{+}06, WWO^{+}08, YJJ^{+}09]. Synapsin1a [LOWTL^{+}08]. synapsis [JGY09, YDL^{+}06, YEL^{+}08]. Synaptic [IDTT^{+}09, Wel06-60, AEJ^{+}06, BRKR06, BJML09, DXC^{+}09, FDB^{+}09, GSW^{+}08, HJ06, JWG^{+}07, KKM^{+}08, KKW^{+}07, LOWTL^{+}08, LJY^{+}07, LYV^{+}08, MSLS07, MPG^{+}09, NFI^{+}09, OVL^{+}07, Tum06-40, Wel05p, XLH^{+}07, YJJ^{+}09]. synaptogenesis [MAG05, NBS^{+}08, Tum05b]. synaptonemal [BFCS^{+}07, YDL^{+}06]. Synaptotagmin [AOT06, BJML09, CSB^{+}06, MSLS07]. Synaptotagmin-12 [MSLS07]. Synchronizing [MPLF^{+}09]. syncytial [FML^{+}06, PMW05]. syndecan [BMR^{+}08, DLBSD^{+}05, ERM^{+}09b, HTKR06, MBR^{+}06, BRM^{+}07, LLHH07]. syndecan-1 [DLBSD^{+}05, MBR^{+}06]. Syndecan-2 [LLHH07]. syndecan-4 [BMR^{+}08]. Syndecan-4-dependent [BRM^{+}07]. syndrome [CMK06, PPBF08, YMTC05]. synectin [ERM^{+}09b]. Synip [YOS^{+}05]. Syntabulin [CGS05]. Syntabulin-mediated [CGS05]. syntaxin [GEP08, HWSJ08, SLM^{+}06, VBSM06]. Syntaxin5a [SSP^{+}07]. Syntaxin5a/Sed5 [SSP^{+}07]. syntelic [YKBR09]. synthase [BYF^{+}05, FHFB^{+}06, VTT^{+}09a]. synthase-related [VTT^{+}09a]. synthesis [DTW06, GEE07, HNvD^{+}07, KC09, Les09-27, LCM^{+}08, PKJS^{+}08, SFP^{+}05, SGK^{+}07, SFB09, TPA^{+}08, iYOW^{+}06]. synthesized [MG08a]. synthetic [HS09a]. Syntrophin [LCL^{+}05]. synuclein [FCDM^{+}07]. system [BTP^{+}07, BMT^{+}07, CWNF^{+}09, DR05, FML^{+}06, GKCM^{+}08, HDR09, LeB06-55, MGW^{+}08, MK07, PBB^{+}09, TM06, WWO^{+}08, Wel07f, YGW06, YBK^{+}06, ZTM^{+}08]. Systematic [DZW^{+}08, LeB05-98, BMBD^{+}09]. systems [HS09a, HR05, PvdGS06, RM05b, SSZR07, Tum06-34]. Syt [LeB07-60]. Syt-12 [LeB07-60].

T [LeB06-77, Rob08-47, BFF^{+}06, FBD^{+}05b, FBY^{+}06, GCF^{+}05, GMvdKC07, KDV^{+}09, LeB05z, LeB08q, MCRVC^{+}08, MDW05, NNS^{+}08, SCS^{+}05, WSS^{+}06, Wil09j, ZDMD06, SSP^{+}07, VBSM06]. TACC [BKHR05, PSS^{+}05]. TACC3 [KNP^{+}05]. TACC3/maskin [KNP^{+}05]. tactics [Wil07o]. Tae [Wil08v]. tafazzins [CMK06]. Tag [Les08-55]. Tagging [Les05i]. tail [HM08, Les08v, RW08a, WHC^{+}07]. tailored [IPM^{+}05]. tails [BKBM06]. Taipale [Sed08i]. Take [Rob08-48, LeB05-61, LeB05-60, LeB06-56, Les07-36, Sho09-47, Wel05-43]. takes [LCGRB09, Les08-42, Les09-32, Sed09d, Sho09-30, Sho09-54, Tum06v]. Taking [Kak08, Les06-27, EGL^{+}09, Sho09o, Sho09-34, Wil07-45]. talin [HWS^{+}07, LVL^{+}05, SMH^{+}09, SCS^{+}05, WBP^{+}08]. talin-dependent [SCS^{+}05]. talk [BPL^{+}05, YTH^{+}09]. Tam41 [KRG^{+}08, THY^{+}06]. tamalin [EYB^{+}06]. Tandem [LeB06-78]. tango [LCGRB09]. Tankyrase [Wil09q]. Tanning [Wil07-63]. target [BBG^{+}05, GCSG^{+}05, NFC^{+}08, PMB^{+}09, PCMSB09, SN05, SWB^{+}07]. targeted [ACsHS08, HCN06, HIK^{+}08, KS08, PPCZ06]. Targeting [HAAHD05, tKJCH06, ASH06, BH08a, CLA^{+}05, GBNS07, HHT^{+}07, JKW^{+}09, KFR^{+}08, LSNY08, LS06, LPB^{+}08, MDW05, NFI^{+}09, OVL^{+}07, RAP08, SLM^{+}06, SPK^{+}09, VCD^{+}06, WBT^{+}08, WF05, WXM^{+}05, XLH^{+}07, YWK^{+}08, YHD^{+}09]. targets [FCDM^{+}07, HCRT08, HAT^{+}05b, LCvD^{+}08, MOS^{+}09, OLC^{+}07, SCVW09, SDD^{+}07, TAK^{+}09, TST08]. Tarun [Sed09m]. tasks [Les07-73]. Tat [BM07, DR05, Tum05-58]. Tau [KKB^{+}06]. Tau-dependent [KKB^{+}06]. Taxol [YKBR09]. TB [Wel05-63, Wel06g, SWM^{+}06]. Tbx5 [CBK^{+}06]. TCF [qGyWX^{+}08]. TCR [MDW05, LeB05f, MDW05]. TCRs [Tum05-72]. team [Les08-55]. teamwork [BW06]. tears [Tum05-77]. Tektin [DHR^{+}08]. Telomerase [SCFKB07]. Telomere [BMSJ07, hKDH^{+}08, TFC^{+}06b, BGJ^{+}07, CS09, TJC06, TSALS05, Wil09q]. Telomeres [LeB05-99, AJR^{+}09, CYO^{+}09, HS09b, JZN^{+}09, LeB07m, LeB08-33, Les07s, Les08-55, Pow06a, SCFKB07]. telomeres-associated [JZN^{+}09]. telomeric [DMBF^{+}07]. Tem1 [GSdBS09]. template [HKP^{+}09, LeB05-78, Yu06]. Temporal [KKM09b, MTV^{+}05, GHK^{+}09, LWG^{+}05, NTY^{+}05]. Temporally [LGH^{+}05, EVH^{+}07, EVH^{+}08, GAS06]. temporary [SFS^{+}07]. tendons [Les06r]. tense [Les07-78]. Tension [Bon05j, Wil07-64, GPC^{+}06, HHZ^{+}08, MD09, OLH^{+}08, Rob08k, SRN^{+}07, WL09a, WL09b, WF06]. tension-dependent [GPC^{+}06]. term [LLL^{+}06, RCS^{+}09, SCFKB07]. terminal [BWJ05, BEN^{+}05, KNKH06, MM05, VBSM06, WPWF08, CWM^{+}09, HWN^{+}07, IPAR05, KVS^{+}09, SDS^{+}05, TdVM^{+}07, WEL^{+}07n]. terminals [JBL^{+}07, XLH^{+}07]. Terminating [TG05]. Termination [GRC^{+}05, BPL^{+}08]. terminus [BHG^{+}05, MBJF^{+}08]. ternary [SMM^{+}08c]. Terry [LeB07-61]. test [RAP08]. testing [Wil07-64]. Tetanus [DBW^{+}06]. Tether [Wel06-61, FSS^{+}09, KHWC08]. tetherin [RKH^{+}09]. tethering [LU05, LMZ^{+}05, STBL09, TFC^{+}06b, YSP^{+}06]. tethers [HSK^{+}05]. Tetrahymena [KPR^{+}07a, KPR^{+}07b]. tetramer [BLC05]. tetraploid [CGK07, Wel05x]. tetraspanin [BZYM^{+}08, EMD^{+}08, LZ08, MWSK06, NKK^{+}06, SKvdBW^{+}06]. tetraspanin-enriched [LZ08, NKK^{+}06]. tetraspanins [DPMB^{+}07]. TEX15 [YEL^{+}08]. TGF [ADC^{+}07, CTGVGFF08, CCM^{+}07, GG06, JGAS08, KKRP05, KLV^{+}08, LD07, Les07-38, NPV^{+}07, SBC^{+}05, SOC^{+}07, WRMH07, WWC^{+}07, XYC^{+}07, YMD^{+}07, BFG^{+}06b]. TGF- [ADC^{+}07, CTGVGFF08, KKRP05, LD07, NPV^{+}07, SBC^{+}05, SOC^{+}07, WRMH07, XYC^{+}07]. TGN [BBPM07, NFCF08]. thaliana [PST^{+}06, VCN^{+}05]. Their [Les08x, BFB^{+}08, CF07, CK07, ÇB09, FFS06, GFC^{+}08, HMM^{+}05, Kat06g, LeB06q, LeB08q, LSM^{+}05, Les07-28, Les07-79, Les08d, Les08-40, Les08-37, Les08-66, Les09b, NTSY08, Sho09k, Sho09x, Wil07-38, Wil07-39, Wil09j]. theme [RMV06]. themselves [Les08-51, Les08-64, Sho09y]. therapy [Tum06t]. there [BH08a, Tum05-73, Tum06-49, Wel05-64]. thick [CBG^{+}09, Les09x, MQL^{+}06]. thickens [KC09]. thin [BLL^{+}06, CBG^{+}09, Les09x, MSV^{+}05]. things [LeB06-65]. think [Ter06]. Thinner [Wel05-65]. Thinning [LeB05-100]. thiol [TRS^{+}08]. Thomson [RVH08]. those [Wel05-64]. thought [Sho09a]. Three [AXM^{+}05, KOM07, MMC08, MFM^{+}06, NPBP07, OHK07, ZRBS07, DWMY09, Les07z, MSW^{+}07, MSW^{+}17, RAS^{+}07, RAS^{+}08b, RLH^{+}09, SSHW09]. Three-dimensional [AXM^{+}05, KOM07, MMC08, MFM^{+}06, NPBP07, OHK07, DWMY09, RLH^{+}09, SSHW09]. three-step [MSW^{+}07, MSW^{+}17]. threonine [WHC^{+}07]. threshold [PMK^{+}05, RKT^{+}07]. thrombospondin [SDN^{+}05b]. thrombospondin-1 [SDN^{+}05b]. thromboxane [PWM^{+}06a]. throughout [BGP^{+}06]. Thrown [Les07-69]. thylakoid [DR05]. Tiam1 [HOM^{+}05, MRO^{+}05]. Tid1p [KKGD05]. tide [Wel07m]. Tie [WKC^{+}09]. Tie-2 [WKC^{+}09]. Tie1 [SKE^{+}05]. Tie2 [CNM^{+}05, SKE^{+}05]. ties [Sho09t, Sho09-42]. TIF [GEE07]. TIF1 [TPZG06]. Tight [DG08, Les08-56, MMT^{+}05, Bon05c, CM06, Les09a, MRO^{+}05, PCN^{+}05, Pow05t, SWT08, SSM05, Tak06]. tighten [Tum05-36]. tightly [PBD^{+}06]. Tilted [Nyb05l]. Tim [SFB09]. TIM23 [THY^{+}06, THS^{+}09]. Tim50 [THS^{+}09]. Tim54p [HCL^{+}07]. Time [Kat06i, Les07-70, Sho09-52, Wel05-66, BMB^{+}07, PH07, Sho09-48, WNF^{+}05b, YTH^{+}09, YM06]. timely [Les07t]. timer [RGL^{+}09]. times [BEH^{+}06, MSO09]. Timing [Les08-57, CLL07, GCM09, GLH^{+}05, HMG^{+}08, SLCB05, Tum05f, WSG06]. TIMP [SBF^{+}06]. TIMP-2 [SBF^{+}06]. TIMP-3 [SBF^{+}06]. TIN2 [hKDH^{+}08]. TIN2-containing [hKDH^{+}08]. tiniest [Rob08o]. tiny [LeB05-33]. tip [GFD^{+}05, LSNY08, Wil09b, Vau05, Xia06]. TIP47 [ASP06, BDW^{+}09]. Tip60 [KDD08]. Tipin [SFB09]. tips [LeB05k, MRvD^{+}05]. tiRNAs [Wil09c]. tissue [BLH^{+}09, CLL07, CG06, JSC08, PGP^{+}07, SBM^{+}07, Sho09-36, TTK^{+}06]. tissue-specific [BLH^{+}09]. titin [WBL^{+}06]. Titinless [LeB06-79]. TLR [BSH^{+}07, WPC^{+}08]. TLR3 [BSH^{+}07]. TLR9 [SRL^{+}06]. TLR9-independent [SRL^{+}06]. TNF [GPN^{+}06, MKH^{+}07, NEF^{+}06, PDM^{+}08, VCC^{+}08]. TNF- [PDM^{+}08, NEF^{+}06]. TNFR1 [ZTKK05]. Tob55 [HWN^{+}07]. toes [Kat06g]. together [GNL^{+}06, Les07-46, Les08u, Les09y, Rex09, Rob08s, Rob08y]. Togetherness [Les06-28]. told [LeB06-77]. Toll [MLC^{+}06, SRR^{+}08, Les06-27]. Toll-like [MLC^{+}06, SRR^{+}08]. TOM [Rap05, SGKP^{+}07a, SGKP^{+}07b]. tomographic [PGL^{+}09]. tomography [MFM^{+}06, NPBP07]. tomosyn [BSMT05, SYM^{+}08a]. tone [HGJ^{+}08]. Tony [Wil08w]. Too [Rob08-50, Bon05d, LeB06t, Wel05d]. tool [Wil06g]. toolkit [RSL06]. Tooth [TT09]. top [Sho09-34]. TopBP1 [YM09]. topogenesis [KMS07, SDI^{+}06]. topography [DWMY09]. topological [JLB^{+}05]. topology [BXHD08, Wel06b]. TOR [GEE07, HCN06, JHY^{+}08, WLE09]. TOR-mediated [WLE09]. Toronto [Wel06o]. torsinA [GD05]. Tortoise [Wel06-62, LeB05-89]. total [MYO^{+}05, Sed08e]. Totally [Nyb05m]. touch [Les07-52, Rob07m, Sho09-38]. toxicity [IPC09, MSH^{+}05, SNG^{+}05, WZM^{+}08]. toxin [ALvdG06, DBW^{+}06, FHS^{+}07a, GMG^{+}07, SZL^{+}09]. toxins [RHMS05]. TPX2 [BH08b, OW06]. traced [WPDZ^{+}06]. track [LDVV06, WTH05]. tracking [SFEK07, SFS^{+}07, Sed09i, Sed09h]. Tracks [Wel06-63, BKLT^{+}08, LeB08-27, Les07-76, Wel06-29]. Traction [GSJ^{+}08]. TRAF2 [VCC^{+}08]. Traffic [HFP^{+}05, Les08-58, BFG^{+}06b, BBF^{+}07, CZP^{+}05, DP08, GOM^{+}09, GT08, LeB08o, Les09q, LGV08, PAV^{+}06, TTK^{+}06, Wel06q, Wil09r]. Trafficking [LeB08-34, AWR^{+}07, BHY^{+}08, BCR^{+}08, CWC^{+}08, DGZ^{+}05, FDB^{+}09, FN06, FCB^{+}09, FCB^{+}19, GMG^{+}07, JLC08, KP08, KBR^{+}08, LBC^{+}07, MW06, NHK^{+}08, NFCF08, PLLH05, SS06a, Sed09l, TMH^{+}05, Tum05-78, Tum06-32, VLFA^{+}07, WRP^{+}06, WBE^{+}08, Wel05v, WP08, WSB^{+}09, XGX^{+}09, ZL05]. traffics [GDR^{+}05]. trail [Les09f, Sed08c, Wil09h]. Trailer [DP06]. trailing [KBR^{+}08]. trains [PGL^{+}09]. trans [GEP08, KVS^{+}09, KES^{+}09, LeB05-46, NSM^{+}09, Poo09, WHOS06, vBAS^{+}07, vBDF^{+}09, SM09, Sho09-32]. trans-binding [KVS^{+}09]. trans-endothelial [vBAS^{+}07]. trans-Golgi [GEP08, KES^{+}09, WHOS06, vBDF^{+}09]. trans-membrane [Poo09]. transactivates [SWH^{+}06]. Transcribing [Rob08-51]. transcript [LeB06-78, PS08]. Transcription [LeB08-35, MTRD^{+}06, McN09, AG09, BYF^{+}05, BZA^{+}06, BBG^{+}05, CMD^{+}08, CDW^{+}05, CVACF07, DKR^{+}06, GGE^{+}09, HMS^{+}08, HKP^{+}09, JBYW06, KBN^{+}06, LG07, LG08, MMS07, MOM^{+}06, MRK^{+}07, NTB^{+}06, PS08, PV06, RGA^{+}06, SSH^{+}05, TPZG06, Tum06m, WBG^{+}08, Wil07-30, XC08, YAV^{+}05]. Transcriptional [AWE^{+}05, HlQY^{+}06, Wel06-64, BMB^{+}07, DTR^{+}05, KTI^{+}05, KS08, LeB06-81, LTDH05, MGR^{+}05, SLMK08, STL^{+}07, Sho09-35]. transcriptionally [HHY^{+}08]. transcripts [LWHC08, PNB07]. Transcytosis [AMS^{+}05]. transcytotic [TMY^{+}09]. transduce [KDV^{+}09]. transducing [KNT^{+}07]. transduction [BKE^{+}07, CKX^{+}08, Les06o, SMH07, SL05]. transepithelial [KSR^{+}08]. transfer [BUJ^{+}08, FHS^{+}07b, Lev07, SCR^{+}09, Sho09-37, VCN^{+}05, WRKC^{+}09]. transferase [OVL^{+}07]. transferases [FHFB^{+}06]. transfers [Lev07]. transform [KGDR^{+}07]. transformation [BHJ^{+}05, PVJJ07, YMDL09]. Transformed [LeB05-101, DHML05]. Transforming [TVP^{+}06, Wil06o]. transforms [GCSG^{+}05]. transgenic [RP05, YAV^{+}05]. Transient [CTY^{+}06, Wil07-65, BDR^{+}06, WJM05]. transiently [SFEK07, SFS^{+}07]. transients [PMW05]. transit [CCC^{+}06]. transit-peptide [CCC^{+}06]. Transition [RM05a, GHK^{+}09, HHY^{+}08, IPG^{+}05, LD07, LDKT06, LJM^{+}07, MSO09, MMS07, TVP^{+}06, WWC^{+}07, ZNP^{+}09, ZMH^{+}09]. transitions [HJ06, PWD^{+}06]. translated [LeB08-35]. translates [MPL^{+}07a]. Translation [LeB07-63, SGK^{+}07, KII^{+}08, LeB06-78, LSC^{+}07, Les05d, MGBH^{+}06, MKSJ06, PMB^{+}09, RSSC06, WGKA^{+}08]. Translational [NMZ^{+}09, WIL^{+}05, XKYW07, YIfHA09, dPMF^{+}09]. translationally [HCC^{+}07]. translocase [KGM^{+}07, Poo09, WvdLH^{+}07]. translocating [COC^{+}05, KMS07]. Translocation [DDH09, BYMB06, COC^{+}05, CJMG05, DGG^{+}08, DR05, FM05, HNvD^{+}07, IBS^{+}05, JCMG08, KBN^{+}06, LWC09, MCRVC^{+}08, MDET08, NYI^{+}05b, Rob08-50, oSCW07, TW07, yYMM05, vKST^{+}07]. translocational [HK08]. translocator [KRG^{+}08, THY^{+}06]. translocators [THS^{+}09]. translocon [ASB^{+}05b, JCMG08, Joh09, LeB07a, LeB08-32, Ska07]. Transmembrane [RCFB08, MNH05, OTH^{+}07, SS08a, XMA^{+}09]. transmigration [DDLC^{+}06]. transmission [AEJ^{+}06]. transmitted [NSM05]. transmitter [BSF^{+}08, CFH^{+}08, Rob08o]. transmitter-induced [BSF^{+}08]. Transport [HHFC06, PGB^{+}06, ARKS^{+}08, AGL^{+}08, ALB^{+}09, AM05b, BM07, BBPM07, CGS05, CKKG05, Den05, DSH^{+}07, ELM09, FHS^{+}07b, GS07, GEP08, GMSS06, GMS^{+}08, HFM^{+}05, HR05, HQF^{+}07, IEGV06, KHB^{+}05, KLR^{+}07, KGH^{+}05, LLH^{+}09, MSR09, MS06, MM05, NRW^{+}08, POCS^{+}06, PGL^{+}09, SRX^{+}07, Sch08, Sho09-43, SSZR07, SBJ^{+}07, SHH^{+}08, TKKT07, TMY^{+}09, UHM^{+}08, VVMS05, WHOS06, Wel06p, Wil06p, YM06, ZKN^{+}06, dPMF^{+}09]. transportation [Les07-36]. transported [MF08, RW05, WGKA^{+}08]. transporter [GS07, JLC08, JGL^{+}06, WOP^{+}06]. transporters [FA08]. transports [HLM^{+}08]. trap [KRP^{+}09, WLS^{+}09]. trapping [HK07]. traps [EMPK^{+}07, FAG^{+}07, LeB07-66]. trash [Sho09-54]. TRC8 [STvdB^{+}09]. TrCP [SCD^{+}08]. Tre1 [KSR^{+}08, Rob08-52]. treat [Les07-49]. treatment [Wil06o]. TREM [NBB^{+}09]. TREM-2 [NBB^{+}09]. trespass [WC07a]. TRF1 [ZMH^{+}09]. TRF2 [DMBF^{+}07]. tricellular [IFF^{+}05]. Tricellulin [IFF^{+}05]. trigger [ALB^{+}09, Les08-32, WWK^{+}06]. triggered [HJ06, MSC^{+}07, SDS^{+}05, VP05, YCR06, YTB07]. triggering [NBB^{+}09]. triggers [DNG^{+}06, GGMBF08, KKW^{+}08, LMMHP^{+}05, MBJF^{+}08, NMV^{+}07, Poo09, TDG^{+}06, ZKN^{+}06]. trilaminar [LRJY06]. TRIM5 [CPAH08]. trimers [SWL^{+}05]. trimethylation [NLG^{+}09]. Trimming [Tum06-50]. Trip [Wel06-65, LeB07l, Les08-42, Les09j, Rob08-36, Xia06]. TRIP6 [TSN^{+}06]. TRIP6/ZRP [TSN^{+}06]. TRIP6/ZRP-1 [TSN^{+}06]. triphosphatase [TTO^{+}08]. triphosphate [LMC^{+}09]. Tripin [HFF^{+}07a]. Tripin/ [HFF^{+}07a]. trisphosphate [HPH^{+}05, MuMI^{+}06, dCD^{+}07]. TrkB [HLM^{+}08]. TrkC [EYB^{+}06]. tRNA [Rob07n, Wil06p, YIfHA09]. tRNAs [TP09, Tum05-57]. tropicalis [BBM^{+}07]. Tropomodulin1 [NFZ^{+}09]. tropomyosin [HLP06, NHK^{+}08]. Trotman [Sed08k]. Trouble [Rob08-31, Les07-33]. TRP [COTAE^{+}05]. TRPL [COTAE^{+}05]. TRPP2 [KBGG^{+}08, Rob08-53]. TRPV4 [AFV^{+}05, FLA^{+}08, KBGG^{+}08, Rob08-53]. trs120 [CZP^{+}05]. true [LC08a]. Truth [Ros07a]. try [Les07x]. Trypanosoma [LN07, dGHW08]. trypanosome [DBD^{+}09]. trypanosomes [LBVN09, SWM^{+}06]. trypsinogen [HOH^{+}08]. TSC [GPN^{+}06, HMS^{+}08]. TSC1 [GMW^{+}05]. TSC2 [CTS^{+}06a, GMW^{+}05]. Tsien [Wil07-55]. Tsukita [Tak06]. Tuba [OIAT06]. Tube [Les08-63, MAZ^{+}08, SMSPK08, SBF^{+}06, Wel06-49]. tubes [GFD^{+}05, Les05h, Wel05-65]. tubular [Nyb05m]. tubulation [TSS^{+}06]. tubule [Les06y, PVJJ07]. tubules [Nyb05l]. Tubulin [BVS^{+}09, PTF^{+}06, ABvBHH06, DMDO08, HRB^{+}06, JSPT05, PWL^{+}09, STG05, Tum06-27, VCD^{+}06, Wel05q, WEL^{+}07n, Wil07-59, ZCJ^{+}08]. tubulin-binding [BKLT^{+}08]. tubulins [SSE05]. tucked [Wil07-38]. tug [LeB08p, Tum06-52]. tug-of-war [LeB08p]. Tumor [Rob08-54, Sed07g, Wil07-66, BvOM^{+}07, BIH^{+}05, BIH^{+}07, Di 08, FWC^{+}07, HZL^{+}07, HF06, KWY^{+}06, LPG^{+}07, Les06d, Les07-49, Les07-67, LBS^{+}08, SGMPV07, SYRL^{+}08, SGB^{+}06, SYM^{+}08b, VCC^{+}08, Wel06m, ZHS06]. tumorigenesis [CSW08, JMB^{+}07, LFW^{+}09a, LFW^{+}09b, MM08, MDWH^{+}08, Wel06v]. Tumors [LeB05-102, DDO^{+}08, LJN^{+}05, Les09-32, Rob08-46, WMS^{+}06, WC09]. tunes [Rob08n]. Tuning [Les08-59, WSG06]. tunnel [Poo09]. Tunneling [Tum06-51]. TuRC [Sho09a, VCD^{+}06]. Turn [Les09-35, Pow05p, SLZ^{+}06]. Turning [LeB07-64, Les08-60, Nyb05n, Wel05-67, SWM^{+}07]. turnover [BDS^{+}08, CEY^{+}05, DvCvH^{+}08, DTW06, EYB09, EHWL06, GMS^{+}08, KAL^{+}06, MTH^{+}06, OPT^{+}07, SHC^{+}05, SCK^{+}07, WWvD07]. turnover-dependent [MTH^{+}06]. turns [LeB05-60, LeB07-63, LeB08z, Les08i]. TWEAK [VCC^{+}08]. TWEAK-FN14 [VCC^{+}08]. TWEAKing [Ash08]. twist [Les07k, Rob07e]. Twisting [Wel06-66, Tum05-70]. Two [BM07, BCJ^{+}07, FSS^{+}09, GTO^{+}07, Kat06j, Kat07e, Kat07f, KMS07, KIKI^{+}07, LeB05-103, LeB05-104, LeB06-82, RBC05, Rob08-55, TPNS07a, TPNS07b, Tum05-74, WPL^{+}06, AWR^{+}07, AZP^{+}05, ABSK06, BFL^{+}05, BZA^{+}06, BCC^{+}09, CCS^{+}06, GEP08, GFD^{+}05, HL06, Joh09, KKRP05, LCGRB09, Les07-75, LNL^{+}08, LGBD^{+}07, Mar06, PCN^{+}05, QHB^{+}07, RGCWG09, Rob08m, Sed09d, Tum05q, Tum06v, Tum06-34, Tum06-48, Yu06]. two-pore [BCC^{+}09]. Two-speed [Kat06j]. two-state [Yu06]. Two-step [Tum05-74, Tum05q]. Type [LPA^{+}07, LBC^{+}07, SLWA07, AWR^{+}07, AG09, ADC^{+}07, BLC05, CM07, CZD^{+}09, CHCB05, CC09, DSLY^{+}05, DFA05, HHT^{+}07, HLP06, HSK^{+}05, HOZ^{+}09, IBS^{+}05, JGAS08, JLB^{+}05, JTH09, KKRP05, KLV^{+}08, MFO09, OFY^{+}05, RRA^{+}09, SDN^{+}05b, SBJ^{+}07, SBC^{+}07, TDT^{+}08, VCSG07, WBG^{+}08, dCD^{+}07, HCRT08]. type-1 [SDN^{+}05b]. types [LDL^{+}09b, PR05]. Typhimurium [SHB^{+}07]. tyrosination [PTF^{+}06, PWL^{+}09]. tyrosine [CNM^{+}05, FBO^{+}07, GJL^{+}08, GQSHO05, LPN^{+}08, NYN^{+}06, SKE^{+}05, SRL^{+}06, SWM^{+}06, WWC^{+}07, dCD^{+}07]. tyrosine-phosphorylated [GJL^{+}08].

U [YOSM06]. Ubc13 [AZP^{+}05]. Ube2j2 [WHR^{+}09]. UBF [SPS^{+}08]. Ubiquilin [LDL^{+}09a]. ubiquitin [AZP^{+}05, CNW^{+}05, DGSN06, EYB09, GGH^{+}08, HDR09, KNY07, LJA09, LJY^{+}07, MOS^{+}09, MSECB05, MG08a, MKC^{+}05, MG09, PGB^{+}06, RRR^{+}09, SOW^{+}09, TTE^{+}07, WWO^{+}08, YGW06, Cho09a, LM05, LZM^{+}06, Tum06-52, Wil09r]. ubiquitin-binding [SOW^{+}09]. ubiquitin-conjugating [AZP^{+}05]. ubiquitin-dependent [MG08a, LZM^{+}06]. Ubiquitin/ [LM05]. ubiquitinates [STvdB^{+}09, WHR^{+}09]. Ubiquitination [PES^{+}07, WHC^{+}07, WMT07, ALvdG06, KBC^{+}07b, MBJM^{+}09, PGB^{+}06, TdVM^{+}07]. ubiquitylation [CBG^{+}09, FD09, PGW^{+}08]. Ubp6 [Wil06q]. Uev1A [AZP^{+}05]. Ule [Wil08n]. ULK [HTK^{+}08]. ULK-interacting [HTK^{+}08]. Ultrastructural [KFC^{+}05]. Ultraviolet [NMV^{+}07, KDD08]. Unassisted [BYMB06]. unattached [Pow06l]. UNC [BMSJ07, JMG^{+}06, LLH^{+}07, MQL^{+}06, SHL^{+}09]. UNC-108 [SHL^{+}09]. UNC-108/ [SHL^{+}09]. UNC-45 [LLH^{+}07]. UNC-98 [MQL^{+}06]. Unc45b [ERS08]. Unc5B [HKIY09]. UNC93B [BSH^{+}07]. uncapping [Wel06h]. uncoated [KFC^{+}05]. uncoating [SSM^{+}08]. uncoiled [Pow06h]. unconventional [SWT^{+}05, TdVM^{+}07]. Uncoupling [WSG06]. underappreciated [MFC06]. undercover [Wel06-33]. undergo [BEH^{+}06]. undergoes [ABvBHH06, FFS06, PVJJ07, SWT08]. underlies [DWMY09, PRB^{+}09, SSD06, STBL09, dGSWW^{+}06]. underlying [ASB^{+}05b]. understanding [lWHMH06b]. unequal [Les07-61]. uneven [BLMS06]. unexpected [LF07]. unfold [Rob08-49]. unfolded [BBFN07, KIKI^{+}07, OOY^{+}06, SPT^{+}09, SPT^{+}21, YOSM06]. Unfolding [Les09-36, Sav09a, SWS^{+}08]. Ungermann [Sho09o]. unicellular [IH05]. unidirectional [SS09]. unifying [CBIO05]. Unique [Les07-73, LKL^{+}05, MYZ^{+}09, RWB^{+}05, WSS^{+}06]. uniquely [DSH^{+}07]. unit [LWBL06]. unite [CZK06]. Uniting [Les08-61, Les08-62]. units [FML^{+}06]. universal [MBJM^{+}09, WWK^{+}06]. unlikely [Wil07c]. unnerving [Sed09f]. unperturbed [TST08]. Unpleasant [Wel06-67]. unreplicated [OLH^{+}08]. unscheduled [WCR^{+}08]. Unscrambling [Tum06-53]. Unspliced [Tum06-54]. unstuck [Les09o]. unsynapsed [MBdR^{+}08]. Untangling [Les06-29]. untransformed [ULN^{+}07]. Untying [Les08-63]. unused [Tum06-55]. unveiled [Wel05b]. Unzipped [Ski05]. up-regulated [NJS^{+}06, PSL^{+}06]. up-regulates [GDF^{+}05]. up-regulation [CYY^{+}08, SFC^{+}08, TRC^{+}08]. uPA [SVS^{+}07a, SVS^{+}07b]. uPAR [CZM^{+}07, Les07-75, LJM^{+}07, MFA^{+}07, SMM08b]. uPAR-induced [MFA^{+}07]. uPARAP [CEY^{+}05]. uPARAP/Endo180 [CEY^{+}05]. updated [TRB09]. upon [Nyb05e, TJC06]. UPR [LeB07-65]. Ups1p [SDI^{+}06, TEIS09]. Ups2p [TEIS09]. upstream [DBDD06, LR05, LGG^{+}05, VKR^{+}09]. Uptake [BSF^{+}08, CTS^{+}06b, FHS^{+}07a, KKM^{+}08, SHjC^{+}07]. urchin [WNF^{+}05b]. urea [LeB06-88]. urokinase [WCR^{+}05]. uroplakin [MWSK06]. urothelial [LBH^{+}05]. used [KP08, WGL^{+}06]. uses [CGJ09, KDV^{+}09, LeB09a]. Usher [DTY^{+}05b]. using [LJS^{+}09b, TMBL^{+}08]. USP7 [FBJB^{+}09]. UTF1 [vdBKB^{+}07]. Utp7 [JhKC08]. Utrn [RGA^{+}06]. utrophin [PMC08]. UV [CBML06, KNKH06]. UV-induced [KNKH06, CBML06]. UXT [STL^{+}07].

V [AXM^{+}05, XMA^{+}09, CSL^{+}05, FSP^{+}07, WCN07, ZKR^{+}07, AFN^{+}08, CMF^{+}09, JTH09, LSR07, PCN^{+}05, VW08]. v-SNARE [FSP^{+}07]. V0 [LBG^{+}06]. V0-ATPase [LBG^{+}06]. Va [WTH05]. Vac14p [DTW06]. Vac7p [DTW06]. Vac8p [SSSL05]. VacA [GMG^{+}07]. vacuolar [BPL^{+}08, LU05, LeB05-70, MSECB05, SSSL05]. vacuole [FCDM^{+}07, JTH09, PLLH05, RM05a, Sho09o]. vacuoles [MES^{+}08, Miz05]. Valerie [Sho09-53]. Valosin [JFM^{+}09]. Valosin-containing [JFM^{+}09]. valves [Wel06w]. VAP [RKvdK^{+}09]. variant [HGJ06, LPWP^{+}07]. variants [AZP^{+}05, GER^{+}07]. variation [YCM^{+}08]. varicella [Wel05-58]. varicosities [DNG^{+}06]. Variety [LeB08-36, LeB06p]. various [LDL^{+}09b]. Vascular [BSK07, LOG^{+}06, Wil07-67, BHK^{+}06, HLA^{+}06, JKCW07, KWS^{+}08, LJN^{+}05, LR05, OSS^{+}07, SBF^{+}06, SYA^{+}08, SPA^{+}05]. VASP [BBM^{+}08, FSK^{+}07, HJK^{+}06, LLHH07, TSP08]. VCP [JFM^{+}09, ASK^{+}09, JFM^{+}09, Sho09-54, LDL^{+}09a]. VE [LR05, WKC^{+}09]. VE-cadherin [LR05]. VE-PTP [WKC^{+}09]. Vectorial [HSTM06]. VEGF [KZC^{+}08, LJN^{+}05, Les07-37, Les07-77, MCD^{+}07, ZXM^{+}06]. VEGF-induced [ZXM^{+}06]. VEGF111 [MCD^{+}07]. VEGFR [LOG^{+}06]. VEGFR-2 [LOG^{+}06]. VEGFR2 [KWS^{+}08]. VEGFs [LeB05-103]. vehicles [KFC^{+}05]. vein [Wel06n]. velocity [MS06]. ventral [STK^{+}05]. ventrally [WMSB^{+}06]. ventricular [PRB^{+}09]. verprolin [SBMP05]. versatile [RSL06]. versus [HBPL09, LeB05-63, MDWH^{+}08, PEER^{+}05, SSHW09, Wil09p, XLH^{+}07, Yu06]. vertebrate [DPP^{+}07, HDS^{+}07, YCX^{+}08, YWM^{+}08]. Very [HFF07b]. Very-KIND [HFF07b]. Vesicle [SAM^{+}07, BMK^{+}07, BHG^{+}05, DXC^{+}09, ECR^{+}09, FBO^{+}07, GSW^{+}08, JWG^{+}07, KKW^{+}07, LeB05-104, LOWTL^{+}08, LYH^{+}05, LSNY08, LWL08a, LWL08b, MSLS07, MHS^{+}09, SHH^{+}08, SHL^{+}09, Wel06-60, YSP^{+}06, ZL05]. vesicle-generating [MHS^{+}09]. vesicle-mediated [ZL05]. Vesicles [ACsHS08, LeB07-66, RW08d, ARKS^{+}08, BRKR06, BHH^{+}06, CFL^{+}05, EMPK^{+}07, HWSJ08, HXZ^{+}07, KES^{+}09, LeB06d, LMZ^{+}05, OVL^{+}07, RBC05, SSM^{+}08, Sho09-39, SORM05, Tum06-40, Wel05p, YOS^{+}05]. vesicular [HFP^{+}05, HR05, KBR^{+}08]. vessel [KZC^{+}08, KSE^{+}07, WKC^{+}09]. vessels [JKCW07, LeB05-103, NIC^{+}09b, Wel05-30]. VI [API^{+}07, MCSY07, OWM05, SRA^{+}05, SS09]. via [ACsHS08, BWL^{+}08, BKE^{+}07, CCS^{+}06, CTS^{+}05, DVF^{+}06, FFW^{+}07, FMD^{+}09, GG06, GFH^{+}06, GRC^{+}05, GGMBF08, GDR^{+}05, GFD^{+}05, HFP^{+}05, HDS^{+}07, KRZ^{+}08, KNKH06, KFR^{+}08, KBC^{+}08, LeB05y, LLHH07, LBC^{+}07, LNPC06, LFL^{+}05, MLV09, MMS07, MOM^{+}06, NMV^{+}07, NBM^{+}08, NNS^{+}08, OLC^{+}07, OFY^{+}05, dJPAW^{+}09, Pow05a, PKJS^{+}08, SBM^{+}07, SRL^{+}06, SSD^{+}09, SLZ^{+}06, SWI06, SWH^{+}06, TKB^{+}06, TMZ^{+}09, ToMY^{+}08, VSR^{+}08, WWX^{+}06, YKH^{+}08, ZKN^{+}06, ZCJ^{+}08]. viability [TMAL^{+}06]. view [LeB05-105, Ped08, Sho09-46]. viewing [Hil08]. VII [CSB^{+}06]. VII-dependent [CSB^{+}06]. vimentin [YGY^{+}05]. Vinculin [HWS^{+}07, CCC^{+}05, INB06, MCSY07]. Viral [RHMS05, KDD^{+}09, MSECB05, WHC^{+}07]. Virtual [Wel06-68]. virulence [CBG^{+}06, PUG^{+}06]. Virus [LeB06-83, Pow06m, CTS^{+}06b, DNG^{+}06, DHML05, FFS06, LK05]. Viruses [IH05, LeB07-67, Pow05w, Tum05-75, LSM^{+}05, Wil08d]. visceral [YBT^{+}09]. viscosity [AFV^{+}05]. Visualization [CPAH08, RSSC06, CM07, NTHN^{+}08]. visualized [SD07]. Visualizing [HTYNK09]. vitamin [MRD^{+}08, OMLP^{+}08, WJM07]. vitro [CGK07, KMK^{+}05b, Les06n, MK07, ONW^{+}07, SSE05, Wel05-40]. vitronectin [MFA^{+}07]. Vivek [Sav09b]. vivo [AVHH09a, AVHH09b, BJZ^{+}09, BLA^{+}09, CM07, CCSK08, DPL^{+}08, DR05, DOS^{+}07, EFG^{+}05, FSK^{+}07, GSC^{+}07, HOH^{+}07, KTN07, LZL^{+}05, MMU^{+}09, NLS^{+}05, ONW^{+}07, PVMH05, RMP06, SRBL06, SKR^{+}07, SWG^{+}05, TTNA^{+}06, VBSM06, YTH^{+}09, YMD^{+}07, YNL^{+}08]. vMIA [PPS^{+}06]. Voeltz [Sed08h]. Voltage [LPB^{+}08, BH08a, LYV^{+}08, Nyb05j]. Voltage-gated [LPB^{+}08, BH08a, LYV^{+}08]. volume [HG07]. Vps34 [JHY^{+}08, MES^{+}08]. Vps4 [ADD^{+}06, NWO06b]. Vps4-mediated [NWO06b]. Vps45p [CCS^{+}06]. Vps74 [WSB^{+}09]. Vps74/GOLPH3 [WSB^{+}09]. VSG [LN07, LBVN09]. VSL [ADD^{+}06]. Vta1 [ADD^{+}06].

VWF [Wel06-69].

W [Wil07-68]. Wac [MFC^{+}09]. wages [YD08]. Wait [Pow06l, Wel05-53]. waiting [Rob08-30]. Walhout [Sho09-35]. Walk [Wel05-68, Rob07e]. Walking [Pow09d]. wall [CJW07, Rob08-42, STK^{+}05]. Wallerian [ASK^{+}09]. walls [Wel05a]. wander [Wel06-53]. wandered [Wel06-67]. war [BEH^{+}06, LeB08p, Tum06-52]. warfare [LeB05-109]. Washington [LW08]. WASP [SWD^{+}08, CHCB05, KCOS06, OYM^{+}09, SBMP05]. WASp- [KCOS06]. WASp-binding [CHCB05]. waste [Wel06-34]. wave [Les09k, Tum05-75]. WAVE2 [NNS^{+}08, SKO^{+}06]. waves [Kat06h]. way [CF07, LeB05-36, LeB08q, Les05e, Les07r, Les07-28, Les07-34, Les07-49, Rob08g, Rob08-36, Rob08-55, Tum05-39, Tum05-43, Wel05-63, Wel05-68]. ways [HdM09, Kat07e, Kat07f, LeB06-82, Sho09-29]. wayward [Les09-35]. WD40 [GGC05]. WDR12 [HRS^{+}05]. weak [KP08, RW08e]. weaken [LeB06k]. Weaker [Nyb05o]. weakest [MZ08]. Weakness [LeB05-106]. web [EMD^{+}08]. webby [Wel07g]. webs [Wel07g]. Weibel [LRCH^{+}05]. Weidinger [Wil07w]. weight [Wel06u]. Weissman [LeB09b]. welcomes [Les07-57]. well [VWK^{+}06]. went [Wel05-47]. Were [Pow05x, Wil06b]. Werner [MGW^{+}08, PPBF08, YMTC05]. Werner/Bloom [MGW^{+}08]. Westermann [Sed07e]. wheel [PWW^{+}09]. where [LeB07-62, Rob08-52]. whereas [FFS06]. which [Les07-34, Wel05-63]. Who [LeB06-84]. whose [DCA^{+}06]. wide [BFWH07, EMB^{+}08, FMA^{+}09, SKM^{+}08, YSB^{+}07]. widespread [GFA^{+}09]. will [Rob08r]. winch [Tum06f]. wing [LeB05-29]. Wingless [SB06]. WIP [KCOS06]. WIP- [KCOS06]. Wiping [Wel05-69]. wisdom [Pow07e]. wise [HG07]. within [DBW^{+}06, DR05, GZBM^{+}06, GMZ^{+}08, HOH^{+}08, Les07-29, NS07, Rob07b, RCFB08, SRBL06, SSJ^{+}07]. without [FSN^{+}06, GAK^{+}05, LeB05-36, LeB06y, LeB07z, LeB07-34, LCvD^{+}08, Les07p, Les07y, MD09, Tum06o, VBSM06, Wel05-36, Wil07r]. Wld [ASK^{+}09, CWM^{+}09, Wil09k]. WNK1 [ZPGB^{+}07]. Wnt [LeB05-108, LeB05-107, LeB06-85, Les07-79, AACS^{+}06, DXC^{+}08, Hor09, KLW^{+}05, LCB^{+}08, MAH^{+}05, MG08b, OMLP^{+}08, dJPAW^{+}09, Pol08, SMH07, TG05]. Wnt- [dJPAW^{+}09]. Wnt/ [LCB^{+}08]. Wnt/Notch [MG08b]. Wnt11 [KHH08, WZCB^{+}06]. Wnt11-mediated [KHH08]. Wnt5a [NYN^{+}06]. Wnt5a-induced [NYN^{+}06]. Wnts [LVG^{+}09]. words [Les06a]. work [Les08-48, Pow05l]. Working [Tum05-76, LeB09b]. workout [Les08g]. world [CZK06, Mis09a]. Woronin [LNL^{+}08]. worth [Wil08j]. Worthylake [LeB07-53]. wound [BFG^{+}06b, CBM^{+}07, CTP^{+}09, LeB05-85, LeB06-64, SWG^{+}05, TPNS07a, TPNS07b, THN^{+}06, WHM^{+}07, WFJ06]. wounded [KSE^{+}07]. wounds [BB05, LeB07k, OTM^{+}07, Sho09-52]. wrestles [Sed09b]. WRN [FPP^{+}08]. WRN-deficient [FPP^{+}08]. wrong [Les07-67, Les08-31]. wrote [HR08]. Wunen [SRL05].

X [GLH^{+}05, KPM05, LW07b, PRK^{+}07, PRK^{+}09, TMI07, Wil07-51, BbCQ^{+}09, PAM^{+}07]. XBP1 [YOSM06]. Xenopus [BBM^{+}07, HH06a, KHH08, LM05, MFH05, OW06, UHM^{+}08, YMTC05]. XIAP [PVM^{+}05]. XIST [HBPL09, Sho09e]. XIST-ential [Sho09e]. Xmus101 [YLM06a]. Xnf7 [CNW^{+}05]. Xorbit [HH06a]. Xorbit/CLASP [HH06a]. Xorbiting [LeB06-86]. Xpo1p [SCVW09]. XRCC3 [OSI^{+}07]. XVI [Wel06o]. XX [LeB06-87].

Y14 [SRBL06]. yang [Kat07a]. YAP [HlQY^{+}06]. Yasushi [Wil08x]. Yb [SRQL09]. Yck3 [LU05]. years [Mel05]. Yeast [LeB05-109, MH07, Pow05z, SHDT05, ARKS^{+}08, ABA^{+}06, BJKF05, BBF^{+}07, BMSJ07, BMBD^{+}09, BEH^{+}06, BLvH^{+}05, CRLP07, CN06, CGGT09, CGJ09, DJV^{+}05, DSK^{+}06, FSG^{+}08, FN06, FDV^{+}06, GRGP07, GAG^{+}06, GMS^{+}08, HGJ06, HSY^{+}06, HEW07, HYB08, HLP06, JMG08, JGY09, JWSS05, KOM07, KKD^{+}07, LSDFS08, Les08o, LSC^{+}06b, LYTL07, LGBD^{+}07, LFS^{+}05, MTRJA05, MPL^{+}07a, MSW^{+}07, MSW^{+}17, MHFH06, MLP06, NNK^{+}08, NN07b, NdCF^{+}05, OSM^{+}09, PYR08, PKM^{+}05, PKB^{+}06, RHMS05, RWM^{+}05, SON06, SHC07, SCVW09, Sho09l, SBMP05, TJC06, Tum05k, Tum05-53, TS06, VP05, VDA^{+}08, WHOS06, WSK^{+}06, ZKN^{+}06, ZSMW09, LeB06-88]. YGR198w [FCDM^{+}07]. yields [KDD^{+}09]. yin [Kat07a]. Yme1L [SCF^{+}07]. yolk [Pow05f]. young [Pow09a, Les08-66]. yourself [LeB06s]. youth [LeB05p, Wel06z]. youthful [Wil07v]. YPP1 [FCDM^{+}07, BSA^{+}08]. Yps1 [VDA^{+}08]. Yuan [LeB07z]. Yvh1 [KOVP09, LLW^{+}09].

Z [ERS08, KFR^{+}08, LeB08s]. Z-disks [KFR^{+}08]. Zacheus [Wil07-68]. Zallen [Wil08m]. ZAP [IKZ^{+}07]. ZAP-70 [IKZ^{+}07]. Zasp [JS07]. ZBP1 [SLR^{+}06]. Zds1 [QU08]. Zds2 [QU08]. zebrafish [CULS^{+}07, FTB^{+}09, LSC^{+}05, PVMH05, YKP^{+}08]. zero [GHX^{+}07]. zero/P0 [GHX^{+}07]. Zhong [Sed08m]. zigzags [Tum06h]. Zinc [Pow07k, YSSH^{+}07, HOZ^{+}09]. zipped [SM09]. zipper [NPB^{+}05]. Zn [MSH^{+}05]. ZO [IUT^{+}07]. ZO-1 [IUT^{+}07]. zone [BBvD05, FvD08, FOW^{+}09, JAV^{+}07, PFD06, SCS^{+}05, TGH^{+}07]. zones [BB05, YN07]. ZRP-1 [TSN^{+}06]. ZW10 [KKW^{+}05, VDSV06]. ZYG [ZRH^{+}09]. ZYG-12 [ZRH^{+}09]. zygote [DPD^{+}09, GCS^{+}07, TPZG06]. zyxin [HJK^{+}06, YHJ^{+}05].

References

Ahmad-Annuar:2006:SAS

[AACS^{+}06] Azlina Ahmad-Annuar, Lorenza Ciani, Iordanis Simeonidis, Judit Herreros, Naila Ben Fredj, Silvana B. Rosso, Anita Hall, Stephen Brickley, and Patricia C. Salinas. Signaling across the synapse: a role for Wnt and Dishevelled in presynaptic assembly and neurotransmitter release. Journal of Cell Biology, 174 (1):127–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/1/127.

Anathy:2009:RAA

[AAG^{+}09] Vikas Anathy, Scott W. Aesif, Amy S. Guala, Marije Havermans, Niki L. Reynaert, Ye-Shih Ho, Ralph C. Budd, and Yvonne M. W. Janssen-Heininger. Redox amplification of apoptosis by caspase-dependent cleavage of glutaredoxin 1 and Sglutathionylation of Fas. Journal of Cell Biology, 184(2):241–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
184/2/241.

Allen:2006:IQN

[ABA^{+}06] Chris Allen, Sabrina Büttner, Anthony D. Aragon, Jason A. Thomas, Osorio Meirelles, Jason E. Jaetao, Don Benn, Stephanie W. Ruby, Marten Veenhuis, Frank Madeo, and Margaret Werner-Washburne. Isolation of quiescent and nonquiescent cells from yeast stationary-phase cultures. Journal of Cell Biology, 174(1):89–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/1/89.

Angelini:2006:MBB

[ABSK06] Sandra Angelini, Diana Boy, Emile Schiltz, and Hans-Georg Koch. Membrane binding of the bacterial signal recognition particle receptor involves two distinct binding sites. Journal of Cell Biology, 174(5):715–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/174/5/715.

Al-Bassam:2006:SBT

[ABvBHH06] Jawdat Al-Bassam, Mark van Breugel, Stephen C. Harrison, and Anthony Hyman. Stu2p binds tubulin and undergoes an open-to-closed conformational change. Journal of Cell Biology, 172(7):1009–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/7/1009.

Aucott:2008:HRD

[ABY^{+}08] Rebecca Aucott, Jörn Bullwinkel, Yang Yu, Wei Shi, Mustafa Billur, Jeremy P. Brown, Ursula Menzel, Dimitris Kioussis, Guozheng Wang, Ingrid Reisert, Jörg Weimer, Raj K. Pandita, Girdhar G. Sharma, Tej K. Pandita, Reinald Fundele, and Prim B. Singh. HP1-β is required for development of the cerebral neocortex and neuromuscular junctions. Journal of Cell Biology, 183(4):597–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/4/597.

Albertson:2008:VAT

[ACsHS08] Roger Albertson, Jian Cao, Tao shih Hsieh, and William Sullivan. Vesicles and actin are targeted to the cleavage furrow via furrow microtubules and the central spindle. Journal of Cell Biology, 181(5):777–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/5/777.

Atfi:2007:DMA

[ADC^{+}07] Azeddine Atfi, Emmanuelle Dumont, Frédéric Colland, Dominique Bonnier, Annie L’Helgoualc’h, Céline Prunier, Nathalie Ferrand, Bruno Clément, Ulla M. Wewer, and Nathalie Théret. The disintegrin and metalloproteinase ADAM12 contributes to TGF-β signaling through interaction with the type II receptor. Journal of Cell Biology, 178(2):201–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/178/2/201.

Azmi:2006:REA

[ADD^{+}06] Ishara Azmi, Brian Davies, Christian Dimaano, Johanna Payne, Debra Eckert, Markus Babst, and David J. Katzmann. Recycling of ESCRTs by the AAA–ATPase Vps4 is regulated by a conserved VSL region in Vta1. Journal of Cell Biology, 172(5): 705–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/5/705.

Audhya:2007:RRSa

[ADO07a] Anjon Audhya, Arshad Desai, and Karen Oegema. A role for Rab5 in structuring the endoplasmic reticulum. Journal of Cell Biology, 178(1):43–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/1/43.

Audhya:2007:RRSb

[ADO07b] Anjon Audhya, Arshad Desai, and Karen Oegema. A role for Rab5 in structuring the endoplasmic reticulum. Journal of Cell Biology, 178(7):1309–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/178/7/1309.

Acharya:2006:DMS

[AEJ^{+}06] Usha Acharya, Michael Beth Edwards, Ramon A. Jorquera, Hugo Silva, Kunio Nagashima, Pedro Labarca, and Jairaj K. Acharya. Drosophila melanogaster Scramblases modulate synaptic transmission. Journal of Cell Biology, 173(1):69–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
173/1/69.

Alon:2005:DAS

[AFM^{+}05] Ronen Alon, Sara W. Feigelson, Eugenia Manevich, David M. Rose, Julia Schmitz, Darryl R. Overby, Eitan Winter, Valentin Grabovsky, Vera Shinder, Benjamin D. Matthews, Maya Sokolovsky-Eisenberg, Donald E. Ingber, Martin Benoit, and Mark H. Ginsberg. α_{4}β_{1}-dependent adhesion strengthening under mechanical strain is regulated by paxillin association with the α_{4}-cytoplasmic domain. Journal of Cell Biology, 171(6): 1073–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/6/1073.

Altmann:2008:CVM

[AFN^{+}08] Katrin Altmann, Martina Frank, Daniel Neumann, Stefan Jakobs, and Benedikt Westermann. The class V myosin motor protein, Myo2, plays a major role in mitochondrial motility in Saccharomyces cerevisiae. Journal of Cell Biology, 181 (1):119–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/1/119.

Andrade:2005:TCI

[AFV^{+}05] Yaniré N. Andrade, Jacqueline Fernandes, Esther Vázquez, José M. Fernández-Fernández, Maite Arniges, Trinidad M. Sánchez, Manuel Villalón, and Miguel A. Valverde. TRPV4 channel is involved in the coupling of fluid viscosity changes to epithelial ciliary activity. Journal of Cell Biology, 168(6): 869–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/6/869.

Andres:2009:RTL

[AG09] Vicente Andrés and José M. González. Role of A-type lamins in signaling, transcription, and chromatin organization. Journal of Cell Biology, 187(7):945–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/187/7/945.

Ahmed:2008:OAA

[AGL^{+}08] Noveera T. Ahmed, Chunlei Gao, Ben F. Lucker, Douglas G. Cole, and David R. Mitchell. ODA16 aids axonemal outer row dynein assembly through an interaction with the intraflagellar transport machinery. Journal of Cell Biology, 183(2):313–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
183/2/313.

Alexandrow:2005:CDP

[AH05] Mark G. Alexandrow and Joyce L. Hamlin. Chromatin decondensation in S-phase involves recruitment of Cdk2 by Cdc45 and histone H1 phosphorylation. Journal of Cell Biology, 168 (6):875–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/6/875.

Anderson:2008:RER

[AH08] Daniel J. Anderson and Martin W. Hetzer. Reshaping of the endoplasmic reticulum limits the rate for nuclear envelope formation. Journal of Cell Biology, 182(5):911–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/182/5/911.

Azimzadeh:2009:HCB

[AHD^{+}09] Juliette Azimzadeh, Polla Hergert, Annie Delouvée, Ursula Euteneuer, Etienne Formstecher, Alexey Khodjakov, and Michel Bornens. hPOC5 is a centrin-binding protein required for assembly of full-length centrioles. Journal of Cell Biology, 185 (1):101–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/1/101.

Audhya:2005:CCS

[AHM^{+}05] Anjon Audhya, Francie Hyndman, Ian X. McLeod, Amy S. Maddox, John R. Yates, Arshad Desai, and Karen Oegema. A complex containing the Sm protein CAR-1 and the RNA helicase CGH-1 is required for embryonic cytokinesis in Caenorhabditis elegans. Journal of Cell Biology, 171(2):267–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/171/2/
267.

Adelfalk:2009:CSP

[AJR^{+}09] Caroline Adelfalk, Johannes Janschek, Ekaterina Revenkova, Cornelia Blei, Bodo Liebe, Eva Göb, Manfred Alsheimer, Ricardo Benavente, Esther de Boer, Ivana Novak, Christer Höög, Harry Scherthan, and Rolf Jessberger. Cohesin SMC1β protects telomeres in meiocytes. Journal of Cell Biology, 187(2):185–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
187/2/185.

Anderson:2006:RG

[AK06] Paul Anderson and Nancy Kedersha. RNA granules. Journal of Cell Biology, 172(6):803–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/172/6/803.

Almsherqi:2006:CML

[AKD06] Zakaria A. Almsherqi, Sepp D. Kohlwein, and Yuru Deng. Cubic membranes: a legend beyond the Flatland* of cell membrane organization. Journal of Cell Biology, 173(6):839–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
173/6/839.

Ally:2009:OPM

[ALB^{+}09] Shabeen Ally, Adam G. Larson, Kari Barlan, Sarah E. Rice, and Vladimir I. Gelfand. Opposite-polarity motors activate one another to trigger cargo transport in live cells. Journal of Cell Biology, 187(7):1071–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/187/7/1071.

Abrami:2006:RPU

[ALvdG06] Laurence Abrami, Stephen H. Leppla, and F. Gisou van der Goot. Receptor palmitoylation and ubiquitination regulate anthrax toxin endocytosis. Journal of Cell Biology, 172(2):309–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
172/2/309.

Anel:2005:C

[AM05a] Alberto Marcelo Diaz Añel and Vivek Malhotra. Correction. Journal of Cell Biology, 169(3):539–??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/169/3/539.

Anel:2005:PRP

[AM05b] Alberto Marcelo Díaz Añel and Vivek Malhotra. PKCη is required for β1γ2/β3γ2- and PKD-mediated transport to the cell surface and the organization of the Golgi apparatus. Journal of Cell Biology, 169(1):83–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/169/1/83.

Alexander:2007:DBH

[AMERT^{+}07] Mary Kate Alexander, Susanna Mlynarczyk-Evans, Morgan Royce-Tolland, Alex Plocik, Sundeep Kalantry, Terry Magnuson, and Barbara Panning. Differences between homologous alleles of olfactory receptor genes require the Polycomb Group protein Eed. Journal of Cell Biology, 179(2):269–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/2/
269.

Anderson:2005:TNE

[AMS^{+}05] Eric Anderson, Sandra Maday, Jeff Sfakianos, Michael Hull, Bettina Winckler, David Sheff, Heike Fölsch, and Ira Mellman. Transcytosis of NgCAM in epithelial cells reflects differential signal recognition on the endocytic and secretory pathways. Journal of Cell Biology, 170(4):595–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/170/4/595.

Aoki:2007:ERS

[ANI^{+}07] Kazuhiro Aoki, Takeshi Nakamura, Takanari Inoue, Tobias Meyer, and Michiyuki Matsuda. An essential role for the SHIP2-dependent negative feedback loop in neuritogenesis of nerve growth factor–stimulated PC12 cells. Journal of Cell Biology, 177(5):817–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/5/817.

Alonso:2005:SLG

[AOP^{+}05] Laura Alonso, Hitoshi Okada, Hilda Amalia Pasolli, Andrew Wakeham, Annick Itie You-Ten, Tak W. Mak, and Elaine Fuchs. Sgk3 links growth factor signaling to maintenance of progenitor cells in the hair follicle. Journal of Cell Biology, 170 (4):559–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/4/559.

Ahras:2006:SIN

[AOT06] Malika Ahras, Grant P. Otto, and Sharon A. Tooze. Synaptotagmin IV is necessary for the maturation of secretory granules in PC12 cells. Journal of Cell Biology, 173(2):241–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/173/2/
241.

Au:2007:MVR

[API^{+}07] Josephine Sui-Yan Au, Claudia Puri, Gudrun Ihrke, John Kendrick-Jones, and Folma Buss. Myosin VI is required for sorting of AP-1B-dependent cargo to the basolateral domain in polarized MDCK cells. Journal of Cell Biology, 177(1):103–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
177/1/103.

Aguilera-Romero:2008:YPC

[ARKS^{+}08] Auxiliadora Aguilera-Romero, Joanna Kaminska, Anne Spang, Howard Riezman, and Manuel Muñiz. The yeast p24 complex is required for the formation of COPI retrograde transport vesicles from the Golgi apparatus. Journal of Cell Biology, 180(4): 713–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/4/713.

Akhtar:2006:RLI

[AS06] Nasreen Akhtar and Charles H. Streuli. Rac1 links integrinmediated adhesion to the control of lactational differentiation in mammary epithelia. Journal of Cell Biology, 173(5):781–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
173/5/781.

Ausoni:2009:FAM

[AS09] Simonetta Ausoni and Saverio Sartore. From fish to amphibians to mammals: in search of novel strategies to optimize cardiac regeneration. Journal of Cell Biology, 184(3):357–??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/184/3/
357.

Alder:2005:C

[ASB^{+}05a] Nathan N. Alder, Ying Shen, Jeffrey L. Brodsky, Linda M. Hendershot, and Arthur E. Johnson. Correction. Journal of Cell Biology, 169(1):203–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/1/203.

Alder:2005:MMU

[ASB^{+}05b] Nathan N. Alder, Ying Shen, Jeffrey L. Brodsky, Linda M. Hendershot, and Arthur E. Johnson. The molecular mechanisms underlying BiP-mediated gating of the Sec61 translocon of the endoplasmic reticulum. Journal of Cell Biology, 168(3):389–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
168/3/389.

Amerik:2006:CLE

[ASH06] Alexander Amerik, Nazia Sindhi, and Mark Hochstrasser. A conserved late endosome–targeting signal required for Doa4 deubiquitylating enzyme function. Journal of Cell Biology, 175 (5):825–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/5/825.

Ashwell:2008:TD

[Ash08] Jonathan D. Ashwell. TWEAKing death. Journal of Cell Biology, 182(1):15–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/1/15.

Amano:2009:CCE

[ASH^{+}09] Miho Amano, Aussie Suzuki, Tetsuya Hori, Chelsea Backer, Katsuya Okawa, Iain M. Cheeseman, and Tatsuo Fukagawa. The CENP–S complex is essential for the stable assembly of outer kinetochore structure. Journal of Cell Biology, 186 (2):173–??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/186/2/173.

Arias-Salgado:2005:PEP

[ASHD^{+}05] Elena Garcia Arias-Salgado, Fawaz Haj, Christophe Dubois, Barry Moran, Ana Kasirer-Friede, Barbara C. Furie, Bruce Furie, Benjamin G. Neel, and Sanford J. Shattil. PTP-1B is an essential positive regulator of platelet integrin signaling. Journal of Cell Biology, 170(5):837–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/170/5/837.

Avery:2009:WRN

[ASK^{+}09] Michelle A. Avery, Amy E. Sheehan, Kimberly S. Kerr, Jing Wang, and Marc R. Freeman. Wld^{S} requires Nmnat1 enzymatic activity and N16–VCP interactions to suppress Wallerian degeneration. Journal of Cell Biology, 184(4):501–??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/184/4/
501.

Aivazian:2006:TKE

[ASP06] Dikran Aivazian, Ramon L. Serrano, and Suzanne Pfeffer. TIP47 is a key effector for Rab9 localization. Journal of Cell Biology, 173(6):917–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/6/917.

Artal-Sanz:2006:LBF

[ASSST06] Marta Artal-Sanz, Chrysanthi Samara, Popi Syntichaki, and Nektarios Tavernarakis. Lysosomal biogenesis and function is critical for necrotic cell death in Caenorhabditis elegans. Journal of Cell Biology, 173(2):231–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/173/2/231.

Anderson:2009:RFDa

[AVHH09a] Daniel J. Anderson, Jesse D. Vargas, Joshua P. Hsiao, and Martin W. Hetzer. Recruitment of functionally distinct membrane proteins to chromatin mediates nuclear envelope formation in vivo. Journal of Cell Biology, 186(2):183–??, July 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/186/2/183.

Anderson:2009:RFDb

[AVHH09b] Daniel J. Anderson, Jesse D. Vargas, Joshua P. Hsiao, and Martin W. Hetzer. Recruitment of functionally distinct membrane proteins to chromatin mediates nuclear envelope formation in vivo. Journal of Cell Biology, 186(6):929–??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/186/6/929.

Albert:2005:TRM

[AWE^{+}05] Y’vonne Albert, Jennifer Whitehead, Laurie Eldredge, John Carter, Xiaoguang Gao, and Warren G. Tourtellotte. Transcriptional regulation of myotube fate specification and intrafusal muscle fiber morphogenesis. Journal of Cell Biology, 169 (2):257–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/2/257.

Axe:2008:AFM

[AWM^{+}08] Elizabeth L. Axe, Simon A. Walker, Maria Manifava, Priya Chandra, H. Llewelyn Roderick, Anja Habermann, Gareth Griffiths, and Nicholas T. Ktistakis. Autophagosome formation from membrane compartments enriched in phosphatidylinositol 3-phosphate and dynamically connected to the endoplasmic reticulum. Journal of Cell Biology, 182(4):685–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/182/4/685.

Alto:2007:TIE

[AWR^{+}07] Neal M. Alto, Andrew W. Weflen, Matthew J. Rardin, Defne Yarar, Cheri S. Lazar, Raffi Tonikian, Antonius Koller, Susan S. Taylor, Charles Boone, Sachdev S. Sidhu, Sandra L. Schmid, Gail A. Hecht, and Jack E. Dixon. The type III effector EspF coordinates membrane trafficking by the spatiotemporal activation of two eukaryotic signaling pathways. Journal of Cell Biology, 178(7):1265–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/178/7/1265.

Adair:2005:TDE

[AXM^{+}05] Brian D. Adair, Jian-Ping Xiong, Catherine Maddock, Simon L. Goodman, M. Amin Arnaout, and Mark Yeager. Threedimensional EM structure of the ectodomain of integrin αVβ3 in a complex with fibronectin. Journal of Cell Biology, 168 (7):1109–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/7/1109.

Andersen:2005:DRU

[AZP^{+}05] Parker L. Andersen, Honglin Zhou, Landon Pastushok, Trevor Moraes, Sean McKenna, Barry Ziola, Michael J. Ellison, Vishva M. Dixit, and Wei Xiao. Distinct regulation of Ubc13 functions by the two ubiquitin-conjugating enzyme variants Mms2 and Uev1A. Journal of Cell Biology, 170(5):745–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
170/5/745.

Beraldo:2005:CAC

[BAdSG05] Flávio H. Beraldo, Fabiana M. Almeida, Aline M. da Silva, and Célia R. S. Garcia. Cyclic AMP and calcium interplay as second messengers in melatonin-dependent regulation of Plasmodium falciparum cell cycle. Journal of Cell Biology, 170(4):551–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
170/4/551.

Bankaitis:2009:CDS

[Ban09] Vytas A. Bankaitis. The Cirque du Soleil of Golgi membrane dynamics. Journal of Cell Biology, 186(2):169–??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/186/2/169.

Barnich:2005:MRN

[BAR^{+}05] Nicolas Barnich, Jose E. Aguirre, Hans-Christian Reinecker, Ramnik Xavier, and Daniel K. Podolsky. Membrane recruitment of NOD2 in intestinal epithelial cells is essential for nuclear factor–κB activation in muramyl dipeptide recognition. Journal of Cell Biology, 170(1):21–??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/170/1/21.

Boag:2008:PSM

[BAR^{+}08] Peter R. Boag, Arzu Atalay, Stacey Robida, Valerie Reinke, and T. Keith Blackwell. Protection of specific maternal messenger RNAs by the P body protein CGH-1 (Dhh1/RCK) during Caenorhabditis elegans oogenesis. Journal of Cell Biology, 182 (3):543–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/3/543.

Bashyam:2007:DCM

[Bas07a] Hema Bashyam. Dying cells make a joint. Journal of Cell Biology, 176(1):4–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/1/4.

Bashyam:2007:SEE

[Bas07b] Hema Bashyam. Self-eating ERs. Journal of Cell Biology, 176(1):??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/1/5b.

Bashyam:2007:SS

[Bas07c] Hema Bashyam. Shrinking spines. Journal of Cell Biology, 176 (1):5–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/1/5.

Bashyam:2007:SBI

[Bas07d] Hema Bashyam. Skin-based immunity. Journal of Cell Biology, 176(1):??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/1/5a.

Bashyam:2007:SFS

[Bas07e] Hema Bashyam. Stretchable force sensors. Journal of Cell Biology, 176(1):??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/1/4a.

Benink:2005:CZA

[BB05] Hélène A. Benink and William M. Bement. Concentric zones of active RhoA and Cdc42 around single cell wounds. Journal of Cell Biology, 168(3):429–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/168/3/429.

Berman:2009:BXI

[BbCQ^{+}09] Sarah B. Berman, Ying bei Chen, Bing Qi, J. Michael McCaffery, Edmund B. Rucker, Sandra Goebbels, Klaus-Armin Nave, Beth A. Arnold, Elizabeth A. Jonas, Fernando J. Pineda, and J. Marie Hardwick. Bcl-x_{L} increases mitochondrial fission, fusion, and biomass in neurons. Journal of Cell Biology, 184 (5):707–??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/184/5/707.

Behnia:2007:YOG

[BBF^{+}07] Rudy Behnia, Francis A. Barr, John J. Flanagan, Charles Barlowe, and Sean Munro. The yeast orthologue of GRASP65 forms a complex with a coiled-coil protein that contributes to ER to Golgi traffic. Journal of Cell Biology, 176(3):255–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/176/3/
255.

Bicknell:2007:NRC

[BBFN07] Alicia A. Bicknell, Anna Babour, Christine M. Federovitch, and Maho Niwa. A novel role in cytokinesis reveals a housekeeping function for the unfolded protein response. Journal of Cell Biology, 177(6):1017–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/6/1017.

Busca:2005:HIF

[BBG^{+}05] Roser Buscà, Edurne Berra, Cédric Gaggioli, Mehdi Khaled, Karine Bille, Barbara Marchetti, Raphaël Thyss, Giorgos Fitsialos, Lionel Larribère, Corine Bertolotto, Thierry Virolle, Pascal Barbry, Jacques Pouysségur, Gilles Ponzio, and Robert Ballotti. Hypoxia-inducible factor 1α is a new target of microphthalmia-associated transcription factor (MITF) in melanoma cells. Journal of Cell Biology, 170(1):49–??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/170/1/
49.

Buttrick:2008:ARC

[BBL^{+}08] Graham J. Buttrick, Luke M. A. Beaumont, Jessica Leitch, Christopher Yau, Julian R. Hughes, and James G. Wakefield. Akt regulates centrosome migration and spindle orientation in the early Drosophila melanogaster embryo. Journal of Cell Biology, 180(3):537–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/3/537.

Brown:2007:XTE

[BBM^{+}07] Katherine S. Brown, Michael D. Blower, Thomas J. Maresca, Timothy C. Grammer, Richard M. Harland, and Rebecca Heald. Xenopus tropicalis egg extracts provide insight into scaling of the mitotic spindle. Journal of Cell Biology, 176 (6):765–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/6/765.

Benz:2008:CAE

[BBM^{+}08] Peter M. Benz, Constanze Blume, Jan Moebius, Chris Oschatz, Kai Schuh, Albert Sickmann, Ulrich Walter, Stephan M. Feller, and Thomas Renné. Cytoskeleton assembly at endothelial cell–cell contacts is regulated by αII–spectrin–VASP complexes. Journal of Cell Biology, 180(1):205–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/180/1/205.

Bossard:2007:DPR

[BBPM07] Carine Bossard, Damien Bresson, Roman S. Polishchuk, and Vivek Malhotra. Dimeric PKD regulates membrane fission to form transport carriers at the TGN. Journal of Cell Biology, 179(6):1123–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/6/1123.

Bement:2005:MDZ

[BBvD05] William M. Bement, Hélène A. Benink, and George von Dassow. A microtubule-dependent zone of active RhoA during cleavage plane specification. Journal of Cell Biology, 170(1):91–??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/170/1/
91.

Bompard:2005:C

[BC05] Guillaume Bompard and Emmanuelle Caron. Correction. Journal of Cell Biology, 169(6):987–??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/169/6/987.

Brailoiu:2009:ERT

[BCC^{+}09] Eugen Brailoiu, Dev Churamani, Xinjiang Cai, Michael G. Schrlau, G. Cristina Brailoiu, Xin Gao, Robert Hooper, Michael J. Boulware, Nae J. Dun, Jonathan S. Marchant, and Sandip Patel. Essential requirement for two-pore channel 1 in NAADP-mediated calcium signaling. Journal of Cell Biology, 186(2):201–??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/186/2/201.

Blasius:2007:TBP

[BCJ^{+}07] T. Lynne Blasius, Dawen Cai, Gloria T. Jih, Christopher P. Toret, and Kristen J. Verhey. Two binding partners cooperate to activate the molecular motor Kinesin-1. Journal of Cell Biology, 176(1):11–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/1/11.

Blagoveshchenskaya:2008:IGT

[BCR^{+}08] Anastasia Blagoveshchenskaya, Fei Ying Cheong, Holger M. Rohde, Greta Glover, Andreas Knödler, Teresa Nicolson, Guido Boehmelt, and Peter Mayinger. Integration of Golgi trafficking and growth factor signaling by the lipid phosphatase SAC1. Journal of Cell Biology, 180(4):803–??, February 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/180/4/803.

Bertotti:2006:IAS

[BCT06] Andrea Bertotti, Paolo M. Comoglio, and Livio Trusolino. β4 integrin activates a Shp2–Src signaling pathway that sustains HGF-induced anchorage-independent growth. Journal of Cell Biology, 175(6):993–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/175/6/993.

Bastide:2007:SRC

[BDP^{+}07] Pauline Bastide, Charbel Darido, Julie Pannequin, Ralf Kist, Sylvie Robine, Christiane Marty-Double, Frédéric Bibeau, Gerd Scherer, Dominique Joubert, Frédéric Hollande, Philippe Blache, and Philippe Jay. Sox9 regulates cell proliferation and is required for Paneth cell differentiation in the intestinal epithelium. Journal of Cell Biology, 178(4):635–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/178/4/635.

Boyer:2006:ITM

[BDR^{+}06] Laurent Boyer, Anne Doye, Monica Rolando, Gilles Flatau, Patrick Munro, Pierre Gounon, René Clément, Céline Pulcini, Michel R. Popoff, Amel Mettouchi, Luce Landraud, Olivier Dussurget, and Emmanuel Lemichez. Induction of transient macroapertures in endothelial cells through RhoA inhibition by Staphylococcus aureus factors. Journal of Cell Biology, 173 (5):809–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/5/809.

Bonetti:2008:NAA

[BDS^{+}08] Paola Bonetti, Teresa Davoli, Cristina Sironi, Bruno Amati, Pier Giuseppe Pelicci, and Emanuela Colombo. Nucleophosmin and its AML-associated mutant regulate c–Myc turnover through Fbw7γ. Journal of Cell Biology, 182(1):19–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/182/1/
19.

Bulankina:2009:TFB

[BDW^{+}09] Anna V. Bulankina, Anke Deggerich, Dirk Wenzel, Kudzai Mutenda, Julia G. Wittmann, Markus G. Rudolph, Koert N. J. Burger, and Stefan Höning. TIP47 functions in the biogenesis of lipid droplets. Journal of Cell Biology, 185(4):641–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/185/4/
641.

Barr:2005:GP

[BE05] Francis A. Barr and Johannes Egerer. Golgi positioning. Journal of Cell Biology, 168(7):993–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/168/7/993.

Beaudoin:2006:CNA

[Bea06] Gerard M. J. Beaudoin. Con-nectin axons and dendrites. Journal of Cell Biology, 174(1):7–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/174/1/7.

Buttner:2006:WYC

[BEH^{+}06] Sabrina Büttner, Tobias Eisenberg, Eva Herker, Didac Carmona-Gutierrez, Guido Kroemer, and Frank Madeo. Why yeast cells can undergo apoptosis: death in times of peace, love, and war. Journal of Cell Biology, 175(4):521–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/175/4/
521.

Brero:2005:MCB

[BEN^{+}05] Alessandro Brero, Hariharan P. Easwaran, Danny Nowak, Ingrid Grunewald, Thomas Cremer, Heinrich Leonhardt, and M. Cristina Cardoso. Methyl CpG–binding proteins induce large-scale chromatin reorganization during terminal differentiation. Journal of Cell Biology, 169(5):733–??, June 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/169/5/733.

Bond:2009:PNB

[BF09] Charles S. Bond and Archa H. Fox. Paraspeckles: nuclear bodies built on long noncoding RNA. Journal of Cell Biology, 186(5):637–??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/186/5/637.

Brechet:2008:PKC

[BFB^{+}08] Aline Bréchet, Marie-Pierre Fache, Anna Brachet, Géraldine Ferracci, Agnés Baude, Marie Irondelle, Sandrine Pereira, Christophe Leterrier, and Bénédicte Dargent. Protein kinase CK2 contributes to the organization of sodium channels in axonal membranes by regulating their interactions with ankyrin G. Journal of Cell Biology, 183(6):1101–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/183/6/
1101.

Bolcun-Filas:2007:SRS

[BFCS^{+}07] Ewelina Bolcun-Filas, Yael Costa, Robert Speed, Mary Taggart, Ricardo Benavente, Dirk G. De Rooij, and Howard J. Cooke. SYCE2 is required for synaptonemal complex assembly, double strand break repair, and homologous recombination. Journal of Cell Biology, 176(6):741–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/176/6/741.

Booth:2006:EHG

[BFF^{+}06] Amy M. Booth, Yi Fang, Jonathan K. Fallon, Jr-Ming Yang, James E. K. Hildreth, and Stephen J. Gould. Exosomes and HIV Gag bud from endosome-like domains of the T cell plasma membrane. Journal of Cell Biology, 172(6):923–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/172/6/923.

Bandyopadhyay:2006:C

[BFG^{+}06a] Balaji Bandyopadhyay, Jianhua Fan, Shengxi Guan, Yong Li, Mei Chen, David T. Woodley, and Wei Li. Correction. Journal of Cell Biology, 173(2):311–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/173/2/311.

Bandyopadhyay:2006:TCR

[BFG^{+}06b] Balaji Bandyopadhyay, Jianhua Fan, Shengxi Guan, Yong Li, Mei Chen, David T. Woodley, and Wei Li. A “traffic control” role for TGFβ3: orchestrating dermal and epidermal cell motility during wound healing. Journal of Cell Biology, 172 (7):1093–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/7/1093.

Berliocchi:2005:BNC

[BFL^{+}05] Laura Berliocchi, Eugenio Fava, Marcel Leist, Volker Horvat, David Dinsdale, David Read, and Pierluigi Nicotera. Botulinum neurotoxin C initiates two different programs for neurite degeneration and neuronal apoptosis. Journal of Cell Biology, 168(4): 607–??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/4/607.

Bialucha:2007:PNM

[BFP^{+}07] Carl U. Bialucha, Emma C. Ferber, Franck Pichaud, Sew Y. Peak-Chew, and Yasuyuki Fujita. p32 is a novel mammalian Lgl binding protein that enhances the activity of protein kinase Cζ and regulates cell polarity. Journal of Cell Biology, 178 (4):575–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/4/575.

Blower:2007:GWA

[BFWH07] Michael D. Blower, Elma Feric, Karsten Weis, and Rebecca Heald. Genome-wide analysis demonstrates conserved localization of messenger RNAs to mitotic microtubules. Journal of Cell Biology, 179(7):1365–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/7/1365.

Brown:2008:ABA

[BGdN^{+}08] Jill M. Brown, Joanne Green, Ricardo Pires das Neves, Helen A. C. Wallace, Andrew J. H. Smith, Jim Hughes, Nicki Gray, Steve Taylor, William G. Wood, Douglas R. Higgs, Francisco J. Iborra, and Veronica J. Buckle. Association between active genes occurs at nuclear speckles and is modulated by chromatin environment. Journal of Cell Biology, 182(6):1083–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/182/6/
1083.

Benetti:2007:SDR

[BGJ^{+}07] Roberta Benetti, Susana Gonzalo, Isabel Jaco, Gunnar Schotta, Peter Klatt, Thomas Jenuwein, and María A. Blasco. Suv4-20h deficiency results in telomere elongation and derepression of telomere recombination. Journal of Cell Biology, 178(6):

925–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/6/925.

Burbank:2006:NMR

[BGP^{+}06] Kendra S. Burbank, Aaron C. Groen, Zachary E. Perlman, Daniel S. Fisher, and Timothy J. Mitchison. A new method reveals microtubule minus ends throughout the meiotic spindle. Journal of Cell Biology, 175(3):369–??, November 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/175/3/369.

Baumgartner:2007:PCS

[BGZ^{+}07] Petra Baumgärtner, Margarethe Geiger, Susanne Zieseniss, Julia Malleier, James A. Huntington, Karin Hochrainer, Edith Bielek, Mechthild Stoeckelhuber, Kirsten Lauber, Dag Scherfeld, Petra Schwille, Katja Wäldele, Klaus Beyer, and Bernd Engelmann. Phosphatidylethanolamine critically supports internalization of cell-penetrating protein C inhibitor. Journal of Cell Biology, 179(4):793–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/4/793.

Bennett:2008:BTC

[BH08a] Vann Bennett and Jane Healy. Being there: cellular targeting of voltage-gated sodium channels in the heart. Journal of Cell Biology, 180(1):13–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/1/13.

Bird:2008:BSC

[BH08b] Alexander W. Bird and Anthony A. Hyman. Building a spindle of the correct length in human cells requires the interaction between TPX2 and Aurora A. Journal of Cell Biology, 182(2):289–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/2/289.

Bielli:2005:RSN

[BHG^{+}05] Anna Bielli, Charles J. Haney, Gavin Gabreski, Simon C. Watkins, Sergei I. Bannykh, and Meir Aridor. Regulation of Sar1 NH_{2} terminus by GTP binding and hydrolysis promotes membrane deformation to control COPII vesicle fission. Journal of Cell Biology, 171(6):919–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/171/6/919.

Borner:2006:CPC

[BHH^{+}06] Georg H. H. Borner, Michael Harbour, Svenja Hester, Kathryn S. Lilley, and Margaret S. Robinson. Comparative proteomics of clathrin-coated vesicles. Journal of Cell Biology, 175(4):571–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/4/571.

Benlimame:2005:FSC

[BHJ^{+}05] Naciba Benlimame, Qiang He, Su Jie, Dingzhang Xiao, Ying Jie Xu, Martin Loignon, David D. Schlaepfer, and Moulay A. Alaoui-Jamali. FAK signaling is critical for ErbB-2/ErbB-3 receptor cooperation for oncogenic transformation and invasion. Journal of Cell Biology, 171(3):505–??, November 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/171/3/505.

Braren:2006:EFE

[BHK^{+}06] Rickmer Braren, Huiqing Hu, Yung Hae Kim, Hilary E. Beggs, Louis F. Reichardt, and Rong Wang. Endothelial FAK is essential for vascular network stability, cell survival, and lamellipodial formation. Journal of Cell Biology, 172(1):151–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/172/1/
151.

Baker:2008:OSS

[BHY^{+}08] Sheila A. Baker, Mohammad Haeri, Peter Yoo, Sidney M. Gospe, Nikolai P. Skiba, Barry E. Knox, and Vadim Y. Arshavsky. The outer segment serves as a default destination for the trafficking of membrane proteins in photoreceptors. Journal of Cell Biology, 183(3):485–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/183/3/485.

Bois:2005:FAS

[BIH^{+}05] Philippe R. J. Bois, Kamel Izeradjene, Peter J. Houghton, John L. Cleveland, Janet A. Houghton, and Gerard C.

Grosveld. FOXO1a acts as a selective tumor suppressor in alveolar rhabdomyosarcoma. Journal of Cell Biology, 170(6): 903–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/6/903.

Bois:2007:FAS

[BIH^{+}07] Philippe R. J. Bois, Kamel Izeradjene, Peter J. Houghton, John L. Cleveland, Janet A. Houghton, and Gerard C. Grosveld. FOXO1a acts as a selective tumor suppressor in alveolar rhabdomyosarcoma. Journal of Cell Biology, 177(3):563– ??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
177/3/563.

Binns:2006:ICB

[BJC^{+}06] Derk Binns, Tom Januszewski, Yue Chen, Justin Hill, Vladislav S. Markin, Yingming Zhao, Christopher Gilpin, Kent D. Chapman, Richard G. W. Anderson, and Joel M. Goodman. An intimate collaboration between peroxisomes and lipid bodies. Journal of Cell Biology, 173(5):719–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/173/5/719.

Bachant:2005:YPC

[BJKF05] Jeff Bachant, Shannon R. Jessen, Sarah E. Kavanaugh, and Candida S. Fielding. The yeast S phase checkpoint enables replicating chromosomes to bi-orient and restrain spindle extension during S phase distress. Journal of Cell Biology, 168 (7):999–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/7/999.

Bekker-Jensen:2006:SOM

[BJLK^{+}06] Simon Bekker-Jensen, Claudia Lukas, Risa Kitagawa, Fredrik Melander, Michael B. Kastan, Jiri Bartek, and Jiri Lukas. Spatial organization of the mammalian genome surveillance machinery in response to DNA strand breaks. Journal of Cell Biology, 173(2):195–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/2/195.

Bekker-Jensen:2005:DAS

[BJLM^{+}05] Simon Bekker-Jensen, Claudia Lukas, Fredrik Melander, Jiri Bartek, and Jiri Lukas. Dynamic assembly and sustained retention of 53BP1 at the sites of DNA damage are controlled by Mdc1/NFBD1. Journal of Cell Biology, 170(2):201–??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/170/2/
201.

Baker:2006:EAA

[BJM^{+}06] Darren J. Baker, Karthik B. Jeganathan, Liviu Malureanu, Carmen Perez-Terzic, Andre Terzic, and Jan M. A. van Deursen. Early aging–associated phenotypes in Bub3/Rae1 haploinsufficient mice. Journal of Cell Biology, 172(4):529–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/172/4/
529.

Barber:2009:PRS

[BJML09] Cynthia F. Barber, Ramon A. Jorquera, Jan E. Melom, and J. Troy Littleton. Postsynaptic regulation of synaptic plasticity by synaptotagmin 4 requires both C2 domains. Journal of Cell Biology, 187(2):295–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/2/295.

Bjork:2009:SCS

[BJZ^{+}09] Petra Björk, ShaoBo Jin, Jian Zhao, Om Prakash Singh, Jan-Olov Persson, Ulf Hellman, and Lars Wieslander. Specific combinations of SR proteins associate with single pre-messenger RNAs in vivo and contribute different functions. Journal of Cell Biology, 184(4):555–??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/184/4/555.

Brieher:2006:RAM

[BKBM06] William M. Brieher, Hao Yuan Kueh, Bryan A. Ballif, and Timothy J. Mitchison. Rapid actin monomer–insensitive depolymerization of Listeria actin comet tails by cofilin, coronin, and Aip1. Journal of Cell Biology, 175(2):315–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/175/2/315.

Bell:2007:QSM

[BKE^{+}07] Eric L. Bell, Tatyana A. Klimova, James Eisenbart, Carlos T. Moraes, Michael P. Murphy, G. R. Scott Budinger, and Navdeep S. Chandel. The Q_{o} site of the mitochondrial complex III is required for the transduction of hypoxic signaling via reactive oxygen species production. Journal of Cell Biology, 177(6):1029–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/6/1029.

Barros:2005:AAD

[BKHR05] Teresa P. Barros, Kazuhisa Kinoshita, Anthony A. Hyman, and Jordan W. Raff. Aurora A activates D–TACC–Msps complexes exclusively at centrosomes to stabilize centrosomal microtubules. Journal of Cell Biology, 170(7):1039–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/170/7/
1039.

Bieling:2008:CTG

[BKLT^{+}08] Peter Bieling, Stefanie Kandels-Lewis, Ivo A. Telley, Juliette van Dijk, Carsten Janke, and Thomas Surrey. CLIP-170 tracks growing microtubule ends by dynamically recognizing composite EB1/tubulin-binding sites. Journal of Cell Biology, 183(7): 1223–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/7/1223.

Brangwynne:2008:CDM

[BKMW08] Clifford P. Brangwynne, Gijsje H. Koenderink, Frederick C. MacKintosh, and David A. Weitz. Cytoplasmic diffusion: molecular motors mix it up. Journal of Cell Biology, 183(4): 583–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/4/583.

Burgess:2009:LRP

[BLA^{+}09] Rebecca C. Burgess, Michael Lisby, Veronika Altmannova, Lumir Krejci, Patrick Sung, and Rodney Rothstein. Localization of recombination proteins and Srs2 reveals anti-recombinase function in vivo. Journal of Cell Biology, 185(6):969–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/185/6/
969.

Bjorkoy:2005:PSF

[BLB^{+}05] Geir Bjørkøy, Trond Lamark, Andreas Brech, Heidi Outzen, Maria Perander, Aud Øvervatn, Harald Stenmark, and Terje Johansen. p62/SQSTM1 forms protein aggregates degraded by autophagy and has a protective effect on huntingtin-induced cell death. Journal of Cell Biology, 171(4):603–??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/171/4/
603.

Bernot:2005:SSH

[BLC05] Kelsie M. Bernot, Chang-Hun Lee, and Pierre A. Coulombe. A small surface hydrophobic stripe in the coiled-coil domain of type I keratins mediates tetramer stability. Journal of Cell Biology, 168(6):965–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/6/965.

Bhagatji:2009:SSS

[BLC^{+}09] Pinkesh Bhagatji, Rania Leventis, Jonathan Comeau, Mohammad Refaei, and John R. Silvius. Steric and not structure-specific factors dictate the endocytic mechanism of glycosylphosphatidylinositol-anchored proteins. Journal of Cell Biology, 186(4):615–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/186/4/615.

Bastiani:2009:MCC

[BLH^{+}09] Michele Bastiani, Libin Liu, Michelle M. Hill, Mark P. Jedrychowski, Susan J. Nixon, Harriet P. Lo, Daniel Abankwa, Robert Luetterforst, Manuel Fernandez-Rojo, Michael R. Breen, Steven P. Gygi, Jorgen Vinten, Piers J. Walser, Kathryn N. North, John F. Hancock, Paul F. Pilch, and Robert G. Parton. MURC/ cavin-4 and cavin family members form tissue-specific caveolar complexes. Journal of Cell Biology, 185(7):1259–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/7/1259.

Bang:2006:NDM

[BLL^{+}06] Marie-Louise Bang, Xiaodong Li, Ryan Littlefield, Shannon Bremner, Andrea Thor, Kirk U. Knowlton, Richard L. Lieber, and Ju Chen. Nebulin-deficient mice exhibit shorter thin filament lengths and reduced contractile function in skeletal muscle. Journal of Cell Biology, 173(6):905–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/173/6/905.

Blank:2006:CRU

[BLMS06] Michael Blank, Yaniv Lerenthal, Leonid Mittelman, and Yosef Shiloh. Condensin I recruitment and uneven chromatin condensation precede mitotic cell death in response to DNA damage. Journal of Cell Biology, 174(2):195–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/174/2/195.

Beronja:2005:EFD

[BLP^{+}05] Slobodan Beronja, Patrick Laprise, Ophelia Papoulas, Milena Pellikka, John Sisson, and Ulrich Tepass. Essential function of Drosophila Sec6 in apical exocytosis of epithelial photoreceptor cells. Journal of Cell Biology, 169(4):635–??, May 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/169/4/635.

Brown:2006:CHG

[BLR^{+}06] Jill M. Brown, Joanne Leach, Joyce E. Reittie, Ann Atzberger, Jane Lee-Prudhoe, William G. Wood, Douglas R. Higgs, Francisco J. Iborra, and Veronica J. Buckle. Coregulated human globin genes are frequently in spatial proximity when active. Journal of Cell Biology, 172(2):177–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/172/2/177.

Bodrikov:2005:REN

[BLS^{+}05] Vsevolod Bodrikov, Iryna Leshchyns’ka, Vladimir Sytnyk, John Overvoorde, Jeroen den Hertog, and Melitta Schachner. RPTPα is essential for NCAM-mediated p59^{fyn} activation and neurite elongation. Journal of Cell Biology, 168(1):127–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
168/1/127.

Badie:2009:RFC

[BLT^{+}09a] Sophie Badie, Chunyan Liao, Maria Thanasoula, Paul Barber, Mark A. Hill, and Madalena Tarsounas. RAD51C facilitates checkpoint signaling by promoting CHK2 phosphorylation. Journal of Cell Biology, 185(4):587–??, May 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/185/4/587.

Bellon:2009:ENR

[BLT09b] Ellen Bellon, Frank P. Luyten, and Przemko Tylzanowski. δ-EF1 is a negative regulator of Ihh in the developing growth plate. Journal of Cell Biology, 187(5):685–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/187/5/685.

Bystricky:2005:CLY

[BLvH^{+}05] Kerstin Bystricky, Thierry Laroche, Griet van Houwe, Marek Blaszczyk, and Susan M. Gasser. Chromosome looping in yeast. Journal of Cell Biology, 168(3):375–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/168/3/375.

Bageshwar:2007:TEP

[BM07] Umesh K. Bageshwar and Siegfried M. Musser. Two electrical potential-dependent steps are required for transport by the Escherichia coli Tat machinery. Journal of Cell Biology, 179 (1):87–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/1/87.

Boireau:2007:TCH

[BMB^{+}07] Stéphanie Boireau, Paolo Maiuri, Eugenia Basyuk, Manuel de la Mata, Anna Knezevich, Bérangère Pradet-Balade, Volker Bäcker, Alberto Kornblihtt, Alessandro Marcello, and Edouard Bertrand. The transcriptional cycle of HIV-1 in real-time and live cells. Journal of Cell Biology, 179(2):291–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/179/2/291.

Burston:2009:RYE

[BMBD^{+}09] Helen E. Burston, Lymarie Maldonado-Báez, Michael Davey, Benjamen Montpetit, Cayetana Schluter, Beverly Wendland, and Elizabeth Conibear. Regulators of yeast endocytosis identified by systematic quantitative analysis. Journal of Cell Biology, 185(6):1097–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/6/1097.

Brandt:2007:DII

[BMG^{+}07] Dominique T. Brandt, Sabrina Marion, Gareth Griffiths, Takashi Watanabe, Kozo Kaibuchi, and Robert Grosse. Dia1 and IQGAP1 interact in cell migration and phagocytic cup formation. Journal of Cell Biology, 178(2):193–??, July 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/178/2/193.

Boulon:2008:HCC

[BMGPB^{+}08] Séverine Boulon, Nathalie Marmier-Gourrier, Bérengère Pradet-Balade, Laurence Wurth, Céline Verheggen, Beáta E. Jády, Benjamin Rothé, Christina Pescia, Marie-Cécile Robert, Tamás Kiss, Barbara Bardoni, Alain Krol, Christiane Branlant, Christine Allmang, Edouard Bertrand, and Bruno Charpentier. The Hsp90 chaperone controls the biogenesis of L7Ae RNPs through conserved machinery. Journal of Cell Biology, 180(3): 579–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/3/579.

Brangwynne:2006:MCB

[BMK^{+}06] Clifford P. Brangwynne, Frederick C. MacKintosh, Sanjay Kumar, Nicholas A. Geisse, Jennifer Talbot, L. Mahadevan, Kevin K. Parker, Donald E. Ingber, and David A. Weitz. Microtubules can bear enhanced compressive loads in living cells because of lateral reinforcement. Journal of Cell Biology, 173 (5):733–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/5/733.

Besse:2007:ICA

[BMK^{+}07] Florence Besse, Sara Mertel, Robert J. Kittel, Carolin Wichmann, Tobias M. Rasse, Stephan J. Sigrist, and Anne Ephrussi. The Ig cell adhesion molecule Basigin controls compartmentalization and vesicle release at Drosophila melanogaster synapses. Journal of Cell Biology, 177(5):843–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/177/5/843.

Buchan:2008:PBP

[BMP08] J. Ross Buchan, Denise Muhlrad, and Roy Parker. P bodies promote stress granule assembly in Saccharomyces cerevisiae. Journal of Cell Biology, 183(3):441–??, November 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/183/3/441.

Bass:2008:PCP

[BMR^{+}08] Mark D. Bass, Mark R. Morgan, Kirsty A. Roach, Jeffrey Settleman, Andrew B. Goryachev, and Martin J. Humphries. p190RhoGAP is the convergence point of adhesion signals from α_{5}β_{1} integrin and syndecan-4. Journal of Cell Biology, 181 (6):1013–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/6/1013.

Bourgin:2007:ERR

[BMRP07] Caroline Bourgin, Keith K. Murai, Melanie Richter, and Elena B. Pasquale. The EphA4 receptor regulates dendritic spine remodeling by affecting β1-integrin signaling pathways. Journal of Cell Biology, 178(7):1295–??, September 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/178/7/1295.

Bartolini:2008:FMS

[BMS^{+}08] Francesca Bartolini, James B. Moseley, Jan Schmoranzer, Lynne Cassimeris, Bruce L. Goode, and Gregg G. Gundersen. The formin mDia2 stabilizes microtubules independently of its actin nucleation activity. Journal of Cell Biology, 181(3):523– ??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
181/3/523.

Bupp:2007:TAN

[BMSJ07] Jennifer M. Bupp, Adriana E. Martin, Elizabeth S. Stensrud, and Sue L. Jaspersen. Telomere anchoring at the nuclear periphery requires the budding yeast Sad1-UNC-84 domain protein Mps3. Journal of Cell Biology, 179(5):845–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/5/
845.

Bihlmaier:2007:DRS

[BMT^{+}07] Karl Bihlmaier, Nikola Mesecke, Nadia Terziyska, Melanie Bien, Kai Hell, and Johannes M. Herrmann. The disulfide relay system of mitochondria is connected to the respiratory chain. Journal of Cell Biology, 179(3):389–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/3/389.

Bonetta:2005:AL

[Bon05a] Laura Bonetta. Actin in locomotion. Journal of Cell Biology, 169(6):838–??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/6/838.

Bonetta:2005:ANC

[Bon05b] Laura Bonetta. Actin in nonmuscle cells. Journal of Cell Biology, 169(6):??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/6/839.1.

Bonetta:2005:ETJ

[Bon05c] Laura Bonetta. Endothelial tight junctions form the blood– brain barrier. Journal of Cell Biology, 169(3):378–??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/169/3/378.

Bonetta:2005:GCM

[Bon05d] Laura Bonetta. Growth cones make proteins, too. Journal of Cell Biology, 169(6):??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/6/839.2.

Bonetta:2005:HML

[Bon05e] Laura Bonetta. How to make a lysosome. Journal of Cell Biology, 169(4):??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/4/553.2.

Bonetta:2005:MSC

[Bon05f] Laura Bonetta. Microtubules shape the cell. Journal of Cell Biology, 169(4):??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/4/553.1.

Bonetta:2005:AMI

[Bon05g] Laura Bonetta. Not actin, not myosin, but intermediate. Journal of Cell Biology, 169(5):??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/169/5/704.2.

Bonetta:2005:SPa

[Bon05h] Laura Bonetta. Seeing peroxisomes. Journal of Cell Biology, 169(5):705–??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/5/705.

Bonetta:2005:SPb

[Bon05i] Laura Bonetta. Spectrin is peripheral. Journal of Cell Biology, 170(1):??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/1/12.1.

Bonetta:2005:TGC

[Bon05j] Laura Bonetta. Tension gets chromosomes oriented. Journal of Cell Biology, 169(5):??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/5/704.1.

Bauer:2005:CCA

[BP05] Sylvian Bauer and Paul H. Patterson. The cell cycle–apoptosis connection revisited in the adult brain. Journal of Cell Biology, 171(4):641–??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/4/641.

Bektas:2005:NAK

[BPL^{+}05] Meryem Bektas, Shawn G. Payne, Hong Liu, Sravan Goparaju, Sheldon Milstien, and Sarah Spiegel. A novel acylglycerol kinase that produces lysophosphatidic acid modulates cross talk with EGFR in prostate cancer cells. Journal of Cell Biology, 169(5):801–??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/5/801.

Brett:2008:ETV

[BPL^{+}08] Christopher L. Brett, Rachael L. Plemel, Braden T. Lobingier, Marissa Vignali, Stanley Fields, and Alexey J. Merz. Efficient termination of vacuolar Rab GTPase signaling requires coordinated action by a GAP and a protein kinase. Journal of Cell Biology, 182(6):1141–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/182/6/1141.

Barmchi:2005:DRA

[BRH05] Mojgan Padash Barmchi, Stephen Rogers, and Udo Häcker. DRhoGEF2 regulates actin organization and contractility in the Drosophila blastoderm embryo. Journal of Cell Biology, 168(4): 575–??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/4/575.

Bamji:2006:BMS

[BRKR06] Shernaz X. Bamji, Beatriz Rico, Nikole Kimes, and Louis F. Reichardt. BDNF mobilizes synaptic vesicles and enhances synapse formation by disrupting cadherin–β-catenin interactions. Journal of Cell Biology, 174(2):289–??, July 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/174/2/289.

Bass:2007:SDR

[BRM^{+}07] Mark D. Bass, Kirsty A. Roach, Mark R. Morgan, Zohreh Mostafavi-Pour, Tobias Schoen, Takashi Muramatsu, Ulrike Mayer, Christoph Ballestrem, Joachim P. Spatz, and Martin J. Humphries. Syndecan-4-dependent Rac1 regulation determines directional migration in response to the extracellular matrix. Journal of Cell Biology, 177(3):527–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/177/3/527.

Brunelle:2009:MDL

[BRY^{+}09] Joslyn K. Brunelle, Jeremy Ryan, Derek Yecies, Joseph T. Opferman, and Anthony Letai. MCL-1-dependent leukemia cells are more sensitive to chemotherapy than BCL-2-dependent counterparts. Journal of Cell Biology, 187(3):429–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/187/3/
429.

Baird:2008:APK

[BSA^{+}08] Dan Baird, Chris Stefan, Anjon Audhya, Sabine Weys, and Scott D. Emr. Assembly of the PtdIns 4-kinase Stt4 complex at the plasma membrane requires Ypp1 and Efr3. Journal of Cell Biology, 183(6):1061–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/183/6/1061.

Boyarchuk:2007:BEA

[BSDA07] Yekaterina Boyarchuk, Adrian Salic, Mary Dasso, and Alexei Arnaoutov. Bub1 is essential for assembly of the functional inner centromere. Journal of Cell Biology, 176(7):919–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/176/7/
919.

Bergami:2008:URP

[BSF^{+}08] Matteo Bergami, Spartaco Santi, Elena Formaggio, Cinzia Cagnoli, Claudia Verderio, Robert Blum, Benedikt Berninger, Michela Matteoli, and Marco Canossa. Uptake and recycling of pro–BDNF for transmitter-induced secretion by cortical astrocytes. Journal of Cell Biology, 183(2):213–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/183/2/213.

Brinkmann:2007:IBM

[BSH^{+}07] Melanie M. Brinkmann, Eric Spooner, Kasper Hoebe, Bruce Beutler, Hidde L. Ploegh, and You-Me Kim. The interaction between the ER membrane protein UNC93B and TLR3, 7, and 9 is crucial for TLR signaling. Journal of Cell Biology, 177 (2):265–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/2/265.

Ball:2007:VEG

[BSK07] Stephen G. Ball, C. Adrian Shuttleworth, and Cay M. Kielty. Vascular endothelial growth factor can signal through plateletderived growth factor receptors. Journal of Cell Biology, 177 (3):489–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/3/489.

Bodrikov:2008:NIC

[BSL^{+}08] Vsevolod Bodrikov, Vladimir Sytnyk, Iryna Leshchyns’ka, Jeroen den Hertog, and Melitta Schachner. NCAM induces CaMKIIα-mediated RPTPα phosphorylation to enhance its catalytic activity and neurite outgrowth. Journal of Cell Biology, 182(6):1185–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/6/1185.

Bui:2008:MAI

[BSM^{+}08] Khanh Huy Bui, Hitoshi Sakakibara, Tandis Movassagh, Kazuhiro Oiwa, and Takashi Ishikawa. Molecular architecture of inner dynein arms in situ in Chlamydomonas reinhardtii flagella. Journal of Cell Biology, 183(5):923–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/183/5/923.

Bui:2009:AID

[BSM^{+}09] Khanh Huy Bui, Hitoshi Sakakibara, Tandis Movassagh, Kazuhiro Oiwa, and Takashi Ishikawa. Asymmetry of inner dynein arms and inter-doublet links in Chlamydomonas flagella. Journal of Cell Biology, 186(3):437–??, August 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/186/3/437.

Baba:2005:PCP

[BSMT05] Takeshi Baba, Toshiaki Sakisaka, Sumiko Mochida, and Yoshimi Takai. PKA-catalyzed phosphorylation of tomosyn and its implication in Ca^{2+}-dependent exocytosis of neurotransmitter. Journal of Cell Biology, 170(7):1113–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/170/7/
1113.

Banerjee:2008:MPG

[BSO^{+}08] Soma Banerjee, Stephanie Smith, Ji-Hyun Oum, Hung-Jiun Liaw, Ji-Young Hwang, Nilabja Sikdar, Akira Motegi, Sang Eun Lee, and Kyungjae Myung. Mph1p promotes gross chromosomal rearrangement through partial inhibition of homologous recombination. Journal of Cell Biology, 181(7):1083–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/181/7/
1083.

Bahe:2005:RFC

[BSW^{+}05] Susanne Bahe, York-Dieter Stierhof, Christopher J. Wilkinson, Florian Leiss, and Erich A. Nigg. Rootletin forms centrioleassociated filaments and functions in centrosome cohesion. Journal of Cell Biology, 171(1):27–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/171/1/27.

Barbaro:2007:CER

[BTD^{+}07] Vanessa Barbaro, Anna Testa, Enzo Di Iorio, Fulvio Mavilio, Graziella Pellegrini, and Michele De Luca. C/ EBPδ regulates cell cycle and self-renewal of human limbal stem cells. Journal of Cell Biology, 177(6):1037–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/177/6/1037.

Benninger:2007:EDR

[BTP^{+}07] Yves Benninger, Tina Thurnherr, Jorge A. Pereira, Sven Krause, Xunwei Wu, Anna Chrostek-Grashoff, Dominik Herzog, Klaus-Armin Nave, Robin J. M. Franklin, Dies Meijer, Cord Brakebusch, Ueli Suter, and João B. Relvas. Essential and distinct roles for cdc42 and rac1 in the regulation of Schwann cell biology during peripheral nervous system development. Journal of Cell Biology, 177(6):1051–??, June 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/177/6/1051.

Bloom:2008:SPI

[BUJ^{+}08] Ona Bloom, Julia J. Unternaehrer, Aimin Jiang, Jeong-Sook Shin, Lélia Delamarre, Patrick Allen, and Ira Mellman. Spinophilin participates in information transfer at immunological synapses. Journal of Cell Biology, 181(2):203–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/181/2/
203.

Boukh-Viner:2005:DEC

[BVGA^{+}05] Tatiana Boukh-Viner, Tong Guo, Alex Alexandrian, André Cerracchio, Christopher Gregg, Sandra Haile, Robert Kyskan, Svetlana Milijevic, Daniel Oren, Jonathan Solomon, Vivianne Wong, Jean-Marc Nicaud, Richard A. Rachubinski, Ann M. English, and Vladimir I. Titorenko. Dynamic ergosterol- and ceramide-rich domains in the peroxisomal membrane serve as an organizing platform for peroxisome fusion. Journal of Cell Biology, 168(5):761–??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/5/761.

Blais:2007:RTS

[BvOM^{+}07] Alexandre Blais, Chris J. C. van Oevelen, Raphaël Margueron, Diego Acosta-Alvear, and Brian David Dynlacht. Retinoblastoma tumor suppressor protein-dependent methylation of histone H3 lysine 27 is associated with irreversible cell cycle exit. Journal of Cell Biology, 179(7):1399–??, December 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/179/7/1399.

Bennett:2005:IAM

[BvRJ^{+}05] Clare L. Bennett, Erwin van Rijn, Steffen Jung, Kayo Inaba, Ralph M. Steinman, Martien L. Kapsenberg, and Björn E. Clausen. Inducible ablation of mouse Langerhans cells diminishes but fails to abrogate contact hypersensitivity. Journal of Cell Biology, 169(4):569–??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/169/4/569.

Bouissou:2009:TRC

[BVS^{+}09] Anaïs Bouissou, Christel Vérollet, Aureliana Sousa, Paula Sampaio, Michel Wright, Claudio E. Sunkel, Andreas Merdes, and Brigitte Raynaud-Messina. γ-tubulin ring complexes regulate microtubule plus end dynamics. Journal of Cell Biology, 187 (3):327–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/187/3/327.

Boiko:2006:MUC

[BW06] Tatiana Boiko and Bettina Winckler. Myelin under construction — teamwork required. Journal of Cell Biology, 172(6):799–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
172/6/799.

Belting:2008:NEN

[BW08] Mattias Belting and Anders Wittrup. Nanotubes, exosomes, and nucleic acid–binding peptides provide novel mechanisms of intercellular communication in eukaryotic cells: implications in health and disease. Journal of Cell Biology, 183(7):1187–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
183/7/1187.

Buonomo:2009:MRC

[BWFdL09] Sara B. C. Buonomo, Yipin Wu, David Ferguson, and Titia de Lange. Mammalian Rif1 contributes to replication stress survival and homology-directed repair. Journal of Cell Biology, 187(3):385–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/3/385.

Besirli:2005:LRN

[BWJ05] Cagri G. Besirli, Erwin F. Wagner, and Eugene M. Johnson. The limited role of NH_{2}-terminal c–Jun phosphorylation in neuronal apoptosis. Journal of Cell Biology, 170(3):401–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/170/3/
401.

Bakkar:2008:INR

[BWL^{+}08] Nadine Bakkar, Jingxin Wang, Katherine J. Ladner, Huating Wang, Jason M. Dahlman, Micheal Carathers, Swarnali Acharyya, Michael A. Rudnicki, Andrew D. Hollenbach, and Denis C. Guttridge. IKK/ NF-κB regulates skeletal myogenesis via a signaling switch to inhibit differentiation and promote mitochondrial biogenesis. Journal of Cell Biology, 180(4): 787–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/4/787.

Bisel:2008:ERG

[BWW^{+}08] Blaine Bisel, Yanzhuang Wang, Jen-Hsuan Wei, Yi Xiang, Danming Tang, Miguel Miron-Mendoza, Shin ichiro Yoshimura, Nobuhiro Nakamura, and Joachim Seemann. ERK regulates Golgi and centrosome orientation towards the leading edge through GRASP65. Journal of Cell Biology, 182(5): 837–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/5/837.

Bogdanov:2008:FFL

[BXHD08] Mikhail Bogdanov, Jun Xie, Phil Heacock, and William Dowhan. To flip or not to flip: lipid–protein charge interactions are a determinant of final membrane protein topology. Journal of Cell Biology, 182(5):925–??, September 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/182/5/925.

Bridges:2006:AIS

[BXN^{+}06] James P. Bridges, Yan Xu, Cheng-Lun Na, Hector R. Wong, and Timothy E. Weaver. Adaptation and increased susceptibility to infection associated with constitutive expression of misfolded SP–C. Journal of Cell Biology, 172(3):395–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/172/3/
395.

Ball:2009:RRH

[BY09] Alexander R. Ball and Kyoko Yokomori. Revisiting the role of heterochromatin protein 1 in DNA repair. Journal of Cell Biology, 185(4):573–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/4/573.

Bachelder:2005:GSK

[BYF^{+}05] Robin E. Bachelder, Sang-Oh Yoon, Clara Franci, Antonio García de Herreros, and Arthur M. Mercurio. Glycogen synthase kinase-3 is an endogenous inhibitor of Snail transcription. Journal of Cell Biology, 168(1):29–??, January 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/168/1/29.

Ben-Yair:2008:NBM

[BYK08] Raz Ben-Yair and Chaya Kalcheim. Notch and bone morphogenetic protein differentially act on dermomyotome cells to generate endothelium, smooth, and striated muscle. Journal of Cell Biology, 180(3):607–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/3/607.

Brambillasca:2006:UTL

[BYMB06] Silvia Brambillasca, Monica Yabal, Marja Makarow, and Nica Borgese. Unassisted translocation of large polypeptide domains across phospholipid bilayers. Journal of Cell Biology, 175(5): 767–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/5/767.

Brito:2008:MDP

[BYR08] Daniela A. Brito, Zhenye Yang, and Conly L. Rieder. Microtubules do not promote mitotic slippage when the spindle assembly checkpoint cannot be satisfied. Journal of Cell Biology, 182(4):623–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/4/623.

Boban:2006:AIN

[BZA^{+}06] Mirta Boban, Arezou Zargari, Claes Andréasson, Stijn Heessen, Johan Thyberg, and Per O. Ljungdahl. Asi1 is an inner nuclear membrane protein that restricts promoter access of two latent transcription factors. Journal of Cell Biology, 173(5):695–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
173/5/695.

Bass-Zubek:2008:PCS

[BZHA^{+}08] Amanda E. Bass-Zubek, Ryan P. Hobbs, Evangeline V. Amargo, Nicholas J. Garcia, Sherry N. Hsieh, Xinyu Chen, James K. Wahl, Mitchell F. Denning, and Kathleen J. Green. Plakophilin 2: a critical scaffold for PKCα that regulates intercellular junction assembly. Journal of Cell Biology, 181(4):605– ??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
181/4/605.

Barreiro:2008:EAR

[BZYM^{+}08] Olga Barreiro, Moreno Zamai, María Yáñez-Mó, Emilio Tejera, Pedro López-Romero, Peter N. Monk, Enrico Gratton, Valeria R. Caiolfa, and Francisco Sánchez-Madrid. Endothelial adhesion receptors are recruited to adherent leukocytes by inclusion in preformed tetraspanin nanoplatforms. Journal of Cell Biology, 183(3):527–??, November 2008. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/183/3/527.

Cullen:2007:HMG

[CAL^{+}07] Sean P. Cullen, Colin Adrain, Alexander U. Lüthi, Patrick J. Duriez, and Seamus J. Martin. Human and murine granzyme B exhibit divergent substrate preferences. Journal of Cell Biology, 176(4):435–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/4/435.

Canman:2009:CA

[Can09] Julie C. Canman. Cytokinetic astralogy. Journal of Cell Biology, 187(6):757–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/6/757.

Cao:2008:NRE

[CAR^{+}08] Jian Cao, Roger Albertson, Blake Riggs, Christine M. Field, and William Sullivan. Nuf, a Rab11 effector, maintains cytokinetic furrow integrity by promoting local actin polymerization. Journal of Cell Biology, 182(2):301–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/182/2/301.

Cai:2008:PDI

[CB08] Liang Cai and James E. Bear. Peering deeply inside the branch. Journal of Cell Biology, 180(5):853–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/180/5/853.

Colakoglu:2009:IFE

[ÇB09] Gülsen Çolakoğlu and Anthony Brown. Intermediate filaments exchange subunits along their length and elongate by end-to-end annealing. Journal of Cell Biology, 185(5):769–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/185/5/
769.

Cooke:2006:MCA

[CBG^{+}06] Brian M. Cooke, Donna W. Buckingham, Fiona K. Glenister, Kate M. Fernandez, Lawrence H. Bannister, Matthias Marti, Narla Mohandas, and Ross L. Coppel. A Maurer’s cleft-associated protein is essential for expression of the major malaria virulence antigen on the surface of infected red blood cells. Journal of Cell Biology, 172(6):899–??, March 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/172/6/899.

Cohen:2009:DMA

[CBG^{+}09] Shenhav Cohen, Jeffrey J. Brault, Steven P. Gygi, David J. Glass, David M. Valenzuela, Carlos Gartner, Esther Latres, and Alfred L. Goldberg. During muscle atrophy, thick, but not thin, filament components are degraded by MuRF1-dependent ubiquitylation. Journal of Cell Biology, 185(6):1083–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/185/6/
1083.

Cullen:2005:CKN

[CBIO05] C. Fiona Cullen, Amy L. Brittle, Takashi Ito, and Hiroyuki Ohkura. The conserved kinase NHK-1 is essential for mitotic progression and unifying acentrosomal meiotic spindles in Drosophila melanogaster. Journal of Cell Biology, 171(4): 593–??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/4/593.

Camarata:2006:LRT

[CBK^{+}06] Troy Camarata, Benjamin Bimber, Andre Kulisz, Teng-Leong Chew, Jennifer Yeung, and Hans-Georg Simon. LMP4 regulates Tbx5 protein subcellular localization and activity. Journal of Cell Biology, 174(3):339–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/3/339.

Chmielowiec:2007:CME

[CBM^{+}07] Jolanta Chmielowiec, Malgorzata Borowiak, Markus Morkel, Theresia Stradal, Barbara Munz, Sabine Werner, Jürgen Wehland, Carmen Birchmeier, and Walter Birchmeier. c-met is essential for wound healing in the skin. Journal of Cell Biology, 177(1):151–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/1/151.

Cheerambathur:2008:DPM

[CBMCSS08] Dhanya K. Cheerambathur, Ingrid Brust-Mascher, Gul Civelekoglu-Scholey, and Jonathan M. Scholey. Dynamic partitioning of mitotic kinesin-5 cross-linkers between microtubule-bound and freely diffusing states. Journal of Cell Biology, 182(3):429–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
182/3/429.

Cioce:2006:UIF

[CBML06] Mario Cioce, Séverine Boulon, A. Gregory Matera, and Angus I. Lamond. UV-induced fragmentation of Cajal bodies. Journal of Cell Biology, 175(3):401–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/175/3/401.

Chiron:2008:CRM

[CBZY08] Stéphane Chiron, Alyona Bobkova, Haowen Zhou, and Michael P. Yaffe. CLASP regulates mitochondrial distribution in Schizosaccharomyces pombe. Journal of Cell Biology, 182(1):41–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/1/41.

Cianciola:2009:ARA

[CC09] Nicholas L. Cianciola and Cathleen R. Carlin. Adenovirus RIDα activates an autonomous cholesterol regulatory mechanism that rescues defects linked to Niemann–Pick disease type C. Journal of Cell Biology, 187(4):537–??, November 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/187/4/537.

Chen:2005:SDF

[CCC^{+}05] Hui Chen, Daniel M. Cohen, Dilshad M. Choudhury, Noriyuki Kioka, and Susan W. Craig. Spatial distribution and functional significance of activated vinculin in living cells. Journal of Cell Biology, 169(3):459–??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/3/459.

Chou:2006:STP

[CCC^{+}06] Ming-Lun Chou, Chiung-Chih Chu, Lih-Jen Chen, Mitsuru Akita, and Hsou min Li. Stimulation of transit-peptide release and ATP hydrolysis by a cochaperone during protein import into chloroplasts. Journal of Cell Biology, 175(6):893–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
175/6/893.

Chan:2009:FAI

[CCH09] Keefe T. Chan, Christa L. Cortesio, and Anna Huttenlocher. FAK alters invadopodia and focal adhesion composition and dynamics to regulate breast cancer invasion. Journal of Cell Biology, 185(2):357–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/2/357.

Croft:2005:AMB

[CCL^{+}05] Daniel R. Croft, Mathew L. Coleman, Shuixing Li, David Robertson, Teresa Sullivan, Colin L. Stewart, and Michael F. Olson. Actin-myosin–based contraction is responsible for apoptotic nuclear disintegration. Journal of Cell Biology, 168(2): 245–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/2/245.

Curto:2007:CDI

[CCL^{+}07] Marcello Curto, Banumathi K. Cole, Dominique Lallemand, Ching-Hui Liu, and Andrea I. McClatchey. Contact-dependent inhibition of EGFR signaling by Nf2/Merlin. Journal of Cell Biology, 177(5):893–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/5/893.

Caswell:2008:RCP

[CCL^{+}08] Patrick T. Caswell, May Chan, Andrew J. Lindsay, Mary W. McCaffrey, David Boettiger, and Jim C. Norman. Rab-coupling protein coordinates recycling of α5β1 integrin and EGFR1 to promote cell migration in 3D microenvironments. Journal of Cell Biology, 183(1):143–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/183/1/143.

Chaudhry:2007:FRB

[CCM^{+}07] Shazia S. Chaudhry, Stuart A. Cain, Amanda Morgan, Sarah L. Dallas, C. Adrian Shuttleworth, and Cay M. Kielty. Fibrillin-1 regulates the bioavailability of TGFβ1. Journal of Cell Biology, 176(3):355–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/3/355.

Cortesio:2008:CPF

[CCP^{+}08] Christa L. Cortesio, Keefe T. Chan, Benjamin J. Perrin, Nicholas O. Burton, Sheng Zhang, Zhong-Yin Zhang, and Anna Huttenlocher. Calpain 2 and PTP1B function in a novel pathway with Src to regulate invadopodia dynamics and breast cancer cell invasion. Journal of Cell Biology, 180(5):957–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/180/5/
957.

Carpp:2006:SMP

[CCS^{+}06] Lindsay N. Carpp, Leonora F. Ciufo, Scott G. Shanks, Alan Boyd, and Nia J. Bryant. The Sec1p/Munc18 protein Vps45p binds its cognate SNARE proteins via two distinct modes. Journal of Cell Biology, 173(6):927–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/173/6/927.

Caldwell:2008:OPD

[CCS^{+}08] Julie M. Caldwell, Yinhuai Chen, Kaila L. Schollaert, James F. Theis, George F. Babcock, Carol S. Newlon, and Yolanda Sanchez. Orchestration of the S-phase and DNA damage checkpoint pathways by replication forks from early origins. Journal of Cell Biology, 180(6):1073–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/6/1073.

Cheerambathur:2007:QAA

[CCSBM^{+}07] Dhanya K. Cheerambathur, Gul Civelekoglu-Scholey, Ingrid Brust-Mascher, Patrizia Sommi, Alex Mogilner, and Jonathan M. Scholey. Quantitative analysis of an anaphase B switch: predicted role for a microtubule catastrophe gradient. Journal of Cell Biology, 177(6):995–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/177/6/995.

Cao:2008:VRA

[CCSK08] Yang Cao, Heesun Cheong, Hui Song, and Daniel J. Klionsky. In vivo reconstitution of autophagy in Saccharomyces cerevisiae. Journal of Cell Biology, 182(4):703–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/182/4/703.

Cech:2007:RH

[CDM^{+}07] Tom Cech, Jack Dixon, Avice Meehan, Josie Briggs, and Carl Rhodes. A reply from HHMI. Journal of Cell Biology, 178 (3):343–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/3/343.

Chen:2005:CMC

[CDW^{+}05] Danyang Chen, Miroslav Dundr, Chen Wang, Anthony Leung, Angus Lamond, Tom Misteli, and Sui Huang. Condensed mitotic chromatin is accessible to transcription factors and chromatin structural proteins. Journal of Cell Biology, 168(1): 41–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/1/41.

Curino:2005:ICD

[CEY^{+}05] Alejandro C. Curino, Lars H. Engelholm, Susan S. Yamada, Kenn Holmbeck, Leif R. Lund, Alfredo A. Molinolo, Niels Behrendt, Boye Schnack Nielsen, and Thomas H. Bugge. Intracellular collagen degradation mediated by uPARAP/Endo180 is a major pathway of extracellular matrix turnover during malignancy. Journal of Cell Biology, 169(6):977–??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/169/6/977.

Carmo-Fonseca:2007:HGF

[CF07] Maria Carmo-Fonseca. How genes find their way inside the cell nucleus. Journal of Cell Biology, 179(6):1093–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/6/
1093.

Craske:2005:SNB

[CFBM05] Madeleine L. Craske, Marc Fivaz, Nizar N. Batada, and Tobias Meyer. Spines and neurite branches function as geometric attractors that enhance protein kinase C action. Journal of Cell Biology, 170(7):1147–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/170/7/1147.

Chen:2008:DSS

[CFH^{+}08] Xiao-Wei Chen, Ya-Qin Feng, Chan-Juan Hao, Xiao-Li Guo, Xin He, Zhi-Yong Zhou, Ning Guo, Hong-Ping Huang, Wei Xiong, Hui Zheng, Pan-Li Zuo, Claire Xi Zhang, Wei Li, and Zhuan Zhou. DTNBP1, a schizophrenia susceptibility gene, affects kinetics of transmitter release. Journal of Cell Biology, 181(5):791–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/5/791.

Chen:2005:CBC

[CFL^{+}05] Ji-Long Chen, Raymond V. Fucini, Lynne Lacomis, Hediye Erdjument-Bromage, Paul Tempst, and Mark Stamnes. Coatomer-bound Cdc42 regulates dynein recruitment to COPI vesicles. Journal of Cell Biology, 169(3):383–??, May 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/169/3/383.

Chan:2009:MCK

[CFU^{+}09] Ying Wai Chan, Luca L. Fava, Andreas Uldschmid, Michael H. A. Schmitz, Daniel W. Gerlich, Erich A. Nigg, and Anna Santamaria. Mitotic control of kinetochore-associated dynein and spindle orientation by human Spindly. Journal of Cell Biology, 185(5):859–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/5/859.

Cumming:2007:EBE

[CFV07] Geoff Cumming, Fiona Fidler, and David L. Vaux. Error bars in experimental biology. Journal of Cell Biology, 177(1):7–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
177/1/7.

Chen:2006:PPM

[CG06] Xuejun Chen and Barry M. Gumbiner. Paraxial protocadherin mediates cell sorting and tissue morphogenesis by regulating Ccadherin adhesion activity. Journal of Cell Biology, 174(2):301– ??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
174/2/301.

Courtheoux:2009:APD

[CGGT09] Thibault Courtheoux, Guillaume Gay, Yannick Gachet, and Sylvie Tournier. Ase1/ prc1-dependent spindle elongation corrects merotely during anaphase in fission yeast. Journal of Cell Biology, 187(3):399–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/187/3/399.

Cui:2009:SYP

[CGJ09] Hong Cui, Santanu K. Ghosh, and Makkuni Jayaram. The selfish yeast plasmid uses the nuclear motor Kip1p but not Cin8p for its localization and equal segregation. Journal of Cell Biology, 185(2):251–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/2/251.

Caldwell:2007:AML

[CGK07] Christine M. Caldwell, Rebecca A. Green, and Kenneth B. Kaplan. APC mutations lead to cytokinetic failures in vitro and tetraploid genotypes in Min mice. Journal of Cell Biology, 178(7):1109–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/7/1109.

Campa:2008:NAC

[CGLC^{+}08] Víctor M. Campa, Raquel Gutiérrez-Lanza, Fabio Cerignoli, Ramón Díaz-Trelles, Brandon Nelson, Toshiya Tsuji, Maria Barcova, Wei Jiang, and Mark Mercola. Notch activates cell cycle reentry and progression in quiescent cardiomyocytes. Journal of Cell Biology, 183(1):129–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/183/1/129.

Cai:2005:SMA

[CGS05] Qian Cai, Claudia Gerwin, and Zu-Hang Sheng. Syntabulinmediated anterograde transport of mitochondria along neuronal processes. Journal of Cell Biology, 170(6):959–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/170/6/
959.

Campbell:2005:DCE

[CHA05] Susan G. Campbell, Nathaniel P. Hoyle, and Mark P. Ashe. Dynamic cycling of eIF2 through a large eIF2B-containing cytoplasmic body. Journal of Cell Biology, 170(6):925–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/170/6/
925.

Chan:2007:MAA

[Cha07] Jonah R. Chan. Myelination: all about Rac ‘n’ roll. Journal of Cell Biology, 177(6):953–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/177/6/953.

Chang:2005:WBT

[CHCB05] Fanny S. Chang, Gil-Soo Han, George M. Carman, and Kendall J. Blumer. A WASp-binding type II phosphatidylinositol 4-kinase required for actin polymerization-driven endosome motility. Journal of Cell Biology, 171(1):133–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/171/1/133.

Charras:2006:RCA

[CHCM06] Guillaume T. Charras, Chi-Kuo Hu, Margaret Coughlin, and Timothy J. Mitchison. Reassembly of contractile actin cortex in cell blebs. Journal of Cell Biology, 175(3):477–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/175/3/
477.

Chao:2008:SMP

[CHH^{+}08] Hsu-Wen Chao, Chen-Jei Hong, Tzyy-Nan Huang, Yi-Ling Lin, and Yi-Ping Hsueh. SUMOylation of the MAGUK protein CASK regulates dendritic spinogenesis. Journal of Cell Biology, 182(1):141–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/1/141.

Choi:2009:UDC

[Cho09a] Eun Choi. Ubiquitin drives cilia shortening. Journal of Cell Biology, 186(4):??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/186/4/448.3.

Clerc:2009:HPG

[CHO^{+}09b] Simone Clerc, Christian Hirsch, Daniela Maria Oggier, Paola Deprez, Claude Jakob, Thomas Sommer, and Markus Aebi. Htm1 protein generates the N-glycan signal for glycoprotein degradation in the endoplasmic reticulum. Journal of Cell Biology, 184(1):159–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/1/159.

Cai:2007:KSO

[CHSV07] Dawen Cai, Adam D. Hoppe, Joel A. Swanson, and Kristen J. Verhey. Kinesin-1 structural organization and conformational changes revealed by FRET stoichiometry in live cells. Journal of Cell Biology, 176(1):51–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/176/1/51.

Cheng:2005:ICR

[CJMG05] Zhiliang Cheng, Ying Jiang, Elisabet C. Mandon, and Reid Gilmore. Identification of cytoplasmic residues of Sec61p involved in ribosome binding and cotranslational translocation. Journal of Cell Biology, 168(1):67–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/168/1/67.

Cabeen:2007:SBB

[CJW07] Matthew T. Cabeen and Christine Jacobs-Wagner. Skin and bones: the bacterial cytoskeleton, cell wall, and cell morphogenesis. Journal of Cell Biology, 179(3):381–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/179/3/381.

Cho:2007:LGP

[CK07] Hyeseon Cho and John H. Kehrl. Localization of G_{iα} proteins in the centrosomes and at the midbody: implication for their role in cell division. Journal of Cell Biology, 178(2):245–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/178/2/

Carreno:2008:MAK

[CKG^{+}08] Sébastien Carreno, Ilektra Kouranti, Edith Szafer Glusman, Margaret T. Fuller, Arnaud Echard, and François Payre. Moesin and its activating kinase Slik are required for cortical stability and microtubule organization in mitotic cells. Journal of Cell Biology, 180(4):739–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/180/4/739.

Cavalli:2005:SDL

[CKKG05] Valeria Cavalli, Pekka Kujala, Judith Klumperman, and Lawrence S. B. Goldstein. Sunday Driver links axonal transport to damage signaling. Journal of Cell Biology, 168(5): 775–??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/5/775.

Chen:2008:RRS

[CKX^{+}08] Kai Chen, Michael T. Kirber, Hui Xiao, Yu Yang, and John F. Keaney. Regulation of ROS signal transduction by NADPH oxidase 4 localization. Journal of Cell Biology, 181(7):1129–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
181/7/1129.

Colombo:2005:MDD

[CLA^{+}05] Sara Colombo, Renato Longhi, Stefano Alcaro, Francesco Ortuso, Teresa Sprocati, Adriano Flora, and Nica Borgese. N-myristoylation determines dual targeting of mammalian NADH-cytochrome b(5) reductase to ER and mitochondrial outer membranes by a mechanism of kinetic partitioning. Journal of Cell Biology, 168(5):735–??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/168/5/735.

Couwenbergs:2007:HGP

[CLG^{+}07] Claudia Couwenbergs, Jean-Claude Labbé, Morgan Goulding, Thomas Marty, Bruce Bowerman, and Monica Gotta. Heterotrimeric G protein signaling functions with dynein to promote spindle positioning in C. elegans. Journal of Cell Biology, 179(1):15–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/1/15.

Chung:2007:CCS

[CLK^{+}07] S. Clare Chung, Andre Limnander, Tomohiro Kurosaki, Arthur Weiss, and Juan I. Korenbrot. Coupling Ca^{2+} store release to Icrac channel activation in B lymphocytes requires the activity of Lyn and Syk kinases. Journal of Cell Biology, 177(2):317–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
177/2/317.

Cao:2007:FHC

[CLL07] Chike Cao, Yanling Liu, and Michael Lehmann. Fork head controls the timing and tissue selectivity of steroid-induced developmental cell death. Journal of Cell Biology, 176(6):843–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
176/6/843.

Crisp:2006:CNC

[CLR^{+}06] Melissa Crisp, Qian Liu, Kyle Roux, J. B. Rattner, Catherine Shanahan, Brian Burke, Phillip D. Stahl, and Didier Hodzic. Coupling of the nucleus and cytoplasm: Role of the LINC complex. Journal of Cell Biology, 172(1):41–??, January 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/172/1/41.

Chapman:2006:HLN

[CLS^{+}06] Gavin Chapman, Lining Liu, Cecilia Sahlgren, Camilla Dahlqvist, and Urban Lendahl. High levels of Notch signaling down-regulate Numb and Numblike. Journal of Cell Biology, 175(4):535–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/4/535.

Chen:2006:PMI

[CM06] Xinyu Chen and Ian G. Macara. Par-3 mediates the inhibition of LIM kinase 2 to regulate cofilin phosphorylation and tight junction assembly. Journal of Cell Biology, 172(5): 671–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/5/671.

Campbell:2007:VVT

[CM07] Christopher S. Campbell and R. Dyche Mullins. In vivo visualization of type II plasmid segregation: bacterial actin filaments pushing plasmids. Journal of Cell Biology, 179(5):1059–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
179/5/1059.

Cook:2009:EOI

[CM09] Peter R. Cook and Davide Marenduzzo. Entropic organization of interphase chromosomes. Journal of Cell Biology, 186(6): 825–??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/186/6/825.

Cakouros:2008:DSR

[CMD^{+}08] Dimitrios Cakouros, Kathryn Mills, Donna Denton, Alicia Paterson, Tasman Daish, and Sharad Kumar. dLKR/ SDH regulates hormone-mediated histone arginine methylation and transcription of cell death genes. Journal of Cell Biology, 182(3): 481–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/3/481.

Chiu:2008:MEC

[CMF08] Yi-Jen Chiu, Elena McBeath, and Keigi Fujiwara. Mechanotransduction in an extracted cell model: Fyn drives stretchand flow-elicited PECAM-1 phosphorylation. Journal of Cell Biology, 182(4):753–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/4/753.

Chang:2009:PPB

[CMF^{+}09] Jinlan Chang, Fred D. Mast, Andrei Fagarasanu, Dorian A. Rachubinski, Gary A. Eitzen, Joel B. Dacks, and Richard A. Rachubinski. Pex3 peroxisome biogenesis proteins function in peroxisome inheritance as class V myosin receptors. Journal of Cell Biology, 187(2):233–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/187/2/233.

Claypool:2006:MMA

[CMK06] Steven M. Claypool, J. Michael McCaffery, and Carla M. Koehler. Mitochondrial mislocalization and altered assembly of a cluster of Barth syndrome mutant tafazzins. Journal of Cell Biology, 174(3):379–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/3/379.

Czuchra:2006:GAI

[CML^{+}06] Aleksandra Czuchra, Hannelore Meyer, Kyle R. Legate, Cord Brakebusch, and Reinhard Fässler. Genetic analysis of β1 integrin “activation motifs” in mice. Journal of Cell Biology, 174(6): 889–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/6/889.

Cowling:2008:IFR

[CMN^{+}08] Belinda S. Cowling, Meagan J. McGrath, Mai-Anh Nguyen, Denny L. Cottle, Anthony J. Kee, Susan Brown, Joachim Schessl, Yaqun Zou, Josephine Joya, Carsten G. Bönnemann, Edna C. Hardeman, and Christina A. Mitchell. Identification of FHL1 as a regulator of skeletal muscle mass: implications for human myopathy. Journal of Cell Biology, 183(6):1033–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
183/6/1033.

Cooper:2006:CCM

[CN06] John A. Cooper and Scott A. Nelson. Checkpoint control of mitotic exit — do budding yeast mind the GAP? Journal of Cell Biology, 172(3):331–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/172/3/331.

Cascone:2005:SIB

[CNM^{+}05] Ilaria Cascone, Lucia Napione, Fabrizio Maniero, Guido Serini, and Federico Bussolino. Stable interaction between α5β1 integrin and Tie2 tyrosine kinase receptor regulates endothelial cell response to Ang-1. Journal of Cell Biology, 170(6): 993–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/6/993.

Corbett-Nelson:2006:SDI

[CNMM^{+}06] Elaine F. Corbett-Nelson, David Mason, John G. Marshall, Yves Collette, and Sergio Grinstein. Signaling-dependent immobilization of acylated proteins in the inner monolayer of the plasma membrane. Journal of Cell Biology, 174(2):255–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/174/2/
255.

Casaletto:2005:IAP

[CNW^{+}05] Jessica B. Casaletto, Leta K. Nutt, Qiju Wu, Jonathan D. Moore, Laurence D. Etkin, Peter K. Jackson, Tim Hunt, and Sally Kornbluth. Inhibition of the anaphase-promoting complex by the Xnf7 ubiquitin ligase. Journal of Cell Biology, 169(1):61–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/1/61.

Claypool:2008:CDI

[COB^{+}08] Steven M. Claypool, Yavuz Oktay, Pinmanee Boontheung, Joseph A. Loo, and Carla M. Koehler. Cardiolipin defines the interactome of the major ADP/ATP carrier protein of the mitochondrial inner membrane. Journal of Cell Biology, 182(5): 937–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/5/937.

Cannon:2005:DBF

[COC^{+}05] Kurt S. Cannon, Eran Or, William M. Clemons, Yoko Shibata, and Tom A. Rapoport. Disulfide bridge formation between SecY and a translocating polypeptide localizes the translocation pore to the center of SecY. Journal of Cell Biology, 169 (2):219–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/2/219.

Compton:2007:CO

[Com07] Duane A. Compton. Chromosome orientation. Journal of Cell Biology, 179(2):179–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/2/179.

Chorna-Ornan:2005:LRI

[COTAE^{+}05] Irit Chorna-Ornan, Vered Tzarfaty, Galit Ankri-Eliahoo, Tamar Joel-Almagor, Nina E. Meyer, Armin Huber, François Payre, and Baruch Minke. Light-regulated interaction of Dmoesin with TRP and TRPL channels is required for maintenance of photoreceptors. Journal of Cell Biology, 171(1): 143–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/1/143.

Campbell:2008:VPI

[CPAH08] Edward M. Campbell, Omar Perez, Jenny L. Anderson, and Thomas J. Hope. Visualization of a proteasome-independent intermediate during restriction of HIV-1 by rhesus TRIM5α. Journal of Cell Biology, 180(3):549–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/180/3/549.

Coloff:2006:MRA

[CR06] Jonathan L. Coloff and Jeffrey C. Rathmell. Metabolic regulation of Akt: roles reversed. Journal of Cell Biology, 175(6): 845–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/6/845.

Chiroli:2007:BYP

[CRLP07] Elena Chiroli, Valentina Rossio, Giovanna Lucchini, and Simonetta Piatti. The budding yeast PP2A^{Cdc55} protein phosphatase prevents the onset of anaphase in response to morphogenetic defects. Journal of Cell Biology, 177(4):599–??, May 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/177/4/599.

Chun:2006:CP

[CRS^{+}06] Soo Jin Chun, Matthew N. Rasband, Richard L. Sidman, Amyn A. Habib, and Timothy Vartanian. Correction 163, p.

397. Journal of Cell Biology, 174(2):315–??, July 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/174/2/315.

Csordas:2006:SFF

[CRV^{+}06] György Csordás, Christian Renken, Péter Várnai, Ludivine Walter, David Weaver, Karolyn F. Buttle, Tamás Balla, Carmen A. Mannella, and György Hajnóczky. Structural and functional features and significance of the physical linkage between ER and mitochondria. Journal of Cell Biology, 174(7): 915–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/7/915.

Carroll:2007:CCG

[CS07] Christopher W. Carroll and Aaron F. Straight. Centromeric chromatin gets loaded. Journal of Cell Biology, 176(6):735–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
176/6/735.

Canudas:2009:DRT

[CS09] Silvia Canudas and Susan Smith. Differential regulation of telomere and centromere cohesion by the Scc3 homologues SA1 and SA2, respectively, in human cells. Journal of Cell Biology, 187(2):165–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/2/165.

Czibener:2006:CSV

[CSB^{+}06] Cecilia Czibener, Nathan M. Sherer, Steven M. Becker, Marc Pypaert, Enfu Hui, Edwin R. Chapman, Walther Mothes, and Norma W. Andrews. Ca^{2+} and synaptotagmin VII-dependent delivery of lysosomal membrane to nascent phagosomes. Journal of Cell Biology, 174(7):997–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/174/7/997.

Cluzel:2005:MDI

[CSL^{+}05] Caroline Cluzel, Frédéric Saltel, Jost Lussi, Frédérique Paulhe, Beat A. Imhof, and Bernhard Wehrle-Haller. The mechanisms and dynamics of α vβ3 integrin clustering in living cells. Journal of Cell Biology, 171(2):383–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/171/2/383.

Chiba:2006:NRH

[CSM^{+}06] Hideki Chiba, Naoyuki Sakai, Masaki Murata, Makoto Osanai, Takafumi Ninomiya, Takashi Kojima, and Norimasa Sawada. The nuclear receptor hepatocyte nuclear factor 4α acts as a morphogen to induce the formation of microvilli. Journal of Cell Biology, 175(6):971–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/175/6/971.

Cerqueira:2009:OCA

[CSMP^{+}09] Antonio Cerqueira, David Santamaría, Bárbara Martínez-Pastor, Miriam Cuadrado, Oscar Fernández-Capetillo, and Mariano Barbacid. Overall Cdk activity modulates the DNA damage response in mammalian cells. Journal of Cell Biology, 187(6):773–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/6/773.

Chu:2006:NCE

[CSSE06] Tony Chu, Ji Sun, Suraj Saksena, and Scott D. Emr. New component of ESCRT–I regulates endosomal sorting complex assembly. Journal of Cell Biology, 175(5):815–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/175/5/
815.

Chen:2008:NPC

[CSSvM08] Min Chen, Lena Singer, Andrea Scharf, and Anna von Mikecz. Nuclear polyglutamine-containing protein aggregates as active proteolytic centers. Journal of Cell Biology, 180(4): 697–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/4/697.

Chang:2006:AIF

[CSTT^{+}06] Lynne Chang, Yaron Shav-Tal, Tatjana Trcek, Robert H. Singer, and Robert D. Goldman. Assembling an intermediate filament network by dynamic cotranslation. Journal of Cell Biology, 172(5):747–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/5/747.

Chia:2008:DNA

[CSW08] William Chia, W. Gregory Somers, and Hongyan Wang. Drosophila neuroblast asymmetric divisions: cell cycle regulators, asymmetric protein localization, and tumorigenesis. Journal of Cell Biology, 180(2):267–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/180/2/267.

Casas-Tinto:2008:FMT

[CTGVGFF08] Sergio Casas-Tinto, Melisa Gomez-Velazquez, Begoña Granadino, and Pedro Fernandez-Funez. FoxK mediates TGF-β signalling during midgut differentiation in flies. Journal of Cell Biology, 183(6):1049–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/6/1049.

Culjkovic:2008:ERR

[CTO^{+}08] Biljana Culjkovic, Keith Tan, Slobodanka Orolicki, Abdellatif Amri, Sylvain Meloche, and Katherine L. B. Borden. The eIF4E RNA regulon promotes the Akt signaling pathway. Journal of Cell Biology, 181(1):51–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/1/51.

Chong:2009:REM

[CTP^{+}09] Han Chung Chong, Ming Jie Tan, Virginie Philippe, Siew Hwey Tan, Chek Kun Tan, Chee Wai Ku, Yan Yih Goh, Walter Wahli, Liliane Michalik, and Nguan Soon Tan. Regulation of epithelial–mesenchymal IL-1 signaling by PPARβ /δ is essential for skin homeostasis and wound healing. Journal of Cell Biology, 184(6):817–??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/6/817.

Culjkovic:2005:EPN

[CTS^{+}05] Biljana Culjkovic, Ivan Topisirovic, Lucy Skrabanek, Melisa Ruiz-Gutierrez, and Katherine L. B. Borden. eIF4E promotes nuclear export of cyclin D1 mRNAs via an element in the 3’UTR. Journal of Cell Biology, 169(2):245–??, April 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/169/2/245.

Cai:2006:ATI

[CTS^{+}06a] Sheng-Li Cai, Andrew R. Tee, John D. Short, Judith M. Bergeron, Jinhee Kim, Jianjun Shen, Ruifeng Guo, Charles L. Johnson, Kaoru Kiguchi, and Cheryl Lyn Walker. Activity of TSC2 is inhibited by AKT-mediated phosphorylation and membrane partitioning. Journal of Cell Biology, 173(2):279–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/173/2/279.

Clement:2006:NRP

[CTS^{+}06b] Christian Clement, Vaibhav Tiwari, Perry M. Scanlan, Tibor Valyi-Nagy, Beatrice Y. J. T. Yue, and Deepak Shukla. A novel role for phagocytosis-like uptake in herpes simplex virus entry. Journal of Cell Biology, 174(7):1009–??, September 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/174/7/1009.

Culjkovic:2006:ECN

[CTS^{+}06c] Biljana Culjkovic, Ivan Topisirovic, Lucy Skrabanek, Melisa Ruiz-Gutierrez, and Katherine L. B. Borden. eIF4E is a central node of an RNA regulon that governs cellular proliferation. Journal of Cell Biology, 175(3):415–??, November 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/175/3/415.

Chen:2006:TAC

[CTY^{+}06] Yun Chen, William R. Thelin, Bing Yang, Sharon L. Milgram, and Ken Jacobson. Transient anchorage of cross-linked glycosylphosphatidylinositol–anchored proteins depends on cholesterol, Src family kinases, caveolin, and phosphoinositides. Journal of Cell Biology, 175(1):169–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/175/1/169.

Cibrian-Uhalte:2007:IZK

[CULS^{+}07] Elena Cibrián-Uhalte, Adam Langenbacher, Xiaodong Shu, Jau-Nian Chen, and Salim Abdelilah-Seyfried. Involvement of zebrafish Na^{+} ,K^{+} ATPase in myocardial cell junction maintenance. Journal of Cell Biology, 176(2):223–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/176/2/223.

Custodio:2007:SCI

[CVACF07] Noélia Custódio, Maria Vivo, Michael Antoniou, and Maria Carmo-Fonseca. Splicing- and cleavage-independent requirement of RNA polymerase II CTD for mRNA release from the transcription site. Journal of Cell Biology, 179(2):199–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/2/
199.

Cao:2005:PNI

[CVD^{+}05] Jiong Cao, Jenni I. Viholainen, Caroline Dart, Helen K. Warwick, Mark L. Leyland, and Michael J. Courtney. The PSD95– nNOS interface. Journal of Cell Biology, 168(1):117–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
168/1/117.

Cude:2007:RGM

[CWC^{+}07] Kelly Cude, Yupeng Wang, Hyun-Jung Choi, Shih-Ling Hsuan, Honglai Zhang, Cun-Yu Wang, and Zhengui Xia. Regulation of the G2–M cell cycle progression by the ERK5-NFκB signaling pathway. Journal of Cell Biology, 177(2):253–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/177/2/253.

Cai:2008:RCM

[CWC^{+}08] Ting Cai, Haojie Wang, Yiliang Chen, Lijun Liu, William T. Gunning, Luis Eduardo M. Quintas, and Zi-Jian Xie. Regulation of caveolin-1 membrane trafficking by the Na/K–ATPase. Journal of Cell Biology, 182(6):1153–??, September 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/182/6/1153.

Conforti:2009:WPR

[CWM^{+}09] Laura Conforti, Anna Wilbrey, Giacomo Morreale, Lucie Janeckova, Bogdan Beirowski, Robert Adalbert, Francesca Mazzola, Michele Di Stefano, Robert Hartley, Elisabetta Babetto, Trevor Smith, Jonathan Gilley, Richard A. Billington, Armando A. Genazzani, Richard R. Ribchester, Giulio Magni, and Michael Coleman. Wld^{S} protein requires Nmnat activity and a short N-terminal sequence to protect axons in mice. Journal of Cell Biology, 184(4):491–??, February 2009. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/184/4/491.

Camara:2009:IMA

[CWNF^{+}09] Joana Câmara, Zhen Wang, Cristina Nunes-Fonseca, Hana C. Friedman, Matthew Grove, Diane L. Sherman, Noboru H. Komiyama, Seth G. Grant, Peter J. Brophy, Alan Peterson, and Charles ffrench Constant. Integrin-mediated axoglial interactions initiate myelination in the central nervous system. Journal of Cell Biology, 185(4):699–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/185/4/699.

Clifford:2008:CCP

[CWRG^{+}08] Dawn M. Clifford, Benjamin A. Wolfe, Rachel H. Roberts-Galbraith, W. Hayes McDonald, John R. Yates, and Kathleen L. Gould. The Clp1/Cdc14 phosphatase contributes to the robustness of cytokinesis by association with anillin-related Mid1. Journal of Cell Biology, 181(1):79–??, April 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/181/1/79.

Cameron:2006:KIP

[CYC^{+}06] Lisa A. Cameron, Ge Yang, Daniela Cimini, Julie C. Canman, Olga Kisurina-Evgenieva, Alexey Khodjakov, Gaudenz Danuser, and E. D. Salmon. Kinesin 5–independent poleward flux of kinetochore microtubules in PtK1 cells. Journal of Cell Biology, 173(2):173–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/2/173.

Chikashige:2009:MPB

[CYO^{+}09] Yuji Chikashige, Miho Yamane, Kasumi Okamasa, Chihiro Tsutsumi, Tomoko Kojidani, Mamiko Sato, Tokuko Haraguchi, and Yasushi Hiraoka. Membrane proteins Bqt3 and -4 anchor telomeres to the nuclear envelope to ensure chromosomal bouquet formation. Journal of Cell Biology, 187(3):413–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/187/3/

Chen:2008:PFL

[CYY^{+}08] Zu-Lin Chen, Huaxu Yu, Wei-Ming Yu, Robert Pawlak, and Sidney Strickland. Proteolytic fragments of laminin promote excitotoxic neurodegeneration by up-regulation of the KA1 subunit of the kainate receptor. Journal of Cell Biology, 183(7): 1299–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/7/1299.

Catalucci:2009:ART

[CZD^{+}09] Daniele Catalucci, Deng-Hong Zhang, Jaime DeSantiago, Franck Aimond, Guillaume Barbara, Jean Chemin, Désiré Bonci, Eckard Picht, Francesca Rusconi, Nancy D. Dalton, Kirk L. Peterson, Sylvain Richard, Donald M. Bers, Joan Heller Brown, and Gianluigi Condorelli. Akt regulates L-type Ca^{2+} channel activity by modulating Ca_{v} α1 protein stability. Journal of Cell Biology, 184(6):923–??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/184/6/923.

Chernomordik:2006:MWU

[CZK06] Leonid V. Chernomordik, Joshua Zimmerberg, and Michael M. Kozlov. Membranes of the world unite! Journal of Cell Biology, 175(2):201–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/2/201.

Caiolfa:2007:MDD

[CZM^{+}07] Valeria R. Caiolfa, Moreno Zamai, Gabriele Malengo, Annapaola Andolfo, Chris D. Madsen, Jason Sutin, Michelle A. Digman, Enrico Gratton, Francesco Blasi, and Nicolai Sidenius. Monomer–dimer dynamics and distribution of GPI-anchored uPAR are determined by cell surface protein assemblies. Journal of Cell Biology, 179(5):1067–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/5/1067.

Cai:2005:MTD

[CZP^{+}05] Huaqing Cai, Yueyi Zhang, Marc Pypaert, Lee Walker, and Susan Ferro-Novick. Mutants in trs120 disrupt traffic from the early endosome to the late Golgi. Journal of Cell Biology, 171(5):823–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/5/823.

Collesi:2008:NSS

[CZSG08] Chiara Collesi, Lorena Zentilin, Gianfranco Sinagra, and Mauro Giacca. Notch1 signaling stimulates proliferation of immature cardiomyocytes. Journal of Cell Biology, 183(1):117–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/183/1/
117.

Dalva:2009:NAM

[Dal09] Matthew B. Dalva. Neuronal activity moves protein palmitoylation into the synapse. Journal of Cell Biology, 186(1):7–??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
186/1/7.

Demarchi:2006:CRM

[DBC^{+}06] Francesca Demarchi, Cosetta Bertoli, Tamara Copetti, Isei Tanida, Claudio Brancolini, Eeva-Liisa Eskelinen, and Claudio Schneider. Calpain is required for macroautophagy in mammalian cells. Journal of Cell Biology, 175(4):595–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/175/4/
595.

Demonchy:2009:KFM

[DBD^{+}09] Raphaël Demonchy, Thierry Blisnick, Caroline Deprez, Géraldine Toutirais, Céline Loussert, William Marande, Philippe Grellier, Philippe Bastin, and Linda Kohl. Kinesin 9 family members perform separate functions in the trypanosome flagellum. Journal of Cell Biology, 187(5):615–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/187/5/615.

Das:2006:SFU

[DBDD06] Amlan Das, Christine Base, Srilakshmi Dhulipala, and Ronald R. Dubreuil. Spectrin functions upstream of ankyrin in a spectrin cytoskeleton assembly pathway. Journal of Cell Biology, 175(2):325–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/2/325.

Deckbar:2007:CBA

[DBK^{+}07] Dorothee Deckbar, Julie Birraux, Andrea Krempler, Leopoldine Tchouandong, Andrea Beucher, Sarah Walker, Tom Stiff, Penny Jeggo, and Markus Löbrich. Chromosome breakage after G2 checkpoint release. Journal of Cell Biology, 176(6):749–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
176/6/749.

Dohn:2009:ERP

[DBR09] Michael R. Dohn, Meredith V. Brown, and Albert B. Reynolds. An essential role for p120-catenin in Src- and Rac1-mediated anchorage-independent cell growth. Journal of Cell Biology, 184 (3):437–??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/184/3/437.

Diks:2006:NGA

[DBvdW^{+}06] Sander H. Diks, Robert J. Bink, Sandra van de Water, Jos Joore, Carina van Rooijen, Fons J. Verbeek, Jeroen den Hertog, Maikel P. Peppelenbosch, and Danica Zivkovic. The novel gene asb11: a regulator of the size of the neural progenitor compartment. Journal of Cell Biology, 174(4):581–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/174/4/
581.

Deinhardt:2006:TTI

[DBW^{+}06] Katrin Deinhardt, Otto Berninghausen, Hugh J. Willison, Colin R. Hopkins, and Giampietro Schiavo. Tetanus toxin is internalized by a sequential clathrin-dependent mechanism initiated within lipid microdomains and independent of epsin1. Journal of Cell Biology, 174(3):459–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/174/3/459.

Detmer:2007:CBM

[DC07] Scott A. Detmer and David C. Chan. Complementation between mouse Mfn1 and Mfn2 protects mitochondrial fusion defects caused by CMT2A disease mutations. Journal of Cell Biology, 176(4):405–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/4/405.

Dellaire:2006:PLN

[DCA^{+}06] Graham Dellaire, Reagan W. Ching, Kashif Ahmed, Farid Jalali, Kenneth C. K. Tse, Robert G. Bristow, and David P. Bazett-Jones. Promyelocytic leukemia nuclear bodies behave as DNA damage sensors whose response to DNA double-strand breaks is regulated by NBS1 and the kinases ATM, Chk2, and ATR. Journal of Cell Biology, 175(1):55–??, October 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/175/1/55.

Dayer:2005:NGI

[DCAC05] Alexandre G. Dayer, Kathryn M. Cleaver, Thamara Abouantoun, and Heather A. Cameron. New GABAergic interneurons in the adult neocortex and striatum are generated from different precursors. Journal of Cell Biology, 168(3):415–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/168/3/
415.

deSouza:2007:FTP

[dCD^{+}07] Nikhil deSouza, Jie Cui, Miroslav Dura, Thomas V. McDonald, and Andrew R. Marks. A function for tyrosine phosphorylation of type 1 inositol 1,4,5-trisphosphate receptor in lymphocyte activation. Journal of Cell Biology, 179(5):923–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/5/
923.

Dupin:2009:CCC

[DCEM09] Isabelle Dupin, Emeline Camand, and Sandrine Etienne-Manneville. Classical cadherins control nucleus and centrosome position and cell polarity. Journal of Cell Biology, 185(5):779– ??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
185/5/779.

Durand:2007:INM

[DCMB^{+}07] Sébastien Durand, Nicolas Cougot, Florence Mahuteau-Betzer, Chi-Hung Nguyen, David S. Grierson, Edouard Bertrand, Jamal Tazi, and Fabrice Lejeune. Inhibition of nonsense-mediated mRNA decay (NMD) by a new chemical molecule reveals the dynamic of NMD factors in P-bodies. Journal of Cell Biology, 178(7):1145–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/7/1145.

Drenan:2005:PRP

[DDB^{+}05] Ryan M. Drenan, Craig A. Doupnik, Maureen P. Boyle, Louis J. Muglia, James E. Huettner, Maurine E. Linder, and Kendall J. Blumer. Palmitoylation regulates plasma membrane–nuclear shuttling of R7BP, a novel membrane anchor for the RGS7 family. Journal of Cell Biology, 169(4):623–??, May 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/169/4/623.

Dewitt:2009:TJL

[DDH09] Sharon Dewitt, Richard L. Darley, and Maurice B. Hallett. Translocation or just location? Pseudopodia affect fluorescent signals. Journal of Cell Biology, 184(2):197–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/184/2/197.

Doulet:2006:NMI

[DDLC^{+}06] Nicolas Doulet, Emmanuel Donnadieu, Marie-Pierre Laran-Chich, Florence Niedergang, Xavier Nassif, Pierre Olivier Couraud, and Sandrine Bourdoulous. Neisseria meningitidis infection of human endothelial cells interferes with leukocyte transmigration by preventing the formation of endothelial docking structures. Journal of Cell Biology, 173(4):627–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/173/4/
627.

Degtyarev:2008:AIP

[DDO^{+}08] Michael Degtyarev, Ann De Mazière, Christine Orr, Jie Lin, Brian B. Lee, Janet Y. Tien, Wei W. Prior, Suzanne van Dijk, Hong Wu, Daniel C. Gray, David P. Davis, Howard M. Stern, Lesley J. Murray, Klaus P. Hoeflich, Judith Klumperman, Lori S. Friedman, and Kui Lin. Akt inhibition promotes autophagy and sensitizes PTEN-null tumors to lysosomotropic agents. Journal of Cell Biology, 183(1):101–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/183/1/101.

DeVay:2009:CMI

[DDRL^{+}09] Rachel M. DeVay, Lenin Dominguez-Ramirez, Laura L. Lackner, Suzanne Hoppins, Henning Stahlberg, and Jodi Nunnari. Coassembly of Mgm1 isoforms requires cardiolipin and mediates mitochondrial inner membrane fusion. Journal of Cell Biology, 186(6):793–??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/186/6/793.

deCathelineau:2009:SDR

[deC09] Aimee deCathelineau. Sculpting the dynein regulatory complex. Journal of Cell Biology, 187(6):753–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/187/6/753.

Dentler:2005:ITI

[Den05] William Dentler. Intraflagellar transport (IFT) during assembly and disassembly of Chlamydomonas flagella. Journal of Cell Biology, 170(4):649–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/4/649.

DiBiase:2005:EST

[DFA05] Valentina Di Biase and Clara Franzini-Armstrong. Evolution of skeletal type e–c coupling. Journal of Cell Biology, 171(4): 695–??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/4/695.

Deponti:2007:NMS

[DFB^{+}07] Daniela Deponti, Stéphanie François, Silvia Baesso, Clara Sciorati, Anna Innocenzi, Vania Broccoli, Françoise Muscatelli, Raffaella Meneveri, Emilio Clementi, Giulio Cossu, and Silvia Brunelli. Necdin mediates skeletal muscle regeneration by promoting myoblast survival and differentiation. Journal of Cell Biology, 179(2):305–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/2/305.

DiVirgilio:2005:RDS

[DG05] Michela Di Virgilio and Jean Gautier. Repair of double-strand breaks by nonhomologous end joining in the absence of Mre11.

Journal of Cell Biology, 171(5):765–??, December 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/171/5/765.

Devaux:2008:TJP

[DG08] Jerome Devaux and Alexander Gow. Tight junctions potentiate the insulative properties of small CNS myelinated axons. Journal of Cell Biology, 183(5):909–??, December 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/183/5/909.

Dange:2008:ART

[DGG^{+}08] Thomas Dange, David Grünwald, Antje Grünwald, Reiner Peters, and Ulrich Kubitscheck. Autonomy and robustness of translocation through the nuclear pore complex: a singlemolecule study. Journal of Cell Biology, 183(1):77–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/183/1/
77.

deGraffenried:2008:PLK

[dGHW08] Christopher L. de Graffenried, Helen H. Ho, and Graham Warren. Polo-like kinase is required for Golgi and bilobe biogenesis in Trypanosoma brucei. Journal of Cell Biology, 181(3):431–??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
181/3/431.

Dobrowolny:2005:MEL

[DGP^{+}05] Gabriella Dobrowolny, Cristina Giacinti, Laura Pelosi, Carmine Nicoletti, Nadine Winn, Laura Barberi, Mario Molinaro, Nadia Rosenthal, and Antonio Musarò. Muscle expression of a local Igf-1 isoform protects motor neurons in an ALS mouse model. Journal of Cell Biology, 168(2):193–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/168/2/193.

Dantuma:2006:DUE

[DGSN06] Nico P. Dantuma, Tom A. M. Groothuis, Florian A. Salomons, and Jacques Neefjes. A dynamic ubiquitin equilibrium couples proteasomal activity to chromatin remodeling. Journal of Cell Biology, 173(1):19–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/1/19.

Strong:2006:LDU

[dGSWW^{+}06] Cristina de Guzman Strong, Philip W. Wertz, Chenwei Wang, Fan Yang, Paul S. Meltzer, Thomas Andl, Sarah E. Millar, I-Cheng Ho, Sung-Yun Pai, and Julia A. Segre. Lipid defect underlies selective skin barrier impairment of an epidermalspecific deletion of Gata-3. Journal of Cell Biology, 175(4): 661–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/4/661.

Delacour:2005:GSA

[DGZ^{+}05] Delphine Delacour, Valérie Gouyer, Jean-Pierre Zanetta, Hervé Drobecq, Emmanuelle Leteurtre, Georges Grard, Odile Moreau-Hannedouche, Emmanuel Maes, Alexandre Pons, Sabine André, André Le Bivic, Hans Joachim Gabius, Aki Manninen, Kai Simons, and Guillemette Huet. Galectin-4 and sulfatides in apical membrane trafficking in enterocyte-like cells. Journal of Cell Biology, 169(3):491–??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/3/491.

Duelli:2005:PVG

[DHML05] Dominik M. Duelli, Stephen Hearn, Michael P. Myers, and Yuri Lazebnik. A primate virus generates transformed human cells by fusion. Journal of Cell Biology, 171(3):493–??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/171/3/
493.

Durcan:2008:TRC

[DHR^{+}08] Thomas M. Durcan, Elizabeth S. Halpin, Trisha Rao, Nicholas S. Collins, Emily K. Tribble, Jessica E. Hornick, and Edward H. Hinchcliffe. Tektin 2 is required for central spindle microtubule organization and the completion of cytokinesis. Journal of Cell Biology, 181(4):595–??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/181/4/595.

Delevoye:2009:AKC

[DHT^{+}09] Cédric Delevoye, Ilse Hurbain, Danièle Tenza, Jean-Baptiste Sibarita, Stéphanie Uzan-Gafsou, Hiroshi Ohno, Willie J. C. Geerts, Arie J. Verkleij, Jean Salamero, Michael S. Marks, and Graça Raposo. AP-1 and KIF13A coordinate endosomal sorting and positioning during melanosome biogenesis. Journal of Cell Biology, 187(2):247–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/2/247.

DiFiore:2008:PBS

[Di 08] Pier Paolo Di Fiore. Playing both sides: nucleophosmin between tumor suppression and oncogenesis. Journal of Cell Biology, 182(1):7–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/1/7.

Perez:2009:BMP

[dJPAW^{+}09] Vinicio A. de Jesus Perez, Tero-Pekka Alastalo, Jenny C. Wu, Jeffrey D. Axelrod, John P. Cooke, Manuel Amieva, and Marlene Rabinovitch. Bone morphogenetic protein 2 induces pulmonary angiogenesis via Wnt-β-catenin and Wnt–RhoA–Rac1 pathways. Journal of Cell Biology, 184(1):83–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/184/1/83.

Dimmer:2005:MMI

[DJV^{+}05] Kai Stefan Dimmer, Stefan Jakobs, Frank Vogel, Katrin Altmann, and Benedikt Westermann. Mdm31 and Mdm32 are inner membrane proteins required for maintenance of mitochondrial shape and stability of mitochondrial DNA nucleoids in yeast. Journal of Cell Biology, 168(1):103–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/168/1/103.

Darzacq:2006:SRA

[DKR^{+}06] Xavier Darzacq, Nupur Kittur, Sujayita Roy, Yaron Shav-Tal, Robert H. Singer, and U. Thomas Meier. Stepwise RNP assembly at the site of H/ACA RNA transcription in human cells. Journal of Cell Biology, 173(2):207–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/173/2/207.

Dansereau:2008:RRC

[DL08] David A. Dansereau and Paul Lasko. RanBPM regulates cell shape, arrangement, and capacity of the female germline stem cell niche in Drosophila melanogaster. Journal of Cell Biology, 182(5):963–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/5/963.

Ding:2005:GFI

[DLBSD^{+}05] Kan Ding, Martha Lopez-Burks, José Antonio Sánchez-Duran, Murray Korc, and Arthur D. Lander. Growth factor–induced shedding of syndecan-1 confers glypican-1 dependence on mitogenic responses of cancer cells. Journal of Cell Biology, 171(4): 729–??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/4/729.

Dinant:2009:AMC

[DLH^{+}09] Christoffel Dinant, Martijn S. Luijsterburg, Thomas Höfer, Gesa von Bornstaedt, Wim Vermeulen, Adriaan B. Houtsmuller, and Roel van Driel. Assembly of multiprotein complexes that control genome function. Journal of Cell Biology, 185(1):21–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
185/1/21.

Dawson:2009:MBP

[DLHW09] T. Renee Dawson, Michelle D. Lazarus, Martin W. Hetzer, and Susan R. Wente. ER membrane–bending proteins are necessary for de novo nuclear pore formation. Journal of Cell Biology, 184(5):659–??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/5/659.

DAndrea:2008:CPP

[DLL^{+}08] Daniela D’Andrea, Giovanna L. Liguori, J. Ann Le Good, Enza Lonardo, Olov Andersson, Daniel B. Constam, Maria G. Persico, and Gabriella Minchiotti. Cripto promotes A–P axis specification independently of its stimulatory effect on Nodal autoinduction. Journal of Cell Biology, 180(3):597–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/180/3/

Deasy:2007:RCS

[DLT^{+}07] Bridget M. Deasy, Aiping Lu, Jessica C. Tebbets, Joseph M. Feduska, Rebecca C. Schugar, Jonathan B. Pollett, Bin Sun, Kenneth L. Urish, Burhan M. Gharaibeh, Baohong Cao, Robert T. Rubin, and Johnny Huard. A role for cell sex in stem cellmediated skeletal muscle regeneration: female cells have higher muscle regeneration efficiency. Journal of Cell Biology, 177 (1):73–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/1/73.

DAlcontres:2007:LTA

[DMBF^{+}07] Martina Stagno D’Alcontres, Aaron Mendez-Bermudez, Jennifer L. Foxon, Nicola J. Royle, and Paolo Salomoni. Lack of TRF2 in ALT cells causes PML-dependent p53 activation and loss of telomeric DNA. Journal of Cell Biology, 179(5): 855–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/5/855.

Dubreuil:2007:MPC

[DMC^{+}07] Véronique Dubreuil, Anne-Marie Marzesco, Denis Corbeil, Wieland B. Huttner, and Michaela Wilsch-Bräuninger. Midbody and primary cilium of neural progenitors release extracellular membrane particles enriched in the stem cell marker prominin-1. Journal of Cell Biology, 176(4):483–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/176/4/
483.

Dammermann:2008:SRD

[DMDO08] Alexander Dammermann, Paul S. Maddox, Arshad Desai, and Karen Oegema. SAS-4 is recruited to a dynamic structure in newly forming centrioles that is stabilized by the γ-tubulin-mediated addition of centriolar microtubules. Journal of Cell Biology, 180(4):771–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/180/4/771.

DeRegge:2006:HGD

[DNG^{+}06] Nick De Regge, Hans J. Nauwynck, Kristin Geenen, Claude Krummenacher, Gary H. Cohen, Roselyn J. Eisenberg, Thomas C. Mettenleiter, and Herman W. Favoreel. αherpesvirus glycoprotein D interaction with sensory neurons triggers formation of varicosities that serve as virus exit sites. Journal of Cell Biology, 174(2):267–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/174/2/267.

Dermaut:2005:ALC

[DNK^{+}05] Bart Dermaut, Koenraad K. Norga, Artur Kania, Patrik Verstreken, Hongling Pan, Yi Zhou, Patrick Callaerts, and Hugo J. Bellen. Aberrant lysosomal carbohydrate storage accompanies endocytic defects and neurodegeneration in Drosophila benchwarmer. Journal of Cell Biology, 170(1):127–??, July 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/170/1/127.

Dundr:2007:ADI

[DOS^{+}07] Miroslav Dundr, Jason K. Ospina, Myong-Hee Sung, Sam John, Madhvi Upender, Thomas Ried, Gordon L. Hager, and A. Gregory Matera. Actin-dependent intranuclear repositioning of an active gene locus in vivo. Journal of Cell Biology, 179(6): 1095–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/6/1095.

Denecker:2008:CRS

[DOVD08] Geertrui Denecker, Petra Ovaere, Peter Vandenabeele, and Wim Declercq. Caspase-14 reveals its secrets. Journal of Cell Biology, 180(3):451–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/3/451.

Decker:2006:CTH

[DP06] Carolyn J. Decker and Roy Parker. CAR-1 and Trailer hitch: driving mRNP granule function at the ER? Journal of Cell Biology, 173(2):159–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/2/159.

DiFiore:2007:ENC

[DP07] Barbara Di Fiore and Jonathon Pines. Emi1 is needed to couple DNA replication with mitosis but does not regulate activation of the mitotic APC/C. Journal of Cell Biology, 177(3):425–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
177/3/425.

Duncan:2008:RPD

[DP08] Mara C. Duncan and Mark Peifer. Regulating polarity by directing traffic: Cdc42 prevents adherens junctions from Crumblin’ aPart. Journal of Cell Biology, 183(6):971–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/183/6/
971.

Daniels:2009:AEP

[DPD^{+}09] Brian R. Daniels, Edward M. Perkins, Terrence M. Dobrowsky, Sean X. Sun, and Denis Wirtz. Asymmetric enrichment of PIE-1 in the Caenorhabditis elegans zygote mediated by binary counterdiffusion. Journal of Cell Biology, 184(4):473–??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/184/4/
473.

Damm:2005:CCI

[DPK^{+}05] Eva-Maria Damm, Lucas Pelkmans, Jürgen Kartenbeck, Anna Mezzacasa, Teymuras Kurzchalia, and Ari Helenius. Clathrinand caveolin-1–independent endocytosis. Journal of Cell Biology, 168(3):477–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/3/477.

Daniel:2008:MAS

[DPL^{+}08] Jeremy A. Daniel, Manuela Pellegrini, Ji-Hoon Lee, Tanya T. Paull, Lionel Feigenbaum, and André Nussenzweig. Multiple autophosphorylation sites are dispensable for murine ATM activation in vivo. Journal of Cell Biology, 183(5):777–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/183/5/
777.

Deneka:2007:MHA

[DPMB^{+}07] Magdalena Deneka, Annegret Pelchen-Matthews, Rahel Byland, Ezequiel Ruiz-Mateos, and Mark Marsh. In macrophages, HIV-1 assembles into an intracellular plasma membrane domain containing the tetraspanins CD81, CD9, and CD53. Journal of Cell Biology, 177(2):329–??, April 2007. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/177/2/329.

diPenta:2009:DLC

[dPMF^{+}09] Alessandra di Penta, Valentina Mercaldo, Fulvio Florenzano, Sebastian Munck, M. Teresa Ciotti, Francesca Zalfa, Delio Mercanti, Marco Molinari, Claudia Bagni, and Tilmann Achsel. Dendritic LSm1/CBP80-mRNPs mark the early steps of transport commitment and translational control. Journal of Cell Biology, 184(3):423–??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/3/423.

Dumont:2007:CRI

[DPP^{+}07] Julien Dumont, Sebastian Petri, Franz Pellegrin, Marie-Emilie Terret, Markus T. Bohnsack, Pascale Rassinier, Virginie Georget, Petr Kalab, Oliver J. Gruss, and Marie-Hélène Verlhac. A centriole- and RanGTP-independent spindle assembly pathway in meiosis I of vertebrate oocytes. Journal of Cell Biology, 176(3):295–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/3/295.

DiCola:2005:LST

[DR05] Alessandra Di Cola and Colin Robinson. Large-scale translocation reversal within the thylakoid Tat system in vivo. Journal of Cell Biology, 171(2):281–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/171/2/281.

Dikeakos:2007:SPD

[DR07] Jimmy D. Dikeakos and Timothy L. Reudelhuber. Sending proteins to dense core secretory granules: still a lot to sort out. Journal of Cell Biology, 177(2):191–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/177/2/191.

Defilippi:2005:ICR

[DRD^{+}05] Paola Defilippi, Arturo Rosso, Patrizia Dentelli, Cristina Calvi, Giovanni Garbarino, Guido Tarone, Luigi Pegoraro, and Maria Felice Brizzi. β1 integrin and IL-3R coordinately regulate STAT5 activation and anchorage-dependent proliferation. Journal of Cell Biology, 168(7):1099–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/168/7/1099.

deRooij:2005:IDA

[dRKD^{+}05] Johan de Rooij, Andre Kerstens, Gaudenz Danuser, Martin A. Schwartz, and Clare M. Waterman-Storer. Integrindependent actomyosin contraction regulates epithelial cell scattering. Journal of Cell Biology, 171(1):153–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/171/1/153.

Dymek:2007:CCR

[DS07] Erin E. Dymek and Elizabeth F. Smith. A conserved CaM- and radial spoke–associated complex mediates regulation of flagellar dynein activity. Journal of Cell Biology, 179(3):515–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/3/
515.

Dumesic:2009:EMK

[DSBK09] Phillip A. Dumesic, Florence A. Scholl, Deborah I. Barragan, and Paul A. Khavari. Erk1/2 MAP kinases are required for epidermal G2/M progression. Journal of Cell Biology, 185(3):409– ??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
185/3/409.

Dunn:2007:MMM

[DSH^{+}07] Brian D. Dunn, Takeshi Sakamoto, Myoung-Soon S. Hong, James R. Sellers, and Peter A. Takizawa. Myo4p is a monomeric myosin with motility uniquely adapted to transport mRNA. Journal of Cell Biology, 178(7):1193–??, September 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/178/7/1193.

Ding:2006:MCM

[DSK^{+}06] Da-Qiao Ding, Nobuko Sakurai, Yuki Katou, Takehiko Itoh, Katsuhiko Shirahige, Tokuko Haraguchi, and Yasushi Hiraoka. Meiotic cohesins modulate chromosome compaction during meiotic prophase in fission yeast. Journal of Cell Biology, 174 (4):499–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/4/499.

DeVries-Seimon:2005:CIM

[DSLY^{+}05] Tracie DeVries-Seimon, Yankun Li, Pin Mei Yao, Elizabeth Stone, Yibin Wang, Roger J. Davis, Richard Flavell, and Ira Tabas. Cholesterol-induced macrophage apoptosis requires ER stress pathways and engagement of the type A scavenger receptor. Journal of Cell Biology, 171(1):61–??, October 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/171/1/61.

Dikovskaya:2007:CP

[DSN^{+}07a] Dina Dikovskaya, David Schiffmann, Ian P. Newton, Abigail Oakley, Karin Kroboth, Owen Sansom, Thomas J. Jamieson, Valerie Meniel, Alan Clarke, and Inke S. Näthke. Correction 176, p. 183. Journal of Cell Biology, 176(3):369–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/176/3/
369.

Dikovskaya:2007:LAI

[DSN^{+}07b] Dina Dikovskaya, David Schiffmann, Ian P. Newton, Abigail Oakley, Karin Kroboth, Owen Sansom, Thomas J. Jamieson, Valerie Meniel, Alan Clarke, and Inke S. Näthke. Loss of APC induces polyploidy as a result of a combination of defects in mitosis and apoptosis. Journal of Cell Biology, 176(2):183–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
176/2/183.

deCathelineau:2009:ASC

[dSW09] Aimee deCathelineau, Ben Short, and Ruth Williams. The American Society for Cell Biology, 48th Annual Meeting San Francisco, CA, December 13–17, 2008. Journal of Cell Biology, 184(2):190–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/2/190.

Dong:2007:MBN

[DTL^{+}07] Min Dong, William H. Tepp, Huisheng Liu, Eric A. Johnson, and Edwin R. Chapman. Mechanism of botulinum neurotoxin B and G entry into hippocampal neurons. Journal of Cell Biology, 179(7):1511–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/7/1511.

Decker:2007:EGA

[DTP07] Carolyn J. Decker, Daniela Teixeira, and Roy Parker. Edc3p and a glutamine/asparagine-rich domain of Lsm4p function in processing body assembly in Saccharomyces cerevisiae. Journal of Cell Biology, 179(3):437–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/3/437.

Dilworth:2005:MNN

[DTR^{+}05] David J. Dilworth, Alan J. Tackett, Richard S. Rogers, Eugene C. Yi, Rowan H. Christmas, Jennifer J. Smith, Andrew F. Siegel, Brian T. Chait, Richard W. Wozniak, and John D. Aitchison. The mobile nucleoporin Nup2p and chromatinbound Prp20p function in endogenous NPC-mediated transcriptional control. Journal of Cell Biology, 171(6):955–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
171/6/955.

Duex:2006:VFC

[DTW06] Jason E. Duex, Fusheng Tang, and Lois S. Weisman. The Vac14p–Fig4p complex acts independently of Vac7p and couples PI3,5P_{2} synthesis and turnover. Journal of Cell Biology, 172(5):693–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/5/693.

DAlterio:2005:C

[DTY^{+}05a] Cecilia D’Alterio, Dao D. D. Tran, Maggie W. Y. Au Yeung, Michael S. H. Hwang, Michelle A. Li, Claudia J. Arana, Vikram K. Mulligan, Mary Kubesh, Praveer Sharma, Maretta Chase, Ulrich Tepass, and Dorothea Godt. Correction. Journal of Cell Biology, 171(6):1085–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/171/6/1085.

DAlterio:2005:DMC

[DTY^{+}05b] Cecilia D’Alterio, Dao D. D. Tran, Maggie W. Y. Au Yeung, Michael S. H. Hwang, Michelle A. Li, Claudia J. Arana, Vikram K. Mulligan, Mary Kubesh, Praveer Sharma, Maretta Chase, Ulrich Tepass, and Dorothea Godt. Drosophila melanogaster Cad99C, the orthologue of human Usher cadherin PCDH15, regulates the length of microvilli. Journal of Cell Biology, 171(3):549–??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/3/549.

Dumont:2005:PIC

[DUR^{+}05] Julien Dumont, Muriel Umbhauer, Pascale Rassinier, André Hanauer, and Marie-Hélène Verlhac. p90rsk is not involved in cytostatic factor arrest in mouse oocytes. Journal of Cell Biology, 169(2):227–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/2/227.

Dragestein:2008:DBG

[DvCvH^{+}08] Katharina A. Dragestein, Wiggert A. van Cappellen, Jeffrey van Haren, George D. Tsibidis, Anna Akhmanova, Tobias A. Knoch, Frank Grosveld, and Niels Galjart. Dynamic behavior of GFP–CLIP–170 reveals fast protein turnover on microtubule plus ends. Journal of Cell Biology, 180(4):729–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/180/4/
729.

Dorner:2006:LAP

[DVF^{+}06] Daniela Dorner, Sylvia Vlcek, Nicole Foeger, Andreas Gajewski, Christian Makolm, Josef Gotzmann, Christopher J. Hutchison, and Roland Foisner. Lamina-associated polypeptide 2α regulates cell cycle progression and differentiation via the retinoblastoma–E2F pathway. Journal of Cell Biology, 173 (1):83–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/1/83.

Danen:2005:C

[DvRF^{+}05a] Erik H. J. Danen, Jacco van Rheenen, Willeke Franken, Stephan Huveneers, Petra Sonneveld, Kees Jalink, and Arnoud Sonnenberg. Correction. Journal of Cell Biology, 170(3):497–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
170/3/497.

Danen:2005:ICM

[DvRF^{+}05b] Erik H. J. Danen, Jacco van Rheenen, Willeke Franken, Stephan Huveneers, Petra Sonneveld, Kees Jalink, and Arnoud Sonnenberg. Integrins control motile strategy through a Rho– cofilin pathway. Journal of Cell Biology, 169(3):515–??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/169/3/
515.

Doyle:2009:ODT

[DWMY09] Andrew D. Doyle, Francis W. Wang, Kazue Matsumoto, and Kenneth M. Yamada. One-dimensional topography underlies three-dimensional fibrillar cell migration. Journal of Cell Biology, 184(4):481–??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/4/481.

Ding:2008:CEC

[DXC^{+}08] Yu Ding, Ying Xi, Ting Chen, Ji yong Wang, Dong lei Tao, Zhi-Li Wu, Yi ping Li, Chen Li, Rong Zeng, and Lin Li. Caprin-2 enhances canonical Wnt signaling through regulating LRP5/6 phosphorylation. Journal of Cell Biology, 182(5): 865–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/5/865.

Deak:2009:MBN

[DXC^{+}09] Ferenc Deák, Yi Xu, Wen-Pin Chang, Irina Dulubova, Mikhail Khvotchev, Xinran Liu, Thomas C. Südhof, and Josep Rizo. Munc18-1 binding to the neuronal SNARE complex controls synaptic vesicle priming. Journal of Cell Biology, 184(5): 751–??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/184/5/751.

Dzhashiashvili:2007:NRA

[DZG^{+}07] Yulia Dzhashiashvili, Yanqing Zhang, Jolanta Galinska, Isabel Lam, Martin Grumet, and James L. Salzer. Nodes of Ranvier and axon initial segments are ankyrin G-dependent domains that assemble by distinct mechanisms. Journal of Cell Biology, 177(5):857–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/5/857.

Dultz:2008:SKA

[DZW^{+}08] Elisa Dultz, Esther Zanin, Claudia Wurzenberger, Marion Braun, Gwénaël Rabut, Lucia Sironi, and Jan Ellenberg. Systematic kinetic analysis of mitotic dis- and reassembly of the nuclear pore in living cells. Journal of Cell Biology, 180(5):857–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
180/5/857.

Ezratty:2009:CMI

[EBMG09] Ellen J. Ezratty, Claire Bertaux, Eugene E. Marcantonio, and Gregg G. Gundersen. Clathrin mediates integrin endocytosis for focal adhesion disassembly in migrating cells. Journal of Cell Biology, 187(5):733–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/187/5/733.

Edwards:2009:IDC

[ECR^{+}09] Stacey L. Edwards, Nicole K. Charlie, Janet E. Richmond, Jan Hegermann, Stefan Eimer, and Kenneth G. Miller. Impaired dense core vesicle maturation in Caenorhabditis elegans mutants lacking Rab2. Journal of Cell Biology, 186(6): 881–??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/186/6/881.

Eshed:2007:SGP

[EFCP07] Yael Eshed, Konstantin Feinberg, David J. Carey, and Elior Peles. Secreted gliomedin is a perinodal matrix component of peripheral nerves. Journal of Cell Biology, 177(3):551–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/177/3/
551.

Elia:2005:EGG

[EFG^{+}05] Natalie Elia, Shahar Frechter, Yinon Gedi, Baruch Minke, and Zvi Selinger. Excess of Gβ_{e} over Gqα_{e} in vivo prevents dark, spontaneous activity of Drosophila photoreceptors. Journal of Cell Biology, 171(3):517–??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/171/3/517.

Epp:2007:LDD

[EFM^{+}07] Nikolas Epp, Gerhard Fürstenberger, Karsten Müller, Silvia de Juanes, Michael Leitges, Ingrid Hausser, Florian Thieme, Gerhard Liebisch, Gerd Schmitz, and Peter Krieg. 12rlipoxygenase deficiency disrupts epidermal barrier function. Journal of Cell Biology, 177(1):173–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/177/1/173.

Emr:2009:JTG

[EGL^{+}09] Scott Emr, Benjamin S. Glick, Adam D. Linstedt, Jennifer Lippincott-Schwartz, Alberto Luini, Vivek Malhotra, Brad J. Marsh, Akihiko Nakano, Suzanne R. Pfeffer, Catherine Rabouille, James E. Rothman, Graham Warren, and Felix T. Wieland. Journeys through the Golgi — taking stock in a new era. Journal of Cell Biology, 187(4):449–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/187/4/449.

Evans:2005:LAD

[EGR^{+}05] Katia J. Evans, Edgar R. Gomes, Steven M. Reisenweber, Gregg G. Gundersen, and Brett P. Lauring. Linking axonal degeneration to microtubule remodeling by Spastin-mediated microtubule severing. Journal of Cell Biology, 168(4):599– ??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/4/599.

Enserink:2009:CCP

[EHHK09] Jorrit M. Enserink, Hans Hombauer, Meng-Er Huang, and Richard D. Kolodner. Cdc28/ cdk1 positively and negatively affects genome stability in S. cerevisiae. Journal of Cell Biology, 185(3):423–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/3/423.

Ehlers:2005:DD

[Ehl05] Michael D. Ehlers. Dendrite development. Journal of Cell Biology, 170(4):517–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/4/517.

Escobar-Henriques:2006:RMF

[EHWL06] Mafalda Escobar-Henriques, Benedikt Westermann, and Thomas Langer. Regulation of mitochondrial fusion by the Fbox protein Mdm30 involves proteasome-independent turnover of Fzo1. Journal of Cell Biology, 173(5):645–??, June 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/173/5/645.

ElChami:2005:ADA

[EIK^{+}05] Nisrine El Chami, Fouziha Ikhlef, Krisztian Kaszas, Sadok Yakoub, Eric Tabone, Benazir Siddeek, Stéphanie Cunha, Claude Beaudoin, Laurent Morel, Mohamed Benahmed, and Daniel C. Régnier. Androgen-dependent apoptosis in male germ cells is regulated through the proto-oncoprotein Cbl. Journal of Cell Biology, 171(4):651–??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/171/4/651.

Ellis:2008:SMS

[ELD^{+}08] Jonathan D. Ellis, David Llères, Marco Denegri, Angus I. Lamond, and Javier F. Cáceres. Spatial mapping of splicing factor complexes involved in exon and intron definition. Journal of Cell Biology, 181(6):921–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/181/6/921.

Emanuele:2008:ABK

[ELJ^{+}08] Michael J. Emanuele, Weijie Lan, Miri Jwa, Stephanie A. Miller, Clarence S. M. Chan, and P. Todd Stukenberg. Aurora B kinase and protein phosphatase 1 have opposing roles in modulating kinetochore assembly. Journal of Cell Biology, 181 (2):241–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/2/241.

Engel:2009:ITP

[ELM09] Benjamin D. Engel, William B. Ludington, and Wallace F. Marshall. Intraflagellar transport particle size scales inversely with flagellar length: revisiting the balance-point length control model. Journal of Cell Biology, 187(1):81–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/187/1/81.

Erhardt:2008:GWA

[EMB^{+}08] Sylvia Erhardt, Barbara G. Mellone, Craig M. Betts, Weiguo Zhang, Gary H. Karpen, and Aaron F. Straight. Genomewide analysis reveals a cell cycle-dependent mechanism controlling centromere propagation. Journal of Cell Biology, 183(5): 805–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/5/805.

Espenel:2008:SMA

[EMD^{+}08] Cedric Espenel, Emmanuel Margeat, Patrice Dosset, Cécile Arduise, Christian Le Grimellec, Catherine A. Royer, Claude Boucheix, Eric Rubinstein, and Pierre-Emmanuel Milhiet. Single-molecule analysis of CD9 dynamics and partitioning reveals multiple modes of interaction in the tetraspanin web. Journal of Cell Biology, 182(4):765–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/182/4/765.

Etienne-Manneville:2005:CPP

[EMMN^{+}05] Sandrine Etienne-Manneville, Jean-Baptiste Manneville, Sarah Nicholls, Michael A. Ferenczi, and Alan Hall. Cdc42 and Par6-PKCζ regulate the spatially localized association of Dlg1 and APC to control cell polarization. Journal of Cell Biology, 170(6):895–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/6/895.

Erez:2007:FMB

[EMPK^{+}07] Hadas Erez, Guy Malkinson, Masha Prager-Khoutorsky, Chris I. De Zeeuw, Casper C. Hoogenraad, and Micha E. Spira. Formation of microtubule-based traps controls the sorting and concentration of vesicles to restricted sites of regenerating neurons after axotomy. Journal of Cell Biology, 176(4): 497–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/4/497.

Ehrenreiter:2005:RRR

[EPV^{+}05] Karin Ehrenreiter, Daniela Piazzolla, Vanishree Velamoor, Izabela Sobczak, J. Victor Small, Junji Takeda, Thomas Leung, and Manuela Baccarini. Raf-1 regulates Rho signaling and cell migration. Journal of Cell Biology, 168(6):955–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/168/6/955.

Ehses:2009:ROP

[ERM^{+}09a] Sarah Ehses, Ines Raschke, Giuseppe Mancuso, Andrea Bernacchia, Stefan Geimer, Daniel Tondera, Jean-Claude Martinou, Benedikt Westermann, Elena I. Rugarli, and Thomas Langer. Regulation of OPA1 processing and mitochondrial fusion by m -AAA protease isoenzymes and OMA1. Journal of Cell Biology, 187(7):1023–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/7/1023.

Elfenbein:2009:SRA

[ERM^{+}09b] Arye Elfenbein, John M. Rhodes, Julia Meller, Martin A. Schwartz, Michiyuki Matsuda, and Michael Simons. Suppression of RhoG activity is mediated by a syndecan 4–synectin– RhoGDI1 complex and is reversed by PKCα in a Rac1 activation pathway. Journal of Cell Biology, 186(1):75–??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/186/1/75.

Etard:2008:SCU

[ERS08] Christelle Etard, Urmas Roostalu, and Uwe Strähle. Shuttling of the chaperones Unc45b and Hsp90a between the A band and the Z line of the myofibril. Journal of Cell Biology, 180 (6):1163–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/6/1163.

Evans:2006:C

[ESG^{+}06a] James E. Evans, Joshua J. Snow, Amy L. Gunnarson, Guangshuo Ou, Henning Stahlberg, Kent L. McDonald, and Jonathan M. Scholey. Correction. Journal of Cell Biology, 172 (6):949–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/6/949.

Evans:2006:CP

[ESG^{+}06b] James E. Evans, Joshua J. Snow, Amy L. Gunnarson, Guangshuo Ou, Henning Stahlberg, Kent L. McDonald, and Jonathan M. Scholey. Correction 172, p. 663. Journal of Cell Biology, 175(5):837–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/175/5/837.

Evans:2006:FMI

[ESG^{+}06c] James E. Evans, Joshua J. Snow, Amy L. Gunnarson, Guangshuo Ou, Henning Stahlberg, Kent L. McDonald, and Jonathan M. Scholey. Functional modulation of IFT kinesins extends the sensory repertoire of ciliated neurons in Caenorhabditis elegans. Journal of Cell Biology, 172(5):663–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/172/5/
663.

Enserink:2006:CPC

[ESZK06] Jorrit M. Enserink, Marcus B. Smolka, Huilin Zhou, and Richard D. Kolodner. Checkpoint proteins control morphogenetic events during DNA replication stress in Saccharomyces cerevisiae. Journal of Cell Biology, 175(5):729–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/175/5/
729.

Espada:2008:NED

[EVF^{+}08] Jesús Espada, Ignacio Varela, Ignacio Flores, Alejandro P. Ugalde, Juan Cadiñanos, Alberto M. Pendás, Colin L. Stewart, Karl Tryggvason, María A. Blasco, José M. P. Freije, and Carlos López-Otín. Nuclear envelope defects cause stem cell dysfunction in premature-aging mice. Journal of Cell Biology, 181(1):27–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/1/27.

Eshghi:2007:IEM

[EVH^{+}07] Shawdee Eshghi, Mariette G. Vogelezang, Richard O. Hynes, Linda G. Griffith, and Harvey F. Lodish. α_{4}β_{1} integrin and erythropoietin mediate temporally distinct steps in erythropoiesis: integrins in red cell development. Journal of Cell Biology, 177 (5):871–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/5/871.

Eshghi:2008:IEM

[EVH^{+}08] Shawdee Eshghi, Mariette G. Vogelezang, Richard O. Hynes, Linda G. Griffith, and Harvey F. Lodish. α4β1 integrin and erythropoietin mediate temporally distinct steps in erythropoiesis: integrins in red cell development. Journal of Cell Biology, 181 (2):395–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/2/395.

Engel:2008:MLD

[EW08] Alex Engel and Peter Walter. Membrane lysis during biological membrane fusion: collateral damage by misregulated fusion machines. Journal of Cell Biology, 183(2):181–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/183/2/181.

Esteban:2006:KDT

[EYB^{+}06] Pedro F. Esteban, Hye-Young Yoon, Jodi Becker, Susan G. Dorsey, Paola Caprari, Mary Ellen Palko, Vincenzo Coppola, H. Uri Saragovi, Paul A. Randazzo, and Lino Tessarollo. A kinase-deficient TrkC receptor isoform activates Arf6–Rac1 signaling through the scaffold protein tamalin. Journal of Cell Biology, 173(2):291–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/2/291.

Eastman:2009:RUL

[EYB09] Scott W. Eastman, Mina Yassaee, and Paul D. Bieniasz. A role for ubiquitin ligases and Spartin/SPG20 in lipid droplet turnover. Journal of Cell Biology, 184(6):881–??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/184/6/881.

Filippi:2008:NRL

[FA08] Beatrice Maria Filippi and Dario R. Alessi. Novel role for the LKB1 pathway in controlling monocarboxylate fuel transporters. Journal of Cell Biology, 183(1):7–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/183/1/7.

Fuchs:2007:NCD

[FAG^{+}07] Tobias A. Fuchs, Ulrike Abed, Christian Goosmann, Robert Hurwitz, Ilka Schulze, Volker Wahn, Yvette Weinrauch, Volker Brinkmann, and Arturo Zychlinsky. Novel cell death program leads to neutrophil extracellular traps. Journal of Cell Biology, 176(2):231–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/2/231.

Ferraiuolo:2005:C

[FBD^{+}05a] Maria A. Ferraiuolo, Sanjukta Basak, Josee Dostie, Elizabeth L. Murray, Daniel R. Schoenberg, and Nahum Sonenberg. Correction. Journal of Cell Biology, 171(1):175–??, October 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/171/1/175.

Ferraiuolo:2005:REB

[FBD^{+}05b] Maria A. Ferraiuolo, Sanjukta Basak, Josee Dostie, Elizabeth L. Murray, Daniel R. Schoenberg, and Nahum Sonenberg. A role for the eIF4E-binding protein 4E–T in P-body formation and mRNA decay. Journal of Cell Biology, 170(6):913–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/170/6/
913.

Frantz:2008:CPS

[FBD^{+}08] Christian Frantz, Gabriela Barreiro, Laura Dominguez, Xiaoming Chen, Robert Eddy, John Condeelis, Mark J. S. Kelly, Matthew P. Jacobson, and Diane L. Barber. Cofilin is a pH sensor for actin free barbed end formation: role of phosphoinositide binding. Journal of Cell Biology, 183(5):865–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/183/5/
865.

Faustrup:2009:UCS

[FBJB^{+}09] Helene Faustrup, Simon Bekker-Jensen, Jiri Bartek, Jiri Lukas, and Niels Mailand. USP7 counteracts SCF^{βTrCP}— but not APC^{Cdh1}-mediated proteolysis of Claspin. Journal of Cell Biology, 184(1):13–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/1/13.

Fang:2007:VST

[FBO^{+}07] Jun Fang, Joseph A. Brzostowski, Stephen Ou, Nilgun Isik, Vinod Nair, and Tian Jin. A vesicle surface tyrosine kinase regulates phagosome maturation. Journal of Cell Biology, 178 (3):411–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/3/411.

Fornaro:2006:SAS

[FBY^{+}06] Mara Fornaro, Peter M. Burch, Wentian Yang, Lei Zhang, Claire E. Hamilton, Jung H. Kim, Benjamin G. Neel, and Anton M. Bennett. SHP-2 activates signaling of the nuclear factor of activated T cells to promote skeletal muscle growth. Journal of Cell Biology, 175(1):87–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/175/1/87.

Francavilla:2009:BNF

[FCB^{+}09] Chiara Francavilla, Paola Cattaneo, Vladimir Berezin, Elisabeth Bock, Diletta Ami, Ario de Marco, Gerhard Christofori, and Ugo Cavallaro. The binding of NCAM to FGFR1 induces a specific cellular response mediated by receptor trafficking. Journal of Cell Biology, 187(7):1101–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/187/7/1101. See correc-

tion [FCB^{+}19].

Francavilla:2019:CBN

[FCB^{+}19] Chiara Francavilla, Paola Cattaneo, Vladimir Berezin, Elisabeth Bock, Diletta Ami, Ario de Marco, Gerhard Christofori, and Ugo Cavallaro. Correction: The binding of NCAM to FGFR1 induces a specific cellular response mediated by receptor trafficking. Journal of Cell Biology, 218(4):1422–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/4/

1422. See [FCB^{+}09].

Flower:2007:YYT

[FCDM^{+}07] Todd R. Flower, Cheryl Clark-Dixon, Cheynita Metoyer, Hui Yang, Runhua Shi, Zhaojie Zhang, and Stephan N. Witt. YGR198w (YPP1) targets A30P α-synuclein to the vacuole for degradation. Journal of Cell Biology, 177(6):1091–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/177/6/
1091.

Foster:2009:MHK

[FD09] Elinor R. Foster and Jessica A. Downs. Methylation of H3 K4 and K79 is not strictly dependent on H2B K123 ubiquitylation. Journal of Cell Biology, 184(5):631–??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/184/5/631.

Fejtova:2009:DLC

[FDB^{+}09] Anna Fejtova, Daria Davydova, Ferdinand Bischof, Vesna Lazarevic, Wilko D. Altrock, Stefano Romorini, Cornelia Schöne, Werner Zuschratter, Michael R. Kreutz, Craig C. Garner, Noam E. Ziv, and Eckart D. Gundelfinger. Dynein light chain regulates axonal trafficking and synaptic levels of Bassoon. Journal of Cell Biology, 185(2):341–??, April 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/185/2/341.

Fielding:2008:ILK

[FDM^{+}08] Andrew B. Fielding, Iveta Dobreva, Paul C. McDonald, Leonard J. Foster, and Shoukat Dedhar. Integrin-linked kinase localizes to the centrosome and regulates mitotic spindle organization. Journal of Cell Biology, 180(4):681–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/180/4/
681.

Fraschini:2006:DBY

[FDV^{+}06] Roberta Fraschini, Claudio D’Ambrosio, Marianna Venturetti, Giovanna Lucchini, and Simonetta Piatti. Disappearance of the budding yeast Bub2–Bfa1 complex from the mother-bound spindle pole contributes to mitotic exit. Journal of Cell Biology, 172(3):335–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/3/335.

Fok:2006:EBV

[FFS06] Victor Fok, Kyle Friend, and Joan A. Steitz. Epstein-Barr virus noncoding RNAs are confined to the nucleus, whereas their partner, the human La protein, undergoes nucleocytoplasmic shuttling. Journal of Cell Biology, 173(3):319–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/173/3/319.

Fagarasanu:2005:IPM

[FFT^{+}05] Monica Fagarasanu, Andrei Fagarasanu, Yuen Yi C. Tam, John D. Aitchison, and Richard A. Rachubinski. Inp1p is a peroxisomal membrane protein required for peroxisome inheritance in Saccharomyces cerevisiae. Journal of Cell Biology, 169(5):765–??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/5/765.

Fan:2007:NCI

[FFW^{+}07] Shuling Fan, Vanessa Fogg, Qian Wang, Xiao-Wei Chen, Chia-Jen Liu, and Ben Margolis. A novel Crumbs3 isoform regulates cell division and ciliogenesis via importin β interactions. Journal of Cell Biology, 178(3):387–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/178/3/387.

Fackler:2008:CMT

[FG08] Oliver T. Fackler and Robert Grosse. Cell motility through plasma membrane blebbing. Journal of Cell Biology, 181 (6):879–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/6/879.

Fernandez-Hernando:2006:IGL

[FHFB^{+}06] Carlos Fernández-Hernando, Masaki Fukata, Pascal N. Bernatchez, Yuko Fukata, Michelle I. Lin, David S. Bredt, and William C. Sessa. Identification of Golgi-localized acyl transferases that palmitoylate and regulate endothelial nitric oxide synthase. Journal of Cell Biology, 174(3):369–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/174/3/369.

Farr:2009:MPF

[FHMC09] Glen A. Farr, Michael Hull, Ira Mellman, and Michael J. Caplan. Membrane proteins follow multiple pathways to the basolateral cell surface in polarized epithelial cells. Journal of Cell Biology, 186(2):269–??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/186/2/269.

Fagerholm:2005:SIC

[FHNG05] Susanna C. Fagerholm, Tiina J. Hilden, Susanna M. Nurmi, and Carl G. Gahmberg. Specific integrin α and β chain phosphorylations regulate LFA-1 activation through affinity-dependent and -independent mechanisms. Journal of Cell Biology, 171(4): 705–??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/4/705.

Faul:2005:PIM

[FHO^{+}05] Christian Faul, Stefan Hüttelmaier, Jun Oh, Virginie Hachet, Robert H. Singer, and Peter Mundel. Promotion of importin α-mediated nuclear import by the phosphorylation-dependent binding of cargo protein to 14-3-3. Journal of Cell Biology, 169(3):415–??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/3/415.

Fuchs:2007:SRG

[FHS^{+}07a] Evelyn Fuchs, Alexander K. Haas, Robert A. Spooner, Shin ichiro Yoshimura, J. Michael Lord, and Francis A. Barr. Specific Rab GTPase-activating proteins define the Shiga toxin and epidermal growth factor uptake pathways. Journal of Cell Biology, 177(6):1133–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/6/1133.

Fugmann:2007:RST

[FHS^{+}07b] Tim Fugmann, Angelika Hausser, Patrik Schöffler, Simone Schmid, Klaus Pfizenmaier, and Monilola A. Olayioye. Regulation of secretory transport by protein kinase D-mediated phosphorylation of the ceramide transfer protein. Journal of Cell Biology, 178(1):15–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/1/15.

Fujito:2005:ICM

[FIK^{+}05] Tsutomu Fujito, Wataru Ikeda, Shigeki Kakunaga, Yukiko Minami, Mihoko Kajita, Yasuhisa Sakamoto, Morito Monden, and Yoshimi Takai. Inhibition of cell movement and proliferation by cell–cell contact-induced interaction of Necl-5 with nectin-3. Journal of Cell Biology, 171(1):165–??, October 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/171/1/165. See cor-

rection [FIK^{+}20].

Fujito:2020:CIC

[FIK^{+}20] Tsutomu Fujito, Wataru Ikeda, Shigeki Kakunaga, Yukiko Minami, Mihoko Kajita, Yasuhisa Sakamoto, Morito Monden, and Yoshimi Takai. Correction: Inhibition of cell movement and proliferation by cell–cell contact-induced interaction of Necl-5 with nectin-3. Journal of Cell Biology, 219(6):??, June 1, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/6/
jcb.20050109005122020c/151782/Correction-Inhibition-
of-cell-movement-and. See [FIK^{+}05].

Faulhammer:2005:CGD

[FKB^{+}05] Frank Faulhammer, Gerlinde Konrad, Ben Brankatschk, Sabina Tahirovic, Andreas Knödler, and Peter Mayinger. Cell growthdependent coordination of lipid signaling and glycosylation is mediated by interactions between Sac1p and Dpm1p. Journal of Cell Biology, 168(2):185–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/168/2/185.

Faza:2009:SFC

[FKJ^{+}09] Marius Boulos Faza, Stefan Kemmler, Sonia Jimeno, Cristina González-Aguilera, Andrés Aguilera, Ed Hurt, and Vikram Govind Panse. Sem1 is a functional component of the nuclear pore complex–associated messenger RNA export machinery. Journal of Cell Biology, 184(6):833–??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/184/6/833.

Fukunaga-Kalabis:2006:CCS

[FKML^{+}06] Mizuho Fukunaga-Kalabis, Gabriela Martinez, Zhao-Jun Liu, Jiri Kalabis, Paul Mrass, Wolfgang Weninger, Sue M. Firth, Nathalie Planque, Bernard Perbal, and Meenhard Herlyn. CCN3 controls 3D spatial localization of melanocytes in the human skin through DDR1. Journal of Cell Biology, 175(4): 563–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/4/563.

Frantz:2007:PFB

[FKN^{+}07] Christian Frantz, Anastasios Karydis, Perihan Nalbant, Klaus M. Hahn, and Diane L. Barber. Positive feedback between Cdc42 activity and H^{+} efflux by the Na–H exchanger NHE1 for polarity of migrating cells. Journal of Cell Biology, 179(3):403–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/3/403.

Frank:2006:RMP

[FKS^{+}06] Natasha Y. Frank, Alvin T. Kho, Tobias Schatton, George F. Murphy, Michael J. Molloy, Qian Zhan, Marco F. Ramoni, Markus H. Frank, Isaac S. Kohane, and Emanuela Gussoni. Regulation of myogenic progenitor proliferation in human fetal skeletal muscle by BMP4 and its antagonist Gremlin. Journal of Cell Biology, 175(1):99–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/175/1/99.

Fernandes:2008:IST

[FLA^{+}08] Jacqueline Fernandes, Ivan M. Lorenzo, Yaniré N. Andrade, Anna Garcia-Elias, Selma A. Serra, José M. Fernández-Fernández, and Miguel A. Valverde. IP_{3} sensitizes TRPV4 channel to the mechano- and osmotransducing messenger 5’- 6’-epoxyeicosatrienoic acid. Journal of Cell Biology, 181(1): 143–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/1/143.

Fivaz:2005:RIT

[FM05] Marc Fivaz and Tobias Meyer. Reversible intracellular translocation of KRas but not HRas in hippocampal neurons regulated by Ca^{2+} /calmodulin. Journal of Cell Biology, 170(3):429–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
170/3/429.

Frohlich:2009:GWS

[FMA^{+}09] Florian Fröhlich, Karen Moreira, Pablo S. Aguilar, Nina C. Hubner, Matthias Mann, Peter Walter, and Tobias C. Walther.

A genome-wide screen for genes affecting eisosomes reveals Nce102 function in sphingolipid signaling. Journal of Cell Biology, 185(7):1227–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/7/1227.

Fugger:2009:HFH

[FMD^{+}09] Kasper Fugger, Martin Mistrik, Jannie Rendtlew Danielsen, Christoffel Dinant, Jacob Falck, Jiri Bartek, Jiri Lukas, and Niels Mailand. Human Fbh1 helicase contributes to genome maintenance via pro- and anti-recombinase activities. Journal of Cell Biology, 186(5):655–??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/186/5/655.

Fagarasanu:2009:MDP

[FMK^{+}09] Andrei Fagarasanu, Fred D. Mast, Barbara Knoblach, Yui Jin, Matthew J. Brunner, Michael R. Logan, J. N. Mark Glover, Gary A. Eitzen, John D. Aitchison, Lois S. Weisman, and Richard A. Rachubinski. Myosin-driven peroxisome partitioning in S. cerevisiae. Journal of Cell Biology, 186(4):541–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
186/4/541.

Frescas:2006:SMS

[FML^{+}06] David Frescas, Manos Mavrakis, Holger Lorenz, Robert De-Lotto, and Jennifer Lippincott-Schwartz. The secretory membrane system in the Drosophila syncytial blastoderm embryo exists as functionally compartmentalized units around individual nuclei. Journal of Cell Biology, 173(2):219–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/173/2/219.

Feng:2008:ARD

[FMM^{+}08] Weiyi Feng, N. Patrick McCabe, Ganapati H. Mahabeleshwar, Payaningal R. Somanath, David R. Phillips, and Tatiana V. Byzova. The angiogenic response is dictated by β_{3} integrin on bone marrow–derived cells. Journal of Cell Biology, 183(6): 1145–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/6/1145.

Foote:2006:CAC

[FN06] Christopher Foote and Steven F. Nothwehr. The clathrin adaptor complex 1 directly binds to a sorting signal in Ste13p to reduce the rate of its trafficking to the late endosome of yeast. Journal of Cell Biology, 173(4):615–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/173/4/615.

Freudzon:2005:RMP

[FNH^{+}05] Leon Freudzon, Rachael P. Norris, Arthur R. Hand, Shigeru Tanaka, Yoshinaga Saeki, Teresa L. Z. Jones, Mark M. Rasenick, Catherine H. Berlot, Lisa M. Mehlmann, and Laurinda A. Jaffe. Regulation of meiotic prophase arrest in mouse oocytes by GPR3, a constitutive activator of the G_{s} G protein. Journal of Cell Biology, 171(2):255–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/171/2/255.

Fairn:2009:ESD

[FOB^{+}09] Gregory D. Fairn, Koji Ogata, Roberto J. Botelho, Philip D. Stahl, Richard A. Anderson, Pietro De Camilli, Tobias Meyer, Shoshana Wodak, and Sergio Grinstein. An electrostatic switch displaces phosphatidylinositol phosphate kinases from the membrane during phagocytosis. Journal of Cell Biology, 187 (5):701–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/187/5/701.

Fouquet:2009:MAZ

[FOW^{+}09] Wernher Fouquet, David Owald, Carolin Wichmann, Sara Mertel, Harald Depner, Marcus Dyba, Stefan Hallermann, Robert J. Kittel, Stefan Eimer, and Stephan J. Sigrist. Maturation of active zone assembly by Drosophila Bruchpilot. Journal of Cell Biology, 186(1):129–??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/186/1/129.

Franchitto:2008:RFS

[FPP^{+}08] Annapaola Franchitto, Livia Maria Pirzio, Ennio Prosperi, Orazio Sapora, Margherita Bignami, and Pietro Pichierri. Replication fork stalling in WRN-deficient cells is overcome by prompt activation of a MUS81-dependent pathway. Journal of Cell Biology, 183(2):241–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/183/2/241.

Flevaris:2007:MSC

[FSG^{+}07] Panagiotis Flevaris, Aleksandra Stojanovic, Haixia Gong, Athar Chishti, Emily Welch, and Xiaoping Du. A molecular switch that controls cell spreading and retraction. Journal of Cell Biology, 179(3):553–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/3/553.

Fei:2008:FFH

[FSG^{+}08] Weihua Fei, Guanghou Shui, Bruno Gaeta, Ximing Du, Lars Kuerschner, Peng Li, Andrew J. Brown, Markus R. Wenk, Robert G. Parton, and Hongyuan Yang. Fld1p, a functional homologue of human seipin, regulates the size of lipid droplets in yeast. Journal of Cell Biology, 180(3):473–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/180/3/473.

Fritz:2008:ISH

[FSH08] Richard Fritz, Karin Stiasny, and Franz X. Heinz. Identification of specific histidines as pH sensors in flavivirus membrane fusion. Journal of Cell Biology, 183(2):353–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/183/2/353.

Furman:2007:EVR

[FSK^{+}07] Craig Furman, Alisha L. Sieminski, Adam V. Kwiatkowski, Douglas A. Rubinson, Eliza Vasile, Roderick T. Bronson, Reinhard Fässler, and Frank B. Gertler. Ena/ VASP is required for endothelial barrier function in vivo. Journal of Cell Biology, 179(4):761–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/4/761.

Farjo:2006:RFN

[FSN^{+}06] Rafal Farjo, Jeff S. Skaggs, Barbara A. Nagel, Alexander B. Quiambao, Zack A. Nash, Steven J. Fliesler, and Muna I. Naash. Retention of function without normal disc morphogenesis occurs in cone but not rod photoreceptors. Journal of Cell Biology, 173(1):59–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/1/59.

Forster:2006:PDI

[FSnP^{+}06] Michele L. Forster, Kelsey Sivick, Young nam Park, Peter Arvan, Wayne I. Lencer, and Billy Tsai. Protein disulfide isomerase-like proteins play opposing roles during retrotranslocation. Journal of Cell Biology, 173(6):853–??, June 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/173/6/853.

Fields:2007:VSC

[FSP^{+}07] Ian C. Fields, Elina Shteyn, Marc Pypaert, Véronique Proux-Gillardeaux, Richard S. Kang, Thierry Galli, and Heike Fölsch. v-SNARE cellubrevin is required for basolateral sorting of AP-1B-dependent cargo in polarized epithelial cells. Journal of Cell Biology, 177(3):477–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/3/477.

Filimonenko:2007:FMB

[FSR^{+}07] Maria Filimonenko, Susanne Stuffers, Camilla Raiborg, Ai Yamamoto, Lene Malerød, Elizabeth M. C. Fisher, Adrian Isaacs, Andreas Brech, Harald Stenmark, and Anne Simonsen. Functional multivesicular bodies are required for autophagic clearance of protein aggregates associated with neurodegenerative disease. Journal of Cell Biology, 179(3):485–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/179/3/485.

Flemming:2009:TSD

[FSS^{+}09] Dirk Flemming, Phillip Sarges, Philipp Stelter, Andrea Hellwig, Bettina Böttcher, and Ed Hurt. Two structurally distinct domains of the nucleoporin Nup170 cooperate to tether a subset of nucleoporins to nuclear pores. Journal of Cell Biology, 185(3):387–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/3/387.

Funayama:2006:LLH

[FSTI06] Ryo Funayama, Motoki Saito, Hiroko Tanobe, and Fuyuki Ishikawa. Loss of linker histone H1 in cellular senescence. Journal of Cell Biology, 175(6):869–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/175/6/869.

Flynt:2009:MMR

[FTB^{+}09] Alex S. Flynt, Elizabeth J. Thatcher, Kristopher Burkewitz, Nan Li, Yinzi Liu, and James G. Patton. miR-8 microRNAs regulate the response to osmotic stress in zebrafish embryos. Journal of Cell Biology, 185(1):115–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/185/1/115.

Frazier:2006:MIR

[FTM^{+}06] Ann E. Frazier, Rebecca D. Taylor, David U. Mick, Bettina Warscheid, Nadine Stoepel, Helmut E. Meyer, Michael T. Ryan, Bernard Guiard, and Peter Rehling. Mdm38 interacts with ribosomes and is a component of the mitochondrial protein export machinery. Journal of Cell Biology, 172(4):553–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/172/4/
553.

Fukata:2005:MCN

[FTT^{+}05] Yuko Fukata, Anastassios V. Tzingounis, Jonathan C. Trinidad, Masaki Fukata, Alma L. Burlingame, Roger A. Nicoll, and David S. Bredt. Molecular constituents of neuronal AMPA receptors. Journal of Cell Biology, 169(3):399–??, May 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/169/3/399.

Fuchs:2008:SSC

[Fuc08] Elaine Fuchs. Skin stem cells: rising to the surface. Journal of Cell Biology, 180(2):273–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/2/273.

Foe:2008:SDM

[FvD08] Victoria E. Foe and George von Dassow. Stable and dynamic microtubules coordinately shape the myosin activation zone during cytokinetic furrow formation. Journal of Cell Biology, 183(3):457–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/3/457.

Famulski:2008:SHK

[FVSC08] Jakub K. Famulski, Larissa Vos, Xuejun Sun, and Gordon Chan. Stable hZW10 kinetochore residency, mediated by hZwint-1 interaction, is essential for the mitotic checkpoint. Journal of Cell Biology, 180(3):507–??, February 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/180/3/507.

Fuster:2007:GAE

[FWC^{+}07] Mark M. Fuster, Lianchun Wang, Janice Castagnola, Lyudmila Sikora, Krisanavane Reddi, Phillip H. A. Lee, Katherine A. Radek, Manuela Schuksz, Joseph R. Bishop, Richard L. Gallo, P. Sriramarao, and Jeffrey D. Esko. Genetic alteration of endothelial heparan sulfate selectively inhibits tumor angiogenesis. Journal of Cell Biology, 177(3):539–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/177/3/539.

Feral:2007:CSP

[FZT^{+}07] Chloé C. Féral, Andries Zijlstra, Eugene Tkachenko, Gerald Prager, Margaret L. Gardel, Marina Slepak, and Mark H. Ginsberg. CD98hc (SLC3A2) participates in fibronectin matrix assembly by mediating integrin signaling. Journal of Cell Biology, 178(4):701–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/4/701.

Grosshans:2006:YLF

[GAG^{+}06] Bianka L. Grosshans, Anna Andreeva, Akanksha Gangar, Sherry Niessen, John R. Yates, Patrick Brennwald, and Peter Novick. The yeast lgl family member Sro7p is an effector of the secretory Rab GTPase Sec4p. Journal of Cell Biology, 172(1):55–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/1/55.

Gupton:2005:CML

[GAK^{+}05] Stephanie L. Gupton, Karen L. Anderson, Thomas P. Kole, Robert S. Fischer, Aaron Ponti, Sarah E. Hitchcock-DeGregori, Gaudenz Danuser, Velia M. Fowler, Denis Wirtz, Dorit Hanein, and Clare M. Waterman-Storer. Cell migration without a lamellipodium. Journal of Cell Biology, 168(4):619–??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/168/4/
619.

Gill:2006:SPB

[GAS06] Tina Gill, Jason Aulds, and Mark E. Schmitt. A specialized processing body that is temporally and asymmetrically regulated during the cell cycle in Saccharomyces cerevisiae. Journal of Cell Biology, 173(1):35–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/173/1/35.

Geng:2008:QAAa

[GBNK08a] Jiefei Geng, Misuzu Baba, Usha Nair, and Daniel J. Klionsky. Quantitative analysis of autophagy-related protein stoichiometry by fluorescence microscopy. Journal of Cell Biology, 182 (1):129–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/1/129.

Geng:2008:QAAb

[GBNK08b] Jiefei Geng, Misuzu Baba, Usha Nair, and Daniel J. Klionsky. Quantitative analysis of autophagy-related protein stoichiometry by fluorescence microscopy. Journal of Cell Biology, 183(6):1175–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/6/1175.

Gu:2007:SRE

[GBNS07] Qingyang Gu, G. Tim Bowden, Daniel Normolle, and Yi Sun. SAG/ ROC2 E3 ligase regulates skin carcinogenesis by stagedependent targeting of c–Jun/AP1 and IκB-α /NF-κB. Journal of Cell Biology, 178(6):1009–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/178/6/1009.

Golub:2005:PPD

[GC05] Tamara Golub and Pico Caroni. PI(4,5)P_{2}-dependent microdomain assemblies capture microtubules to promote and control leading edge motility. Journal of Cell Biology, 169 (1):151–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/1/151.

Greaves:2007:PDP

[GC07] Jennifer Greaves and Luke H. Chamberlain. Palmitoylationdependent protein sorting. Journal of Cell Biology, 176(3): 249–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/3/249.

Gaus:2005:CPM

[GCF^{+}05] Katharina Gaus, Elena Chklovskaia, Barbara Fazekas de St.Groth, Wendy Jessup, and Thomas Harder. Condensation of the plasma membrane at the site of T lymphocyte activation. Journal of Cell Biology, 171(1):121–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/171/1/121.

Garcia-Cortes:2009:PTC

[GCM09] Juan Carlos García-Cortés and Dannel McCollum. Proper timing of cytokinesis is regulated by Schizosaccharomyces pombe Etd1. Journal of Cell Biology, 186(5):739–??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/186/5/739.

Goulding:2007:CNC

[GCS^{+}07] Morgan B. Goulding, Julie C. Canman, Eric N. Senning, Andrew H. Marcus, and Bruce Bowerman. Control of nuclear centration in the C. elegans zygote by receptor-independent Gα signaling and myosin II. Journal of Cell Biology, 178(7): 1177–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/7/1177.

Gavert:2005:LNT

[GCSG^{+}05] Nancy Gavert, Maralice Conacci-Sorrell, Daniela Gast, Annette Schneider, Peter Altevogt, Thomas Brabletz, and Avri Ben-Ze’ev. L1, a novel target of β-catenin signaling, transforms cells and is expressed at the invasive front of colon cancers. Journal of Cell Biology, 168(4):633–??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/168/4/633.

Goodchild:2005:APT

[GD05] Rose E. Goodchild and William T. Dauer. The AAA+ protein torsinA interacts with a conserved domain present in LAP1 and a novel ER protein. Journal of Cell Biology, 168(6):855–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
168/6/855.

Green:2008:SPA

[GDA^{+}08] Kim N. Green, Angelo Demuro, Yama Akbari, Brian D. Hitt, Ian F. Smith, Ian Parker, and Frank M. LaFerla. SERCA pump activity is physiologically regulated by presenilin and regulates amyloid β production. Journal of Cell Biology, 181(7):1107–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
181/7/1107.

Gondcaille:2005:PRA

[GDF^{+}05] Catherine Gondcaille, Marianne Depreter, Stéphane Fourcade, Maria Rita Lecca, Sabrina Leclercq, Pascal G. P. Martin, Thierry Pineau, Françoise Cadepond, Martine ElEtr, Nathalie Bertrand, Alain Beley, Sandrine Duclos, Dirk De Craemer, Frank Roels, Stéphane Savary, and Maurice Bugaut. Phenylbutyrate up-regulates the adrenoleukodystrophy-related gene as a nonclassical peroxisome proliferator. Journal of Cell Biology, 169(1):93–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/1/93.

Goodwin:2005:DRT

[GDR^{+}05] J. Shawn Goodwin, Kimberly R. Drake, Carl Rogers, Latasha Wright, Jennifer Lippincott-Schwartz, Mark R. Philips, and Anne K. Kenworthy. Depalmitoylated Ras traffics to and from the Golgi complex via a nonvesicular pathway. Journal of Cell Biology, 170(2):261–??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/2/261.

Gadea:2007:LPP

[GdTAR07] Gilles Gadea, Marion de Toledo, Christelle Anguille, and Pierre Roux. Loss of p53 promotes RhoA–ROCK-dependent cell migration and invasion in 3D matrices. Journal of Cell Biology, 178(1):23–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/1/23.

Gentry:2007:PLC

[GDW^{+}07] Matthew S. Gentry, Robert H. Dowen, Carolyn A. Worby, Seema Mattoo, Joseph R. Ecker, and Jack E. Dixon. The phosphatase laforin crosses evolutionary boundaries and links carbohydrate metabolism to neuronal disease. Journal of Cell Biology, 178(3):477–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/3/477.

Grewal:2007:DTI

[GEE07] Savraj S. Grewal, Justin R. Evans, and Bruce A. Edgar. Drosophila TIF–IA is required for ribosome synthesis and cell growth and is regulated by the TOR pathway. Journal of Cell Biology, 179(6):1105–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/6/1105.

Gillissen:2007:MDB

[GEH^{+}07] Bernhard Gillissen, Frank Essmann, Philipp G. Hemmati, Antje Richter, Anja Richter, Ilker Öztop, Govindaswamy Chinnadurai, Bernd Dörken, and Peter T. Daniel. Mcl-1 determines the Bax dependency of Nbk/Bik-induced apoptosis. Journal of Cell Biology, 179(4):701–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/4/701.

Gerhart:2006:MPE

[GEN^{+}06] Jacquelyn Gerhart, Justin Elder, Christine Neely, Jared Schure, Tage Kvist, Karen Knudsen, and Mindy George-Weinstein. MyoD-positive epiblast cells regulate skeletal muscle differentiation in the embryo. Journal of Cell Biology, 175(2):283–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
175/2/283.

Ganley:2008:SSC

[GEP08] Ian G. Ganley, Eric Espinosa, and Suzanne R. Pfeffer. A syntaxin 10-SNARE complex distinguishes two distinct transport routes from endosomes to the trans-Golgi in human cells. Journal of Cell Biology, 180(1):159–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/180/1/159.

Govin:2007:PHR

[GER^{+}07] Jérôme Govin, Emmanuelle Escoffier, Sophie Rousseaux, Lauriane Kuhn, Myriam Ferro, Julien Thévenon, Raffaella Catena, Irwin Davidson, Jérôme Garin, Saadi Khochbin, and Cécile Caron. Pericentric heterochromatin reprogramming by new histone variants during mouse spermiogenesis. Journal of Cell Biology, 176(3):283–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/3/283.

Gilljam:2009:INW

[GFA^{+}09] Karin M. Gilljam, Emadoldin Feyzi, Per A. Aas, Mirta M. L. Sousa, Rebekka Müller, Cathrine B. Vågbø, Tara C. Catterall, Nina B. Liabakk, Geir Slupphaug, Finn Drabløs, Hans E. Krokan, and Marit Otterlei. Identification of a novel, widespread, and functionally important PCNA-binding motif. Journal of Cell Biology, 186(5):645–??, September 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/186/5/645.

Grund:2008:INM

[GFC^{+}08] Stefanie E. Grund, Tamás Fischer, Ghislain G. Cabal, Oreto Antúnez, José E. Pérez-Ortín, and Ed Hurt. The inner nuclear membrane protein Src1 associates with subtelomeric genes and alters their regulated gene expression. Journal of Cell Biology, 182(5):897–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/5/897.

Gu:2005:RFG

[GFD^{+}05] Ying Gu, Ying Fu, Peter Dowd, Shundai Li, Vanessa Vernoud, Simon Gilroy, and Zhenbiao Yang. A Rho family GTPase controls actin dynamics and tip growth via two counteracting downstream pathways in pollen tubes. Journal of Cell Biology, 169(1):127–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/1/127.

Gebhardt:2006:MRK

[GFH^{+}06] Anneli Gebhardt, Michaela Frye, Steffi Herold, Salvador Aznar Benitah, Kristin Braun, Birgit Samans, Fiona M. Watt, Hans-Peter Elsässer, and Martin Eilers. Myc regulates keratinocyte adhesion and differentiation via complex formation with Miz1. Journal of Cell Biology, 172(1):139–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/172/1/139.

Ge:2006:BCT

[GG06] Gaoxiang Ge and Daniel S. Greenspan. BMP1 controls TGFβ1 activation via cleavage of latent TGFβ-binding protein. Journal of Cell Biology, 175(1):111–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/175/1/111.

Gohil:2009:MMB

[GG09] Vishal M. Gohil and Miriam L. Greenberg. Mitochondrial membrane biogenesis: phospholipids and proteins go hand in hand. Journal of Cell Biology, 184(4):469–??, February 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/184/4/469.

Gao:2008:ACP

[GGB^{+}08] Shan Gao, Maria Grazia Giansanti, Graham J. Buttrick, Sharada Ramasubramanyan, Adam Auton, Maurizio Gatti, and James G. Wakefield. Australin: a chromosomal passenger protein required specifically for Drosophila melanogaster male meiosis. Journal of Cell Biology, 180(3):521–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/180/3/521.

Guo:2007:SIP

[GGBV^{+}07] Tong Guo, Christopher Gregg, Tatiana Boukh-Viner, Pavlo Kyryakov, Alexander Goldberg, Simon Bourque, Farhana Banu, Sandra Haile, Svetlana Milijevic, Karen Hung Yeung San, Jonathan Solomon, Vivianne Wong, and Vladimir I. Titorenko. A signal from inside the peroxisome initiates its division by promoting the remodeling of the peroxisomal membrane. Journal of Cell Biology, 177(2):289–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/177/2/289.

Griffin:2005:WPC

[GGC05] Erik E. Griffin, Johannes Graumann, and David C. Chan. The WD40 protein Caf4p is a component of the mitochondrial fission machinery and recruits Dnm1p to mitochondria. Journal of Cell Biology, 170(2):237–??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/2/237.

Gontan:2009:EMN

[GGE^{+}09] Cristina Gontan, Thomas Güttler, Erik Engelen, Jeroen Demmers, Maarten Fornerod, Frank G. Grosveld, Dick Tibboel, Dirk Görlich, Raymond A. Poot, and Robbert J. Rottier. Exportin 4 mediates a novel nuclear import pathway for Sox family transcription factors. Journal of Cell Biology, 185(1):27–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
185/1/27.

Grande-Garcia:2007:CRC

[GGEd^{+}07] Araceli Grande-García, Asier Echarri, Johan de Rooij, Nazilla B. Alderson, Clare M. Waterman-Storer, José M. Valdivielso, and Miguel A. del Pozo. Caveolin-1 regulates cell polarization and directional migration through Src kinase and Rho GTPases. Journal of Cell Biology, 177(4):683–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/177/4/
683.

Grim:2008:ICC

[GGH^{+}08] Jonathan E. Grim, Michael P. Gustafson, Roli K. Hirata, Amanda C. Hagar, Jherek Swanger, Markus Welcker, Harry C. Hwang, Johan Ericsson, David W. Russell, and Bruce E. Clurman. Isoform- and cell cycle-dependent substrate degradation by the Fbw7 ubiquitin ligase. Journal of Cell Biology, 181 (6):913–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/6/913.

Ghosh:2008:AGT

[GGMBF08] Pradipta Ghosh, Mikel Garcia-Marcos, Scott J. Bornheimer, and Marilyn G. Farquhar. Activation of Gα i3 triggers cell migration via regulation of GIV. Journal of Cell Biology, 182 (2):381–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/2/381.

Gallenne:2009:BAB

[GGO^{+}09] Tristan Gallenne, Fabien Gautier, Lisa Oliver, Eric Hervouet, Belinda Noël, John A. Hickman, Olivier Geneste, Pierre-François Cartron, François M. Vallette, Stephen Manon, and Philippe Juin. Bax activation by the BH3-only protein Puma promotes cell dependence on antiapoptotic Bcl-2 family members. Journal of Cell Biology, 185(2):279–??, April 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/185/2/279.

Gechev:2005:HPS

[GH05] Tsanko S. Gechev and Jacques Hille. Hydrogen peroxide as a signal controlling plant programmed cell death. Journal of Cell Biology, 168(1):17–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/1/17.

Godsel:2005:DAD

[GHA^{+}05] Lisa M. Godsel, Sherry N. Hsieh, Evangeline V. Amargo, Amanda E. Bass, Lauren T. Pascoe-McGillicuddy, Arthur C. Huen, Meghan E. Thorne, Claire A. Gaudry, Jung K. Park, Kyunghee Myung, Robert D. Goldman, Teng-Leong Chew, and Kathleen J. Green. Desmoplakin assembly dynamics in four dimensions. Journal of Cell Biology, 171(6):1045–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/171/6/
1045.

Gil-Henn:2007:DMD

[GHDS^{+}07] Hava Gil-Henn, Olivier Destaing, Natalie A. Sims, Kazuhiro Aoki, Neil Alles, Lynn Neff, Archana Sanjay, Angela Bruzzaniti, Pietro De Camilli, Roland Baron, and Joseph Schlessinger. Defective microtubule-dependent podosome organization in osteoclasts leads to increased bone density in Pyk2^{−/−} mice. Journal of Cell Biology, 178(6):1053–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/178/6/1053.

Guan:2009:DRO

[GHK^{+}09] Zeqiang Guan, Christina M. Hughes, Settapong Kosiyatrakul, Paolo Norio, Ranjan Sen, Steven Fiering, C. David Allis, Eric E. Bouhassira, and Carl L. Schildkraut. Decreased replication origin activity in temporal transition regions. Journal of Cell Biology, 187(5):623–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/5/623.

Gardner:2008:MBM

[GHM^{+}08] Melissa K. Gardner, Julian Haase, Karthikeyan Mythreye, Jeffrey N. Molk, MaryBeth Anderson, Ajit P. Joglekar, Eileen T. O’Toole, Mark Winey, E. D. Salmon, David J. Odde, and Kerry Bloom. The microtubule-based motor Kar3 and plus end–binding protein Bim1 provide structural support for the anaphase spindle. Journal of Cell Biology, 180(1):91–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
180/1/91.

Geserick:2009:CII

[GHM^{+}09] Peter Geserick, Mike Hupe, Maryline Moulin, W. Wei-Lynn Wong, Maria Feoktistova, Beate Kellert, Harald Gollnick, John Silke, and Martin Leverkus. Cellular IAPs inhibit a cryptic CD95-induced cell death by limiting RIP1 kinase recruitment. Journal of Cell Biology, 187(7):1037–??, December 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/187/7/1037.

Goldoni:2009:DNA

[GHN^{+}09] Silvia Goldoni, Ashley Humphries, Alexander Nyström, Sampurna Sattar, Rick T. Owens, David J. McQuillan, Keith Ireton, and Renato V. Iozzo. Decorin is a novel antagonistic ligand of the Met receptor. Journal of Cell Biology, 185(4):743–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/185/4/
743.

Gladfelter:2006:AND

[GHP06] Amy S. Gladfelter, A. Katrin Hungerbuehler, and Peter Philippsen. Asynchronous nuclear division cycles in multinucleated cells. Journal of Cell Biology, 172(3):347–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/172/3/
347.

Glading:2007:KCR

[GHSG07] Angela Glading, Jaewon Han, Rebecca A. Stockton, and Mark H. Ginsberg. KRIT-1/ CCM1 is a Rap1 effector that regulates endothelial cell–cell junctions. Journal of Cell Biology, 179(2):247–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/2/247.

Gaboreanu:2007:MPZ

[GHX^{+}07] Ana-Maria Gaboreanu, Ronald Hrstka, Wenbo Xu, Michael Shy, John Kamholz, Jack Lilien, and Janne Balsamo. Myelin protein zero/P0 phosphorylation and function require an adaptor protein linking it to RACK1 and PKCα. Journal of Cell Biology, 177(4):707–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/4/707.

Giraudo:2005:SCP

[GHY^{+}05] Claudio G. Giraudo, Chuan Hu, Daoqi You, Avram M. Slovic, Eugene V. Mosharov, David Sulzer, Thomas J. Melia, and James E. Rothman. SNAREs can promote complete fusion and hemifusion as alternative outcomes. Journal of Cell Biology, 170(2):249–??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/2/249.

Goetz:2008:CRF

[GJL^{+}08] Jacky G. Goetz, Bharat Joshi, Patrick Lajoie, Scott S. Strugnell, Trevor Scudamore, Liliana D. Kojic, and Ivan R. Nabi. Concerted regulation of focal adhesion dynamics by galectin-3 and tyrosine-phosphorylated caveolin-1. Journal of Cell Biology, 180(6):1261–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/6/1261.

Galabova-Kovacs:2008:ERB

[GKCM^{+}08] Gergana Galabova-Kovacs, Federica Catalanotti, Dana Matzen, Gloria X. Reyes, Jürgen Zezula, Ruth Herbst, Alcino Silva, Ingrid Walter, and Manuela Baccarini. Essential role of B–Raf in oligodendrocyte maturation and myelination during postnatal central nervous system development. Journal of Cell Biology, 180(5):947–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/5/947.

Grove:2007:FRA

[GKN^{+}07] Matthew Grove, Noboru H. Komiyama, Klaus-Armin Nave, Seth G. Grant, Diane L. Sherman, and Peter J. Brophy. FAK is required for axonal sorting by Schwann cells. Journal of Cell Biology, 176(3):277–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/3/277.

Griparic:2007:RMD

[GKvdB07] Lorena Griparic, Takayuki Kanazawa, and Alexander M. van der Bliek. Regulation of the mitochondrial dynamin-like protein Opa1 by proteolytic cleavage. Journal of Cell Biology, 178(5):757–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/5/757.

Guo:2006:CGP

[GL06] Yusong Guo and Adam D. Linstedt. COPII–Golgi protein interactions regulate COPII coat assembly and Golgi size. Journal of Cell Biology, 174(1):53–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/174/1/53.

Gaus:2006:IMA

[GLBS06] Katharina Gaus, Soazig Le Lay, Nagaraj Balasubramanian, and Martin A. Schwartz. Integrin-mediated adhesion regulates membrane order. Journal of Cell Biology, 174(5):725–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
174/5/725.

Glembotski:2007:GGR

[Gle07] Christopher C. Glembotski. Getting a G–RRP on regulated exocytosis in the heart. Journal of Cell Biology, 179(3):371– ??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/3/371.

Gribnau:2005:XCC

[GLH^{+}05] Joost Gribnau, Sandra Luikenhuis, Konrad Hochedlinger, Kim Monkhorst, and Rudolf Jaenisch. X chromosome choice occurs independently of asynchronous replication timing. Journal of Cell Biology, 168(3):365–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/168/3/365.

Gillespie:2009:PDG

[GLS^{+}09] Mark A. Gillespie, Fabien Le Grand, Anthony Scimè, Shihuan Kuang, Julia von Maltzahn, Vanessa Seale, Ana Cuenda, Jeffrey A. Ranish, and Michael A. Rudnicki. p38-γ-dependent gene silencing restricts entry into the myogenic differentiation program. Journal of Cell Biology, 187(7):991–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/187/7/991.

Gard:2009:CMDa

[GLX^{+}09a] Scarlett Gard, William Light, Bo Xiong, Tania Bose, Adrian J. McNairn, Bethany Harris, Brian Fleharty, Chris Seidel, Jason H. Brickner, and Jennifer L. Gerton. Cohesinopathy mutations disrupt the subnuclear organization of chromatin. Journal of Cell Biology, 187(4):455–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/187/4/455.

Gard:2009:CMDb

[GLX^{+}09b] Scarlett Gard, William Light, Bo Xiong, Tania Bose, Adrian J. McNairn, Bethany Harris, Brian Fleharty, Chris Seidel, Jason H. Brickner, and Jennifer L. Gerton. Cohesinopathy mutations disrupt the subnuclear organization of chromatin. Journal of Cell Biology, 187(5):749–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/187/5/749.

Gauthier:2007:EEA

[GMG^{+}07] Nils C. Gauthier, Pascale Monzo, Teresa Gonzalez, Anne Doye, Amanda Oldani, Pierre Gounon, Vittorio Ricci, Mireille Cormont, and Patrice Boquet. Early endosomes associated with dynamic F-actin structures are required for late trafficking of H. pylori VacA toxin. Journal of Cell Biology, 177(2):343–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
177/2/343.

Gomi:2005:GMD

[GMK^{+}05] Hiroshi Gomi, Shin Mizutani, Kazuo Kasai, Shigeyoshi Itohara, and Tetsuro Izumi. Granuphilin molecularly docks insulin granules to the fusion machinery. Journal of Cell Biology, 171(1):99–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/1/99.

Giacobini:2008:SRG

[GMM^{+}08] Paolo Giacobini, Andrea Messina, Francesca Morello, Nicoletta Ferraris, Simona Corso, Junia Penachioni, Silvia Giordano, Luca Tamagnone, and Aldo Fasolo. Semaphorin 4D regulates gonadotropin hormone-releasing hormone-1 neuronal migration through PlexinB1–Met complex. Journal of Cell Biology, 183(3):555–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/3/555.

Grossmann:2008:PMM

[GMS^{+}08] Guido Grossmann, Jan Malinsky, Wiebke Stahlschmidt, Martin Loibl, Ina Weig-Meckl, Wolf B. Frommer, Miroslava Opekarová, and Widmar Tanner. Plasma membrane microdomains regulate turnover of transport proteins in yeast. Journal of Cell Biology, 183(6):1075–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/6/1075.

Glater:2006:ATM

[GMSS06] Elizabeth E. Glater, Laura J. Megeath, R. Steven Stowers, and Thomas L. Schwarz. Axonal transport of mitochondria requires milton to recruit kinesin heavy chain and is light chain independent. Journal of Cell Biology, 173(4):545–??, May 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/173/4/545.

Gerard:2007:PPC

[GMvdKC07] Audrey Gérard, Alexander E. E. Mertens, Rob A. van der Kammen, and John G. Collard. The Par polarity complex regulates Rap1- and chemokine-induced T cell polarization. Journal of Cell Biology, 176(6):863–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/176/6/863.

Gan:2005:IFI

[GMW^{+}05] Boyi Gan, Zara K. Melkoumian, Xiaoyang Wu, Kun-Liang Guan, and Jun-Lin Guan. Identification of FIP200 interaction with the TSC1–TSC2 complex and its role in regulation of cell size control. Journal of Cell Biology, 170(3):379–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/170/3/
379.

Goshima:2008:APC

[GMZ^{+}08] Gohta Goshima, Mirjam Mayer, Nan Zhang, Nico Stuurman, and Ronald D. Vale. Augmin: a protein complex required for centrosome-independent microtubule generation within the spindle. Journal of Cell Biology, 181(3):421–??, May 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/181/3/421.

Gerhart:2007:CEM

[GNE^{+}07] Jacquelyn Gerhart, Christine Neely, Justin Elder, Jessica Pfautz, Jordanna Perlman, Luis Narciso, Kersti K. Linask, Karen Knudsen, and Mindy George-Weinstein. Cells that express MyoD mRNA in the epiblast are stably committed to the skeletal muscle lineage. Journal of Cell Biology, 178(4):649–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
178/4/649.

Gruneberg:2006:KCK

[GNL^{+}06] Ulrike Gruneberg, Rüdiger Neef, Xiuling Li, Eunice H. Y. Chan, Ravindra B. Chalamalasetty, Erich A. Nigg, and Francis A. Barr. KIF14 and citron kinase act together to promote efficient cytokinesis. Journal of Cell Biology, 172(3):363–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/172/3/
363.

Gonzalez:2008:FRA

[GNPC^{+}08] José María González, Ana Navarro-Puche, Berta Casar, Piero Crespo, and Vicente Andrés. Fast regulation of AP-1 activity through interaction of lamin A/C, ERK1/2, and c–Fos at the nuclear envelope. Journal of Cell Biology, 183(4): 653–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/4/653.

Goshima:2005:MFM

[GNV05] Gohta Goshima, François Nédélec, and Ronald D. Vale. Mechanisms for focusing mitotic spindle poles by minus end-directed motor proteins. Journal of Cell Biology, 171(2):229–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/171/2/
229.

Glozman:2009:GDD

[GOM^{+}09] Rina Glozman, Tsukasa Okiyoneda, Cory M. Mulvihill, James M. Rini, Herve Barriere, and Gergely L. Lukacs. Nglycans are direct determinants of CFTR folding and stability in secretory and endocytic membrane traffic. Journal of Cell Biology, 184(6):847–??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/6/847.

Goffin:2006:FAS

[GPC^{+}06] Jérôme M. Goffin, Philippe Pittet, Gabor Csucs, Jost W. Lussi, Jean-Jacques Meister, and Boris Hinz. Focal adhesion size controls tension-dependent recruitment of α-smooth muscle actin to stress fibers. Journal of Cell Biology, 172(2):259–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/172/2/
259.

Goldfinger:2006:RRR

[GPJ^{+}06] Lawrence E. Goldfinger, Celeste Ptak, Erin D. Jeffery, Jeffrey Shabanowitz, Donald F. Hunt, and Mark H. Ginsberg. RLIP76 (RalBP1) is an R–Ras effector that mediates adhesiondependent Rac activation and cell migration. Journal of Cell Biology, 174(6):877–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/174/6/877.

Gan:2006:RFC

[GPN^{+}06] Boyi Gan, Xu Peng, Tamas Nagy, Ana Alcaraz, Hua Gu, and Jun-Lin Guan. Role of FIP200 in cardiac and liver development and its regulation of TNFα and TSC–mTOR signaling pathways. Journal of Cell Biology, 175(1):121–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/175/1/121.

Gonzalez-Quevedo:2005:RTP

[GQSHO05] Rosa Gonzalez-Quevedo, Marina Shoffer, Lily Horng, and Anthony E. Oro. Receptor tyrosine phosphatase-dependent cytoskeletal remodeling by the hedgehog-responsive gene MIM/ BEG4. Journal of Cell Biology, 168(3):453–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/168/3/453.

Grammont:2007:AJR

[Gra07] Muriel Grammont. Adherens junction remodeling by the Notch pathway in Drosophila melanogaster oogenesis. Journal of Cell Biology, 177(1):139–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/1/139.

Gerbino:2005:TCS

[GRC^{+}05] Andrea Gerbino, Warren C. Ruder, Silvana Curci, Tullio Pozzan, Manuela Zaccolo, and Aldebaran M. Hofer. Termination of cAMP signals by Ca^{2+} and Gα_{i} via extracellular Ca^{2+} sensors. Journal of Cell Biology, 171(2):303–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/171/2/303.

Garcia-Rodriguez:2007:PPF

[GRGP07] Luis J. García-Rodríguez, Anna Card Gay, and Liza A. Pon. Puf3p, a Pumilio family RNA binding protein, localizes to mitochondria and regulates mitochondrial biogenesis and motility in budding yeast. Journal of Cell Biology, 176(2):197–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/176/2/
197.

Galante:2007:RST

[GS07] Leontine L. Galante and Jean E. Schwarzbauer. Requirements for sulfate transport and the diastrophic dysplasia sulfate transporter in fibronectin matrix assembly. Journal of Cell Biology, 179(5):999–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/5/999.

Galvez:2006:CRD

[GSB^{+}06a] Beatriz G. Galvez, Maurilio Sampaolesi, Silvia Brunelli, Diego Covarello, Manuela Gavina, Barbara Rossi, Gabriela Constantin, Yvan Torrente, and Giulio Cossu. Complete repair of dystrophic skeletal muscle by mesoangioblasts with enhanced migration ability. Journal of Cell Biology, 174(2):231–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/174/2/
231.

Galvez:2006:CPb

[GSB^{+}06b] Beatriz G. Galvez, Maurilio Sampaolesi, Silvia Brunelli, Diego Covarello, Manuela Gavina, Barbara Rossi, Gabriela Constantin, Yvan Torrente, and Giulio Cossu. Correction 174, p.

231. Journal of Cell Biology, 175(2):361–??, October 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/175/2/361.

Galvez:2006:CPa

[GSB^{+}06c] Beatriz G. Galvez, Maurilio Sampaolesi, Silvia Brunelli, Diego Covarello, Manuela Gavina, Barbara Rossi, Gabriela Costantin, Yvan Torrente, and Giulio Cossu. Correction 174, p. 231. Journal of Cell Biology, 174(4):605–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/174/4/605.

Geffers:2007:DFD

[GSC^{+}07] Insa Geffers, Katrin Serth, Gavin Chapman, Robert Jaekel, Karin Schuster-Gossler, Ralf Cordes, Duncan B. Sparrow, Elisabeth Kremmer, Sally L. Dunwoodie, Thomas Klein, and Achim Gossler. Divergent functions and distinct localization of the Notch ligands DLL1 and DLL3 in vivo. Journal of Cell Biology, 178(3):465–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/3/465.

Geymonat:2009:LCB

[GSdBS09] Marco Geymonat, Adonis Spanos, Geoffroy de Bettignies, and Steven G. Sedgwick. Lte1 contributes to Bfa1 localization rather than stimulating nucleotide exchange by Tem1. Journal of Cell Biology, 187(4):497–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/187/4/497.

Gonzalvez:2008:CPE

[GSH^{+}08] Francois Gonzalvez, Zachary T. Schug, Riekelt H. Houtkooper, Elaine D. MacKenzie, David G. Brooks, Ronald J. A. Wanders, Patrice X. Petit, Frédéric M. Vaz, and Eyal Gottlieb. Cardiolipin provides an essential activating platform for caspase-8 on mitochondria. Journal of Cell Biology, 183(4):681–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/183/4/
681.

Gilkerson:2008:MNM

[GSHD08] Robert W. Gilkerson, Eric A. Schon, Evelyn Hernandez, and Mercy M. Davidson. Mitochondrial nucleoids maintain genetic autonomy but allow for functional complementation. Journal of Cell Biology, 181(7):1117–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/181/7/1117.

Getsios:2009:DDS

[GSiK^{+}09] Spiro Getsios, Cory L. Simpson, Shin ichiro Kojima, Robert Harmon, Linda J. Sheu, Rachel L. Dusek, Mona Cornwell, and Kathleen J. Green. Desmoglein 1-dependent suppression of EGFR signaling promotes epidermal differentiation and morphogenesis. Journal of Cell Biology, 185(7):1243–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/185/7/
1243.

Gardel:2008:TSF

[GSJ^{+}08] Margaret L. Gardel, Benedikt Sabass, Lin Ji, Gaudenz Danuser, Ulrich S. Schwarz, and Clare M. Waterman. Traction stress in focal adhesions correlates biphasically with actin retrograde flow speed. Journal of Cell Biology, 183(6):999–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/183/6/

Graser:2007:CNC

[GSL^{+}07] Susanne Graser, York-Dieter Stierhof, Sébastien B. Lavoie, Oliver S. Gassner, Stefan Lamla, Mikael Le Clech, and Erich A. Nigg. Cep164, a novel centriole appendage protein required for primary cilium formation. Journal of Cell Biology, 179(2): 321–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/2/321.

Griffis:2007:SNP

[GSV07] Eric R. Griffis, Nico Stuurman, and Ronald D. Vale. Spindly, a novel protein essential for silencing the spindle assembly checkpoint, recruits dynein to the kinetochore. Journal of Cell Biology, 177(6):1005–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/6/1005.

Gu:2008:AFE

[GSW^{+}08] Mingyu Gu, Kim Schuske, Shigeki Watanabe, Qiang Liu, Paul Baum, Gian Garriga, and Erik M. Jorgensen. μ2 adaptin facilitates but is not essential for synaptic vesicle recycling in Caenorhabditis elegans. Journal of Cell Biology, 183(5): 881–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/5/881.

Gu:2009:PEM

[GSY^{+}09] Bingnan Gu, Peng Sun, Yuanyang Yuan, Ricardo C. Moraes, Aihua Li, Andy Teng, Anshu Agrawal, Catherine Rhéaume, Virginia Bilanchone, Jacqueline M. Veltmaat, Ken-Ichi Takemaru, Sarah Millar, Eva Y.-H. P. Lee, Michael T. Lewis, Boan Li, and Xing Dai. Pygo2 expands mammary progenitor cells by facilitating histone H3 K4 methylation. Journal of Cell Biology, 185(5):811–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/5/811.

Goss:2008:BDC

[GT08] John W. Goss and Derek K. Toomre. Both daughter cells traffic and exocytose membrane at the cleavage furrow during mammalian cytokinesis. Journal of Cell Biology, 181(7):1047–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
181/7/1047.

Gilbert:2007:DMA

[GTB^{+}07] Nick Gilbert, Inga Thomson, Shelagh Boyle, James Allan, Bernard Ramsahoye, and Wendy A. Bickmore. DNA methylation affects nuclear organization, histone modifications, and linker histone binding but not chromatin compaction. Journal of Cell Biology, 177(3):401–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/177/3/401.

Gillis:2005:NRC

[GTHK05] Amethyst N. Gillis, Scott Thomas, Scott D. Hansen, and Kenneth B. Kaplan. A novel role for the CBF3 kinetochore–scaffold complex in regulating septin dynamics and cytokinesis. Journal of Cell Biology, 171(5):773–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/171/5/773.

Gonsalvez:2007:TDA

[GTO^{+}07] Graydon B. Gonsalvez, Liping Tian, Jason K. Ospina, François-Michel Boisvert, Angus I. Lamond, and A. Gregory Matera. Two distinct arginine methyltransferases are required for biogenesis of Sm-class ribonucleoproteins. Journal of Cell Biology, 178(5):733–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/5/733.

Goetze:2006:BSD

[GTX^{+}06] Bernhard Goetze, Fabian Tuebing, Yunli Xie, Mario M. Dorostkar, Sabine Thomas, Ulrich Pehl, Stefan Boehm, Paolo Macchi, and Michael A. Kiebler. The brain-specific doublestranded RNA-binding protein Staufen2 is required for dendritic spine morphogenesis. Journal of Cell Biology, 172(2): 221–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/2/221.

Gokhale:2009:RDD

[GWS09] Avanti Gokhale, Maureen Wirschell, and Winfield S. Sale. Regulation of dynein-driven microtubule sliding by the axonemal protein kinase CK1 in Chlamydomonas flagella. Journal of Cell Biology, 186(6):817–??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/186/6/817.

Ge:2007:CRE

[GXJF07] Chunxi Ge, Guozhi Xiao, Di Jiang, and Renny T. Franceschi. Critical role of the extracellular signal-regulated kinase-MAPK pathway in osteoblast differentiation and skeletal development. Journal of Cell Biology, 176(5):709–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/176/5/709.

Garvalov:2006:LPW

[GZBM^{+}06] Boyan K. Garvalov, Benoît Zuber, Cédric Bouchet-Marquis, Mikhail Kudryashev, Manuela Gruska, Martin Beck, Andrew Leis, Friedrich Frischknecht, Frank Bradke, Wolfgang Baumeister, Jacques Dubochet, and Marek Cyrklaff. Luminal particles within cellular microtubules. Journal of Cell Biology, 174(6): 759–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/6/759.

Honda:2005:TAE

[HAAHD05] Akira Honda, Omayma S. Al-Awar, Jesse C. Hay, and Julie G. Donaldson. Targeting of Arf-1 to the early Golgi by membrin, an ER–Golgi SNARE. Journal of Cell Biology, 168(7):1039–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
168/7/1039.

Hackney:2007:JSK

[Hac07] David D. Hackney. Jump-starting kinesin. Journal of Cell Biology, 176(1):7–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/1/7.

Hofmann:2009:SNA

[HAN^{+}09] Wilma A. Hofmann, Alessandro Arduini, Samantha M. Nicol, Carlos J. Camacho, James L. Lessard, Frances V. Fuller-Pace, and Primal de Lanerolle. SUMOylation of nuclear actin. Journal of Cell Biology, 186(2):193–??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/186/2/193.

Harrison:2008:PSF

[Har08] Stephen C. Harrison. The pH sensor for flavivirus membrane fusion. Journal of Cell Biology, 183(2):177–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/183/2/177.

Hatten:2005:LLN

[Hat05a] Mary E. Hatten. LIS-less neurons don’t even make it to the starting gate. Journal of Cell Biology, 170(6):867–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/170/6/
867.

Higashino:2005:AET

[HAT^{+}05b] Fumihiro Higashino, Mariko Aoyagi, Akiko Takahashi, Masaho Ishino, Masato Taoka, Toshiaki Isobe, Masanobu Kobayashi, Yasunori Totsuka, Takao Kohgo, and Masanobu Shindoh. Adenovirus E4orf6 targets pp32/LANP to control the fate of AREcontaining mRNAs by perturbing the CRM1-dependent mechanism. Journal of Cell Biology, 170(1):15–??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/170/1/15.

Helenius:2008:FSD

[HB08] Iiro Taneli Helenius and Greg J. Beitel. The first “Slit” is the deepest: the secret to a hollow heart. Journal of Cell Biology, 182(2):221–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/2/221.

Henkler:2005:EDF

[HBD^{+}05] Frank Henkler, Eva Behrle, Kevin M. Dennehy, Andreas Wicovsky, Nathalie Peters, Clemens Warnke, Klaus Pfizenmaier, and Harald Wajant. The extracellular domains of FasL and Fas are sufficient for the formation of supramolecular FasL–Fas clusters of high stability. Journal of Cell Biology, 168(7):1087–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/168/7/

Hagedorn:2006:DMA

[HBK^{+}06] Elliott J. Hagedorn, Jennifer L. Bayraktar, Vasundhara R. Kandachar, Ting Bai, Dane M. Englert, and Henry C. Chang. Drosophila melanogaster auxilin regulates the internalization of Delta to control activity of the Notch signaling pathway. Journal of Cell Biology, 173(3):443–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/173/3/443.

Hussain:2009:NRM

[HBN^{+}09] Shobbir Hussain, Sandra Blanco Benavente, Elisabete Nascimento, Ilaria Dragoni, Agata Kurowski, Astrid Gillich, Peter Humphreys, and Michaela Frye. The nucleolar RNA methyltransferase Misu (NSun2) is required for mitotic spindle stability. Journal of Cell Biology, 186(1):27–??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/186/1/27.

Hall:2009:AME

[HBPL09] Lisa L. Hall, Meg Byron, Gayle Pageau, and Jeanne B. Lawrence. AURKB-mediated effects on chromatin regulate binding versus release of XIST RNA to the inactive chromosome. Journal of Cell Biology, 186(4):491–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/186/4/491.

Hou:2008:ECD

[HCB^{+}08] Ying-Chen Claire Hou, Suganthi Chittaranjan, Sharon González Barbosa, Kimberly McCall, and Sharon M. Gorski. Effector caspase Dcp-1 and IAP protein Bruce regulate starvation-induced autophagy during Drosophila melanogaster oogenesis. Journal of Cell Biology, 182(6):1127–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/182/6/1127.

Hoyle:2007:SDR

[HCC^{+}07] Nathaniel P. Hoyle, Lydia M. Castelli, Susan G. Campbell, Leah E. A. Holmes, and Mark P. Ashe. Stress-dependent relocalization of translationally primed mRNPs to cytoplasmic granules that are kinetically and spatially distinct from Pbodies. Journal of Cell Biology, 179(1):65–??, October 2007.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/179/1/65.

Hu:2008:CPD

[HCFM08] Chi-Kuo Hu, Margaret Coughlin, Christine M. Field, and Timothy J. Mitchison. Cell polarization during monopolar cytokinesis. Journal of Cell Biology, 181(2):195–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/181/2/195.

Hou:2007:AAS

[HCK^{+}07] Fajian Hou, Chih-Wen Chu, Xiangduo Kong, Kyoko Yokomori, and Hui Zou. The acetyltransferase activity of San stabilizes the mitotic cohesin at the centromeres in a shugoshin-independent manner. Journal of Cell Biology, 177(4):587–??, May 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/177/4/587.

Hwang:2007:TCI

[HCL^{+}07] David K. Hwang, Steven M. Claypool, Danielle Leuenberger, Heather L. Tienson, and Carla M. Koehler. Tim54p connects inner membrane assembly and proteolytic pathways in the mitochondrion. Journal of Cell Biology, 178(7):1161–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/178/7/
1161.

Hennig:2006:TCB

[HCN06] Krista M. Hennig, Julien Colombani, and Thomas P. Neufeld. TOR coordinates bulk and targeted endocytosis in the Drosophila melanogaster fat body to regulate cell growth. Journal of Cell Biology, 173(6):963–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/173/6/963.

Hancock:2008:PTI

[HCRT08] Melissa L. Hancock, Sarah E. Canetta, Lorna W. Role, and David A. Talmage. Presynaptic Type III Neuregulin1-ErbB signaling targets α7 nicotinic acetylcholine receptors to axons. Journal of Cell Biology, 181(3):511–??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/181/3/511.

Hofmann:2006:NAP

[HdL06] Wilma A. Hofmann and Primal de Lanerolle. Nuclear actin: to polymerize or not to polymerize. Journal of Cell Biology, 172 (4):495–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/4/495.

Huang:2007:FDP

[HDM^{+}07] Kaiyao Huang, Dennis R. Diener, Aaron Mitchell, Gregory J. Pazour, George B. Witman, and Joel L. Rosenbaum. Function and dynamics of PKD2 in Chlamydomonas reinhardtii flagella. Journal of Cell Biology, 179(3):501–??, November 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/179/3/501.

Hill:2009:NSN

[HdM09] Emma Hill, Aimee deCathelineau, and Tom Misteli. New science — new ways to communicate it. Journal of Cell Biology, 185(2):183–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/2/183.

Huang:2009:UCS

[HDR09] Kaiyao Huang, Dennis R. Diener, and Joel L. Rosenbaum. The ubiquitin conjugation system is involved in the disassembly of cilia and flagella. Journal of Cell Biology, 186(4):601–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
186/4/601.

Hochegger:2007:ERC

[HDS^{+}07] Helfrid Hochegger, Donniphat Dejsuphong, Eiichiro Sonoda, Alihossein Saberi, Eeson Rajendra, Jane Kirk, Tim Hunt, and Shunichi Takeda. An essential role for Cdk1 in S phase control is revealed via chemical genetics in vertebrate cells. Journal of Cell Biology, 178(2):257–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/178/2/257.

Heuser:2005:DEE

[Heu05] John Heuser. Deep-etch EM reveals that the early poxvirus envelope is a single membrane bilayer stabilized by a geodetic “honeycomb” surface coat. Journal of Cell Biology, 169 (2):269–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/2/269.

Heiman:2007:GRP

[HEW07] Maxwell G. Heiman, Alex Engel, and Peter Walter. The Golgiresident protease Kex2 acts in conjunction with Prm1 to facilitate cell fusion during yeast mating. Journal of Cell Biology, 176(2):209–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/2/209.

Hughes:2006:PAT

[HF06] Sarah C. Hughes and Richard G. Fehon. Phosphorylation and activity of the tumor suppressor Merlin and the ERM protein Moesin are coordinately regulated by the Slik kinase. Journal of Cell Biology, 175(2):305–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/175/2/305.

Huang:2007:THR

[HFF^{+}07a] Haomin Huang, Jie Feng, Jakub Famulski, Jerome B. Rattner, Song Tao Liu, Gary D. Kao, Ruth Muschel, Gordon K. T. Chan, and Tim J. Yen. Tripin/ hSgo2 recruits MCAK to the inner centromere to correct defective kinetochore attachments. Journal of Cell Biology, 177(3):413–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/177/3/413.

Huang:2007:VKK

[HFF07b] Jinhong Huang, Asako Furuya, and Teiichi Furuichi. Very-KIND, a KIND domain–containing RasGEF, controls dendrite growth by linking Ras small GTPases and MAP2. Journal of Cell Biology, 179(3):539–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/3/539.

He:2005:RCD

[HFM^{+}05] Yan He, Franto Francis, Kenneth A. Myers, Wenqian Yu, Mark M. Black, and Peter W. Baas. Role of cytoplasmic dynein in the axonal transport of microtubules and neurofilaments. Journal of Cell Biology, 168(5):697–??, February 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/168/5/697.

Hasdemir:2005:TKK

[HFP^{+}05] Burcu Hasdemir, Daniel J. Fitzgerald, Ian A. Prior, Alexei V. Tepikin, and Robert D. Burgoyne. Traffic of Kv4 K^{+} channels mediated by KChIP1 is via a novel post–ER vesicular pathway. Journal of Cell Biology, 171(3):459–??, November 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/171/3/459.

Hendriksen:2005:REN

[HFvdV^{+}05] Jolita Hendriksen, Francois Fagotto, Hella van der Velde, Martijn van Schie, Jasprien Noordermeer, and Maarten Fornerod. RanBP3 enhances nuclear export of active β-catenin independently of CRM1. Journal of Cell Biology, 171(5):785–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
171/5/785.

Hata:2006:RIP

[HFY^{+}06] Katsuhiko Hata, Masashi Fujitani, Yuichi Yasuda, Hideo Doya, Tomoko Saito, Satoru Yamagishi, Bernhard K. Mueller, and Toshihide Yamashita. RGMa inhibition promotes axonal growth and recovery after spinal cord injury. Journal of Cell Biology, 173(1):47–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/1/47.

Huber:2007:SWN

[HG07] Michael D. Huber and Larry Gerace. The size-wise nucleus: nuclear volume control in eukaryotes. Journal of Cell Biology, 179(4):583–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/4/583.

Head:2009:IPI

[HGA^{+}09] Brian Head, Lorena Griparic, Mandana Amiri, Shilpa Gandre-Babbe, and Alexander M. van der Bliek. Inducible proteolytic inactivation of OPA1 mediated by the OMA1 protease in mammalian cells. Journal of Cell Biology, 187(7):959–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/187/7/
959.

Hajra:2006:CSH

[HGJ06] Sujata Hajra, Santanu Kumar Ghosh, and Makkuni Jayaram. The centromere-specific histone variant Cse4p (CENP–A) is essential for functional chromatin architecture at the yeast 2-μ m circle partitioning locus and promotes equal plasmid segregation. Journal of Cell Biology, 174(6):779–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/174/6/779.

Hinoi:2008:STM

[HGJ^{+}08] Eiichi Hinoi, Nan Gao, Dae Young Jung, Vijay Yadav, Tatsuya Yoshizawa, Martin G. Myers, Streamson C. Chua, Jason K. Kim, Klaus H. Kaestner, and Gerard Karsenty. The sympathetic tone mediates leptin’s inhibition of insulin secretion by modulating osteocalcin bioactivity. Journal of Cell Biology, 183(7):1235–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/7/1235.

Hoffman:2006:BAS

[HGK^{+}06] Lisa M. Hoffman, Kamal Garcha, Konstantina Karamboulas, Matthew F. Cowan, Linsay M. Drysdale, William A. Horton, and T. Michael Underhill. BMP action in skeletogenesis involves attenuation of retinoid signaling. Journal of Cell Biology, 174(1):101–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/1/101.

Hanono:2006:ERM

[HGR^{+}06] Abraham Hanono, Damien Garbett, David Reczek, David N. Chambers, and Anthony Bretscher. EPI64 regulates microvillar subdomains and structure. Journal of Cell Biology, 175(5): 803–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/5/803.

Hannak:2006:XCL

[HH06a] Eva Hannak and Rebecca Heald. Xorbit/CLASP links dynamic microtubules to chromosomes in the Xenopus meiotic spindle. Journal of Cell Biology, 172(1):19–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/172/1/19.

Huang:2006:DME

[HH06b] Baoying Huang and Tim C. Huffaker. Dynamic microtubules are essential for efficient chromosome capture and biorientation in S. cerevisiae. Journal of Cell Biology, 175(1):17–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/175/1/
17.

Horiguchi:2006:TPG

[HHFC06] Kaori Horiguchi, Toshihiko Hanada, Yasuhisa Fukui, and Athar H. Chishti. Transport of PIP_{3} by GAKIN, a kinesin-3 family protein, regulates neuronal cell polarity. Journal of Cell Biology, 174(3):425–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/3/425.

Hoppins:2009:MOI

[HHS^{+}09] Suzanne Hoppins, Jennifer Horner, Cheng Song, J. Michael Mc-Caffery, and Jodi Nunnari. Mitochondrial outer and inner membrane fusion requires a modified carrier protein. Journal of Cell Biology, 184(4):569–??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/184/4/569.

Han:2007:STT

[HHT^{+}07] Jianzhong Han, Liang Han, Priyanka Tiwari, Zhexing Wen, and James Q. Zheng. Spatial targeting of type II protein kinase A to filopodia mediates the regulation of growth cone guidance by cAMP. Journal of Cell Biology, 176(1):101–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/176/1/101.

Hirano:2008:EFP

[HHY^{+}08] Mariko Hirano, Shigeru Hashimoto, Shigenobu Yonemura, Hisataka Sabe, and Shinichi Aizawa. EPB41L5 functions to post-transcriptionally regulate cadherin and integrin during epithelial–mesenchymal transition. Journal of Cell Biology, 182(6):1217–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/6/1217.

Huang:2008:PSB

[HHZ^{+}08] Haomin Huang, James Hittle, Francesca Zappacosta, Roland S. Annan, Avram Hershko, and Timothy J. Yen. Phosphorylation sites in BubR1 that regulate kinetochore attachment, tension, and mitotic exit. Journal of Cell Biology, 183(4): 667–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/4/667.

Hieda:2008:MAG

[HIK^{+}08] Miki Hieda, Mayumi Isokane, Michiko Koizumi, Chiduru Higashi, Taro Tachibana, Masachika Shudou, Tomohiko Taguchi, Yohki Hieda, and Shigeki Higashiyama. Membrane-anchored growth factor, HB–EGF, on the cell surface targeted to the inner nuclear membrane. Journal of Cell Biology, 180(4): 763–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/4/763.

Hill:2007:JCA

[Hil07] Emma Hill. JCB content automatically deposited in PubMed Central (PMC). Journal of Cell Biology, 179(1):9–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/1/
9.

Hill:2008:AJD

[Hil08] Emma Hill. Announcing the JCB DataViewer, a browser-based application for viewing original image files. Journal of Cell Biology, 183(6):969–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/6/969.

Han:2006:STS

[HJ06] Xue Han and Meyer B. Jackson. Structural transitions in the synaptic SNARE complex during Ca^{2+}-triggered exocytosis. Journal of Cell Biology, 172(2):281–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/172/2/281.

Hammarlund:2007:ABA

[HJB07] Marc Hammarlund, Erik M. Jorgensen, and Michael J. Bastiani. Axons break in animals lacking β-spectrin. Journal of Cell Biology, 176(3):269–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/176/3/269.

Hoffman:2006:GAZ

[HJK^{+}06] Laura M. Hoffman, Christopher C. Jensen, Susanne Kloeker, C.-L. Albert Wang, Masaaki Yoshigi, and Mary C. Beckerle. Genetic ablation of zyxin causes Mena/VASP mislocalization, increased motility, and deficits in actin remodeling. Journal of Cell Biology, 172(5):771–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/172/5/771.

Howard:2005:COS

[HK05] Martin Howard and Karsten Kruse. Cellular organization by self-organization. Journal of Cell Biology, 168(4):533– ??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/4/533.

Hesselson:2006:GCE

[HK06] Daniel Hesselson and Judith Kimble. Growth control by EGF repeats of the C. elegans Fibulin-1C isoform. Journal of Cell Biology, 175(2):217–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/2/217.

Herrmann:2007:CMI

[HK07] Johannes M. Herrmann and Roman Köhl. Catch me if you can! Oxidative protein trapping in the intermembrane space of mitochondria. Journal of Cell Biology, 176(5):559–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/176/5/
559.

Hegde:2008:CTR

[HK08] Ramanujan S. Hegde and Sang-Wook Kang. The concept of translocational regulation. Journal of Cell Biology, 182(2):225– ??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
182/2/225.

Kim:2008:TDC

[hKDH^{+}08] Sahn ho Kim, Albert R. Davalos, Seok-Jin Heo, Francis Rodier, Ying Zou, Christian Beausejour, Patrick Kaminker, Steven M. Yannone, and Judith Campisi. Telomere dysfunction and cell survival: roles for distinct TIN2-containing complexes. Journal of Cell Biology, 181(3):447–??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/181/3/447.

Hata:2009:UAL

[HKIY09] Katsuhiko Hata, Kozo Kaibuchi, Shinobu Inagaki, and Toshihide Yamashita. Unc5B associates with LARG to mediate the action of repulsive guidance molecule. Journal of Cell Biology, 184(5):737–??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/184/5/737.

Holway:2006:CSD

[HKLM06] Antonia H. Holway, Seung-Hwan Kim, Adriana La Volpe, and W. Matthew Michael. Checkpoint silencing during the DNA damage response in Caenorhabditis elegans embryos. Journal of Cell Biology, 172(7):999–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/172/7/999.

Holway:2007:CSD

[HKLM07] Antonia H. Holway, Seung-Hwan Kim, Adriana La Volpe, and W. Matthew Michael. Checkpoint silencing during the DNA damage response in Caenorhabditis elegans embryos. Journal of Cell Biology, 178(5):887–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/178/5/887.

Hayashi:2009:DGR

[HKN09] Yoshiki Hayashi, Satoru Kobayashi, and Hiroshi Nakato. Drosophila glypicans regulate the germline stem cell niche. Journal of Cell Biology, 187(4):473–??, November 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/187/4/473.

Hu:2009:LSC

[HKP^{+}09] Yan Hu, Igor Kireev, Matt Plutz, Nazanin Ashourian, and Andrew S. Belmont. Large-scale chromatin structure of inducible genes: transcription on a condensed, linear template. Journal of Cell Biology, 185(1):87–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/185/1/87.

Hess:2005:QEM

[HKV^{+}05] Samuel T. Hess, Mukesh Kumar, Anil Verma, Jane Farrington, Anne Kenworthy, and Joshua Zimmerberg. Quantitative electron microscopy and fluorescence spectroscopy of the membrane distribution of influenza hemagglutinin. Journal of Cell Biology, 169(6):965–??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/6/965.

Hotulainen:2006:SFG

[HL06] Pirta Hotulainen and Pekka Lappalainen. Stress fibers are generated by two distinct actin assembly mechanisms in motile cells. Journal of Cell Biology, 173(3):383–??, May 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/173/3/383.

Hou:2006:FCI

[HLA^{+}06] Rong Hou, Liming Liu, Syed Anees, Shungo Hiroyasu, and Nicholas E. S. Sibinga. The Fat1 cadherin integrates vascular smooth muscle cell growth and migration signals. Journal of Cell Biology, 173(3):417–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/173/3/417.

Howell:2007:ARG

[HLJ^{+}07] Gareth R. Howell, Richard T. Libby, Tatjana C. Jakobs, Richard S. Smith, F. Campbell Phalan, Joseph W. Barter, Jessica M. Barbay, Jeffrey K. Marchant, Nagaraju Mahesh, Vittorio Porciatti, Alan V. Whitmore, Richard H. Masland, and Simon W. M. John. Axons of retinal ganglion cells are insulted in the optic nerve early in DBA/2J glaucoma. Journal of Cell Biology, 179(7):1523–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/7/1523.

Holterman:2007:MRP

[HLK^{+}07] Chet E. Holterman, Fabien Le Grand, Shihuan Kuang, Patrick Seale, and Michael A. Rudnicki. Megf10 regulates the progression of the satellite cell myogenic program. Journal of Cell Biology, 179(5):911–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/5/911.

Hao:2009:PCM

[HLK^{+}09] Jian-Jiang Hao, Yin Liu, Michael Kruhlak, Karen E. Debell, Barbara L. Rellahan, and Stephen Shaw. Phospholipase C-mediated hydrolysis of PIP2 releases ERM proteins from lymphocyte membrane. Journal of Cell Biology, 184(3): 451–??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/184/3/451.

Han:2008:CKI

[HLL^{+}08] Xiao-Jian Han, Yun-Fei Lu, Shun-Ai Li, Taku Kaitsuka, Yasufumi Sato, Kazuhito Tomizawa, Angus C. Nairn, Kohji Takei, Hideki Matsui, and Masayuki Matsushita. CaM kinase Iαinduced phosphorylation of Drp1 regulates mitochondrial morphology. Journal of Cell Biology, 182(3):573–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/182/3/573.

Hollien:2009:RID

[HLL^{+}09] Julie Hollien, Jonathan H. Lin, Han Li, Nicole Stevens, Peter Walter, and Jonathan S. Weissman. Regulated Ire1-dependent decay of messenger RNAs in mammalian cells. Journal of Cell Biology, 186(3):323–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/186/3/323.

Ha:2008:NSC

[HLM^{+}08] Junghoon Ha, Kevin W.-H. Lo, Kenneth R. Myers, Tiffany M. Carr, Michael K. Humsi, Bareza A. Rasoul, Rosalind A. Segal, and K. Kevin Pfister. A neuron-specific cytoplasmic dynein isoform preferentially transports TrkB signaling endosomes. Journal of Cell Biology, 181(6):1027–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/181/6/1027.

Huckaba:2006:RTI

[HLP06] Thomas M. Huckaba, Thomas Lipkin, and Liza A. Pon. Roles of type II myosin and a tropomyosin isoform in retrograde actin flow in budding yeast. Journal of Cell Biology, 175(6): 957–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/6/957.

Hoshino:2006:TRI

[HlQY^{+}06] Masataka Hoshino, Mei ling Qi, Natsue Yoshimura, Tomoyuki Miyashita, Kazuhiko Tagawa, Yo ichi Wada, Yasushi Enokido, Shigeki Marubuchi, Phoebe Harjes, Nobutaka Arai, Kiyomitsu Oyanagi, Giovanni Blandino, Marius Sudol, Tina Rich, Ichiro Kanazawa, Erich E. Wanker, Minoru Saitoe, and Hitoshi Okazawa. Transcriptional repression induces a slowly progressive atypical neuronal death associated with changes of YAP isoforms and p73. Journal of Cell Biology, 172(4): 589–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/4/589.

Hotulainen:2009:DMA

[HLS^{+}09] Pirta Hotulainen, Olaya Llano, Sergei Smirnov, Kimmo Tanhuanpää, Jan Faix, Claudio Rivera, and Pekka Lappalainen. Defining mechanisms of actin polymerization and depolymerization during dendritic spine morphogenesis. Journal of Cell Biology, 185(2):323–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/2/323.

Hao:2008:RCB

[HM08] Yi Hao and Ian G. Macara. Regulation of chromatin binding by a conformational switch in the tail of the Ran exchange factor RCC1. Journal of Cell Biology, 182(5):827–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/182/5/
827.

Hebbar:2008:LNI

[HMG^{+}08] Sachin Hebbar, Mariano T. Mesngon, Aimee M. Guillotte, Bhavim Desai, Ramses Ayala, and Deanna S. Smith. Lis1 and Ndel1 influence the timing of nuclear envelope breakdown in neural stem cells. Journal of Cell Biology, 182(6): 1063–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/6/1063.

Hishikawa:2005:MME

[HMM^{+}05] Keiichi Hishikawa, Takeshi Marumo, Shigeki Miura, Asato Nakanishi, Yumi Matsuzaki, Katsunori Shibata, Tomoko Ichiyanagi, Hiroko Kohike, Takuya Komori, Ichiro Takahashi, Osamu Takase, Naohiko Imai, Masahiro Yoshikawa, Toshihiko Inowa, Matsuhiko Hayashi, Toshio Nakaki, Hiromitsu Nakauchi, Hideyuki Okano, and Toshiro Fujita. Musculin/ MyoR is expressed in kidney side population cells and can regulate their function. Journal of Cell Biology, 169(6):921–??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/169/6/
921.

He:2007:APA

[HMR^{+}07] Jiuya He, Chih-Chieh Mao, Aurelio Reyes, Hiroshi Sembongi, Miriam Di Re, Caroline Granycome, Andrew B. Clippingdale, Ian M. Fearnley, Michael Harbour, Alan J. Robinson, Stefanie Reichelt, Johannes N. Spelbrink, John E. Walker, and Ian J. Holt. The AAA^{+} protein ATAD3 has displacement loop binding properties and is involved in mitochondrial nucleoid organization. Journal of Cell Biology, 176(2):141–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/176/2/141.

Harvey:2008:FRT

[HMS^{+}08] Kieran F. Harvey, Jaakko Mattila, Avi Sofer, F. Christian Bennett, Matthew R. Ramsey, Leif W. Ellisen, Oscar Puig, and Iswar K. Hariharan. FOXO-regulated transcription restricts overgrowth of Tsc mutant organs. Journal of Cell Biology, 180 (4):691–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/4/691.

Halter:2007:PPG

[HNvD^{+}07] David Halter, Sylvia Neumann, Suzanne M. van Dijk, Jasja Wolthoorn, Ann M. de Mazière, Otilia V. Vieira, Peter Mattjus, Judith Klumperman, Gerrit van Meer, and Hein Sprong. Preand post-Golgi translocation of glucosylceramide in glycosphingolipid synthesis. Journal of Cell Biology, 179(1):101–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/1/

Hickson:2008:RDC

[HO08] Gilles R. X. Hickson and Patrick H. O’Farrell. Rho-dependent control of anillin behavior during cytokinesis. Journal of Cell Biology, 180(2):285–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/2/285.

Hirai:2007:FDE

[HOH^{+}07] Maretoshi Hirai, Tetsuya Ohbayashi, Masahito Horiguchi, Katsuya Okawa, Akari Hagiwara, Kenneth R. Chien, Toru Kita, and Tomoyuki Nakamura. Fibulin-5/ DANCE has an elastogenic organizer activity that is abrogated by proteolytic cleavage in vivo. Journal of Cell Biology, 176(7):1061–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/176/7/
1061.

Hashimoto:2008:IAT

[HOH^{+}08] Daisuke Hashimoto, Masaki Ohmuraya, Masahiko Hirota, Akitsugu Yamamoto, Koichi Suyama, Satoshi Ida, Yuushi Okumura, Etsuhisa Takahashi, Hiroshi Kido, Kimi Araki, Hideo Baba, Noboru Mizushima, and Ken ichi Yamamura. Involvement of autophagy in trypsinogen activation within the pancreatic acinar cells. Journal of Cell Biology, 181(7):1065–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/181/7/
1065.

Hamelers:2005:RAT

[HOM^{+}05] Irene H. L. Hamelers, Cristina Olivo, Alexander E. E. Mertens, D. Michiel Pegtel, Rob A. van der Kammen, Arnoud Sonnenberg, and John G. Collard. The Rac activator Tiam1 is required for α3β1-mediated laminin-5 deposition, cell spreading, and cell migration. Journal of Cell Biology, 171(5):871–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/171/5/
871.

Hedstrom:2008:ARM

[HOR08] Kristian L. Hedstrom, Yasuhiro Ogawa, and Matthew N. Rasband. AnkyrinG is required for maintenance of the axon initial segment and neuronal polarity. Journal of Cell Biology, 183(4): 635–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/4/635.

Horsley:2009:EWS

[Hor09] Valerie Horsley. Epigenetics, Wnt signaling, and stem cells: the Pygo2 connection. Journal of Cell Biology, 185(5):761–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/185/5/
761.

Hurt:2009:CCT

[HOZ^{+}09] Jessica A. Hurt, Robert A. Obar, Bo Zhai, Natalie G. Farny, Steven P. Gygi, and Pamela A. Silver. A conserved CCCH-type zinc finger protein regulates mRNA nuclear adenylation and export. Journal of Cell Biology, 185(2):265–??, April 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/185/2/265.

Harris:2005:PSA

[HP05] Tony J. C. Harris and Mark Peifer. The positioning and segregation of apical cues during epithelial polarity establishment in Drosophila. Journal of Cell Biology, 170(5):813–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/170/5/
813.

Hur:2005:JMI

[HPH^{+}05] Eun-Mi Hur, Yong-Soo Park, Yang Hoon Huh, Seung Hyun Yoo, Kyung-Chul Woo, Bo-Hwa Choi, and Kyong-Tai Kim. Junctional membrane inositol 1,4,5-trisphosphate receptor complex coordinates sensitization of the silent EGF-induced Ca^{2+} signaling. Journal of Cell Biology, 169(4):657–??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/169/4/657.

Habib:2007:MSB

[HPT^{+}07] Tania Habib, Heon Park, Mark Tsang, Ignacio Moreno de Alborán, Andrea Nicks, Leslie Wilson, Paul S. Knoepfler, Sarah Andrews, David J. Rawlings, Robert N. Eisenman, and Brian M. Iritani. Myc stimulates B lymphocyte differentiation and amplifies calcium signaling. Journal of Cell Biology, 179 (4):717–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/4/717.

Hou:2007:FAI

[HQF^{+}07] Yuqing Hou, Hongmin Qin, John A. Follit, Gregory J. Pazour, Joel L. Rosenbaum, and George B. Witman. Functional analysis of an individual IFT protein: IFT46 is required for transport of outer dynein arms into flagella. Journal of Cell Biology, 176(5):653–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/5/653.

Heinrich:2005:GNC

[HR05] Reinhart Heinrich and Tom A. Rapoport. Generation of nonidentical compartments in vesicular transport systems. Journal of Cell Biology, 168(2):271–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/168/2/271.

Hill:2008:YWI

[HR08] Emma Hill and Mike Rossner. You wrote it; you own it! Journal of Cell Biology, 181(3):405–??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/181/3/405.

Hurley:2009:CCF

[HR09] James H. Hurley and Xuefeng Ren. The circuitry of cargo flux in the ESCRT pathway. Journal of Cell Biology, 185(2):185–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
185/2/185.

Haren:2006:NDR

[HRB^{+}06] Laurence Haren, Marie-Hélène Remy, Ingrid Bazin, Isabelle Callebaut, Michel Wright, and Andreas Merdes. NEDD1dependent recruitment of the γ-tubulin ring complex to the centrosome is necessary for centriole duplication and spindle assembly. Journal of Cell Biology, 172(4):505–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/172/4/505.

Heuser:2009:DRC

[HRK^{+}09] Thomas Heuser, Milen Raytchev, Jeremy Krell, Mary E. Porter, and Daniela Nicastro. The dynein regulatory complex is the nexin link and a major regulatory node in cilia and flagella.

Journal of Cell Biology, 187(6):921–??, December 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/187/6/921.

Hanson:2008:PMD

[HRLH08] Phyllis I. Hanson, Robyn Roth, Yuan Lin, and John E. Heuser. Plasma membrane deformation by circular arrays of ESCRT– III protein filaments. Journal of Cell Biology, 180(2):389–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
180/2/389.

Holzel:2005:MWN

[HRS^{+}05] Michael Hölzel, Michaela Rohrmoser, Martin Schlee, Thomas Grimm, Thomas Harasim, Anastassia Malamoussi, Anita Gruber-Eber, Elisabeth Kremmer, Wolfgang Hiddemann, Georg W. Bornkamm, and Dirk Eick. Mammalian WDR12 is a novel member of the Pes1–Bop1 complex and is required for ribosome biogenesis and cell proliferation. Journal of Cell Biology, 170(3):367–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/3/367.

Haynes:2009:ESB

[HS09a] Karmella A. Haynes and Pamela A. Silver. Eukaryotic systems broaden the scope of synthetic biology. Journal of Cell Biology, 187(5):589–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/5/589.

Hsiao:2009:STR

[HS09b] Susan J. Hsiao and Susan Smith. Sister telomeres rendered dysfunctional by persistent cohesion are fused by NHEJ. Journal of Cell Biology, 184(4):515–??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/184/4/515.

Hisata:2007:RPG

[HSB^{+}07] Shu Hisata, Toshiaki Sakisaka, Takeshi Baba, Tomohiro Yamada, Kazuhiro Aoki, Michiyuki Matsuda, and Yoshimi Takai. Rap1-PDZ-GEF1 interacts with a neurotrophin receptor at late endosomes, leading to sustained activation of Rap1 and ERK and neurite outgrowth. Journal of Cell Biology, 178(5):843–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
178/5/843.

Hudmon:2005:CTT

[HSK^{+}05] Andy Hudmon, Howard Schulman, James Kim, Janet M. Maltez, Richard W. Tsien, and Geoffrey S. Pitt. CaMKII tethers to L-type Ca^{2+} channels, establishing a local and dedicated integrator of Ca^{2+} signals for facilitation. Journal of Cell Biology, 171(3):537–??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/3/537.

Hammond:2009:RBR

[HSLI09] Gerald R. V. Hammond, Yirong Sim, Leon Lagnado, and Robin F. Irvine. Reversible binding and rapid diffusion of proteins in complex with inositol lipids serves to coordinate free movement with spatial information. Journal of Cell Biology, 184(2):297–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/2/297.

Hua:2006:VIA

[HSTM06] Wei Hua, David Sheff, Derek Toomre, and Ira Mellman. Vectorial insertion of apical and basolateral membrane proteins in polarized epithelial cells revealed by quantitative 3D live cell imaging. Journal of Cell Biology, 172(7):1035–??, March 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/172/7/1035.

He:2006:RAP

[HSY^{+}06] Congcong He, Hui Song, Tomohiro Yorimitsu, Iryna Monastyrska, Wei-Lien Yen, Julie E. Legakis, and Daniel J. Klionsky. Recruitment of Atg9 to the preautophagosomal structure by Atg11 is essential for selective autophagy in budding yeast. Journal of Cell Biology, 175(6):925–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/175/6/925.

Harris:2008:CPP

[HT08] Kathryn P. Harris and Ulrich Tepass. Cdc42 and Par proteins stabilize dynamic adherens junctions in the Drosophila neuroectoderm through regulation of apical endocytosis. Journal of Cell Biology, 183(6):1129–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/183/6/1129.

Hara:2008:FUI

[HTK^{+}08] Taichi Hara, Akito Takamura, Chieko Kishi, Shun ichiro Iemura, Tohru Natsume, Jun-Lin Guan, and Noboru Mizushima. FIP200, a ULK-interacting protein, is required for autophagosome formation in mammalian cells. Journal of Cell Biology, 181(3):497–??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/3/497.

Hienola:2006:SDI

[HTKR06] Anni Hienola, Sarka Tumova, Evgeny Kulesskiy, and Heikki Rauvala. N-syndecan deficiency impairs neural migration in brain. Journal of Cell Biology, 174(4):569–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/174/4/569.

Hayashi-Takanaka:2009:VHM

[HTYNK09] Yoko Hayashi-Takanaka, Kazuo Yamagata, Naohito Nozaki, and Hiroshi Kimura. Visualizing histone modifications in living cells: spatiotemporal dynamics of H3 phosphorylation during interphase. Journal of Cell Biology, 187(6):781–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/187/6/
781.

Hill:2009:SCM

[HW09] Emma Hill and Ruth Williams. Super-coil me: Sizing up centromeric nucleosomes. Journal of Cell Biology, 186(4):453–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
186/4/453.

Huang:2008:ICR

[HWJ^{+}08] XiaoDong Huang, Diana Wu, Hua Jin, Dwayne Stupack, and Jean Y. J. Wang. Induction of cell retraction by the combined actions of Abl–CrkII and Rho–ROCK1 signaling. Journal of Cell Biology, 183(4):711–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/183/4/711.

Habib:2007:TDT

[HWN^{+}07] Shukry J. Habib, Thomas Waizenegger, Agathe Niewienda, Stefan A. Paschen, Walter Neupert, and Doron Rapaport. The N-terminal domain of Tob55 has a receptor-like function in the biogenesis of mitochondrial β-barrel proteins. Journal of Cell Biology, 176(1):77–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/1/77.

Hemmerich:2008:DIK

[HWPH^{+}08] Peter Hemmerich, Stefanie Weidtkamp-Peters, Christian Hoischen, Lars Schmiedeberg, Indri Erliandri, and Stephan Diekmann. Dynamics of inner kinetochore assembly and maintenance in living cells. Journal of Cell Biology, 180(6):1101–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
180/6/1101.

Humphries:2007:VCF

[HWS^{+}07] Jonathan D. Humphries, Pengbo Wang, Charles Streuli, Benny Geiger, Martin J. Humphries, and Christoph Ballestrem. Vinculin controls focal adhesion formation by direct interactions with talin and actin. Journal of Cell Biology, 179(5):1043– ??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/5/1043.

Hammarlund:2008:CSD

[HWSJ08] Marc Hammarlund, Shigeki Watanabe, Kim Schuske, and Erik M. Jorgensen. CAPS and syntaxin dock dense core vesicles to the plasma membrane in neurons. Journal of Cell Biology, 180(3):483–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/3/483.

Hedstrom:2007:NAS

[HXO^{+}07] Kristian L. Hedstrom, Xiaorong Xu, Yasuhiro Ogawa, Renato Frischknecht, Constanze I. Seidenbecher, Peter Shrager, and Matthew N. Rasband. Neurofascin assembles a specialized extracellular matrix at the axon initial segment. Journal of Cell Biology, 178(5):875–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/5/875.

He:2007:EMS

[HXZ^{+}07] Bing He, Fengong Xi, Jian Zhang, Daniel TerBush, Xiaoyu Zhang, and Wei Guo. Exo70p mediates the secretion of specific exocytic vesicles at early stages of the cell cycle for polarized cell growth. Journal of Cell Biology, 176(6):771–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/176/6/
771.

Huang:2008:ANA

[HYB08] Yinyi Huang, Hongyan Yan, and Mohan K. Balasubramanian. Assembly of normal actomyosin rings in the absence of Mid1p and cortical nodes in fission yeast. Journal of Cell Biology, 183 (6):979–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/6/979.

He:2007:DTN

[HZL^{+}07] Ping He, Zhenyu Zhong, Kristina Lindholm, Lilian Berning, Wendy Lee, Cynthia Lemere, Matthias Staufenbiel, Rena Li, and Yong Shen. Deletion of tumor necrosis factor death receptor inhibits amyloid β generation and prevents learning and memory deficits in Alzheimer’s mice. Journal of Cell Biology, 178(5):829–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/5/829.

Houben:2005:EEN

[HZOL05] Edith N. G. Houben, Raz Zarivach, Bauke Oudega, and Joen Luirink. Early encounters of a nascent membrane protein. Journal of Cell Biology, 170(1):27–??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/170/1/27.

Huang:2005:ORC

[HZR05] Zhen Huang, Keling Zang, and Louis F. Reichardt. The origin recognition core complex regulates dendrite and spine development in postmitotic neurons. Journal of Cell Biology, 170 (4):527–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/4/527.

Inoue:2005:CMP

[IBS^{+}05] Kentaro Inoue, Amy J. Baldwin, Rebecca L. Shipman, Kyoko Matsui, Steven M. Theg, and Masaru Ohme-Takagi. Complete maturation of the plastid protein translocation channel requires a type I signal peptidase. Journal of Cell Biology, 171(3): 425–??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/3/425.

Inoue:2009:SAP

[IDTT^{+}09] Eiji Inoue, Maki Deguchi-Tawarada, Aki Togawa, Chiyuki Matsui, Kohei Arita, Sayaka Katahira-Tayama, Toshitaka Sato, Emiko Yamauchi, Yoshiya Oda, and Yoshimi Takai. Synaptic activity prompts γ-secretase-mediated cleavage of EphA4 and dendritic spine formation. Journal of Cell Biology, 185(3):551– ??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
185/3/551.

Imanishi:2006:ARM

[IEGV06] Miki Imanishi, Nicholas F. Endres, Arne Gennerich, and Ronald D. Vale. Autoinhibition regulates the motility of the C. elegans intraflagellar transport motor OSM-3. Journal of Cell Biology, 174(7):931–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/174/7/931.

Ikenouchi:2005:TCN

[IFF^{+}05] Junichi Ikenouchi, Mikio Furuse, Kyoko Furuse, Hiroyuki Sasaki, Sachiko Tsukita, and Shoichiro Tsukita. Tricellulin constitutes a novel barrier at tricellular contacts of epithelial cells. Journal of Cell Biology, 171(6):939–??, December 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/171/6/939.

Idrissi:2008:DAM

[IGFG^{+}08] Fatima-Zahra Idrissi, Helga Grötsch, Isabel M. Fernández-Golbano, Cristina Presciatto-Baschong, Howard Riezman, and María-Isabel Geli. Distinct acto/myosin–I structures associate with endocytic profiles at the plasma membrane. Journal of Cell Biology, 180(6):1219–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/6/1219.

Ivanovska:2005:VAG

[IH05] Iva Ivanovska and J. Marie Hardwick. Viruses activate a genetically conserved cell death pathway in a unicellular organism. Journal of Cell Biology, 170(3):391–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/170/3/391.

Hiraga:2008:HHL

[iHBD08] Shin ichiro Hiraga, Sotirios Botsios, and Anne D. Donaldson. Histone H3 lysine 56 acetylation by Rtt109 is crucial for chromosome positioning. Journal of Cell Biology, 183(4): 641–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/4/641.

Ilani:2007:ISF

[IKZ^{+}07] Tal Ilani, Chand Khanna, Ming Zhou, Timothy D. Veenstra, and Anthony Bretscher. Immune synapse formation requires ZAP-70 recruitment by ezrin and CD43 removal by moesin. Journal of Cell Biology, 179(4):733–??, November 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/179/4/733.

Ishiuchi:2009:MFD

[IMY^{+}09] Takashi Ishiuchi, Kazuyo Misaki, Shigenobu Yonemura, Masatoshi Takeichi, and Takuji Tanoue. Mammalian Fat and Dachsous cadherins regulate apical membrane organization in the embryonic cerebral cortex. Journal of Cell Biology, 185 (6):959–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/6/959.

Izard:2006:SAM

[INB06] Tina Izard, Guy Tran Van Nhieu, and Philippe R. J. Bois. Shigella applies molecular mimicry to subvert vinculin and invade host cells. Journal of Cell Biology, 175(3):465–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/175/3/
465.

Ishibashi:2005:MIM

[IPAR05] Jeff Ishibashi, Robert L. Perry, Atsushi Asakura, and Michael A. Rudnicki. MyoD induces myogenic differentiation through cooperation of its NH_{2}- and COOH-terminal regions. Journal of Cell Biology, 171(3):471–??, November 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/171/3/471.

Ilieva:2009:NCA

[IPC09] Hristelina Ilieva, Magdalini Polymenidou, and Don W. Cleveland. Non-cell autonomous toxicity in neurodegenerative disorders: ALS and beyond. Journal of Cell Biology, 187(6): 761–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/187/6/761.

Irie:2005:DRA

[IPG^{+}05] Hanna Y. Irie, Rachel V. Pearline, Dorre Grueneberg, Maximilian Hsia, Preethi Ravichandran, Nayantara Kothari, Sridaran Natesan, and Joan S. Brugge. Distinct roles of Akt1 and Akt2 in regulating cell migration and epithelial–mesenchymal transition. Journal of Cell Biology, 171(6):1023–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/171/6/
1023.

Ingerman:2005:DFS

[IPM^{+}05] Elena Ingerman, Edward M. Perkins, Michael Marino, Jason A. Mears, J. Michael McCaffery, Jenny E. Hinshaw, and Jodi Nunnari. Dnm1 forms spirals that are structurally tailored to fit mitochondria. Journal of Cell Biology, 170(7):1021–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/170/7/
1021.

Idone:2008:RIP

[ITG^{+}08] Vincent Idone, Christina Tam, John W. Goss, Derek Toomre, Marc Pypaert, and Norma W. Andrews. Repair of injured plasma membrane by rapid Ca^{2+}-dependent endocytosis. Journal of Cell Biology, 180(5):905–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/180/5/905.

Ikenouchi:2007:RZF

[IUT^{+}07] Junichi Ikenouchi, Kazuaki Umeda, Sachiko Tsukita, Mikio Furuse, and Shoichiro Tsukita. Requirement of ZO-1 for the formation of belt-like adherens junctions during epithelial cell polarization. Journal of Cell Biology, 176(6):779–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/176/6/779.

Wakabayashi:2006:MCR

[iWK06] Ken ichi Wakabayashi and Stephen M. King. Modulation of Chlamydomonas reinhardtii flagellar motility by redox poise. Journal of Cell Biology, 173(5):743–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/173/5/743.

Yoshimura:2007:FDR

[iYEF^{+}07] Shin ichiro Yoshimura, Johannes Egerer, Evelyn Fuchs, Alexander K. Haas, and Francis A. Barr. Functional dissection of Rab GTPases involved in primary cilium formation. Journal of Cell Biology, 178(3):363–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/3/363.

Yamashita:2006:PSP

[iYOW^{+}06] Shun ichi Yamashita, Masahide Oku, Yuko Wasada, Yoshitaka Ano, and Yasuyoshi Sakai. PI4P-signaling pathway for the synthesis of a nascent membrane structure in selective autophagy. Journal of Cell Biology, 173(5):709–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/173/5/709.

Ju:2008:RPG

[JASC08] Tongzhong Ju, Rajindra P. Aryal, Caleb J. Stowell, and Richard D. Cummings. Regulation of protein O-glycosylation by the endoplasmic reticulum–localized molecular chaperone Cosmc. Journal of Cell Biology, 182(3):531–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/182/3/531.

Jablonska:2007:CCP

[JAV^{+}07] Beata Jablonska, Adan Aguirre, Renaud Vandenbosch, Shibeshih Belachew, Cyril Berthet, Philipp Kaldis, and Vittorio Gallo. Cdk2 is critical for proliferation and self-renewal of neural progenitor cells in the adult subventricular zone. Journal of Cell Biology, 179(6):1231–??, December 2007. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/6/1231.

Joglekar:2008:MAK

[JBF^{+}08] Ajit P. Joglekar, David Bouck, Ken Finley, Xingkun Liu, Yakun Wan, Judith Berman, Xiangwei He, E. D. Salmon, and Kerry S. Bloom. Molecular architecture of the kinetochore-microtubule attachment site is conserved between point and regional centromeres. Journal of Cell Biology, 181(4):587–??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/181/4/587.

Jansen:2007:PCC

[JBFC07] Lars E. T. Jansen, Ben E. Black, Daniel R. Foltz, and Don W. Cleveland. Propagation of centromeric chromatin requires exit from mitosis. Journal of Cell Biology, 176(6):795–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/176/6/
795.

Jablonka:2007:DCC

[JBL^{+}07] Sibylle Jablonka, Marcus Beck, Barbara Dorothea Lechner, Christine Mayer, and Michael Sendtner. Defective Ca^{2+} channel clustering in axon terminals disturbs excitability in motoneurons in spinal muscular atrophy. Journal of Cell Biology, 179 (1):139–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/1/139.

Jansen:2006:PCC

[JBYW06] Jaclyn M. Jansen, Margaret F. Barry, Charles K. Yoo, and Eric L. Weiss. Phosphoregulation of Cbk1 is critical for RAM network control of transcription and morphogenesis. Journal of Cell Biology, 175(5):755–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/175/5/755.

Jiang:2008:IBS

[JCMG08] Ying Jiang, Zhiliang Cheng, Elisabet C. Mandon, and Reid Gilmore. An interaction between the SRP receptor and the translocon is critical during cotranslational protein translocation. Journal of Cell Biology, 180(6):1149–??, March 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/180/6/1149.

Jaco:2008:CMR

[JCVB08] Isabel Jaco, Andrés Canela, Elsa Vera, and Maria A. Blasco. Centromere mitotic recombination in mammalian cells. Journal of Cell Biology, 181(6):885–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/181/6/885.

Janetopoulos:2006:PSP

[JD06] Chris Janetopoulos and Peter Devreotes. Phosphoinositide signaling plays a key role in cytokinesis. Journal of Cell Biology, 174(4):485–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/4/485.

Juliano:2005:EMG

[JDK^{+}05] Rudy L. Juliano, Vidula R. Dixit, Hyunmin Kang, Tai Young Kim, Yuko Miyamoto, and Dong Xu. Epigenetic manipulation of gene expression. Journal of Cell Biology, 169(6):847–??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/169/6/
847.

Jorgensen:2007:HMS

[JET^{+}07] Stine Jørgensen, Ingegerd Elvers, Morten Beck Trelle, Tobias Menzel, Morten Eskildsen, Ole Nørregaard Jensen, Thomas Helleday, Kristian Helin, and Claus Storgaard Sørensen. The histone methyltransferase SET8 is required for S-phase progression. Journal of Cell Biology, 179(7):1337–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/7/
1337.

Ju:2009:VCP

[JFM^{+}09] Jeong-Sun Ju, Rodrigo A. Fuentealba, Sara E. Miller, Erin Jackson, David Piwnica-Worms, Robert H. Baloh, and Conrad C. Weihl. Valosin-containing protein (VCP) is required for autophagy and is disrupted in VCP disease. Journal of Cell Biology, 187(6):875–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/6/875.

Jacob:2008:ELS

[JGAS08] Claire Jacob, Henrik Grabner, Suzana Atanasoski, and Ueli Suter. Expression and localization of Ski determine cell type– specific TGFβ signaling effects on the cell cycle. Journal of Cell Biology, 182(3):519–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/3/519.

Jehle:2006:ABC

[JGL^{+}06] Andreas W. Jehle, Shyra J. Gardai, Suzhao Li, Patrick Linsel-Nitschke, Konosuke Morimoto, William J. Janssen, R. William Vandivier, Nan Wang, Steven Greenberg, Benjamin M. Dale, Chunbo Qin, Peter M. Henson, and Alan R. Tall. ATP-binding cassette transporter A7 enhances phagocytosis of apoptotic cells and associated ERK signaling in macrophages. Journal of Cell Biology, 174(4):547–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/4/547.

Jin:2009:PRH

[JGY09] Hui Jin, Vincent Guacci, and Hong-Guo Yu. Pds5 is required for homologue pairing and inhibits synapsis of sister chromatids during yeast meiosis. Journal of Cell Biology, 186(5): 713–??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/186/5/713.

Jwa:2008:RSI

[JhKC08] Miri Jwa, Jae hyun Kim, and Clarence S. M. Chan. Regulation of Sli15/INCENP, kinetochore, and Cdc14 phosphatase functions by the ribosome biogenesis protein Utp7. Journal of Cell Biology, 182(6):1099–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/182/6/1099.

Juhasz:2008:CIP

[JHY^{+}08] Gábor Juhász, Jahda H. Hill, Ying Yan, Miklós Sass, Eric H. Baehrecke, Jonathan M. Backer, and Thomas P. Neufeld. The class III PI(3)K Vps34 promotes autophagy and endocytosis but not TOR signaling in Drosophila. Journal of Cell Biology, 181(4):655–??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/4/655.

Jakobsson:2007:BBV

[JKCW07] Lars Jakobsson, Johan Kreuger, and Lena Claesson-Welsh. Building blood vessels — stem cell models in vascular biology. Journal of Cell Biology, 177(5):751–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/177/5/751.

Jaffe:2008:CCS

[JKDH08] Aron B. Jaffe, Noriko Kaji, Joanne Durgan, and Alan Hall. Cdc42 controls spindle orientation to position the apical surface during epithelial morphogenesis. Journal of Cell Biology, 183(4):625–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/4/625.

James:2008:PBR

[JKKM08] Declan J. James, Chuenchanok Khodthong, Judith A. Kowalchyk, and Thomas F. J. Martin. Phosphatidylinositol 4,5-bisphosphate regulates SNARE-dependent membrane fusion. Journal of Cell Biology, 182(2):355–??, July 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/182/2/355.

Johswich:2009:GTD

[JKW^{+}09] Anita Johswich, Benjamin Kraft, Manfred Wuhrer, Monika Berger, André M. Deelder, Cornelis H. Hokke, Rita Gerardy-Schahn, and Hans Bakker. Golgi targeting of Drosophila melanogaster β4GalNAcTB requires a DHHC protein family– related protein as a pilot. Journal of Cell Biology, 184(1): 173–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/184/1/173.

Jakobs:2005:RGC

[JLB^{+}05] Tatjana C. Jakobs, Richard T. Libby, Yixin Ben, Simon W. M. John, and Richard H. Masland. Retinal ganglion cell degeneration is topological but not cell type specific in DBA/2J mice. Journal of Cell Biology, 171(2):313–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/171/2/313.

Jang:2008:LIA

[JLC08] Cholsoon Jang, Gina Lee, and Jongkyeong Chung. LKB1 induces apical trafficking of Silnoon, a monocarboxylate transporter, in Drosophila melanogaster. Journal of Cell Biology, 183 (1):11–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/1/11.

Jones:2006:CDI

[JLG^{+}06] Robert G. Jones, Xiufen Li, Phillip D. Gray, Jinqiu Kuang, Frederic Clayton, Wade S. Samowitz, Blair B. Madison, Deborah L. Gumucio, and Scott K. Kuwada. Conditional deletion of β1 integrins in the intestinal epithelium causes a loss of Hedgehog expression, intestinal hyperplasia, and early postnatal lethality. Journal of Cell Biology, 175(3):505–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/175/3/
505.

Jeon:2007:RRA

[JLL^{+}07] Taeck J. Jeon, Dai-Jen Lee, Susan Lee, Gerald Weeks, and Richard A. Firtel. Regulation of Rap1 activity by RapGAP1 controls cell adhesion at the front of chemotaxing cells. Journal of Cell Biology, 179(5):833–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/5/833.

Jeon:2007:RCC

[JLM^{+}07] Taeck J. Jeon, Dai-Jen Lee, Sylvain Merlot, Gerald Weeks, and Richard A. Firtel. Rap1 controls cell adhesion and cell motility through the regulation of myosin II. Journal of Cell Biology, 176(7):1021–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/7/1021.

Jeganathan:2007:BMC

[JMB^{+}07] Karthik Jeganathan, Liviu Malureanu, Darren J. Baker, Susan C. Abraham, and Jan M. van Deursen. Bub1 mediates cell death in response to chromosome missegregation and acts to suppress spontaneous tumorigenesis. Journal of Cell Biology, 179(2):255–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/2/255.

Jaspersen:2006:SUH

[JMG^{+}06] Sue L. Jaspersen, Adriana E. Martin, Galina Glazko, Thomas H. Giddings, Garry Morgan, Arcady Mushegian, and Mark Winey. The Sad1-UNC-84 homology domain in Mps3 interacts with Mps2 to connect the spindle pole body with the nuclear envelope. Journal of Cell Biology, 174(5):665–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/174/5/
665.

Jin:2008:EPP

[JMG08] Hui Jin, J. Michael McCaffery, and Eric Grote. Ergosterol promotes pheromone signaling and plasma membrane fusion in mating yeast. Journal of Cell Biology, 180(4):813–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/180/4/
813.

Janer:2006:PCP

[JMM^{+}06] Alexandre Janer, Elodie Martin, Marie-Paule Muriel, Morwena Latouche, Hiroto Fujigasaki, Merle Ruberg, Alexis Brice, Yvon Trottier, and Annie Sittler. PML clastosomes prevent nuclear accumulation of mutant ataxin-7 and other polyglutamine proteins. Journal of Cell Biology, 174(1):65–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/174/1/65.

Jauregui:2008:CEN

[JNHB08] Andrew R. Jauregui, Ken C. Q. Nguyen, David H. Hall, and Maureen M. Barr. The Caenorhabditis elegans nephrocystins act as global modifiers of cilium structure. Journal of Cell Biology, 180(5):973–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/5/973.

Johnson:2009:SFC

[Joh09] Arthur E. Johnson. The structural and functional coupling of two molecular machines, the ribosome and the translocon. Journal of Cell Biology, 185(5):765–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/185/5/765.

Jansen:2006:MRR

[JP06] Katie M. Jansen and Grace K. Pavlath. Mannose receptor regulates myoblast motility and muscle growth. Journal of Cell Biology, 174(3):403–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/3/403.

Johansson:2007:AED

[JRZ^{+}07] Marie Johansson, Nuno Rocha, Wilbert Zwart, Ingrid Jordens, Lennert Janssen, Coenraad Kuijl, Vesa M. Olkkonen, and Jacques Neefjes. Activation of endosomal dynein motors by stepwise assembly of Rab7–RILP–p150^{Glued}, ORP1L, and the receptor β lll spectrin. Journal of Cell Biology, 176(4): 459–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/4/459.

Jani:2007:ZRA

[JS07] Klodiana Jani and Frieder Schöck. Zasp is required for the assembly of functional integrin adhesion sites. Journal of Cell Biology, 179(7):1583–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/7/1583.

Johnson:2008:DCC

[JSC08] Ruth I. Johnson, Midori J. Seppa, and Ross L. Cagan. The Drosophila CD2AP/CIN85 orthologue Cindr regulates junctions and cytoskeleton dynamics during tissue patterning. Journal of Cell Biology, 180(6):1191–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/180/6/1191.

Jonassen:2008:DIM

[JSFP08] Julie A. Jonassen, Jovenal San Agustin, John A. Follit, and Gregory J. Pazour. Deletion of IFT20 in the mouse kidney causes misorientation of the mitotic spindle and cystic kidney disease. Journal of Cell Biology, 183(3):377–??, November 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/183/3/377.

Janson:2005:EFB

[JSPT05] Marcel E. Janson, Thanuja Gangi Setty, Anne Paoletti, and P. T. Tran. Efficient formation of bipolar microtubule bundles requires microtubule-bound γ-tubulin complexes. Journal of Cell Biology, 169(2):297–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/169/2/297.

Jung:2009:DTV

[JTH09] Goeh Jung, Margaret A. Titus, and John A. Hammer. The Dictyostelium type V myosin MyoJ is responsible for the cortical association and motility of contractile vacuole membranes. Journal of Cell Biology, 186(4):555–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/186/4/555.

Jones:2005:PMF

[JTN^{+}05] Nathan C. Jones, Kristina J. Tyner, Lisa Nibarger, Heather M. Stanley, Dawn D. W. Cornelison, Yuri V. Fedorov, and Bradley B. Olwin. The p38α /β MAPK functions as a molecular switch to activate the quiescent satellite cell. Journal of Cell Biology, 169(1):105–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/1/105.

Jang:2008:DRM

[JWC^{+}08] Chang-Young Jang, Jim Wong, Judith A. Coppinger, Akiko Seki, John R. Yates, and Guowei Fang. DDA3 recruits microtubule depolymerase Kif2a to spindle poles and controls spindle dynamics and mitotic chromosome movement. Journal of Cell Biology, 181(2):255–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/2/255.

Jung:2007:MBS

[JWG^{+}07] Nadja Jung, Martin Wienisch, Mingyu Gu, James B. Rand, Sebastian L. Müller, Gerd Krause, Erik M. Jorgensen, Jürgen Klingauf, and Volker Haucke. Molecular basis of synaptic vesicle cargo recognition by the endocytic sorting adaptor stonin 2. Journal of Cell Biology, 179(7):1497–??, December 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/179/7/1497.

Juschke:2005:SNI

[JWSS05] Christoph Jüschke, Andrea Wächter, Blanche Schwappach, and Matthias Seedorf. SEC18 /nsf-independent, protein-sorting pathway from the yeast cortical ER to the plasma membrane. Journal of Cell Biology, 169(4):613–??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/169/4/613.

Je:2006:NIS

[JYZL06] Hyun-Soo Je, Feng Yang, Jiangzheng Zhou, and Bai Lu. Neurotrophin 3 induces structural and functional modification of synapses through distinct molecular mechanisms. Journal of Cell Biology, 175(6):1029–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/175/6/1029.

Jiang:2009:IAL

[JZN^{+}09] Wei-Qin Jiang, Ze-Huai Zhong, Akira Nguyen, Jeremy D. Henson, Christian D. Toouli, Antony W. Braithwaite, and Roger R. Reddel. Induction of alternative lengthening of telomeresassociated PML bodies by p53/p21 requires HP1 proteins. Journal of Cell Biology, 185(5):797–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/185/5/797.

Kiprilov:2008:HES

[KAD^{+}08] Enko N. Kiprilov, Aashir Awan, Romain Desprat, Michelle Velho, Christian A. Clement, Anne Grete Byskov, Claus Y. Andersen, Peter Satir, Eric E. Bouhassira, Søren T. Christensen, and Rhoda Elison Hirsch. Human embryonic stem cells in culture possess primary cilia with hedgehog signaling machinery. Journal of Cell Biology, 180(5):897–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/180/5/897.

Kaksonen:2008:TAE

[Kak08] Marko Kaksonen. Taking apart the endocytic machinery. Journal of Cell Biology, 180(6):1059–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/180/6/1059.

Kim:2006:PRC

[KAL^{+}06] Ikjin Kim, Jungmi Ahn, Chang Liu, Kaori Tanabe, Jennifer Apodaca, Tadashi Suzuki, and Hai Rao. The Png1–Rad23 complex regulates glycoprotein turnover. Journal of Cell Biology, 172(2):211–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/2/211.

Katsnelson:2006:AC

[Kat06a] Alla Katsnelson. Aging on the clock. Journal of Cell Biology, 172(3):??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/3/327b.

Katsnelson:2006:BCS

[Kat06b] Alla Katsnelson. B cells spread to affinity. Journal of Cell Biology, 173(5):642–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/5/642.

Katsnelson:2006:CCF

[Kat06c] Alla Katsnelson. Cell clusters follow the leaders. Journal of Cell Biology, 173(5):??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/5/642a.

Katsnelson:2006:CSA

[Kat06d] Alla Katsnelson. E. coli squeezed into action. Journal of Cell Biology, 172(3):??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/3/327a.

Katsnelson:2006:CCH

[Kat06e] Alla Katsnelson. ES cells clean house. Journal of Cell Biology, 173(5):??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/5/643b.

Katsnelson:2006:GP

[Kat06f] Alla Katsnelson. Genome punctuation. Journal of Cell Biology, 173(5):??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/5/643a.

Katsnelson:2006:HPK

[Kat06g] Alla Katsnelson. Hyperdynamic proteins keep stem cells on their toes. Journal of Cell Biology, 172(3):326–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/172/3/
326.

Katsnelson:2006:MHW

[Kat06h] Alla Katsnelson. Making HoxD waves. Journal of Cell Biology, 172(3):??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/3/326a.

Katsnelson:2006:TN

[Kat06i] Alla Katsnelson. Time out for neurons. Journal of Cell Biology, 173(5):643–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/5/643.

Katsnelson:2006:TSC

[Kat06j] Alla Katsnelson. Two-speed cell specification. Journal of Cell Biology, 172(3):327–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/3/327.

Katsnelson:2007:CYY

[Kat07a] Alla Katsnelson. Checkpoint yin and yang. Journal of Cell Biology, 177(4):??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/4/569a.

Katsnelson:2007:LLD

[Kat07b] Alla Katsnelson. Limiting layer diffusion. Journal of Cell Biology, 177(4):??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/4/569b.

Katsnelson:2007:MRM

[Kat07c] Alla Katsnelson. Membrane recycling for mitosis. Journal of Cell Biology, 177(4):??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/4/568a.

Katsnelson:2007:MME

[Kat07d] Alla Katsnelson. MicroRNAs with macro effects. Journal of Cell Biology, 177(4):568–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/4/568.

Katsnelson:2007:TWCa

[Kat07e] Alla Katsnelson. Two ways to condense. Journal of Cell Biology, 177(4):569–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/4/569.

Katsnelson:2007:TWCb

[Kat07f] Alla Katsnelson. Two ways to condense. Journal of Cell Biology, 178(4):713–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/4/713.

Kelleher:2007:DLO

[KBC^{+}07a] Daniel J. Kelleher, Sulagna Banerjee, Anthony J. Cura, John Samuelson, and Reid Gilmore. Dolichol-linked oligosaccharide selection by the oligosaccharyltransferase in protist and fungal organisms. Journal of Cell Biology, 177(1):29–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/177/1/29.

Kumar:2007:SSU

[KBC^{+}07b] K. G. Suresh Kumar, Hervé Barriere, Christopher J. Carbone, Jianghuai Liu, Gayathri Swaminathan, Ping Xu, Ying Li, Darren P. Baker, Junmin Peng, Gergely L. Lukacs, and Serge Y. Fuchs. Site-specific ubiquitination exposes a linear motif to promote interferon-α receptor endocytosis. Journal of Cell Biology, 179(5):935–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/5/935.

Kumari:2008:NAR

[KBC^{+}08] Sudha Kumari, Virginia Borroni, Ashutosh Chaudhry, Baron Chanda, Ramiro Massol, Satyajit Mayor, and Francisco J. Barrantes. Nicotinic acetylcholine receptor is internalized via a Racdependent, dynamin-independent endocytic pathway. Journal of Cell Biology, 181(7):1179–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/181/7/1179.

Kottgen:2008:TTF

[KBGG^{+}08] Michael Köttgen, Björn Buchholz, Miguel A. Garcia-Gonzalez, Fruzsina Kotsis, Xiao Fu, Mara Doerken, Christopher Boehlke, Daniel Steffl, Robert Tauber, Tomasz Wegierski, Roland Nitschke, Makoto Suzuki, Albrecht Kramer-Zucker, Gregory G. Germino, Terry Watnick, Jean Prenen, Bernd Nilius, E. Wolfgang Kuehn, and Gerd Walz. TRPP2 and TRPV4 form a polymodal sensory channel complex. Journal of Cell Biology, 182 (3):437–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/3/437.

Kahn:2008:CNH

[KBI^{+}08] Richard A. Kahn, Elspeth Bruford, Hiroki Inoue, John M. Logsdon, Zhongzhen Nie, Richard T. Premont, Paul A. Randazzo, Masanobu Satake, Anne B. Theibert, Maria L. Zapp, and Dan Cassel. Consensus nomenclature for the human ArfGAP domain-containing proteins. Journal of Cell Biology, 182(6):1039–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/6/1039.

Khaled:2005:CDC

[KBK^{+}05] Annette R. Khaled, Dmitry V. Bulavin, Christina Kittipatarin, Wen Qing Li, Michelle Alvarez, Kyungjae Kim, Howard A. Young, Albert J. Fornace, and Scott K. Durum. Cytokinedriven cell cycling is mediated through Cdc25A. Journal of Cell Biology, 169(5):755–??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/5/755.

Kawashima:2006:RGA

[KBN^{+}06] Toshiyuki Kawashima, Ying Chun Bao, Yasushi Nomura, Yuseok Moon, Yukio Tonozuka, Yukinori Minoshima, Tomonori Hatori, Akiho Tsuchiya, Mari Kiyono, Tetsuya Nosaka, Hideaki Nakajima, David A. Williams, and Toshio Kitamura. Rac1 and a GTPase-activating protein, MgcRacGAP, are required for nuclear translocation of STAT transcription factors. Journal of Cell Biology, 175(6):937–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/175/6/937.

Kriebel:2008:CCM

[KBR^{+}08] Paul W. Kriebel, Valarie A. Barr, Erin C. Rericha, Guofeng Zhang, and Carole A. Parent. Collective cell migration requires vesicular trafficking for chemoattractant delivery at the trailing edge. Journal of Cell Biology, 183(5):949–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/183/5/949.

Kaiserman:2006:MHM

[KBS^{+}06] Dion Kaiserman, Catherina H. Bird, Jiuru Sun, Antony Matthews, Kheng Ung, James C. Whisstock, Philip E. Thompson, Joseph A. Trapani, and Phillip I. Bird. The major human and mouse granzymes are structurally and functionally divergent. Journal of Cell Biology, 175(4):619–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/175/4/619.

Kearns:2006:PNF

[KBW^{+}06] Jeffrey D. Kearns, Soumen Basak, Shannon L. Werner, Christine S. Huang, and Alexander Hoffmann. IκBϵ provides negative feedback to control NF-κB oscillations, signaling dynamics, and inflammatory gene expression. Journal of Cell Biology, 173(5):659–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/5/659.

Kang:2008:CBC

[KBY^{+}08] Jong-Sun Kang, Gyu-Un Bae, Min-Jeong Yi, Youn-Joo Yang, Ji-Eun Oh, Giichi Takaesu, Yi Ting Zhou, Boon Chuan Low, and Robert S. Krauss. A Cdo–Bnip–2–Cdc42 signaling pathway regulates p38α /β MAPK activity and myogenic differentiation. Journal of Cell Biology, 182(3):497–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/182/3/497.

Kellner:2009:KPS

[KC09] Juliane C. Kellner and Pierre A. Coulombe. Keratins and protein synthesis: the plot thickens. Journal of Cell Biology, 187 (2):157–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/187/2/157.

Karren:2005:RFM

[KCAS05] Mary Anne Karren, Emily M. Coonrod, Teresa K. Anderson, and Janet M. Shaw. The role of Fis1p–Mdv1p interactions in mitochondrial fission complex assembly. Journal of Cell Biology, 171(2):291–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/2/291.

Knights:2006:DPA

[KCD^{+}06a] Chad D. Knights, Jason Catania, Simone Di Giovanni, Selen Muratoglu, Ricardo Perez, Amber Swartzbeck, Andrew A. Quong, Xiaojing Zhang, Terry Beerman, Richard G. Pestell, and Maria Laura Avantaggiati. Distinct p53 acetylation cassettes differentially influence gene-expression patterns and cell fate. Journal of Cell Biology, 173(4):533–??, May 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/173/4/533.

Kruhlak:2006:CCS

[KCD^{+}06b] Michael J. Kruhlak, Arkady Celeste, Graham Dellaire, Oscar Fernandez-Capetillo, Waltraud G. Müller, James G. McNally, David P. Bazett-Jones, and André Nussenzweig. Changes in chromatin structure and mobility in living cells at sites of DNA double-strand breaks. Journal of Cell Biology, 172(6):823–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
172/6/823.

Kahn:2006:NHA

[KCE^{+}06] Richard A. Kahn, Jacqueline Cherfils, Marek Elias, Ruth C. Lovering, Sean Munro, and Annette Schurmann. Nomenclature for the human Arf family of GTP-binding proteins: ARF, ARL, and SAR proteins. Journal of Cell Biology, 172(5): 645–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/5/645.

Kline:2006:HMC

[KCH^{+}06] Susan L. Kline, Iain M. Cheeseman, Tetsuya Hori, Tatsuo Fukagawa, and Arshad Desai. The human Mis12 complex is required for kinetochore assembly and proper chromosome segregation. Journal of Cell Biology, 173(1):9–??, April 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/173/1/9.

Kueh:2008:ADC

[KCMB08] Hao Yuan Kueh, Guillaume T. Charras, Timothy J. Mitchison, and William M. Brieher. Actin disassembly by cofilin, coronin, and Aip1 occurs in bursts and is inhibited by barbed-end cappers. Journal of Cell Biology, 182(2):341–??, July 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/182/2/341.

Krzewski:2006:FWW

[KCOS06] Konrad Krzewski, Xi Chen, Jordan S. Orange, and Jack L. Strominger. Formation of a WIP-, WASp-, actin-, and myosin IIA–containing multiprotein complex in activated NK cells and its alteration by KIR inhibitory signaling. Journal of Cell Biology, 173(1):121–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/1/121.

Kuang:2006:DRP

[KCS^{+}06] Shihuan Kuang, Sophie B. Chargé, Patrick Seale, Michael Huh, and Michael A. Rudnicki. Distinct roles for Pax7 and Pax3 in adult regenerative myogenesis. Journal of Cell Biology, 172 (1):103–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/1/103.

Kranz:2008:BTD

[KDD08] Dominique Kranz, Christoph Dohmesen, and Matthias Dobbelstein. BRCA1 and Tip60 determine the cellular response to ultraviolet irradiation through distinct pathways. Journal of Cell Biology, 182(1):197–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/1/197.

Kyei:2009:API

[KDD^{+}09] George B. Kyei, Christina Dinkins, Alexander S. Davis, Esteban Roberts, Sudha B. Singh, Chunsheng Dong, Li Wu, Eiki Kominami, Takashi Ueno, Akitsugu Yamamoto, Maurizio Federico, Antonito Panganiban, Isabelle Vergne, and Vojo Deretic. Autophagy pathway intersects with HIV-1 biosynthesis and regulates viral yields in macrophages. Journal of Cell Biology, 186(2):255–??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/186/2/255.

Komarova:2009:MEB

[KDG^{+}09] Yulia Komarova, Christian O. De Groot, Ilya Grigoriev, Susana Montenegro Gouveia, E. Laura Munteanu, Joseph M. Schober, Srinivas Honnappa, Rubén M. Buey, Casper C. Hoogenraad, Marileen Dogterom, Gary G. Borisy, Michel O. Steinmetz, and Anna Akhmanova. Mammalian end binding proteins control persistent microtubule growth. Journal of Cell Biology, 184(5):691–??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/5/691.

Kaizuka:2009:CCU

[KDV^{+}09] Yoshihisa Kaizuka, Adam D. Douglass, Santosh Vardhana, Michael L. Dustin, and Ronald D. Vale. The coreceptor CD2 uses plasma membrane microdomains to transduce signals in T cells. Journal of Cell Biology, 185(3):521–??, May 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/185/3/521.

Klemm:2009:SSS

[KES^{+}09] Robin W. Klemm, Christer S. Ejsing, Michal A. Surma, Hermann-Josef Kaiser, Mathias J. Gerl, Julio L. Sampaio, Quentin de Robillard, Charles Ferguson, Tomasz J. Proszynski, Andrej Shevchenko, and Kai Simons. Segregation of sphingolipids and sterols during formation of secretory vesicles at the trans-Golgi network. Journal of Cell Biology, 185(4):601–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
185/4/601.

Kirkham:2005:UIU

[KFC^{+}05] Matthew Kirkham, Akikazu Fujita, Rahul Chadda, Susan J. Nixon, Teymuras V. Kurzchalia, Deepak K. Sharma, Richard E. Pagano, John F. Hancock, Satyajit Mayor, and Robert G. Parton. Ultrastructural identification of uncoated caveolinindependent early endocytic vehicles. Journal of Cell Biology, 168(3):465–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/3/465.

Kose:2005:KHS

[KFK^{+}05] Shingo Kose, Maiko Furuta, Makiko Koike, Yoshihiro Yoneda, and Naoko Imamoto. The 70-kD heat shock cognate protein (hsc70) facilitates the nuclear export of the import receptors. Journal of Cell Biology, 171(1):19–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/171/1/19.

Konieczny:2008:MID

[KFR^{+}08] Patryk Konieczny, Peter Fuchs, Siegfried Reipert, Wolfram S. Kunz, Anikó Zeöld, Irmgard Fischer, Denise Paulin, Rolf Schröder, and Gerhard Wiche. Myofiber integrity depends on desmin network targeting to Z-disks and costameres via distinct plectin isoforms. Journal of Cell Biology, 181(4):667–??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/181/4/
667.

Kidane:2005:DFR

[KG05] Dawit Kidane and Peter L. Graumann. Dynamic formation of RecA filaments at DNA double strand break repair centers in live cells. Journal of Cell Biology, 170(3):357–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/170/3/
357.

Kierbel:2007:PAE

[KGDR^{+}07] Arlinet Kierbel, Ama Gassama-Diagne, Claudia Rocha, Lilliana Radoshevich, Joan Olson, Keith Mostov, and Joanne Engel. Pseudomonas aeruginosa exploits a PIP3-dependent pathway to transform apical into basolateral membrane. Journal of Cell Biology, 177(1):21–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/1/21.

Kubitscheck:2005:NTS

[KGH^{+}05] Ulrich Kubitscheck, David Grünwald, Andreas Hoekstra, Daniel Rohleder, Thorsten Kues, Jan Peter Siebrasse, and Reiner Peters. Nuclear transport of single molecules. Journal of Cell Biology, 168(2):233–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/168/2/233.

Keizer-Gunnink:2007:CCA

[KGKV07] Ineke Keizer-Gunnink, Arjan Kortholt, and Peter J. M. Van Haastert. Chemoattractants and chemorepellents act by inducing opposite polarity in phospholipase C and PI3-kinase signaling. Journal of Cell Biology, 177(4):579–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/177/4/579.

Kota:2007:MCS

[KGL07] Jhansi Kota, C. Fredrik Gilstring, and Per O. Ljungdahl. Membrane chaperone Shr3 assists in folding amino acid permeases preventing precocious ERAD. Journal of Cell Biology, 176(5): 617–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/5/617.

Kutik:2007:CTC

[KGM^{+}07] Stephan Kutik, Bernard Guiard, Helmut E. Meyer, Nils Wiedemann, and Nikolaus Pfanner. Cooperation of translocase complexes in mitochondrial protein import. Journal of Cell Biology, 179(4):585–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/4/585.

Kong:2009:DCB

[KGM^{+}09] Fang Kong, Andrés J. García, A. Paul Mould, Martin J. Humphries, and Cheng Zhu. Demonstration of catch bonds between an integrin and its ligand. Journal of Cell Biology, 185 (7):1275–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/7/1275.

Kieran:2005:MDR

[KHB^{+}05] Dairin Kieran, Majid Hafezparast, Stephanie Bohnert, James R. T. Dick, Joanne Martin, Giampietro Schiavo, Elizabeth M. C. Fisher, and Linda Greensmith. A mutation in dynein rescues axonal transport defects and extends the life span of ALS mice. Journal of Cell Biology, 169(4):561–??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/169/4/561.

Kiesler:2005:HHM

[KHBV05] Eva Kiesler, Manuela E. Hase, David Brodin, and Neus Visa. Hrp59, an hnRNP M protein in Chironomus and Drosophila, binds to exonic splicing enhancers and is required for expression of a subset of mRNAs. Journal of Cell Biology, 168(7):1013–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
168/7/1013.

Kim:2008:RCF

[KHH08] Gun-Hwa Kim, Jung-Hyun Her, and Jin-Kwan Han. Ryk cooperates with Frizzled 7 to promote Wnt11-mediated endocytosis and is essential for Xenopus laevis convergent extension movements. Journal of Cell Biology, 182(6):1073–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/182/6/
1073.

Kwong:2007:MRC

[KHSM07] Jennifer Q. Kwong, Matthew S. Henning, Anatoly A. Starkov, and Giovanni Manfredi. The mitochondrial respiratory chain is a modulator of apoptosis. Journal of Cell Biology, 179(6): 1163–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/6/1163.

Kim:2007:SPM

[KHW^{+}07] Seung-Hwan Kim, Antonia H. Holway, Suzanne Wolff, Andrew Dillin, and W. Matthew Michael. SMK-1/ PPH-4.1-mediated silencing of the CHK-1 response to DNA damage in early C. elegans embryos. Journal of Cell Biology, 179(1):41–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/1/
41.

Kim:2009:SPM

[KHW^{+}09] Seung-Hwan Kim, Antonia H. Holway, Suzanne Wolff, Andrew Dillin, and W. Matthew Michael. SMK-1/ PPH-4.1-mediated silencing of the CHK-1 response to DNA damage in early C. elegans embryos. Journal of Cell Biology, 184(4):613–??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/184/4/

Kim:2008:CCS

[KHWC08] Yumi Kim, John E. Heuser, Clare M. Waterman, and Don W. Cleveland. CENP-e combines a slow, processive motor and a flexible coiled coil to produce an essential motile kinetochore tether. Journal of Cell Biology, 181(3):411–??, May 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/181/3/411.

Kawahara:2008:NRB

[KII^{+}08] Hironori Kawahara, Takao Imai, Hiroaki Imataka, Masafumi Tsujimoto, Ken Matsumoto, and Hideyuki Okano. Neural RNA-binding protein Musashi1 inhibits translation initiation by competing with eIF4G for PABP. Journal of Cell Biology, 181(4):639–??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/4/639.

Kimata:2007:TRS

[KIKI^{+}07] Yukio Kimata, Yuki Ishiwata-Kimata, Tatsuhiko Ito, Aiko Hirata, Tomohide Suzuki, Daisuke Oikawa, Masato Takeuchi, and Kenji Kohno. Two regulatory steps of ER-stress sensor Ire1 involving its cluster formation and interaction with unfolded proteins. Journal of Cell Biology, 179(1):75–??, October 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/179/1/75.

King:2006:TDM

[KKB^{+}06] Michelle E. King, Ho-Man Kan, Peter W. Baas, Alev Erisir, Charles G. Glabe, and George S. Bloom. Tau-dependent microtubule disassembly initiated by prefibrillar β-amyloid. Journal of Cell Biology, 175(4):541–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/175/4/541.

Kim:2007:BSE

[KKC^{+}07] Jinoh Kim, Bertrand Kleizen, Regina Choy, Gopal Thinakaran, Sangram S. Sisodia, and Randy W. Schekman. Biogenesis of γ-secretase early in the secretory pathway. Journal of Cell Biology, 179(5):951–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/5/951.

Kim:2007:HRH

[KKD^{+}07] Jung-Ae Kim, Michael Kruhlak, Farokh Dotiwala, André Nussenzweig, and James E. Haber. Heterochromatin is refractory to γ-H2AX modification in yeast and mammals. Journal of Cell Biology, 178(2):209–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/178/2/209.

Kishi:2005:L

[KKES05] Masashi Kishi, Terrance T. Kummer, Stephen J. Eglen, and Joshua R. Sanes. LL5β. Journal of Cell Biology, 169(2):355–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
169/2/355.

Kateneva:2005:RPT

[KKGD05] Anna V. Kateneva, Anton A. Konovchenko, Vincent Guacci, and Michael E. Dresser. Recombination protein Tid1p controls resolution of cohesin-dependent linkages in meiosis in Saccharomyces cerevisiae. Journal of Cell Biology, 171(2):241–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
171/2/241.

Kamimura:2006:SFR

[KKH^{+}06] Keisuke Kamimura, Takashi Koyama, Hiroko Habuchi, Ryu Ueda, Masayuki Masu, Koji Kimata, and Hiroshi Nakato. Specific and flexible roles of heparan sulfate modifications in Drosophila FGF signaling. Journal of Cell Biology, 174(6): 773–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/6/773.

Kim:2005:ISR

[KKK^{+}05] Jong-Soo Kim, Tatiana B. Krasieva, Hitoshi Kurumizaka, David J. Chen, A. Malcolm R. Taylor, and Kyoko Yokomori. Independent and sequential recruitment of NHEJ and HR factors to DNA damage sites in mammalian cells. Journal of Cell Biology, 170(3):341–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/3/341.

Kasprowicz:2008:ICH

[KKM^{+}08] Jaroslaw Kasprowicz, Sabine Kuenen, Katarzyna Miskiewicz, Ron L. P. Habets, Liesbet Smitz, and Patrik Verstreken. Inactivation of clathrin heavy chain inhibits synaptic recycling but allows bulk membrane uptake. Journal of Cell Biology, 182(5): 1007–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/5/1007.

Klebig:2009:BRC

[KKM09a] Christiane Klebig, Dirk Korinth, and Patrick Meraldi. Bub1 regulates chromosome segregation in a kinetochore-independent manner. Journal of Cell Biology, 185(5):841–??, June 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/185/5/841.

Koto:2009:TRD

[KKM09b] Akiko Koto, Erina Kuranaga, and Masayuki Miura. Temporal regulation of Drosophila IAP1 determines caspase functions in sensory organ development. Journal of Cell Biology, 187 (2):219–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/187/2/219.

Konig:2005:TAT

[KKRP05] Hans-Georg König, Donat Kögel, Abdelhaq Rami, and Jochen H. M. Prehn. TGF-β1 activates two distinct type I receptors in neurons. Journal of Cell Biology, 168(7):1077–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/168/7/
1077.

Kops:2005:ZLM

[KKW^{+}05] Geert J. P. L. Kops, Yumi Kim, Beth A. A. Weaver, Yinghui Mao, Ian McLeod, John R. Yates, Mitsuo Tagaya, and Don W. Cleveland. ZW10 links mitotic checkpoint signaling to the structural kinetochore. Journal of Cell Biology, 169(1):49–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
169/1/49.

Koh:2007:EDC

[KKW^{+}07] Tong-Wey Koh, Viktor I. Korolchuk, Yogesh P. Wairkar, Wei Jiao, Emma Evergren, Hongling Pan, Yi Zhou, Koen J. T. Venken, Oleg Shupliakov, Iain M. Robinson, Cahir J. O’Kane, and Hugo J. Bellen. Eps15 and Dap160 control synaptic vesicle membrane retrieval and synapse development. Journal of Cell Biology, 178(2):309–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/2/309.

Kapitein:2008:MCL

[KKW^{+}08] Lukas C. Kapitein, Benjamin H. Kwok, Joshua S. Weinger, Christoph F. Schmidt, Tarun M. Kapoor, and Erwin J. G. Peterman. Microtubule cross-linking triggers the directional motility of kinesin-5. Journal of Cell Biology, 182(3):421–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
182/3/421.

Kim:2009:IDC

[KKW^{+}09] Young Kim, Matthias C. Kugler, Ying Wei, Kevin K. Kim, Xiaopeng Li, Alexis N. Brumwell, and Harold A. Chapman. Integrin α3β1-dependent β-catenin phosphorylation links epithelial Smad signaling to cell contacts. Journal of Cell Biology, 184(2):309–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/2/309.

Kim:2005:NFB

[KKY^{+}05] Hong-Kyung Kim, Jeong-Gu Kang, Shigehiko Yumura, Charles J. Walsh, Jin Won Cho, and JooHun Lee. De novo formation of basal bodies in Naegleria gruberi. Journal of Cell Biology, 169(5):719–??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/5/719.

Kota:2005:SML

[KL05] Jhansi Kota and Per O. Ljungdahl. Specialized membranelocalized chaperones prevent aggregation of polytopic proteins in the ER. Journal of Cell Biology, 168(1):79–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/168/1/79.

Kumar:2009:GPM

[KLG^{+}09] Praveen Kumar, Karen S. Lyle, Sarah Gierke, Alexandre Matov, Gaudenz Danuser, and Torsten Wittmann. GSK3β phosphorylation modulates CLASP–microtubule association and lamella microtubule attachment. Journal of Cell Biology, 184 (6):895–??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/184/6/895.

Kim:2007:MBD

[KLR^{+}07] Hwajin Kim, Shuo-Chien Ling, Gregory C. Rogers, Comert Kural, Paul R. Selvin, Stephen L. Rogers, and Vladimir I. Gelfand. Microtubule binding by dynactin is required for microtubule organization but not cargo transport. Journal of Cell Biology, 176(5):641–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/5/641.

Khmelinskii:2007:CRM

[KLRS07] Anton Khmelinskii, Clare Lawrence, Johanna Roostalu, and Elmar Schiebel. Cdc14-regulated midzone assembly controls anaphase B. Journal of Cell Biology, 177(6):981–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/177/6/981.

Kim:2006:MSM

[KLS^{+}06] Hak Kyun Kim, Yong Sun Lee, Umasundari Sivaprasad, Ankit Malhotra, and Anindya Dutta. Muscle-specific microRNA miR-206 promotes muscle differentiation. Journal of Cell Biology, 174(5):677–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/5/677.

Kowanetz:2008:TIS

[KLV^{+}08] Marcin Kowanetz, Peter Lönn, Michael Vanlandewijck, Katarzyna Kowanetz, Carl-Henrik Heldin, and Aristidis Moustakas. TGFβ induces SIK to negatively regulate type I receptor kinase signaling. Journal of Cell Biology, 182(4):655–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/182/4/

Kleber:2005:NCS

[KLW^{+}05] Maurice Kléber, Hye-Youn Lee, Heiko Wurdak, Johanna Buchstaller, Martin M. Riccomagno, Lars M. Ittner, Ueli Suter, Douglas J. Epstein, and Lukas Sommer. Neural crest stem cell maintenance by combinatorial Wnt and BMP signaling. Journal of Cell Biology, 169(2):309–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/169/2/309.

Karpowicz:2005:C

[KMK^{+}05a] Phillip Karpowicz, Cindi Morshead, Angela Kam, Eric Jervis, John Ramunas, Vincent Cheng, and Derek van der Kooy. Correction. Journal of Cell Biology, 170(7):1169–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/170/7/
1169.

Karpowicz:2005:SIS

[KMK^{+}05b] Phillip Karpowicz, Cindi Morshead, Angela Kam, Eric Jervis, John Ramunas, Vincent Cheng, and Derek van der Kooy. Support for the immortal strand hypothesis: neural stem cells partition DNA asymmetrically in vitro. Journal of Cell Biology, 170 (5):721–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/5/721.

Kida:2007:TTH

[KMS07] Yuichiro Kida, Fumiko Morimoto, and Masao Sakaguchi. Two translocating hydrophilic segments of a nascent chain span the ER membrane during multispanning protein topogenesis. Journal of Cell Biology, 179(7):1441–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/7/1441.

Kim:2006:OMM

[KMSLS06] Peter K. Kim, Robert T. Mullen, Uwe Schumann, and Jennifer Lippincott-Schwartz. The origin and maintenance of mammalian peroxisomes involves a de novo PEX16-dependent pathway from the ER. Journal of Cell Biology, 173(4):521–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/173/4/

Katagiri:2006:SSC

[KNKH06] Chika Katagiri, Jotaro Nakanishi, Kuniko Kadoya, and Toshihiko Hibino. Serpin squamous cell carcinoma antigen inhibits UV-induced apoptosis via suppression of c–JUN NH_{2}-terminal kinase. Journal of Cell Biology, 172(7):983–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/172/7/983.

Kinoshita:2005:APT

[KNP^{+}05] Kazuhisa Kinoshita, Tim L. Noetzel, Laurence Pelletier, Karl Mechtler, David N. Drechsel, Anne Schwager, Mike Lee, Jordan W. Raff, and Anthony A. Hyman. Aurora A phosphorylation of TACC3/maskin is required for centrosome-dependent microtubule assembly in mitosis. Journal of Cell Biology, 170(7):1047–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/7/1047.

Kunisaki:2006:DRA

[KNT^{+}06] Yuya Kunisaki, Akihiko Nishikimi, Yoshihiko Tanaka, Ryosuke Takii, Mayuko Noda, Ayumi Inayoshi, Ken ichi Watanabe, Fumiyuki Sanematsu, Takehiko Sasazuki, Takehiko Sasaki, and Yoshinori Fukui. DOCK2 is a Rac activator that regulates motility and polarity during neutrophil chemotaxis. Journal of Cell Biology, 174(5):647–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/174/5/647.

Kuribayashi:2007:ERP

[KNT^{+}07] Kageaki Kuribayashi, Kiminori Nakamura, Maki Tanaka, Tsutomu Sato, Junji Kato, Katsunori Sasaki, Rishu Takimoto, Katsuhisa Kogawa, Takeshi Terui, Tetsuji Takayama, Takayuki Onuma, Takuya Matsunaga, and Yoshiro Niitsu. Essential role of protein kinase C ζ in transducing a motility signal induced by superoxide and a chemotactic peptide, fMLP. Journal of Cell Biology, 176(7):1049–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/7/1049.

Karbowski:2007:MEU

[KNY07] Mariusz Karbowski, Albert Neutzner, and Richard J. Youle. The mitochondrial E3 ubiquitin ligase MARCH5 is required for Drp1 dependent mitochondrial division. Journal of Cell Biology, 178(1):71–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/1/71.

Ku:2006:DPR

[KO06] Nam-On Ku and M. Bishr Omary. A disease- and phosphorylationrelated nonmechanical function for keratin 8. Journal of Cell Biology, 174(1):115–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/1/115.

Kimura:2007:LCP

[KO07] Akatsuki Kimura and Shuichi Onami. Local cortical pullingforce repression switches centrosomal centration and posterior displacement in C. elegans. Journal of Cell Biology, 179(7): 1347–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/7/1347.

Kiuchi:2007:CPS

[KOKM07] Tai Kiuchi, Kazumasa Ohashi, Souichi Kurita, and Kensaku Mizuno. Cofilin promotes stimulus-induced lamellipodium formation by generating an abundant supply of actin monomers. Journal of Cell Biology, 177(3):465–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/177/3/465.

Kamasaki:2007:TDA

[KOM07] Tomoko Kamasaki, Masako Osumi, and Issei Mabuchi. Threedimensional arrangement of F-actin in the contractile ring of fission yeast. Journal of Cell Biology, 178(5):765–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/178/5/
765.

Kemmler:2009:YRL

[KOVP09] Stefan Kemmler, Laura Occhipinti, Maria Veisu, and Vikram Govind Panse. Yvh1 is required for a late maturation step in the 60S biogenesis pathway. Journal of Cell Biology, 186(6):863– ??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/186/6/863.

Kermorgant:2008:RTC

[KP08] Stéphanie Kermorgant and Peter J. Parker. Receptor trafficking controls weak signal delivery: a strategy used by c–Met for STAT3 nuclear accumulation. Journal of Cell Biology, 182(5): 855–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/5/855.

Klee:2005:BXS

[KPM05] Martina Klee and Felipe X. Pimentel-Muiños. Bcl-X_{L} specifically activates Bak to induce swelling and restructuring of the endoplasmic reticulum. Journal of Cell Biology, 168(5): 723–??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/5/723.

Kilburn:2007:NTBa

[KPR^{+}07a] Chandra L. Kilburn, Chad G. Pearson, Edwin P. Romijn, Janet B. Meehl, Thomas H. Giddings, Brady P. Culver, John R. Yates, and Mark Winey. New Tetrahymena basal body protein components identify basal body domain structure. Journal of Cell Biology, 178(6):905–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/178/6/905.

Kilburn:2007:NTBb

[KPR^{+}07b] Chandra L. Kilburn, Chad G. Pearson, Edwin P. Romijn, Janet B. Meehl, Thomas H. Giddings, Brady P. Culver, John R. Yates, and Mark Winey. New Tetrahymena basal body protein components identify basal body domain structure. Journal of Cell Biology, 179(1):167–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/1/167.

Kuznetsov:2007:RDM

[KPS^{+}07] Sergey Kuznetsov, Manuela Pellegrini, Kristy Shuda, Oscar Fernandez-Capetillo, Yilun Liu, Betty K. Martin, Sandra Burkett, Eileen Southon, Debananda Pati, Lino Tessarollo, Stephen C. West, Peter J. Donovan, Andre Nussenzweig, and Shyam K. Sharan. RAD51C deficiency in mice results in early prophase I arrest in males and sister chromatid separation at metaphase II in females. Journal of Cell Biology, 176(5):

581–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/5/581.

Kim:2006:CCR

[KR06] Dae Young Kim and Richard Roy. Cell cycle regulators control centrosome elimination during oogenesis in Caenorhabditis elegans. Journal of Cell Biology, 174(6):751–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/174/6/751.

Kang:2007:MDB

[KR07] Jihong Kang and Serge Rivest. MyD88-deficient bone marrow cells accelerate onset and reduce survival in a mouse model of amyotrophic lateral sclerosis. Journal of Cell Biology, 179(6): 1219–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/6/1219.

Kutik:2008:TMP

[KRG^{+}08] Stephan Kutik, Michael Rissler, Xue Li Guan, Bernard Guiard, Guanghou Shui, Natalia Gebert, Philip N. Heacock, Peter Rehling, William Dowhan, Markus R. Wenk, Nikolaus Pfanner, and Nils Wiedemann. The translocator maintenance protein Tam41 is required for mitochondrial cardiolipin biosynthesis. Journal of Cell Biology, 183(7):1213–??, December 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/183/7/1213.

Kelley:2009:CDE

[KRP^{+}09] Rusty Kelley, Rongqin Ren, Xinchun Pi, Yaxu Wu, Isabel Moreno, Monte Willis, Martin Moser, Malcolm Ross, Monika Podkowa, Liliana Attisano, and Cam Patterson. A concentration-dependent endocytic trap and sink mechanism converts Bmper from an activator to an inhibitor of Bmp signaling. Journal of Cell Biology, 184(4):597–??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/184/4/597.

Kressler:2008:AAR

[KRPH08] Dieter Kressler, Daniela Roser, Brigitte Pertschy, and Ed Hurt. The AAA ATPase Rix7 powers progression of ribosome biogenesis by stripping Nsa1 from pre-60S particles. Journal of Cell Biology, 181(6):935–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/6/935.

Kakinuma:2008:KRR

[KRZ^{+}08] Naoto Kakinuma, Badal Chandra Roy, Yun Zhu, Yong Wang, and Ryoiti Kiyama. Kank regulates RhoA-dependent formation of actin stress fibers and cell migration via 14-3-3 in PI3K–Akt signaling. Journal of Cell Biology, 181(3):537–??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/181/3/537.

Kumaran:2008:GLT

[KS08] R. Ileng Kumaran and David L. Spector. A genetic locus targeted to the nuclear periphery in living cells maintains its transcriptional competence. Journal of Cell Biology, 180(1):51–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
180/1/51.

Kedersha:2005:SGP

[KSA^{+}05] Nancy Kedersha, Georg Stoecklin, Maranatha Ayodele, Patrick Yacono, Jens Lykke-Andersen, Marvin J. Fritzler, Donalyn Scheuner, Randal J. Kaufman, David E. Golan, and Paul Anderson. Stress granules and processing bodies are dynamically linked sites of mRNP remodeling. Journal of Cell Biology, 169 (6):871–??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/6/871.

Kumin:2007:PRB

[KSE^{+}07] Angelika Kümin, Matthias Schäfer, Nikolas Epp, Philippe Bugnon, Christiane Born-Berclaz, Annette Oxenius, Anke Klippel, Wilhelm Bloch, and Sabine Werner. Peroxiredoxin 6 is required for blood vessel integrity in wounded skin. Journal of Cell Biology, 179(4):747–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/4/747.

Khanday:2006:SMA

[KSK^{+}06] Firdous A. Khanday, Lakshmi Santhanam, Kenji Kasuno, Tohru Yamamori, Asma Naqvi, Jeremy DeRicco, Artem Bugayenko, Ilwola Mattagajasingh, Andrea Disanza, Giorgio Scita, and Kaikobad Irani. Sos-mediated activation of rac1 by p66shc. Journal of Cell Biology, 172(6):817–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/172/6/817.

Kolas:2005:LMP

[KSL^{+}05] Nadine K. Kolas, Anton Svetlanov, Michelle L. Lenzi, Frank P. Macaluso, Steven M. Lipkin, R. Michael Liskay, John Greally, Winfried Edelmann, and Paula E. Cohen. Localization of MMR proteins on meiotic chromosomes in mice indicates distinct functions during prophase I. Journal of Cell Biology, 171(3):447–??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/3/447.

Kreuger:2006:IBH

[KSpLL06] Johan Kreuger, Dorothe Spillmann, Jin ping Li, and Ulf Lindahl. Interactions between heparan sulfate and proteins: the concept of specificity. Journal of Cell Biology, 174(3):323–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
174/3/323.

Kunwar:2008:TGI

[KSR^{+}08] Prabhat S. Kunwar, Hiroko Sano, Andrew D. Renault, Vitor Barbosa, Naoyuki Fuse, and Ruth Lehmann. Tre1 GPCR initiates germ cell transepithelial migration by regulating Drosophila melanogaster E-cadherin. Journal of Cell Biology, 183(1):157–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/1/157.

Keck:2007:COI

[KST^{+}07] Jamie M. Keck, Matthew K. Summers, Donato Tedesco, Susanna Ekholm-Reed, Li-Chiou Chuang, Peter K. Jackson, and Steven I. Reed. Cyclin E overexpression impairs progression through mitosis by inhibiting APC^{Cdh1}. Journal of Cell Biology, 178(3):371–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/3/371.

Kimura:2006:NHE

[KTA^{+}06] Hiroshi Kimura, Nanako Takizawa, Eric Allemand, Tetsuya Hori, Francisco J. Iborra, Naohito Nozaki, Michiko Muraki, Masatoshi Hagiwara, Adrian R. Krainer, Tatsuo Fukagawa, and Katsuya Okawa. A novel histone exchange factor, protein phosphatase 2Cγ, mediates the exchange and dephosphorylation of H2A–H2B. Journal of Cell Biology, 175(3):389–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/175/3/
389.

Kino:2005:GPI

[KTI^{+}05] Tomoshige Kino, Anatoly Tiulpakov, Takamasa Ichijo, Ly Chheng, Tohru Kozasa, and George P. Chrousos. G protein β interacts with the glucocorticoid receptor and suppresses its transcriptional activity in the nucleus. Journal of Cell Biology, 169 (6):885–??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/6/885.

Kohsaka:2007:VIP

[KTN07] Hiroshi Kohsaka, Etsuko Takasu, and Akinao Nose. In vivo induction of postsynaptic molecular assembly by the cell adhesion molecule Fasciclin2. Journal of Cell Biology, 179(6): 1289–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/6/1289.

Kotani:2005:R

[KTYT05] Shuji Kotani, Hirofumi Tanaka, Hideyo Yasuda, and Kazuo Todokoro. Retraction. Journal of Cell Biology, 169(1):205–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
169/1/205.

Karr:2009:RGR

[KVC^{+}09] Julie Karr, Vasia Vagin, Kaiyun Chen, Subhashree Ganesan, Oxana Olenkina, Vladimir Gvozdev, and David E. Featherstone. Regulation of glutamate receptor subunit availability by microRNAs. Journal of Cell Biology, 185(4):685–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/185/4/

Klaile:2009:CTI

[KVS^{+}09] Esther Klaile, Olga Vorontsova, Kristmundur Sigmundsson, Mario M. Müller, Bernhard B. Singer, Lars-Göran Öfverstedt, Stina Svensson, Ulf Skoglund, and Björn Öbrink. The CEA-CAM1 N-terminal Ig domain mediates cis- and trans-binding and is essential for allosteric rearrangements of CEACAM1 microclusters. Journal of Cell Biology, 187(4):553–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/187/4/
553.

Krouwels:2005:GHM

[KWA^{+}05] Ilke M. Krouwels, Karien Wiesmeijer, Tsion E. Abraham, Chris Molenaar, Nico P. Verwoerd, Hans J. Tanke, and Roeland W. Dirks. A glue for heterochromatin maintenance. Journal of Cell Biology, 170(4):537–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/4/537.

Kawasaki:2008:RSD

[KWS^{+}08] Kyoko Kawasaki, Tetsuro Watabe, Hitoshi Sase, Masanori Hirashima, Hiroshi Koide, Yasuyuki Morishita, Keiko Yuki, Toshikuni Sasaoka, Toshio Suda, Motoya Katsuki, Kohei Miyazono, and Keiji Miyazawa. Ras signaling directs endothelial specification of VEGFR2^{+} vascular progenitor cells. Journal of Cell Biology, 181(1):131–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/181/1/131.

Komatsu:2005:ISI

[KWU^{+}05] Masaaki Komatsu, Satoshi Waguri, Takashi Ueno, Junichi Iwata, Shigeo Murata, Isei Tanida, Junji Ezaki, Noboru Mizushima, Yoshinori Ohsumi, Yasuo Uchiyama, Eiki Kominami, Keiji Tanaka, and Tomoki Chiba. Impairment of starvation-induced and constitutive autophagy in Atg7deficient mice. Journal of Cell Biology, 169(3):425–??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/169/3/
425.

Kuo:2006:TSD

[KWY^{+}06] Jean-Cheng Kuo, Won-Jing Wang, Chung-Chen Yao, Pei-Rung Wu, and Ruey-Hwa Chen. The tumor suppressor DAPK inhibits cell motility by blocking the integrin-mediated polarity pathway. Journal of Cell Biology, 172(4):619–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/172/4/
619.

Kumada:2006:SCL

[KYK^{+}06] Kazuki Kumada, Ryoji Yao, Tokuichi Kawaguchi, Mika Karasawa, Yutaka Hoshikawa, Koji Ichikawa, Yoshinobu Sugitani, Issei Imoto, Johji Inazawa, Minoru Sugawara, Mitsuhiro Yanagida, and Tetsuo Noda. The selective continued linkage of centromeres from mitosis to interphase in the absence of mammalian separase. Journal of Cell Biology, 172(6):835–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/172/6/
835.

Kappas:2008:VRF

[KZC^{+}08] Nicholas C. Kappas, Gefei Zeng, John C. Chappell, Joseph B. Kearney, Surovi Hazarika, Kimberly G. Kallianos, Cam Patterson, Brian H. Annex, and Victoria L. Bautch. The VEGF receptor Flt-1 spatially modulates Flk-1 signaling and blood vessel branching. Journal of Cell Biology, 181(5):847–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/181/5/
847.

Kothapalli:2007:HCA

[KZH^{+}07] Devashish Kothapalli, Liang Zhao, Elizabeth A. Hawthorne, Yan Cheng, Eric Lee, Ellen Puré, and Richard K. Assoian. Hyaluronan and CD44 antagonize mitogen-dependent cyclin D1 expression in mesenchymal cells. Journal of Cell Biology, 176 (4):535–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/4/535.

Kruger:2007:ISR

[KZS07] Tim Krüger, Hanswalter Zentgraf, and Ulrich Scheer. Intranucleolar sites of ribosome biogenesis defined by the localization of early binding ribosomal proteins. Journal of Cell Biology, 177(4):573–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/4/573.

Khosravani:2008:PPA

[KZT^{+}08] Houman Khosravani, Yunfeng Zhang, Shigeki Tsutsui, Shahid Hameed, Christophe Altier, Jawed Hamid, Lina Chen, Michelle Villemaire, Zenobia Ali, Frank R. Jirik, and Gerald W. Zamponi. Prion protein attenuates excitotoxicity by inhibiting NMDA receptors. Journal of Cell Biology, 181(3):551–??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/181/3/
551.

Khosravani:2009:PPA

[KZT^{+}09] Houman Khosravani, Yunfeng Zhang, Shigeki Tsutsui, Shahid Hameed, Christophe Altier, Jawed Hamid, Lina Chen, Michelle Villemaire, Zenobia Ali, Frank R. Jirik, and Gerald W. Zamponi. Prion protein attenuates excitotoxicity by inhibiting NMDA receptors. Journal of Cell Biology, 185(6):1127–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/185/6/
1127.

Lassar:2009:PMF

[Las09] Andrew B. Lassar. The p38 MAPK family, a pushmi-pullyu of skeletal muscle differentiation. Journal of Cell Biology, 187(7): 941–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/187/7/941.

Loo:2008:SPP

[LB08] Tsui-Han Loo and Mohan Balasubramanian. Schizosaccharomyces pombe Pak-related protein, Pak1p/Orb2p, phosphorylates myosin regulatory light chain to inhibit cytokinesis. Journal of Cell Biology, 183(5):785–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/183/5/785.

Ling:2007:TPP

[LBC^{+}07] Kun Ling, Shawn F. Bairstow, Chateen Carbonara, Dmitry A. Turbin, David G. Huntsman, and Richard A. Anderson. Type Iγ phosphatidylinositol phosphate kinase modulates adherens junction and E-cadherin trafficking via a direct interaction with μ1B adaptin. Journal of Cell Biology, 176(3):343–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/176/3/
343.

Liegeois:2006:VAM

[LBG^{+}06] Samuel Liégeois, Alexandre Benedetto, Jean-Marie Garnier, Yannick Schwab, and Michel Labouesse. The V0-ATPase mediates apical secretion of exosomes containing Hedgehog-related proteins in Caenorhabditis elegans. Journal of Cell Biology, 173(6):949–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/6/949.

Liang:2005:CBU

[LBH^{+}05] Feng-Xia Liang, Maarten C. Bosland, Hongying Huang, Rok Romih, Solange Baptiste, Fang-Ming Deng, Xue-Ru Wu, Ellen Shapiro, and Tung-Tien Sun. Cellular basis of urothelial squamous metaplasia. Journal of Cell Biology, 171(5):835–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
171/5/835.

LeClaire:2008:PAC

[LBI^{+}08] Lawrence L. LeClaire, Martin Baumgartner, Janet H. Iwasa, R. Dyche Mullins, and Diane L. Barber. Phosphorylation of the Arp2/3 complex is necessary to nucleate actin filaments. Journal of Cell Biology, 182(4):647–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/182/4/647.

Li:2008:HTS

[LBS^{+}08] Haiyang Li, Adayabalam S. Balajee, Tao Su, Bo Cen, Tom K. Hei, and I. Bernard Weinstein. The HINT1 tumor suppressor regulates both γ-H2AX and ATM in response to DNA damage. Journal of Cell Biology, 183(2):253–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/183/2/253.

Landeira:2009:CRV

[LBVN09] David Landeira, Jean-Mathieu Bart, Daria Van Tyne, and Miguel Navarro. Cohesin regulates VSG monoallelic expression in trypanosomes. Journal of Cell Biology, 186(2):243–??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/186/2/
243.

Lee:2007:JRR

[LC07] Sangil Lee and Lynn Cooley. Jagunal is required for reorganizing the endoplasmic reticulum during Drosophila oogenesis. Journal of Cell Biology, 176(7):941–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/176/7/941.

Lawrence:2008:GAT

[LC08a] Jeanne B. Lawrence and Christine M. Clemson. Gene associations: true romance or chance meeting in a nuclear neighborhood? Journal of Cell Biology, 182(6):1035–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/182/6/
1035.

Lee:2008:RRD

[LC08b] Hsiao-Hui Lee and Zee-Fen Chang. Regulation of RhoAdependent ROCKII activation by Shp2. Journal of Cell Biology, 181(6):999–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/6/999.

Lee:2009:SOK

[LC09] Chang-Hun Lee and Pierre A. Coulombe. Self-organization of keratin intermediate filaments into cross-linked networks. Journal of Cell Biology, 186(3):409–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/186/3/409.

Liebner:2008:WCS

[LCB^{+}08] Stefan Liebner, Monica Corada, Thorsten Bangsow, Jane Babbage, Andrea Taddei, Cathrin J. Czupalla, Marco Reis, Angelina Felici, Hartwig Wolburg, Marcus Fruttiger, Makoto M. Taketo, Harald von Melchner, Karl Heinz Plate, Holger Gerhardt, and Elisabetta Dejana. Wnt/ β-catenin signaling controls development of the blood–brain barrier. Journal of Cell Biology, 183(3):409–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/183/3/409.

Lakkaraju:2009:ITT

[LCGRB09] Aparna Lakkaraju, Jose Maria Carvajal-Gonzalez, and Enrique Rodriguez-Boulan. It takes two to tango to the melanosome. Journal of Cell Biology, 187(2):161–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/187/2/161.

Luo:2005:SRA

[LCL^{+}05] Shuo Luo, Yu Chen, Kwok-On Lai, Juan Carlos Arévalo, Stanley C. Froehner, Marvin E. Adams, Moses V. Chao, and Nancy Y. Ip. α-syntrophin regulates ARMS localization at the neuromuscular junction and enhances EphA4 signaling in an ARMS-dependent manner. Journal of Cell Biology, 169(5):813– ??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
169/5/813.

Link:2007:CFC

[LCL^{+}07] Nichole Link, Po Chen, Wan-Jin Lu, Kristi Pogue, Amy Chuong, Miguel Mata, Joshua Checketts, and John M. Abrams. A collective form of cell death requires homeodomain interacting protein kinase. Journal of Cell Biology, 178(4):567–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
178/4/567.

Lombardi:2008:HDN

[LCM^{+}08] Maria Lombardi, Gabriella Castoria, Antimo Migliaccio, Maria Vittoria Barone, Rosina Di Stasio, Alessandra Ciociola, Daniela Bottero, Hiroshi Yamaguchi, Ettore Appella, and Ferdinando Auricchio. Hormone-dependent nuclear export of estradiol receptor and DNA synthesis in breast cancer cells. Journal of Cell Biology, 182(2):327–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/182/2/327.

Lancaster:2007:NPB

[LCO07] Oscar M. Lancaster, C. Fiona Cullen, and Hiroyuki Ohkura. NHK-1 phosphorylates BAF to allow karyosome formation in the Drosophila oocyte nucleus. Journal of Cell Biology, 179(5): 817–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/5/817.

Leyvraz:2005:EBF

[LCR^{+}05] Céline Leyvraz, Roch-Philippe Charles, Isabelle Rubera, Marjorie Guitard, Samuel Rotman, Bernadette Breiden, Konrad Sandhoff, and Edith Hummler. The epidermal barrier function is dependent on the serine protease CAP1/ Prss8. Journal of Cell Biology, 170(3):487–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/170/3/487.

Lin:2009:GRE

[LCS^{+}09] Fang Lin, Songhai Chen, Diane S. Sepich, Jennifer Ray Panizzi, Sherry G. Clendenon, James A. Marrs, Heidi E. Hamm, and Lilianna Solnica-Krezel. Gα_{12/13} regulate epiboly by inhibiting E-cadherin activity and modulating the actin cytoskeleton. Journal of Cell Biology, 184(6):909–??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/184/6/909.

Li:2009:PAM

[LCT^{+}09] Jingxia Li, Haobin Chen, Moon-Shong Tang, Xianglin Shi, Shantu Amin, Dhimant Desai, Max Costa, and Chuanshu Huang. PI-3K and Akt are mediators of AP-1 induction by 5-MCDE in mouse epidermal Cl41 cells. Journal of Cell Biology, 186(2):297–??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/186/2/297.

Lee:2008:NBL

[LCvD^{+}08] Erinna F. Lee, Peter E. Czabotar, Mark F. van Delft, Ewa M. Michalak, Michelle J. Boyle, Simon N. Willis, Hamsa Puthalakath, Philippe Bouillet, Peter M. Colman, David C. S. Huang, and W. Douglas Fairlie. A novel BH3 ligand that selectively targets Mcl-1 reveals that apoptosis can proceed without Mcl-1 degradation. Journal of Cell Biology, 180(2):341–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/180/2/
341.

Lamouille:2007:CSI

[LD07] Samy Lamouille and Rik Derynck. Cell size and invasion in TGF-β-induced epithelial to mesenchymal transition is regulated by activation of the mTOR pathway. Journal of Cell Biology, 178(3):437–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/3/437.

Latham:2009:H

[LD09] John A. Latham and Sharon Y. R. Dent. The H2BK123Rgument. Journal of Cell Biology, 186(3):313–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/186/3/313.

Lee:2008:BCS

[LDGK08] Intaek Lee, Balraj Doray, Jennifer Govero, and Stuart Kornfeld. Binding of cargo sorting signals to AP-1 enhances its association with ADP ribosylation factor 1–GTP. Journal of Cell Biology, 180(3):467–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/3/467.

Li:2009:EBP

[LDGO^{+}09] Wei Li, Leah R. DeBella, Tugba Guven-Ozkan, Rueyling Lin, and Lesilee S. Rose. An eIF4E-binding protein regulates katanin protein levels in C. elegans embryos. Journal of Cell Biology, 187(1):33–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/1/33.

Lee:2006:EMT

[LDKT06] Jonathan M. Lee, Shoukat Dedhar, Raghu Kalluri, and Erik W. Thompson. The epithelial–mesenchymal transition: new insights in signaling, development, and disease. Journal of Cell Biology, 172(7):973–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/7/973.

Lim:2009:UPV

[LDL^{+}09a] Precious J. Lim, Rebecca Danner, Jing Liang, Howard Doong, Christine Harman, Deepa Srinivasan, Cara Rothenberg, Hongmin Wang, Yihong Ye, Shengyun Fang, and Mervyn J. Monteiro. Ubiquilin and p97/VCP bind erasin, forming a complex involved in ERAD. Journal of Cell Biology, 187(2):201–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
187/2/201.

Luijsterburg:2009:HPR

[LDL^{+}09b] Martijn S. Luijsterburg, Christoffel Dinant, Hannes Lans, Jan Stap, Elzbieta Wiernasz, Saskia Lagerwerf, Daniël O. Warmerdam, Michael Lindh, Maartje C. Brink, Jurek W. Dobrucki, Jacob A. Aten, Maria I. Fousteri, Gert Jansen, Nico P. Dantuma, Wim Vermeulen, Leon H. F. Mullenders, Adriaan B. Houtsmuller, Pernette J. Verschure, and Roel van Driel. Heterochromatin protein 1 is recruited to various types of DNA damage. Journal of Cell Biology, 185(4):577–??, May 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/185/4/577.

Liu:2005:ADR

[LDPLS05] Wei Liu, Rainer Duden, Robert D. Phair, and Jennifer Lippincott-Schwartz. ArfGAP1 dynamics and its role in COPI coat assembly on Golgi membranes of living cells. Journal of Cell Biology, 168(7):1053–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/168/7/1053.

Lechtreck:2008:MHI

[LDR^{+}08] Karl-Ferdinand Lechtreck, Philippe Delmotte, Michael L. Robinson, Michael J. Sanderson, and George B. Witman. Mutations in Hydin impair ciliary motility in mice. Journal of Cell Biology, 180(3):633–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/3/633.

Lam:2006:PDP

[LDS^{+}06] Karen K. Y. Lam, Michael Davey, Beimeng Sun, Amy F. Roth, Nicholas G. Davis, and Elizabeth Conibear. Palmitoylation by the DHHC protein Pfa4 regulates the ER exit of Chs3. Journal of Cell Biology, 174(1):19–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/174/1/19.

Lamkanfi:2006:CLB

[LDVV06] Mohamed Lamkanfi, Wim Declercq, Tom Vanden Berghe, and Peter Vandenabeele. Caspases leave the beaten track: caspasemediated activation of NF-κB. Journal of Cell Biology, 173 (2):165–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/2/165.

LeBrasseur:2005:AAC

[LeB05a] Nicole LeBrasseur. Actin adhesion controls bacterial movement. Journal of Cell Biology, 171(4):??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/171/4/582.2.

LeBrasseur:2005:AR

[LeB05b] Nicole LeBrasseur. An actin ruler. Journal of Cell Biology, 170 (6):??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/6/861.3.

LeBrasseur:2005:ACM

[LeB05c] Nicole LeBrasseur. Acyl cycles move Ras. Journal of Cell Biology, 168(6):??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/6/850.2.

LeBrasseur:2005:APO

[LeB05d] Nicole LeBrasseur. APC proteolysis not overdone. Journal of Cell Biology, 169(1):??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/1/16.3.

LeBrasseur:2005:AIR

[LeB05e] Nicole LeBrasseur. APP for injury repair. Journal of Cell Biology, 170(5):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/5/699.1.

LeBrasseur:2005:ATS

[LeB05f] Nicole LeBrasseur. Autoimmune TCR structure. Journal of Cell Biology, 169(3):??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/3/376.1.

LeBrasseur:2005:BBI

[LeB05g] Nicole LeBrasseur. Binding a bent integrin. Journal of Cell Biology, 168(7):??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/7/985.1.

LeBrasseur:2005:BCE

[LeB05h] Nicole LeBrasseur. BiP chaperones ER entry. Journal of Cell Biology, 168(3):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/3/346.2.

LeBrasseur:2005:BPF

[LeB05i] Nicole LeBrasseur. Blood pressure and fertility. Journal of Cell Biology, 168(4):??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/4/522.1.

LeBrasseur:2005:BCB

[LeB05j] Nicole LeBrasseur. Breaking cell bonds. Journal of Cell Biology, 171(1):??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/1/8.2.

LeBrasseur:2005:BPR

[LeB05k] Nicole LeBrasseur. Bringing PIPs to root tips. Journal of Cell Biology, 168(5):??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/5/673.1.

LeBrasseur:2005:BGM

[LeB05l] Nicole LeBrasseur. Building glial microtubules. Journal of Cell Biology, 170(4):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/4/511.3.

LeBrasseur:2005:CRM

[LeB05m] Nicole LeBrasseur. Cadherin rules microvilli. Journal of Cell Biology, 171(3):??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/3/405.3.

LeBrasseur:2005:CSS

[LeB05n] Nicole LeBrasseur. Calcium says stop. Journal of Cell Biology, 168(4):??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/4/522.2.

LeBrasseur:2005:CPH

[LeB05o] Nicole LeBrasseur. Cdk5 protects huntingtin. Journal of Cell Biology, 169(4):??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/4/548.3.

LeBrasseur:2005:CDR

[LeB05p] Nicole LeBrasseur. Cells do not relive youth. Journal of Cell Biology, 168(6):??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/6/851.2.

LeBrasseur:2005:CLM

[LeB05q] Nicole LeBrasseur. Centromeric links in meiosis. Journal of Cell Biology, 169(5):??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/5/703.3.

LeBrasseur:2005:CFB

[LeB05r] Nicole LeBrasseur. Centrosome-free basal MT organization. Journal of Cell Biology, 171(5):??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/171/5/754.1.

LeBrasseur:2005:CCC

[LeB05s] Nicole LeBrasseur. Complexes to coordinate calcium. Journal of Cell Biology, 169(4):??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/4/549.2.

LeBrasseur:2005:CNA

[LeB05t] Nicole LeBrasseur. Contrasts in neuronal aggregates. Journal of Cell Biology, 171(1):??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/171/1/8.1.

LeBrasseur:2005:CK

[LeB05u] Nicole LeBrasseur. Cooperative kinesin. Journal of Cell Biology, 168(5):??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/5/672.1.

LeBrasseur:2005:CBS

[LeB05v] Nicole LeBrasseur. Crossovers break SC symmetry. Journal of Cell Biology, 168(5):??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/168/5/672.2.

LeBrasseur:2005:CMF

[LeB05w] Nicole LeBrasseur. Cycles of muscle fusion. Journal of Cell Biology, 169(6):??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/6/835.1.

LeBrasseur:2005:DD

[LeB05x] Nicole LeBrasseur. Deacetylate to differentiate. Journal of Cell Biology, 169(4):??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/4/548.1.

LeBrasseur:2005:DM

[LeB05y] Nicole LeBrasseur. Dividing via miRNA. Journal of Cell Biology, 170(1):??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/1/10.1.

LeBrasseur:2005:DDL

[LeB05z] Nicole LeBrasseur. DNA damage lures T cells. Journal of Cell Biology, 170(3):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/3/337.2.

LeBrasseur:2005:DPS

[LeB05-27] Nicole LeBrasseur. Docked to prevent secretion. Journal of Cell Biology, 171(1):??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/1/9.3.

LeBrasseur:2005:DKC

[LeB05-28] Nicole LeBrasseur. Does the Kai clock rotate? Journal of Cell Biology, 169(5):??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/5/702.1.

LeBrasseur:2005:DSW

[LeB05-29] Nicole LeBrasseur. DPP shapes up wing cells. Journal of Cell Biology, 169(1):??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/1/17.2.

LeBrasseur:2005:ENG

[LeB05-30] Nicole LeBrasseur. The early neurite gets the axon. Journal of Cell Biology, 170(5):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/5/698.2.

LeBrasseur:2005:EAS

[LeB05-31] Nicole LeBrasseur. eIF2 activated in spots. Journal of Cell Biology, 170(6):??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/6/861.1.

LeBrasseur:2005:D

[LeB05-32] Nicole LeBrasseur. ER domains. Journal of Cell Biology, 169(6):??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/6/834.1.

LeBrasseur:2005:TCD

[LeB05-33] Nicole LeBrasseur. ER for tiny calcium domains. Journal of Cell Biology, 171(1):??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/1/9.2.

LeBrasseur:2005:SSM

[LeB05-34] Nicole LeBrasseur. ER stress shapes muscle. Journal of Cell Biology, 169(4):??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/4/548.2.

LeBrasseur:2005:EH

[LeB05-35] Nicole LeBrasseur. The evolution of hubs. Journal of Cell Biology, 170(3):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/3/337.1.

LeBrasseur:2005:EFW

[LeB05-36] Nicole LeBrasseur. Exocyst finds way without Rho. Journal of Cell Biology, 170(4):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/4/511.2.

LeBrasseur:2005:FGY

[LeB05-37] Nicole LeBrasseur. FGF gets in your bones. Journal of Cell Biology, 168(7):??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/7/985.2.

LeBrasseur:2005:FCb

[LeB05-38] Nicole LeBrasseur. Folding in a crowd. Journal of Cell Biology, 169(1):??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/1/16.2.

LeBrasseur:2005:FSS

[LeB05-39] Nicole LeBrasseur. Formin steps and slips. Journal of Cell Biology, 170(6):??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/6/860.2.

LeBrasseur:2005:FCa

[LeB05-40] Nicole LeBrasseur. Freedom in condensation. Journal of Cell Biology, 168(1):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/1/9.3.

LeBrasseur:2005:GPL

[LeB05-41] Nicole LeBrasseur. G proteins live in excess. Journal of Cell Biology, 171(3):??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/3/405.2.

LeBrasseur:2005:GB

[LeB05-42] Nicole LeBrasseur. Generous bacteria. Journal of Cell Biology, 170(3):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/3/336.1.

LeBrasseur:2005:GRC

[LeB05-43] Nicole LeBrasseur. Golgi relies on centrin. Journal of Cell Biology, 171(4):??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/4/583.2.

LeBrasseur:2005:HK

[LeB05-44] Nicole LeBrasseur. A hairy kinase. Journal of Cell Biology, 170(4):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/4/511.1.

LeBrasseur:2005:IME

[LeB05-45] Nicole LeBrasseur. Inequality made equal. Journal of Cell Biology, 169(5):??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/5/703.2.

LeBrasseur:2005:IT

[LeB05-46] Nicole LeBrasseur. Infection in trans. Journal of Cell Biology, 170(1):??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/1/10.2.

LeBrasseur:2005:LSS

[LeB05-47] Nicole LeBrasseur. Lattice supports single bilayer. Journal of Cell Biology, 169(2):??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/2/213.1.

LeBrasseur:2005:LM

[LeB05-48] Nicole LeBrasseur. The learning matrix. Journal of Cell Biology, 168(6):??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/6/851.1.

LeBrasseur:2005:LS

[LeB05-49] Nicole LeBrasseur. Links that separate. Journal of Cell Biology, 169(6):??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/6/835.2.

LeBrasseur:2005:LLN

[LeB05-50] Nicole LeBrasseur. LIS-less neurons fall behind. Journal of Cell Biology, 170(6):??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/6/860.1.

LeBrasseur:2005:LPS

[LeB05-51] Nicole LeBrasseur. Local power for synapses. Journal of Cell Biology, 170(5):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/5/699.3.

LeBrasseur:2005:LO

[LeB05-52] Nicole LeBrasseur. Localized oxidation. Journal of Cell Biology, 171(5):??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/5/755.3.

LeBrasseur:2005:LSM

[LeB05-53] Nicole LeBrasseur. Loopy Sec61 mutants. Journal of Cell Biology, 168(1):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/1/8.1.

LeBrasseur:2005:LCC

[LeB05-54] Nicole LeBrasseur. Lumen cues calcium release. Journal of Cell Biology, 168(4):??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/4/523.1.

LeBrasseur:2005:MDN

[LeB05-55] Nicole LeBrasseur. Marks of death on nuclei. Journal of Cell Biology, 171(2):??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/2/193.3.

LeBrasseur:2005:MD

[LeB05-56] Nicole LeBrasseur. Matrix of death. Journal of Cell Biology, 171(3):??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/3/404.2.

LeBrasseur:2005:MBM

[LeB05-57] Nicole LeBrasseur. Microtubules born on microtubules. Journal of Cell Biology, 169(2):??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/2/212.2.

LeBrasseur:2005:MLM

[LeB05-58] Nicole LeBrasseur. MITF links maturation and quiescence. Journal of Cell Biology, 168(1):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/168/1/8.2.

LeBrasseur:2005:MRD

[LeB05-59] Nicole LeBrasseur. MMPs rearrange DNA. Journal of Cell Biology, 170(3):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/3/336.2.

LeBrasseur:2005:MTT

[LeB05-60] Nicole LeBrasseur. Motors take turns. Journal of Cell Biology, 169(3):??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/3/377.3.

LeBrasseur:2005:MCT

[LeB05-61] Nicole LeBrasseur. Muscle cells take heart. Journal of Cell Biology, 169(1):??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/1/16.1.

LeBrasseur:2005:MRN

[LeB05-62] Nicole LeBrasseur. Myosin reins in neurites. Journal of Cell Biology, 169(5):??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/5/703.1.

LeBrasseur:2005:MVS

[LeB05-63] Nicole LeBrasseur. Myristoylation versus SRP. Journal of Cell Biology, 168(5):??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/5/672.3.

LeBrasseur:2005:NEE

[LeB05-64] Nicole LeBrasseur. Need an EGO to escape arrest. Journal of Cell Biology, 170(3):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/3/336.3.

LeBrasseur:2005:NDS

[LeB05-65] Nicole LeBrasseur. Nerve damage surveillance. Journal of Cell Biology, 168(5):??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/5/673.2.

LeBrasseur:2005:NIC

[LeB05-66] Nicole LeBrasseur. The nesprin–IF connection. Journal of Cell Biology, 171(5):??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/5/755.1.

LeBrasseur:2005:NBN

[LeB05-67] Nicole LeBrasseur. Neurons born in new places. Journal of Cell Biology, 168(3):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/3/347.1.

LeBrasseur:2005:NMP

[LeB05-68] Nicole LeBrasseur. A new mRNA polymerase. Journal of Cell Biology, 170(5):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/5/699.2.

LeBrasseur:2005:NR

[LeB05-69] Nicole LeBrasseur. New route to Aβ. Journal of Cell Biology, 171(1):??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/1/9.1.

LeBrasseur:2005:NVR

[LeB05-70] Nicole LeBrasseur. No vacuolar reunion. Journal of Cell Biology, 168(3):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/3/346.1.

LeBrasseur:2005:NBS

[LeB05-71] Nicole LeBrasseur. Nonadhesion-based sorting. Journal of Cell Biology, 170(4):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/4/510.1.

LeBrasseur:2005:NO

[LeB05-72] Nicole LeBrasseur. A Notch outlier. Journal of Cell Biology, 170(6):??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/6/861.2.

LeBrasseur:2005:OSE

[LeB05-73] Nicole LeBrasseur. One SecY is enough. Journal of Cell Biology, 169(2):??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/2/213.2.

LeBrasseur:2005:OEC

[LeB05-74] Nicole LeBrasseur. Orienting early cleavage. Journal of Cell Biology, 169(1):??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/1/17.1.

LeBrasseur:2005:AC

[LeB05-75] Nicole LeBrasseur. Out with the active β-catenin. Journal of Cell Biology, 171(5):??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/171/5/755.2.

LeBrasseur:2005:PSP

[LeB05-76] Nicole LeBrasseur. p107 Skps from S phase. Journal of Cell Biology, 168(1):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/1/9.1.

LeBrasseur:2005:PA

[LeB05-77] Nicole LeBrasseur. Patterned by adhesion. Journal of Cell Biology, 170(1):??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/1/11.1.

LeBrasseur:2005:PPT

[LeB05-78] Nicole LeBrasseur. Polymerase with a protein template. Journal of Cell Biology, 171(2):??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/171/2/192.2.

LeBrasseur:2005:PLA

[LeB05-79] Nicole LeBrasseur. Presenting lipid antigens. Journal of Cell Biology, 171(2):??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/2/193.2.

LeBrasseur:2005:QSL

[LeB05-80] Nicole LeBrasseur. Quick stop for lymphocytes. Journal of Cell Biology, 169(3):??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/3/376.2.

LeBrasseur:2005:RR

[LeB05-81] Nicole LeBrasseur. Rec’d recombination. Journal of Cell Biology, 170(1):??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/1/10.3.

LeBrasseur:2005:RCN

[LeB05-82] Nicole LeBrasseur. Replication complexes in neuronal morphology. Journal of Cell Biology, 170(4):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/170/4/510.2.

LeBrasseur:2005:RSD

[LeB05-83] Nicole LeBrasseur. Restraining spindles in distress. Journal of Cell Biology, 168(7):??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/7/984.1.

LeBrasseur:2005:RDI

[LeB05-84] Nicole LeBrasseur. Ribosome does it all. Journal of Cell Biology, 168(4):??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/4/522.3.

LeBrasseur:2005:RAW

[LeB05-85] Nicole LeBrasseur. Ring around the wound. Journal of Cell Biology, 168(3):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/3/347.2.

LeBrasseur:2005:SCa

[LeB05-86] Nicole LeBrasseur. Secretion for cytokinesis. Journal of Cell Biology, 171(2):??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/2/192.1.

LeBrasseur:2005:SE

[LeB05-87] Nicole LeBrasseur. Selectivity for eIF4E. Journal of Cell Biology, 169(2):??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/2/213.3.

LeBrasseur:2005:SRL

[LeB05-88] Nicole LeBrasseur. Separating receptor and ligand. Journal of Cell Biology, 169(3):??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/3/377.1.

LeBrasseur:2005:STH

[LeB05-89] Nicole LeBrasseur. Signaling with tortoise and hare. Journal of Cell Biology, 171(4):??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/171/4/583.3.

LeBrasseur:2005:SSF

[LeB05-90] Nicole LeBrasseur. Simple sensing in fibroblasts. Journal of Cell Biology, 171(5):??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/5/754.2.

LeBrasseur:2005:SH

[LeB05-91] Nicole LeBrasseur. A Sisyphean helicase. Journal of Cell Biology, 171(4):??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/4/582.1.

LeBrasseur:2005:SMO

[LeB05-92] Nicole LeBrasseur. Sorting the mRNA overflow. Journal of Cell Biology, 169(6):??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/6/834.2.

LeBrasseur:2005:SR

[LeB05-93] Nicole LeBrasseur. Spines reach out. Journal of Cell Biology, 169(3):??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/3/377.2.

LeBrasseur:2005:SN

[LeB05-94] Nicole LeBrasseur. Stability in numbers. Journal of Cell Biology, 170(1):??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/1/11.2.

LeBrasseur:2005:SUP

[LeB05-95] Nicole LeBrasseur. Stiffening under pressure. Journal of Cell Biology, 169(5):??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/5/702.2.

LeBrasseur:2005:SCb

[LeB05-96] Nicole LeBrasseur. Stuck on a channel. Journal of Cell Biology, 171(3):??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/3/405.1.

LeBrasseur:2005:SPS

[LeB05-97] Nicole LeBrasseur. Supersensitive phospholipid sensing. Journal of Cell Biology, 168(4):??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/168/4/523.2.

LeBrasseur:2005:SI

[LeB05-98] Nicole LeBrasseur. Systematic import. Journal of Cell Biology, 168(7):??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/7/984.2.

LeBrasseur:2005:TAC

[LeB05-99] Nicole LeBrasseur. Telomeres avoid checkpoints. Journal of Cell Biology, 171(4):??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/171/4/583.1.

LeBrasseur:2005:TM

[LeB05-100] Nicole LeBrasseur. Thinning microtubules. Journal of Cell Biology, 168(1):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/1/9.2.

LeBrasseur:2005:TF

[LeB05-101] Nicole LeBrasseur. Transformed fusions. Journal of Cell Biology, 171(3):??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/3/404.1.

LeBrasseur:2005:TFC

[LeB05-102] Nicole LeBrasseur. Tumors feed on collagen. Journal of Cell Biology, 169(6):??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/6/835.3.

LeBrasseur:2005:TVB

[LeB05-103] Nicole LeBrasseur. Two VEGFs for big or branchy vessels. Journal of Cell Biology, 169(4):??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/169/4/549.1.

LeBrasseur:2005:TVP

[LeB05-104] Nicole LeBrasseur. Two vesicle pools for neurons. Journal of Cell Biology, 168(6):??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/6/850.3.

LeBrasseur:2005:VPO

[LeB05-105] Nicole LeBrasseur. A view of primitive organelles. Journal of Cell Biology, 170(5):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/5/698.1.

LeBrasseur:2005:WS

[LeB05-106] Nicole LeBrasseur. Weakness as a strength. Journal of Cell Biology, 168(6):??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/6/850.1.

LeBrasseur:2005:WBS

[LeB05-107] Nicole LeBrasseur. Wnt + BMP = stem cell. Journal of Cell Biology, 169(2):??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/2/212.1.

LeBrasseur:2005:WA

[LeB05-108] Nicole LeBrasseur. Wnt for adhesion. Journal of Cell Biology, 171(2):??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/2/193.1.

LeBrasseur:2005:YW

[LeB05-109] Nicole LeBrasseur. Yeast warfare. Journal of Cell Biology, 168(3):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/3/346.3.

LeBrasseur:2006:AKS

[LeB06a] Nicole LeBrasseur. Acetylate to kill or save. Journal of Cell Biology, 173(4):455–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/4/455.

LeBrasseur:2006:ASG

[LeB06b] Nicole LeBrasseur. Alien sensing by GC content. Journal of Cell Biology, 174(1):??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/1/4a.

LeBrasseur:2006:AOB

[LeB06c] Nicole LeBrasseur. APC’s order of business. Journal of Cell Biology, 172(4):??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/4/483b.

LeBrasseur:2006:AVE

[LeB06d] Nicole LeBrasseur. Axon’s vesicles in excess. Journal of Cell Biology, 174(1):??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/1/5b.

LeBrasseur:2006:BBP

[LeB06e] Nicole LeBrasseur. Big bones in p53 mutants. Journal of Cell Biology, 172(1):??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/1/2.2.

LeBrasseur:2006:BR

[LeB06f] Nicole LeBrasseur. BRCA1 in replication. Journal of Cell Biology, 175(5):676–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/5/676.

LeBrasseur:2006:CSR

[LeB06g] Nicole LeBrasseur. Cadherin stops receptor endocytosis. Journal of Cell Biology, 174(4):??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/174/4/475a.

LeBrasseur:2006:CF

[LeB06h] Nicole LeBrasseur. Cage folding. Journal of Cell Biology, 173(6):??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/6/826a.

LeBrasseur:2006:CCR

[LeB06i] Nicole LeBrasseur. Cell cycle regulators kick out centrosomes. Journal of Cell Biology, 174(6):??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/174/6/736a.

LeBrasseur:2006:CLD

[LeB06j] Nicole LeBrasseur. Cohesin loops DNA. Journal of Cell Biology, 174(4):475–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/4/475.

LeBrasseur:2006:CSW

[LeB06k] Nicole LeBrasseur. Coping strategies weaken cells. Journal of Cell Biology, 172(3):??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/172/3/325b.

LeBrasseur:2006:CC

[LeB06l] Nicole LeBrasseur. Cortex construction. Journal of Cell Biology, 175(3):??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/3/365b.

LeBrasseur:2006:DSB

[LeB06m] Nicole LeBrasseur. Damage stops bud growth. Journal of Cell Biology, 175(5):??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/5/677b.

LeBrasseur:2006:DN

[LeB06n] Nicole LeBrasseur. Defaulty neurons. Journal of Cell Biology, 172(1):??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/1/2.1.

LeBrasseur:2006:DC

[LeB06o] Nicole LeBrasseur. Direct is correct. Journal of Cell Biology, 172(7):956–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/7/956.

LeBrasseur:2006:DV

[LeB06p] Nicole LeBrasseur. Disorder for variety. Journal of Cell Biology, 173(6):826–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/6/826.

LeBrasseur:2006:DTO

[LeB06q] Nicole LeBrasseur. Dividing to their own beat. Journal of Cell Biology, 172(3):??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/3/324a.

LeBrasseur:2006:DRN

[LeB06r] Nicole LeBrasseur. DNA recruits nuclear envelope. Journal of Cell Biology, 173(4):??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/4/454a.

LeBrasseur:2006:DIY

[LeB06s] Nicole LeBrasseur. Do-it-yourself insertion. Journal of Cell Biology, 175(5):??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/5/677a.

LeBrasseur:2006:DCT

[LeB06t] Nicole LeBrasseur. Don’t cut too soon. Journal of Cell Biology, 173(3):316–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/3/316.

LeBrasseur:2006:EEP

[LeB06u] Nicole LeBrasseur. eIF4E exports proliferation mRNAs. Journal of Cell Biology, 175(3):??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/175/3/364a.

LeBrasseur:2006:EGP

[LeB06v] Nicole LeBrasseur. Entropy guards pore. Journal of Cell Biology, 174(1):5–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/1/5.

LeBrasseur:2006:EC

[LeB06w] Nicole LeBrasseur. Escaping a checkpoint. Journal of Cell Biology, 172(7):957–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/7/957.

LeBrasseur:2006:FIO

[LeB06x] Nicole LeBrasseur. Fight infection with oxidation. Journal of Cell Biology, 175(4):??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/175/4/518a.

LeBrasseur:2006:FS

[LeB06y] Nicole LeBrasseur. Flux without sliding. Journal of Cell Biology, 173(2):??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/2/146a.

LeBrasseur:2006:GPA

[LeB06z] Nicole LeBrasseur. Genes pulse with activity. Journal of Cell Biology, 173(6):827–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/6/827.

LeBrasseur:2006:HRC

[LeB06-27] Nicole LeBrasseur. An H1 to relax chromatin? Journal of Cell Biology, 175(6):??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/6/843a.

LeBrasseur:2006:HMG

[LeB06-28] Nicole LeBrasseur. Hedgehog makes giant brains. Journal of Cell Biology, 173(1):??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/1/5a.

LeBrasseur:2006:HP

[LeB06-29] Nicole LeBrasseur. Hidden particles. Journal of Cell Biology, 174(6):737–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/6/737.

LeBrasseur:2006:HEP

[LeB06-30] Nicole LeBrasseur. Histone exchange program. Journal of Cell Biology, 175(3):365–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/175/3/365.

LeBrasseur:2006:HB

[LeB06-31] Nicole LeBrasseur. Hoarding bug. Journal of Cell Biology, 173(4):454–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/4/454.

LeBrasseur:2006:IPC

[LeB06-32] Nicole LeBrasseur. Insulin pathway clears aggregates. Journal of Cell Biology, 172(5):??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/172/5/639b.

LeBrasseur:2006:IM

[LeB06-33] Nicole LeBrasseur. Integrins for mitosis. Journal of Cell Biology, 174(4):??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/4/474a.

LeBrasseur:2006:IGI

[LeB06-34] Nicole LeBrasseur. An integrin’s gut instincts. Journal of Cell Biology, 175(3):??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/3/365a.

LeBrasseur:2006:KPE

[LeB06-35] Nicole LeBrasseur. Kar’ted to plus ends. Journal of Cell Biology, 172(1):??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/1/3.3.

LeBrasseur:2006:KSN

[LeB06-36] Nicole LeBrasseur. Kinesin switch in neurites. Journal of Cell Biology, 173(1):5–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/1/5.

LeBrasseur:2006:LG

[LeB06-37] Nicole LeBrasseur. Lipids be gone. Journal of Cell Biology, 175 (3):364–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/3/364.

LeBrasseur:2006:MC

[LeB06-38] Nicole LeBrasseur. Malignancy from cooperation. Journal of Cell Biology, 174(7):??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/174/7/908a.

LeBrasseur:2006:MSM

[LeB06-39] Nicole LeBrasseur. A MAP’s stabilizing mechanism. Journal of Cell Biology, 172(7):??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/7/956a.

LeBrasseur:2006:MID

[LeB06-40] Nicole LeBrasseur. Migrating interphase DNA. Journal of Cell Biology, 173(3):??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/3/317a.

LeBrasseur:2006:MMM

[LeB06-41] Nicole LeBrasseur. Milton for motile mitochondria. Journal of Cell Biology, 173(4):??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/4/455a.

LeBrasseur:2006:MEC

[LeB06-42] Nicole LeBrasseur. A minus end checkpoint. Journal of Cell Biology, 175(4):519–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/4/519.

LeBrasseur:2006:MM

[LeB06-43] Nicole LeBrasseur. Monooriented in the middle. Journal of Cell Biology, 172(4):??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/4/482a.

LeBrasseur:2006:MLE

[LeB06-44] Nicole LeBrasseur. Motherly loss ends cell cycle. Journal of Cell Biology, 172(3):324–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/172/3/324.

LeBrasseur:2006:MHB

[LeB06-45] Nicole LeBrasseur. mRNAs held back. Journal of Cell Biology, 172(3):??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/3/325a.

LeBrasseur:2006:MRP

[LeB06-46] Nicole LeBrasseur. A myosin for RNA Pol II. Journal of Cell Biology, 174(7):909–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/174/7/909.

LeBrasseur:2006:MNS

[LeB06-47] Nicole LeBrasseur. Myosin’s need for speed. Journal of Cell Biology, 175(4):??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/4/519b.

LeBrasseur:2006:NC

[LeB06-48] Nicole LeBrasseur. Neuronal corridors. Journal of Cell Biology, 173(3):317–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/3/317.

LeBrasseur:2006:NSN

[LeB06-49] Nicole LeBrasseur. Neurons survive with NCS-1. Journal of Cell Biology, 172(7):??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/7/957b.

LeBrasseur:2006:NFM

[LeB06-50] Nicole LeBrasseur. A new function for mitochondria. Journal of Cell Biology, 173(1):??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/1/4a.

LeBrasseur:2006:NIR

[LeB06-51] Nicole LeBrasseur. Newt insight into regeneration. Journal of Cell Biology, 172(3):325–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/172/3/325.

LeBrasseur:2006:NFF

[LeB06-52] Nicole LeBrasseur. The Nuclear Family: Functional Organization of the Cell Nucleus Prague, Czech Republic, May 5–8, 2006. Journal of Cell Biology, 174(4):478–??, August 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/174/4/478.

LeBrasseur:2006:PCC

[LeB06-53] Nicole LeBrasseur. Par-3 for the cofilin course. Journal of Cell Biology, 172(5):??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/5/639a.

LeBrasseur:2006:PE

[LeB06-54] Nicole LeBrasseur. Partitioning endosomes. Journal of Cell Biology, 175(6):??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/6/843b.

LeBrasseur:2006:PDS

[LeB06-55] Nicole LeBrasseur. Peroxide detection system. Journal of Cell Biology, 175(5):??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/5/676a.

LeBrasseur:2006:PGT

[LeB06-56] Nicole LeBrasseur. Phosphate give and take. Journal of Cell Biology, 174(1):??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/1/5a.

LeBrasseur:2006:PDC

[LeB06-57] Nicole LeBrasseur. PKC decodes Ca^{2+}. Journal of Cell Biology, 174(4):474–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/4/474.

LeBrasseur:2006:PGE

[LeB06-58] Nicole LeBrasseur. Pro and Gly for elasticity. Journal of Cell Biology, 175(6):843–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/175/6/843.

LeBrasseur:2006:PEN

[LeB06-59] Nicole LeBrasseur. Promoting ESCRT–I narcissism. Journal of Cell Biology, 175(5):677–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/175/5/677.

LeBrasseur:2006:PMR

[LeB06-60] Nicole LeBrasseur. Putting mitosis in reverse. Journal of Cell Biology, 173(3):??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/3/316a.

LeBrasseur:2006:RRM

[LeB06-61] Nicole LeBrasseur. Repair road map. Journal of Cell Biology, 173(2):147–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/2/147.

LeBrasseur:2006:RMB

[LeB06-62] Nicole LeBrasseur. The Repo–Man bringeth. Journal of Cell Biology, 172(5):??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/5/638a.

LeBrasseur:2006:SCB

[LeB06-63] Nicole LeBrasseur. Selective calcium blockade. Journal of Cell Biology, 172(1):??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/1/3.2.

LeBrasseur:2006:SKO

[LeB06-64] Nicole LeBrasseur. Serum keeps order in the wound. Journal of Cell Biology, 172(7):??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/7/957a.

LeBrasseur:2006:STI

[LeB06-65] Nicole LeBrasseur. The shape of things to ingest. Journal of Cell Biology, 173(1):4–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/1/4.

LeBrasseur:2006:SCG

[LeB06-66] Nicole LeBrasseur. Sights on cone geometry. Journal of Cell Biology, 174(1):4–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/1/4.

LeBrasseur:2006:SKb

[LeB06-67] Nicole LeBrasseur. Signaling from keratins. Journal of Cell Biology, 173(6):??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/6/827b.

LeBrasseur:2006:SAF

[LeB06-68] Nicole LeBrasseur. Sizing up actin fences. Journal of Cell Biology, 174(6):??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/6/737a.

LeBrasseur:2006:SC

[LeB06-69] Nicole LeBrasseur. Specialized chaperones. Journal of Cell Biology, 172(4):483–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/4/483.

LeBrasseur:2006:SKa

[LeB06-70] Nicole LeBrasseur. Spill and kill. Journal of Cell Biology, 173 (2):146–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/2/146.

LeBrasseur:2006:SSD

[LeB06-71] Nicole LeBrasseur. Spines slow dendrite diffusion. Journal of Cell Biology, 175(6):842–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/175/6/842.

LeBrasseur:2006:SGC

[LeB06-72] Nicole LeBrasseur. Spitz on growth control. Journal of Cell Biology, 175(4):518–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/4/518.

LeBrasseur:2006:SM

[LeB06-73] Nicole LeBrasseur. Spreading mitochondria. Journal of Cell Biology, 172(4):482–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/4/482.

LeBrasseur:2006:SH

[LeB06-74] Nicole LeBrasseur. Stealing histones. Journal of Cell Biology, 174(6):??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/6/737b.

LeBrasseur:2006:SCI

[LeB06-75] Nicole LeBrasseur. STIMulating Ca^{2+} influx. Journal of Cell Biology, 174(6):736–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/174/6/736.

LeBrasseur:2006:SFI

[LeB06-76] Nicole LeBrasseur. Stomata fight infection. Journal of Cell Biology, 174(7):??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/7/909b.

LeBrasseur:2006:CTK

[LeB06-77] Nicole LeBrasseur. T cells told to keep rolling. Journal of Cell Biology, 174(7):??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/7/909a.

LeBrasseur:2006:TTT

[LeB06-78] Nicole LeBrasseur. Tandem transcript translation. Journal of Cell Biology, 172(5):638–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/172/5/638.

LeBrasseur:2006:TM

[LeB06-79] Nicole LeBrasseur. Titinless mice. Journal of Cell Biology, 173(4):??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/4/455b.

LeBrasseur:2006:ENO

[LeB06-80] Nicole LeBrasseur. To each nucleus its own ER. Journal of Cell Biology, 173(2):??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/2/147a.

LeBrasseur:2006:TBH

[LeB06-81] Nicole LeBrasseur. A transcriptional black hole. Journal of Cell Biology, 174(7):908–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/174/7/908.

LeBrasseur:2006:TWI

[LeB06-82] Nicole LeBrasseur. Two ways for insertion. Journal of Cell Biology, 173(3):??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/3/317b.

LeBrasseur:2006:VSS

[LeB06-83] Nicole LeBrasseur. Virus squeezed shut. Journal of Cell Biology, 173(6):??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/6/827a.

LeBrasseur:2006:WPRa

[LeB06-84] Nicole LeBrasseur. Who Pix up Rac1? Journal of Cell Biology, 172(5):639–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/5/639.

LeBrasseur:2006:WPRb

[LeB06-85] Nicole LeBrasseur. Wnt polarity relay. Journal of Cell Biology, 175(6):??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/6/842a.

LeBrasseur:2006:XS

[LeB06-86] Nicole LeBrasseur. Xorbiting the spindle. Journal of Cell Biology, 172(1):??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/1/3.1.

LeBrasseur:2006:XM

[LeB06-87] Nicole LeBrasseur. XX meetings. Journal of Cell Biology, 172 (4):??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/4/483a.

LeBrasseur:2006:YPU

[LeB06-88] Nicole LeBrasseur. Yeast’s primitive urea. Journal of Cell Biology, 175(4):??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/4/519a.

LeBrasseur:2007:AFT

[LeB07a] Nicole LeBrasseur. Acrobatic flexibility of the translocon. Journal of Cell Biology, 179(7):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/7/1326.1.

LeBrasseur:2007:AMH

[LeB07b] Nicole LeBrasseur. Actin mesh hinders microtubules. Journal of Cell Biology, 179(4):??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/4/570.1.

LeBrasseur:2007:APO

[LeB07c] Nicole LeBrasseur. Apoptosis prevents osteoporosis. Journal of Cell Biology, 179(1):??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/1/5.3.

LeBrasseur:2007:AIM

[LeB07d] Nicole LeBrasseur. Apoptotic inaccessibility with maturity. Journal of Cell Biology, 179(5):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/5/806.1.

LeBrasseur:2007:BRM

[LeB07e] Nicole LeBrasseur. Bug redecorates membrane. Journal of Cell Biology, 177(1):??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/1/3a.

LeBrasseur:2007:CGD

[LeB07f] Nicole LeBrasseur. Catastrophe gradient drives anaphase B. Journal of Cell Biology, 177(6):??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/177/6/944a.

LeBrasseur:2007:CFD

[LeB07g] Nicole LeBrasseur. Cell fate depends on Golgi. Journal of Cell Biology, 177(3):??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/3/373b.

LeBrasseur:2007:CI

[LeB07h] Nicole LeBrasseur. Centriolar inequality. Journal of Cell Biology, 177(1):3–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/1/3.

LeBrasseur:2007:CLE

[LeB07i] Nicole LeBrasseur. Chromosomes leave envelope for karyosome. Journal of Cell Biology, 179(5):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/5/807.1.

LeBrasseur:2007:DD

[LeB07j] Nicole LeBrasseur. Death by default. Journal of Cell Biology, 176(6):??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/6/732a.

LeBrasseur:2007:DNM

[LeB07k] Nicole LeBrasseur. A dual need for myosin in wounds. Journal of Cell Biology, 176(1):3–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/176/1/3.

LeBrasseur:2007:DST

[LeB07l] Nicole LeBrasseur. Dynein’s spindly trip. Journal of Cell Biology, 177(6):??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/6/945b.

LeBrasseur:2007:ERS

[LeB07m] Nicole LeBrasseur. Early replication for short telomeres. Journal of Cell Biology, 177(2):187–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/177/2/187.

LeBrasseur:2007:EHC

[LeB07n] Nicole LeBrasseur. Emerin hooks centrosome to nucleus. Journal of Cell Biology, 178(6):??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/178/6/891.2.

LeBrasseur:2007:EPN

[LeB07o] Nicole LeBrasseur. Endocytosis pulls Notch apart. Journal of Cell Biology, 176(4):372–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/176/4/372.

LeBrasseur:2007:FRA

[LeB07p] Nicole LeBrasseur. Females’ regenerating advantage. Journal of Cell Biology, 177(1):2–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/1/2.

LeBrasseur:2007:GGS

[LeB07q] Nicole LeBrasseur. Glycan growth switch. Journal of Cell Biology, 177(3):373–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/3/373.

LeBrasseur:2007:GHS

[LeB07r] Nicole LeBrasseur. Golgi helps shape dendrites. Journal of Cell Biology, 178(7):??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/7/1096.1.

LeBrasseur:2007:GFA

[LeB07s] Nicole LeBrasseur. Growth factors attract mRNAs. Journal of Cell Biology, 178(6):??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/178/6/890.1.

LeBrasseur:2007:HPI

[LeB07t] Nicole LeBrasseur. Hidde Ploegh: Immunologist, journeyman. Journal of Cell Biology, 179(3):364–??, November 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/179/3/364.

LeBrasseur:2007:HHP

[LeB07u] Nicole LeBrasseur. Howard Hang: Posttranslational pathogenesis. Journal of Cell Biology, 178(3):340–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/178/3/340.

LeBrasseur:2007:HFF

[LeB07v] Nicole LeBrasseur. Hydin flips flagella. Journal of Cell Biology, 176(4):??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/4/373b.

LeBrasseur:2007:HRR

[LeB07w] Nicole LeBrasseur. Hypoxic reaction to reactive oxygen. Journal of Cell Biology, 177(6):??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/177/6/945a.

LeBrasseur:2007:IRC

[LeB07x] Nicole LeBrasseur. Immunity raises cholesterol. Journal of Cell Biology, 177(3):??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/3/373a.

LeBrasseur:2007:IMP

[LeB07y] Nicole LeBrasseur. Independent membrane protein clustering. Journal of Cell Biology, 178(7):??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/178/7/1096.2.

LeBrasseur:2007:JYC

[LeB07z] Nicole LeBrasseur. Junying Yuan: Changing avenues without losing focus. Journal of Cell Biology, 179(2):174–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/2/
174.

LeBrasseur:2007:KCS

[LeB07-27] Nicole LeBrasseur. Katanin cuts some, spares others. Journal of Cell Biology, 178(6):??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/178/6/891.1.

LeBrasseur:2007:KRA

[LeB07-28] Nicole LeBrasseur. Kinesin-5 reins in axon branching. Journal of Cell Biology, 178(6):??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/178/6/890.2.

LeBrasseur:2007:LDG

[LeB07-29] Nicole LeBrasseur. Linking division and growth. Journal of Cell Biology, 177(3):??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/3/372a.

LeBrasseur:2007:MAE

[LeB07-30] Nicole LeBrasseur. MHC and antigen in exosomes. Journal of Cell Biology, 179(1):??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/1/4.1.

LeBrasseur:2007:MSD

[LeB07-31] Nicole LeBrasseur. Mirror symmetry in division. Journal of Cell Biology, 177(2):186–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/2/186.

LeBrasseur:2007:MMI

[LeB07-32] Nicole LeBrasseur. Mitochondria move to immune synapse. Journal of Cell Biology, 178(7):??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/178/7/1097.1.

LeBrasseur:2007:MSA

[LeB07-33] Nicole LeBrasseur. Motility from self-antigens. Journal of Cell Biology, 177(3):372–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/3/372.

LeBrasseur:2007:MI

[LeB07-34] Nicole LeBrasseur. Motility without invasion. Journal of Cell Biology, 179(7):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/7/1327.1.

LeBrasseur:2007:MBS

[LeB07-35] Nicole LeBrasseur. A myosin for basolateral sorting. Journal of Cell Biology, 177(1):??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/1/3b.

LeBrasseur:2007:MPM

[LeB07-36] Nicole LeBrasseur. Myosin paces on microtubules. Journal of Cell Biology, 176(6):??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/6/733a.

LeBrasseur:2007:NBC

[LeB07-37] Nicole LeBrasseur. Nervous blood cells. Journal of Cell Biology, 179(1):??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/1/5.2.

LeBrasseur:2007:NNA

[LeB07-38] Nicole LeBrasseur. New nucleator for actin. Journal of Cell Biology, 179(4):??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/4/571.3.

LeBrasseur:2007:NRJ

[LeB07-39] Nicole LeBrasseur. Notch remodels junctions. Journal of Cell Biology, 177(1):??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/1/2a.

LeBrasseur:2007:PPP

[LeB07-40] Nicole LeBrasseur. ParM pushes plasmids apart. Journal of Cell Biology, 179(5):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/5/806.2.

LeBrasseur:2007:PKA

[LeB07-41] Nicole LeBrasseur. Partners in kinesin activation. Journal of Cell Biology, 176(1):2–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/176/1/2.

LeBrasseur:2007:PFR

[LeB07-42] Nicole LeBrasseur. Peptidase frees receptors in endosomes. Journal of Cell Biology, 179(5):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/5/807.2.

LeBrasseur:2007:PM

[LeB07-43] Nicole LeBrasseur. Pinched off membrane. Journal of Cell Biology, 176(4):??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/4/372a.

LeBrasseur:2007:PSG

[LeB07-44] Nicole LeBrasseur. PKA steers growth cones. Journal of Cell Biology, 176(1):??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/1/2a.

LeBrasseur:2007:PHH

[LeB07-45] Nicole LeBrasseur. Plant hormone is human cytokine. Journal of Cell Biology, 177(2):??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/2/187b.

LeBrasseur:2007:PPL

[LeB07-46] Nicole LeBrasseur. Platelets’ preset lifespan. Journal of Cell Biology, 177(2):??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/2/186a.

LeBrasseur:2007:PSO

[LeB07-47] Nicole LeBrasseur. The pore slides open. Journal of Cell Biology, 177(2):??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/2/187a.

LeBrasseur:2007:PDM

[LeB07-48] Nicole LeBrasseur. Pro-diversity mitofusins. Journal of Cell Biology, 176(4):??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/4/373a.

LeBrasseur:2007:PEP

[LeB07-49] Nicole LeBrasseur. Promoters enter plant–pathogen fray. Journal of Cell Biology, 179(4):??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/4/570.2.

LeBrasseur:2007:PSD

[LeB07-50] Nicole LeBrasseur. Protein shields for dehydration. Journal of Cell Biology, 179(4):??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/4/571.2.

LeBrasseur:2007:PES

[LeB07-51] Nicole LeBrasseur. Proteins on the edge of stability. Journal of Cell Biology, 179(1):??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/1/4.2.

LeBrasseur:2007:PPB

[LeB07-52] Nicole LeBrasseur. Pyk2 for perfect bones. Journal of Cell Biology, 178(6):??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/6/891.3.

LeBrasseur:2007:RWF

[LeB07-53] Nicole LeBrasseur. Rebecca Worthylake: The flooding of science. Journal of Cell Biology, 178(1):6–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/178/1/6.

LeBrasseur:2007:SDR

[LeB07-54] Nicole LeBrasseur. Scarred by a death receptor. Journal of Cell Biology, 177(6):944–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/6/944.

LeBrasseur:2007:SMA

[LeB07-55] Nicole LeBrasseur. Slow-moving Alzheimer’s. Journal of Cell Biology, 179(5):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/5/807.3.

LeBrasseur:2007:SSC

[LeB07-56] Nicole LeBrasseur. Special skin cells get pigments. Journal of Cell Biology, 179(1):??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/1/5.1.

LeBrasseur:2007:SGJ

[LeB07-57] Nicole LeBrasseur. Stick with gap junctions. Journal of Cell Biology, 178(7):??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/7/1097.2.

LeBrasseur:2007:SFM

[LeB07-58] Nicole LeBrasseur. Sticky fingers for migration. Journal of Cell Biology, 176(6):733–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/6/733.

LeBrasseur:2007:SSO

[LeB07-59] Nicole LeBrasseur. Stress stunts organ growth. Journal of Cell Biology, 176(6):732–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/6/732.

LeBrasseur:2007:SSR

[LeB07-60] Nicole LeBrasseur. Syt-12 for spontaneous release. Journal of Cell Biology, 176(1):??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/176/1/3a.

LeBrasseur:2007:TLC

[LeB07-61] Nicole LeBrasseur. Terry Lechler: The cytoskeleton is skin deep. Journal of Cell Biology, 178(4):546–??, August 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/178/4/546.

LeBrasseur:2007:BWN

[LeB07-62] Nicole LeBrasseur. To bud where no bud has gone before. Journal of Cell Biology, 179(7):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/7/1326.2.

LeBrasseur:2007:TTD

[LeB07-63] Nicole LeBrasseur. Translation turns on dendritic cells. Journal of Cell Biology, 179(7):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/7/1327.2.

LeBrasseur:2007:TMA

[LeB07-64] Nicole LeBrasseur. Turning motor into anchor. Journal of Cell Biology, 179(4):??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/4/571.1.

LeBrasseur:2007:UC

[LeB07-65] Nicole LeBrasseur. The UPR for cytokinesis. Journal of Cell Biology, 177(6):945–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/6/945.

LeBrasseur:2007:VFT

[LeB07-66] Nicole LeBrasseur. Vesicles fall for traps. Journal of Cell Biology, 176(4):373–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/4/373.

LeBrasseur:2007:VIF

[LeB07-67] Nicole LeBrasseur. Viruses infect filopodia. Journal of Cell Biology, 176(6):??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/6/733b.

LeBrasseur:2008:AKC

[LeB08a] Nicole LeBrasseur. An adhesion kinase on centrosomes. Journal of Cell Biology, 180(4):??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/4/649.1.

LeBrasseur:2008:AAM

[LeB08b] Nicole LeBrasseur. Alan Aderem: From molecules to megabytes. Journal of Cell Biology, 180(2):252–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/180/2/

LeBrasseur:2008:CCC

[LeB08c] Nicole LeBrasseur. Channel changes customers. Journal of Cell Biology, 181(3):??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/3/401.2.

LeBrasseur:2008:CKF

[LeB08d] Nicole LeBrasseur. Cleaning the kidney’s filter. Journal of Cell Biology, 180(3):??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/3/447.1.

LeBrasseur:2008:DSY

[LeB08e] Nicole LeBrasseur. DNA spring in your spindle. Journal of Cell Biology, 180(3):??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/3/446.1.

LeBrasseur:2008:EOS

[LeB08f] Nicole LeBrasseur. Early origins sound S phase alarm. Journal of Cell Biology, 180(6):??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/6/1053.2.

LeBrasseur:2008:EP

[LeB08g] Nicole LeBrasseur. An eIF for proliferation. Journal of Cell Biology, 181(2):??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/2/175.2.

LeBrasseur:2008:ESM

[LeB08h] Nicole LeBrasseur. Enzyme shapes mitochondria. Journal of Cell Biology, 181(1):??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/1/5.2.

LeBrasseur:2008:ERA

[LeB08i] Nicole LeBrasseur. ESCRT rings around the membrane. Journal of Cell Biology, 180(2):??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/2/248.2.

LeBrasseur:2008:HCE

[LeB08j] Nicole LeBrasseur. Hungry cells eat ribosomes. Journal of Cell Biology, 181(3):??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/3/400.2.

LeBrasseur:2008:ISM

[LeB08k] Nicole LeBrasseur. ICAP for sensing matrix density. Journal of Cell Biology, 180(2):??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/2/249.1.

LeBrasseur:2008:LFB

[LeB08l] Nicole LeBrasseur. Lamellipodia fight back formin. Journal of Cell Biology, 180(6):??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/6/1052.1.

LeBrasseur:2008:LPP

[LeB08m] Nicole LeBrasseur. Lamins position PCNA. Journal of Cell Biology, 181(2):??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/2/174.2.

LeBrasseur:2008:LOC

[LeB08n] Nicole LeBrasseur. Losing odd chromosomes. Journal of Cell Biology, 180(4):??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/4/649.2.

LeBrasseur:2008:MTG

[LeB08o] Nicole LeBrasseur. Matching traffic and growth. Journal of Cell Biology, 180(4):??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/4/649.3.

LeBrasseur:2008:MTW

[LeB08p] Nicole LeBrasseur. Motor tug-of-war is productive. Journal of Cell Biology, 181(1):??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/1/4.2.

LeBrasseur:2008:MSC

[LeB08q] Nicole LeBrasseur. mTOR sends T cells on their way. Journal of Cell Biology, 181(3):??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/3/401.3.

LeBrasseur:2008:MKC

[LeB08r] Nicole LeBrasseur. Myosin and kinesin collaborate. Journal of Cell Biology, 181(1):??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/1/5.3.

LeBrasseur:2008:MCZ

[LeB08s] Nicole LeBrasseur. Myosin chaperones: from A to Z and back. Journal of Cell Biology, 180(6):??, March 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/180/6/1052.2.

LeBrasseur:2008:NSH

[LeB08t] Nicole LeBrasseur. Nanotubes spread HIV. Journal of Cell Biology, 180(3):??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/3/446.2.

LeBrasseur:2008:PBR

[LeB08u] Nicole LeBrasseur. Phosphorylation for break repair. Journal of Cell Biology, 181(2):??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/2/175.1.

LeBrasseur:2008:PMG

[LeB08v] Nicole LeBrasseur. Plasma membrane in Golgi costume. Journal of Cell Biology, 180(5):??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/5/848.1.

LeBrasseur:2008:PF

[LeB08w] Nicole LeBrasseur. Polarizing for furrows. Journal of Cell Biology, 181(2):??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/2/174.1.

LeBrasseur:2008:PAA

[LeB08x] Nicole LeBrasseur. Pseudokinase is active after all. Journal of Cell Biology, 181(3):??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/3/401.1.

LeBrasseur:2008:QHP

[LeB08y] Nicole LeBrasseur. Quick hops on and off plus ends. Journal of Cell Biology, 180(4):??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/4/648.2.

LeBrasseur:2008:RTM

[LeB08z] Nicole LeBrasseur. Rb turns up mitochondria. Journal of Cell Biology, 180(5):??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/5/848.2.

LeBrasseur:2008:SCH

[LeB08-27] Nicole LeBrasseur. Septins clear high speed tracks. Journal of Cell Biology, 180(2):??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/2/248.1.

LeBrasseur:2008:SNR

[LeB08-28] Nicole LeBrasseur. Shuffling of neuronal receptors. Journal of Cell Biology, 181(3):??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/3/400.1.

LeBrasseur:2008:SC

[LeB08-29] Nicole LeBrasseur. Sopping up chemokine. Journal of Cell Biology, 180(5):??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/5/849.1.

LeBrasseur:2008:SAM

[LeB08-30] Nicole LeBrasseur. Spine actin for memory. Journal of Cell Biology, 181(1):??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/1/4.1.

LeBrasseur:2008:SCF

[LeB08-31] Nicole LeBrasseur. Spokes coordinate flagella. Journal of Cell Biology, 180(2):??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/2/249.2.

LeBrasseur:2008:SFT

[LeB08-32] Nicole LeBrasseur. The SR finds the translocon. Journal of Cell Biology, 180(6):??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/6/1053.1.

LeBrasseur:2008:SCS

[LeB08-33] Nicole LeBrasseur. Stem cells sport longest telomeres. Journal of Cell Biology, 181(1):??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/1/5.1.

LeBrasseur:2008:TCS

[LeB08-34] Nicole LeBrasseur. Trafficking from cell surface to nuclear envelope. Journal of Cell Biology, 180(4):??, February 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/180/4/648.1.

LeBrasseur:2008:TFT

[LeB08-35] Nicole LeBrasseur. Transcription factor translated in axons. Journal of Cell Biology, 180(3):??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/180/3/447.2.

LeBrasseur:2008:VLE

[LeB08-36] Nicole LeBrasseur. Variety in leading edge actin. Journal of Cell Biology, 180(5):??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/5/849.2.

LeBrasseur:2009:HUA

[LeB09a] Nicole LeBrasseur. HIV uses autophagy for its own means. Journal of Cell Biology, 186(2):165–??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/186/2/165.

LeBrasseur:2009:IWW

[LeB09b] Nicole LeBrasseur. Irving Weissman: Working for regenerative medicine. Journal of Cell Biology, 187(3):316–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/187/3/
316.

LeBrasseur:2009:LRL

[LeB09c] Nicole LeBrasseur. Louis Reichardt: the long climb to science’s summits. Journal of Cell Biology, 186(5):634–??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/186/5/
634.

Lechler:2008:LLN

[Lec08] Terry Lechler. Limiting lumens: a new role for Cdc42. Journal of Cell Biology, 183(4):575–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/183/4/575.

LaTerra:2005:NCA

[LEH^{+}05] Sabrina La Terra, Christopher N. English, Polla Hergert, Bruce F. McEwen, Greenfield Sluder, and Alexey Khodjakov. The de novo centriole assembly pathway in HeLa cells. Journal of Cell Biology, 168(5):713–??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/168/5/713.

Leslie:2005:EID

[Les05a] Mitch Leslie. EGF is internalized and degraded. Journal of Cell Biology, 170(3):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/3/339.2.

Leslie:2005:IS

[Les05b] Mitch Leslie. Isolating SRP. Journal of Cell Biology, 171 (1):13–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/1/13.

Leslie:2005:INL

[Les05c] Mitch Leslie. The isolation of the nuclear lamina. Journal of Cell Biology, 170(6):865–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/170/6/865.

Leslie:2005:LT

[Les05d] Mitch Leslie. Lost in translation. Journal of Cell Biology, 170 (3):338–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/3/338.

Leslie:2005:MTO

[Les05e] Mitch Leslie. More than one way to attach. Journal of Cell Biology, 171(3):??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/3/409.1.

Leslie:2005:PNP

[Les05f] Mitch Leslie. A portrait of the nuclear pore complex. Journal of Cell Biology, 171(2):??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/171/2/194.2.

Leslie:2005:PG

[Les05g] Mitch Leslie. Powered by gel. Journal of Cell Biology, 170(3):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
170/3/339.1.

Leslie:2005:RYO

[Les05h] Mitch Leslie. Roll-your-own endothelial tubes. Journal of Cell Biology, 171(2):??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/2/194.1.

Leslie:2005:TO

[Les05i] Mitch Leslie. Tagging an organelle. Journal of Cell Biology, 171 (2):195–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/2/195.

Leslie:2006:ABW

[Les06a] Mitch Leslie. The action behind the words: embryonic stem cell research marches on. Journal of Cell Biology, 174(6):

743–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/6/743.

Leslie:2006:AND

[Les06b] Mitch Leslie. Active neuronal death. Journal of Cell Biology, 172(4):??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/4/485a.

Leslie:2006:BE

[Les06c] Mitch Leslie. Brief encounter. Journal of Cell Biology, 175 (1):3–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/1/3.

Leslie:2006:CTS

[Les06d] Mitch Leslie. Cadherin as a tumor suppressor. Journal of Cell Biology, 172(4):485–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/172/4/485.

Leslie:2006:CC

[Les06e] Mitch Leslie. Centrosome choreography. Journal of Cell Biology, 172(3):328–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/3/328.

Leslie:2006:CLC

[Les06f] Mitch Leslie. Cyclin localization controls activity. Journal of Cell Biology, 172(7):961–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/172/7/961.

Leslie:2006:DL

[Les06g] Mitch Leslie. Division of labor. Journal of Cell Biology, 175 (1):??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/1/3a.

Leslie:2006:DC

[Les06h] Mitch Leslie. Dying on cue. Journal of Cell Biology, 173 (2):150–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/2/150.

Leslie:2006:EDF

[Les06i] Mitch Leslie. ECM determines fate. Journal of Cell Biology, 173(1):??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/1/6a.

Leslie:2006:EBR

[Les06j] Mitch Leslie. Endosomes bring up the rear. Journal of Cell Biology, 175(2):??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/2/194a.

Leslie:2006:GES

[Les06k] Mitch Leslie. Getting the embryo into shape. Journal of Cell Biology, 175(2):195–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/2/195.

Leslie:2006:HDL

[Les06l] Mitch Leslie. Hold on for dear life. Journal of Cell Biology, 173 (2):151–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/2/151.

Leslie:2006:HS

[Les06m] Mitch Leslie. The Hunting of the snRNP. Journal of Cell Biology, 173(1):6–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/1/6.

Leslie:2006:VNI

[Les06n] Mitch Leslie. In vitro nuclear import. Journal of Cell Biology, 172(6):789–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/6/789.

Leslie:2006:IST

[Les06o] Mitch Leslie. Integrin signal transduction. Journal of Cell Biology, 173(2):??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/2/150a.

Leslie:2006:KPF

[Les06p] Mitch Leslie. Karyopherins play the field. Journal of Cell Biology, 175(4):??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/4/516a.

Leslie:2006:LM

[Les06q] Mitch Leslie. Lining up microtubules. Journal of Cell Biology, 175(4):??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/4/517a.

Leslie:2006:MT

[Les06r] Mitch Leslie. Making tendons. Journal of Cell Biology, 172 (2):167–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/2/167.

Leslie:2006:MBb

[Les06s] Mitch Leslie. Melanin’s backups. Journal of Cell Biology, 175(2):??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/2/195b.

Leslie:2006:MBa

[Les06t] Mitch Leslie. Membrane backtracking. Journal of Cell Biology, 175(2):??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/2/195a.

Leslie:2006:MMR

[Les06u] Mitch Leslie. Molecular mothballs for RNA. Journal of Cell Biology, 175(4):??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/4/517b.

Leslie:2006:C

[Les06v] Mitch Leslie. Not so close. Journal of Cell Biology, 175(1):??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
175/1/2a.

Leslie:2006:PPC

[Les06w] Mitch Leslie. Passenger proteins check in. Journal of Cell Biology, 172(3):329–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/3/329.

Leslie:2006:PLS

[Les06x] Mitch Leslie. A protein long shot. Journal of Cell Biology, 175 (1):2–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/1/2.

Leslie:2006:RAT

[Les06y] Mitch Leslie. Ring around the tubule. Journal of Cell Biology, 175(1):??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/1/3b.

Leslie:2006:SC

[Les06z] Mitch Leslie. Skeleton crew. Journal of Cell Biology, 172 (7):960–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/7/960.

Leslie:2006:TTN

[Les06-27] Mitch Leslie. Taking a Toll on neurons. Journal of Cell Biology, 175(2):194–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/2/194.

Leslie:2006:TMS

[Les06-28] Mitch Leslie. Togetherness, molecular style. Journal of Cell Biology, 175(4):516–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/4/516.

Leslie:2006:UDC

[Les06-29] Mitch Leslie. Untangling a deadly conspiracy. Journal of Cell Biology, 175(4):517–??, November 2006. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/175/4/517.

Leslie:2007:ANM

[Les07a] Mitch Leslie. AIS neurofascin’d to matrix. Journal of Cell Biology, 178(5):??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/5/717.1.

Leslie:2007:APS

[Les07b] Mitch Leslie. Apoptosis promoters spur cell division. Journal of Cell Biology, 178(6):??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/178/6/893.2.

Leslie:2007:BOM

[Les07c] Mitch Leslie. Blinded by old macrophages. Journal of Cell Biology, 179(5):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/5/809.1.

Leslie:2007:BAR

[Les07d] Mitch Leslie. Breaking actin to rebuild it. Journal of Cell Biology, 177(3):??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/3/371a.

Leslie:2007:CKC

[Les07e] Mitch Leslie. Caching killer carbs. Journal of Cell Biology, 178(3):338–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/3/338.

Leslie:2007:CCS

[Les07f] Mitch Leslie. Cancer cells straighten out. Journal of Cell Biology, 179(4):??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/4/569.3.

Leslie:2007:CFM

[Les07g] Mitch Leslie. Caspase-free mitotic death. Journal of Cell Biology, 178(2):181–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/2/181.

Leslie:2007:CNM

[Les07h] Mitch Leslie. Casting NETs for microbes. Journal of Cell Biology, 176(2):128–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/2/128.

Leslie:2007:CM

[Les07i] Mitch Leslie. Caught in the middle. Journal of Cell Biology, 176(3):??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/3/244a.

Leslie:2007:CCd

[Les07j] Mitch Leslie. Centering the centriole. Journal of Cell Biology, 178(5):??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/5/717.2.

Leslie:2007:CTS

[Les07k] Mitch Leslie. A chromatin twist to silencing choice. Journal of Cell Biology, 179(2):??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/2/170.1.

Leslie:2007:CCL

[Les07l] Mitch Leslie. Chromosomes can’t let go. Journal of Cell Biology, 176(4):??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/4/374a.

Leslie:2007:CCC

[Les07m] Mitch Leslie. Closing the connexin connection. Journal of Cell Biology, 177(5):??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/5/745a.

Leslie:2007:CCb

[Les07n] Mitch Leslie. Coddling cancer. Journal of Cell Biology, 176(4): 374–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/4/374.

Leslie:2007:CCR

[Les07o] Mitch Leslie. Coilin to the centromere’s rescue. Journal of Cell Biology, 177(5):745–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/5/745.

Leslie:2007:CCa

[Les07p] Mitch Leslie. Cycling on without centrosomes. Journal of Cell Biology, 176(2):??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/2/129b.

Leslie:2007:DCS

[Les07q] Mitch Leslie. Death comes for the salivary gland. Journal of Cell Biology, 178(1):??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/1/2a.

Leslie:2007:AW

[Les07r] Mitch Leslie. Die another way. Journal of Cell Biology, 176(6):??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/6/731b.

Leslie:2007:ETG

[Les07s] Mitch Leslie. Embryonic telomeres go long. Journal of Cell Biology, 179(5):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/5/808.1.

Leslie:2007:EET

[Les07t] Mitch Leslie. Emi1 ensures timely mitosis. Journal of Cell Biology, 177(3):370–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/3/370.

Leslie:2007:FAD

[Les07u] Mitch Leslie. Fighting Alzheimer’s by defusing a death receptor. Journal of Cell Biology, 178(5):??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/178/5/716.1.

Leslie:2007:FMA

[Les07v] Mitch Leslie. Filopodia motor ahead. Journal of Cell Biology, 179(2):??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/2/171.2.

Leslie:2007:COE

[Les07w] Mitch Leslie. For cdks, one is enough. Journal of Cell Biology, 178(6):??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/6/892.2.

Leslie:2007:LLC

[Les07x] Mitch Leslie. For long, lush cilia, try Par3. Journal of Cell Biology, 179(6):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/6/1085.2.

Leslie:2007:FB

[Les07y] Mitch Leslie. Fusing without breaking. Journal of Cell Biology, 178(6):??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/6/893.1.

Leslie:2007:FTE

[Les07z] Mitch Leslie. Fusion in three easy steps. Journal of Cell Biology, 179(4):??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/4/569.1.

Leslie:2007:GFM

[Les07-27] Mitch Leslie. Go forth or multiply. Journal of Cell Biology, 178(1):2–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/1/2.

Leslie:2007:GTO

[Les07-28] Mitch Leslie. Going their own way in the NPC. Journal of Cell Biology, 178(7):??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/7/1095.3.

Leslie:2007:HSW

[Les07-29] Mitch Leslie. Healing skin from within. Journal of Cell Biology, 179(4):??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/4/568.2.

Leslie:2007:HCA

[Les07-30] Mitch Leslie. Helping cells achieve oneness. Journal of Cell Biology, 176(2):??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/2/128a.

Leslie:2007:HDG

[Les07-31] Mitch Leslie. Hey, DNA, get over here. Journal of Cell Biology, 179(6):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/6/1085.1.

Leslie:2007:HCR

[Les07-32] Mitch Leslie. How CENP–A reaches the centromere. Journal of Cell Biology, 176(6):??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/6/730a.

Leslie:2007:HKD

[Les07-33] Mitch Leslie. How kinases dodge trouble. Journal of Cell Biology, 176(3):??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/3/245a.

Leslie:2007:HCF

[Les07-34] Mitch Leslie. Hungry cells forget which way is up. Journal of Cell Biology, 177(3):??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/3/370a.

Leslie:2007:KSC

[Les07-35] Mitch Leslie. Keeping sisters close. Journal of Cell Biology, 176(4):??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/4/375a.

Leslie:2007:KTA

[Les07-36] Mitch Leslie. Kinetochores take alternative transportation. Journal of Cell Biology, 178(2):??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/178/2/180a.

Leslie:2007:LMV

[Les07-37] Mitch Leslie. A lean and mean VEGF for cancer cells. Journal of Cell Biology, 179(6):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/6/1084.1.

Leslie:2007:LT

[Les07-38] Mitch Leslie. Liberating TGFβ. Journal of Cell Biology, 176 (3):245–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/3/245.

Leslie:2007:LSM

[Les07-39] Mitch Leslie. The long and short of mitochondrial fusion. Journal of Cell Biology, 178(5):??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/178/5/716.2.

Leslie:2007:MRM

[Les07-40] Mitch Leslie. Making room for muscle. Journal of Cell Biology, 179(6):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/6/1085.3.

Leslie:2007:MCS

[Les07-41] Mitch Leslie. Merlin casts spell on proliferation protein. Journal of Cell Biology, 177(5):??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/5/744a.

Leslie:2007:MMa

[Les07-42] Mitch Leslie. Meting out mitochondria. Journal of Cell Biology, 176(2):129–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/2/129.

Leslie:2007:MR

[Les07-43] Mitch Leslie. Microglia to the rescue. Journal of Cell Biology, 179(6):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/6/1084.2.

Leslie:2007:MMb

[Les07-44] Mitch Leslie. Mitochondria on the MARCH. Journal of Cell Biology, 178(1):3–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/1/3.

Leslie:2007:MFN

[Les07-45] Mitch Leslie. Muscle first in neurodegenerative disease? Journal of Cell Biology, 176(6):??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/176/6/731a.

Leslie:2007:MPT

[Les07-46] Mitch Leslie. Myosins pull together to move cargo. Journal of Cell Biology, 178(7):??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/178/7/1094.2.

Leslie:2007:NB

[Les07-47] Mitch Leslie. Nervous breakdown. Journal of Cell Biology, 176 (3):244–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/3/244.

Leslie:2007:NMM

[Les07-48] Mitch Leslie. New mRNA modification? Journal of Cell Biology, 179(2):??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/2/171.3.

Leslie:2007:NWT

[Les07-49] Mitch Leslie. No way to treat a tumor. Journal of Cell Biology, 178(1):??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/1/3a.

Leslie:2007:PHB

[Les07-50] Mitch Leslie. p53 hits the brakes. Journal of Cell Biology, 178(1):??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/1/3b.

Leslie:2007:PIP

[Les07-51] Mitch Leslie. Pol II piece makes and breaks mRNA. Journal of Cell Biology, 178(7):??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/178/7/1095.1.

Leslie:2007:PPT

[Les07-52] Mitch Leslie. Polarizing power of touch. Journal of Cell Biology, 178(2):180–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/2/180.

Leslie:2007:PED

[Les07-53] Mitch Leslie. The price of excess DNA. Journal of Cell Biology, 178(6):??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/6/893.3.

Leslie:2007:PDP

[Les07-54] Mitch Leslie. Putting Dia1 in its place. Journal of Cell Biology, 178(2):??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/2/181a.

Leslie:2007:RDL

[Les07-55] Mitch Leslie. A receptor with divided loyalties. Journal of Cell Biology, 179(2):??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/2/170.2.

Leslie:2007:RDD

[Les07-56] Mitch Leslie. RECQ1’s double duty. Journal of Cell Biology, 176(4):375–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/4/375.

Leslie:2007:RWL

[Les07-57] Mitch Leslie. RhoG welcomes leukocytes. Journal of Cell Biology, 178(7):??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/7/1095.2.

Leslie:2007:RPCb

[Les07-58] Mitch Leslie. RNA polymerase caught loitering. Journal of Cell Biology, 179(5):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/5/809.2.

Leslie:2007:RPCa

[Les07-59] Mitch Leslie. RNA polymerase on the clock. Journal of Cell Biology, 179(2):??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/2/171.1.

Leslie:2007:SSS

[Les07-60] Mitch Leslie. Self-starting slime mold. Journal of Cell Biology, 178(2):??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/2/181b.

Leslie:2007:SUD

[Les07-61] Mitch Leslie. Separate and unequal Deltas. Journal of Cell Biology, 178(3):??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/3/339a.

Leslie:2007:SRC

[Les07-62] Mitch Leslie. A simple reentry into the cell cycle. Journal of Cell Biology, 176(6):730–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/176/6/730.

Leslie:2007:SNJ

[Les07-63] Mitch Leslie. Sizing the nucleus just right. Journal of Cell Biology, 179(4):??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/4/568.1.

Leslie:2007:SC

[Les07-64] Mitch Leslie. A sloppy checkpoint. Journal of Cell Biology, 176 (6):731–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/6/731.

Leslie:2007:SCC

[Les07-65] Mitch Leslie. Small cancer cells don’t stray. Journal of Cell Biology, 178(3):339–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/3/339.

Leslie:2007:SDP

[Les07-66] Mitch Leslie. Splitting the difference on peroxisomes. Journal of Cell Biology, 178(3):??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/3/338a.

Leslie:2007:STW

[Les07-67] Mitch Leslie. Starting a tumor off wrong. Journal of Cell Biology, 178(7):??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/7/1094.1.

Leslie:2007:SCS

[Les07-68] Mitch Leslie. Stem cell side show. Journal of Cell Biology, 176(3):??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/3/245b.

Leslie:2007:TDL

[Les07-69] Mitch Leslie. Thrown for a D-loop. Journal of Cell Biology, 176(2):??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/2/129a.

Leslie:2007:TS

[Les07-70] Mitch Leslie. Time to split. Journal of Cell Biology, 179 (5):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/5/809.3.

Leslie:2007:BN

[Les07-71] Mitch Leslie. To build a node. Journal of Cell Biology, 177(5): ??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
177/5/745b.

Leslie:2007:MNP

[Les07-72] Mitch Leslie. To make a nuclear pore, just relax. Journal of Cell Biology, 178(5):??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/5/717.3.

Leslie:2007:UTR

[Les07-73] Mitch Leslie. Unique tasks for ribosomal paralogues. Journal of Cell Biology, 179(5):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/5/808.2.

Leslie:2007:SGS

[Les07-74] Mitch Leslie. The up side of a gas shortage. Journal of Cell Biology, 177(3):371–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/3/371.

Leslie:2007:UST

[Les07-75] Mitch Leslie. uPAR’s signaling two step. Journal of Cell Biology, 177(5):744–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/5/744.

Leslie:2007:WCJ

[Les07-76] Mitch Leslie. When cargos jump tracks. Journal of Cell Biology, 179(4):??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/4/569.2.

Leslie:2007:WVS

[Les07-77] Mitch Leslie. When VEGF strays. Journal of Cell Biology, 177(3):??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/3/371b.

Leslie:2007:WCT

[Les07-78] Mitch Leslie. Why are cells so tense? Journal of Cell Biology, 176(4):??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/4/375b.

Leslie:2007:WMS

[Les07-79] Mitch Leslie. Wnt makes stem cells act their age. Journal of Cell Biology, 178(6):??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/178/6/892.1.

Leslie:2008:AHR

[Les08a] Mitch Leslie. Actin hitches a ride to the cleavage furrow. Journal of Cell Biology, 181(5):??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/181/5/713.2.

Leslie:2008:APC

[Les08b] Mitch Leslie. Alzheimer’s protein controls calcium’s ins and outs. Journal of Cell Biology, 181(7):??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/181/7/1042.2.

Leslie:2008:APK

[Les08c] Mitch Leslie. Arp2/ 3 phosphorylation kickstarts cells. Journal of Cell Biology, 182(4):??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/182/4/617.2.

Leslie:2008:ALT

[Les08d] Mitch Leslie. Axons lose their identity. Journal of Cell Biology, 183(4):??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/4/568.3.

Leslie:2008:BF

[Les08e] Mitch Leslie. Backtracking on FAK. Journal of Cell Biology, 180(1):??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/1/2.1.

Leslie:2008:BBD

[Les08f] Mitch Leslie. Bad breaks don’t doom meiotic cells. Journal of Cell Biology, 182(2):??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/2/217.2.

Leslie:2008:CLW

[Les08g] Mitch Leslie. Cardio(lipin) workout beefs up mitochondrial imports. Journal of Cell Biology, 183(7):??, December 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/183/7/1178.1.

Leslie:2008:CCM

[Les08h] Mitch Leslie. Career change for a mitotic protein. Journal of Cell Biology, 181(1):??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/1/2.2.

Leslie:2008:CTS

[Les08i] Mitch Leslie. Cdc42 turns the spindle. Journal of Cell Biology, 183(4):??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/4/568.1.

Leslie:2008:CPR

[Les08j] Mitch Leslie. Cells’ puncture repair kit. Journal of Cell Biology, 180(5):??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/5/847.2.

Leslie:2008:CSE

[Les08k] Mitch Leslie. Cellular self-eating promotes pancreatitis. Journal of Cell Biology, 181(7):??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/7/1042.1.

Leslie:2008:CGF

[Les08l] Mitch Leslie. CENP-e goes fishing for microtubules. Journal of Cell Biology, 181(3):??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/3/399.3.

Leslie:2008:CC

[Les08m] Mitch Leslie. Cutting the cord. Journal of Cell Biology, 181(4): ??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
181/4/569.1.

Leslie:2008:DCS

[Les08n] Mitch Leslie. Daughter cells share duties. Journal of Cell Biology, 181(7):??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/7/1043.3.

Leslie:2008:DYC

[Les08o] Mitch Leslie. Dieting yeast clean house. Journal of Cell Biology, 181(5):??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/5/713.1.

Leslie:2008:DH

[Les08p] Mitch Leslie. DNA healer. Journal of Cell Biology, 183 (4):??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/4/569.2.

Leslie:2008:DJC

[Les08q] Mitch Leslie. DNA, just chill. Journal of Cell Biology, 183 (7):??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/7/1178.2.

Leslie:2008:DBG

[Les08r] Mitch Leslie. Driven to the brink by a G protein. Journal of Cell Biology, 182(2):??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/2/216.2.

Leslie:2008:FP

[Les08s] Mitch Leslie. Feeling the pull. Journal of Cell Biology, 182 (4):??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/4/616.2.

Leslie:2008:GFS

[Les08t] Mitch Leslie. Genetic feng shui for cancer cells? Journal of Cell Biology, 180(1):??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/1/3.2.

Leslie:2008:GTC

[Les08u] Mitch Leslie. A get-together for caspases. Journal of Cell Biology, 183(4):??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/4/568.2.

Leslie:2008:HT

[Les08v] Mitch Leslie. Hauling tail. Journal of Cell Biology, 181(1):??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
181/1/3.2.

Leslie:2008:HAG

[Les08w] Mitch Leslie. Helping an axon go long. Journal of Cell Biology, 180(3):??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/3/445.3.

Leslie:2008:HCL

[Les08x] Mitch Leslie. Helping cells lose their inhibitions. Journal of Cell Biology, 182(2):??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/2/217.1.

Leslie:2008:HCP

[Les08y] Mitch Leslie. How a cell puts itself on the menu. Journal of Cell Biology, 182(4):??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/4/617.1.

Leslie:2008:HCM

[Les08z] Mitch Leslie. How cells make local calls. Journal of Cell Biology, 181(7):??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/7/1043.1.

Leslie:2008:HEM

[Les08-27] Mitch Leslie. How to escape from a macrophage. Journal of Cell Biology, 180(5):??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/5/846.1.

Leslie:2008:HDP

[Les08-28] Mitch Leslie. Huntington’s disease protein extends its reach. Journal of Cell Biology, 181(5):??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/181/5/712.1.

Leslie:2008:LRA

[Les08-29] Mitch Leslie. Laminin reaches across the synapse. Journal of Cell Biology, 182(6):??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/182/6/1031.1.

Leslie:2008:LLD

[Les08-30] Mitch Leslie. Leptin’s long-distance call to the pancreas. Journal of Cell Biology, 183(7):??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/183/7/1178.3.

Leslie:2008:MWC

[Les08-31] Mitch Leslie. Making the wrong connections in epilepsy. Journal of Cell Biology, 180(5):??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/5/846.2.

Leslie:2008:MFT

[Les08-32] Mitch Leslie. Matrix fragments trigger fatal excitement. Journal of Cell Biology, 183(7):??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/183/7/1179.1.

Leslie:2008:MDS

[Les08-33] Mitch Leslie. Mitochondrial DNA stays home. Journal of Cell Biology, 181(7):??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/7/1043.2.

Leslie:2008:NLS

[Les08-34] Mitch Leslie. Neurons on life support. Journal of Cell Biology, 181(1):??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/1/3.1.

Leslie:2008:NSS

[Les08-35] Mitch Leslie. New spin on spindle size. Journal of Cell Biology, 182(2):??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/2/216.3.

Leslie:2008:OMI

[Les08-36] Mitch Leslie. Origins of mitotic inequality. Journal of Cell Biology, 180(5):??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/5/847.1.

Leslie:2008:PST

[Les08-37] Mitch Leslie. Prions show their good side. Journal of Cell Biology, 181(3):??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/3/398.1.

Leslie:2008:PGH

[Les08-38] Mitch Leslie. Proteins gang up on HIV. Journal of Cell Biology, 180(3):??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/3/444.2.

Leslie:2008:PC

[Les08-39] Mitch Leslie. Pumping up caveolae. Journal of Cell Biology, 182(6):??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/6/1031.2.

Leslie:2008:PCT

[Les08-40] Mitch Leslie. Putting channels in their place. Journal of Cell Biology, 180(1):??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/1/3.1.

Leslie:2008:RBR

[Les08-41] Mitch Leslie. Rab(ble) rouser. Journal of Cell Biology, 182 (6):??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/6/1030.2.

Leslie:2008:RTT

[Les08-42] Mitch Leslie. Rac1 takes a trip. Journal of Cell Biology, 181(3): ??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
181/3/399.1.

Leslie:2008:RR

[Les08-43] Mitch Leslie. Ribosomes rebuffed. Journal of Cell Biology, 181(4):??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/4/569.2.

Leslie:2008:RPD

[Les08-44] Mitch Leslie. RNA polymerase doesn’t make deliveries. Journal of Cell Biology, 181(4):??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/4/569.3.

Leslie:2008:RGA

[Les08-45] Mitch Leslie. Rules of gene attraction. Journal of Cell Biology, 182(6):??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/6/1030.3.

Leslie:2008:SPG

[Les08-46] Mitch Leslie. Short proteins got no reason to leave. Journal of Cell Biology, 181(1):??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/1/3.3.

Leslie:2008:SSC

[Les08-47] Mitch Leslie. SIRT2 stops cells. Journal of Cell Biology, 180(5):??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/5/847.3.

Leslie:2008:SEW

[Les08-48] Mitch Leslie. Slow down, enzymes at work. Journal of Cell Biology, 183(7):??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/7/1179.2.

Leslie:2008:SSS

[Les08-49] Mitch Leslie. Slow synapses in schizophrenia? Journal of Cell Biology, 181(5):??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/5/712.2.

Leslie:2008:SCSa

[Les08-50] Mitch Leslie. Stem cell slowdown. Journal of Cell Biology, 181(1):??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/1/2.1.

Leslie:2008:SCSb

[Les08-51] Mitch Leslie. Stem cells stand up for themselves. Journal of Cell Biology, 182(4):??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/4/616.1.

Leslie:2008:SM

[Les08-52] Mitch Leslie. Stuck on the membrane. Journal of Cell Biology, 180(3):??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/3/444.1.

Leslie:2008:SLD

[Les08-53] Mitch Leslie. Supersized lipid droplets. Journal of Cell Biology, 180(3):??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/3/445.1.

Leslie:2008:SCC

[Les08-54] Mitch Leslie. Synapse crowd control. Journal of Cell Biology, 181(5):??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/5/713.3.

Leslie:2008:TTT

[Les08-55] Mitch Leslie. Tag team at the telomeres. Journal of Cell Biology, 181(3):??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/3/399.2.

Leslie:2008:TJL

[Les08-56] Mitch Leslie. Tight junctions loosen up. Journal of Cell Biology, 181(4):??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/4/568.2.

Leslie:2008:TB

[Les08-57] Mitch Leslie. Timing the breakdown. Journal of Cell Biology, 182(6):??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/6/1030.1.

Leslie:2008:TC

[Les08-58] Mitch Leslie. Traffic control in the ER. Journal of Cell Biology, 180(1):??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/1/2.2.

Leslie:2008:TMA

[Les08-59] Mitch Leslie. Tuning up mitochondria for angiogenesis. Journal of Cell Biology, 180(1):??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/1/3.3.

Leslie:2008:TBC

[Les08-60] Mitch Leslie. Turning back the clock for Schwann cells. Journal of Cell Biology, 181(4):??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/4/568.1.

Leslie:2008:USCa

[Les08-61] Mitch Leslie. Uniting to save cancer cells. Journal of Cell Biology, 183(4):??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/4/569.1.

Leslie:2008:USCb

[Les08-62] Mitch Leslie. Uniting to save cancer cells. Journal of Cell Biology, 183(7):1315–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/183/7/1315.

Leslie:2008:UTF

[Les08-63] Mitch Leslie. Untying tube formation. Journal of Cell Biology, 182(2):??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/2/216.1.

Leslie:2008:WCM

[Les08-64] Mitch Leslie. When cells make a meal of themselves. Journal of Cell Biology, 181(3):??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/3/398.2.

Leslie:2008:WCL

[Les08-65] Mitch Leslie. When cilia lose the beat. Journal of Cell Biology, 180(3):??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/3/445.2.

Leslie:2008:YPC

[Les08-66] Mitch Leslie. Young proteins cut down in their prime. Journal of Cell Biology, 182(4):??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/182/4/616.3.

Leslie:2009:AHC

[Les09a] Mitch Leslie. Aki1 helps centrioles stay tight. Journal of Cell Biology, 187(5):??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/5/584.1.

Leslie:2009:ACL

[Les09b] Mitch Leslie. Apoptotic cells let down their guard. Journal of Cell Biology, 184(2):??, January 2009. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/184/2/187.1.

Leslie:2009:AMB

[Les09c] Mitch Leslie. Autophagosomes meet before they eat. Journal of Cell Biology, 185(2):??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/2/178.2.

Leslie:2009:BCC

[Les09d] Mitch Leslie. Birth control for centrioles. Journal of Cell Biology, 184(2):??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/2/186.2.

Leslie:2009:BRCb

[Les09e] Mitch Leslie. Both routes to cell death necessary. Journal of Cell Biology, 186(3):??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/186/3/300.1.

Leslie:2009:BRCa

[Les09f] Mitch Leslie. Brain-repairing cells follow trail of GABA. Journal of Cell Biology, 186(1):??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/186/1/2.2.

Leslie:2009:BMA

[Les09g] Mitch Leslie. Building memories with actin. Journal of Cell Biology, 186(1):??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/186/1/2.1.

Leslie:2009:CCR

[Les09h] Mitch Leslie. Calcium channel roundup. Journal of Cell Biology, 186(1):??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/186/1/2.3.

Leslie:2009:CRF

[Les09i] Mitch Leslie. Calcium’s route to freedom. Journal of Cell Biology, 186(2):??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/186/2/164.2.

Leslie:2009:CNR

[Les09j] Mitch Leslie. CD44’s nuclear road trip. Journal of Cell Biology, 185(6):??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/6/930.2.

Leslie:2009:CDL

[Les09k] Mitch Leslie. Chlamydomonas does the lopsided wave. Journal of Cell Biology, 186(3):??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/186/3/300.3.

Leslie:2009:CSI

[Les09l] Mitch Leslie. Cytomegalovirus sticks it to MHC. Journal of Cell Biology, 186(5):??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/186/5/632.1.

Leslie:2009:DMS

[Les09m] Mitch Leslie. The downside of microtubule stability. Journal of Cell Biology, 185(6):??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/6/930.1.

Leslie:2009:DMN

[Les09n] Mitch Leslie. Dystrophin makes a new connection. Journal of Cell Biology, 186(3):??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/186/3/300.2.

Leslie:2009:FAC

[Les09o] Mitch Leslie. Focal adhesions come unstuck. Journal of Cell Biology, 187(5):585–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/5/585.

Leslie:2009:CEM

[Les09p] Mitch Leslie. For chromosomes, equality means segregation. Journal of Cell Biology, 186(1):3–??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/186/1/3.

Leslie:2009:GMT

[Les09q] Mitch Leslie. From gene manager to traffic cop. Journal of Cell Biology, 186(3):301–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/186/3/301.

Leslie:2009:HMG

[Les09r] Mitch Leslie. How a membrane gets the bends. Journal of Cell Biology, 185(6):??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/6/930.3.

Leslie:2009:KCM

[Les09s] Mitch Leslie. Killing the β-catenin messenger. Journal of Cell Biology, 186(2):??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/186/2/164.3.

Leslie:2009:LRG

[Les09t] Mitch Leslie. Lipid rafts get connected. Journal of Cell Biology, 184(5):??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/184/5/617.1.

Leslie:2009:MB

[Les09u] Mitch Leslie. Managing the breakup. Journal of Cell Biology, 186(5):633–??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/186/5/633.

Leslie:2009:MPS

[Les09v] Mitch Leslie. Membrane proteins stick around. Journal of Cell Biology, 184(2):??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/2/186.1.

Leslie:2009:MGS

[Les09w] Mitch Leslie. Mitochondrial growth spurt. Journal of Cell Biology, 184(5):??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/184/5/616.3.

Leslie:2009:MAT

[Les09x] Mitch Leslie. Muscle atrophy through thick but not thin. Journal of Cell Biology, 185(6):931–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/185/6/931.

Leslie:2009:MPT

[Les09y] Mitch Leslie. Myosins pull together — separately. Journal of Cell Biology, 187(1):3–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/187/1/3.

Leslie:2009:NCA

[Les09z] Mitch Leslie. No copying allowed. Journal of Cell Biology, 187 (5):??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/187/5/584.2.

Leslie:2009:NPS

[Les09-27] Mitch Leslie. Nuclear pore synthesis on a roll. Journal of Cell Biology, 184(5):??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/5/616.2.

Leslie:2009:PGM

[Les09-28] Mitch Leslie. A plasmid goes motoring. Journal of Cell Biology, 185(2):179–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/2/179.

Leslie:2009:PNE

[Les09-29] Mitch Leslie. Preventing the nuclear envelope from getting on a roll. Journal of Cell Biology, 186(2):??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/186/2/164.1.

Leslie:2009:PSA

[Les09-30] Mitch Leslie. Putting the squeeze on apoptotic cells. Journal of Cell Biology, 186(5):??, September 2009. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/186/5/632.2.

Leslie:2009:RPA

[Les09-31] Mitch Leslie. Rules of phagocytic attraction. Journal of Cell Biology, 187(5):??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/5/584.3.

Leslie:2009:STT

[Les09-32] Mitch Leslie. SIRT1 takes down tumors. Journal of Cell Biology, 185(2):??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/2/178.1.

Leslie:2009:SHC

[Les09-33] Mitch Leslie. Spindle’s hidden collaborator exposed. Journal of Cell Biology, 184(5):??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/5/616.1.

Leslie:2009:SSS

[Les09-34] Mitch Leslie. Stopping sister synapses. Journal of Cell Biology, 186(5):??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/186/5/632.3.

Leslie:2009:TBW

[Les09-35] Mitch Leslie. Turn back, wayward axon. Journal of Cell Biology, 184(5):??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/184/5/617.2.

Leslie:2009:UCN

[Les09-36] Mitch Leslie. Unfolding chromatin nets. Journal of Cell Biology, 184(2):??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/184/2/187.2.

Leslie:2009:WCC

[Les09-37] Mitch Leslie. When cancer cells can’t let go. Journal of Cell Biology, 185(2):??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/2/178.3.

Letai:2009:PSB

[Let09] Anthony Letai. Puma strikes Bax. Journal of Cell Biology, 185 (2):189–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/2/189.

Levine:2007:LTP

[Lev07] Tim P. Levine. A lipid transfer protein that transfers lipid. Journal of Cell Biology, 179(1):11–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/1/11.

Lechler:2007:DUR

[LF07] Terry Lechler and Elaine Fuchs. Desmoplakin: an unexpected regulator of microtubule organization in the epidermis. Journal of Cell Biology, 176(2):147–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/176/2/147.

Lewis:2007:NEP

[LFH07] Alaron Lewis, Rachael Felberbaum, and Mark Hochstrasser. A nuclear envelope protein linking nuclear pore basket assembly, SUMO protease regulation, and mRNA surveillance. Journal of Cell Biology, 178(5):813–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/178/5/813.

Luo:2005:PRS

[LFL^{+}05] Ji Luo, Seth J. Field, Jennifer Y. Lee, Jeffrey A. Engelman, and Lewis C. Cantley. The p85 regulatory subunit of phosphoinositide 3-kinase down-regulates IRS-1 signaling via the formation of a sequestration complex. Journal of Cell Biology, 170(3):455–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/3/455.

Lince-Faria:2009:SCM

[LFMO^{+}09] Mariana Lince-Faria, Stefano Maffini, Bernard Orr, Yun Ding, Cláudia Florindo, Claudio E. Sunkel, Álvaro Tavares, Jørgen Johansen, Kristen M. Johansen, and Helder Maiato. Spatiotemporal control of mitosis by the conserved spindle matrix protein Megator. Journal of Cell Biology, 184(5):647–??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/184/5/647.

Luedeke:2005:SDC

[LFS^{+}05] Cosima Luedeke, Stéphanie Buvelot Frei, Ivo Sbalzarini, Heinz Schwarz, Anne Spang, and Yves Barral. Septin-dependent compartmentalization of the endoplasmic reticulum during yeast polarized growth. Journal of Cell Biology, 169(6):897–??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/169/6/
897.

Li:2009:LSAa

[LFW^{+}09a] Min Li, Xiao Fang, Zhubo Wei, J. Philippe York, and Pumin Zhang. Loss of spindle assembly checkpoint-mediated inhibition of Cdc20 promotes tumorigenesis in mice. Journal of Cell Biology, 185(6):983–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/6/983.

Li:2009:LSAb

[LFW^{+}09b] Min Li, Xiao Fang, Zhubo Wei, J. Philippe York, and Pumin Zhang. Loss of spindle assembly checkpoint-mediated inhibition of Cdc20 promotes tumorigenesis in mice. Journal of Cell Biology, 186(1):161–??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/186/1/161.

Lu:2007:PDC

[LG07] Junjie Lu and David M. Gilbert. Proliferation-dependent and cell cycle–regulated transcription of mouse pericentric heterochromatin. Journal of Cell Biology, 179(3):411–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/3/
411.

Lu:2008:PDC

[LG08] Junjie Lu and David M. Gilbert. Proliferation-dependent and cell cycle–regulated transcription of mouse pericentric heterochromatin. Journal of Cell Biology, 181(1):171–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/181/1/
171.

Lotan:2007:RSY

[LGBD^{+}07] Rona Lotan, Vicky Goler-Baron, Lea Duek, Gal Haimovich, and Mordechai Choder. The Rpb7p subunit of yeast RNA polymerase II plays roles in the two major cytoplasmic mRNA decay mechanisms. Journal of Cell Biology, 178(7):1133–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/178/7/
1133.

Lajoie:2009:LRS

[LGDN09] Patrick Lajoie, Jacky G. Goetz, James W. Dennis, and Ivan R. Nabi. Lattices, rafts, and scaffolds: domain regulation of receptor signaling at the plasma membrane. Journal of Cell Biology, 185(3):381–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/3/381.

Lynch:2005:CSU

[LGG^{+}05] Jeffrey Lynch, Lei Guo, Pascal Gelebart, Kaari Chilibeck, Jian Xu, Jeffery D. Molkentin, Luis B. Agellon, and Marek Michalak. Calreticulin signals upstream of calcineurin and MEF2C in a critical Ca^{2+}-dependent signaling cascade. Journal of Cell Biology, 170(1):37–??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/1/37.

Larson:2005:TRI

[LGH^{+}05] Daniel R. Larson, Julie A. Gosse, David A. Holowka, Barbara A. Baird, and Watt W. Webb. Temporally resolved interactions between antigen-stimulated IgE receptors and Lyn kinase on living cells. Journal of Cell Biology, 171(3):527– ??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/3/527.

Lorenz:2007:ILK

[LGT^{+}07] Katrin Lorenz, Carsten Grashoff, Robert Torka, Takao Sakai, Lutz Langbein, Wilhelm Bloch, Monique Aumailley, and Reinhard Fässler. Integrin-linked kinase is required for epidermal and hair follicle morphogenesis. Journal of Cell Biology, 177 (3):501–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/3/501.

Lien:2008:CBD

[LGV08] Wen-Hui Lien, Vladimir I. Gelfand, and Valeri Vasioukhin. α– e-catenin binds to dynamitin and regulates dynactin-mediated intracellular traffic. Journal of Cell Biology, 183(6):989–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
183/6/989.

Lee:2005:POO

[LGZ^{+}05] John Lee, Michel Gravel, Rulin Zhang, Pierre Thibault, and Peter E. Braun. Process outgrowth in oligodendrocytes is mediated by CNP, a novel microtubule assembly myelin protein. Journal of Cell Biology, 170(4):661–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/170/4/661.

Lalli:2005:RGR

[LH05] Giovanna Lalli and Alan Hall. Ral GTPases regulate neurite branching through GAP-43 and the exocyst complex. Journal of Cell Biology, 171(5):857–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/171/5/857.

Lewkowicz:2008:MBP

[LHL^{+}08] Elodie Lewkowicz, Floriane Herit, Christophe Le Clainche, Pierre Bourdoncle, Franck Perez, and Florence Niedergang. The microtubule-binding protein CLIP-170 coordinates mDia1 and actin reorganization during CR3-mediated phagocytosis. Journal of Cell Biology, 183(7):1287–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/183/7/1287.

Liu:2009:NSC

[LHM^{+}09] Xiaolin Liu, Sizhou Huang, Jun Ma, Chun Li, Yaoguang Zhang, and Lingfei Luo. NF-κB and Snail1a coordinate the cell cycle with gastrulation. Journal of Cell Biology, 184(6):805–??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
184/6/805.

Lin:2005:ERS

[LHRP05] Wensheng Lin, Heather P. Harding, David Ron, and Brian Popko. Endoplasmic reticulum stress modulates the response of myelinating oligodendrocytes to the immune cytokine interferon-γ. Journal of Cell Biology, 169(4):603–??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/169/4/603.

Lammerding:2005:ANS

[LHS^{+}05] Jan Lammerding, Janet Hsiao, P. Christian Schulze, Serguei Kozlov, Colin L. Stewart, and Richard T. Lee. Abnormal nuclear shape and impaired mechanotransduction in emerindeficient cells. Journal of Cell Biology, 170(5):781–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/170/5/
781.

Li:2008:CPR

[LHZ^{+}08] Xueni Li, Mei Huang, Huiling Zheng, Yinyin Wang, Fangli Ren, Yu Shang, Yonggong Zhai, David M. Irwin, Yuguang Shi, Di Chen, and Zhijie Chang. CHIP promotes Runx2 degradation and negatively regulates osteoblast differentiation. Journal of Cell Biology, 181(6):959–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/181/6/959.

Lin:2008:CBS

[Lin08] Haifan Lin. Cell biology of stem cells: an enigma of asymmetry and self-renewal. Journal of Cell Biology, 180(2):257–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
180/2/257.

Lauwers:2009:KLU

[LJA09] Elsa Lauwers, Christophe Jacob, and Bruno André. K63-linked ubiquitin chains as a specific signal for protein sorting into the multivesicular body pathway. Journal of Cell Biology, 185 (3):493–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/3/493.

Lester:2007:UIE

[LJM^{+}07] Robin D. Lester, Minji Jo, Valérie Montel, Shinako Takimoto, and Steven L. Gonias. uPAR induces epithelial–mesenchymal transition in hypoxic breast cancer cells. Journal of Cell Biology, 178(3):425–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/3/425.

Lee:2005:PVM

[LJN^{+}05] Sunyoung Lee, Shahla M. Jilani, Ganka V. Nikolova, Darren Carpizo, and M. Luisa Iruela-Arispe. Processing of VEGF–A by matrix metalloproteinases regulates bioavailability and vascular patterning in tumors. Journal of Cell Biology, 169(4):681–??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
169/4/681.

LaLonde:2005:RPR

[LJR^{+}05] Mary M. LaLonde, Hilde Janssens, Erica Rosenbaum, Seok-Yong Choi, J. Peter Gergen, Nansi J. Colley, William S. Stark, and Michael A. Frohman. Regulation of phototransduction responsiveness and retinal degeneration by a phospholipase D– generated signaling lipid. Journal of Cell Biology, 169(3):471– ??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
169/3/471.

Lechtreck:2009:CRB

[LJS^{+}09a] Karl-Ferdinand Lechtreck, Eric C. Johnson, Tsuyoshi Sakai, Deborah Cochran, Bryan A. Ballif, John Rush, Gregory J. Pazour, Mitsuo Ikebe, and George B. Witman. The Chlamydomonas reinhardtii BBSome is an IFT cargo required for export of specific signaling proteins from flagella. Journal of Cell Biology, 187(7):1117–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/187/7/1117.

Lleres:2009:QAC

[LJS^{+}09b] David Llères, John James, Sam Swift, David G. Norman, and Angus I. Lamond. Quantitative analysis of chromatin compaction in living cells using FLIM–FRET. Journal of Cell Biology, 187(4):481–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/4/481.

Lu:2007:RSG

[LJY^{+}07] Zhonghua Lu, Hyun-Soo Je, Paul Young, Jimmy Gross, Bai Lu, and Guoping Feng. Regulation of synaptic growth and maturation by a synapse-associated E3 ubiquitin ligase at the neuromuscular junction. Journal of Cell Biology, 177(6):1077– ??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
177/6/1077.

Liao:2005:DIC

[LK05] Maofu Liao and Margaret Kielian. Domain III from class II fusion proteins functions as a dominant-negative inhibitor of virus membrane fusion. Journal of Cell Biology, 171(1):111–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
171/1/111.

LeBrasseur:2006:GGC

[LK06] Nicole LeBrasseur and Alla Katsnelson. Golden Gate cells: The American Society for Cell Biology San Francisco, CA December 10–14, 2005. Journal of Cell Biology, 172(4):486–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/172/4/
486.

Lee:2005:OMD

[LKC05] Wei-Lih Lee, Michelle A. Kaiser, and John A. Cooper. The offloading model for dynein function. Journal of Cell Biology, 168(2):201–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/2/201.

Lindqvist:2005:CCC

[LKL^{+}05] Arne Lindqvist, Helena Källström, Andreas Lundgren, Emad Barsoum, and Christina Karlsson Rosenthal. Cdc25B cooperates with Cdc25A to induce mitosis but has a unique role in activating cyclin B1–Cdk1 at the centrosome. Journal of Cell Biology, 171(1):35–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/1/35.

Liu:2007:AMP

[LLC^{+}07] Enbo Liu, Alan Yueh-Luen Lee, Takuya Chiba, Erin Olson, Peiqing Sun, and Xiaohua Wu. The ATR-mediated S phase checkpoint prevents rereplication in mammalian cells when licensing control is disrupted. Journal of Cell Biology, 179(4): 643–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/4/643.

Luissint:2008:JML

[LLC^{+}08] Anny-Claude Luissint, Pierre G. Lutz, David A. Calderwood, Pierre-Olivier Couraud, and Sandrine Bourdoulous. JAM-L-mediated leukocyte adhesion to endothelial cells is regulated in cis by α4β1 integrin activation. Journal of Cell Biology, 183(6): 1159–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/6/1159.

Landsverk:2007:UCM

[LLH^{+}07] Megan L. Landsverk, Shumin Li, Alex H. Hutagalung, Ayaz Najafov, Thorsten Hoppe, José M. Barral, and Henry F. Epstein. The UNC-45 chaperone mediates sarcomere assembly through myosin degradation in Caenorhabditis elegans. Journal of Cell Biology, 177(2):205–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/2/205.

Li:2009:HHH

[LLH^{+}09] Yan Li, Seung Lim, David Hoffman, Pontus Aspenstrom, Howard J. Federoff, and David A. Rempe. HUMMR, a hypoxiaand HIF-1α-inducible protein, alters mitochondrial distribution and transport. Journal of Cell Biology, 185(6):1065–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/185/6/
1065.

Lin:2007:SIF

[LLHH07] Yi-Ling Lin, Ya-Ting Lei, Chen-Jei Hong, and Yi-Ping Hsueh. Syndecan-2 induces filopodia and dendritic spine formation via the neurofibromin-PKA-Ena/VASP pathway. Journal of Cell Biology, 177(5):829–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/5/829.

Leisner:2005:ERC

[LLJ^{+}05] Tina M. Leisner, Mingjuan Liu, Zahara M. Jaffer, Jonathan Chernoff, and Leslie V. Parise. Essential role of CIB1 in regulating PAK1 activation and cell migration. Journal of Cell Biology, 170(3):465–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/3/465.

Lee:2006:PAA

[LLL^{+}06] Jin-A Lee, Sue-Hyun Lee, Changhoon Lee, Deok-Jin Chang, Yong Lee, Hyoung Kim, Ye-Hwang Cheang, Hyoung-Gon Ko, Yong-Seok Lee, Heejung Jun, Dusan Bartsch, Eric R. Kandel, and Bong-Kiun Kaang. PKA-activated ApAF-ApC/EBP heterodimer is a key downstream effector of ApCREB and is necessary and sufficient for the consolidation of long-term facilitation. Journal of Cell Biology, 174(6):827–??, September 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/174/6/827.

Li:2005:LSB

[LLM^{+}05] Shaohua Li, Patricia Liquari, Karen K. McKee, David Harrison, Raj Patel, Sean Lee, and Peter D. Yurchenco. Laminin– sulfatide binding initiates basement membrane assembly and enables receptor signaling in Schwann cells and fibroblasts. Journal of Cell Biology, 169(1):179–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/169/1/179.

Lidke:2005:RS

[LLR^{+}05] Diane S. Lidke, Keith A. Lidke, Bernd Rieger, Thomas M. Jovin, and Donna J. Arndt-Jovin. Reaching out for signals. Journal of Cell Biology, 170(4):619–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/170/4/619.

Loo:2009:HPS

[LLS^{+}09] Lit-Hsin Loo, Hai-Jui Lin, Dinesh K. Singh, Kathleen M. Lyons, Steven J. Altschuler, and Lani F. Wu. Heterogeneity in the physiological states and pharmacological responses of differentiating 3T3-L1 preadipocytes. Journal of Cell Biology, 187(3): 375–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/187/3/375.

Lim:2008:PFC

[LLT^{+}08] Yangmi Lim, Ssang-Taek Lim, Alok Tomar, Margaret Gardel, Joie A. Bernard-Trifilo, Xiao Lei Chen, Sean A. Uryu, Rafaela Canete-Soler, Jinbin Zhai, Hong Lin, William W. Schlaepfer, Perihan Nalbant, Gary Bokoch, Dusko Ilic, Clare Waterman-Storer, and David D. Schlaepfer. PyK2 and FAK connections to p190Rho guanine nucleotide exchange factor regulate RhoA activity, focal adhesion formation, and cell motility. Journal of Cell Biology, 180(1):187–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/1/187.

Lo:2009:RSA

[LLW^{+}09] Kai-Yin Lo, Zhihua Li, Feng Wang, Edward M. Marcotte, and Arlen W. Johnson. Ribosome stalk assembly requires the dualspecificity phosphatase Yvh1 for the exchange of Mrt4 with P0. Journal of Cell Biology, 186(6):849–??, September 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/186/6/849.

Leach:2005:USM

[LM05] Craig A. Leach and W. Matthew Michael. Ubiquitin/ SUMO modification of PCNA promotes replication fork progression in Xenopus laevis egg extracts. Journal of Cell Biology, 171(6): 947–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/6/947.

Liu:2006:DMC

[LMB^{+}06] Ji-Long Liu, Christine Murphy, Michael Buszczak, Sarah Clatterbuck, Robyn Goodman, and Joseph G. Gall. The Drosophila melanogaster Cajal body. Journal of Cell Biology, 172(6): 875–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/6/875.

Li:2009:REM

[LMC^{+}09] Gang Li, Marco Mongillo, King-Tung Chin, Heather Harding, David Ron, Andrew R. Marks, and Ira Tabas. Role of ERO1-α-mediated stimulation of inositol 1,4,5-triphosphate receptor activity in endoplasmic reticulum stress–induced apoptosis. Journal of Cell Biology, 186(6):783–??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/186/6/783.

Li:2008:CCR

[LMHT08] Feng-Qian Li, Adaobi Mofunanya, Kimberley Harris, and Ken-Ichi Takemaru. Chibby cooperates with 14-3-3 to regulate βcatenin subcellular distribution and signaling activity. Journal of Cell Biology, 181(7):1141–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/181/7/1141.

Lapetina:2009:AIC

[LMM^{+}09] Stefanie Lapetina, Christopher C. Mader, Kazuya Machida, Bruce J. Mayer, and Anthony J. Koleske. Arg interacts with cortactin to promote adhesion-dependent cell edge protrusion. Journal of Cell Biology, 185(3):503–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/185/3/503.

Laurent-Matha:2005:CIH

[LMMHP^{+}05] Valérie Laurent-Matha, Sharon Maruani-Herrmann, Christine Prébois, Mélanie Beaujouin, Murielle Glondu, Agnès Noël, Marie Luz Alvarez-Gonzalez, Sylvia Blacher, Peter Coopman, Stephen Baghdiguian, Christine Gilles, Jadranka Loncarek, Gilles Freiss, Françoise Vignon, and Emmanuelle Liaudet-Coopman. Catalytically inactive human cathepsin D triggers fibroblast invasive growth. Journal of Cell Biology, 168(3): 489–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/3/489.

Lin:2005:CS

[LMS05] Joseph Lin, Mark J. Miller, and Andrey S. Shaw. The c–SMAC. Journal of Cell Biology, 170(2):177–??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/170/2/177.

Lamkanfi:2009:GIC

[LMV^{+}09] Mohamed Lamkanfi, James L. Mueller, Alberto C. Vitari, Shahram Misaghi, Anna Fedorova, Kurt Deshayes, Wyne P.

Lee, Hal M. Hoffman, and Vishva M. Dixit. Glyburide inhibits the Cryopyrin/Nalp3 inflammasome. Journal of Cell Biology, 187(1):61–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/1/61.

Lizunov:2005:ISH

[LMZ^{+}05] Vladimir A. Lizunov, Hideko Matsumoto, Joshua Zimmerberg, Samuel W. Cushman, and Vadim A. Frolov. Insulin stimulates the halting, tethering, and fusion of mobile GLUT4 vesicles in rat adipose cells. Journal of Cell Biology, 169(3):481–??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/169/3/
481.

Landeira:2007:NRV

[LN07] David Landeira and Miguel Navarro. Nuclear repositioning of the VSG promoter during developmental silencing in Trypanosoma brucei. Journal of Cell Biology, 176(2):133–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
176/2/133.

Ladi:2005:DDL

[LNG^{+}05] Ena Ladi, James T. Nichols, Weihong Ge, Alison Miyamoto, Christine Yao, Liang-Tung Yang, Jim Boulter, Yi E. Sun, Chris Kintner, and Gerry Weinmaster. The divergent DSL ligand Dll3 does not activate Notch signaling but cell autonomously attenuates signaling induced by other DSL ligands. Journal of Cell Biology, 170(6):983–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/170/6/983.

Liu:2008:MTO

[LNL^{+}08] Fangfang Liu, Seng Kah Ng, Yanfen Lu, Wilson Low, Julian Lai, and Gregory Jedd. Making two organelles from one: Woronin body biogenesis by peroxisomal protein sorting. Journal of Cell Biology, 180(2):325–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/180/2/325.

Liu:2006:CES

[LNPC06] Wendy F. Liu, Celeste M. Nelson, Dana M. Pirone, and Christopher S. Chen. E-cadherin engagement stimulates proliferation via Rac1. Journal of Cell Biology, 173(3):431–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/173/3/431.

Lampugnani:2006:VEC

[LOG^{+}06] Maria Grazia Lampugnani, Fabrizio Orsenigo, Maria Cristina Gagliani, Carlo Tacchetti, and Elisabetta Dejana. Vascular endothelial cadherin controls VEGFR-2 internalization and signaling from intracellular compartments. Journal of Cell Biology, 174(4):593–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/4/593.

Leal-Ortiz:2008:PMS

[LOWTL^{+}08] Sergio Leal-Ortiz, Clarissa L. Waites, Ryan Terry-Lorenzo, Pedro Zamorano, Eckart D. Gundelfinger, and Craig C. Garner. Piccolo modulation of Synapsin1a dynamics regulates synaptic vesicle exocytosis. Journal of Cell Biology, 181(5):831–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/181/5/
831.

LeBrasseur:2005:CDC

[LP05] Nicole LeBrasseur and Kendall Powell. Cells in D.C.: The American Society for Cell Biology. Journal of Cell Biology, 168 (4):526–??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/4/526.

Lacalle:2007:TPP

[LPA^{+}07] Rosa Ana Lacalle, Rosa M. Peregil, Juan Pablo Albar, Ernesto Merino, Carlos Martínez-A, Isabel Mérida, and Santos Mañes. Type I phosphatidylinositol 4-phosphate 5-kinase controls neutrophil polarity and directional movement. Journal of Cell Biology, 179(7):1539–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/7/1539.

Li:2007:ACC

[LPB^{+}07] Jian Li, Peter J. Peters, Ming Bai, Jun Dai, Erik Bos, Tomas Kirchhausen, Konstantin V. Kandror, and Victor W. Hsu. An ACAP1-containing clathrin coat complex for endocytic recycling. Journal of Cell Biology, 178(3):453–??, July 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/178/3/453.

Lowe:2008:VGC

[LPB^{+}08] John S. Lowe, Oleg Palygin, Naina Bhasin, Thomas J. Hund, Penelope A. Boyden, Erwin Shibata, Mark E. Anderson, and Peter J. Mohler. Voltage-gated Na_{v} channel targeting in the heart requires an ankyrin–G-dependent cellular pathway. Journal of Cell Biology, 180(1):173–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/180/1/173.

Licht:2006:JRE

[LPF^{+}06] Alexander H. Licht, Oliver T. Pein, Lore Florin, Bettina Hartenstein, Hendrik Reuter, Bernd Arnold, Peter Lichter, Peter Angel, and Marina Schorpp-Kistner. JunB is required for endothelial cell morphogenesis by regulating core-binding factor β. Journal of Cell Biology, 175(6):981–??, December 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/175/6/981.

Lajoie:2007:PMD

[LPG^{+}07] Patrick Lajoie, Emily A. Partridge, Ginette Guay, Jacky G. Goetz, Judy Pawling, Annick Lagana, Bharat Joshi, James W. Dennis, and Ivan R. Nabi. Plasma membrane domain organization regulates EGFR signaling in tumor cells. Journal of Cell Biology, 179(2):341–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/2/341.

Liu:2007:FAS

[LPM^{+}07] Qian Liu, Nelly Pante, Tom Misteli, Mohamed Elsagga, Melissa Crisp, Didier Hodzic, Brian Burke, and Kyle J. Roux. Functional association of Sun1 with nuclear pore complexes. Journal of Cell Biology, 178(5):785–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/178/5/785.

Lee:2008:SRC

[LPN^{+}08] Seung-Hye Lee, I.-Feng Peng, Yu Gie Ng, Masahiro Yanagisawa, Shernaz X. Bamji, Lisa P. Elia, Janne Balsamo, Jack Lilien, Panos Z. Anastasiadis, Erik M. Ullian, and Louis F. Reichardt. Synapses are regulated by the cytoplasmic tyrosine kinase Fer in a pathway mediated by p120catenin, Fer, SHP-2, and β-catenin. Journal of Cell Biology, 183(5):893–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/183/5/
893.

Lara-Pezzi:2007:NOC

[LPWP^{+}07] Enrique Lara-Pezzi, Nadine Winn, Angelika Paul, Karl McCullagh, Esfir Slominsky, Maria Paola Santini, Foteini Mourkioti, Padmini Sarathchandra, Satsuki Fukushima, Ken Suzuki, and Nadia Rosenthal. A naturally occurring calcineurin variant inhibits FoxO activity and enhances skeletal muscle regeneration. Journal of Cell Biology, 179(6):1205–??, December 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/179/6/1205.

Luo:2005:CAU

[LR05] Yang Luo and Glenn L. Radice. N-cadherin acts upstream of VE-cadherin in controlling vascular morphogenesis. Journal of Cell Biology, 169(1):29–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/1/29.

Lucas:2007:MPC

[LR07] Eliana P. Lucas and Jordan W. Raff. Maintaining the proper connection between the centrioles and the pericentriolar matrix requires Drosophila Centrosomin. Journal of Cell Biology, 178(5):725–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/5/725.

Lindqvist:2009:DEM

[LRBM09] Arne Lindqvist, Verónica Rodríguez-Bravo, and René H. Medema. The decision to enter mitosis: feedback and redundancy in the mitotic entry network. Journal of Cell Biology, 185(2):193–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/2/193.

Lui-Roberts:2005:ACC

[LRCH^{+}05] Winnie W. Y. Lui-Roberts, Lucy M. Collinson, Lindsay J. Hewlett, Grégoire Michaux, and Daniel F. Cutler. An AP-1/clathrin coat plays a novel and essential role in forming the Weibel–Palade bodies of endothelial cells. Journal of Cell Biology, 170(4):627–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/4/627.

Liu:2006:MAP

[LRJY06] Song-Tao Liu, Jerome B. Rattner, Sandra A. Jablonski, and Tim J. Yen. Mapping the assembly pathways that specify formation of the trilaminar kinetochore plates in human cells. Journal of Cell Biology, 175(1):41–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/175/1/41.

Liu:2005:ADI

[LRS05] Yewei Liu, William R. Randall, and Martin F. Schneider. Activity-dependent and -independent nuclear fluxes of HDAC4 mediated by different kinases in adult skeletal muscle. Journal of Cell Biology, 168(6):887–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/168/6/887.

Li:2006:RPT

[LS06] Ming Li and Danny J. Schnell. Reconstitution of protein targeting to the inner envelope membrane of chloroplasts. Journal of Cell Biology, 175(2):249–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/175/2/249.

Lake:2009:SRA

[LS09] Blue B. Lake and Sergei Y. Sokol. Strabismus regulates asymmetric cell divisions and cell fate determination in the mouse brain. Journal of Cell Biology, 185(1):59–??, April 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/185/1/59.

Lin:2005:ERG

[LSC^{+}05] Fang Lin, Diane S. Sepich, Songhai Chen, Jacek Topczewski, Chunyue Yin, Lilianna Solnica-Krezel, and Heidi Hamm. Essential roles of Gα_{12/13} signaling in distinct cell behaviors driving zebrafish convergence and extension gastrulation movements. Journal of Cell Biology, 169(5):777–??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/169/5/777.

Levy:2006:MND

[LSC^{+}06a] Jennifer R. Levy, Charlotte J. Sumner, Juliane P. Caviston, Mariko K. Tokito, Srikanth Ranganathan, Lee A. Ligon, Karen E. Wallace, Bernadette H. LaMonte, George G. Harmison, Imke Puls, Kenneth H. Fischbeck, and Erika L. F. Holzbaur. A motor neuron disease–associated mutation in p150^{Glued} perturbs dynactin function and induces protein aggregation. Journal of Cell Biology, 172(5):733–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/172/5/
733.

Li:2006:SRS

[LSC^{+}06b] Sam Li, Alan M. Sandercock, Paul Conduit, Carol V. Robinson, Roger L. Williams, and John V. Kilmartin. Structural role of Sfi1p–centrin filaments in budding yeast spindle pole body duplication. Journal of Cell Biology, 173(6):867–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/173/6/867.

Lelouard:2007:RTR

[LSC^{+}07] Hugues Lelouard, Enrico K. Schmidt, Voahirana Camosseto, Giovanna Clavarino, Maurizio Ceppi, Hsiang-Ting Hsu, and Philippe Pierre. Regulation of translation is required for dendritic cell function and survival during activation. Journal of Cell Biology, 179(7):1427–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/7/1427.

Lebreton:2006:FNI

[LSD^{+}06] Alice Lebreton, Cosmin Saveanu, Laurence Decourty, Jean-Christophe Rain, Alain Jacquier, and Micheline Fromont-Racine. A functional network involved in the recycling of nucleocytoplasmic pre-60S factors. Journal of Cell Biology, 173 (3):349–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/3/349.

Laporte:2008:RCL

[LSDFS08] Damien Laporte, Bénédicte Salin, Bertrand Daignan-Fornier, and Isabelle Sagot. Reversible cytoplasmic localization of the proteasome in quiescent yeast cells. Journal of Cell Biology, 181(5):737–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/5/737.

Liao:2007:PII

[LSdWL07] Yi-Chun Liao, Lizhen Si, Ralph W. deVere White, and Su Hao Lo. The phosphotyrosine-independent interaction of DLC-1 and the SH2 domain of cten regulates focal adhesion localization and growth suppression activity of DLC-1. Journal of Cell Biology, 176(1):43–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/1/43.

Lengner:2006:ODS

[LSG^{+}06] Christopher J. Lengner, Heather A. Steinman, James Gagnon, Thomas W. Smith, Janet E. Henderson, Barbara E. Kream, Gary S. Stein, Jane B. Lian, and Stephen N. Jones. Osteoblast differentiation and skeletal development are regulated by Mdm2–p53 signaling. Journal of Cell Biology, 172(6):909–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
172/6/909.

Liu:2008:SCS

[LSIT^{+}08] Yunhao Liu, Daniel Timothy Sweet, Mohamad Irani-Tehrani, Nobuyo Maeda, and Ellie Tzima. Shc coordinates signals from intercellular junctions and integrins to regulate flow-induced inflammation. Journal of Cell Biology, 182(1):185–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/182/1/185.

Lehmann:2005:AMD

[LSM^{+}05] Maik J. Lehmann, Nathan M. Sherer, Carolyn B. Marks, Marc Pypaert, and Walther Mothes. Actin- and myosin-driven movement of viruses along filopodia precedes their entry into cells. Journal of Cell Biology, 170(2):317–??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/170/2/317.

Lee:2008:RGD

[LSNY08] Yong Jik Lee, Amy Szumlanski, Erik Nielsen, and Zhenbiao Yang. Rho-GTPase-dependent filamentous actin dynamics coordinate vesicle targeting and exocytosis during tip growth.

Journal of Cell Biology, 181(7):1155–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/181/7/1155.

Li:2007:MVR

[LSR07] Bingbing X. Li, Akiko K. Satoh, and Donald F. Ready. Myosin V, Rab11, and dRip11 direct apical secretion and cellular morphogenesis in developing Drosophila photoreceptors. Journal of Cell Biology, 177(4):659–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/177/4/659.

Lu:2009:CDP

[LTAM^{+}09] Albert Lu, Francesc Tebar, Blanca Alvarez-Moya, Cristina López-Alcalá, Maria Calvo, Carlos Enrich, Neus Agell, Takeshi Nakamura, Michiyuki Matsuda, and Oriol Bachs. A clathrindependent pathway leads to KRas signaling on late endosomes en route to lysosomes. Journal of Cell Biology, 184(6):863–??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
184/6/863.

Loercher:2005:MLD

[LTDH05] Amy E. Loercher, Elizabeth M. H. Tank, Rachel B. Delston, and J. William Harbour. MITF links differentiation with cell cycle arrest in melanocytes by transcriptional activation of INK4A. Journal of Cell Biology, 168(1):35–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/168/1/35.

Lim:2009:PMP

[LTK09] Ai Khim Lim, Liheng Tao, and Toshie Kai. piRNAs mediate posttranscriptional retroelement silencing and localization to pi-bodies in the Drosophila germline. Journal of Cell Biology, 186(3):333–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/186/3/333.

LaGrassa:2005:VKY

[LU05] Tracy J. LaGrassa and Christian Ungermann. The vacuolar kinase Yck3 maintains organelle fragmentation by regulating the HOPS tethering complex. Journal of Cell Biology, 168(3): 401–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/3/401.

Luo:2005:CPH

[LVDR05] Shouqing Luo, Coralie Vacher, Janet E. Davies, and David C. Rubinsztein. Cdk5 phosphorylation of huntingtin reduces its cleavage by caspases. Journal of Cell Biology, 169(4):647–??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
169/4/647.

Liu:2009:CWF

[LVG^{+}09] Guizhong Liu, Sapna Vijayakumar, Luca Grumolato, Randy Arroyave, HuiFang Qiao, Gal Akiri, and Stuart A. Aaronson. Canonical Wnts function as potent regulators of osteogenesis by human mesenchymal stem cells. Journal of Cell Biology, 185(1):67–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/1/67.

Lee:2005:RIB

[LVL^{+}05] Sang Yoon Lee, Sergey Voronov, Kresimir Letinic, Angus C. Nairn, Gilbert Di Paolo, and Pietro De Camilli. Regulation of the interaction between PIPKIγ and talin by proline-directed protein kinases. Journal of Cell Biology, 168(5):789–??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/168/5/
789.

Lechtreck:2007:CRH

[LW07a] Karl-Ferdinand Lechtreck and George B. Witman. Chlamydomonas reinhardtii hydin is a central pair protein required for flagellar motility. Journal of Cell Biology, 176(4):473–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/176/4/
473.

Leeb:2007:RCR

[LW07b] Martin Leeb and Anton Wutz. Ring1B is crucial for the regulation of developmental control genes and PRC1 proteins but not X inactivation in embryonic cells. Journal of Cell Biology, 178(2):219–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/2/219.

LeBrasseur:2008:CCA

[LW08] Nicole LeBrasseur and Ruth Williams. Capital cells: The American Society for Cell Biology, Washington, D.C., December 1–5, 2007. Journal of Cell Biology, 180(4):654–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/180/4/654.

Luik:2006:EUS

[LWBL06] Riina M. Luik, Minnie M. Wu, JoAnn Buchanan, and Richard S. Lewis. The elementary unit of store-operated Ca^{2+} entry: local activation of CRAC channels by STIM1 at ER– plasma membrane junctions. Journal of Cell Biology, 174(6): 815–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/6/815.

Lee:2009:AAS

[LWC09] Jia-Lin Lee, Mei-Jung Wang, and Jeou-Yuan Chen. Acetylation and activation of STAT3 mediated by nuclear translocation of CD44. Journal of Cell Biology, 185(6):949–??, June 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/185/6/949.

Liu:2005:CRI

[LWG^{+}05] Qing Liu, Simon A. Walker, Dingcheng Gao, James A. Taylor, Yan-Feng Dai, Rebecca S. Arkell, Martin D. Bootman, H. Llewelyn Roderick, Peter J. Cullen, and Peter J. Lockyer. CAPRI and RASAL impose different modes of information processing on Ras due to contrasting temporal filtering of Ca^{2+}. Journal of Cell Biology, 170(2):183–??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/170/2/183.

Li:2008:DMA

[LWHC08] Zhouhua Li, Liwei Wang, Thomas S. Hays, and Yu Cai. Dyneinmediated apical localization of crumbs transcripts is required for Crumbs activity in epithelial polarity. Journal of Cell Biology, 180(1):31–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/1/31.

Wang:2006:CP

[lWHMH06a] Yu li Wang, Klaus M. Hahn, Robert F. Murphy, and Alan F. Horwitz. Correction 174, p. 481. Journal of Cell Biology, 174(6): 903–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/6/903.

Wang:2006:IUF

[lWHMH06b] Yu li Wang, Klaus M. Hahn, Robert F. Murphy, and Alan F. Horwitz. From imaging to understanding: Frontiers in Live Cell Imaging, Bethesda, MD, April 19–21, 2006. Journal of Cell Biology, 174(4):481–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/174/4/481.

Llobet:2008:MDCa

[LWL08a] Artur Llobet, Minnie Wu, and Leon Lagnado. The mouth of a dense-core vesicle opens and closes in a concerted action regulated by calcium and amphiphysin. Journal of Cell Biology, 182(5):1017–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/5/1017.

Llobet:2008:MDCb

[LWL08b] Artur Llobet, Minnie Wu, and Leon Lagnado. The mouth of a dense-core vesicle opens and closes in a concerted action regulated by calcium and amphiphysin. Journal of Cell Biology, 183 (1):169–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/1/169.

Lee:2005:APC

[LYH^{+}05] Stella Y. Lee, Jia-Shu Yang, Wanjin Hong, Richard T. Premont, and Victor W. Hsu. ARFGAP1 plays a central role in coupling COPI cargo sorting with vesicle formation. Journal of Cell Biology, 168(2):281–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/2/281.

Lou:2006:FEA

[LYK^{+}06] Jizhong Lou, Tadayuki Yago, Arkadiusz G. Klopocki, Padmaja Mehta, Wei Chen, Veronika I. Zarnitsyna, Nicolai V. Bovin, Cheng Zhu, and Rodger P. McEver. Flow-enhanced adhesion regulated by a selectin interdomain hinge. Journal of Cell Biology, 174(7):1107–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/7/1107.

Loewen:2007:ICY

[LYTL07] Christopher J. R. Loewen, Barry P. Young, Shabnam Tavassoli, and Timothy P. Levine. Inheritance of cortical ER in yeast is required for normal septin organization. Journal of Cell Biology, 179(3):467–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/3/467.

Ly:2008:SRS

[LYV^{+}08] Cindy V. Ly, Chi-Kuang Yao, Patrik Verstreken, Tomoko Ohyama, and Hugo J. Bellen. straightjacket is required for the synaptic stabilization of cacophony, a voltage-gated calcium channel α_{1} subunit. Journal of Cell Biology, 181(1):157–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
181/1/157.

Leonard:2008:SHS

[LYWK08] Jennifer M. Leonard, Hong Ye, Cynthia Wetmore, and Larry M. Karnitz. Sonic Hedgehog signaling impairs ionizing radiation– induced checkpoint activation and induces genomic instability. Journal of Cell Biology, 183(3):385–??, November 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/183/3/385.

Li:2008:FRP

[LYY08] Juying Li, Helen L. Yin, and Junying Yuan. Flightless-i regulates proinflammatory caspases by selectively modulating intracellular localization and caspase activity. Journal of Cell Biology, 181(2):321–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/2/321.

Ley:2008:DLH

[LZ08] Klaus Ley and Hong Zhang. Dances with leukocytes: how tetraspanin-enriched microdomains assemble to form endothelial adhesive platforms. Journal of Cell Biology, 183(3):375– ??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/3/375.

Liu:2009:REM

[LZ09] Zhonghua Liu and Yixian Zheng. A requirement for epsin in mitotic membrane and spindle organization. Journal of Cell Biology, 186(4):473–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/186/4/473.

Lee:2005:BOA

[LZL^{+}05] Edward B. Lee, Bin Zhang, Kangning Liu, Eric A. Greenbaum, Robert W. Doms, John Q. Trojanowski, and Virginia M.-Y. Lee. BACE overexpression alters the subcellular processing of APP and inhibits Aβ deposition in vivo. Journal of Cell Biology, 168(2):291–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/2/291.

Leon:2006:DPI

[LZM^{+}06] Sébastien Léon, Lan Zhang, W. Hayes McDonald, John Yates, James M. Cregg, and Suresh Subramani. Dynamics of the peroxisomal import cycle of PpPex20p: Ubiquitin-dependent localization and regulation. Journal of Cell Biology, 172(1):67–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
172/1/67.

Martin:2005:AMR

[MAG05] Agnès O. Martin, Gérard Alonso, and Nathalie C. Guérineau. Agrin mediates a rapid switch from electrical coupling to chemical neurotransmission during synaptogenesis. Journal of Cell Biology, 169(3):503–??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/3/503.

Michaelson:2008:RAN

[MAG^{+}08] David Michaelson, Wasif Abidi, Daniele Guardavaccaro, Mo Zhou, Ian Ahearn, Michele Pagano, and Mark R. Philips. Rac1 accumulates in the nucleus during the G2 phase of the cell cycle and promotes cell division. Journal of Cell Biology, 181(3):485–??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/3/485.

Mansukhani:2005:SIF

[MAH^{+}05] Alka Mansukhani, Davide Ambrosetti, Greg Holmes, Lizbeth Cornivelli, and Claudio Basilico. Sox2 induction by FGF and FGFR2 activating mutations inhibits Wnt signaling and osteoblast differentiation. Journal of Cell Biology, 168(7): 1065–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/7/1065.

Mukhopadhyay:2006:SAF

[MAK^{+}06] Debaditya Mukhopadhyay, Ferhan Ayaydin, Nagamalleswari Kolli, Shyh-Han Tan, Tadashi Anan, Ai Kametaka, Yoshiaki Azuma, Keith D. Wilkinson, and Mary Dasso. SUSP1 antagonizes formation of highly SUMO2/3-conjugated species. Journal of Cell Biology, 174(7):939–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/174/7/939.

Malik:2006:HGS

[Mal06] Harmit S. Malik. A hitchhiker’s guide to survival finally makes CENs. Journal of Cell Biology, 174(6):747–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/174/6/747.

McNally:2006:KCM

[MAOM06] Karen McNally, Anjon Audhya, Karen Oegema, and Francis J. McNally. Katanin controls mitotic and meiotic spindle length. Journal of Cell Biology, 175(6):881–??, December 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/175/6/881.

Marks:2006:DDT

[Mar06] Michael S. Marks. Darkness descends with two Rabs. Journal of Cell Biology, 175(2):199–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/175/2/199.

Marshall:2008:CBB

[Mar08] Wallace F. Marshall. The cell biological basis of ciliary disease. Journal of Cell Biology, 180(1):17–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/180/1/17.

Mathur:2005:CBE

[Mat05] Jaideep Mathur. Conservation of boundary extension mechanisms between plants and animals. Journal of Cell Biology, 168 (5):679–??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/5/679.

Matera:2006:DCB

[Mat06] A. Gregory Matera. Drosophila Cajal bodies: accessories not included. Journal of Cell Biology, 172(6):791–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/172/6/791.

Medioni:2008:GCC

[MAZ^{+}08] Caroline Medioni, Martine Astier, Monika Zmojdzian, Krzysztof Jagla, and Michel Sémériva. Genetic control of cell morphogenesis during Drosophila melanogaster cardiac tube formation. Journal of Cell Biology, 182(2):249–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/182/2/249.

Myers:2007:KRG

[MB07] Kenneth A. Myers and Peter W. Baas. Kinesin-5 regulates the growth of the axon by acting as a brake on its microtubule array. Journal of Cell Biology, 178(6):1081–??, September 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/178/6/1081.

Mahadevaiah:2008:EMA

[MBdR^{+}08] Shantha K. Mahadevaiah, Déborah Bourc’his, Dirk G. de Rooij, Timothy H. Bestor, James M. A. Turner, and Paul S. Burgoyne. Extensive meiotic asynapsis in mice antagonises meiotic silencing of unsynapsed chromatin and consequently disrupts meiotic sex chromosome inactivation. Journal of Cell Biology, 182(2):263–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/2/263.

Muller:2007:DAP

[MBF^{+}07] Berndt Müller, Jane Blackburn, Carmen Feijoo, Xiujie Zhao, and Carl Smythe. DNA-activated protein kinase functions in a newly observed S phase checkpoint that links histone mRNA abundance with DNA replication. Journal of Cell Biology, 179(7):1385–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/7/1385.

Muller:2008:DAP

[MBF^{+}08] Berndt Müller, Jane Blackburn, Carmen Feijoo, Xiujie Zhao, and Carl Smythe. DNA-activated protein kinase functions in a newly observed S phase checkpoint that links histone mRNA abundance with DNA replication. Journal of Cell Biology, 180 (4):843–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/4/843.

Melander:2008:PSR

[MBJF^{+}08] Fredrik Melander, Simon Bekker-Jensen, Jacob Falck, Jiri Bartek, Niels Mailand, and Jiri Lukas. Phosphorylation of SDT repeats in the MDC1 N terminus triggers retention of NBS1 at the DNA damage–modified chromatin. Journal of Cell Biology, 181(2):213–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/2/213.

Marteijn:2009:NER

[MBJM^{+}09] Jurgen A. Marteijn, Simon Bekker-Jensen, Niels Mailand, Hannes Lans, Petra Schwertman, Audrey M. Gourdin, Nico P. Dantuma, Jiri Lukas, and Wim Vermeulen. Nucleotide excision repair–induced H2A ubiquitination is dependent on MDC1 and RNF8 and reveals a universal DNA damage response. Journal of Cell Biology, 186(6):835–??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/186/6/835.

Meinen:2007:LMB

[MBL^{+}07] Sarina Meinen, Patrizia Barzaghi, Shuo Lin, Hanns Lochmüller, and Markus A. Ruegg. Linker molecules between laminins and dystroglycan ameliorate laminin-α2–deficient muscular dystrophy at all disease stages. Journal of Cell Biology, 176(7):979–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
176/7/979.

Marti:2005:SME

[MBR^{+}05] Matthias Marti, Jake Baum, Melanie Rug, Leann Tilley, and Alan F. Cowman. Signal-mediated export of proteins from the malaria parasite to the host erythrocyte. Journal of Cell Biology, 171(4):587–??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/4/587.

Ma:2006:HDS

[MBR^{+}06] Peisong Ma, Shannon L. Beck, Ronald W. Raab, Robert L. McKown, George L. Coffman, Atsushi Utani, William J. Chirico, Alan C. Rapraeger, and Gordon W. Laurie. Heparanase deglycanation of syndecan-1 is required for binding of the epithelial-restricted prosecretory mitogen lacritin. Journal of Cell Biology, 174(7):1097–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/174/7/1097.

Mondal:2008:LRM

[MBS^{+}08] Subhanjan Mondal, Deenadayalan Bakthavatsalam, Paul Steimle, Berthold Gassen, Francisco Rivero, and Angelika A. Noegel. Linking Ras to myosin function: RasGEF Q, a Dictyostelium exchange factor for RasB, affects myosin II functions. Journal of Cell Biology, 181(5):747–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/181/5/747.

McDonald:2006:NAE

[MCA^{+}06] Darin McDonald, Gustavo Carrero, Christi Andrin, Gerda de Vries, and Michael J. Hendzel. Nucleoplasmic β-actin exists in a dynamic equilibrium between low-mobility polymeric species and rapidly diffusing populations. Journal of Cell Biology, 172(4):541–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/4/541.

McBride:2008:PMA

[McB08] Heidi M. McBride. Parkin mitochondria in the autophagosome. Journal of Cell Biology, 183(5):757–??, December 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/183/5/757.

Mineur:2007:NIB

[MCD^{+}07] Pierre Mineur, Alain C. Colige, Christophe. F. Deroanne, Johanne Dubail, Frédéric Kesteloot, Yvette Habraken, Agnès Noël, Stefan Vöö, Johannes Waltenberger, Charles M. Lapière, Betty V. Nusgens, and Charles A. Lambert. Newly identified biologically active and proteolysis-resistant VEGF–A isoform VEGF111 is induced by genotoxic agents. Journal of Cell Biology, 179(6):1261–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/6/1261.

Maffucci:2005:CIP

[MCF^{+}05] Tania Maffucci, Frank T. Cooke, Fiona M. Foster, Colin J. Traer, Michael J. Fry, and Marco Falasca. Class II phosphoinositide 3-kinase defines a novel signaling pathway in cell migration. Journal of Cell Biology, 169(5):789–??, June 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/169/5/789.

McNally:2009:TCC

[McN09] James G. McNally. Transcription, chromatin condensation, and gene migration. Journal of Cell Biology, 185(1):7–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/185/1/
7.

Martin-Cofreces:2008:MTM

[MCRVC^{+}08] Noa B. Martín-Cófreces, Javier Robles-Valero, J. Román Cabrero, María Mittelbrunn, Mónica Gordón-Alonso, Ching-Hwa Sung, Balbino Alarcón, Jesús Vázquez, and Francisco Sánchez-Madrid. MTOC translocation modulates IS formation and controls sustained T cell signaling. Journal of Cell Biology, 182(5):951–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/5/951.

Maddugoda:2007:MVV

[MCSY07] Madhavi P. Maddugoda, Matthew S. Crampton, Annette M. Shewan, and Alpha S. Yap. Myosin VI and vinculin cooperate during the morphogenesis of cadherin cell–cell contacts in mammalian epithelial cells. Journal of Cell Biology, 178(3):529–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
178/3/529.

McEwen:2009:RSA

[MD09] Bruce F. McEwen and Yimin Dong. Releasing the spindle assembly checkpoint without tension. Journal of Cell Biology, 184 (3):355–??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/184/3/355.

Mao:2005:MCC

[MDC05] Yinghui Mao, Arshad Desai, and Don W. Cleveland. Microtubule capture by CENP–E silences BubR1-dependent mitotic checkpoint signaling. Journal of Cell Biology, 170(6): 873–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/6/873.

Mazroui:2008:CMC

[MDC^{+}08] Rachid Mazroui, Sergio Di Marco, Eveline Clair, Christopher von Roretz, Scott A. Tenenbaum, Jack D. Keene, Maya Saleh, and Imed-Eddine Gallouzi. Caspase-mediated cleavage of HuR in the cytoplasm contributes to pp32/PHAP–I regulation of apoptosis. Journal of Cell Biology, 180(1):113–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/180/1/
113.

Mebratu:2008:BOP

[MDET08] Yohannes A. Mebratu, Burton F. Dickey, Chris Evans, and Yohannes Tesfaigzi. The BH3-only protein Bik/Blk/Nbk inhibits nuclear translocation of activated ERK1/2 to mediate IFNγ-induced cell death. Journal of Cell Biology, 183(3): 429–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/3/429.

Mills:2006:DMA

[MDH^{+}06] Kathryn Mills, Tasman Daish, Kieran F. Harvey, Cathie M. Pfleger, Iswar K. Hariharan, and Sharad Kumar. The Drosophila melanogaster Apaf-1 homologue ARK is required for most, but not all, programmed cell death. Journal of Cell Biology, 172(6):809–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/6/809.

Mejat:2009:LCM

[MDL^{+}09] Alexandre Méjat, Valérie Decostre, Juan Li, Laure Renou, Akanchha Kesari, Daniel Hantaï, Colin L. Stewart, Xiao Xiao, Eric Hoffman, Gisèle Bonne, and Tom Misteli. Lamin A/Cmediated neuromuscular junction defects in Emery–Dreifuss muscular dystrophy. Journal of Cell Biology, 184(1):31–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
184/1/31.

Montembault:2007:PSA

[MDPG07] Emilie Montembault, Stéphanie Dutertre, Claude Prigent, and Régis Giet. PRP4 is a spindle assembly checkpoint protein required for MPS1, MAD1, and MAD2 localization to the kinetochores. Journal of Cell Biology, 179(4):601–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/179/4/601.

Myers:2005:SLA

[MDW05] Margaret D. Myers, Leonard L. Dragone, and Arthur Weiss. Src-like adaptor protein down-regulates T cell receptor (TCR)– CD3 expression by targeting TCRζ for degradation. Journal of Cell Biology, 170(2):285–??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/170/2/285.

Minter-Dykhouse:2008:DVO

[MDWH^{+}08] Katherine Minter-Dykhouse, Irene Ward, Michael S. Y. Huen, Junjie Chen, and Zhenkun Lou. Distinct versus overlapping functions of MDC1 and 53BP1 in DNA damage response and tumorigenesis. Journal of Cell Biology, 181(5):727–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/181/5/
727.

Maurel:2007:NLP

[MEG^{+}07] Patrice Maurel, Steven Einheber, Jolanta Galinska, Pratik Thaker, Isabel Lam, Marina B. Rubin, Steven S. Scherer, Yoshinuri Murakami, David H. Gutmann, and James L. Salzer. Nectin-like proteins mediate axon–Schwann cell interactions along the internode and are essential for myelination. Journal of Cell Biology, 178(5):861–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/178/5/861.

Mellman:2005:FYC

[Mel05] Ira Mellman. Fifty years of cell biology. Journal of Cell Biology, 168(1):15–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/1/15.

Melzer:2006:WSG

[Mel06] Werner Melzer. When sparks get old. Journal of Cell Biology, 174(5):613–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/5/613.

Meldolesi:2008:IAN

[Mel08a] Jacopo Meldolesi. Inhibition of adipogenesis: a new job for the ER Ca^{2+} pool. Journal of Cell Biology, 182(1):11–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/182/1/
11.

Mellman:2008:GPC

[Mel08b] Ira Mellman. George E. Palade, cell biology and The JCB. Journal of Cell Biology, 183(3):365–??, November 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/183/3/365.

Mellman:2009:SS

[Mel09] Ira Mellman. By the scientists, for the scientists. Journal of Cell Biology, 184(1):7–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/184/1/7.

Mallo:2008:SPP

[MES^{+}08] Gustavo V. Mallo, Marianela Espina, Adam C. Smith, Mauricio R. Terebiznik, Ainel Alemán, B. Brett Finlay, Lucia E. Rameh, Sergio Grinstein, and John H. Brumell. SopB promotes phosphatidylinositol 3-phosphate formation on Salmonella vacuoles by recruiting Rab5 and Vps34. Journal of Cell Biology, 182(4):741–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/4/741.

Matsuzaki:2008:PMP

[MF08] Takashi Matsuzaki and Yukio Fujiki. The peroxisomal membrane protein import receptor Pex3p is directly transported to peroxisomes by a novel Pex19p- and Pex16p-dependent pathway. Journal of Cell Biology, 183(7):1275–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/183/7/
1275.

Madsen:2007:UIC

[MFA^{+}07] Chris D. Madsen, Gian Maria Sarra Ferraris, Annapaola Andolfo, Orla Cunningham, and Nicolai Sidenius. uPAR-induced cell adhesion and migration: vitronectin provides the key. Journal of Cell Biology, 177(5):927–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/177/5/927.

Millon-Fremillon:2008:CAR

[MFBG^{+}08] Angélique Millon-Frémillon, Daniel Bouvard, Alexei Grichine, Sandra Manet-Dupé, Marc R. Block, and Corinne Albiges-Rizo. Cell adaptive response to extracellular matrix density is controlled by ICAP-1-dependent β_{1}-integrin affinity. Journal of Cell Biology, 180(2):427–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/2/427.

Marenduzzo:2006:DAU

[MFC06] Davide Marenduzzo, Kieran Finan, and Peter R. Cook. The depletion attraction: an underappreciated force driving cellular organization. Journal of Cell Biology, 175(5):681–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/175/5/
681.

Meireles:2009:WNA

[MFC^{+}09] Ana M. Meireles, Katherine H. Fisher, Nathalie Colombié, James G. Wakefield, and Hiroyuki Ohkura. Wac: a new Augmin subunit required for chromosome alignment but not for acentrosomal microtubule assembly in female meiosis. Journal of Cell Biology, 184(6):777–??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/184/6/777.

Maresca:2005:HHE

[MFH05] Thomas J. Maresca, Benjamin S. Freedman, and Rebecca Heald. Histone H1 is essential for mitotic chromosome architecture and segregation in Xenopus laevis egg extracts. Journal of Cell Biology, 169(6):859–??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/169/6/859.

Morone:2006:TDR

[MFM^{+}06] Nobuhiro Morone, Takahiro Fujiwara, Kotono Murase, Rinshi S. Kasai, Hiroshi Ike, Shigeki Yuasa, Jiro Usukura, and Akihiro Kusumi. Three-dimensional reconstruction of the membrane skeleton at the plasma membrane interface by electron tomography. Journal of Cell Biology, 174(6):851–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/174/6/
851.

McCleland:2009:ICT

[MFO09] Mark L. McCleland, Jeffrey A. Farrell, and Patrick H. O’Farrell. Influence of cyclin type and dose on mitotic entry and progression in the early Drosophila embryo. Journal of Cell Biology, 184(5):639–??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/5/639.

Maher:2009:ACP

[MFS^{+}09] Meghan T. Maher, Annette S. Flozak, Adam M. Stocker, Anjen Chenn, and Cara J. Gottardi. Activity of the β-catenin phosphodestruction complex at cell–cell contacts is enhanced by cadherin-based adhesion. Journal of Cell Biology, 186(2):219– ??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
186/2/219.

Medicherla:2008:HSO

[MG08a] Balasubrahmanyam Medicherla and Alfred L. Goldberg. Heat shock and oxygen radicals stimulate ubiquitin-dependent degradation mainly of newly synthesized proteins. Journal of Cell Biology, 182(4):663–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/4/663.

Meshorer:2008:GWN

[MG08b] Eran Meshorer and Yosef Gruenbaum. Gone with the Wnt/ Notch: stem cells in laminopathies, progeria, and aging. Journal of Cell Biology, 181(1):9–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/181/1/9.

Messick:2009:ULD

[MG09] Troy E. Messick and Roger A. Greenberg. The ubiquitin landscape at DNA double-strand breaks. Journal of Cell Biology, 187(3):319–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/3/319.

Marciniak:2006:ADS

[MGBH^{+}06] Stefan J. Marciniak, Lidia Garcia-Bonilla, Junjie Hu, Heather P. Harding, and David Ron. Activation-dependent substrate recruitment by the eukaryotic translation initiation factor 2 kinase PERK. Journal of Cell Biology, 172(2):201–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/172/2/
201.

Mulder:2009:NMB

[MGBM^{+}09] Anke M. Mulder, Alex Glavis-Bloom, Carolyn A. Moores, Michael Wagenbach, Bridget Carragher, Linda Wordeman, and Ronald A. Milligan. A new model for binding of kinesin 13 to curved microtubule protofilaments. Journal of Cell Biology, 185(1):51–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/1/51.

Merino:2009:RBO

[MGH^{+}09] Delphine Mérino, Maybelline Giam, Peter D. Hughes, Owen M. Siggs, Klaus Heger, Lorraine A. O’Reilly, Jerry M. Adams, Andreas Strasser, Erinna F. Lee, Walter D. Fairlie, and Philippe Bouillet. The role of BH3-only protein Bim extends beyond inhibiting Bcl-2-like prosurvival proteins. Journal of Cell Biology, 186(3):355–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/186/3/355.

Meaburn:2009:DSG

[MGK^{+}09] Karen J. Meaburn, Prabhakar R. Gudla, Sameena Khan, Stephen J. Lockett, and Tom Misteli. Disease-specific gene repositioning in breast cancer. Journal of Cell Biology, 187(6): 801–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/187/6/801.

Mills:2005:HIP

[MGR^{+}05] Ian G. Mills, Luke Gaughan, Craig Robson, Theodora Ross, Stuart McCracken, John Kelly, and David E. Neal. Huntingtin interacting protein 1 modulates the transcriptional activity of nuclear hormone receptors. Journal of Cell Biology, 170(2):191– ??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
170/2/191.

Madia:2008:LMS

[MGW^{+}08] Federica Madia, Cristina Gattazzo, Min Wei, Paola Fabrizio, William C. Burhans, Martin Weinberger, Abdoulaye Galbani, Jesse R. Smith, Christopher Nguyen, Selina Huey, Lucio Comai, and Valter D. Longo. Longevity mutation in SCH9 prevents recombination errors and premature genomic instability in a Werner/Bloom model system. Journal of Cell Biology, 180 (1):67–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/1/67.

Motley:2007:YPM

[MH07] Alison M. Motley and Ewald H. Hettema. Yeast peroxisomes multiply by growth and division. Journal of Cell Biology, 178 (3):399–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/3/399.

Mohl:2009:DMD

[MHC^{+}09] Dane A. Mohl, Michael J. Huddleston, Therese S. Collingwood, Roland S. Annan, and Raymond J. Deshaies. Dbf2– Mob1 drives relocalization of protein phosphatase Cdc14 to the cytoplasm during exit from mitosis. Journal of Cell Biology, 184(4):527–??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/4/527.

Montpetit:2006:SBY

[MHFH06] Ben Montpetit, Tony R. Hazbun, Stanley Fields, and Philip Hieter. Sumoylation of the budding yeast kinetochore protein Ndc10 is required for Ndc10 spindle localization and regulation of anaphase spindle elongation. Journal of Cell Biology, 174 (5):653–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/5/653.

Madgwick:2006:MER

[MHL^{+}06] Suzanne Madgwick, David V. Hansen, Mark Levasseur, Peter K. Jackson, and Keith T. Jones. Mouse Emi2 is required to enter meiosis II by reestablishing cyclin B1 during interkinesis. Journal of Cell Biology, 174(6):791–??, September 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/174/6/791.

Madesh:2005:SRS

[MHM^{+}05] Muniswamy Madesh, Brian J. Hawkins, Tatyana Milovanova, Cunnigaiper D. Bhanumathy, Suresh K. Joseph, Satish P. RamachandraRao, Kumar Sharma, Tomohiro Kurosaki, and Aron B. Fisher. Selective role for superoxide in InsP_{3} receptormediated mitochondrial dysfunction and endothelial apoptosis. Journal of Cell Biology, 170(7):1079–??, September 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/170/7/1079.

Maddox:2007:FGI

[MHM^{+}07] Paul S. Maddox, Francie Hyndman, Joost Monen, Karen Oegema, and Arshad Desai. Functional genomics identifies a Myb domain–containing protein family required for assembly of CENP–A chromatin. Journal of Cell Biology, 176(6):757–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
176/6/757.

McConnell:2009:EMV

[MHS^{+}09] Russell E. McConnell, James N. Higginbotham, David A. Shifrin, David L. Tabb, Robert J. Coffey, and Matthew J. Tyska. The enterocyte microvillus is a vesicle-generating organelle. Journal of Cell Biology, 185(7):1285–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/185/7/
1285.

Muslimov:2006:SCD

[MIBT06] Ilham A. Muslimov, Anna Iacoangeli, Jürgen Brosius, and Henri Tiedge. Spatial codes in dendritic BC1 RNA. Journal of Cell Biology, 175(3):427–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/175/3/427.

Misteli:2005:GGC

[Mis05] Tom Misteli. Going in GTP cycles in the nucleolus. Journal of Cell Biology, 168(2):177–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/168/2/177.

Misteli:2009:CWM

[Mis09a] Tom Misteli. The changing world of modern cell biology. Journal of Cell Biology, 184(1):11–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/184/1/11.

Misteli:2009:JMC

[Mis09b] Tom Misteli. JCB meeting: Cell Biology of Disease. Journal of Cell Biology, 187(2):155–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/187/2/155.

Mizushima:2005:GAV

[Miz05] Noboru Mizushima. Aβ generation in autophagic vacuoles. Journal of Cell Biology, 171(1):15–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/171/1/15.

Murga:2007:GCC

[MJF^{+}07] Matilde Murga, Isabel Jaco, Yuhong Fan, Rebeca Soria, Barbara Martinez-Pastor, Myriam Cuadrado, Seung-Min Yang, Maria A. Blasco, Arthur I. Skoultchi, and Oscar Fernandez-Capetillo. Global chromatin compaction limits the strength of the DNA damage response. Journal of Cell Biology, 178(7): 1101–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/7/1101.

Moller-Jensen:2007:PSS

[MJG07] Jakob Møller-Jensen and Kenn Gerdes. Plasmid segregation: spatial awareness at the molecular level. Journal of Cell Biology, 179(5):813–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/5/813.

Muhlhausser:2007:VND

[MK07] Petra Mühlhäusser and Ulrike Kutay. An in vitro nuclear disassembly system reveals a role for the RanGTPase system and microtubule-dependent steps in nuclear envelope breakdown. Journal of Cell Biology, 178(4):595–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/178/4/595.

Manzl:2009:CAA

[MKB^{+}09] Claudia Manzl, Gerhard Krumschnabel, Florian Bock, Benedicte Sohm, Verena Labi, Florian Baumgartner, Emmanuelle Logette, Jürg Tschopp, and Andreas Villunger. Caspase-2 activation in the absence of PIDDosome formation. Journal of Cell Biology, 185(2):291–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/2/291.

Mekhail:2005:RUL

[MKC^{+}05] Karim Mekhail, Mireille Khacho, Amanda Carrigan, Robert R. J. Hache, Lakshman Gunaratnam, and Stephen Lee. Regulation of ubiquitin ligase dynamics by the nucleolus. Journal of Cell Biology, 170(5):733–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/170/5/733.

Mimori-Kiyosue:2005:CCB

[MKGL^{+}05] Yuko Mimori-Kiyosue, Ilya Grigoriev, Gideon Lansbergen, Hiroyuki Sasaki, Chiyuki Matsui, Fedor Severin, Niels Galjart, Frank Grosveld, Ivan Vorobjev, Shoichiro Tsukita, and Anna Akhmanova. CLASP1 and CLASP2 bind to EB1 and regulate microtubule plus-end dynamics at the cell cortex. Journal of Cell Biology, 168(1):141–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/168/1/141.

Manderson:2007:SMR

[MKH^{+}07] Anthony P. Manderson, Jason G. Kay, Luke A. Hammond, Darren L. Brown, and Jennifer L. Stow. Subcompartments of the macrophage recycling endosome direct the differential secretion of IL-6 and TNFα. Journal of Cell Biology, 178(1):57–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/178/1/
57.

Munro:2006:RIB

[MKSJ06] Trent P. Munro, Sunjong Kwon, Bruce J. Schnapp, and Daniel St Johnston. A repeated IMP-binding motif controls oskar mRNA translation and anchoring independently of Drosophila melanogaster IMP. Journal of Cell Biology, 172(4): 577–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/4/577.

Muller:2009:HAC

[MKV^{+}09] Mario M. Müller, Esther Klaile, Olga Vorontsova, Bernhard B. Singer, and Björn Öbrink. Homophilic adhesion and CEACAM1-S regulate dimerization of CEACAM1-L and recruitment of SHP-2 and c–Src. Journal of Cell Biology, 187(4): 569–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/187/4/569.

Mao:2005:NRG

[ML05] Yingwei Mao and Angel W.-M. Lee. A novel role for Gab2 in bFGF-mediated cell survival during retinoic acid–induced neuronal differentiation. Journal of Cell Biology, 170(2):305–??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
170/2/305.

Ma:2006:TLR

[MLC^{+}06] Yinghua Ma, Jianxue Li, Isaac Chiu, Yawen Wang, Jacob A. Sloane, Jining Lü, Bela Kosaras, Richard L. Sidman, Joseph J.

Volpe, and Timothy Vartanian. Toll-like receptor 8 functions as a negative regulator of neurite outgrowth and inducer of neuronal apoptosis. Journal of Cell Biology, 175(2):209–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/175/2/
209.

Mandal:2007:CDR

[MLC^{+}07] Atin K. Mandal, Paul Lee, Jennifer A. Chen, Nadinath Nillegoda, Alana Heller, Susan DiStasio, Handy Oen, Jacob Victor, Devi M. Nair, Jeffrey L. Brodsky, and Avrom J. Caplan. Cdc37 has distinct roles in protein kinase quality control that protect nascent chains from degradation and promote posttranslational maturation. Journal of Cell Biology, 176(3):319–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/176/3/
319.

Mukhopadhyay:2006:DMP

[MLH^{+}06] Sudit S. Mukhopadhyay, Kathryn S. Leung, M. John Hicks, Philip J. Hastings, Hagop Youssoufian, and Sharon E. Plon. Defective mitochondrial peroxiredoxin-3 results in sensitivity to oxidative stress in Fanconi anemia. Journal of Cell Biology, 175(2):225–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/2/225.

Morrison:2006:SLR

[MLHS06] Jamie I. Morrison, Sara Lööf, Pingping He, and András Simon. Salamander limb regeneration involves the activation of a multipotent skeletal muscle satellite cell population. Journal of Cell Biology, 172(3):433–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/3/433.

Mueller:2006:SHY

[MLP06] Britta Mueller, Brendan N. Lilley, and Hidde L. Ploegh. SEL1L, the homologue of yeast Hrd3p, is involved in protein dislocation from the mammalian ER. Journal of Cell Biology, 175 (2):261–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/2/261.

Malhas:2007:DLB

[MLS^{+}07] Ashraf Malhas, Chiu Fan Lee, Rebecca Sanders, Nigel J. Saunders, and David J. Vaux. Defects in lamin B1 expression or processing affect interphase chromosome position and gene expression. Journal of Cell Biology, 176(5):593–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/176/5/593.

Malhas:2009:LBC

[MLV09] Ashraf N. Malhas, Chiu Fan Lee, and David J. Vaux. Lamin B1 controls oxidative stress responses via Oct-1. Journal of Cell Biology, 184(1):45–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/1/45.

Muresan:2005:CTP

[MM05] Zoia Muresan and Virgil Muresan. Coordinated transport of phosphorylated amyloid-β precursor protein and c–Jun NH_{2}- terminal kinase–interacting protein-1. Journal of Cell Biology, 171(4):615–??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/4/615.

Meaburn:2008:LSA

[MM08] Karen J. Meaburn and Tom Misteli. Locus-specific and activityindependent gene repositioning during early tumorigenesis. Journal of Cell Biology, 180(1):39–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/180/1/39.

Marshall:2008:TDL

[MMC08] Owen J. Marshall, Alan T. Marshall, and K. H. Andy Choo. Three-dimensional localization of CENP–A suggests a complex higher order structure of centromeric chromatin. Journal of Cell Biology, 183(7):1193–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/183/7/1193.

Madrid:2006:RIM

[MMCW06] Alexis S. Madrid, Joel Mancuso, W. Zacheus Cande, and Karsten Weis. The role of the integral membrane nucleoporins Ndc1p and Pom152p in nuclear pore complex assembly and function. Journal of Cell Biology, 173(3):361–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/173/3/361.

McGill:2009:ICC

[MMH09] Melanie A. McGill, R. F. Andrew McKinley, and Tony J. C. Harris. Independent cadherin–catenin and Bazooka clusters interact to assemble adherens junctions. Journal of Cell Biology, 185(5):787–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/5/787.

Matsushita:2005:HPR

[MMM^{+}05] Kenji Matsushita, Craig N. Morrell, Rebecca J. A. Mason, Munekazu Yamakuchi, Firdous A. Khanday, Kaikobad Irani, and Charles J. Lowenstein. Hydrogen peroxide regulation of endothelial exocytosis by inhibition of N-ethylmaleimide sensitive factor. Journal of Cell Biology, 170(1):73–??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/170/1/73.

Munck:2009:DFP

[MMNH09] Joanne M. Munck, Alison M. Motley, James M. Nuttall, and Ewald H. Hettema. A dual function for Pex3p in peroxisome formation and inheritance. Journal of Cell Biology, 187(4): 463–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/187/4/463.

Morita:2007:DRM

[MMS07] Tsuyoshi Morita, Taira Mayanagi, and Kenji Sobue. Dual roles of myocardin-related transcription factors in epithelial– mesenchymal transition via slug induction and actin remodeling. Journal of Cell Biology, 179(5):1027–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/5/
1027.

Miyamoto:2005:TJS

[MMT^{+}05] Tatsuo Miyamoto, Kazumasa Morita, Daisuke Takemoto, Kosei Takeuchi, Yuka Kitano, Tsuyoshi Miyakawa, Kiyomi Nakayama, Yasushi Okamura, Hiroyuki Sasaki, Yoshiki Miyachi, Mikio Furuse, and Shoichiro Tsukita. Tight junctions in Schwann cells of peripheral myelinated axons. Journal of Cell Biology, 169(3):527–??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/3/527.

Mizoguchi:2009:ICC

[MMU^{+}09] Toshihide Mizoguchi, Akinori Muto, Nobuyuki Udagawa, Atsushi Arai, Teruhito Yamashita, Akihiro Hosoya, Tadashi Ninomiya, Hiroaki Nakamura, Yohei Yamamoto, Saya Kinugawa, Midori Nakamura, Yuko Nakamichi, Yasuhiro Kobayashi, Sakae Nagasawa, Kimimitsu Oda, Hirofumi Tanaka, Mitsuo Tagaya, Josef M. Penninger, Michio Ito, and Naoyuki Takahashi. Identification of cell cycle–arrested quiescent osteoclast precursors in vivo. Journal of Cell Biology, 184(4):541–??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/184/4/541.

Ma:2007:PMA

[MMW^{+}07] Zhenyi Ma, David P. Myers, Ru Feng Wu, Fiemu E. Nwariaku, and Lance S. Terada. p66^{Shc} mediates anoikis through RhoA. Journal of Cell Biology, 179(1):23–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/1/23.

Margalit:2007:BAF

[MNF^{+}07] Ayelet Margalit, Esther Neufeld, Naomi Feinstein, Katherine L. Wilson, Benjamin Podbilewicz, and Yosef Gruenbaum. Barrier to autointegration factor blocks premature cell fusion and maintains adult muscle integrity in C. elegans. Journal of Cell Biology, 178(4):661–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/4/661.

Meier:2005:PRT

[MNH05] Stephan Meier, Walter Neupert, and Johannes M. Herrmann. Proline residues of transmembrane domains determine the sorting of inner membrane proteins in mitochondria. Journal of Cell Biology, 170(6):881–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/170/6/881.

Menon:2005:PFL

[MOCC05] Sree Devi Menon, Zalina Osman, Kho Chenchill, and William Chia. A positive feedback loop between Dumbfounded and Rolling pebbles leads to myotube enlargement in Drosophila. Journal of Cell Biology, 169(6):909–??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/169/6/909.

Moriyama:2006:NSH

[MOM^{+}06] Mariko Moriyama, Masatake Osawa, Siu-Shan Mak, Toshiyuki Ohtsuka, Norio Yamamoto, Hua Han, Véronique Delmas, Ryoichiro Kageyama, Friedrich Beermann, Lionel Larue, and Shin-Ichi Nishikawa. Notch signaling via Hes1 transcription factor maintains survival of melanoblasts and melanocyte stem cells. Journal of Cell Biology, 173(3):333–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/173/3/333.

Maerki:2009:CKE

[MOS^{+}09] Sarah Maerki, Michael H. Olma, Titu Staubli, Patrick Steigemann, Daniel W. Gerlich, Manfredo Quadroni, Izabela Sumara, and Matthias Peter. The Cul3–KLHL21 E3 ubiquitin ligase targets Aurora B to midzone microtubules in anaphase and is required for cytokinesis. Journal of Cell Biology, 187(6): 791–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/187/6/791.

Massaro:2009:MME

[MPD09] Catherine M. Massaro, Jan Pielage, and Graeme W. Davis. Molecular mechanisms that enhance synapse stability despite persistent disruption of the spectrin/ankyrin/microtubule cytoskeleton. Journal of Cell Biology, 187(1):101–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/187/1/
101.

Merino:2009:NKI

[MPG^{+}09] Carlos Merino, Jay Penney, Miranda González, Kazuya Tsurudome, Myriam Moujahidine, Michael B. O’Connor, Esther M. Verheyen, and Pejmun Haghighi. Nemo kinase interacts with Mad to coordinate synaptic growth at the Drosophila neuromuscular junction. Journal of Cell Biology, 185(4):713–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/185/4/
713.

Maekawa:2007:YCT

[MPL^{+}07a] Hiromi Maekawa, Claire Priest, Johannes Lechner, Gislene Pereira, and Elmar Schiebel. The yeast centrosome translates the positional information of the anaphase spindle into a cell cycle signal. Journal of Cell Biology, 179(3):423–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/3/
423.

Makhnevych:2007:RKR

[MPL^{+}07b] Taras Makhnevych, Christopher Ptak, C. Patrick Lusk, John D. Aitchison, and Richard W. Wozniak. The role of karyopherins in the regulated sumoylation of septins. Journal of Cell Biology, 177(1):39–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/1/39.

Matos:2009:SCS

[MPLF^{+}09] Irina Matos, António J. Pereira, Mariana Lince-Faria, Lisa A. Cameron, Edward D. Salmon, and Helder Maiato. Synchronizing chromosome segregation by flux-dependent force equalization at kinetochores. Journal of Cell Biology, 186(1):11–??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/186/1/
11.

Mattila:2007:MMI

[MPS^{+}07] Pieta K. Mattila, Anette Pykäläinen, Juha Saarikangas, Ville O. Paavilainen, Helena Vihinen, Eija Jokitalo, and Pekka Lappalainen. Missing-in-metastasis and IRSp53 deform PI(4,5)P_{2}-rich membranes by an inverse BAR domainlike mechanism. Journal of Cell Biology, 176(7):953–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/176/7/
953.

Miller:2006:ULI

[MQL^{+}06] Rachel K. Miller, Hiroshi Qadota, Megan L. Landsverk, Kristina B. Mercer, Henry F. Epstein, and Guy M. Benian. UNC-98 links an integrin-associated complex to thick filaments in Caenorhabditis elegans muscle. Journal of Cell Biology, 175 (6):853–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/6/853.

Ma:2008:KMC

[MQWP08] Yan-Qing Ma, Jun Qin, Chuanyue Wu, and Edward F. Plow. Kindlin-2 (Mig-2): a co-activator of β_{3} integrins. Journal of Cell Biology, 181(3):439–??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/3/439.

Medici:2008:FKS

[MRD^{+}08] Damian Medici, Mohammed S. Razzaque, Stephelynn DeLuca, Trent L. Rector, Bo Hou, Kihwa Kang, Regina Goetz, Moosa Mohammadi, Makoto Kuro-o, Bjorn R. Olsen, and Beate Lanske. FGF-23–klotho signaling stimulates proliferation and prevents vitamin D–induced apoptosis. Journal of Cell Biology, 182(3):459–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/3/459.

Muller:2007:OCH

[MRK^{+}07] Waltraud G. Müller, Dietmar Rieder, Tatiana S. Karpova, Sam John, Zlatko Trajanoski, and James G. McNally. Organization of chromatin and histone modifications at a transcription site. Journal of Cell Biology, 177(6):957–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/177/6/957.

Muenzner:2005:CEH

[MRKH05] Petra Muenzner, Manfred Rohde, Susanne Kneitz, and Christof R. Hauck. CEACAM engagement by human pathogens enhances cell adhesion and counteracts bacteria-induced detachment of epithelial cells. Journal of Cell Biology, 170(5): 825–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/5/825.

Mortusewicz:2008:RRP

[MRL^{+}08] Oliver Mortusewicz, Wera Roth, Na Li, M. Cristina Cardoso, Michael Meisterernst, and Heinrich Leonhardt. Recruitment of RNA polymerase II cofactor PC4 to DNA damage sites. Journal of Cell Biology, 183(5):769–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/183/5/769.

Mertens:2005:RAT

[MRO^{+}05] Alexander E. E. Mertens, Tomasz P. Rygiel, Cristina Olivo, Rob van der Kammen, and John G. Collard. The Rac activator Tiam1 controls tight junction biogenesis in keratinocytes through binding to and activation of the Par polarity complex. Journal of Cell Biology, 170(7):1029–??, September 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/170/7/1029.

Moore:2005:MAT

[MRvD^{+}05] Ayana T. Moore, Kathleen E. Rankin, George von Dassow, Leticia Peris, Michael Wagenbach, Yulia Ovechkina, Annie Andrieux, Didier Job, and Linda Wordeman. MCAK associates with the tips of polymerizing microtubules. Journal of Cell Biology, 169(3):391–??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/3/391.

Miller:2006:DEC

[MS06] Kyle E. Miller and Michael P. Sheetz. Direct evidence for coherent low velocity axonal transport of mitochondria. Journal of Cell Biology, 173(3):373–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/173/3/373.

Motegi:2007:RRM

[MS07] Fumio Motegi and Geraldine Seydoux. Revisiting the role of microtubules in C. elegans polarity. Journal of Cell Biology, 179 (3):367–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/3/367.

Miaczynska:2008:MFE

[MS08] Marta Miaczynska and Harald Stenmark. Mechanisms and functions of endocytosis. Journal of Cell Biology, 180(1):7–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
180/1/7.

Maresca:2009:ISA

[MS09] Thomas J. Maresca and Edward D. Salmon. Intrakinetochore stretch is associated with changes in kinetochore phosphorylation and spindle assembly checkpoint activity. Journal of Cell Biology, 184(3):373–??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/184/3/373.

Molk:2006:NCD

[MSB06] Jeffrey N. Molk, E. D. Salmon, and Kerry Bloom. Nuclear congression is driven by cytoplasmic microtubule plus end interactions in S. cerevisiae. Journal of Cell Biology, 172(1):27–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
172/1/27.

Morency:2007:NCR

[MSC^{+}07] Eric Morency, Mirna Sabra, Frédéric Catez, Pascale Texier, and Patrick Lomonte. A novel cell response triggered by interphase centromere structural instability. Journal of Cell Biology, 177 (5):757–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/5/757.

Martin-Serrano:2005:HUL

[MSECB05] Juan Martin-Serrano, Scott W. Eastman, Wayne Chung, and Paul D. Bieniasz. HECT ubiquitin ligases link viral and cellular PPXY motifs to the vacuolar protein-sorting pathway. Journal of Cell Biology, 168(1):89–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/168/1/89.

Matsumoto:2005:SPN

[MSH^{+}05] Gen Matsumoto, Aleksandar Stojanovic, Carina I. Holmberg, Soojin Kim, and Richard I. Morimoto. Structural properties and neuronal toxicity of amyotrophic lateral sclerosis-associated Cu/Zn superoxide dismutase 1 aggregates. Journal of Cell Biology, 171(1):75–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/1/75.

Muto:2005:LRC

[MSK05] Etsuko Muto, Hiroyuki Sakai, and Kuniyoshi Kaseda. Longrange cooperative binding of kinesin to a microtubule in the presence of ATP. Journal of Cell Biology, 168(5):691–??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/168/5/
691.

Mirouse:2007:LAM

[MSK^{+}07] Vincent Mirouse, Lance L. Swick, Nevzat Kazgan, Daniel St Johnston, and Jay E. Brenman. LKB1 and AMPK maintain epithelial cell polarity under energetic stress. Journal of Cell Biology, 177(3):387–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/3/387.

Makio:2009:NNN

[MSL^{+}09] Tadashi Makio, Leslie H. Stanton, Cheng-Chao Lin, David S. Goldfarb, Karsten Weis, and Richard W. Wozniak. The nucleoporins Nup170p and Nup157p are essential for nuclear pore complex assembly. Journal of Cell Biology, 185(3):459–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/185/3/
459.

Maximov:2007:SSV

[MSLS07] Anton Maximov, Ok-Ho Shin, Xinran Liu, and Thomas C. Südhof. Synaptotagmin-12, a synaptic vesicle phosphoprotein that modulates spontaneous neurotransmitter release. Journal of Cell Biology, 176(1):113–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/176/1/113.

Motegi:2006:HSS

[MSM^{+}06] Akira Motegi, Raman Sood, Helen Moinova, Sanford D. Markowitz, Pu Paul Liu, and Kyungjae Myung. Human SH-PRH suppresses genomic instability through proliferating cell nuclear antigen polyubiquitination. Journal of Cell Biology, 175 (5):703–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/5/703.

Michelsen:2007:NCB

[MSM^{+}07] Kai Michelsen, Volker Schmid, Jutta Metz, Katja Heusser, Urban Liebel, Torsten Schwede, Anne Spang, and Blanche Schwappach. Novel cargo-binding site in the β and δ subunits of coatomer. Journal of Cell Biology, 179(2):209–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/2/
209.

McCleland:2009:DRT

[MSO09] Mark L. McCleland, Antony W. Shermoen, and Patrick H. O’Farrell. DNA replication times the cell cycle and contributes to the mid-blastula transition in Drosophila embryos. Journal of Cell Biology, 187(1):7–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/187/1/7.

Milenkovic:2009:LTS

[MSR09] Ljiljana Milenkovic, Matthew P. Scott, and Rajat Rohatgi. Lateral transport of Smoothened from the plasma membrane to the membrane of the cilium. Journal of Cell Biology, 187(3): 365–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/187/3/365.

Meder:2005:GPN

[MSSF05] Doris Meder, Anna Shevchenko, Kai Simons, and Joachim Füllekrug. Gp135/ podocalyxin and NHERF-2 participate in the formation of a preapical domain during polarization of MDCK cells. Journal of Cell Biology, 168(2):303–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/168/2/
303.

McElhinny:2005:NRA

[MSV^{+}05] Abigail S. McElhinny, Catherine Schwach, Melinda Valichnac, Sarah Mount-Patrick, and Carol C. Gregorio. Nebulin regulates the assembly and lengths of the thin filaments in striated muscle. Journal of Cell Biology, 170(6):947–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/170/6/947.

Melloy:2007:NFD

[MSW^{+}07] Patricia Melloy, Shu Shen, Erin White, J. Richard McIntosh, and Mark D. Rose. Nuclear fusion during yeast mating occurs by a three-step pathway. Journal of Cell Biology, 179(4): 659–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress. org/content/179/4/659. See correction [MSW^{+}17].

Melloy:2017:CNF

[MSW^{+}17] Patricia Melloy, Shu Shen, Erin White, J. Richard McIntosh, and Mark D. Rose. Correction: Nuclear fusion during yeast mating occurs by a three-step pathway. Journal of Cell Biology, 216(11):3883–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/11/3883. See [MSW^{+}07].

McConnell:2007:MPS

[MT07] Russell E. McConnell and Matthew J. Tyska. Myosin-1a powers the sliding of apical membrane along microvillar actin bundles. Journal of Cell Biology, 177(4):671–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/177/4/671.

Mennella:2009:MDP

[MTB^{+}09] Vito Mennella, Dong-Yan Tan, Daniel W. Buster, Ana B. Asenjo, Uttama Rath, Ao Ma, Hernando J. Sosa, and David J. Sharp. Motor domain phosphorylation and regulation of the Drosophila kinesin 13, KLP10A. Journal of Cell Biology, 186 (4):481–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/186/4/481.

Miyoshi:2006:ATD

[MTH^{+}06] Takushi Miyoshi, Takahiro Tsuji, Chiharu Higashida, Maud Hertzog, Akiko Fujita, Shuh Narumiya, Giorgio Scita, and Naoki Watanabe. Actin turnover-dependent fast dissociation of capping protein in the dendritic nucleation actin network: evidence of frequent filament severing. Journal of Cell Biology, 175(6):947–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/6/947.

Maas:2006:NCG

[MTL^{+}06] Christoph Maas, Nadia Tagnaouti, Sven Loebrich, Bardo Behrend, Corinna Lappe-Siefke, and Matthias Kneussel. Neuronal cotransport of glycine receptor and the scaffold protein gephyrin. Journal of Cell Biology, 172(3):441–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/172/3/441.

Machin:2006:TRG

[MTRD^{+}06] Felix Machín, Jordi Torres-Rosell, Giacomo De Piccoli, Jesús A. Carballo, Rita S. Cha, Adam Jarmuz, and Luis Aragón. Transcription of ribosomal genes can cause nondisjunction. Journal of Cell Biology, 173(6):893–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/173/6/893.

Machin:2005:SIC

[MTRJA05] Félix Machín, Jordi Torres-Rosell, Adam Jarmuz, and Luis Aragón. Spindle-independent condensation-mediated segregation of yeast ribosomal DNA in late anaphase. Journal of Cell Biology, 168(2):209–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/2/209.

Meister:2005:TSR

[MTV^{+}05] Peter Meister, Angela Taddei, Laurence Vernis, Mickaël Poidevin, Susan M. Gasser, and Giuseppe Baldacci. Temporal separation of replication and recombination requires the intra– S checkpoint. Journal of Cell Biology, 168(4):537–??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/168/4/
537.

Muallem:2005:DCS

[Mua05] Shmuel Muallem. Decoding Ca^{2+} signals. Journal of Cell Biology, 170(2):173–??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/2/173.

Matsu-ura:2006:CIT

[MuMI^{+}06] Toru Matsu-ura, Takayuki Michikawa, Takafumi Inoue, Atsushi Miyawaki, Manabu Yoshida, and Katsuhiko Mikoshiba. Cytosolic inositol 1,4,5-trisphosphate dynamics during intracellular calcium oscillations in living cells. Journal of Cell Biology, 173(5):755–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/5/755.

Marangos:2007:PAP

[MVC^{+}07] Petros Marangos, Emmy W. Verschuren, Ruby Chen, Peter K. Jackson, and John Carroll. Prophase I arrest and progression to metaphase I in mouse oocytes are controlled by Emi1-dependent regulation of APC^{Cdh1}. Journal of Cell Biology, 176(1):65–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
176/1/65.

Millard:2006:RDR

[MW06] Susan M. Millard and Stephen A. Wood. Riding the DUBway: regulation of protein trafficking by deubiquitylating enzymes. Journal of Cell Biology, 173(4):463–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/173/4/463.

Makagiansar:2007:DPN

[MWMS07] Irwan T. Makagiansar, Scott Williams, Tomas Mustelin, and William B. Stallcup. Differential phosphorylation of NG2 proteoglycan by ERK and PKCα helps balance cell proliferation and migration. Journal of Cell Biology, 178(1):155–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/178/1/
155.

Min:2006:SBT

[MWSK06] Guangwei Min, Huaibin Wang, Tung-Tien Sun, and Xiang-Peng Kong. Structural basis for tetraspanin functions as revealed by the cryo–EM structure of uroplakin complexes at 6-Å resolution. Journal of Cell Biology, 173(6):975–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/173/6/975.

Madia:2009:OHS

[MWY^{+}09] Federica Madia, Min Wei, Valerie Yuan, Jia Hu, Cristina Gattazzo, Phuong Pham, Myron F. Goodman, and Valter D.

Longo. Oncogene homologue Sch9 promotes age-dependent mutations by a superoxide and Rev1/Polζ-dependent mechanism. Journal of Cell Biology, 186(4):509–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/186/4/509.

Martin:2005:EAA

[MXR^{+}05] Adam C. Martin, Xiao-Ping Xu, Isabelle Rouiller, Marko Kaksonen, Yidi Sun, Lisa Belmont, Niels Volkmann, Dorit Hanein, Matthew Welch, and David G. Drubin. Effects of Arp2 and Arp3 nucleotide-binding pocket mutations on Arp2/3 complex function. Journal of Cell Biology, 168(2):315–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/168/2/315.

Mardakheh:2009:SIL

[MYMH09] Faraz K. Mardakheh, Mona Yekezare, Laura M. Machesky, and John K. Heath. Spred2 interaction with the late endosomal protein NBR1 down-regulates fibroblast growth factor receptor signaling. Journal of Cell Biology, 187(2):265–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/187/2/
265.

Mangahas:2008:SGR

[MYMZ08] Paolo M. Mangahas, Xiaomeng Yu, Kenneth G. Miller, and Zheng Zhou. The small GTPase Rab2 functions in the removal of apoptotic cells in Caenorhabditis elegans. Journal of Cell Biology, 180(2):357–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/2/357.

Maehara:2005:RTE

[MYO^{+}05] Kayoko Maehara, Kimi Yamakoshi, Naoko Ohtani, Yoshiaki Kubo, Akiko Takahashi, Seiji Arase, Nic Jones, and Eiji Hara. Reduction of total E2F/DP activity induces senescence-like cell cycle arrest in cancer cells lacking functional pRB and p53. Journal of Cell Biology, 168(4):553–??, February 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/168/4/553.

Mao:2009:EUR

[MYZ^{+}09] Yuntao S. Mao, Masaki Yamaga, Xiaohui Zhu, Yongjie Wei, Hui-Qiao Sun, Jing Wang, Mia Yun, Yanfeng Wang, Gilbert Di Paolo, Michael Bennett, Ira Mellman, Charles S. Abrams, Pietro De Camilli, Christopher Y. Lu, and Helen L. Yin. Essential and unique roles of PIP5K-γ and -α in Fcγ receptormediated phagocytosis. Journal of Cell Biology, 184(2):281–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
184/2/281.

McShane:2008:SWS

[MZ08] Marisa P. McShane and Marino Zerial. Survival of the weakest: signaling aided by endosomes. Journal of Cell Biology, 182(5): 823–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/5/823.

Mudhasani:2008:LMB

[MZH^{+}08] Rajini Mudhasani, Zhiqing Zhu, Gyorgy Hutvagner, Christine M. Eischen, Stephen Lyle, Lisa L. Hall, Jeanne B. Lawrence, Anthony N. Imbalzano, and Stephen N. Jones. Loss of miRNA biogenesis induces p19^{Arf}-p53 signaling and senescence in primary cells. Journal of Cell Biology, 181(7):1055–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
181/7/1055.

Nakamura:2009:CAC

[NAF09] Akito Nakamura, Hiroyuki Arai, and Naoya Fujita. Centrosomal Aki1 and cohesin function in separase-regulated centriole disengagement. Journal of Cell Biology, 187(5):607–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/187/5/
607.

Noble:2008:MMR

[NAG^{+}08] Scott L. Noble, Brittany L. Allen, Lai Kuan Goh, Kristen Nordick, and Thomas C. Evans. Maternal mRNAs are regulated by diverse P body–related mRNP granules during early Caenorhabditis elegans development. Journal of Cell Biology, 182(3):559–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/3/559.

Nemethova:2008:BAC

[NAS08] Maria Nemethova, Sonja Auinger, and J. Victor Small. Building the actin cytoskeleton: filopodia contribute to the construction of contractile bundles in the lamella. Journal of Cell Biology, 180(6):1233–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/6/1233.

NDiaye:2009:TTR

[NBB^{+}09] Elsa-Noah N’Diaye, Catherine S. Branda, Steven S. Branda, Lisette Nevarez, Marco Colonna, Clifford Lowell, Jessica A. Hamerman, and William E. Seaman. TREM-2 (triggering receptor expressed on myeloid cells 2) is a phagocytic receptor for bacteria. Journal of Cell Biology, 184(2):215–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/184/2/
215.

Nath:2008:LAE

[NBM^{+}08] Anjali K. Nath, Rachel M. Brown, Michael Michaud, M. Rocio Sierra-Honigmann, Michael Snyder, and Joseph A. Madri. Leptin affects endocardial cushion formation by modulating EMT and migration via Akt signaling cascades. Journal of Cell Biology, 181(2):367–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/2/367.

Nikonenko:2008:PPS

[NBS^{+}08] Irina Nikonenko, Bernadett Boda, Sylvain Steen, Graham Knott, Egbert Welker, and Dominique Muller. PSD-95 promotes synaptogenesis and multiinnervated spine formation through nitric oxide signaling. Journal of Cell Biology, 183(6): 1115–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/6/1115.

Nagao:2008:CCS

[NCB^{+}08] Motoshi Nagao, Kenneth Campbell, Kevin Burns, Chia-Yi Kuan, Andreas Trumpp, and Masato Nakafuku. Coordinated control of self-renewal and differentiation of neural stem cells by Myc and the p19^{ARF}-p53 pathway. Journal of Cell Biology, 183(7):1243–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/7/1243.

Nagao:2009:CCS

[NCB^{+}09] Motoshi Nagao, Kenneth Campbell, Kevin Burns, Chia-Yi Kuan, Andreas Trumpp, and Masato Nakafuku. Coordinated control of self-renewal and differentiation of neural stem cells by Myc and the p19^{ARF}-p53 pathway. Journal of Cell Biology, 184(2):335–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/2/335.

Niepel:2005:NPC

[NdCF^{+}05] Mario Niepel, Caterina Strambio de Castillia, Joseph Fasolo, Brian T. Chait, and Michael P. Rout. The nuclear pore complex–associated protein, Mlp2p, binds to the yeast spindle pole body and promotes its efficient assembly. Journal of Cell Biology, 170(2):225–??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/2/225.

Niikura:2007:BMC

[NDS^{+}07] Yohei Niikura, Amruta Dixit, Ray Scott, Guy Perkins, and Katsumi Kitagawa. BUB1 mediation of caspase-independent mitotic death determines cell fate. Journal of Cell Biology, 178(2):283–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/2/283.

Nakamori:2006:MMM

[NEF^{+}06] Yoshitaka Nakamori, Masahiro Emoto, Naofumi Fukuda, Akihiko Taguchi, Shigeru Okuya, Michiko Tajiri, Makoto Miyagishi, Kazunari Taira, Yoshinao Wada, and Yukio Tanizawa. Myosin motor Myo1c and its receptor NEMO/IKK-γ promote TNF-α-induced serine^{307} phosphorylation of IRS-1. Journal of Cell Biology, 173(5):665–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/173/5/665.

Niessen:2008:SDN

[NFC^{+}08] Kyle Niessen, YangXin Fu, Linda Chang, Pamela A. Hoodless, Deborah McFadden, and Aly Karsan. Slug is a direct Notch target required for initiation of cardiac cushion cellularization. Journal of Cell Biology, 182(2):315–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/182/2/315.

Nokes:2008:RRM

[NFCF08] Rita L. Nokes, Ian C. Fields, Ruth N. Collins, and Heike Fölsch. Rab13 regulates membrane trafficking between TGN and recycling endosomes in polarized epithelial cells. Journal of Cell Biology, 182(5):845–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/182/5/845.

Noritake:2009:MDP

[NFI^{+}09] Jun Noritake, Yuko Fukata, Tsuyoshi Iwanaga, Naoki Hosomi, Ryouhei Tsutsumi, Naoto Matsuda, Hideki Tani, Hiroko Iwanari, Yasuhiro Mochizuki, Tatsuhiko Kodama, Yoshiharu Matsuura, David S. Bredt, Takao Hamakubo, and Masaki Fukata. Mobile DHHC palmitoylating enzyme mediates activity-sensitive synaptic targeting of PSD-95. Journal of Cell Biology, 186(1):147–??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/186/1/147.

Nowak:2009:TRM

[NFZ^{+}09] Roberta B. Nowak, Robert S. Fischer, Rebecca K. Zoltoski, Jerome R. Kuszak, and Velia M. Fowler. Tropomodulin1 is required for membrane skeleton organization and hexagonal geometry of fiber cells in the mouse lens. Journal of Cell Biology, 186(6):915–??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/186/6/915.

Noda:2008:RIG

[NHK^{+}08] Yumi Noda, Saburo Horikawa, Eiichiro Kanda, Maho Yamashita, Hu Meng, Kayoko Eto, Yuhua Li, Michio Kuwahara, Keiji Hirai, Changi Pack, Masataka Kinjo, Shigeo Okabe, and Sei Sasaki. Reciprocal interaction with G-actin and tropomyosin is essential for aquaporin-2 trafficking. Journal of Cell Biology, 182(3):587–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/182/3/587.

Nichols:2009:ELA

[Nic09a] Ben Nichols. Endocytosis of lipid-anchored proteins: excluding GEECs from the crowd. Journal of Cell Biology, 186(4):457–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
186/4/457.

Norrmen:2009:FCF

[NIC^{+}09b] Camilla Norrmén, Konstantin I. Ivanov, Jianpin Cheng, Nadine Zangger, Mauro Delorenzi, Muriel Jaquet, Naoyuki Miura, Pauli Puolakkainen, Valerie Horsley, Junhao Hu, Hellmut G. Augustin, Seppo Ylä-Herttuala, Kari Alitalo, and Tatiana V. Petrova. FOXC2 controls formation and maturation of lymphatic collecting vessels through cooperation with NFATc1. Journal of Cell Biology, 185(3):439–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/185/3/439.

Nakamura:2006:NRN

[NJS^{+}06] Tomoe Y. Nakamura, Andreas Jeromin, George Smith, Hideaki Kurushima, Hitoshi Koga, Yusaku Nakabeppu, Shigeo Wakabayashi, and Junichi Nabekura. Novel role of neuronal Ca^{2+} sensor-1 as a survival factor up-regulated in injured neurons. Journal of Cell Biology, 172(7):1081–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/172/7/1081.

Nydegger:2006:MTE

[NKK^{+}06] Sascha Nydegger, Sandhya Khurana, Dimitry N. Krementsov, Michelangelo Foti, and Markus Thali. Mapping of tetraspaninenriched microdomains that can function as gateways for HIV-1. Journal of Cell Biology, 173(5):795–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/173/5/795.

Naylor:2005:AIM

[NLC^{+}05] Matthew J. Naylor, Na Li, Julia Cheung, Emma T. Lowe, Elise Lambert, Rebecca Marlow, Pengbo Wang, Franziska Schatzmann, Timothy Wintermantel, Günther Schüetz, Alan R.

Clarke, Ulrich Mueller, Nancy E. Hynes, and Charles H. Streuli. Ablation of β1 integrin in mammary epithelium reveals a key role for integrin in glandular morphogenesis and differentiation. Journal of Cell Biology, 171(4):717–??, November 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/171/4/717.

Nakanishi:2009:HHM

[NLG^{+}09] Shima Nakanishi, Jung Shin Lee, Kathryn E. Gardner, Jennifer M. Gardner, Yoh hei Takahashi, Mahesh B. Chandrasekharan, Zu-Wen Sun, Mary Ann Osley, Brian D. Strahl, Sue L. Jaspersen, and Ali Shilatifard. Histone H2BK123 monoubiquitination is the critical determinant for H3K4 and H3K79 trimethylation by COMPASS and Dot1. Journal of Cell Biology, 186(3):371–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/186/3/371.

Ng:2006:SRM

[NLL^{+}06] Dominic Chi Hiung Ng, Bao Hong Lin, Cheh Peng Lim, Guochang Huang, Tong Zhang, Valeria Poli, and Xinmin Cao. Stat3 regulates microtubules by antagonizing the depolymerization activity of stathmin. Journal of Cell Biology, 172(2): 245–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/2/245.

Narayanan:2005:IGD

[NLS^{+}05] Radhakrishnan Narayanan, Marilyn Leonard, Byeong Doo Song, Sandra L. Schmid, and Mani Ramaswami. An internal GAP domain negatively regulates presynaptic dynamin in vivo. Journal of Cell Biology, 169(1):117–??, April 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/169/1/117.

Nichols:2007:DLE

[NMO^{+}07] James T. Nichols, Alison Miyamoto, Samantha L. Olsen, Brendan D’Souza, Christine Yao, and Gerry Weinmaster. DSL ligand endocytosis physically dissociates Notch1 heterodimers before activating proteolysis can occur. Journal of Cell Biology, 176(4):445–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/4/445.

Naik:2007:URT

[NMV^{+}07] Edwina Naik, Ewa M. Michalak, Andreas Villunger, Jerry M. Adams, and Andreas Strasser. Ultraviolet radiation triggers apoptosis of fibroblasts and skin keratinocytes mainly via the BH3-only protein Noxa. Journal of Cell Biology, 176(4): 415–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/4/415.

Nguyen:2009:TCR

[NMZ^{+}09] Chau H. Nguyen, Hong Ming, Peishen Zhao, Lynne Hugendubler, Robert Gros, Scot R. Kimball, and Peter Chidiac. Translational control by RGS2. Journal of Cell Biology, 186(5): 755–??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/186/5/755.

Nejsum:2007:MMD

[NN07a] Lene N. Nejsum and W. James Nelson. A molecular mechanism directly linking E-cadherin adhesion to initiation of epithelial cell surface polarity. Journal of Cell Biology, 178(2):323–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
178/2/323.

Neumann:2007:NSC

[NN07b] Frank R. Neumann and Paul Nurse. Nuclear size control in fission yeast. Journal of Cell Biology, 179(4):593–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/4/
593.

Nakazawa:2008:DES

[NNK^{+}08] Norihiko Nakazawa, Takahiro Nakamura, Aya Kokubu, Masahiro Ebe, Koji Nagao, and Mitsuhiro Yanagida. Dissection of the essential steps for condensin accumulation at kinetochores and rDNAs during fission yeast mitosis. Journal of Cell Biology, 180(6):1115–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/6/1115.

Nolz:2008:WCR

[NNS^{+}08] Jeffrey C. Nolz, Lucas P. Nacusi, Colin M. Segovis, Ricardo B. Medeiros, Jason S. Mitchell, Yoji Shimizu, and Daniel D. Billadeau. The WAVE2 complex regulates T cell receptor signaling to integrins via Abl- and CrkL–C3G-mediated activation of Rap1. Journal of Cell Biology, 182(6):1231–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/182/6/
1231.

Nakamura:2007:SBF

[NOH^{+}07] Fumihiko Nakamura, Teresia M. Osborn, Christopher A. Hartemink, John H. Hartwig, and Thomas P. Stossel. Structural basis of filamin A functions. Journal of Cell Biology, 179(5):1011–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/5/1011.

North:2006:SBB

[Nor06] Alison J. North. Seeing is believing? A beginners’ guide to practical pitfalls in image acquisition. Journal of Cell Biology, 172 (1):9–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/1/9.

Nakano:2005:MEL

[NPB^{+}05] Ichiro Nakano, Andres A. Paucar, Ruchi Bajpai, Joseph D. Dougherty, Amani Zewail, Theresa K. Kelly, Kevin J. Kim, Jing Ou, Matthias Groszer, Tetsuya Imura, William A. Freije, Stanley F. Nelson, Michael V. Sofroniew, Hong Wu, Xin Liu, Alexey V. Terskikh, Daniel H. Geschwind, and Harley I. Kornblum. Maternal embryonic leucine zipper kinase (MELK) regulates multipotent neural progenitor proliferation. Journal of Cell Biology, 170(3):413–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/170/3/413.

Nickell:2007:TDA

[NPBP07] Stephan Nickell, Paul S.-H. Park, Wolfgang Baumeister, and Krzysztof Palczewski. Three-dimensional architecture of murine rod outer segments determined by cryoelectron tomography. Journal of Cell Biology, 177(5):917–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/177/5/917.

Nakao:2008:CDP

[NPHT08] Shinsuke Nakao, Anna Platek, Shinji Hirano, and Masatoshi Takeichi. Contact-dependent promotion of cell migration by the OL–protocadherin–Nap1 interaction. Journal of Cell Biology, 182(2):395–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/2/395.

Nagata:2006:EMS

[NPMZ06] Yosuke Nagata, Terence A. Partridge, Ryoichi Matsuda, and Peter S. Zammit. Entry of muscle satellite cells into the cell cycle requires sphingolipid signaling. Journal of Cell Biology, 174(2):245–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/2/245.

Niimi:2007:NSN

[NPV^{+}07] Hideki Niimi, Katerina Pardali, Michael Vanlandewijck, Carl-Henrik Heldin, and Aristidis Moustakas. Notch signaling is necessary for epithelial growth arrest by TGF-β. Journal of Cell Biology, 176(5):695–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/176/5/695.

Nezi:2006:AMC

[NRD^{+}06] Luigi Nezi, Giulia Rancati, Anna De Antoni, Sebastiano Pasqualato, Simonetta Piatti, and Andrea Musacchio. Accumulation of Mad2–Cdc20 complex during spindle checkpoint activation requires binding of open and closed conformers of Mad2 in Saccharomyces cerevisiae. Journal of Cell Biology, 174(1):39–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/1/39.

Nowak:2009:RSM

[NRJ^{+}09] Kristen J. Nowak, Gianina Ravenscroft, Connie Jackaman, Aleksandra Filipovska, Stefan M. Davies, Esther M. Lim, Sarah E. Squire, Allyson C. Potter, Elizabeth Baker, Sophie Clément, Caroline A. Sewry, Victoria Fabian, Kelly Crawford, James L. Lessard, Lisa M. Griffiths, John M. Papadimitriou, Yun Shen, Grant Morahan, Anthony J. Bakker, Kay E. Davies, and Nigel G. Laing. Rescue of skeletal muscle α-actin–null mice by cardiac (fetal) α-actin. Journal of Cell Biology, 185(5):903– ??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
185/5/903.

Niemann:2005:GID

[NRL^{+}05] Axel Niemann, Marcel Ruegg, Veronica La Padula, Angelo Schenone, and Ueli Suter. Ganglioside-induced differentiation associated protein 1 is a regulator of the mitochondrial network. Journal of Cell Biology, 170(7):1067–??, September 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/170/7/1067.

Nyfeler:2008:IEI

[NRW^{+}08] Beat Nyfeler, Veronika Reiterer, Markus W. Wendeler, Eduard Stefan, Bin Zhang, Stephen W. Michnick, and Hans-Peter Hauri. Identification of ERGIC-53 as an intracellular transport receptor of α_{1}-antitrypsin. Journal of Cell Biology, 180 (4):705–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/4/705.

Nicholson:2006:SIG

[NS06] Thomas B. Nicholson and Clifford P. Stanners. Specific inhibition of GPI-anchored protein function by homing and selfassociation of specific GPI anchors. Journal of Cell Biology, 175 (4):647–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/4/647.

Nicholson:2007:INF

[NS07] Thomas B. Nicholson and Clifford P. Stanners. Identification of a novel functional specificity signal within the GPI anchor signal sequence of carcinoembryonic antigen. Journal of Cell Biology, 177(2):211–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/2/211.

Nakanishi:2005:ERS

[NSM05] Keiko Nakanishi, Tatsuhiko Sudo, and Nobuhiro Morishima. Endoplasmic reticulum stress signaling transmitted by ATF6 mediates apoptosis during muscle development. Journal of Cell Biology, 169(4):555–??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/4/555.

Niault:2009:AIT

[NSM^{+}09] Théodora Niault, Izabela Sobczak, Katrin Meissl, Gregory Weitsman, Daniela Piazzolla, Gabriele Maurer, Florian Kern, Karin Ehrenreiter, Matthias Hamerl, Ismail Moarefi, Thomas Leung, Oliviero Carugo, Tony Ng, and Manuela Baccarini. From autoinhibition to inhibition in trans: the Raf-1 regulatory domain inhibits Rok-α kinase activity. Journal of Cell Biology, 187(3):335–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/3/335.

Nishihama:2009:RII

[NSO^{+}09] Ryuichi Nishihama, Jennifer H. Schreiter, Masayuki Onishi, Elizabeth A. Vallen, Julia Hanna, Katarina Moravcevic, Margaret F. Lippincott, Haesun Han, Mark A. Lemmon, John R. Pringle, and Erfei Bi. Role of Inn1 and its interactions with Hof1 and Cyk3 in promoting cleavage furrow and septum formation in S. cerevisiae. Journal of Cell Biology, 185(6):995–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/185/6/
995.

Nezis:2008:RPD

[NSS^{+}08] Ioannis P. Nezis, Anne Simonsen, Antonia P. Sagona, Kim Finley, Sébastien Gaumer, Didier Contamine, Tor Erik Rusten, Harald Stenmark, and Andreas Brech. Ref(2)P, the Drosophila melanogaster homologue of mammalian p62, is required for the formation of protein aggregates in adult brain. Journal of Cell Biology, 180(6):1065–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/6/1065.

Nair:2006:ORG

[NTB^{+}06] Mahalakshmi Nair, Andy Teng, Virginia Bilanchone, Anshu Agrawal, Baoan Li, and Xing Dai. Ovol1 regulates the growth arrest of embryonic epidermal progenitor cells and represses cmyc transcription. Journal of Cell Biology, 173(2):253–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/173/2/
253.

Nagaya:2008:RMG

[NTHN^{+}08] Hisao Nagaya, Taku Tamura, Arisa Higa-Nishiyama, Koji Ohashi, Mayumi Takeuchi, Hitoshi Hashimoto, Kiyotaka Hatsuzawa, Masataka Kinjo, Tatsuya Okada, and Ikuo Wada. Regulated motion of glycoproteins revealed by direct visualization of a single cargo in the endoplasmic reticulum. Journal of Cell Biology, 180(1):129–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/1/129.

Nakamichi:2005:KOP

[NTS^{+}05] Ikuo Nakamichi, Diana M. Toivola, Pavel Strnad, Sara A. Michie, Robert G. Oshima, Hélène Baribault, and M. Bishr Omary. Keratin 8 overexpression promotes mouse Mallory body formation. Journal of Cell Biology, 171(6):931–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/171/6/
931.

Narendra:2008:PRS

[NTSY08] Derek Narendra, Atsushi Tanaka, Der-Fen Suen, and Richard J. Youle. Parkin is recruited selectively to impaired mitochondria and promotes their autophagy. Journal of Cell Biology, 183(5): 795–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/5/795.

Nishita:2005:STR

[NTY^{+}05] Michiru Nishita, Chinatsu Tomizawa, Masahiro Yamamoto, Yuji Horita, Kazumasa Ohashi, and Kensaku Mizuno. Spatial and temporal regulation of cofilin activity by LIM kinase and Slingshot is critical for directional cell migration. Journal of Cell Biology, 171(2):349–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/171/2/349.

Nabeshima:2005:CCL

[NVC05] Kentaro Nabeshima, Anne M. Villeneuve, and Monica P. Colaiácovo. Crossing over is coupled to late meiotic prophase bivalent differentiation through asymmetric disassembly of the SC.

Journal of Cell Biology, 168(5):683–??, February 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/168/5/683.

Nguyen:2007:DMB

[NVEF07] Hanh T. Nguyen, Francesca Voza, Nader Ezzeddine, and Manfred Frasch. Drosophila mind bomb2 is required for maintaining muscle integrity and survival. Journal of Cell Biology, 179 (2):219–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/2/219.

Nishimune:2008:LPP

[NVJ^{+}08] Hiroshi Nishimune, Gregorio Valdez, George Jarad, Casey L. Moulson, Ulrich Müller, Jeffrey H. Miner, and Joshua R. Sanes. Laminins promote postsynaptic maturation by an autocrine mechanism at the neuromuscular junction. Journal of Cell Biology, 182(6):1201–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/6/1201.

Nayal:2006:PPS

[NWB^{+}06] Anjana Nayal, Donna J. Webb, Claire M. Brown, Erik M. Schaefer, Miguel Vicente-Manzanares, and Alan Rick Horwitz. Paxillin phosphorylation at Ser273 localizes a GIT1–PIX– PAK complex and regulates adhesion and protrusion dynamics. Journal of Cell Biology, 173(4):587–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/173/4/587.

Nickerson:2006:CP

[NWO06a] Daniel P. Nickerson, Matthew West, and Greg Odorizzi. Correction 175, p. 715. Journal of Cell Biology, 175(6):1043–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
175/6/1043.

Nickerson:2006:DCV

[NWO06b] Daniel P. Nickerson, Matthew West, and Greg Odorizzi. Did2 coordinates Vps4-mediated dissociation of ESCRT–III from endosomes. Journal of Cell Biology, 175(5):715–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/175/5/
715.

Novak:2008:CSD

[NWR^{+}08] Ivana Novak, Hong Wang, Ekaterina Revenkova, Rolf Jessberger, Harry Scherthan, and Christer Höög. Cohesin Smc1β determines meiotic chromatin axis loop organization. Journal of Cell Biology, 180(1):83–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/1/83.

Nyberg:2005:ASM

[Nyb05a] Kara A. Nyberg. Axon size matters. Journal of Cell Biology, 170(2):??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/2/164.2.

Nyberg:2005:BS

[Nyb05b] Kara A. Nyberg. Breaking down symmetry. Journal of Cell Biology, 169(6):??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/6/836.1.

Nyberg:2005:DSH

[Nyb05c] Kara A. Nyberg. The death stops here. Journal of Cell Biology, 169(6):??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/6/837.1.

Nyberg:2005:FF

[Nyb05d] Kara A. Nyberg. Fiber flex. Journal of Cell Biology, 170(2):??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
170/2/165.2.

Nyberg:2005:GUG

[Nyb05e] Kara A. Nyberg. Gradient upon gradient. Journal of Cell Biology, 170(7):??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/7/1016.2.

Nyberg:2005:MMA

[Nyb05f] Kara A. Nyberg. Making mature attachments. Journal of Cell Biology, 169(6):??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/6/837.2.

Nyberg:2005:NBB

[Nyb05g] Kara A. Nyberg. New bend on budding. Journal of Cell Biology, 170(7):??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/7/1016.1.

Nyberg:2005:PPC

[Nyb05h] Kara A. Nyberg. Putting the pinch on cell division. Journal of Cell Biology, 169(6):??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/6/836.3.

Nyberg:2005:SC

[Nyb05i] Kara A. Nyberg. Selective clusters. Journal of Cell Biology, 170(2):??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/2/164.1.

Nyberg:2005:SVF

[Nyb05j] Kara A. Nyberg. Sensing voltage for function. Journal of Cell Biology, 170(2):??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/2/165.1.

Nyberg:2005:SD

[Nyb05k] Kara A. Nyberg. Smell’s different. Journal of Cell Biology, 170(2):??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/2/165.3.

Nyberg:2005:TT

[Nyb05l] Kara A. Nyberg. Tilted tubules. Journal of Cell Biology, 170 (7):??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/7/1017.3.

Nyberg:2005:TTC

[Nyb05m] Kara A. Nyberg. Totally tubular chitin. Journal of Cell Biology, 170(7):??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/7/1017.1.

Nyberg:2005:TF

[Nyb05n] Kara A. Nyberg. Turning off flow. Journal of Cell Biology, 170 (7):??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/7/1017.2.

Nyberg:2005:WB

[Nyb05o] Kara A. Nyberg. Weaker is better. Journal of Cell Biology, 169(6):??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/6/836.2.

Nishida:2005:C

[NYI^{+}05a] Keigo Nishida, Satoru Yamasaki, Yukitaka Ito, Koki Kabu, Kotaro Hattori, Tohru Tezuka, Hirofumi Nishizumi, Daisuke Kitamura, Ryo Goitsuka, Raif S. Geha, Tadashi Yamamoto, Takeshi Yagi, and Toshio Hirano. Correction. Journal of Cell Biology, 171(1):177–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/1/177.

Nishida:2005:FMM

[NYI^{+}05b] Keigo Nishida, Satoru Yamasaki, Yukitaka Ito, Koki Kabu, Kotaro Hattori, Tohru Tezuka, Hirofumi Nishizumi, Daisuke Kitamura, Ryo Goitsuka, Raif S. Geha, Tadashi Yamamoto, Takeshi Yagi, and Toshio Hirano. FcϵRI-mediated mast cell degranulation requires calcium-independent microtubule-dependent translocation of granules to the plasma membrane. Journal of Cell Biology, 170(1):115–??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/170/1/115.

Nakajima:2008:MPK

[NYM^{+}08] Hiroyuki Nakajima, Shigenobu Yonemura, Masayuki Murata, Nobuhiro Nakamura, Helen Piwnica-Worms, and Eisuke Nishida. Myt1 protein kinase is essential for Golgi and ER assembly during mitotic exit. Journal of Cell Biology, 181 (1):89–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/1/89.

Nishita:2006:FFM

[NYN^{+}06] Michiru Nishita, Sa Kan Yoo, Akira Nomachi, Shuichi Kani, Nagako Sougawa, Yasutaka Ohta, Shinji Takada, Akira Kikuchi, and Yasuhiro Minami. Filopodia formation mediated by receptor tyrosine kinase Ror2 is required for Wnt5a-induced cell migration. Journal of Cell Biology, 175(4):555–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/175/4/
555.

Nodari:2007:IAR

[NZQ^{+}07] Alessandro Nodari, Desirée Zambroni, Angelo Quattrini, Felipe A. Court, Alessandra D’Urso, Alessandra Recchia, Victor L. J. Tybulewicz, Lawrence Wrabetz, and M. Laura Feltri. β1 integrin activates Rac1 in Schwann cells to generate radial lamellae during axonal sorting and myelination. Journal of Cell Biology, 177(6):1063–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/6/1063.

Olson:2006:INC

[OBY^{+}06] Sara K. Olson, Joseph R. Bishop, John R. Yates, Karen Oegema, and Jeffrey D. Esko. Identification of novel chondroitin proteoglycans in Caenorhabditis elegans: embryonic cell division depends on CPG-1 and CPG-2. Journal of Cell Biology, 173(6):985–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/6/985.

Okreglak:2007:CRF

[OD07] Voytek Okreglak and David G. Drubin. Cofilin recruitment and function during actin-mediated endocytosis dictated by actin nucleotide state. Journal of Cell Biology, 178(7):1251– ??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/7/1251.

Odell:2008:ABM

[OF08] Garrett M. Odell and Victoria E. Foe. An agent-based model contrasts opposite effects of dynamic and stable microtubules on cleavage furrow positioning. Journal of Cell Biology, 183(3): 471–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/3/471.

Olsen:2007:RDA

[OFfL^{+}07] Olav Olsen, Lars Funke, Jia fu Long, Masaki Fukata, Toshinari Kazuta, Jonathan C. Trinidad, Kimberly A. Moore, Hidemi Misawa, Paul A. Welling, Alma L. Burlingame, Mingjie Zhang, and David S. Bredt. Renal defects associated with improper polarization of the CRB and DLG polarity complexes in MALS-3 knockout mice. Journal of Cell Biology, 179(1):151–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/1/
151.

Ooashi:2005:CAM

[OFY^{+}05] Noriko Ooashi, Akira Futatsugi, Fumie Yoshihara, Katsuhiko Mikoshiba, and Hiroyuki Kamiguchi. Cell adhesion molecules regulate Ca^{2+}-mediated steering of growth cones via cyclic AMP and ryanodine receptor type 3. Journal of Cell Biology, 170(7):1159–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/7/1159.

ODonnell:2006:IPM

[OHH^{+}06] Rebecca A. O’Donnell, Fiona Hackett, Steven A. Howell, Moritz Treeck, Nicole Struck, Zita Krnajski, Chrislaine Withers-Martinez, Tim W. Gilberger, and Michael J. Blackman. Intramembrane proteolysis mediates shedding of a key adhesin during erythrocyte invasion by the malaria parasite. Journal of Cell Biology, 174(7):1023–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/174/7/1023.

Oda:2007:TDS

[OHK07] Toshiyuki Oda, Nobutaka Hirokawa, and Masahide Kikkawa. Three-dimensional structures of the flagellar dynein–microtubule complex by cryoelectron microscopy. Journal of Cell Biology, 177(2):243–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/2/243.

Osman:2009:GIP

[OHP^{+}09] Christof Osman, Mathias Haag, Christoph Potting, Jonathan Rodenfels, Phat Vinh Dip, Felix T. Wieland, Britta Brügger, Benedikt Westermann, and Thomas Langer. The genetic interactome of prohibitins: coordinated control of cardiolipin and phosphatidylethanolamine by conserved regulators in mitochondria. Journal of Cell Biology, 184(4):583–??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/184/4/
583.

Otani:2006:CGT

[OIAT06] Tetsuhisa Otani, Tetsuo Ichii, Shinya Aono, and Masatoshi Takeichi. Cdc42 GEF Tuba regulates the junctional configuration of simple epithelial cells. Journal of Cell Biology, 175 (1):135–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/1/135.

Ohara-Imaizumi:2007:IAR

[OIFN^{+}07] Mica Ohara-Imaizumi, Tomonori Fujiwara, Yoko Nakamichi, Tadashi Okamura, Yoshihiro Akimoto, Junko Kawai, Satsuki Matsushima, Hayato Kawakami, Takashi Watanabe, Kimio Akagawa, and Shinya Nagamatsu. Imaging analysis reveals mechanistic differences between first- and second-phase insulin exocytosis. Journal of Cell Biology, 177(4):695–??, May 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/177/4/695.

Oikawa:2008:SST

[OIT08] Tsukasa Oikawa, Toshiki Itoh, and Tadaomi Takenawa. Sequential signals toward podosome formation in NIH-src cells. Journal of Cell Biology, 182(1):157–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/182/1/157.

Oinuma:2006:SPB

[OKN06] Izumi Oinuma, Hironori Katoh, and Manabu Negishi. Semaphorin 4D/Plexin–B1-mediated R–Ras GAP activity inhibits cell migration by regulating β_{1} integrin activity. Journal of Cell Biology, 173(4):601–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/4/601.

Olzmann:2007:PMK

[OLC^{+}07] James A. Olzmann, Lian Li, Maksim V. Chudaev, Jue Chen, Francisco A. Perez, Richard D. Palmiter, and Lih-Shen Chin. Parkin-mediated K63-linked polyubiquitination targets misfolded DJ-1 to aggresomes via binding to HDAC6. Journal of Cell Biology, 178(6):1025–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/178/6/1025.

OConnell:2008:SAC

[OLH^{+}08] Christopher B. O’Connell, Jadranka Lončarek, Polla Hergert, Antonis Kourtidis, Douglas S. Conklin, and Alexey Khodjakov. The spindle assembly checkpoint is satisfied in the absence of interkinetochore tension during mitosis with unreplicated genomes. Journal of Cell Biology, 183(1):29–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/183/1/29.

OConnell:2009:RCC

[OLKK09] Christopher B. O’Connell, Jadranka Lončarek, Petr Kaláb, and Alexey Khodjakov. Relative contributions of chromatin and kinetochores to mitotic spindle assembly. Journal of Cell Biology, 187(1):43–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/1/43.

Okada:2005:MNM

[OLLG05] Tomoyo Okada, Miguel Lopez-Lago, and Filippo G. Giancotti. Merlin/ NF-2 mediates contact inhibition of growth by suppressing recruitment of Rac to the plasma membrane. Journal of Cell Biology, 171(2):361–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/171/2/361.

OReilly:2008:ICP

[OLS08] Alana M. O’Reilly, Hsiu-Hsiang Lee, and Michael A. Simon. Integrins control the positioning and proliferation of follicle stem cells in the Drosophila ovary. Journal of Cell Biology, 182 (4):801–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/4/801.

Olsen:2005:NRR

[OMF^{+}05] Olav Olsen, Kimberly A. Moore, Masaki Fukata, Toshinari Kazuta, Jonathan C. Trinidad, Fred W. Kauer, Michel Streuli, Hidemi Misawa, Alma L. Burlingame, Roger A. Nicoll, and David S. Bredt. Neurotransmitter release regulated by a MALSliprin-α presynaptic complex. Journal of Cell Biology, 170(7): 1127–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/7/1127.

Osmanagic-Myers:2006:PCK

[OMGW^{+}06] Selma Osmanagic-Myers, Martin Gregor, Gernot Walko, Gerald Burgstaller, Siegfried Reipert, and Gerhard Wiche. Plectincontrolled keratin cytoarchitecture affects MAP kinases involved in cellular stress response and migration. Journal of Cell Biology, 174(4):557–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/174/4/557.

Ordonez-Moran:2008:RRP

[OMLP^{+}08] Paloma Ordóñez-Morán, María Jesús Larriba, Héctor G. Pálmer, Ruth A. Valero, Antonio Barbáchano, Mireia Duñach, Antonio García de Herreros, Carlos Villalobos, María Teresa Berciano, Miguel Lafarga, and Alberto Muñoz. RhoA–ROCK and p38MAPK–MSK1 mediate vitamin D effects on gene expression, phenotype, and Wnt pathway in colon cancer cells. Journal of Cell Biology, 183(4):697–??, November 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/183/4/697.

Oyama:2007:CSP

[ONW^{+}07] Tomomi Oyama, Toshio Nagai, Hiroshi Wada, Atsuhiko Thomas Naito, Katsuhisa Matsuura, Koji Iwanaga, Toshinao Takahashi, Motohiro Goto, Yoko Mikami, Noritaka Yasuda, Hiroshi Akazawa, Akiyoshi Uezumi, Shin’ichi Takeda, and Issei Komuro. Cardiac side population cells have a potential to migrate and differentiate into cardiomyocytes in vitro and in vivo. Journal of Cell Biology, 176(3):329–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/176/3/329.

Oishi:2007:RNM

[OOS07] Kumiko Oishi, Hideyuki Okano, and Hitoshi Sawa. RMD-1, a novel microtubule-associated protein, functions in chromosome segregation in Caenorhabditis elegans. Journal of Cell Biology, 179(6):1149–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/6/1149.

Oishi:2009:RNM

[OOS09] Kumiko Oishi, Hideyuki Okano, and Hitoshi Sawa. RMD-1, a novel microtubule-associated protein, functions in chromosome segregation in Caenorhabditis elegans. Journal of Cell Biology, 186(4):629–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/186/4/629.

Oda:2006:DDR

[OOY^{+}06] Yukako Oda, Tetsuya Okada, Hiderou Yoshida, Randal J. Kaufman, Kazuhiro Nagata, and Kazutoshi Mori. Derlin-2 and Derlin-3 are regulated by the mammalian unfolded protein response and are required for ER-associated degradation. Journal of Cell Biology, 172(3):383–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/172/3/383.

Owen:2007:RLP

[OPT^{+}07] Katherine A. Owen, Fiona J. Pixley, Keena S. Thomas, Miguel Vicente-Manzanares, Brianne J. Ray, Alan F. Horwitz, J. Thomas Parsons, Hilary E. Beggs, E. Richard Stanley, and Amy H. Bouton. Regulation of lamellipodial persistence, adhesion turnover, and motility in macrophages by focal adhesion kinase. Journal of Cell Biology, 179(6):1275–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/6/
1275.

Osmundson:2008:HEL

[ORM^{+}08] Evan C. Osmundson, Dipankar Ray, Finola E. Moore, Qingshen Gao, Gerald H. Thomsen, and Hiroaki Kiyokawa. The HECT E3 ligase Smurf2 is required for Mad2-dependent spindle assembly checkpoint. Journal of Cell Biology, 183(2):267–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/183/2/
267.

Orr:2005:SEM

[OSB^{+}05] A. Wayne Orr, John M. Sanders, Melissa Bevard, Elizabeth Coleman, Ian J. Sarembock, and Martin Alexander Schwartz. The subendothelial extracellular matrix modulates NF-κB activation by flow. Journal of Cell Biology, 169(1):191–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/169/1/
191.

Shan:2007:CCG

[oSCW07] Shu ou Shan, Sowmya Chandrasekar, and Peter Walter. Conformational changes in the GTPase modules of the signal reception particle and its receptor drive initiation of protein translocation. Journal of Cell Biology, 178(4):611–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/178/4/611.

Otsuki:2007:FIB

[OSI^{+}07] Makoto Otsuki, Masayuki Seki, Eri Inoue, Akari Yoshimura, Genta Kato, Saki Yamanouchi, Yoh ichi Kawabe, Shusuke Tada, Akira Shinohara, Jun ichiro Komura, Tetsuya Ono, Shunichi Takeda, Yutaka Ishii, and Takemi Enomoto. Functional interactions between BLM and XRCC3 in the cell. Journal of Cell Biology, 179(1):53–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/1/53.

Onischenko:2009:RNI

[OSM^{+}09] Evgeny Onischenko, Leslie H. Stanton, Alexis S. Madrid, Thomas Kieselbach, and Karsten Weis. Role of the Ndc1 interaction network in yeast nuclear pore complex assembly and maintenance. Journal of Cell Biology, 185(3):475–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/185/3/475.

Orr:2007:MSP

[OSS^{+}07] A. Wayne Orr, Rebecca Stockton, Michael B. Simmers, John M. Sanders, Ian J. Sarembock, Brett R. Blackman, and Martin Alexander Schwartz. Matrix-specific p21-activated kinase activation regulates vascular permeability in atherogenesis.

Journal of Cell Biology, 176(5):719–??, February 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/176/5/719.

Otera:2007:NIP

[OTH^{+}07] Hidenori Otera, Yohsuke Taira, Chika Horie, Yurina Suzuki, Hiroyuki Suzuki, Kiyoko Setoguchi, Hiroki Kato, Toshihiko Oka, and Katsuyoshi Mihara. A novel insertion pathway of mitochondrial outer membrane proteins with multiple transmembrane segments. Journal of Cell Biology, 179(7):1355–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/7/
1355.

Osaka:2007:ADR

[OTM^{+}07] Nao Osaka, Takumi Takahashi, Shiori Murakami, Atsushi Matsuzawa, Takuya Noguchi, Takeshi Fujiwara, Hiroyuki Aburatani, Keiji Moriyama, Kohsuke Takeda, and Hidenori Ichijo. ASK1-dependent recruitment and activation of macrophages induce hair growth in skin wounds. Journal of Cell Biology, 176 (7):903–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/7/903.

Ou:2009:MSC

[OV09] Guangshuo Ou and Ronald D. Vale. Molecular signatures of cell migration in C. elegans Q neuroblasts. Journal of Cell Biology, 185(1):77–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/1/77.

Ohyama:2007:HIP

[OVL^{+}07] Tomoko Ohyama, Patrik Verstreken, Cindy V. Ly, Tanja Rosenmund, Akhila Rajan, An-Chi Tien, Claire Haueter, Karen L. Schulze, and Hugo J. Bellen. Huntingtin-interacting protein 14, a palmitoyl transferase required for exocytosis and targeting of CSP to synaptic vesicles. Journal of Cell Biology, 179(7): 1481–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/7/1481.

OBrien:2006:TRP

[OW06] Lori L. O’Brien and Christiane Wiese. TPX2 is required for postmitotic nuclear assembly in cell-free Xenopus laevis egg extracts. Journal of Cell Biology, 173(5):685–??, June 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/173/5/685.

Osterweil:2005:RMV

[OWM05] Emily Osterweil, David G. Wells, and Mark S. Mooseker. A role for myosin VI in postsynaptic structure and glutamate receptor endocytosis. Journal of Cell Biology, 168(2):329–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/168/2/
329.

Oceguera-Yanez:2005:EMR

[OYKY^{+}05] Fabian Oceguera-Yanez, Kazuhiro Kimura, Shingo Yasuda, Chiharu Higashida, Toshio Kitamura, Yasushi Hiraoka, Tokuko Haraguchi, and Shuh Narumiya. Ect2 and MgcRacGAP regulate the activation and function of Cdc42 in mitosis. Journal of Cell Biology, 168(2):221–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/168/2/221.

Oser:2009:CRC

[OYM^{+}09] Matthew Oser, Hideki Yamaguchi, Christopher C. Mader, J. J. Bravo-Cordero, Marianela Arias, Xiaoming Chen, Vera Des-Marais, Jacco van Rheenen, Anthony J. Koleske, and John Condeelis. Cortactin regulates cofilin and N–WASp activities to control the stages of invadopodium assembly and maturation. Journal of Cell Biology, 186(4):571–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/186/4/571.

Olguin:2007:RIB

[OYTO07] Hugo C. Olguin, Zhihong Yang, Stephen J. Tapscott, and Bradley B. Olwin. Reciprocal inhibition between Pax7 and muscle regulatory factors modulates myogenic cell fate determination. Journal of Cell Biology, 177(5):769–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/177/5/769.

Pazour:2005:PAE

[PALW05] Gregory J. Pazour, Nathan Agrin, John Leszyk, and George B. Witman. Proteomic analysis of a eukaryotic cilium. Journal of Cell Biology, 170(1):103–??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/170/1/103.

Paronetto:2007:RBP

[PAM^{+}07] Maria Paola Paronetto, Tilman Achsel, Autumn Massiello, Charles E. Chalfant, and Claudio Sette. The RNA-binding protein Sam68 modulates the alternative splicing of Bcl-x. Journal of Cell Biology, 176(7):929–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/176/7/929.

Pellinen:2006:SGR

[PAV^{+}06] Teijo Pellinen, Antti Arjonen, Karoliina Vuoriluoto, Katja Kallio, Jack A. M. Fransen, and Johanna Ivaska. Small GT-Pase Rab21 regulates cell adhesion and controls endosomal traffic of β1-integrins. Journal of Cell Biology, 173(5):767–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/173/5/
767.

Patel:2005:DRH

[PB05] Hitesh Patel and Diane L. Barber. A developmentally regulated Na–H exchanger in Dictyostelium discoideum is necessary for cell polarity during chemotaxis. Journal of Cell Biology, 169(2):321–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/2/321.

Parkinson:2008:CJN

[PBAF^{+}08] David B. Parkinson, Ambily Bhaskaran, Peter Arthur-Farraj, Luke A. Noon, Ashwin Woodhoo, Alison C. Lloyd, M. Laura Feltri, Lawrence Wrabetz, Axel Behrens, Rhona Mirsky, and Kristján R. Jessen. c-jun is a negative regulator of myelination. Journal of Cell Biology, 181(4):625–??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/181/4/625.

Pereira:2009:ILK

[PBB^{+}09] Jorge A. Pereira, Yves Benninger, Reto Baumann, Ana Filipa Gonçalves, Murat Özçelik, Tina Thurnherr, Nicolas Tricaud, Dies Meijer, Reinhard Fässler, Ueli Suter, and João B. Relvas. Integrin-linked kinase is required for radial sorting of axons and Schwann cell remyelination in the peripheral nervous system. Journal of Cell Biology, 185(1):147–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/185/1/147.

Pisconti:2006:FIN

[PBD^{+}06] Addolorata Pisconti, Silvia Brunelli, Monica Di Padova, Clara De Palma, Daniela Deponti, Silvia Baesso, Vittorio Sartorelli, Giulio Cossu, and Emilio Clementi. Follistatin induction by nitric oxide through cyclic GMP: a tightly regulated signaling pathway that controls myoblast fusion. Journal of Cell Biology, 172(2):233–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/2/233.

Paganetti:2005:SSI

[PCG^{+}05] Paolo Paganetti, Verena Calanca, Carmela Galli, Muriel Stefani, and Maurizio Molinari. β-site specific intrabodies to decrease and prevent generation of Alzheimer’s Aβ peptide. Journal of Cell Biology, 168(6):863–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/168/6/863.

Patel:2009:DDH

[PCMSB09] Prajal H. Patel, Mauro Costa-Mattioli, Karen L. Schulze, and Hugo J. Bellen. The Drosophila deoxyhypusine hydroxylase homologue nero and its target eIF5A are required for cell growth and the regulation of autophagy. Journal of Cell Biology, 185 (7):1181–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/7/1181.

Pashkova:2005:MVA

[PCN^{+}05] Natasha Pashkova, Natalie L. Catlett, Jennifer L. Novak, Guanming Wu, Renne Lu, Robert E. Cohen, and Lois S. Weisman. Myosin V attachment to cargo requires the tight association of two functional subdomains. Journal of Cell Biology, 168(3):

359–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/3/359.

Poliakov:2008:REA

[PCPW08] Alexei Poliakov, Maria L. Cotrina, Andrea Pasini, and David G. Wilkinson. Regulation of EphB2 activation and cell repulsion by feedback control of the MAPK pathway. Journal of Cell Biology, 183(5):933–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/5/933.

Padilla:2007:ECE

[PCR^{+}07] Benjamin E. Padilla, Graeme S. Cottrell, Dirk Roosterman, Stella Pikios, Laurent Muller, Martin Steinhoff, and Nigel W. Bunnett. Endothelin-converting enzyme-1 regulates endosomal sorting of calcitonin receptor-like receptor and β-arrestins. Journal of Cell Biology, 179(5):981–??, December 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/179/5/981.

Paschen:2008:CCL

[PCV^{+}08] Stefan A. Paschen, Jan G. Christian, Juliane Vier, Franziska Schmidt, Axel Walch, David M. Ojcius, and Georg Häcker. Cytopathicity of Chlamydia is largely reproduced by expression of a single chlamydial protease. Journal of Cell Biology, 182 (1):117–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/1/117.

Primo:2007:ERP

[PdBR^{+}07] Luca Primo, Laura di Blasio, Cristina Roca, Sara Droetto, Roberto Piva, Brian Schaffhausen, and Federico Bussolino. Essential role of PDK1 in regulating endothelial cell migration. Journal of Cell Biology, 176(7):1035–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/176/7/1035.

Panagopoulou:2008:DMT

[PDM^{+}08] Panagiota Panagopoulou, Constantinos H. Davos, Derek J. Milner, Emily Varela, JoAnn Cameron, Douglas L. Mann, and Yassemi Capetanaki. Desmin mediates TNF-α-induced aggregate formation and intercalated disk reorganization in heart failure. Journal of Cell Biology, 181(5):761–??, June 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/181/5/761.

Pederson:2008:FNA

[Ped08] Thoru Pederson. As functional nuclear actin comes into view, is it globular, filamentous, or both? Journal of Cell Biology, 180(6):1061–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/6/1061.

Pankov:2005:RSR

[PEER^{+}05] Roumen Pankov, Yukinori Endo, Sharona Even-Ram, Masaru Araki, Katherine Clark, Edna Cukierman, Kazue Matsumoto, and Kenneth M. Yamada. A Rac switch regulates random versus directionally persistent cell migration. Journal of Cell Biology, 170(5):793–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/5/793.

Platta:2007:UPI

[PES^{+}07] Harald W. Platta, Fouzi El Magraoui, Daniel Schlee, Silke Grunau, Wolfgang Girzalsky, and Ralf Erdmann. Ubiquitination of the peroxisomal import receptor Pex5p is required for its recycling. Journal of Cell Biology, 177(2):197–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/177/2/197.

Pratt:2005:LPA

[PEW^{+}05] Stephen J. Pratt, Holly Epple, Michael Ward, Yunfeng Feng, Vania M. Braga, and Gregory D. Longmore. The LIM protein Ajuba influences p130Cas localization and Rac1 activity during cell migration. Journal of Cell Biology, 168(5):813–??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/168/5/
813.

Pielage:2006:PSS

[PFD06] Jan Pielage, Richard D. Fetter, and Graeme W. Davis. A postsynaptic Spectrin scaffold defines active zone size, spacing, and efficacy at the Drosophila neuromuscular junction. Journal of Cell Biology, 175(3):491–??, November 2006. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/175/3/491.

Pfister:2005:CDN

[PFG^{+}05] K. Kevin Pfister, Elizabeth M. C. Fisher, Ian R. Gibbons, Thomas S. Hays, Erika L. F. Holzbaur, J. Richard McIntosh, Mary E. Porter, Trina A. Schroer, Kevin T. Vaughan, George B. Witman, Stephen M. King, and Richard B. Vallee. Cytoplasmic dynein nomenclature. Journal of Cell Biology, 171(3): 411–??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/3/411.

Puklin-Faucher:2006:HHH

[PFGSV06] Eileen Puklin-Faucher, Mu Gao, Klaus Schulten, and Viola Vogel. How the headpiece hinge angle is opened: new insights into the dynamics of integrin activation. Journal of Cell Biology, 175(2):349–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/2/349.

Patel:2006:DAF

[PG06] Jayesh C. Patel and Jorge E. Galán. Differential activation and function of Rho GTPases during Salmonella-host cell interactions. Journal of Cell Biology, 175(3):453–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/175/3/453.

Pak:2006:TLL

[PGB^{+}06] Youngshil Pak, Wioletta K. Glowacka, M. Christine Bruce, Nam Pham, and Daniela Rotin. Transport of LAPTM5 to lysosomes requires association with the ubiquitin ligase Nedd4, but not LAPTM5 ubiquitination. Journal of Cell Biology, 175(4): 631–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/4/631.

Pigino:2009:ETA

[PGL^{+}09] Gaia Pigino, Stefan Geimer, Salvatore Lanzavecchia, Eugenio Paccagnini, Francesca Cantele, Dennis R. Diener, Joel L. Rosenbaum, and Pietro Lupetti. Electron-tomographic analysis of intraflagellar transport particle trains in situ. Journal of Cell Biology, 187(1):135–??, October 2009. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/187/1/135.

Petri:2007:DSM

[PGO^{+}07] Sebastian Petri, Matthias Grimmler, Sabine Over, Utz Fischer, and Oliver J. Gruss. Dephosphorylation of survival motor neurons (SMN) by PPM1G/PP2Cγ governs Cajal body localization and stability of the SMN complex. Journal of Cell Biology, 179(3):451–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/3/451.

Palumbo:2007:CMS

[PGP^{+}07] Roberta Palumbo, Beatriz G. Galvez, Tobias Pusterla, Francesco De Marchis, Giulio Cossu, Kenneth B. Marcu, and Marco E. Bianchi. Cells migrating to sites of tissue damage in response to the danger signal HMGB1 require NF-κB activation. Journal of Cell Biology, 179(1):33–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/179/1/33.

Postow:2008:KRD

[PGW^{+}08] Lisa Postow, Cristina Ghenoiu, Eileen M. Woo, Andrew N. Krutchinsky, Brian T. Chait, and Hironori Funabiki. Ku80 removal from DNA through double strand break–induced ubiquitylation. Journal of Cell Biology, 182(3):467–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/182/3/467.

Pearse:2008:CBR

[PGWH08] Bradley R. Pearse, Luke Gabriel, Ning Wang, and Daniel N. Hebert. A cell-based reglucosylation assay demonstrates the role of GT1 in the quality control of a maturing glycoprotein. Journal of Cell Biology, 181(2):309–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/181/2/309.

Pearson:2007:RTI

[PH07] Gray W. Pearson and Tony Hunter. Real-time imaging reveals that noninvasive mammary epithelial acini can contain motile cells. Journal of Cell Biology, 179(7):1555–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/7/
1555.

Pekovic:2007:NLL

[PHB^{+}07] Vanja Pekovic, Jens Harborth, Jos L. V. Broers, Frans C. S. Ramaekers, Baziel van Engelen, Martin Lammens, Thomas von Zglinicki, Roland Foisner, Chris Hutchison, and Ewa Markiewicz. Nucleoplasmic LAP2α-lamin A complexes are required to maintain a proliferative state in human fibroblasts. Journal of Cell Biology, 176(2):163–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/176/2/163.

Prins:2009:DMA

[PHM^{+}09] Kurt W. Prins, Jill L. Humston, Amisha Mehta, Victoria Tate, Evelyn Ralston, and James M. Ervasti. Dystrophin is a microtubule-associated protein. Journal of Cell Biology, 186 (3):363–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/186/3/363.

Pelicano:2006:MRD

[PhXD^{+}06] Hélène Pelicano, Rui hua Xu, Min Du, Li Feng, Ryohei Sasaki, Jennifer S. Carew, Yumin Hu, Latha Ramdas, Limei Hu, Michael J. Keating, Wei Zhang, William Plunkett, and Peng Huang. Mitochondrial respiration defects in cancer cells cause activation of Akt survival pathway through a redox-mediated mechanism. Journal of Cell Biology, 175(6):913–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/175/6/
913.

Prass:2006:DML

[PJMR06] Marcus Prass, Ken Jacobson, Alex Mogilner, and Manfred Radmacher. Direct measurement of the lamellipodial protrusive force in a migrating cell. Journal of Cell Biology, 174(6): 767–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/6/767.

Proszynski:2006:PMP

[PKB^{+}06] Tomasz J. Proszynski, Robin Klemm, Michel Bagnat, Katharina Gaus, and Kai Simons. Plasma membrane polarization during mating in yeast cells. Journal of Cell Biology, 173(6):861–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
173/6/861.

Prange-Kiel:2008:GRH

[PKJS^{+}08] Janine Prange-Kiel, Hubertus Jarry, Michael Schoen, Patrick Kohlmann, Christina Lohse, Lepu Zhou, and Gabriele M. Rune. Gonadotropin-releasing hormone regulates spine density via its regulatory role in hippocampal estrogen synthesis. Journal of Cell Biology, 180(2):417–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/2/417.

Patel-King:2009:OAD

[PKK09] Ramila S. Patel-King and Stephen M. King. An outer arm dynein light chain acts in a conformational switch for flagellar motility. Journal of Cell Biology, 186(2):283–??, July 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/186/2/283.

Pozniakovsky:2005:RMP

[PKM^{+}05] Andrei I. Pozniakovsky, Dmitry A. Knorre, Olga V. Markova, Anthony A. Hyman, Vladimir P. Skulachev, and Fedor F. Severin. Role of mitochondria in the pheromone- and amiodaroneinduced programmed death of yeast. Journal of Cell Biology, 168(2):257–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/2/257.

Pageau:2006:BFN

[PL06] Gayle J. Pageau and Jeanne B. Lawrence. BRCA1 foci in normal S-phase nuclei are linked to interphase centromeres and replication of pericentric heterochromatin. Journal of Cell Biology, 175(5):693–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/5/693.

Pandithage:2008:RSF

[PLH^{+}08] Ruwin Pandithage, Richard Lilischkis, Kai Harting, Alexandra Wolf, Britta Jedamzik, Juliane Lüscher-Firzlaff, Jörg Vervoorts, Edwin Lasonder, Elisabeth Kremmer, Bernd Knöll, and Bernhard Lüscher. The regulation of SIRT2 function by cyclindependent kinases affects cell motility. Journal of Cell Biology, 180(5):915–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/5/915.

Park:2005:AFC

[PLLH05] Misoon Park, Daeseok Lee, Gil-Je Lee, and Inhwan Hwang. AtRMR1 functions as a cargo receptor for protein trafficking to the protein storage vacuole. Journal of Cell Biology, 170 (5):757–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/5/757.

Pirone:2006:IRF

[PLR^{+}06] Dana M. Pirone, Wendy F. Liu, Sami Alom Ruiz, Lin Gao, Srivatsan Raghavan, Christopher A. Lemmon, Lewis H. Romer, and Christopher S. Chen. An inhibitory role for FAK in regulating proliferation: a link between limited adhesion and RhoA– ROCK signaling. Journal of Cell Biology, 174(2):277–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/174/2/
277.

Paronetto:2009:SRT

[PMB^{+}09] Maria Paola Paronetto, Valeria Messina, Enrica Bianchi, Marco Barchi, Gillian Vogel, Costanzo Moretti, Fioretta Palombi, Mario Stefanini, Raffaele Geremia, Stéphane Richard, and Claudio Sette. Sam68 regulates translation of target mRNAs in male germ cells, necessary for mouse spermatogenesis. Journal of Cell Biology, 185(2):235–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/185/2/235.

Peter:2008:SRD

[PMC08] Angela K. Peter, Jamie L. Marshall, and Rachelle H. Crosbie. Sarcospan reduces dystrophic pathology: stabilization of the utrophin–glycoprotein complex. Journal of Cell Biology, 183 (3):419–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/3/419.

Piazzolla:2005:RST

[PMK^{+}05] Daniela Piazzolla, Katrin Meissl, Lucia Kucerova, Cristina Rubiolo, and Manuela Baccarini. Raf-1 sets the threshold of Fas sensitivity by modulating Rok-α signaling. Journal of Cell Biology, 171(6):1013–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/6/1013.

Porter:2007:BNK

[PMK^{+}07] Iain M. Porter, Sarah E. McClelland, Guennadi A. Khoudoli, Christopher J. Hunter, Jens S. Andersen, Andrew D. McAinsh, J. Julian Blow, and Jason R. Swedlow. Bod1, a novel kinetochore protein required for chromosome biorientation. Journal of Cell Biology, 179(2):187–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/2/187.

Pajalunga:2007:CRC

[PMS^{+}07] Deborah Pajalunga, Alessia Mazzola, Anna Maria Salzano, Maria Grazia Biferi, Gabriele De Luca, and Marco Crescenzi. Critical requirement for cell cycle inhibitors in sustaining nonproliferative states. Journal of Cell Biology, 176(6):807–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
176/6/807.

Parry:2005:MBE

[PMW05] Huw Parry, Alex McDougall, and Michael Whitaker. Microdomains bounded by endoplasmic reticulum segregate cell cycle calcium transients in syncytial Drosophila embryos. Journal of Cell Biology, 171(1):47–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/171/1/47.

Patel:2007:SIR

[PNB07] Snehal Bhikhu Patel, Natalya Novikova, and Michel Bellini. Splicing-independent recruitment of spliceosomal small nuclear RNPs to nascent RNA polymerase II transcripts. Journal of Cell Biology, 178(6):937–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/178/6/937.

Perpelescu:2009:AEC

[PNO^{+}09] Marinela Perpelescu, Naohito Nozaki, Chikashi Obuse, Hua Yang, and Kinya Yoda. Active establishment of centromeric CENP–A chromatin by RSF complex. Journal of Cell Biology, 185(3):397–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/3/397.

Pan:2006:MTI

[POCS^{+}06] Xiaoyu Pan, Guangshuo Ou, Gul Civelekoglu-Scholey, Oliver E. Blacque, Nicholas F. Endres, Li Tao, Alex Mogilner, Michel R. Leroux, Ronald D. Vale, and Jonathan M. Scholey. Mechanism of transport of IFT particles in C. elegans cilia by the concerted action of kinesin–II and OSM-3 motors. Journal of Cell Biology, 174(7):1035–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/7/1035.

Pearson:2009:BBS

[POG^{+}09] Chad G. Pearson, Daniel P. S. Osborn, Thomas H. Giddings, Philip L. Beales, and Mark Winey. Basal body stability and ciliogenesis requires the conserved component Poc1. Journal of Cell Biology, 187(6):905–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/187/6/905.

Paavilainen:2008:SAD

[POGL08] Ville O. Paavilainen, Esko Oksanen, Adrian Goldman, and Pekka Lappalainen. Structure of the actin-depolymerizing factor homology domain in complex with actin. Journal of Cell Biology, 182(1):51–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/1/51.

Polakis:2008:FBB

[Pol08] Paul Polakis. Formation of the blood–brain barrier: Wnt signaling seals the deal. Journal of Cell Biology, 183(3):371– ??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/3/371.

Pool:2009:TMS

[Poo09] Martin R. Pool. A trans-membrane segment inside the ribosome exit tunnel triggers RAMP4 recruitment to the Sec61p translocase. Journal of Cell Biology, 185(5):889–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/185/5/889.

Powell:2005:AMI

[Pow05a] Kendall Powell. Actin and microtubules interact via MAPs. Journal of Cell Biology, 170(5):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/170/5/701.2.

Powell:2005:APB

[Pow05b] Kendall Powell. Actin pushes in bizarre places. Journal of Cell Biology, 170(2):167–??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/2/167.

Powell:2005:BLI

[Pow05c] Kendall Powell. Basal lamina instructs innervation. Journal of Cell Biology, 170(5):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/5/701.1.

Powell:2005:BBR

[Pow05d] Kendall Powell. A big BiP on the radar screen. Journal of Cell Biology, 171(5):759–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/171/5/759.

Powell:2005:BCC

[Pow05e] Kendall Powell. Building a case for the chromosome scaffold. Journal of Cell Biology, 171(5):758–??, December 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/171/5/758.

Powell:2005:CPB

[Pow05f] Kendall Powell. Coated pits bring in the yolk. Journal of Cell Biology, 169(1):18–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/1/18.

Powell:2005:CM

[Pow05g] Kendall Powell. Contacting the matrix. Journal of Cell Biology, 170(4):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/4/515.1.

Powell:2005:CRS

[Pow05h] Kendall Powell. Curbside recycling at the synapse. Journal of Cell Biology, 170(2):166–??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/170/2/166.

Powell:2005:DBF

[Pow05i] Kendall Powell. Dishing up bone formation. Journal of Cell Biology, 171(3):??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/3/409.2.

Powell:2005:ESP

[Pow05j] Kendall Powell. Excess secretory products fuse with lysosomes. Journal of Cell Biology, 169(2):??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/169/2/216.1.

Powell:2005:FS

[Pow05k] Kendall Powell. The first supper. Journal of Cell Biology, 169(2):??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/2/216.2.

Powell:2005:FEE

[Pow05l] Kendall Powell. Frog egg extracts can do a cell’s work. Journal of Cell Biology, 171(4):585–??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/171/4/585.

Powell:2005:MML

[Pow05m] Kendall Powell. A macrophage mystery leads to dendritic discovery. Journal of Cell Biology, 170(1):??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/170/1/12.2.

Powell:2005:MGC

[Pow05n] Kendall Powell. Microtubules as a gyroscope for cells on the go. Journal of Cell Biology, 170(5):700–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/170/5/700.

Powell:2005:MPP

[Pow05o] Kendall Powell. Microtubules park parallel in the half-spindle. Journal of Cell Biology, 171(1):12–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/171/1/12.

Powell:2005:MTR

[Pow05p] Kendall Powell. Microtubules turn over rapidly. Journal of Cell Biology, 171(6):917–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/171/6/917.

Powell:2005:MPC

[Pow05q] Kendall Powell. Myosin powers cytokinesis. Journal of Cell Biology, 170(4):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/4/515.2.

Powell:2005:PML

[Pow05r] Kendall Powell. “porterplasm” and the microtrabecular lattice. Journal of Cell Biology, 170(6):864–??, September 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/170/6/864.

Powell:2005:PM

[Pow05s] Kendall Powell. Pursuing the middleman. Journal of Cell Biology, 171(4):584–??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/4/584.

Powell:2005:SIT

[Pow05t] Kendall Powell. Sticking it out with tight junctions. Journal of Cell Biology, 171(6):916–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/171/6/916.

Powell:2005:SBD

[Pow05u] Kendall Powell. The sticky business of discovering cadherins. Journal of Cell Biology, 170(4):514–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/170/4/514.

Powell:2005:SSP

[Pow05v] Kendall Powell. Sugars sprinkled onto proteins in the Golgi. Journal of Cell Biology, 169(1):19–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/169/1/19.

Powell:2005:VCE

[Pow05w] Kendall Powell. Viruses catch an endocytic ride into the cell. Journal of Cell Biology, 170(7):1018–??, September 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/170/7/1018.

Powell:2005:WMC

[Pow05x] Kendall Powell. Were mitochondrial contractions driving the cellular energy cycle? Journal of Cell Biology, 169(2):217–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
169/2/217.

Powell:2005:WCR

[Pow05y] Kendall Powell. What the cytoskeleton really looks like. Journal of Cell Biology, 170(7):1019–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/170/7/1019.

Powell:2005:YBC

[Pow05z] Kendall Powell. Yeast becomes a cell biologist. Journal of Cell Biology, 171(3):408–??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/171/3/408.

Powell:2006:ATG

[Pow06a] Kendall Powell. Anchored telomeres get longer. Journal of Cell Biology, 172(6):786–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/6/786.

Powell:2006:BSP

[Pow06b] Kendall Powell. BFA sends proteins back. Journal of Cell Biology, 172(4):484–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/4/484.

Powell:2006:CSL

[Pow06c] Kendall Powell. Chromosome snatching by lateral microtubules. Journal of Cell Biology, 172(6):788–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/172/6/788.

Powell:2006:ESE

[Pow06d] Kendall Powell. ECM signals ECM degradation. Journal of Cell Biology, 172(5):642–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/172/5/642.

Powell:2006:FG

[Pow06e] Kendall Powell. Filling in the gap. Journal of Cell Biology, 172 (1):6–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/1/6.

Powell:2006:HBM

[Pow06f] Kendall Powell. Hungry for better memory. Journal of Cell Biology, 172(6):??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/6/787b.

Powell:2006:KKR

[Pow06g] Kendall Powell. Killer K–Ras. Journal of Cell Biology, 172 (6):??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/6/786a.

Powell:2006:KU

[Pow06h] Kendall Powell. The kinetochore uncoiled. Journal of Cell Biology, 173(1):7–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/1/7.

Powell:2006:LRI

[Pow06i] Kendall Powell. Lipid raft idea is floated. Journal of Cell Biology, 172(2):166–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/2/166.

Powell:2006:MM

[Pow06j] Kendall Powell. MAP 1C is a motor. Journal of Cell Biology, 172(1):7–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/1/7.

Powell:2006:SM

[Pow06k] Kendall Powell. Skating on MTs. Journal of Cell Biology, 172 (6):787–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/6/787.

Powell:2006:UKS

[Pow06l] Kendall Powell. An unattached kinetochore screams “Wait!”. Journal of Cell Biology, 173(2):152–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/173/2/152.

Powell:2006:VDC

[Pow06m] Kendall Powell. Virus diverts checkpoint. Journal of Cell Biology, 172(6):??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/6/787a.

Powell:2007:BPA

[Pow07a] Kendall Powell. Bim plays apoptosis offense. Journal of Cell Biology, 177(4):??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/4/566a.

Powell:2007:BMS

[Pow07b] Kendall Powell. A bouquet for meiotic spindle. Journal of Cell Biology, 178(4):??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/4/544.1.

Powell:2007:CPP

[Pow07c] Kendall Powell. Cells’ precise pattern readout. Journal of Cell Biology, 178(4):??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/4/544.2.

Powell:2007:CR

[Pow07d] Kendall Powell. Chemotaxis in reverse. Journal of Cell Biology, 177(4):567–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/4/567.

Powell:2007:CGW

[Pow07e] Kendall Powell. Collective guidance wisdom. Journal of Cell Biology, 178(4):??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/4/545.3.

Powell:2007:FSL

[Pow07f] Kendall Powell. Fluid for a single lumen. Journal of Cell Biology, 178(4):??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/4/545.2.

Powell:2007:MMMa

[Pow07g] Kendall Powell. Motoring microvillar membrane. Journal of Cell Biology, 177(4):??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/4/567b.

Powell:2007:MMMb

[Pow07h] Kendall Powell. Motoring microvillar membrane. Journal of Cell Biology, 177(5):941–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/177/5/941.

Powell:2007:SCB

[Pow07i] Kendall Powell. SINEs create boundaries. Journal of Cell Biology, 178(4):??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/4/545.1.

Powell:2007:SS

[Pow07j] Kendall Powell. STALLing for signaling. Journal of Cell Biology, 177(4):566–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/4/566.

Powell:2007:ZM

[Pow07k] Kendall Powell. Zinc messenger. Journal of Cell Biology, 177(4):??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/4/567a.

Powell:2009:ICM

[Pow09a] Kendall Powell. Irina Conboy: Making the old feel young again. Journal of Cell Biology, 187(1):4–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/187/1/4.

Powell:2009:PPN

[Pow09b] Kendall Powell. PIKfyve promotes neuron self-preservation. Journal of Cell Biology, 187(2):??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/187/2/150.3.

Powell:2009:SNG

[Pow09c] Kendall Powell. Securing the next generation’s peroxisomes. Journal of Cell Biology, 187(2):??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/187/2/150.2.

Powell:2009:WFC

[Pow09d] Kendall Powell. Walking a fine caspase line. Journal of Cell Biology, 187(2):??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/2/150.1.

Pajcini:2008:MMS

[PPA^{+}08] Kostandin V. Pajcini, Jason H. Pomerantz, Ozan Alkan, Regis Doyonnas, and Helen M. Blau. Myoblasts and macrophages share molecular components that contribute to cell–cell fusion. Journal of Cell Biology, 180(5):1005–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/180/5/1005.

Pirzio:2008:WSH

[PPBF08] Livia Maria Pirzio, Pietro Pichierri, Margherita Bignami, and Annapaola Franchitto. Werner syndrome helicase activity is essential in maintaining fragile site stability. Journal of Cell Biology, 180(2):305–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/2/305.

Paladino:2006:GAP

[PPCZ06] Simona Paladino, Thomas Pocard, Maria Agata Catino, and Chiara Zurzolo. GPI-anchored proteins are directly targeted to the apical surface in fully polarized MDCK cells. Journal of Cell Biology, 172(7):1023–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/172/7/1023.

Pal:2008:RMC

[PPK08] Gayatri Pal, Maria T. Z. Paraz, and Douglas R. Kellogg. Regulation of Mih1/Cdc25 by protein phosphatase 2A and casein kinase 1. Journal of Cell Biology, 180(5):931–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/180/5/931.

Poncet:2006:CEC

[PPS^{+}06] Delphine Poncet, Anne-Laure Pauleau, Gyorgy Szabadkai, Angelo Vozza, Sebastian R. Scholz, Morgane Le Bras, Jean-Jacques Brière, Abdelali Jalil, Ronan Le Moigne, Catherine Brenner, Gabriele Hahn, Ilka Wittig, Hermann Schägger, Christophe Lemaire, Katiuscia Bianchi, Sylvie Souquère, Gerard Pierron, Pierre Rustin, Victor S. Goldmacher, Rosario Rizzuto, Ferdinando Palmieri, and Guido Kroemer. Cytopathic effects of the cytomegalovirus-encoded apoptosis inhibitory protein vMIA. Journal of Cell Biology, 174(7):985–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/174/7/
985.

Pacquelet:2005:RMR

[PR05] Anne Pacquelet and Pernille Rørth. Regulatory mechanisms required for DE-cadherin function in cell migration and other types of adhesion. Journal of Cell Biology, 170(5):803–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
170/5/803.

Pilichou:2009:MNU

[PRB^{+}09] Kalliopi Pilichou, Carol Ann Remme, Cristina Basso, Maria E. Campian, Stefania Rizzo, Phil Barnett, Brendon P. Scicluna, Barbara Bauce, Maurice J. B. van den Hoff, Jacques M. T. de Bakker, Hanno L. Tan, Marialuisa Valente, Andrea Nava, Arthur A. M. Wilde, Antoon F. M. Moorman, Gaetano Thiene, and Connie R. Bezzina. Myocyte necrosis underlies progressive myocardial dystrophy in mouse dsg2 -related arrhythmogenic right ventricular cardiomyopathy. Journal of Cell Biology, 186(3):??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/186/3/i5.

Pi:2007:SRM

[PRK^{+}07] Xinchun Pi, Rongqin Ren, Russell Kelley, Chunlian Zhang, Martin Moser, Aparna B. Bohil, Melinda DiVito, Richard E. Cheney, and Cam Patterson. Sequential roles for myosin–X in BMP6-dependent filopodial extension, migration, and activation of BMP receptors. Journal of Cell Biology, 179(7): 1569–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/7/1569.

Pi:2009:SRM

[PRK^{+}09] Xinchun Pi, Rongqin Ren, Russell Kelley, Chunlian Zhang, Martin Moser, Aparna B. Bohil, Melinda DiVito, Richard E. Cheney, and Cam Patterson. Sequential roles for myosin–X in BMP6-dependent filopodial extension, migration, and activation of BMP receptors. Journal of Cell Biology, 184(4): 611–??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/184/4/611.

Panupinthu:2008:PRO

[PRZ^{+}08] Nattapon Panupinthu, Joseph T. Rogers, Lin Zhao, Luis Pastor Solano-Flores, Fred Possmayer, Stephen M. Sims, and S. Jeffrey Dixon. P2X7 receptors on osteoblasts couple to production of lysophosphatidic acid: a signaling axis promoting osteogenesis. Journal of Cell Biology, 181(5):859–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/181/5/859.

Pawlicki:2008:PMT

[PS08] Jan M. Pawlicki and Joan A. Steitz. Primary microRNA transcript retention at sites of transcription leads to enhanced microRNA production. Journal of Cell Biology, 182(1):61–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/182/1/
61.

Pal:2006:HHC

[PSL^{+}06] Arun Pal, Fedor Severin, Barbara Lommer, Anna Shevchenko, and Marino Zerial. Huntingtin-HAP40 complex is a novel Rab5 effector that regulates early endosome motility and is up-regulated in Huntington’s disease. Journal of Cell Biology, 172(4):605–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/4/605.

Peset:2005:FRM

[PSS^{+}05] Isabel Peset, Jeanette Seiler, Teresa Sardon, Luis A. Bejarano, Sonja Rybina, and Isabelle Vernos. Function and regulation of Maskin, a TACC family protein, in microtubule growth during mitosis. Journal of Cell Biology, 170(7):1057–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/170/7/
1057.

Preuss:2006:RRE

[PST^{+}06] Mary L. Preuss, Aaron J. Schmitz, Julie M. Thole, Heather K. S. Bonner, Marisa S. Otegui, and Erik Nielsen. A role for the RabA4b effector protein PI-4Kβ1 in polarized expansion of root hair cells in Arabidopsis thaliana. Journal of Cell Biology, 172(7):991–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/7/991.

Perlin:2007:PGG

[PT07] Julie R. Perlin and William S. Talbot. Putting the glue in glia: Necls mediate Schwann cell–axon adhesion. Journal of Cell Biology, 178(5):721–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/5/721.

Pederson:2009:SNN

[PT09] Thoru Pederson and Robert Y. L. Tsai. In search of nonribosomal nucleolar protein function and regulation. Journal of Cell Biology, 184(6):771–??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/6/771.

Peris:2006:TTM

[PTF^{+}06] Leticia Peris, Manuel Thery, Julien Fauré, Yasmina Saoudi, Laurence Lafanechère, John K. Chilton, Phillip Gordon-Weeks, Niels Galjart, Michel Bornens, Linda Wordeman, Juergen Wehland, Annie Andrieux, and Didier Job. Tubulin tyrosination is a major factor affecting the recruitment of CAP– Gly proteins at microtubule plus ends. Journal of Cell Biology, 174(6):839–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/6/839.

Pante:2005:MIG

[PTL^{+}05] Guido Pante, Jane Thompson, Fabienne Lamballe, Tomoko Iwata, Ingvar Ferby, Francis A. Barr, Alun M. Davies, Flavio Maina, and Rüdiger Klein. Mitogen-inducible gene 6 is an endogenous inhibitor of HGF/Met-induced cell migration and neurite growth. Journal of Cell Biology, 171(2):337–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/171/2/
337.

Tong:2009:CSE

[pTyLZ^{+}09] Xiao ping Tong, Xiang yao Li, Bing Zhou, Wanhua Shen, Zhi jun Zhang, Tian le Xu, and Shumin Duan. Ca^{2+} signaling evoked by activation of Na^{+} channels and Na^{+} /Ca^{2+} exchangers is required for GABA-induced NG2 cell migration. Journal of Cell Biology, 186(1):113–??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/186/1/113.

Pardo:2006:MPB

[PUG^{+}06] Julian Pardo, Christin Urban, Eva M. Galvez, Paul G. Ekert, Uwe Müller, June Kwon-Chung, Mario Lobigs, Arno Müllbacher, Reinhard Wallich, Christoph Borner, and Markus M. Simon. The mitochondrial protein Bak is pivotal for gliotoxininduced apoptosis and a critical host factor of Aspergillus fumigatus virulence in mice. Journal of Cell Biology, 174(4):509–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
174/4/509.

Putney:2005:CCE

[Put05] James W. Putney. Capacitative calcium entry. Journal of Cell Biology, 169(3):381–??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/3/381.

Percipalle:2006:MFN

[PV06] Piergiorgio Percipalle and Neus Visa. Molecular functions of nuclear actin in transcription. Journal of Cell Biology, 172 (7):967–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/7/967.

Paluch:2006:CSB

[PvdGS06] Ewa Paluch, Jasper van der Gucht, and Cécile Sykes. Cracking up: symmetry breaking in cellular systems. Journal of Cell Biology, 175(5):687–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/5/687.

Puhka:2007:ERR

[PVJJ07] Maija Puhka, Helena Vihinen, Merja Joensuu, and Eija Jokitalo. Endoplasmic reticulum remains continuous and undergoes sheet-to-tubule transformation during cell division in mammalian cells. Journal of Cell Biology, 179(5):895–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/5/
895.

Potts:2005:RAL

[PVM^{+}05] Malia B. Potts, Allyson E. Vaughn, Holly McDonough, Cam Patterson, and Mohanish Deshmukh. Reduced Apaf-1 levels in cardiomyocytes engage strict regulation of apoptosis by endogenous XIAP. Journal of Cell Biology, 171(6):925–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/171/6/
925.

Poggi:2005:INL

[PVMH05] Lucia Poggi, Marta Vitorino, Ichiro Masai, and William A. Harris. Influences on neural lineage and mode of division in the zebrafish retina in vivo. Journal of Cell Biology, 171(6):

991–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/6/991.

Phalen:2006:OSG

[PWD^{+}06] Timothy J. Phalen, Kelly Weirather, Paula B. Deming, Vikas Anathy, Alan K. Howe, Albert van der Vliet, Thomas J. Jönsson, Leslie B. Poole, and Nicholas H. Heintz. Oxidation state governs structural transitions in peroxiredoxin II that correlate with cell cycle arrest and recovery. Journal of Cell Biology, 175(5):779–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/5/779.

Peris:2009:MDM

[PWL^{+}09] Leticia Peris, Michael Wagenbach, Laurence Lafanechère, Jacques Brocard, Ayana T. Moore, Frank Kozielski, Didier Job, Linda Wordeman, and Annie Andrieux. Motor-dependent microtubule disassembly driven by tubulin tyrosination. Journal of Cell Biology, 185(7):1159–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/185/7/1159.

Pal:2006:ANB

[PWM^{+}06a] Saumen Pal, Jing Wu, Justin K. Murray, Samuel H. Gellman, Michele A. Wozniak, Patricia J. Keely, Meghan E. Boyer, Timothy M. Gomez, Sean M. Hasso, John F. Fallon, and Emery H. Bresnick. An antiangiogenic neurokinin– B/thromboxane A2 regulatory axis. Journal of Cell Biology, 174(7):1047–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/7/1047.

Pal:2006:CP

[PWM^{+}06b] Saumen Pal, Jing Wu, Justin K. Murray, Samuel H. Gellman, Michele A. Wozniak, Patricia J. Keely, Meghan E. Boyer, Timothy M. Gomez, Sean M. Hasso, John F. Fallon, and Emery H. Bresnick. Correction 174, p. 1047. Journal of Cell Biology, 175 (4):671–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/4/671.

Polevoy:2009:DRD

[PWW^{+}09] Gordon Polevoy, Ho-Chun Wei, Raymond Wong, Zsofia Szentpetery, Yeun Ju Kim, Philip Goldbach, Sarah K. Steinbach, Tamas Balla, and Julie A. Brill. Dual roles for the Drosophila PI 4-kinase Four wheel drive in localizing Rab11 during cytokinesis. Journal of Cell Biology, 187(6):847–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/187/6/847.

Patra:2007:SPE

[PXD^{+}07] Debabrata Patra, Xiaoyun Xing, Sherri Davies, Jennifer Bryan, Carl Franz, Ernst B. Hunziker, and Linda J. Sandell. Site-1 protease is essential for endochondral bone formation in mice. Journal of Cell Biology, 179(4):687–??, November 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/179/4/687.

Paterson:2008:DLY

[PYR08] Joanna Mathis Paterson, Casey A. Ydenberg, and Mark D. Rose. Dynamic localization of yeast Fus2p to an expanding ring at the cell fusion junction during mating. Journal of Cell Biology, 181(4):697–??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/4/697.

Provot:2007:HRD

[PZG^{+}07] Sylvain Provot, Dawn Zinyk, Yasemin Gunes, Richa Kathri, Quynh Le, Henry M. Kronenberg, Randall S. Johnson, Michael T. Longaker, Amato J. Giaccia, and Ernestina Schipani. Hif-1α regulates differentiation of limb bud mesenchyme and joint development. Journal of Cell Biology, 177 (3):451–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/3/451.

Qin:2005:MSP

[QCGM05] Yi Qin, Christopher Capaldo, Barry M. Gumbiner, and Ian G. Macara. The mammalian Scribble polarity protein regulates epithelial cell adhesion and migration through E-cadherin. Journal of Cell Biology, 171(6):1061–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/171/6/1061.

Gan:2008:NDC

[qGyWX^{+}08] Xiao qing Gan, Ji yong Wang, Ying Xi, Zhi li Wu, Yi ping Li, and Lin Li. Nuclear Dvl, c–Jun, β-catenin, and TCF form a complex leading to stabiLization of β–catenin–TCF interaction. Journal of Cell Biology, 180(6):1087–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/180/6/1087.

Quinlan:2007:RIB

[QHB^{+}07] Margot E. Quinlan, Susanne Hilgert, Anaid Bedrossian, R. Dyche Mullins, and Eugen Kerkhoff. Regulatory interactions between two actin nucleators, Spire and Cappuccino. Journal of Cell Biology, 179(1):117–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/1/117.

Qin:2006:NGC

[QIH^{+}06] Qingyu Qin, Ryoko Inatome, Azusa Hotta, Masaki Kojima, Hirohei Yamamura, Hirokazu Hirai, Toshihiro Yoshizawa, Hirofumi Tanaka, Kiyoko Fukami, and Shigeru Yanagi. A novel GTPase, CRAG, mediates promyelocytic leukemia protein– associated nuclear body formation and degradation of expanded polyglutamine protein. Journal of Cell Biology, 172(4): 497–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/4/497.

Quarmby:2005:CCC

[QP05] Lynne M. Quarmby and Jeremy D. K. Parker. Cilia and the cell cycle? Journal of Cell Biology, 169(5):707–??, June 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/169/5/707.

Queralt:2008:SCZ

[QU08] Ethel Queralt and Frank Uhlmann. Separase cooperates with Zds1 and Zds2 to activate Cdc14 phosphatase in early anaphase. Journal of Cell Biology, 182(5):873–??, September 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/182/5/873.

Qiu:2008:RNP

[QWD^{+}08] Runxiang Qiu, Xiuyun Wang, Alice Davy, Chen Wu, Kiyohito Murai, Heying Zhang, John G. Flanagan, Philippe Soriano, and Qiang Lu. Regulation of neural progenitor cell state by ephrin– B. Journal of Cell Biology, 181(6):973–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/181/6/973.

Rapaport:2005:HDT

[Rap05] Doron Rapaport. How does the TOM complex mediate insertion of precursor proteins into the mitochondrial outer membrane? Journal of Cell Biology, 171(3):419–??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/171/3/419.

Rusan:2008:PMT

[RAP08] Nasser M. Rusan, Kathryn Akong, and Mark Peifer. Putting the model to the test: are APC proteins essential for neuronal polarity, axon outgrowth, and axon targeting? Journal of Cell Biology, 183(2):203–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/2/203.

Raykhel:2007:MSC

[RAS^{+}07] Irina Raykhel, Heli Alanen, Kirsi Salo, Jaana Jurvansuu, Van Dat Nguyen, Maria Latva-Ranta, and Lloyd Ruddock. A molecular specificity code for the three mammalian KDEL receptors. Journal of Cell Biology, 179(6):1193–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/6/
1193.

Rasband:2008:CGA

[Ras08a] Matthew N. Rasband. Na^{+} channels get anchored … with a little help. Journal of Cell Biology, 183(6):975–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/183/6/
975.

Raykhel:2008:MSC

[RAS^{+}08b] Irina Raykhel, Heli Alanen, Kirsi Salo, Jaana Jurvansuu, Van Dat Nguyen, Maria Latva-Ranta, and Lloyd Ruddock. A molecular specificity code for the three mammalian KDEL receptors. Journal of Cell Biology, 180(3):645–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/180/3/645.

Richards:2005:TME

[RBC05] David A. Richards, Jihong Bai, and Edwin R. Chapman. Two modes of exocytosis at hippocampal synapses revealed by rate of FM1-43 efflux from individual vesicles. Journal of Cell Biology, 168(6):929–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/6/929.

Reverte:2006:PIF

[RBJL06] Carlos G. Reverte, Angela Benware, Christopher W. Jones, and Susan E. LaFlamme. Perturbing integrin function inhibits microtubule growth from centrosomes, spindle assembly, and cytokinesis. Journal of Cell Biology, 174(4):491–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/174/4/491.

Ronchi:2008:TDD

[RCFB08] Paolo Ronchi, Sara Colombo, Maura Francolini, and Nica Borgese. Transmembrane domain-dependent partitioning of membrane proteins within the endoplasmic reticulum. Journal of Cell Biology, 181(1):105–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/181/1/105.

Roberts:2006:HOR

[RCKD06] Esteban A. Roberts, Jennifer Chua, George B. Kyei, and Vojo Deretic. Higher order Rab programming in phagolysosome biogenesis. Journal of Cell Biology, 174(7):923–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/174/7/
923.

Raghu:2009:RBD

[RCM^{+}09] Padinjat Raghu, Elise Coessens, Maria Manifava, Plamen Georgiev, Trevor Pettitt, Eleanor Wood, Isaac Garcia-Murillas, Hanneke Okkenhaug, Deepti Trivedi, Qifeng Zhang, Azam Razzaq, Ola Zaid, Michael Wakelam, Cahir J. O’Kane, and Nicholas Ktistakis. Rhabdomere biogenesis in Drosophila photoreceptors is acutely sensitive to phosphatidic acid levels. Journal of Cell Biology, 185(1):129–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/185/1/129.

Rex:2009:DRG

[RCS^{+}09] Christopher S. Rex, Lulu Y. Chen, Anupam Sharma, Jihua Liu, Alex H. Babayan, Christine M. Gall, and Gary Lynch. Different Rho GTPase-dependent signaling pathways initiate sequential steps in the consolidation of long-term potentiation. Journal of Cell Biology, 186(1):85–??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/186/1/85.

Razi:2009:EEE

[RCT09] Minoo Razi, Edmond Y. W. Chan, and Sharon A. Tooze. Early endosomes and endosomal coatomer are required for autophagy. Journal of Cell Biology, 185(2):305–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/185/2/305.

Roos:2005:SEC

[RDY^{+}05] Jack Roos, Paul J. DiGregorio, Andriy V. Yeromin, Kari Ohlsen, Maria Lioudyno, Shenyuan Zhang, Olga Safrina, J. Ashot Kozak, Steven L. Wagner, Michael D. Cahalan, Gönül Veliçelebi, and Kenneth A. Stauderman. STIM1, an essential and conserved component of store-operated Ca^{2+} channel function. Journal of Cell Biology, 169(3):435–??, May 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/169/3/435.

Rexach:2009:PTN

[Rex09] Michael Rexach. Piecing together nuclear pore complex assembly during interphase. Journal of Cell Biology, 185(3):377–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
185/3/377.

Rosario:2007:NOM

[RFBB07] Marta Rosário, Renate Franke, Christien Bednarski, and Walter Birchmeier. The neurite outgrowth multiadaptor RhoGAP, NOMA–GAP, regulates neurite extension through SHP2 and Cdc42. Journal of Cell Biology, 178(3):503–??, July 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/178/3/503.

Rhee:2006:PAK

[RG06a] Sangmyung Rhee and Frederick Grinnell. P21-activated kinase 1: convergence point in PDGF- and LPA-stimulated collagen matrix contraction by human fibroblasts. Journal of Cell Biology, 172(3):423–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/3/423.

Rice:2006:PLM

[RG06b] Sarah E. Rice and Vladimir I. Gelfand. Paradigm lost: milton connects kinesin heavy chain to miro on mitochondria. Journal of Cell Biology, 173(4):459–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/173/4/459.

Rosenberg:2006:MIF

[RGA^{+}06] Miriam I. Rosenberg, Sara A. Georges, Amy Asawachaicharn, Erwin Analau, and Stephen J. Tapscott. MyoD inhibits Fstl1 and Utrn expression by inducing transcription of miR-206. Journal of Cell Biology, 175(1):77–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/175/1/77.

Roberts-Galbraith:2009:SDT

[RGCWG09] Rachel H. Roberts-Galbraith, Jun-Song Chen, Jianqiu Wang, and Kathleen L. Gould. The SH3 domains of two PCH family members cooperate in assembly of the Schizosaccharomyces pombe contractile ring. Journal of Cell Biology, 184(1):113–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
184/1/113.

Rahmani:2009:SSC

[RGL^{+}09] Zohra Rahmani, Mary E. Gagou, Christophe Lefebvre, Doruk Emre, and Roger E. Karess. Separating the spindle, checkpoint, and timer functions of BubR1. Journal of Cell Biology, 187(5):597–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/5/597.

Rajendra:2007:DMM

[RGW^{+}07] T. K. Rajendra, Graydon B. Gonsalvez, Michael P. Walker, Karl B. Shpargel, Helen K. Salz, and A. Gregory Matera. A Drosophila melanogaster model of spinal muscular atrophy reveals a function for SMN in striated muscle. Journal of Cell Biology, 176(6):831–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/6/831.

Razafsky:2009:BKU

[RH09] David Razafsky and Didier Hodzic. Bringing KASH under the SUN: the many faces of nucleo-cytoskeletal connections. Journal of Cell Biology, 186(4):461–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/186/4/461.

Reiter:2005:VKT

[RHMS05] Jochen Reiter, Eva Herker, Frank Madeo, and Manfred J. Schmitt. Viral killer toxins induce caspase-mediated apoptosis in yeast. Journal of Cell Biology, 168(3):353–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/168/3/
353.

Reintsch:2005:CCC

[RHMW^{+}05] Wolfgang E. Reintsch, Anette Habring-Mueller, Renee W. Wang, Anne Schohl, and François Fagotto. β-catenin controls cell sorting at the notochord–somite boundary independently of cadherin-mediated adhesion. Journal of Cell Biology, 170 (4):675–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/4/675.

Raychaudhuri:2006:NSM

[RIHP06] Sumana Raychaudhuri, Young Jun Im, James H. Hurley, and William A. Prinz. Nonvesicular sterol movement from plasma membrane to ER requires oxysterol-binding protein–related proteins and phosphoinositides. Journal of Cell Biology, 173 (1):107–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/1/107.

Rybkin:2007:RAN

[RKB^{+}07] Igor I. Rybkin, Mi-Sung Kim, Svetlana Bezprozvannaya, Xiaoxia Qi, James A. Richardson, Craig F. Plato, Joseph A. Hill, Rhonda Bassel-Duby, and Eric N. Olson. Regulation of atrial natriuretic peptide secretion by a novel Ras-like protein. Journal of Cell Biology, 179(3):527–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/3/527.

Rollason:2009:CTR

[RKH^{+}09] Ruth Rollason, Viktor Korolchuk, Clare Hamilton, Mark Jepson, and George Banting. A CD317/tetherin–RICH2 complex plays a critical role in the organization of the subapical actin cytoskeleton in polarized epithelial cells. Journal of Cell Biology, 184(5):721–??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/5/721.

Roy:2009:KAA

[RKK09] Badal Chandra Roy, Naoto Kakinuma, and Ryoiti Kiyama. Kank attenuates actin remodeling by preventing interaction between IRSp53 and Rac1. Journal of Cell Biology, 184(2):253–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
184/2/253.

Rezniczek:2007:PSS

[RKN^{+}07] Günther A. Rezniczek, Patryk Konieczny, Branislav Nikolic, Siegfried Reipert, Doris Schneller, Christina Abrahamsberg, Kay E. Davies, Steve J. Winder, and Gerhard Wiche. Plectin 1f scaffolding at the sarcolemma of dystrophic (mdx) muscle fibers through multiple interactions with β-dystroglycan. Journal of Cell Biology, 176(7):965–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/176/7/965.

Ribeiro:2007:DFM

[RKT^{+}07] Paulo S. Ribeiro, Erina Kuranaga, Tencho Tenev, François Leulier, Masayuki Miura, and Pascal Meier. DIAP2 functions as a mechanism-based regulator of drICE that contributes to the caspase activity threshold in living cells. Journal of Cell Biology, 179(7):1467–??, December 2007. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/7/1467.

Rocha:2009:CSO

[RKvdK^{+}09] Nuno Rocha, Coenraad Kuijl, Rik van der Kant, Lennert Janssen, Diane Houben, Hans Janssen, Wilbert Zwart, and Jacques Neefjes. Cholesterol sensor ORP1L contacts the ER protein VAP to control Rab7–RILP–p150^{Glued} and late endosome positioning. Journal of Cell Biology, 185(7):1209–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/185/7/
1209.

Rolland:2009:BLP

[RLDC09] Stéphane G. Rolland, Yun Lu, Charles N. David, and Barbara Conradt. The BCL-2-like protein CED-9 of C. elegans promotes FZO-1/Mfn1,2- and EAT-3/Opa1-dependent mitochondrial fusion. Journal of Cell Biology, 186(4):525–??, August 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/186/4/525.

Rowe:2009:MCR

[RLH^{+}09] R. Grant Rowe, Xiao-Yan Li, Yuexian Hu, Thomas L. Saunders, Ismo Virtanen, Antonio Garcia de Herreros, Karl-Friedrich Becker, Signe Ingvarsen, Lars H. Engelholm, Guido T. Bommer, Eric R. Fearon, and Stephen J. Weiss. Mesenchymal cells reactivate Snail1 expression to drive three-dimensional invasion programs. Journal of Cell Biology, 184(3):399–??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/184/3/
399.

Reese:2005:THP

[RM05a] Christoph Reese and Andreas Mayer. Transition from hemifusion to pore opening is rate limiting for vacuole membrane fusion. Journal of Cell Biology, 171(6):981–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/171/6/981.

Riddick:2005:SAI

[RM05b] Gregory Riddick and Ian G. Macara. A systems analysis of importin-α–β mediated nuclear protein import. Journal of Cell Biology, 168(7):1027–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/7/1027.

Rossner:2007:HRG

[RM07] Mike Rossner and Ira Mellman. How the rich get richer. Journal of Cell Biology, 177(6):951–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/177/6/951.

Rivers:2008:PPD

[RMAA08] David M. Rivers, Sergio Moreno, Mary Abraham, and Julie Ahringer. PAR proteins direct asymmetry of the cell cycle regulators Polo-like kinase and Cdc25. Journal of Cell Biology, 180 (5):877–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/5/877.

Rodier:2005:PIG

[RMC^{+}05] Geneviève Rodier, Constantin Makris, Philippe Coulombe, Anthony Scime, Keiko Nakayama, Keiichi I. Nakayama, and Sylvain Meloche. p107 inhibits G1 to S phase progression by downregulating expression of the F-box protein Skp2. Journal of Cell Biology, 168(1):55–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/1/55.

Royou:2008:GCP

[RMKS08] Anne Royou, Derek McCusker, Douglas R. Kellogg, and William Sullivan. Grapes(Chk1) prevents nuclear CDK1 activation by delaying cyclin B nuclear accumulation. Journal of Cell Biology, 183(1):63–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/183/1/63.

Rudolf:2006:DVM

[RMP06] Rüdiger Rudolf, Paulo J. Magalhães, and Tullio Pozzan. Direct in vivo monitoring of sarcoplasmic reticulum Ca^{2+} and cytosolic cAMP dynamics in mouse skeletal muscle. Journal of Cell Biology, 173(2):187–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/2/187.

Risson:2009:MIM

[RMR^{+}09] Valérie Risson, Laetitia Mazelin, Mila Roceri, Hervé Sanchez, Vincent Moncollin, Claudine Corneloup, Hélène Richard-Bulteau, Alban Vignaud, Dominique Baas, Aurélia Defour, Damien Freyssenet, Jean-François Tanti, Yannick Le-Marchand-Brustel, Bernard Ferrier, Agnès Conjard-Duplany, Klaas Romanino, Stéphanie Bauché, Daniel Hantaï, Matthias Mueller, Sara C. Kozma, George Thomas, Markus A. Rüegg, Arnaud Ferry, Mario Pende, Xavier Bigard, Nathalie Koulmann, Laurent Schaeffer, and Yann-Gaël Gangloff. Muscle inactivation of mTOR causes metabolic and dystrophin defects leading to severe myopathy. Journal of Cell Biology, 187(6): 859–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/187/6/859.

Roll-Mecak:2006:MMM

[RMV06] Antonina Roll-Mecak and Ronald D. Vale. Making more microtubules by severing: a common theme of noncentrosomal microtubule arrays? Journal of Cell Biology, 175(6):849–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/175/6/
849.

Relaix:2006:PPD

[RMZ^{+}06] Frédéric Relaix, Didier Montarras, Stéphane Zaffran, Barbara Gayraud-Morel, Didier Rocancourt, Shahragim Tajbakhsh, Ahmed Mansouri, Ana Cumano, and Margaret Buckingham. Pax3 and Pax7 have distinct and overlapping functions in adult muscle progenitor cells. Journal of Cell Biology, 172(1):91–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
172/1/91.

Relaix:2007:CP

[RMZ^{+}07] Frédéric Relaix, Didier Montarras, Stéphane Zaffran, Barbara Gayraud-Morel, Didier Rocancourt, Shahragim Tajbakhsh, Ahmed Mansouri, Ana Cumano, and Margaret Buckingham. Correction 172, p. 91. Journal of Cell Biology, 176(1):125–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
176/1/125.

Robinson:2007:BIB

[Rob07a] Richard Robinson. Building the immunological bull’s eye. Journal of Cell Biology, 179(7):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/7/1329.2.

Robinson:2007:CWC

[Rob07b] Richard Robinson. Cells within cells (within cells). Journal of Cell Biology, 179(7):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/7/1328.1.

Robinson:2007:CPS

[Rob07c] Richard Robinson. A chewing proteasome is stabilized. Journal of Cell Biology, 179(6):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/6/1086.2.

Robinson:2007:EML

[Rob07d] Richard Robinson. Enzyme may link inflammation to new cell fate. Journal of Cell Biology, 179(2):??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/2/172.2.

Robinson:2007:FWR

[Rob07e] Richard Robinson. Forward walk with a random twist. Journal of Cell Biology, 179(2):??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/2/173.1.

Robinson:2007:HIN

[Rob07f] Richard Robinson. Hand it to the nucleolus. Journal of Cell Biology, 179(2):??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/2/173.2.

Robinson:2007:LAP

[Rob07g] Richard Robinson. Lipid + Alzheimer’s plaque = problem. Journal of Cell Biology, 179(7):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/7/1329.3.

Robinson:2007:MSM

[Rob07h] Richard Robinson. Minor spliceosome, major surprise: it’s cytoplasmic. Journal of Cell Biology, 179(6):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/6/
1086.1.

Robinson:2007:MMW

[Rob07i] Richard Robinson. Misfolding in muscle when neuron misfires. Journal of Cell Biology, 179(6):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/6/1087.1.

Robinson:2007:MRS

[Rob07j] Richard Robinson. Myosin remains strong as muscle contracts. Journal of Cell Biology, 179(6):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/6/1087.2.

Robinson:2007:ND

[Rob07k] Richard Robinson. The NPC, in detail. Journal of Cell Biology, 179(7):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/7/1328.2.

Robinson:2007:NRA

[Rob07l] Richard Robinson. Nucleus runs ahead of centrosome. Journal of Cell Biology, 179(2):??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/2/173.3.

Robinson:2007:RBC

[Rob07m] Richard Robinson. Red blood cells have a killer touch. Journal of Cell Biology, 179(2):??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/2/172.1.

Robinson:2007:TGA

[Rob07n] Richard Robinson. tRNA’s Los is a gain after damage. Journal of Cell Biology, 179(7):??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/7/1329.1.

Robinson:2008:AIP

[Rob08a] Richard Robinson. Abundance of ITAM prevents autoimmunity. Journal of Cell Biology, 181(5):??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/181/5/715.2.

Robinson:2008:AMM

[Rob08b] Richard Robinson. Actin and myosin in the mitotic spindle. Journal of Cell Biology, 182(1):??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/182/1/3.2.

Robinson:2008:ACO

[Rob08c] Richard Robinson. Aquaporin clears its own path. Journal of Cell Biology, 182(3):??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/3/412.3.

Robinson:2008:ANS

[Rob08d] Richard Robinson. Atherosclerosis needs shc. Journal of Cell Biology, 182(1):??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/1/2.1.

Robinson:2008:AID

[Rob08e] Richard Robinson. Autophagy inhibitors deliver a knock-out blow. Journal of Cell Biology, 183(1):??, October 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/183/1/2.3.

Robinson:2008:BB

[Rob08f] Richard Robinson. BBB building. Journal of Cell Biology, 183 (3):??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/3/366.3.

Robinson:2008:BWS

[Rob08g] Richard Robinson. A better way to see splicing partners. Journal of Cell Biology, 181(6):??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/181/6/875.2.

Robinson:2008:CCL

[Rob08h] Richard Robinson. Centromeres cross over, a lot. Journal of Cell Biology, 181(6):??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/6/874.2.

Robinson:2008:CRC

[Rob08i] Richard Robinson. Complementary receptors control RhoA. Journal of Cell Biology, 181(6):??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/181/6/874.1.

Robinson:2008:EPB

[Rob08j] Richard Robinson. Egg P bodies protect maternal mRNA. Journal of Cell Biology, 182(3):??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/182/3/413.2.

Robinson:2008:EET

[Rob08k] Richard Robinson. Embryos experience tension. Journal of Cell Biology, 181(2):??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/2/176.1.

Robinson:2008:ENM

[Rob08l] Richard Robinson. Embryos need Mom’s nucleolus. Journal of Cell Biology, 180(4):??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/4/650.1.

Robinson:2008:FPT

[Rob08m] Richard Robinson. Four proteins, two new cells. Journal of Cell Biology, 181(5):??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/5/714.2.

Robinson:2008:GFT

[Rob08n] Richard Robinson. GABA fine-tunes the glycine receptor. Journal of Cell Biology, 180(6):??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/6/1054.1.

Robinson:2008:HTT

[Rob08o] Richard Robinson. H^{+}, the tiniest transmitter. Journal of Cell Biology, 180(2):??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/2/251.3.

Robinson:2008:HCR

[Rob08p] Richard Robinson. Histone chaperone regulates replication. Journal of Cell Biology, 180(2):??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/180/2/250.1.

Robinson:2008:HGS

[Rob08q] Richard Robinson. How the ER gets its shape. Journal of Cell Biology, 180(6):??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/6/1055.2.

Robinson:2008:IYB

[Rob08r] Richard Robinson. If you build it, leukocytes will come (in). Journal of Cell Biology, 183(3):??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/183/3/367.1.

Robinson:2008:KPP

[Rob08s] Richard Robinson. Kinesin’s parts pull together. Journal of Cell Biology, 180(4):??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/4/651.3.

Robinson:2008:LMW

[Rob08t] Richard Robinson. Learn more when your neighbors do. Journal of Cell Biology, 180(2):??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/2/251.1.

Robinson:2008:LNS

[Rob08u] Richard Robinson. Lipid is new Star in nuclear regulation. Journal of Cell Biology, 180(6):??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/6/1055.3.

Robinson:2008:MFA

[Rob08v] Richard Robinson. Microtubule failure assists cleavage success. Journal of Cell Biology, 183(3):??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/183/3/367.2.

Robinson:2008:MSC

[Rob08w] Richard Robinson. Mitochondria split with a calcium hit. Journal of Cell Biology, 182(3):??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/182/3/413.1.

Robinson:2008:MC

[Rob08x] Richard Robinson. The mitochondrial connection. Journal of Cell Biology, 181(6):??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/6/875.1.

Robinson:2008:MBG

[Rob08y] Richard Robinson. Motors bring genes together. Journal of Cell Biology, 181(2):??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/2/177.3.

Robinson:2008:MCL

[Rob08z] Richard Robinson. MukBs cooperate to loop DNA. Journal of Cell Biology, 181(2):??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/2/177.2.

Robinson:2008:NEC

[Rob08-27] Richard Robinson. New exit: caspase-1 for secretion. Journal of Cell Biology, 180(6):??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/6/1055.1.

Robinson:2008:NDJ

[Rob08-28] Richard Robinson. No direction? Join the crowd. Journal of Cell Biology, 183(3):??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/183/3/366.1.

Robinson:2008:NSS

[Rob08-29] Richard Robinson. No stretch for spindle assembly checkpoint. Journal of Cell Biology, 183(1):??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/183/1/2.2.

Robinson:2008:NWN

[Rob08-30] Richard Robinson. No waiting at the nuclear pore complex. Journal of Cell Biology, 183(1):??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/183/1/2.1.

Robinson:2008:NMD

[Rob08-31] Richard Robinson. Nucleophosmin mutation is double trouble. Journal of Cell Biology, 182(1):??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/182/1/3.1.

Robinson:2008:OP

[Rob08-32] Richard Robinson. Organelles in parallel. Journal of Cell Biology, 180(2):??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/2/251.2.

Robinson:2008:PLG

[Rob08-33] Richard Robinson. Pack loose, grow long. Journal of Cell Biology, 181(2):??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/2/177.1.

Robinson:2008:PI

[Rob08-34] Richard Robinson. Partner up to invade. Journal of Cell Biology, 181(6):??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/6/875.3.

Robinson:2008:PMP

[Rob08-35] Richard Robinson. Pegging mitochondrial position. Journal of Cell Biology, 182(1):??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/1/2.3.

Robinson:2008:QFO

[Rob08-36] Richard Robinson. Quick flip = one-way proton trip. Journal of Cell Biology, 181(4):??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/4/571.2.

Robinson:2008:RFC

[Rob08-37] Richard Robinson. Random, fast, and out of control. Journal of Cell Biology, 183(1):??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/183/1/3.2.

Robinson:2008:RSS

[Rob08-38] Richard Robinson. rDNA silencing saves starving cells. Journal of Cell Biology, 181(5):??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/5/714.1.

Robinson:2008:RGS

[Rob08-39] Richard Robinson. Reach out and gRab something. Journal of Cell Biology, 180(4):??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/4/651.1.

Robinson:2008:RKA

[Rob08-40] Richard Robinson. Robo keeps axons moving. Journal of Cell Biology, 181(5):??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/5/715.3.

Robinson:2008:SLP

[Rob08-41] Richard Robinson. Sarcospan, a little protein for a big problem. Journal of Cell Biology, 183(3):??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/183/3/366.2.

Robinson:2008:SAC

[Rob08-42] Richard Robinson. Self-assembly of the cell wall. Journal of Cell Biology, 180(4):??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/4/651.2.

Robinson:2008:SSE

[Rob08-43] Richard Robinson. Signaling specifically from the endosome. Journal of Cell Biology, 181(4):??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/181/4/571.1.

Robinson:2008:SFC

[Rob08-44] Richard Robinson. Simple forces, complex shapes. Journal of Cell Biology, 180(2):??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/2/250.2.

Robinson:2008:SKL

[Rob08-45] Richard Robinson. Sumo keeps lamin a in place. Journal of Cell Biology, 182(1):??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/1/2.2.

Robinson:2008:STE

[Rob08-46] Richard Robinson. Sweeter tumors with EGFR. Journal of Cell Biology, 181(5):??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/5/715.1.

Robinson:2008:CCT

[Rob08-47] Richard Robinson. T cells count and then decide. Journal of Cell Biology, 180(4):??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/4/650.2.

Robinson:2008:TBG

[Rob08-48] Richard Robinson. Take a big gulp of pox. Journal of Cell Biology, 181(4):??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/4/571.3.

Robinson:2008:FFU

[Rob08-49] Richard Robinson. To fold, first unfold. Journal of Cell Biology, 180(6):??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/6/1054.2.

Robinson:2008:TLT

[Rob08-50] Richard Robinson. Too long for translocation? Journal of Cell Biology, 181(4):??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/4/570.2.

Robinson:2008:TSD

[Rob08-51] Richard Robinson. Transcribing “silent” DNA resilences it. Journal of Cell Biology, 181(2):??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/181/2/176.2.

Robinson:2008:TSW

[Rob08-52] Richard Robinson. Tre1 says when, and where, to go. Journal of Cell Biology, 183(1):??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/183/1/3.1.

Robinson:2008:TST

[Rob08-53] Richard Robinson. TRPV4 shows TRPP2 the moves. Journal of Cell Biology, 182(3):??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/182/3/412.1.

Robinson:2008:TCS

[Rob08-54] Richard Robinson. Tumor cells share oncogenic receptors. Journal of Cell Biology, 181(4):??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/181/4/570.1.

Robinson:2008:TMM

[Rob08-55] Richard Robinson. Two microtubules means one-way movement. Journal of Cell Biology, 182(3):??, August 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/182/3/412.2.

Reber:2008:CKI

[ROKG08] Simone Reber, Sabine Over, Iva Kronja, and Oliver J. Gruss. CaM kinase II initiates meiotic spindle depolymerization independently of APC/C activation. Journal of Cell Biology, 183(6):1007–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/6/1007.

Rossner:2005:NPE

[Ros05] Mike Rossner. The NIH policy on enhancing public access to publications resulting from NIH-funded research. Journal of Cell Biology, 168(7):991–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/168/7/991.

Rossner:2007:BTP

[Ros07a] Mike Rossner. Bending the Truth with a PRISM. Journal of Cell Biology, 179(2):177–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/2/177.

Rossner:2007:HCR

[Ros07b] Mike Rossner. Hwang case review committee misses the mark. Journal of Cell Biology, 176(2):131–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/176/2/131.

Rossner:2008:FSSa

[Ros08a] Mike Rossner. A false sense of security. Journal of Cell Biology, 183(4):573–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/4/573.

Rossner:2008:FSSb

[Ros08b] Mike Rossner. A false sense of security. Journal of Cell Biology, 183(7):1317–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/7/1317.

Rossner:2008:HJG

[Ros08c] Mike Rossner. Heather Joseph: Getting the message across. Journal of Cell Biology, 183(3):368–??, November 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/183/3/368.

Rossner:2009:CG

[Ros09] Mike Rossner. A challenge to Goliath. Journal of Cell Biology, 185(4):571–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/4/571.

Rout:2008:PPF

[Rou08] Michael P. Rout. The peroxisome: a production in four acts. Journal of Cell Biology, 181(2):185–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/181/2/185.

Ryan:2005:AGP

[RP05] Kathleen A. Ryan and Sanjay W. Pimplikar. Activation of GSK-3 and phosphorylation of CRMP2 in transgenic mice expressing APP intracellular domain. Journal of Cell Biology, 171 (2):327–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/2/327.

Rusan:2007:RNC

[RP07] Nasser M. Rusan and Mark Peifer. A role for a novel centrosome cycle in asymmetric cell division. Journal of Cell Biology, 177(1):13–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/1/13.

Rusan:2008:OCR

[RP08] Nasser M. Rusan and Mark Peifer. Original CIN: reviewing roles for APC in chromosome instability. Journal of Cell Biology, 181(5):719–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/5/719.

Rucci:2009:GBD

[RRA^{+}09] Nadia Rucci, Anna Rufo, Marina Alamanou, Mattia Capulli, Andrea Del Fattore, Emma Åhrman, Daria Capece, Valeria Iansante, Francesca Zazzeroni, Edoardo Alesse, Dick Heinegård, and Anna Teti. The glycosaminoglycan-binding domain of PRELP acts as a cell type–specific NF-κB inhibitor that impairs osteoclastogenesis. Journal of Cell Biology, 187(5):669– ??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/187/5/669.

Roxrud:2008:ELI

[RRP^{+}08] Ingrid Roxrud, Camilla Raiborg, Nina Marie Pedersen, Espen Stang, and Harald Stenmark. An endosomally localized isoform of Eps15 interacts with Hrs to mediate degradation of epidermal growth factor receptor. Journal of Cell Biology, 180(6):1205–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
180/6/1205.

Rogers:2009:SUL

[RRR^{+}09] Gregory C. Rogers, Nasser M. Rusan, David M. Roberts, Mark Peifer, and Stephen L. Rogers. The SCF^{Slimb} ubiquitin ligase regulates Plk4/Sak levels to block centriole reduplication. Journal of Cell Biology, 184(2):225–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/184/2/225.

Reither:2006:PVK

[RSL06] Gregor Reither, Michael Schaefer, and Peter Lipp. PKCα: a versatile key for decoding the cellular calcium toolkit. Journal of Cell Biology, 174(4):521–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/174/4/521.

Ritchie:2008:LAL

[RSM^{+}08] Kieran Ritchie, Claudia Seah, Jana Moulin, Christian Isaac, Frederick Dick, and Nathalie G. Bérubé. Loss of ATRX leads to chromosome cohesion and congression defects. Journal of Cell Biology, 180(2):315–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/2/315.

Rabl:2009:FCC

[RSS^{+}09] Regina Rabl, Vincent Soubannier, Roland Scholz, Frank Vogel, Nadine Mendl, Andreja Vasiljev-Neumeyer, Christian Körner, Ravi Jagasia, Thomas Keil, Wolfgang Baumeister, Marek Cyrklaff, Walter Neupert, and Andreas S. Reichert. Formation of cristae and crista junctions in mitochondria depends on antagonism between Fcj1 and Su e /g. Journal of Cell Biology, 185 (6):1047–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/6/1047.

Rodriguez:2006:VMT

[RSSC06] Alexis J. Rodriguez, Shailesh M. Shenoy, Robert H. Singer, and John Condeelis. Visualization of mRNA translation in living cells. Journal of Cell Biology, 175(1):67–??, October 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/175/1/67.

Ramirez-Valle:2008:ELN

[RVBZ^{+}08] Francisco Ramírez-Valle, Steve Braunstein, Jiri Zavadil, Silvia C. Formenti, and Robert J. Schneider. eIF4GI links nutrient sensing by mTOR to cell proliferation and inhibition of autophagy. Journal of Cell Biology, 181(2):293–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/181/2/293.

Rossner:2007:SMD

[RVH07] Mike Rossner, Heather Van Epps, and Emma Hill. Show me the data. Journal of Cell Biology, 179(6):1091–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/6/
1091.

Rossner:2008:IRR

[RVH08] Mike Rossner, Heather Van Epps, and Emma Hill. Irreproducible results: a response to Thomson Scientific. Journal of Cell Biology, 180(2):254–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/2/254.

Rojas:2008:RRR

[RvVM^{+}08] Raul Rojas, Thijs van Vlijmen, Gonzalo A. Mardones, Yogikala Prabhu, Adriana L. Rojas, Shabaz Mohammed, Albert J. R. Heck, Graça Raposo, Peter van der Sluijs, and Juan S. Bonifacino. Regulation of retromer recruitment to endosomes by sequential action of Rab5 and Rab7. Journal of Cell Biology, 183(3):513–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/3/513.

Rusan:2005:CFM

[RW05] Nasser M. Rusan and Patricia Wadsworth. Centrosome fragments and microtubules are transported asymmetrically away from division plane in anaphase. Journal of Cell Biology, 168 (1):21–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/1/21.

Robinson:2008:BBF

[RW08a] Richard Robinson and Ruth Williams. Better binding with a flick of the tail. Journal of Cell Biology, 182(5):??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/182/5/
818.2.

Robinson:2008:KRC

[RW08b] Richard Robinson and Ruth Williams. Kick out the reticulons to close the envelope. Journal of Cell Biology, 182 (5):??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/5/818.1.

Robinson:2008:SCN

[RW08c] Richard Robinson and Ruth Williams. SRC1, central to the nuclear periphery. Journal of Cell Biology, 182(5):??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/182/5/
819.2.

Robinson:2008:VRD

[RW08d] Richard Robinson and Ruth Williams. Vesicles roar, and don’t collapse. Journal of Cell Biology, 182(5):??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/182/5/818.
3.

Robinson:2008:WSG

[RW08e] Richard Robinson and Ruth Williams. A weak signaler gets in close. Journal of Cell Biology, 182(5):??, September 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/182/5/819.1.

Reim:2005:SFU

[RWB^{+}05] Kerstin Reim, Heike Wegmeyer, Johann Helmut Brandstätter, Mingshan Xue, Christian Rosenmund, Thomas Dresbach, Kay Hofmann, and Nils Brose. Structurally and functionally unique complexins at retinal ribbon synapses. Journal of Cell Biology, 169(4):669–??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/4/669.

Roumanie:2005:RGR

[RWM^{+}05] Olivier Roumanie, Hao Wu, Jeffrey N. Molk, Guendalina Rossi, Kerry Bloom, and Patrick Brennwald. Rho GTPase regulation of exocytosis in yeast is independent of GTP hydrolysis and polarization of the exocyst complex. Journal of Cell Biology, 170(4):583–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/4/583.

Rouiller:2008:SBA

[RXA^{+}08] Isabelle Rouiller, Xiao-Ping Xu, Kurt J. Amann, Coumaran Egile, Stephan Nickell, Daniela Nicastro, Rong Li, Thomas D. Pollard, Niels Volkmann, and Dorit Hanein. The structural basis of actin filament branching by the Arp2/3 complex. Journal of Cell Biology, 180(5):887–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/5/887.

Ren:2008:IRM

[RYD08] Hongxia Ren, Ping Yin, and Cunming Duan. IGFBP-5 regulates muscle cell differentiation by binding to IGF–II and switching on the IGF–II auto-regulation loop. Journal of Cell Biology, 182(5):979–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/182/5/979.

Reilein:2005:SOA

[RYN05] Amy Reilein, Soichiro Yamada, and W. James Nelson. Selforganization of an acentrosomal microtubule network at the basal cortex of polarized epithelial cells. Journal of Cell Biology, 171(5):845–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/5/845.

Stefanovic:2009:IOS

[SAD^{+}09] Sonia Stefanovic, Nesrine Abboud, Stéphanie Désilets, David Nury, Chad Cowan, and Michel Pucéat. Interplay of Oct4 with Sox2 and Sox17: a molecular switch from stem cell pluripotency to specifying a cardiac fate. Journal of Cell Biology, 186(5): 665–??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/186/5/665.

Salzer:2008:SM

[Sal08] James L. Salzer. Switching myelination on and off. Journal of Cell Biology, 181(4):575–??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/181/4/575.

Shnyrova:2007:VFS

[SAM^{+}07] Anna V. Shnyrova, Juan Ayllon, Ilya I. Mikhalyov, Enrique Villar, Joshua Zimmerberg, and Vadim A. Frolov. Vesicle formation by self-assembly of membrane-bound matrix proteins into a fluidlike budding domain. Journal of Cell Biology, 179(4): 627–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/4/627.

Savage:2009:BKF

[Sav09a] Liz Savage. Bryan Krantz: From folding to unfolding proteins. Journal of Cell Biology, 184(5):618–??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/184/5/618.

Savage:2009:VMG

[Sav09b] Liz Savage. Vivek Malhotra: Gaga for the Golgi. Journal of Cell Biology, 184(1):4–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/184/1/4.

Seto:2006:IRP

[SB06] Elaine S. Seto and Hugo J. Bellen. Internalization is required for proper Wingless signaling in Drosophila melanogaster. Journal of Cell Biology, 173(1):95–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/173/1/95.

Schmidt:2009:LAA

[SB09] John A. Schmidt and William J. Brown. Lysophosphatidic acid acyltransferase 3 regulates Golgi complex structure and function. Journal of Cell Biology, 186(2):211–??, July 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/186/2/211.

Salm:2005:TMD

[SBC^{+}05] Sarah N. Salm, Patricia E. Burger, Sandra Coetzee, Ken Goto, David Moscatelli, and E. Lynette Wilson. TGF-β maintains dormancy of prostatic stem cells in the proximal region of ducts. Journal of Cell Biology, 170(1):81–??, July 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/170/1/81.

Schweizer:2006:NCN

[SBC^{+}06] Jürgen Schweizer, Paul E. Bowden, Pierre A. Coulombe, Lutz Langbein, E. Birgitte Lane, Thomas M. Magin, Lois Maltais, M. Bishr Omary, David A. D. Parry, Michael A. Rogers, and Mathew W. Wright. New consensus nomenclature for mammalian keratins. Journal of Cell Biology, 174(2):169–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/174/2/
169.

Suraweera:2007:SDA

[SBC^{+}07] Amila Suraweera, Olivier J. Becherel, Philip Chen, Natalie Rundle, Rick Woods, Jun Nakamura, Magtouf Gatei, Chiara Criscuolo, Alessandro Filla, Luciana Chessa, Markus Fußer, Bernd Epe, Nuri Gueven, and Martin F. Lavin. Senataxin, defective in ataxia oculomotor apraxia type 2, is involved in the defense against oxidative DNA damage. Journal of Cell Biology, 177(6):969–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/6/969.

Saunders:2006:CVT

[SBF^{+}06] W. Brian Saunders, Brenda L. Bohnsack, Jennifer B. Faske, Nicholas J. Anthis, Kayla J. Bayless, Karen K. Hirschi, and George E. Davis. Coregulation of vascular tube stabilization by endothelial cell TIMP-2 and pericyte TIMP-3. Journal of Cell Biology, 175(1):179–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/1/179.

Smith:2007:DES

[SBJ^{+}07] Kelly P. Smith, Meg Byron, Carol Johnson, Yigong Xing, and Jeanne B. Lawrence. Defining early steps in mRNA transport: mutant mRNA in myotonic dystrophy type I is blocked at entry into SC-35 domains. Journal of Cell Biology, 178(6): 951–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/6/951.

Sachs:2007:PNR

[SBM^{+}07] Benjamin D. Sachs, George S. Baillie, Julianne R. McCall, Melissa A. Passino, Christian Schachtrup, Derek A. Wallace, Allan J. Dunlop, Kirsty F. MacKenzie, Enno Klussmann, Martin J. Lynch, Shoana L. Sikorski, Tal Nuriel, Igor Tsigelny, Jin Zhang, Miles D. Houslay, Moses V. Chao, and Katerina Akassoglou. p75 neurotrophin receptor regulates tissue fibrosis through inhibition of plasminogen activation via a PDE4/ cAMP/ PKA pathway. Journal of Cell Biology, 177(6):1119–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
177/6/1119.

Sirotkin:2005:IWM

[SBMP05] Vladimir Sirotkin, Christopher C. Beltzner, Jean-Baptiste Marchand, and Thomas D. Pollard. Interactions of WASp, myosin–I, and verprolin with Arp2/3 complex during actin patch assembly in fission yeast. Journal of Cell Biology, 170 (4):637–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/4/637.

Szabadkai:2006:CMC

[SBV^{+}06] György Szabadkai, Katiuscia Bianchi, Péter Várnai, Diego De Stefani, Mariusz R. Wieckowski, Dario Cavagna, Anikó I. Nagy, Tamás Balla, and Rosario Rizzuto. Chaperone-mediated coupling of endoplasmic reticulum and mitochondrial Ca^{2+} channels. Journal of Cell Biology, 175(6):901–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/175/6/901.

Sharma:2007:KRD

[SBW^{+}07] Neeraj Sharma, Jessica Bryant, Dorota Wloga, Rachel Donaldson, Richard C. Davis, Maria Jerka-Dziadosz, and Jacek Gaertig. Katanin regulates dynamics of microtubules and biogenesis of motile cilia. Journal of Cell Biology, 178(6):1065–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/178/6/
1065.

Seki:2008:PTD

[SCD^{+}08] Akiko Seki, Judith A. Coppinger, Haining Du, Chang-Young Jang, John R. Yates, and Guowei Fang. Plk1- and β-TrCP-dependent degradation of Bora controls mitotic progression. Journal of Cell Biology, 181(1):65–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/181/1/65.

Song:2007:OPC

[SCF^{+}07] Zhiyin Song, Hsiuchen Chen, Maja Fiket, Christiane Alexander, and David C. Chan. OPA1 processing controls mitochondrial fusion and is regulated by mRNA splicing, membrane potential, and Yme1L. Journal of Cell Biology, 178(5):749–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/178/5/
749.

Siegl-Cachedenier:2007:TRE

[SCFKB07] Irene Siegl-Cachedenier, Ignacio Flores, Peter Klatt, and Maria A. Blasco. Telomerase reverses epidermal hair follicle stem cell defects and loss of long-term survival associated with critically short telomeres. Journal of Cell Biology, 179 (2):277–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/2/277.

Scholey:2008:ITM

[Sch08] Jonathan M. Scholey. Intraflagellar transport motors in cilia: moving along the cell’s antenna. Journal of Cell Biology, 180 (1):23–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/1/23.

Sun:2007:PPT

[SCK^{+}07] Yidi Sun, Susheela Carroll, Marko Kaksonen, Junko Y. Toshima, and David G. Drubin. PtdIns(4,5)P_{2} turnover is required for multiple stages during clathrin- and actin-dependent endocytic internalization. Journal of Cell Biology, 177(2): 355–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/2/355.

Shah:2007:ARR

[SCM^{+}07] Ankur H. Shah, Nicholas L. Cianciola, Jeffrey L. Mills, Frank D. Sönnichsen, and Cathleen Carlin. Adenovirus RIDα regulates endosome maturation by mimicking GTP–Rab7. Journal of Cell Biology, 179(5):965–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/5/965.

Stehman:2007:NNR

[SCMV07] Stephanie A. Stehman, Yu Chen, Richard J. McKenney, and Richard B. Vallee. NudE and NudEL are required for mitotic progression and are involved in dynein recruitment to kinetochores. Journal of Cell Biology, 178(4):583–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/178/4/583.

Scorrano:2008:CGC

[Sco08] Luca Scorrano. Caspase-8 goes cardiolipin: a new platform to provide mitochondria with microdomains of apoptotic signals? Journal of Cell Biology, 183(4):579–??, November 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/183/4/579.

Schulz:2009:LRS

[SCR^{+}09] Timothy A. Schulz, Mal-Gi Choi, Sumana Raychaudhuri, Jason A. Mears, Rodolfo Ghirlando, Jenny E. Hinshaw, and William A. Prinz. Lipid-regulated sterol transfer between closely apposed membranes by oxysterol-binding protein homologues. Journal of Cell Biology, 187(6):889–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/187/6/889.

Smith:2005:TDL

[SCS^{+}05] Andrew Smith, Yolanda R. Carrasco, Paula Stanley, Nelly Kieffer, Facundo D. Batista, and Nancy Hogg. A talin-dependent LFA-1 focal zone is formed by rapidly migrating T lymphocytes. Journal of Cell Biology, 170(1):141–??, July 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/170/1/141.

Scott:2009:NEF

[SCVW09] Robert J. Scott, Lucas V. Cairo, David W. Van de Vosse, and Richard W. Wozniak. The nuclear export factor Xpo1p targets Mad1p to kinetochores in yeast. Journal of Cell Biology, 184 (1):21–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/184/1/21.

Sindelar:2007:BKF

[SD07] Charles V. Sindelar and Kenneth H. Downing. The beginning of kinesin’s force-generating cycle visualized at 9-Å resolution. Journal of Cell Biology, 177(3):377–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/177/3/377.

Scott:2005:PBP

[SDB^{+}05] Cameron C. Scott, Wendy Dobson, Roberto J. Botelho, Natasha Coady-Osberg, Philippe Chavrier, David A. Knecht, Colin Heath, Philip Stahl, and Sergio Grinstein. Phosphatidylinositol-4,5- bis phosphate hydrolysis directs actin remodeling during phagocytosis. Journal of Cell Biology, 169(1):139–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/169/1/
139.

So:2009:ASS

[SDC09] Sairei So, Anthony J. Davis, and David J. Chen. Autophosphorylation at serine 1981 stabilizes ATM at DNA damage sites. Journal of Cell Biology, 187(7):977–??, December 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/187/7/977.

Shvartsman:2007:SKA

[SDD^{+}07] Dmitry E. Shvartsman, John C. Donaldson, Begoña Diaz, Orit Gutman, G. Steven Martin, and Yoav I. Henis. Src kinase activity and SH2 domain regulate the dynamics of Src association with lipid and protein targets. Journal of Cell Biology, 178 (4):675–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/4/675.

Sauvonnet:2005:CDR

[SDDV05] Nathalie Sauvonnet, Annick Dujeancourt, and Alice Dautry-Varsat. Cortactin and dynamin are required for the clathrinindependent endocytosis of γ c cytokine receptor. Journal of Cell Biology, 168(1):155–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/168/1/155.

Sesaki:2006:UCI

[SDI^{+}06] Hiromi Sesaki, Cory D. Dunn, Miho Iijima, Kelly A. Shepard, Michael P. Yaffe, Carolyn E. Machamer, and Robert E. Jensen. Ups1p, a conserved intermembrane space protein, regulates mitochondrial shape and alternative topogenesis of Mgm1p. Journal of Cell Biology, 173(5):651–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/173/5/651.

Sacco:2005:IIB

[SDL^{+}05] Alessandra Sacco, Regis Doyonnas, Mark A. LaBarge, Mark M. Hammer, Peggy Kraft, and Helen M. Blau. IGF-i increases bone marrow contribution to adult skeletal muscle and enhances the fusion of myelomonocytic precursors. Journal of Cell Biology, 171(3):483–??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/3/483.

Short:2005:C

[SDN^{+}05a] Sarah M. Short, Alexandrine Derrien, Radha P. Narsimhan, Jack Lawler, Donald E. Ingber, and Bruce R. Zetter. Correction. Journal of Cell Biology, 169(3):541–??, May 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/169/3/541.

Short:2005:IEC

[SDN^{+}05b] Sarah M. Short, Alexandrine Derrien, Radha P. Narsimhan, Jack Lawler, Donald E. Ingber, and Bruce R. Zetter. Inhibition of endothelial cell migration by thrombospondin-1 type-1 repeats is mediated by β_{1} integrins. Journal of Cell Biology, 168(4):643–??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/4/643.

Stuart:2005:RSA

[SDS^{+}05] Lynda M. Stuart, Jiusheng Deng, Jessica M. Silver, Kazue Takahashi, Anita A. Tseng, Elizabeth J. Hennessy, R. Alan B. Ezekowitz, and Kathryn J. Moore. Response to Staphylococcus aureus requires CD36-mediated phagocytosis triggered by the COOH-terminal cytoplasmic domain. Journal of Cell Biology, 170(3):477–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/3/477.

Stevens:2009:ACP

[SDW^{+}09] Naomi R. Stevens, Jeroen Dobbelaere, Alan Wainman, Fanni Gergely, and Jordan W. Raff. Ana3 is a conserved protein required for the structural integrity of centrioles and basal bodies. Journal of Cell Biology, 187(3):355–??, November 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/187/3/355.

Sedwick:2007:CA

[Sed07a] Caitlin Sedwick. Cholesterol and Alzheimer’s. Journal of Cell Biology, 179(3):??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/3/363.3.

Sedwick:2007:EMS

[Sed07b] Caitlin Sedwick. Elizabeth Miller: Sleuthing the details of the secretory pathway. Journal of Cell Biology, 179(4):572– ??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/4/572.

Sedwick:2007:IIR

[Sed07c] Caitlin Sedwick. Imaging the immune response. Journal of Cell Biology, 179(3):??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/3/362.1.

Sedwick:2007:LMN

[Sed07d] Caitlin Sedwick. Localized mRNA is the norm. Journal of Cell Biology, 179(3):??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/3/363.2.

Sedwick:2007:SWC

[Sed07e] Caitlin Sedwick. Stefan Westermann: a close look at kinetochore function. Journal of Cell Biology, 179(5):810–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/5/
810.

Sedwick:2007:SMS

[Sed07f] Caitlin Sedwick. Structure of the meiotic spindle. Journal of Cell Biology, 179(3):??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/3/363.1.

Sedwick:2007:TAI

[Sed07g] Caitlin Sedwick. Tumors’ accomplices in invasion. Journal of Cell Biology, 179(3):??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/3/362.2.

Sedwick:2008:AFB

[Sed08a] Caitlin Sedwick. Alison Frand: Breaking out new ideas on molting. Journal of Cell Biology, 181(6):876–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/181/6/876.

Sedwick:2008:AKA

[Sed08b] Caitlin Sedwick. Amy Kiger: Abuzz on cell biology. Journal of Cell Biology, 182(1):4–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/1/4.

Sedwick:2008:DCL

[Sed08c] Caitlin Sedwick. Dictyostelium cells lay a breadcrumb trail. Journal of Cell Biology, 183(5):??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/183/5/752.3.

Sedwick:2008:CSL

[Sed08d] Caitlin Sedwick. Claudin 11 stops the leaks. Journal of Cell Biology, 183(5):??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/5/752.1.

Sedwick:2008:CAT

[Sed08e] Caitlin Sedwick. Cofilin activity is a total coincidence. Journal of Cell Biology, 183(5):??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/183/5/753.1.

Sedwick:2008:DGA

[Sed08f] Caitlin Sedwick. Dynein gives the “all clear”. Journal of Cell Biology, 183(5):??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/5/753.2.

Sedwick:2008:EMG

[Sed08g] Caitlin Sedwick. Eran Meshorer: getting a chromatin perspective. Journal of Cell Biology, 182(4):618–??, August 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/182/4/618.

Sedwick:2008:GVS

[Sed08h] Caitlin Sedwick. Gia Voeltz: Shaping ideas about ER shape. Journal of Cell Biology, 180(1):4–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/180/1/4.

Sedwick:2008:JTL

[Sed08i] Caitlin Sedwick. Jussi Taipale: For the love of data. Journal of Cell Biology, 182(2):218–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/182/2/218.

Sedwick:2008:KSM

[Sed08j] Caitlin Sedwick. Kai Simons: Membrane master. Journal of Cell Biology, 183(7):1180–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/183/7/1180.

Sedwick:2008:LTM

[Sed08k] Caitlin Sedwick. Lloyd Trotman: Of mice and men, cancer, and PTEN. Journal of Cell Biology, 181(3):402–??, May 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/181/3/402.

Sedwick:2008:PCH

[Sed08l] Caitlin Sedwick. Parkin cleans house. Journal of Cell Biology, 183(5):??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/5/752.2.

Sedwick:2008:QZS

[Sed08m] Caitlin Sedwick. Qing Zhong: Scoring a slam dunk on the autophagy court. Journal of Cell Biology, 183(2):174–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/183/2/
174.

Sedwick:2008:RHP

[Sed08n] Caitlin Sedwick. Rebecca Heald: Passionate about cycles. Journal of Cell Biology, 183(4):570–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/183/4/570.

Sedwick:2009:CLP

[Sed09a] Caitlin Sedwick. CERT loss puts the brakes on growth. Journal of Cell Biology, 184(1):??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/184/1/3.2.

Sedwick:2009:FMH

[Sed09b] Caitlin Sedwick. Frauke Melchior: How SUMO wrestles other proteins. Journal of Cell Biology, 187(5):586–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/187/5/
586.

Sedwick:2009:HGS

[Sed09c] Caitlin Sedwick. Htm1p: getting a “sugar-handle” on misfolded proteins. Journal of Cell Biology, 184(1):??, January 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/184/1/2.3.

Sedwick:2009:ITT

[Sed09d] Caitlin Sedwick. It takes two to party with chromatin. Journal of Cell Biology, 184(1):??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/184/1/2.2.

Sedwick:2009:JNK

[Sed09e] Caitlin Sedwick. Jodi Nunnari: Keeping an eye on mitochondrial inheritance. Journal of Cell Biology, 184(2):188–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
184/2/188.

Sedwick:2009:LCD

[Sed09f] Caitlin Sedwick. Lamin A/C deficiency is unnerving. Journal of Cell Biology, 184(1):??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/184/1/2.1.

Sedwick:2009:LBL

[Sed09g] Caitlin Sedwick. Lamin B locks up Oct-1. Journal of Cell Biology, 184(1):??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/1/3.1.

Sedwick:2009:MKT

[Sed09h] Caitlin Sedwick. Marko Kaksonen: Tracking the inner secrets of endocytosis. Journal of Cell Biology, 184(6):768–??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/184/6/
768.

Sedwick:2009:MBT

[Sed09i] Caitlin Sedwick. Michael Buszczak: Tracking the big game in stem cell identity. Journal of Cell Biology, 187(4):446– ??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/187/4/446.

Sedwick:2009:PFP

[Sed09j] Caitlin Sedwick. Peter Friedl: Painting a picture of cancer immunology. Journal of Cell Biology, 185(2):180–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/185/2/180.

Sedwick:2009:SMN

[Sed09k] Caitlin Sedwick. Senthil Muthuswamy: a new road map for cancer. Journal of Cell Biology, 187(2):152–??, October 2009.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/187/2/152.

Sedwick:2009:SPS

[Sed09l] Caitlin Sedwick. Suzanne Pfeffer: Sorting through membrane trafficking. Journal of Cell Biology, 185(1):4–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/185/1/4.

Sedwick:2009:TKR

[Sed09m] Caitlin Sedwick. Tarun Kapoor: In the right position to study chromosomes. Journal of Cell Biology, 187(6):754–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/187/6/
754.

Sarkar:2005:LIA

[SFB^{+}05] Sovan Sarkar, R. Andres Floto, Zdenek Berger, Sara Imarisio, Axelle Cordenier, Matthieu Pasco, Lynnette J. Cook, and David C. Rubinsztein. Lithium induces autophagy by inhibiting inositol monophosphatase. Journal of Cell Biology, 170(7): 1101–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/7/1101.

Smith:2009:TTD

[SFB09] Kevin D. Smith, Michael A. Fu, and Eric J. Brown. Tim– tipin dysfunction creates an indispensible reliance on the ATR– Chk1 pathway for continued DNA synthesis. Journal of Cell Biology, 187(1):15–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/1/15.

Shintani:2008:CMR

[SFC^{+}08] Yasushi Shintani, Yuri Fukumoto, Nina Chaika, Robert Svoboda, Margaret J. Wheelock, and Keith R. Johnson. Collagen Imediated up-regulation of N-cadherin requires cooperative signals from integrins and discoidin domain receptor 1. Journal of Cell Biology, 180(6):1277–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/6/1277.

Suzuki:2007:DRP

[SFEK07] Kenichi G. N. Suzuki, Takahiro K. Fujiwara, Michael Edidin, and Akihiro Kusumi. Dynamic recruitment of phospholipase Cγ at transiently immobilized GPI-anchored receptor clusters induces IP_{3} –Ca^{2+} signaling: single-molecule tracking study 2. Journal of Cell Biology, 177(4):731–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/177/4/731.

Saelens:2005:PSP

[SFP^{+}05] Xavier Saelens, Nele Festjens, Eef Parthoens, Isabel Vanoverberghe, Michael Kalai, Frank van Kuppeveld, and Peter Vandenabeele. Protein synthesis persists during necrotic cell death. Journal of Cell Biology, 168(4):545–??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/168/4/545.

Suzuki:2007:GAR

[SFS^{+}07] Kenichi G. N. Suzuki, Takahiro K. Fujiwara, Fumiyuki Sanematsu, Ryota Iino, Michael Edidin, and Akihiro Kusumi. GPIanchored receptor clusters transiently recruit Lyn and Gα for temporary cluster immobilization and Lyn activation: singlemolecule tracking study 1. Journal of Cell Biology, 177(4):717– ??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
177/4/717.

Shin:2005:EMH

[SFY^{+}05] Ji-Yeon Shin, Zhi-Hui Fang, Zhao-Xue Yu, Chuan-En Wang, Shi-Hua Li, and Xiao-Jiang Li. Expression of mutant huntingtin in glial cells contributes to neuronal excitotoxicity. Journal of Cell Biology, 171(6):1001–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/171/6/1001.

Shin:2006:C

[SFY^{+}06] Ji-Yeon Shin, Zhi-Hui Fang, Zhao-Xue Yu, Chuan-En Wang, Shi-Hua Li, and Xiao-Jiang Li. Correction. Journal of Cell Biology, 172(6):953–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/6/953.

Shen:2009:APS

[SG09] Wei Shen and Barry Ganetzky. Autophagy promotes synapse development in Drosophila. Journal of Cell Biology, 187(1): 71–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/187/1/71.

Schermer:2006:HLT

[SGB^{+}06] Bernhard Schermer, Cristina Ghenoiu, Malte Bartram, Roman Ulrich Müller, Fruzsina Kotsis, Martin Höhne, Wolfgang Kühn, Manuela Rapka, Roland Nitschke, Hanswalter Zentgraf, Manfred Fliegauf, Heymut Omran, Gerd Walz, and Thomas Benzing. The von Hippel–Lindau tumor suppressor protein controls ciliogenesis by orienting microtubule growth. Journal of Cell Biology, 175(4):547–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/175/4/547.

Srsen:2006:ICP

[SGDM06] Vlastimil Srsen, Nicole Gnadt, Alexander Dammermann, and Andreas Merdes. Inhibition of centrosome protein assembly leads to p53-dependent exit from the cell cycle. Journal of Cell Biology, 174(5):625–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/5/625.

Shang:2007:TAP

[SGK^{+}07] Jie Shang, Ningguo Gao, Randal J. Kaufman, David Ron, Heather P. Harding, and Mark A. Lehrman. Translation attenuation by PERK balances ER glycoprotein synthesis with lipid-linked oligosaccharide flux. Journal of Cell Biology, 176 (5):605–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/5/605.

Sinka:2008:GCC

[SGKM08] Rita Sinka, Alison K. Gillingham, Vangelis Kondylis, and Sean Munro. Golgi coiled-coil proteins contain multiple binding sites for Rab family G proteins. Journal of Cell Biology, 183(4): 607–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/4/607.

Stojanovski:2007:AFMa

[SGKP^{+}07a] Diana Stojanovski, Bernard Guiard, Vera Kozjak-Pavlovic, Nikolaus Pfanner, and Chris Meisinger. Alternative function for the mitochondrial SAM complex in biogenesis of α-helical TOM proteins. Journal of Cell Biology, 179(5):881–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/5/
881.

Stojanovski:2007:AFMb

[SGKP^{+}07b] Diana Stojanovski, Bernard Guiard, Vera Kozjak-Pavlovic, Nikolaus Pfanner, and Chris Meisinger. Alternative function for the mitochondrial SAM complex in biogenesis of α-helical TOM proteins. Journal of Cell Biology, 179(7):1613–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/7/
1613.

Sahai:2007:SRT

[SGMPV07] Erik Sahai, Raquel Garcia-Medina, Jacques Pouysségur, and Emmanuel Vial. Smurf1 regulates tumor cell plasticity and motility through degradation of RhoA leading to localized inhibition of contractility. Journal of Cell Biology, 176(1):35–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
176/1/35.

Schneider:2005:QED

[SH05] Ian C. Schneider and Jason M. Haugh. Quantitative elucidation of a distinct spatial gradient-sensing mechanism in fibroblasts. Journal of Cell Biology, 171(5):883–??, December 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/171/5/883.

Smith:2007:NRG

[SHB^{+}07] Adam C. Smith, Won Do Heo, Virginie Braun, Xiuju Jiang, Chloe Macrae, James E. Casanova, Marci A. Scidmore, Sergio Grinstein, Tobias Meyer, and John H. Brumell. A network of Rab GTPases controls phagosome maturation and is modulated by Salmonella enterica serovar Typhimurium. Journal of Cell Biology, 176(3):263–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/3/263.

Shin:2005:ECR

[SHC^{+}05] Hye-Won Shin, Mitsuko Hayashi, Savvas Christoforidis, Sandra Lacas-Gervais, Sebastian Hoepfner, Markus R. Wenk, Jan Modregger, Sandrine Uttenweiler-Joseph, Matthias Wilm, Arne Nystuen, Wayne N. Frankel, Michele Solimena, Pietro De Camilli, and Marino Zerial. An enzymatic cascade of Rab5 effectors regulates phosphoinositide turnover in the endocytic pathway. Journal of Cell Biology, 170(4):607–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/170/4/607.

Scarcelli:2007:YIM

[SHC07] John J. Scarcelli, Christine A. Hodge, and Charles N. Cole. The yeast integral membrane protein Apq12 potentially links membrane dynamics to assembly of nuclear pore complexes. Journal of Cell Biology, 178(5):799–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/178/5/799.

Silk:2009:RNM

[SHC09] Alain D. Silk, Andrew J. Holland, and Don W. Cleveland. Requirements for NuMA in maintenance and establishment of mammalian spindle poles. Journal of Cell Biology, 184 (5):677–??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/184/5/677.

Smolka:2006:FDM

[ShCM^{+}06] Marcus B. Smolka, Sheng hong Chen, Paul S. Maddox, Jorrit M. Enserink, Claudio P. Albuquerque, Xiao X. Wei, Arshad Desai, Richard D. Kolodner, and Huilin Zhou. An FHA domain-mediated protein interaction network of Rad53 reveals its role in polarized cell growth. Journal of Cell Biology, 175 (5):743–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/5/743.

Strahl:2005:YPK

[SHDT05] Thomas Strahl, Hiroko Hama, Daryll B. DeWald, and Jeremy Thorner. Yeast phosphatidylinositol 4-kinase, Pik1, has essential roles at the Golgi and in the nucleus. Journal of Cell Biology, 171(6):967–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/6/967.

Spiliotis:2008:EPR

[SHH^{+}08] Elias T. Spiliotis, Stephen J. Hunt, Qicong Hu, Makoto Kinoshita, and W. James Nelson. Epithelial polarity requires septin coupling of vesicle transport to polyglutamylated microtubules. Journal of Cell Biology, 180(2):295–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/180/2/295.

Singh:2007:ICU

[SHjC^{+}07] Raman Deep Singh, Eileen L. Holicky, Zhi jie Cheng, Seong-Youl Kim, Christine L. Wheatley, David L. Marks, Robert Bittman, and Richard E. Pagano. Inhibition of caveolar uptake, SV40 infection, and β1-integrin signaling by a nonnatural glycosphingolipid stereoisomer. Journal of Cell Biology, 176 (7):895–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/7/895.

Spada:2007:DIR

[SHK^{+}07] Fabio Spada, Andrea Haemmer, David Kuch, Ulrich Rothbauer, Lothar Schermelleh, Elisabeth Kremmer, Thomas Carell, Gernot Längst, and Heinrich Leonhardt. DNMT1 but not its interaction with the replication machinery is required for maintenance of DNA methylation in human cells. Journal of Cell Biology, 176(5):565–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/5/565.

Sumakovic:2009:URE

[SHL^{+}09] Marija Sumakovic, Jan Hegermann, Ling Luo, Steven J. Husson, Katrin Schwarze, Christian Olendrowitz, Liliane Schoofs, Janet Richmond, and Stefan Eimer. UNC-108/ RAB-2 and its effector RIC-19 are involved in dense core vesicle maturation in Caenorhabditis elegans. Journal of Cell Biology, 186(6): 897–??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/186/6/897.

Short:2008:FAS

[Sho08] Ben Short. FHL1 adds some muscle. Journal of Cell Biology, 183(6):??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/6/965.2.

Short:2009:TGM

[Sho09a] Ben Short. γ-TuRC gives microtubules pause for thought. Journal of Cell Biology, 187(3):??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/187/3/314.1.

Short:2009:AM

[Sho09b] Ben Short. Adhering to the message. Journal of Cell Biology, 187(4):445–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/4/445.

Short:2009:AML

[Sho09c] Ben Short. Alexander Meissner: Learning the reprogramming code. Journal of Cell Biology, 186(4):450–??, August 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/186/4/450.

Short:2009:ACE

[Sho09d] Ben Short. Apratoxin chaperones EGFR to its destruction. Journal of Cell Biology, 185(4):??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/185/4/566.1.

Short:2009:ABA

[Sho09e] Ben Short. Aurora B answers an XIST-ential question. Journal of Cell Biology, 186(4):449–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/186/4/449.

Short:2009:ABH

[Sho09f] Ben Short. Autophagy balances a Highwire act in synapse development. Journal of Cell Biology, 187(1):??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/187/1/2.2.

Short:2009:BPR

[Sho09g] Ben Short. BBS proteins run an export business. Journal of Cell Biology, 187(7):??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/7/936.3.

Short:2009:BSMa

[Sho09h] Ben Short. BDNF sends mixed signals. Journal of Cell Biology, 185(4):??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/4/566.2.

Short:2009:BSS

[Sho09i] Ben Short. Bmper sticker sends Bmp to lysosomes. Journal of Cell Biology, 184(4):??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/184/4/465.2.

Short:2009:BSMb

[Sho09j] Ben Short. Bringing stability to microtubules. Journal of Cell Biology, 185(7):??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/7/1130.1.

Short:2009:CBE

[Sho09k] Ben Short. Cell biologists expand their networks. Journal of Cell Biology, 186(3):305–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/186/3/305.

Short:2009:CCK

[Sho09l] Ben Short. Cell cycle keeps Fus2p in check before yeast mate. Journal of Cell Biology, 184(3):??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/184/3/338.2.

Short:2009:CPC

[Sho09m] Ben Short. Cell packing comes under the lens. Journal of Cell Biology, 186(6):??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/186/6/768.3.

Short:2009:CDD

[Sho09n] Ben Short. Chemosensitivity depends on the dependency. Journal of Cell Biology, 187(3):??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/187/3/314.3.

Short:2009:CUT

[Sho09o] Ben Short. Christian Ungermann: Taking apart vacuole fusion. Journal of Cell Biology, 184(3):340–??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/184/3/340.

Short:2009:CPM

[Sho09p] Ben Short. Cohesin pathway makes the necessary arrangements. Journal of Cell Biology, 187(4):??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/187/4/444.1.

Short:2009:DSI

[Sho09q] Ben Short. David Sherwood: Invasive procedures. Journal of Cell Biology, 185(4):568–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/185/4/568.

Short:2009:DMN

[Sho09r] Ben Short. Decorin has Met a new receptor. Journal of Cell Biology, 185(4):??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/4/566.3.

Short:2009:DDD

[Sho09s] Ben Short. Desmoglein directs differentiation. Journal of Cell Biology, 185(7):??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/7/1130.2.

Short:2009:DTB

[Sho09t] Ben Short. The different ties that bind. Journal of Cell Biology, 187(2):151–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/2/151.

Short:2009:EFL

[Sho09u] Ben Short. Elaine Fuchs: a love for science that’s more than skin deep. Journal of Cell Biology, 187(7):938–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/187/7/
938.

Short:2009:FIS

[Sho09v] Ben Short. Fibroblasts invade at a snail’s pace. Journal of Cell Biology, 184(3):??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/3/339.2.

Short:2009:FGC

[Sho09w] Ben Short. Fibronectin is a good catch for integrin. Journal of Cell Biology, 185(7):??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/7/1130.3.

Short:2009:GMT

[Sho09x] Ben Short. Genes make their position clear. Journal of Cell Biology, 187(6):??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/6/752.1.

Short:2009:GCD

[Sho09y] Ben Short. Germ cells decide for themselves. Journal of Cell Biology, 187(4):??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/4/444.2.

Short:2009:GLC

[Sho09z] Ben Short. GSK3β loosens the CLASP on microtubules. Journal of Cell Biology, 184(6):??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/184/6/766.3.

Short:2009:HSG

[Sho09-27] Ben Short. How ER stress gives cells the CHOP. Journal of Cell Biology, 186(6):??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/186/6/768.1.

Short:2009:ICS

[Sho09-28] Ben Short. Iain Cheeseman: a strong attachment to kinetochores. Journal of Cell Biology, 186(6):770–??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/186/6/770.

Short:2009:ICB

[Sho09-29] Ben Short. IKK cuts both ways in Huntington’s disease. Journal of Cell Biology, 187(7):??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/187/7/936.2.

Short:2009:KTD

[Sho09-30] Ben Short. KRas takes a different route. Journal of Cell Biology, 184(6):??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/184/6/766.1.

Short:2009:LDL

[Sho09-31] Ben Short. Lipid droplets lead a Spartin existence. Journal of Cell Biology, 184(6):??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/6/766.2.

Short:2009:LRI

[Sho09-32] Ben Short. Lipid rafts inflate at the trans–Golgi. Journal of Cell Biology, 185(4):567–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/4/567.

Short:2009:LSM

[Sho09-33] Ben Short. The long and short of mitochondrial fusion. Journal of Cell Biology, 186(6):??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/186/6/768.2.

Short:2009:LPT

[Sho09-34] Ben Short. Lucas Pelkmans: Taking it from the top. Journal of Cell Biology, 185(6):932–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/185/6/932.

Short:2009:MWT

[Sho09-35] Ben Short. Marian Walhout: Transcriptional mapmaker. Journal of Cell Biology, 186(1):4–??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/186/1/4.

Short:2009:MGB

[Sho09-36] Ben Short. Matt Gibson: The blueprints of tissue architecture. Journal of Cell Biology, 186(2):166–??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/186/2/166.

Short:2009:MET

[Sho09-37] Ben Short. Membranes enter transfer negotiations. Journal of Cell Biology, 187(6):??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/187/6/752.3.

Short:2009:MCB

[Sho09-38] Ben Short. Merkel cells bear the touch of epidermis. Journal of Cell Biology, 187(1):??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/187/1/2.3.

Short:2009:MGG

[Sho09-39] Ben Short. Microvilli give gut vesicles the brush-off. Journal of Cell Biology, 185(7):1131–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/185/7/1131.

Short:2009:MBC

[Sho09-40] Ben Short. Mina Bissell: Context is everything. Journal of Cell Biology, 185(3):374–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/3/374.

Short:2009:MPS

[Sho09-41] Ben Short. Mitochondrial proteases strike the right balance. Journal of Cell Biology, 187(7):937–??, December 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/187/7/937.

Short:2009:MST

[Sho09-42] Ben Short. Mitotic spindle ties the Golgi ribbon. Journal of Cell Biology, 184(3):??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/3/339.1.

Short:2009:MTS

[Sho09-43] Ben Short. mRNP transport starts in the nucleus. Journal of Cell Biology, 184(3):??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/184/3/338.1.

Short:2009:MSS

[Sho09-44] Ben Short. A MUG shot of spindle assembly. Journal of Cell Biology, 187(1):??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/1/2.1.

Short:2009:MDB

[Sho09-45] Ben Short. Myosin drives bladder movements. Journal of Cell Biology, 186(4):??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/186/4/448.2.

Short:2009:AVA

[Sho09-46] Ben Short. Not the average view of adipogenesis. Journal of Cell Biology, 187(3):315–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/187/3/315.

Short:2009:OMT

[Sho09-47] Ben Short. Opposing motors take the strain. Journal of Cell Biology, 187(7):??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/7/936.1.

Short:2009:PIR

[Sho09-48] Ben Short. Peroxisome inheritance at the right time and place. Journal of Cell Biology, 186(4):??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/186/4/448.1.

Short:2009:RDS

[Sho09-49] Ben Short. Rab2 directs a stop-loss program. Journal of Cell Biology, 186(6):769–??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/186/6/769.

Short:2009:RST

[Sho09-50] Ben Short. Raf-1 sheds then spreads its inhibitions. Journal of Cell Biology, 187(3):??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/187/3/314.2.

Short:2009:SSR

[Sho09-51] Ben Short. Sizing up the ER stress response. Journal of Cell Biology, 187(4):??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/4/444.3.

Short:2009:TPH

[Sho09-52] Ben Short. Time (and PPARβ /δ) heals all wounds. Journal of Cell Biology, 184(6):767–??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/184/6/767.

Short:2009:VHG

[Sho09-53] Ben Short. Valerie Horsley: Getting under the skin. Journal of Cell Biology, 184(4):466–??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/184/4/466.

Short:2009:VTT

[Sho09-54] Ben Short. VCP takes out the trash. Journal of Cell Biology, 187(6):??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/187/6/752.2.

Stavru:2006:NCM

[SHS^{+}06] Fabrizia Stavru, Bastian B. Hülsmann, Anne Spang, Enno Hartmann, Volker C. Cordes, and Dirk Görlich. NDC1: a crucial membrane-integral nucleoporin of metazoan nuclear pore complexes. Journal of Cell Biology, 173(4):509–??, May 2006.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/173/4/509.

Sawyer:2009:DAH

[SHS^{+}09] Jessica K. Sawyer, Nathan J. Harris, Kevin C. Slep, Ulrike Gaul, and Mark Peifer. The Drosophila afadin homologue Canoe regulates linkage of the actin cytoskeleton to adherens junctions during apical constriction. Journal of Cell Biology, 186(1):57–??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/186/1/57.

Shcherbakova:2007:OAS

[SHX^{+}07] Olga G. Shcherbakova, Carl M. Hurt, Yang Xiang, Mark L. Dell’Acqua, Qi Zhang, Richard W. Tsien, and Brian K. Kobilka. Organization of β-adrenoceptor signaling compartments by sympathetic innervation of cardiac myocytes. Journal of Cell Biology, 176(4):521–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/176/4/521.

Skoufias:2007:MPA

[SIL^{+}07] Dimitrios A. Skoufias, Rose-Laure Indorato, Françoise Lacroix, Andreas Panopoulos, and Robert L. Margolis. Mitosis persists in the absence of Cdk1 activity when proteolysis or protein phosphatase activity is suppressed. Journal of Cell Biology, 179(4):671–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/4/671.

Sugimoto:2008:KBP

[SIS^{+}08] Masahiko Sugimoto, Akihito Inoko, Takashi Shiromizu, Masanori Nakayama, Peng Zou, Shigenobu Yonemura, Yuko Hayashi, Ichiro Izawa, Mikio Sasoh, Yukitaka Uji, Kozo Kaibuchi, Tohru Kiyono, and Masaki Inagaki. The keratinbinding protein Albatross regulates polarization of epithelial cells. Journal of Cell Biology, 183(1):19–??, October 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/183/1/19.

Sugimura:2008:PER

[SiTT^{+}08] Kazuto Sugimura, Shin ichiro Takebayashi, Hiroshi Taguchi, Shunichi Takeda, and Katsuzumi Okumura. PARP-1 ensures regulation of replication fork progression by homologous recombination on damaged DNA. Journal of Cell Biology, 183(7): 1203–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/7/1203.

Sasaki:2007:GPI

[SJL^{+}07] Atsuo T. Sasaki, Chris Janetopoulos, Susan Lee, Pascale G. Charest, Kosuke Takeda, Lauren W. Sundheimer, Ruedi Meili, Peter N. Devreotes, and Richard A. Firtel. G proteinindependent Ras/PI3K/F-actin circuit regulates basic cell motility. Journal of Cell Biology, 178(2):185–??, July 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/178/2/185.

Skach:2007:ERT

[Ska07] William R. Skach. The expanding role of the ER translocon in membrane protein folding. Journal of Cell Biology, 179(7): 1333–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/7/1333.

Saharinen:2005:MAR

[SKE^{+}05] Pipsa Saharinen, Katja Kerkelä, Niklas Ekman, Marie Marron, Nicholas Brindle, Gyun Min Lee, Hellmut Augustin, Gou Young Koh, and Kari Alitalo. Multiple angiopoietin recombinant proteins activate the Tie1 receptor tyrosine kinase and promote its interaction with Tie2. Journal of Cell Biology, 169(2):239–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
169/2/239.

Skibbens:2005:UL

[Ski05] Robert V. Skibbens. Unzipped and loaded. Journal of Cell Biology, 169(6):841–??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/6/841.

Saleem:2008:GWA

[SKM^{+}08] Ramsey A. Saleem, Barbara Knoblach, Fred D. Mast, Jennifer J. Smith, John Boyle, C. Melissa Dobson, Rose Long-O’Donnell, Richard A. Rachubinski, and John D. Aitchison. Genome-wide analysis of signaling networks regulating fatty acid–induced gene expression and organelle biogenesis. Journal of Cell Biology, 181(2):281–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/181/2/281.

Suetsugu:2006:OWC

[SKO^{+}06] Shiro Suetsugu, Shusaku Kurisu, Tsukasa Oikawa, Daisuke Yamazaki, Atsushi Oda, and Tadaomi Takenawa. Optimization of WAVE2 complex–induced actin polymerization by membranebound IRSp53, PIP_{3}, and Rac. Journal of Cell Biology, 173 (4):571–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/4/571.

Sakurai:2008:MMS

[SKO^{+}08] Takashi Sakurai, Kumi Kaneko, Misako Okuno, Koji Wada, Taku Kashiyama, Hideaki Shimizu, Takumi Akagi, Tsutomu Hashikawa, and Nobuyuki Nukina. Membrane microdomain switching: a regulatory mechanism of amyloid precursor protein processing. Journal of Cell Biology, 183(2):339–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/183/2/
339.

Sleeman:2007:DER

[SKR^{+}07] Katherine E. Sleeman, Howard Kendrick, David Robertson, Clare M. Isacke, Alan Ashworth, and Matthew J. Smalley. Dissociation of estrogen receptor expression and in vivo stem cell activity in the mammary gland. Journal of Cell Biology, 176 (1):19–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/1/19.

Sachs:2006:KFM

[SKvdBW^{+}06] Norman Sachs, Maaike Kreft, Marius A. van den Bergh Weerman, Andy J. Beynon, Theo A. Peters, Jan J. Weening, and Arnoud Sonnenberg. Kidney failure in mice lacking the tetraspanin CD151. Journal of Cell Biology, 175(1):33–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
175/1/33.

Song:2005:DAN

[SL05] Jae J. Song and Yong J. Lee. Dissociation of Akt1 from its negative regulator JIP1 is mediated through the ASK1–MEK– JNK signal transduction pathway during metabolic oxidative stress. Journal of Cell Biology, 170(1):61–??, July 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/170/1/61.

Shen:2005:HMA

[SLCB05] Siming Shen, Jiadong Li, and Patrizia Casaccia-Bonnefil. Histone modifications affect timing of oligodendrocyte progenitor differentiation in the developing rat brain. Journal of Cell Biology, 169(4):577–??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/4/577.

Song:2008:EHR

[SLK^{+}08] Sungmin Song, Huikyong Lee, Tae-In Kam, Mei Ling Tai, Joo-Yong Lee, Jee-Yeon Noh, Sang Mi Shim, Soo Jung Seo, Young-Yun Kong, Toshiyuki Nakagawa, Chul-Woong Chung, Deog-Young Choi, Hammou Oubrahim, and Yong-Keun Jung. E2-25K/ hip-2 regulates caspase-12 in ER stress-mediated Aβ neurotoxicity. Journal of Cell Biology, 182(4):675–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/182/4/675.

Sharma:2006:ATS

[SLM^{+}06] Nikunj Sharma, Seng Hui Low, Saurav Misra, Bhattaram Pallavi, and Thomas Weimbs. Apical targeting of syntaxin 3 is essential for epithelial cell polarity. Journal of Cell Biology, 173(6):937–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/6/937.

Stanya:2008:CPN

[SLMK08] Kristopher J. Stanya, Yu Liu, Anthony R. Means, and Hung-Ying Kao. Cdk2 and Pin1 negatively regulate the transcriptional corepressor SMRT. Journal of Cell Biology, 183(1):49–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
183/1/49.

Sytnyk:2006:NPA

[SLNS06] Vladimir Sytnyk, Iryna Leshchyns’ka, Alexander G. Nikonenko, and Melitta Schachner. NCAM promotes assembly and activitydependent remodeling of the postsynaptic signaling complex. Journal of Cell Biology, 174(7):1071–??, September 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/174/7/1071.

Shopland:2006:FOC

[SLP^{+}06] Lindsay S. Shopland, Christopher R. Lynch, Kevin A. Peterson, Kathleen Thornton, Nick Kepper, Johann von Hase, Stefan Stein, Sarah Vincent, Kelly R. Molloy, Gregor Kreth, Christoph Cremer, Carol J. Bult, and Timothy P. O’Brien. Folding and organization of a contiguous chromosome region according to the gene distribution pattern in primary genomic sequence. Journal of Cell Biology, 174(1):27–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/174/1/27.

Stohr:2006:ZRM

[SLR^{+}06] Nadine Stöhr, Marcell Lederer, Claudia Reinke, Sylke Meyer, Mechthild Hatzfeld, Robert H. Singer, and Stefan Hüttelmaier. ZBP1 regulates mRNA stability during cellular stress. Journal of Cell Biology, 175(4):527–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/175/4/527.

Sun:2007:SRS

[SLV^{+}07] Hong Sun, Li Li, Cécile Vercherat, Neriman Tuba Gulbagci, Sujata Acharjee, Jiali Li, Teng-Kai Chung, Tin Htwe Thin, and Reshma Taneja. Stra13 regulates satellite cell activation by antagonizing Notch signaling. Journal of Cell Biology, 177(4):647– ??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
177/4/647.

Sun:2007:TPP

[SLWA07] Yue Sun, Kun Ling, Matthew P. Wagoner, and Richard A. Anderson. Type Iγ phosphatidylinositol phosphate kinase is required for EGF-stimulated directional cell migration. Journal of Cell Biology, 178(2):297–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/178/2/297.

Song:2006:IPT

[SLZ^{+}06] Lun Song, Jingxia Li, Dongyun Zhang, Zheng gang Liu, Jianping Ye, Qimin Zhan, Han-Ming Shen, Matt Whiteman, and Chuanshu Huang. IKKβ programs to turn on the GADD45α-MKK4-JNK apoptotic cascade specifically via p50 NF-κB in arsenite response. Journal of Cell Biology, 175(4):607–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/175/4/
607.

Schwartz:2009:CRP

[SM09] Matthew L. Schwartz and Alexey J. Merz. Capture and release of partially zipped trans–SNARE complexes on intact organelles. Journal of Cell Biology, 185(3):535–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/185/3/535.

Sun:2007:RRD

[SMG07] Chun Xiang Sun, Marco A. O. Magalhães, and Michael Glogauer. Rac1 and Rac2 differentially regulate actin free barbed end formation downstream of the fMLP receptor. Journal of Cell Biology, 179(2):239–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/2/239.

Schlessinger:2007:CNW

[SMH07] Karni Schlessinger, Edward J. McManus, and Alan Hall. Cdc42 and noncanonical Wnt signal transduction pathways cooperate to promote cell polarity. Journal of Cell Biology, 178(3):355–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
178/3/355.

Saltel:2009:NPP

[SMH^{+}09] Frédéric Saltel, Eva Mortier, Vesa P. Hytönen, Marie-Claude Jacquier, Pascale Zimmermann, Viola Vogel, Wei Liu, and Bernhard Wehrle-Haller. New PI(4,5)P_{2}- and membrane proximal integrin–binding motifs in the talin head control β3-integrin clustering. Journal of Cell Biology, 187(5):715–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/187/5/

Smith:2007:HSF

[Smi07] Elizabeth F. Smith. Hydin seek: finding a function in ciliary motility. Journal of Cell Biology, 176(4):403–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/176/4/403.

Sasaki:2008:BDA

[SMM^{+}08a] Toru Sasaki, Edyta Marcon, Tracy McQuire, Yoichi Arai, Peter B. Moens, and Hitoshi Okada. Bat3 deficiency accelerates the degradation of Hsp70-2/HspA2 during spermatogenesis. Journal of Cell Biology, 182(3):449–??, August 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/182/3/449.

Smith:2008:UPF

[SMM08b] Harvey W. Smith, Pierfrancesco Marra, and Christopher J. Marshall. uPAR promotes formation of the p130Cas–Crk complex to activate Rac through DOCK180. Journal of Cell Biology, 182(4):777–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/4/777.

Stojanovski:2008:MPI

[SMM^{+}08c] Diana Stojanovski, Dusanka Milenkovic, Judith M. Müller, Kipros Gabriel, Agnes Schulze-Specking, Michael J. Baker, Michael T. Ryan, Bernard Guiard, Nikolaus Pfanner, and Agnieszka Chacinska. Mitochondrial protein import: precursor oxidation in a ternary complex with disulfide carrier and sulfhydryl oxidase. Journal of Cell Biology, 183(2):195–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
183/2/195.

Slattum:2009:PRM

[SMR09] Gloria Slattum, Karen M. McGee, and Jody Rosenblatt. P115 RhoGEF and microtubules decide the direction apoptotic cells extrude from an epithelium. Journal of Cell Biology, 186(5): 693–??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/186/5/693.

Solan:2007:PSG

[SMRS^{+}07] Joell L. Solan, Lucrecia Marquez-Rosado, Paul L. Sorgen, Perry J. Thornton, Philip R. Gafken, and Paul D. Lampe. Phosphorylation at S365 is a gatekeeper event that changes the structure of Cx43 and prevents down-regulation by PKC. Journal of Cell Biology, 179(6):1301–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/6/1301.

Santiago-Martinez:2008:RSR

[SMSPK08] Edgardo Santiago-Martínez, Nadine H. Soplop, Rajesh Patel, and Sunita G. Kramer. Repulsion by Slit and Roundabout prevents Shotgun/E-cadherin-mediated cell adhesion during Drosophila heart tube lumen formation. Journal of Cell Biology, 182(2):241–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/2/241.

Spycher:2008:CPM

[SMT^{+}08] Christoph Spycher, Edward S. Miller, Kelly Townsend, Lucijana Pavic, Nicholas A. Morrice, Pavel Janscak, Grant S. Stewart, and Manuel Stucki. Constitutive phosphorylation of MDC1 physically links the MRE11–RAD50–NBS1 complex to damaged chromatin. Journal of Cell Biology, 181(2):227–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/181/2/
227.

Sun:2007:JSS

[SMW^{+}07] Luguo Sun, Kewei Ma, Haixia Wang, Fang Xiao, Yan Gao, Wei Zhang, Kepeng Wang, Xiang Gao, Nancy Ip, and Zhenguo Wu. JAK1–STAT1–STAT3, a key pathway promoting proliferation and preventing premature differentiation of myoblasts. Journal of Cell Biology, 179(1):129–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/1/129.

Saint-Marcoux:2009:BCB

[SMWdV09] Denis Saint-Marcoux, Francis-André Wollman, and Catherine de Vitry. Biogenesis of cytochrome b_{6} in photosynthetic membranes. Journal of Cell Biology, 185(7):1195–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/185/7/
1195.

Spear:2005:SCS

[SN05] Eric D. Spear and Davis T. W. Ng. Single, context-specific glycans can target misfolded glycoproteins for ER-associated degradation. Journal of Cell Biology, 169(1):73–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/169/1/73.

Shimohata:2007:SAI

[SNA^{+}07] Nobuyuki Shimohata, Shushi Nagamori, Yoshinori Akiyama, H. Ronald Kaback, and Koreaki Ito. SecY alterations that impair membrane protein folding and generate a membrane stress. Journal of Cell Biology, 176(3):307–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/176/3/307.

Schneider:2006:GCC

[SNC^{+}06] Mary D. Schneider, Nima Najand, Sana Chaker, Justin M. Pare, Julie Haskins, Sarah C. Hughes, Tom C. Hobman, John Locke, and Andrew J. Simmonds. Gawky is a component of cytoplasmic mRNA processing bodies required for early Drosophila development. Journal of Cell Biology, 174(3):349– ??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
174/3/349.

Smith:2005:PPF

[SNG^{+}05] Wanli W. Smith, Darrell D. Norton, Myriam Gorospe, Haibing Jiang, Shino Nemoto, Nikki J. Holbrook, Toren Finkel, and John W. Kusiak. Phosphorylation of p66Shc and forkhead proteins mediates Aβ toxicity. Journal of Cell Biology, 169(2):331–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/2/331.

Sevilla:2007:MDI

[SNG^{+}07] Lisa M. Sevilla, Rachida Nachat, Karen R. Groot, John F. Klement, Jouni Uitto, Philippe Djian, Arto Määttä, and Fiona M. Watt. Mice deficient in involucrin, envoplakin, and periplakin have a defective epidermal barrier. Journal of Cell Biology, 179(7):1599–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/7/1599.

Stavru:2006:NPC

[SNPCG06] Fabrizia Stavru, Gitte Nautrup-Pedersen, Volker C. Cordes, and Dirk Görlich. Nuclear pore complex assembly and maintenance in POM121- and gp210-deficient cells. Journal of Cell Biology, 173(4):477–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/4/477.

Spagnoli:2007:TSE

[SOC^{+}07] Anna Spagnoli, Lynda O’Rear, Ronald L. Chandler, Froilan Granero-Molto, Douglas P. Mortlock, Agnieszka E. Gorska, Jared A. Weis, Lara Longobardi, Anna Chytil, Kimberly Shimer, and Harold L. Moses. TGF-β signaling is essential for joint morphogenesis. Journal of Cell Biology, 177(6):1105–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
177/6/1105.

Saito:2006:MMC

[SON06] Takamune T. Saito, Daisuke Okuzaki, and Hiroshi Nojima. Mcp5, a meiotic cell cortex protein, is required for nuclear movement mediated by dynein and microtubules in fission yeast. Journal of Cell Biology, 173(1):27–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/173/1/27.

Shemesh:2005:NMA

[SOR^{+}05] Tom Shemesh, Takanori Otomo, Michael K. Rosen, Alexander D. Bershadsky, and Michael M. Kozlov. A novel mechanism of actin filament processive capping by formin. Journal of Cell Biology, 170(6):889–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/170/6/889.

Seth:2006:ARC

[SOR06] Abhinav Seth, Chinatsu Otomo, and Michael K. Rosen. Autoinhibition regulates cellular localization and actin assembly activity of the diaphanous-related formins FRLα and mDia1. Journal of Cell Biology, 174(5):701–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/174/5/701.

Sommer:2005:ECS

[SORM05] Bernhard Sommer, Adrian Oprins, Catherine Rabouille, and Sean Munro. The exocyst component Sec5 is present on endocytic vesicles in the oocyte of Drosophila melanogaster. Journal of Cell Biology, 169(6):953–??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/169/6/953.

Shields:2009:EUB

[SOW^{+}09] S. Brookhart Shields, Andrea J. Oestreich, Stanley Winistorfer, Doris Nguyen, Johanna A. Payne, David J. Katzmann, and Robert Piper. ESCRT ubiquitin-binding domains function cooperatively during MVB cargo sorting. Journal of Cell Biology, 185(2):213–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/2/213.

Shen:2005:CKF

[SPA^{+}05] Tang-Long Shen, Ann Y.-J. Park, Ana Alcaraz, Xu Peng, Ihnkyung Jang, Pandelakis Koni, Richard A. Flavell, Hua Gu, and Jun-Lin Guan. Conditional knockout of focal adhesion kinase in endothelial cells reveals its role in angiogenesis and vascular development in late embryogenesis. Journal of Cell Biology, 169(6):941–??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/6/941.

Sideris:2009:NIS

[SPK^{+}09] Dionisia P. Sideris, Nikos Petrakis, Nitsa Katrakili, Despina Mikropoulou, Angelo Gallo, Simone Ciofi-Baffoni, Lucia Banci, Ivano Bertini, and Kostas Tokatlidis. A novel intermembrane space–targeting signal docks cysteines onto Mia40 during mitochondrial oxidative folding. Journal of Cell Biology, 187(7): 1007–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/187/7/1007.

Sanij:2008:ULD

[SPS^{+}08] Elaine Sanij, Gretchen Poortinga, Kerith Sharkey, Sandy Hung, Timothy P. Holloway, Jaclyn Quin, Elysia Robb, Lee H. Wong, Walter G. Thomas, Victor Stefanovsky, Tom Moss, Lawrence Rothblum, Katherine M. Hannan, Grant A. McArthur, Richard B. Pearson, and Ross D. Hannan. UBF levels determine the number of active ribosomal RNA genes in mammals. Journal of Cell Biology, 183(7):1259–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/183/7/
1259.

Schuck:2009:MEA

[SPT^{+}09] Sebastian Schuck, William A. Prinz, Kurt S. Thorn, Christiane Voss, and Peter Walter. Membrane expansion alleviates endoplasmic reticulum stress independently of the unfolded protein response. Journal of Cell Biology, 187(4):525–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/187/4/

525. See correction [SPT^{+}21].

Schuck:2021:CME

[SPT^{+}21] Sebastian Schuck, William A. Prinz, Kurt S. Thorn, Christiane Voss, and Peter Walter. Correction: Membrane expansion alleviates endoplasmic reticulum stress independently of the unfolded protein response. Journal of Cell Biology, 220(4):??, April 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/4/jcb.20090707402092021c/211800/
Correction-Membrane-expansion-alleviates. See [SPT^{+}09].

Szabo:2008:CIC

[SQB^{+}08] Eva Szabo, Yuanyuan Qiu, Shairaz Baksh, Marek Michalak, and Michal Opas. Calreticulin inhibits commitment to adipocyte differentiation. Journal of Cell Biology, 182(1):103– ??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
182/1/103.

Sloboda:2007:MSC

[SR07] Roger D. Sloboda and Joel L. Rosenbaum. Making sense of cilia and flagella. Journal of Cell Biology, 179(4):575–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/4/

Sahlender:2005:OLM

[SRA^{+}05] Daniela A. Sahlender, Rhys C. Roberts, Susan D. Arden, Giulietta Spudich, Marcus J. Taylor, J. Paul Luzio, John Kendrick-Jones, and Folma Buss. Optineurin links myosin VI to the Golgi complex and is involved in Golgi organization and exocytosis. Journal of Cell Biology, 169(2):285–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/169/2/285.

Schmidt:2006:VBA

[SRBL06] Ute Schmidt, Karsten Richter, Axel Bernhard Berger, and Peter Lichter. In vivo BiFC analysis of Y14 and NXF1 mRNA export complexes: preferential localization within and around SC35 domains. Journal of Cell Biology, 172(3):373–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/172/3/
373.

Spasic:2007:RCA

[SRD^{+}07] Dragana Spasic, Tim Raemaekers, Katleen Dillen, Ilse Declerck, Veerle Baert, Lutgarde Serneels, Joachim Füllekrug, and Wim Annaert. Rer1p competes with APH-1 for binding to nicastrin and regulates γ-secretase complex assembly in the early secretory pathway. Journal of Cell Biology, 176(5):629–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/176/5/
629.

Slep:2005:SDE

[SRE^{+}05] Kevin C. Slep, Stephen L. Rogers, Sarah L. Elliott, Hiroyuki Ohkura, Peter A. Kolodziej, and Ronald D. Vale. Structural determinants for EB1-mediated recruitment of APC and spectraplakins to the microtubule plus end. Journal of Cell Biology, 168(4):587–??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/4/587.

Sano:2005:CLM

[SRL05] Hiroko Sano, Andrew D. Renault, and Ruth Lehmann. Control of lateral migration and germ cell elimination by the Drosophila melanogaster lipid phosphate phosphatases Wunen and Wunen 2. Journal of Cell Biology, 171(4):675–??, November 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/171/4/675.

Sanjuan:2006:CIT

[SRL^{+}06] Miguel A. Sanjuan, Navin Rao, Kuei-Tai A. Lai, Yin Gu, Siquan Sun, Anja Fuchs, Wai-Ping Fung-Leung, Marco Colonna, and Lars Karlsson. CpG-induced tyrosine phosphorylation occurs via a TLR9-independent mechanism and is required for cytokine secretion. Journal of Cell Biology, 172(7):1057–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
172/7/1057.

Schober:2007:FAK

[SRN^{+}07] Markus Schober, Srikala Raghavan, Maria Nikolova, Lisa Polak, H. Amalia Pasolli, Hilary E. Beggs, Louis F. Reichardt, and Elaine Fuchs. Focal adhesion kinase modulates tension signaling to control actin and focal adhesion dynamics. Journal of Cell Biology, 176(5):667–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/176/5/667.

Szakmary:2009:YPD

[SRQL09] Akos Szakmary, Mary Reedy, Hongying Qi, and Haifan Lin. The Yb protein defines a novel organelle and regulates male germline stem cell self-renewal in Drosophila melanogaster. Journal of Cell Biology, 185(4):613–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/185/4/613.

Shechter:2008:TLR

[SRR^{+}08] Ravid Shechter, Ayal Ronen, Asya Rolls, Anat London, Sharon Bakalash, Michael J. Young, and Michal Schwartz. Toll-like receptor 4 restricts retinal progenitor cell proliferation. Journal of Cell Biology, 183(3):393–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/183/3/393.

Skaar:2009:ICH

[SRS^{+}09] Jeffrey R. Skaar, Derek J. Richard, Anita Saraf, Alfredo Toschi, Emma Bolderson, Laurence Florens, Michael P. Washburn, Kum Kum Khanna, and Michele Pagano. INTS3 controls the hSSB1-mediated DNA damage response. Journal of Cell Biology, 187(1):25–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/1/25.

Sabri:2007:DFD

[SRX^{+}07] Nafiseh Sabri, Peggy Roth, Nikos Xylourgidis, Fatemeh Sadeghifar, Jeremy Adler, and Christos Samakovlis. Distinct functions of the Drosophila Nup153 and Nup214 FG domains in nuclear protein transport. Journal of Cell Biology, 178(4): 557–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/4/557.

Smukler:2006:ESC

[SRXvdK06] Simon R. Smukler, Susan B. Runciman, Shunbin Xu, and Derek van der Kooy. Embryonic stem cells assume a primitive neural stem cell fate in the absence of extrinsic influences. Journal of Cell Biology, 172(1):79–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/172/1/79.

Schuck:2006:CFT

[SS06a] Sebastian Schuck and Kai Simons. Controversy fuels trafficking of GPI-anchored proteins. Journal of Cell Biology, 172 (7):963–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/7/963.

Shaul:2006:ERG

[SS06b] Yoav D. Shaul and Rony Seger. ERK1c regulates Golgi fragmentation during mitosis. Journal of Cell Biology, 172(6): 885–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/6/885.

Sakaguchi:2007:AMH

[SS07] Ayako Sakaguchi and Ruth Steward. Aberrant monomethylation of histone H4 lysine 20 activates the DNA damage checkpoint in Drosophila melanogaster. Journal of Cell Biology, 176 (2):155–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/2/155.

Scott:2008:RSA

[SS08a] Daniel C. Scott and Randy Schekman. Role of Sec61p in the ER-associated degradation of short-lived transmembrane proteins. Journal of Cell Biology, 181(7):1095–??, June 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/181/7/1095.

Simonsen:2008:SEA

[SS08b] Anne Simonsen and Harald Stenmark. Self-eating from an ERassociated cup. Journal of Cell Biology, 182(4):621–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/182/4/
621.

Sivaramakrishnan:2009:CMV

[SS09] Sivaraj Sivaramakrishnan and James A. Spudich. Coupled myosin VI motors facilitate unidirectional movement on an Factin network. Journal of Cell Biology, 187(1):53–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/187/1/
53.

Sun:2008:REG

[SSAD08] Jianjun Sun, Laila Smith, Alexander Armento, and Wu-Min Deng. Regulation of the endocycle/gene amplification switch by Notch and ecdysone signaling. Journal of Cell Biology, 182(5): 885–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/5/885.

Schubert:2006:LRA

[SSD06] Vanessa Schubert, Jorge Santos Da Silva, and Carlos G. Dotti. Localized recruitment and activation of RhoA underlies dendritic spine morphology in a glutamate receptor-dependent manner. Journal of Cell Biology, 172(3):453–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/172/3/453.

Schneider:2009:HEN

[SSD^{+}09] Linda Schneider, Christian-Martin Stock, Peter Dieterich, Bo Hammer Jensen, Lotte Bang Pedersen, Peter Satir, Albrecht Schwab, Søren Tvorup Christensen, and Stine Falsig Pedersen.

The Na^{+} /H^{+} exchanger NHE1 is required for directional migration stimulated via PDGFR-α in the primary cilium. Journal of Cell Biology, 185(1):163–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/185/1/163.

Sontag:2005:VAG

[SSE05] Christopher A. Sontag, James T. Staley, and Harold P. Erickson. In vitro assembly and GTP hydrolysis by bacterial tubulins BtubA and BtubB. Journal of Cell Biology, 169(2):233–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
169/2/233.

Schmidt:2005:HMG

[SSH^{+}05] Katy Schmidt, Thorsten Schinke, Michael Haberland, Matthias Priemel, Arndt F. Schilling, Cordula Mueldner, Johannes M. Rueger, Elisabeth Sock, Michael Wegner, and Michael Amling. The high mobility group transcription factor Sox8 is a negative regulator of osteoblast differentiation. Journal of Cell Biology, 168(6):899–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/6/899.

Salpingidou:2007:NRN

[SSHP^{+}07] Georgia Salpingidou, Andrei Smertenko, Irena Hausmanowa-Petrucewicz, Patrick J. Hussey, and Chris J. Hutchison. A novel role for the nuclear membrane protein emerin in association of the centrosome to the outer nuclear membrane. Journal of Cell Biology, 178(6):897–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/178/6/897.

Sabeh:2009:PDV

[SSHW09] Farideh Sabeh, Ryoko Shimizu-Hirota, and Stephen J. Weiss. Protease-dependent versus -independent cancer cell invasion programs: three-dimensional amoeboid movement revisited. Journal of Cell Biology, 185(1):11–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/185/1/11.

Schmidt:2007:RGC

[SSJ^{+}07] Hannes Schmidt, Agne Stonkute, René Jüttner, Susanne Schäffer, Jens Buttgereit, Robert Feil, Franz Hofmann, and Fritz G. Rathjen. The receptor guanylyl cyclase Npr2 is essential for sensory axon bifurcation within the spinal cord. Journal of Cell Biology, 179(2):331–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/2/331.

Staiger:2009:AFD

[SSK^{+}09] Christopher J. Staiger, Michael B. Sheahan, Parul Khurana, Xia Wang, David W. McCurdy, and Laurent Blanchoin. Actin filament dynamics are dominated by rapid growth and severing activity in the Arabidopsis cortical array. Journal of Cell Biology, 184(2):269–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/2/269.

Santuccione:2005:PPR

[SSLS05] Antonella Santuccione, Vladimir Sytnyk, Iryna Leshchyns’ka, and Melitta Schachner. Prion protein recruits its neuronal receptor NCAM to lipid rafts to activate p59^{fyn} and to enhance neurite outgrowth. Journal of Cell Biology, 169(2):341–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
169/2/341.

Shin:2005:PRT

[SSM05] Kunyoo Shin, Sam Straight, and Ben Margolis. PATJ regulates tight junction formation and polarity in mammalian epithelial cells. Journal of Cell Biology, 168(5):705–??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/168/5/705.

Semerdjieva:2008:CRA

[SSM^{+}08] Sophia Semerdjieva, Barry Shortt, Emma Maxwell, Sukhdeep Singh, Paul Fonarev, Jonathan Hansen, Giampietro Schiavo, Barth D. Grant, and Elizabeth Smythe. Coordinated regulation of AP2 uncoating from clathrin-coated vesicles by rab5 and hRME-6. Journal of Cell Biology, 183(3):499–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/183/3/

Shestakova:2007:ICO

[SSP^{+}07] Anna Shestakova, Elena Suvorova, Oleksandra Pavliv, Galimat Khaidakova, and Vladimir Lupashin. Interaction of the conserved oligomeric Golgi complex with t–SNARE Syntaxin5a/ Sed5 enhances intra-Golgi SNARE complex stability. Journal of Cell Biology, 179(6):1179–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/6/1179.

Schleicher:2008:PMA

[SSS^{+}08a] Michael Schleicher, Benjamin R. Shepherd, Yajaira Suarez, Carlos Fernandez-Hernando, Jun Yu, Yong Pan, Lisette M. Acevedo, Gerald S. Shadel, and William C. Sessa. Prohibitin-1 maintains the angiogenic capacity of endothelial cells by regulating mitochondrial function and senescence. Journal of Cell Biology, 180(1):101–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/1/101.

Shumaker:2008:HCN

[SSS^{+}08b] Dale K. Shumaker, Liliana Solimando, Kaushik Sengupta, Takeshi Shimi, Stephen A. Adam, Antje Grunwald, Sergei V. Strelkov, Ueli Aebi, M. Cristina Cardoso, and Robert D. Goldman. The highly conserved nuclear lamin Ig-fold binds to PCNA: its role in DNA replication. Journal of Cell Biology, 181(2):269–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/2/269.

Strochlic:2007:GSF

[SSSB07] Todd I. Strochlic, Thanuja Gangi Setty, Anand Sitaram, and Christopher G. Burd. Grd19/ snx3p functions as a cargospecific adapter for retromer-dependent endocytic recycling. Journal of Cell Biology, 177(1):115–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/177/1/115.

Smotrys:2005:VDC

[SSSL05] Jessica E. Smotrys, Marissa J. Schoenfish, Monica A. Stutz, and Maurine E. Linder. The vacuolar DHHC–CRD protein Pfa3p is a protein acyltransferase for Vac8p. Journal of Cell Biology, 170(7):1091–??, September 2005. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/170/7/1091.

Slepchenko:2007:SMO

[SSZR07] Boris M. Slepchenko, Irina Semenova, Ilya Zaliapin, and Vladimir Rodionov. Switching of membrane organelles between cytoskeletal transport systems is determined by regulation of the microtubule-based transport. Journal of Cell Biology, 179 (4):635–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/4/635.

Simonsen:2009:CME

[ST09] Anne Simonsen and Sharon A. Tooze. Coordination of membrane events during autophagy by multiple class III PI3-kinase complexes. Journal of Cell Biology, 186(6):773–??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/186/6/
773.

Sengupta:2009:OTH

[STBL09] Debrup Sengupta, Steven Truschel, Collin Bachert, and Adam D. Linstedt. Organelle tethering by a homotypic PDZ interaction underlies formation of the Golgi membrane network. Journal of Cell Biology, 186(1):41–??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/186/1/41.

Steele:2008:AQN

[Ste08] Andrew D. Steele. All quiet on the neuronal front: NMDA receptor inhibition by prion protein. Journal of Cell Biology, 181(3):407–??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/3/407.

Shang:2005:MAR

[STG05] Yuhua Shang, Che-Chia Tsao, and Martin A. Gorovsky. Mutational analyses reveal a novel function of the nucleotide-binding domain of γ-tubulin in the regulation of basal body biogenesis. Journal of Cell Biology, 171(6):1035–??, December 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/171/6/1035.

Shimizu:2005:RRC

[STK^{+}05] Yoshihiko Shimizu, Dean Thumkeo, Jeongsin Keel, Toshimasa Ishizaki, Hiroko Oshima, Masanobu Oshima, Yoichi Noda, Fumio Matsumura, Makoto M. Taketo, and Shuh Narumiya. ROCK-i regulates closure of the eyelids and ventral body wall by inducing assembly of actomyosin bundles. Journal of Cell Biology, 168(6):941–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/6/941.

Sun:2007:UNE

[STL^{+}07] Shaogang Sun, Yujie Tang, Xiwen Lou, Lianhui Zhu, Kai Yang, Bianhong Zhang, Hexin Shi, and Chen Wang. UXT is a novel and essential cofactor in the NF-κB transcriptional enhanceosome. Journal of Cell Biology, 178(2):231–??, July 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/178/2/231.

Sfakianos:2007:PFB

[STM^{+}07] Jeff Sfakianos, Akashi Togawa, Sandra Maday, Mike Hull, Marc Pypaert, Lloyd Cantley, Derek Toomre, and Ira Mellman. Par3 functions in the biogenesis of the primary cilium in polarized epithelial cells. Journal of Cell Biology, 179(6):1133–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/6/
1133.

Sunayama:2005:JAA

[STMG05] Jun Sunayama, Fuminori Tsuruta, Norihisa Masuyama, and Yukiko Gotoh. JNK antagonizes Akt-mediated survival signals by phosphorylating 14-3-3. Journal of Cell Biology, 170(2):295– ??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
170/2/295.

Sohn:2008:MHF

[STP08] Hae Won Sohn, Pavel Tolar, and Susan K. Pierce. Membrane heterogeneities in the formation of B cell receptor–Lyn kinase microclusters and the immune synapse. Journal of Cell Biology, 182(2):367–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/2/367.

Shimada:2008:SIA

[STU^{+}08] Tadayuki Shimada, Michinori Toriyama, Kaori Uemura, Hiroyuki Kamiguchi, Tadao Sugiura, Naoki Watanabe, and Naoyuki Inagaki. Shootin1 interacts with actin retrograde flow and L1-CAM to promote axon outgrowth. Journal of Cell Biology, 181(5):817–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/5/817.

Stagg:2009:TEL

[STvdB^{+}09] Helen R. Stagg, Mair Thomas, Dick van den Boomen, Emmanuel J. H. J. Wiertz, Harry A. Drabkin, Robert M. Gemmill, and Paul J. Lehner. The TRC8 E3 ligase ubiquitinates MHC class I molecules before dislocation from the ER. Journal of Cell Biology, 186(5):685–??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/186/5/685.

Suelves:2007:UDEa

[SVS^{+}07a] Mònica Suelves, Berta Vidal, Antonio L. Serrano, Marc Tjwa, Josep Roma, Roser López-Alemany, Aernout Luttun, María Martínez de Lagrán, Maria Àngels Díaz, Mercè Jardí, Manuel Roig, Mara Dierssen, Mieke Dewerchin, Peter Carmeliet, and Pura Muñoz-Cánoves. uPA deficiency exacerbates muscular dystrophy in MDX mice. Journal of Cell Biology, 178(6):1039–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/6/1039.

Suelves:2007:UDEb

[SVS^{+}07b] Mònica Suelves, Berta Vidal, Antonio L. Serrano, Marc Tjwa, Josep Roma, Roser López-Alemany, Aernout Luttun, María Martínez de Lagrán, Àngels Díaz-Ramos, Mercè Jardí, Manuel Roig, Mara Dierssen, Mieke Dewerchin, Peter Carmeliet, and Pura Muñoz-Cánoves. uPA deficiency exacerbates muscular dystrophy in MDX mice. Journal of Cell Biology, 179(1):165–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/1/165.

Sutton:2007:RAG

[SWB^{+}07] Vivien R. Sutton, Nigel J. Waterhouse, Kylie A. Browne, Karin Sedelies, Annette Ciccone, Desiree Anthony, Aulikki Koskinen, Arno Mullbacher, and Joseph A. Trapani. Residual active granzyme B in cathepsin C–null lymphocytes is sufficient for perforin-dependent target cell apoptosis. Journal of Cell Biology, 176(4):425–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/4/425.

Sarmiento:2008:WFM

[SWD^{+}08] Corina Sarmiento, Weigang Wang, Athanassios Dovas, Hideki Yamaguchi, Mazen Sidani, Mirvat El-Sibai, Vera DesMarais, Holly A. Holman, Susan Kitchen, Jonathan M. Backer, Art Alberts, and John Condeelis. WASP family members and formin proteins coordinate regulation of cell protrusions in carcinoma cells. Journal of Cell Biology, 180(6):1245–??, March 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/180/6/1245.

Stramer:2005:LIW

[SWG^{+}05] Brian Stramer, Will Wood, Michael J. Galko, Michael J. Redd, Antonio Jacinto, Susan M. Parkhurst, and Paul Martin. Live imaging of wound inflammation in Drosophila embryos reveals key roles for small GTPases during in vivo cell migration. Journal of Cell Biology, 168(4):567–??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/168/4/567.

Sukocheva:2006:ETE

[SWH^{+}06] Olga Sukocheva, Carol Wadham, Andrew Holmes, Nathaniel Albanese, Emily Verrier, Feng Feng, Alex Bernal, Claudia K. Derian, Axel Ullrich, Mathew A. Vadas, and Pu Xia. Estrogen transactivates EGFR via the sphingosine 1-phosphate receptor Edg-3: the role of sphingosine kinase-1. Journal of Cell Biology, 173(2):301–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/2/301.

Sturge:2006:EGL

[SWI06] Justin Sturge, Dirk Wienke, and Clare M. Isacke. Endosomes generate localized Rho–ROCK–MLC2-based contractile signals via Endo180 to promote adhesion disassembly. Journal of Cell Biology, 175(2):337–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/2/337.

Szajner:2005:ESS

[SWL^{+}05] Patricia Szajner, Andrea S. Weisberg, Jacob Lebowitz, John Heuser, and Bernard Moss. External scaffold of spherical immature poxvirus particles is made of protein trimers, forming a honeycomb lattice. Journal of Cell Biology, 170(6): 971–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/6/971.

Szoor:2006:PTP

[SWM^{+}06] Balázs Szöőr, Jude Wilson, Helen McElhinney, Lydia Tabernero, and Keith R. Matthews. Protein tyrosine phosphatase Tb PTP1: a molecular switch controlling life cycle differentiation in trypanosomes. Journal of Cell Biology, 175(2):293–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/175/2/
293.

Sidani:2007:CDM

[SWM^{+}07] Mazen Sidani, Deborah Wessels, Ghassan Mouneimne, Mousumi Ghosh, Sumanta Goswami, Corina Sarmiento, Weigang Wang, Spencer Kuhl, Mirvat El-Sibai, Jonathan M. Backer, Robert Eddy, David Soll, and John Condeelis. Cofilin determines the migration behavior and turning frequency of metastatic cancer cells. Journal of Cell Biology, 179(4):777–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/179/4/777.

Skinner:2008:MPU

[SWS^{+}08] John J. Skinner, Stacey Wood, James Shorter, S. Walter Englander, and Ben E. Black. The Mad2 partial unfolding model: regulating mitosis through Mad2 conformational switching. Journal of Cell Biology, 183(5):761–??, December 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/183/5/761.

Seelenmeyer:2005:CSC

[SWT^{+}05] Claudia Seelenmeyer, Sabine Wegehingel, Ivo Tews, Markus Künzler, Markus Aebi, and Walter Nickel. Cell surface counter receptors are essential components of the unconventional export machinery of galectin-1. Journal of Cell Biology, 171(2):373–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
171/2/373.

Shen:2008:TJP

[SWT08] Le Shen, Christopher R. Weber, and Jerrold R. Turner. The tight junction protein complex undergoes rapid and continuous molecular remodeling at steady state. Journal of Cell Biology, 181(4):683–??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/4/683.

Seed:2008:SIS

[SX08] Brian Seed and Ramnik Xavier. Spinophilin and the immune synapse. Journal of Cell Biology, 181(2):181–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/181/2/181.

Shim:2006:CEL

[SXH^{+}06] Jae-Hyuck Shim, Changchun Xiao, Matthew S. Hayden, Ki-Young Lee, E. Sergio Trombetta, Marc Pypaert, Atsuki Nara, Tamotsu Yoshimori, Bettina Wilm, Hediye Erdjument-Bromage, Paul Tempst, Brigid L. M. Hogan, Ira Mellman, and Sankar Ghosh. CHMP5 is essential for late endosome function and down-regulation of receptor signaling during mouse embryogenesis. Journal of Cell Biology, 172(7):1045–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/172/7/
1045.

Shang:2008:PFI

[SYA^{+}08] Yueting Shang, Tadashi Yoshida, Brad A. Amendt, James F. Martin, and Gary K. Owens. Pitx2 is functionally important in the early stages of vascular smooth muscle cell differentiation. Journal of Cell Biology, 181(3):461–??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/181/3/461.

Schramp:2008:EPS

[SYKM08] Mark Schramp, Olivia Ying, Tai Young Kim, and G. Steven Martin. ERK5 promotes Src-induced podosome formation by limiting Rho activation. Journal of Cell Biology, 181(7):1195– ??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
181/7/1195.

Sakisaka:2008:DIS

[SYM^{+}08a] Toshiaki Sakisaka, Yasunori Yamamoto, Sumiko Mochida, Michiko Nakamura, Kouki Nishikawa, Hiroyoshi Ishizaki, Miki Okamoto-Tanaka, Jun Miyoshi, Yoshinori Fujiyoshi, Toshiya Manabe, and Yoshimi Takai. Dual inhibition of SNARE complex formation by tomosyn ensures controlled neurotransmitter release. Journal of Cell Biology, 183(2):323–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/183/2/323.

Soto:2008:PCI

[SYM^{+}08b] Edwin Soto, Masahiro Yanagisawa, Laura A. Marlow, John A. Copland, Edith A. Perez, and Panos Z. Anastasiadis. p120 catenin induces opposing effects on tumor cell growth depending on E-cadherin expression. Journal of Cell Biology, 183(4): 737–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/4/737.

Sakurai-Yageta:2008:IIE

[SYRL^{+}08] Mika Sakurai-Yageta, Chiara Recchi, Gaëlle Le Dez, Jean-Baptiste Sibarita, Laurent Daviet, Jacques Camonis, Crislyn D’Souza-Schorey, and Philippe Chavrier. The interaction of IQGAP1 with the exocyst complex is required for tumor cell invasion downstream of Cdc42 and RhoA. Journal of Cell Biology, 181(6):985–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/6/985.

Shen:2009:CTP

[SZL^{+}09] Shensi Shen, Pengtao Zhang, Martin A. Lovchik, Ying Li, Liuya Tang, Zhimin Chen, Rong Zeng, Dawei Ma, Junying Yuan, and Qiang Yu. Cyclodepsipeptide toxin promotes the degradation of Hsp90 client proteins through chaperone-mediated autophagy.

Journal of Cell Biology, 185(4):629–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/185/4/629.

Springer:2008:SBD

[SZX08] Timothy A. Springer, Jianghai Zhu, and Tsan Xiao. Structural basis for distinctive recognition of fibrinogen γC peptide by the platelet integrin α_{IIb} β_{3}. Journal of Cell Biology, 182(4):791–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
182/4/791.

Tsai:2007:MIA

[TA07] Miao-Chih Tsai and Julie Ahringer. Microtubules are involved in anterior-posterior axis formation in C. elegans embryos. Journal of Cell Biology, 179(3):397–??, November 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/179/3/397.

Takeichi:2006:STL

[Tak06] Masatoshi Takeichi. Shoichiro Tsukita: a life exploring the molecular architecture of the tight junction. Journal of Cell Biology, 172(3):321–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/3/321.

Thompson:2009:IPH

[TAK^{+}09] Leslie Michels Thompson, Charity T. Aiken, Linda S. Kaltenbach, Namita Agrawal, Katalin Illes, Ali Khoshnan, Marta Martinez-Vincente, Montserrat Arrasate, Jacqueline Gire O’Rourke, Hasan Khashwji, Tamas Lukacsovich, Ya-Zhen Zhu, Alice L. Lau, Ashish Massey, Michael R. Hayden, Scott O. Zeitlin, Steven Finkbeiner, Kim N. Green, Frank M. LaFerla, Gillian Bates, Lan Huang, Paul H. Patterson, Donald C. Lo, Ana Maria Cuervo, J. Lawrence Marsh, and Joan S. Steffan. IKK phosphorylates Huntingtin and targets it for degradation by the proteasome and lysosome. Journal of Cell Biology, 187(7):1083–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/187/7/1083.

Tan:2006:KFR

[TAM^{+}06] Dongyan Tan, Ana B. Asenjo, Vito Mennella, David J. Sharp, and Hernando Sosa. Kinesin-13s form rings around microtubules. Journal of Cell Biology, 175(1):25–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/175/1/25.

Thompson:2008:ELB

[TC08] Sarah L. Thompson and Duane A. Compton. Examining the link between chromosomal instability and aneuploidy in human cells. Journal of Cell Biology, 180(4):665–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/180/4/665.

Todorovicc:2005:MPC

[TCHL05] Viktor Todorovicç, Chih-Chiun Chen, Nissim Hay, and Lester F. Lau. The matrix protein CCN1 (CYR61) induces apoptosis in fibroblasts. Journal of Cell Biology, 171(3):559– ??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/3/559.

Tsai:2005:LRI

[TCKV05] Jin-Wu Tsai, Yu Chen, Arnold R. Kriegstein, and Richard B. Vallee. LIS1 RNA interference blocks neural stem cell division, morphogenesis, and motility at multiple stages. Journal of Cell Biology, 170(6):935–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/170/6/935.

Taggart:2006:SLI

[TCW^{+}06] Clifford C. Taggart, Sally-Ann Cryan, Sinead Weldon, Aileen Gibbons, Catherine M. Greene, Emer Kelly, Teck Boon Low, Shane J. O’Neill, and Noel G. McElvaney. Secretory leucoprotease inhibitor binds to NF-κB binding sites in monocytes and inhibits p65 binding. Journal of Cell Biology, 172(1):??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
172/1/i2.

Tsai:2008:ISE

[TD08] Richard K. Tsai and Dennis E. Discher. Inhibition of “self” engulfment through deactivation of myosin–II at the phagocytic synapse between human cells. Journal of Cell Biology, 180 (5):989–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/5/989.

tomDieck:2005:MDP

[tDAK^{+}05] Susanne tom Dieck, Wilko D. Altrock, Michael M. Kessels, Britta Qualmann, Hanna Regus, Dana Brauner, Anna Fejtová, Oliver Bracko, Eckart D. Gundelfinger, and Johann H. Brandstätter. Molecular dissection of the photoreceptor ribbon synapse. Journal of Cell Biology, 168(5):825–??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/168/5/
825.

Tackett:2005:PGC

[TDD^{+}05] Alan J. Tackett, David J. Dilworth, Megan J. Davey, Michael O’Donnell, John D. Aitchison, Michael P. Rout, and Brian T. Chait. Proteomic and genomic characterization of chromatin complexes at a boundary. Journal of Cell Biology, 169(1):35–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
169/1/35.

Trajkovic:2006:NGS

[TDG^{+}06] Katarina Trajkovic, Ajit Singh Dhaunchak, José T. Goncalves, Dirk Wenzel, Anja Schneider, Gertrude Bunt, Klaus-Armin Nave, and Mikael Simons. Neuron to glia signaling triggers myelin membrane exocytosis from endosomal storage sites. Journal of Cell Biology, 172(6):937–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/172/6/937.

Terrin:2006:PSH

[TDP^{+}06] Anna Terrin, Giulietta Di Benedetto, Vanessa Pertegato, York-Fong Cheung, George Baillie, Martin J. Lynch, Nicola Elvassore, Anke Prinz, Friedrich W. Herberg, Miles D. Houslay, and Manuela Zaccolo. PGE_{1} stimulation of HEK293 cells generates multiple contiguous domains with different [cAMP]: role of compartmentalized phosphodiesterases. Journal of Cell Biology, 175(3):441–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/3/441.

Tovey:2008:SCT

[TDT^{+}08] Stephen C. Tovey, Skarlatos G. Dedos, Emily J. A. Taylor, Jarrod E. Church, and Colin W. Taylor. Selective coupling of type 6 adenylyl cyclase with type 2 IP_{3} receptors mediates direct sensitization of IP_{3} receptors by cAMP. Journal of Cell Biology, 183(2):297–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/2/297.

Tait:2007:AIB

[TdVM^{+}07] Stephen W. G. Tait, Evert de Vries, Chiel Maas, Anna M. Keller, Clive S. D’Santos, and Jannie Borst. Apoptosis induction by Bid requires unconventional ubiquitination and degradation of its N-terminal fragment. Journal of Cell Biology, 179(7):1453–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/7/1453.

Tamura:2009:UUA

[TEIS09] Yasushi Tamura, Toshiya Endo, Miho Iijima, and Hiromi Sesaki. Ups1p and Ups2p antagonistically regulate cardiolipin metabolism in mitochondria. Journal of Cell Biology, 185 (6):1029–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/6/1029.

Terada:2006:SRO

[Ter06] Lance S. Terada. Specificity in reactive oxidant signaling: think globally, act locally. Journal of Cell Biology, 174(5):615–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
174/5/615.

Therizols:2006:C

[TFC^{+}06a] Pierre Therizols, Cécile Fairhead, Ghislain G. Cabal, Auguste Genovesio, Jean-Christophe Olivo-Marin, Bernard Dujon, and Emmanuelle Fabre. Correction. Journal of Cell Biology, 172 (6):951–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/6/951.

Therizols:2006:TTN

[TFC^{+}06b] Pierre Therizols, Cécile Fairhead, Ghislain G. Cabal, Auguste Genovesio, Jean-Christophe Olivo-Marin, Bernard Dujon, and Emmanuelle Fabre. Telomere tethering at the nuclear periphery is essential for efficient DNA double strand break repair in subtelomeric region. Journal of Cell Biology, 172(2):189–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
172/2/189.

Thorne:2005:TWS

[TG05] Meghan E. Thorne and Cara J. Gottardi. Terminating Wnt signals. Journal of Cell Biology, 171(5):761–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/171/5/761.

Tong:2007:APC

[TGH^{+}07] Zongtian Tong, Xiang-Dong Gao, Audrey S. Howell, Indrani Bose, Daniel J. Lew, and Erfei Bi. Adjacent positioning of cellular structures enabled by a Cdc42 GTPase-activating proteinmediated zone of inhibition. Journal of Cell Biology, 179(7): 1375–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/7/1375.

Tsuruta:2009:PRC

[TGRD09] Fuminori Tsuruta, Eric M. Green, Matthieu Rousset, and Ricardo E. Dolmetsch. PIKfyve regulates Ca_{V} 1.2 degradation and prevents excitotoxic cell death. Journal of Cell Biology, 187 (2):279–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/187/2/279.

Thomas:2006:ADS

[THL^{+}06] Steven G. Thomas, Shanjin Huang, Shutian Li, Christopher J. Staiger, and Vernonica E. Franklin-Tong. Actin depolymerization is sufficient to induce programmed cell death in selfincompatible pollen. Journal of Cell Biology, 174(2):221–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
174/2/221.

Tolg:2006:RFD

[THN^{+}06] Cornelia Tolg, Sara R. Hamilton, Kerry-Ann Nakrieko, Fatemeh Kooshesh, Paul Walton, James B. McCarthy, Mina J. Bissell, and Eva A. Turley. Rhamm^{−/−} fibroblasts are defective in CD44-mediated ERK1,2 motogenic signaling, leading to defective skin wound repair. Journal of Cell Biology, 175(6): 1017–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/6/1017.

Thomas:2006:CBI

[Tho06] Wendy Thomas. For catch bonds, it all hinges on the interdomain region. Journal of Cell Biology, 174(7):911–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/174/7/
911.

Tamura:2009:TTP

[THS^{+}09] Yasushi Tamura, Yoshihiro Harada, Takuya Shiota, Koji Yamano, Kazuaki Watanabe, Mihoko Yokota, Hayashi Yamamoto, Hiromi Sesaki, and Toshiya Endo. Tim23–Tim50 pair coordinates functions of translocators and motor proteins in mitochondrial protein import. Journal of Cell Biology, 184(1): 129–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/184/1/129.

Thummel:2007:RFH

[Thu07] Carl S. Thummel. To die or not to die — a role for Fork head. Journal of Cell Biology, 176(6):737–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/176/6/737.

Tamura:2006:ITM

[THY^{+}06] Yasushi Tamura, Yoshihiro Harada, Koji Yamano, Kazuaki Watanabe, Daigo Ishikawa, Chié Ohshima, Shuh ichi Nishikawa, Hayashi Yamamoto, and Toshiya Endo. Identification of Tam41 maintaining integrity of the TIM23 protein translocator complex in mitochondria. Journal of Cell Biology, 174(5):631–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
174/5/631.

Tang:2006:BEI

[TJC06] Xie Tang, Ye Jin, and W. Zacheus Cande. Bqt2p is essential for initiating telomere clustering upon pheromone sensing in fission yeast. Journal of Cell Biology, 173(6):845–??, June 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/173/6/845.

Takaesu:2006:APM

[TKB^{+}06] Giichi Takaesu, Jong-Sun Kang, Gyu-Un Bae, Min-Jeong Yi, Clement M. Lee, E. Premkumar Reddy, and Robert S. Krauss. Activation of p38α /β MAPK in myogenesis via binding of the scaffold protein JLP to the cell surface protein Cdo. Journal of Cell Biology, 175(3):383–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/175/3/383.

Tamura:2005:AEK

[TKH^{+}05] Atsushi Tamura, Shojiro Kikuchi, Masaki Hata, Tatsuya Katsuno, Takeshi Matsui, Hisayoshi Hayashi, Yuichi Suzuki, Tetsuo Noda, Shoichiro Tsukita, and Sachiko Tsukita. Achlorhydria by ezrin knockdown. Journal of Cell Biology, 169(1):21–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/169/1/
21.

tenKlooster:2006:TAR

[tKJCH06] Jean Paul ten Klooster, Zahara M. Jaffer, Jonathan Chernoff, and Peter L. Hordijk. Targeting and activation of Rac1 are mediated by the exchange factor β-Pix. Journal of Cell Biology, 172(5):759–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/5/759.

Taxis:2005:SNC

[TKK^{+}05] Christof Taxis, Philipp Keller, Zaharoula Kavagiou, Lars Juhl Jensen, Julien Colombelli, Peer Bork, Ernst H. K. Stelzer, and Michael Knop. Spore number control and breeding in Saccharomyces cerevisiae. Journal of Cell Biology, 171(4): 627–??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/4/627.

Tanaka:2007:MMM

[TKKT07] Kozo Tanaka, Etsushi Kitamura, Yoko Kitamura, and Tomoyuki U. Tanaka. Molecular mechanisms of microtubuledependent kinetochore transport toward spindle poles. Journal of Cell Biology, 178(2):269–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/178/2/269.

Thein:2007:ARM

[TKSNG07] Kerstin H. Thein, Julia Kleylein-Sohn, Erich A. Nigg, and Ulrike Gruneberg. Astrin is required for the maintenance of sister chromatid cohesion and centrosome integrity. Journal of Cell Biology, 178(3):345–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/3/345.

Tan:2008:MPD

[TLDB08] Shengjiang Tan, Ekaterina Lyulcheva, Jon Dean, and Daimark Bennett. Mars promotes dTACC dephosphorylation on mitotic spindles to ensure spindle stability. Journal of Cell Biology, 182(1):27–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/1/27.

Treanor:2006:MIK

[TLK^{+}06] Bebhinn Treanor, Peter M. P. Lanigan, Sunil Kumar, Chris Dunsby, Ian Munro, Egidijus Auksorius, Fiona J. Culley, Marco A. Purbhoo, David Phillips, Mark A. A. Neil, Deborah N. Burshtyn, Paul M. W. French, and Daniel M. Davis. Microclusters of inhibitory killer immunoglobulin-like receptor signaling at natural killer cell immunological synapses. Journal of Cell Biology, 174(1):153–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/174/1/153.

Takahashi:2007:RMF

[TLM^{+}07] Seiichiro Takahashi, Michael Leiss, Markus Moser, Tomoo Ohashi, Tomoe Kitao, Dominik Heckmann, Alexander Pfeifer, Horst Kessler, Junichi Takagi, Harold P. Erickson, and Reinhard Fässler. The RGD motif in fibronectin is essential for development but dispensable for fibril assembly. Journal of Cell Biology, 178(1):167–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/1/167.

Tsai:2005:MGD

[TM05] Robert Y. L. Tsai and Ronald D. G. McKay. A multistep, GTP-driven mechanism controlling the dynamic cycling of nucleostemin. Journal of Cell Biology, 168(2):179–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/168/2/
179.

Titorenko:2006:PBP

[TM06] Vladimir I. Titorenko and Robert T. Mullen. Peroxisome biogenesis: the peroxisomal endomembrane system and the role of the ER. Journal of Cell Biology, 174(1):11–??, July 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/174/1/11.

Trinkle-Mulcahy:2006:RMR

[TMAL^{+}06] Laura Trinkle-Mulcahy, Jens Andersen, Yun Wah Lam, Greg Moorhead, Matthias Mann, and Angus I. Lamond. Repo-man recruits PP1γ to chromatin and is essential for cell viability. Journal of Cell Biology, 172(5):679–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/172/5/679.

Trinkle-Mulcahy:2008:ISP

[TMBL^{+}08] Laura Trinkle-Mulcahy, Séverine Boulon, Yun Wah Lam, Roby Urcia, François-Michel Boisvert, Franck Vandermoere, Nick A. Morrice, Sam Swift, Ulrich Rothbauer, Heinrich Leonhardt, and Angus Lamond. Identifying specific protein interaction partners using quantitative mass spectrometry and bead proteomes. Journal of Cell Biology, 183(2):223–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/183/2/223.

Tagawa:2005:ATC

[TMH^{+}05] Akiko Tagawa, Anna Mezzacasa, Arnold Hayer, Andrea Longatti, Lucas Pelkmans, and Ari Helenius. Assembly and trafficking of caveolar domains in the cell. Journal of Cell Biology, 170(5):769–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/5/769.

Togashi:2006:IAR

[TMH^{+}06] Hideru Togashi, Jun Miyoshi, Tomoyuki Honda, Toshiaki Sakisaka, Yoshimi Takai, and Masatoshi Takeichi. Interneurite affinity is regulated by heterophilic nectin interactions in concert with the cadherin machinery. Journal of Cell Biology, 174 (1):141–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/1/141.

Tardat:2007:PSD

[TMH^{+}07] Mathieu Tardat, Rabih Murr, Zdenko Herceg, Claude Sardet, and Eric Julien. PR-Set7-dependent lysine methylation ensures genome replication and stability through S phase. Journal of Cell Biology, 179(7):1413–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/7/1413.

Tokuo:2007:MAM

[TMI07] Hiroshi Tokuo, Katsuhide Mabuchi, and Mitsuo Ikebe. The motor activity of myosin–X promotes actin fiber convergence at the cell periphery to initiate filopodia formation. Journal of Cell Biology, 179(2):229–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/2/229.

Tomita:2005:CNC

[TMW^{+}05] Yuichi Tomita, Keisuke Matsumura, Yoshio Wakamatsu, Yumi Matsuzaki, Isao Shibuya, Haruko Kawaguchi, Masaki Ieda, Sachiko Kanakubo, Takuya Shimazaki, Satoshi Ogawa, Noriko Osumi, Hideyuki Okano, and Keiichi Fukuda. Cardiac neural crest cells contribute to the dormant multipotent stem cell in the mammalian heart. Journal of Cell Biology, 170(7): 1135–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/7/1135.

Tzaban:2009:RTP

[TMY^{+}09] Salit Tzaban, Ramiro H. Massol, Elizabeth Yen, Wendy Hamman, Scott R. Frank, Lynne A. Lapierre, Steen H. Hansen, James R. Goldenring, Richard S. Blumberg, and Wayne I. Lencer. The recycling and transcytotic pathways for IgG transport by FcRn are distinct and display an inherent polarity.

Journal of Cell Biology, 185(4):673–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/185/4/673.

Tibelius:2009:MPR

[TMZ^{+}09] Alexandra Tibelius, Joachim Marhold, Hanswalter Zentgraf, Christoph E. Heilig, Heidemarie Neitzel, Bernard Ducommun, Anita Rauch, Anthony D. Ho, Jiri Bartek, and Alwin Krämer. Microcephalin and pericentrin regulate mitotic entry via centrosome-associated Chk1. Journal of Cell Biology, 185 (7):1149–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/7/1149.

Tripathi:2007:BHD

[TNS07] Vivek Tripathi, Tirunelvely Nagarjuna, and Sagar Sengupta. BLM helicase-dependent and –independent roles of 53BP1 during replication stress-mediated homologous recombination. Journal of Cell Biology, 178(1):9–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/178/1/9.

Tararuk:2006:CP

[TÖL^{+}06a] Tatsiana Tararuk, Nina Östman, Wenrui Li, Benny Björkblom, Artur Padzik, Justyna Zdrojewska, Vesa Hongisto, Thomas Herdegen, Witold Konopka, Michael J. Courtney, and Eleanor T. Coffey. Correction 173, p. 265. Journal of Cell Biology, 173(5):821–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/5/821.

Tararuk:2006:JPS

[TÖL^{+}06b] Tatsiana Tararuk, Nina Östman, Wenrui Li, Benny Björkblom, Artur Padzik, Justyna Zdrojewska, Vesa Hongisto, Thomas Herdegen, Witold Konopka, Michael J. Courtney, and Eleanor T. Coffey. JNK1 phosphorylation of SCG10 determines microtubule dynamics and axodendritic length. Journal of Cell Biology, 173(2):265–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/173/2/265.

Tomishige:2008:AMK

[Tom08] Michio Tomishige. Activation of mitotic kinesin by microtubule bundling. Journal of Cell Biology, 182(3):417–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/182/3/417.

Toyo-oka:2008:PPC

[ToMY^{+}08] Kazuhito Toyo-oka, Daisuke Mori, Yoshihisa Yano, Masayuki Shiota, Hiroshi Iwao, Hidemasa Goto, Masaki Inagaki, Noriko Hiraiwa, Masami Muramatsu, Anthony Wynshaw-Boris, Atsushi Yoshiki, and Shinji Hirotsune. Protein phosphatase 4 catalytic subunit regulates Cdk1 activity and microtubule organization via NDEL1 dephosphorylation. Journal of Cell Biology, 180(6):1133–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/6/1133.

Toomre:2005:SIR

[Too05] Derek Toomre. Spying on IgE receptor signaling. Journal of Cell Biology, 171(3):415–??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/171/3/415.

Thompson:2009:RRC

[TP09] Debrah M. Thompson and Roy Parker. The RNase Rny1p cleaves tRNAs and promotes cell death during oxidative stress in Saccharomyces cerevisiae. Journal of Cell Biology, 185 (1):43–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/1/43.

Tian:2008:CSR

[TPA^{+}08] Yong Tian, Caroline Pate, Alberto Andreolotti, Limin Wang, Elaine Tuomanen, Kelli Boyd, Enrique Claro, and Suzanne Jackowski. Cytokine secretion requires phosphatidylcholine synthesis. Journal of Cell Biology, 181(6):945–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/181/6/945.

Tamada:2007:TDMa

[TPNS07a] Masako Tamada, Tomas D. Perez, W. James Nelson, and Michael P. Sheetz. Two distinct modes of myosin assembly and dynamics during epithelial wound closure. Journal of Cell Biology, 176(1):27–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/1/27.

Tamada:2007:TDMb

[TPNS07b] Masako Tamada, Tomas D. Perez, W. James Nelson, and Michael P. Sheetz. Two distinct modes of myosin assembly and dynamics during epithelial wound closure. Journal of Cell Biology, 176(4):545–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/4/545.

Torres-Padilla:2006:RTM

[TPZG06] Maria Elena Torres-Padilla and Magdalena Zernicka-Goetz. Role of TIF1α as a modulator of embryonic transcription in the mouse zygote. Journal of Cell Biology, 174(3):329–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/174/3/
329.

Tien:2009:NU

[TRB09] An-Chi Tien, Akhila Rajan, and Hugo J. Bellen. A Notch updated. Journal of Cell Biology, 184(5):621–??, March 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/184/5/621.

Tang:2008:DDE

[TRC^{+}08] Kai-Fu Tang, Hong Ren, Jia Cao, Gui-Li Zeng, Jing Xie, Min Chen, Lu Wang, and Cai-Xia He. Decreased Dicer expression elicits DNA damage and up-regulation of MICA and MICB. Journal of Cell Biology, 182(2):233–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/182/2/233.

Tien:2008:ETO

[TRS^{+}08] An-Chi Tien, Akhila Rajan, Karen L. Schulze, Hyung Don Ryoo, Melih Acar, Hermann Steller, and Hugo J. Bellen. Ero1L, a thiol oxidase, is required for Notch signaling through cysteine bridge formation of the Lin12-Notch repeats in Drosophila melanogaster. Journal of Cell Biology, 182(6):1113–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/182/6/
1113.

Tytell:2006:AKM

[TS06] Jessica D. Tytell and Peter K. Sorger. Analysis of kinesin motor function at budding yeast kinetochores. Journal of Cell Biology, 172(6):861–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/6/861.

Trelles-Sticken:2005:MTC

[TSALS05] Edgar Trelles-Sticken, Caroline Adelfalk, Josef Loidl, and Harry Scherthan. Meiotic telomere clustering requires actin for its formation and cohesin for its resolution. Journal of Cell Biology, 170(2):213–??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/2/213.

Toriyama:2006:SPI

[TSK^{+}06] Michinori Toriyama, Tadayuki Shimada, Ki Bum Kim, Mari Mitsuba, Eiko Nomura, Kazuhiro Katsuta, Yuichi Sakumura, Peter Roepstorff, and Naoyuki Inagaki. Shootin1: a protein involved in the organization of an asymmetric signal for neuronal polarization. Journal of Cell Biology, 175(1):147–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/175/1/
147.

Takizawa:2006:SMF

[TSN^{+}06] Norio Takizawa, Tara C. Smith, Thomas Nebl, Jessica L. Crowley, Stephen J. Palmieri, Lawrence M. Lifshitz, Anka G. Ehrhardt, Laura M. Hoffman, Mary C. Beckerle, and Elizabeth J. Luna. Supervillin modulation of focal adhesions involving TRIP6/ZRP-1. Journal of Cell Biology, 174(3):447–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/174/3/
447.

Tian:2007:ANR

[TSN^{+}07] Li Tian, Michael Stefanidakis, Lin Ning, Philippe Van Lint, Henrietta Nyman-Huttunen, Claude Libert, Shigeyoshi Itohara, Masayoshi Mishina, Heikki Rauvala, and Carl G. Gahmberg. Activation of NMDA receptors promotes dendritic spine development through MMP-mediated ICAM-5 cleavage. Journal of Cell Biology, 178(4):687–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/178/4/687.

Torrecilla:2007:PRG

[TSP^{+}07] Ignacio Torrecilla, Elizabeth J. Spragg, Benoit Poulin, Phillip J. McWilliams, Sharad C. Mistry, Andree Blaukat, and Andrew B. Tobin. Phosphorylation and regulation of a G protein–coupled receptor by protein kinase CK2. Journal of Cell Biology, 177 (1):127–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/1/127.

Trichet:2008:RAC

[TSP08] Léa Trichet, Cécile Sykes, and Julie Plastino. Relaxing the actin cytoskeleton for adhesion and movement with Ena/VASP. Journal of Cell Biology, 181(1):19–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/181/1/19.

Tsujita:2006:CBA

[TSS^{+}06] Kazuya Tsujita, Shiro Suetsugu, Nobunari Sasaki, Masahiro Furutani, Tsukasa Oikawa, and Tadaomi Takenawa. Coordination between the actin cytoskeleton and membrane deformation by a novel membrane tubulation domain of PCH proteins is involved in endocytosis. Journal of Cell Biology, 172(2):269–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
172/2/269.

Tighe:2008:MKA

[TST08] Anthony Tighe, Oliver Staples, and Stephen Taylor. Mps1 kinase activity restrains anaphase during an unperturbed mitosis and targets Mad2 to kinetochores. Journal of Cell Biology, 181(6):893–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/6/893.

Tanabe:2009:DIM

[TT09] Kenji Tanabe and Kohji Takei. Dynamic instability of microtubules requires dynamin 2 and is impaired in a Charcot–Marie– Tooth mutant. Journal of Cell Biology, 185(6):939–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/185/6/
939.

Tuvia:2007:UEL

[TTE^{+}07] Shmuel Tuvia, Daniel Taglicht, Omri Erez, Iris Alroy, Iris Alchanati, Vivian Bicoviski, Mally Dori-Bachash, Danny Ben-Avraham, and Yuval Reiss. The ubiquitin E3 ligase POSH regulates calcium homeostasis through spatial control of Herp. Journal of Cell Biology, 177(1):51–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/177/1/51.

Teckchandani:2009:QPI

[TTG^{+}09] Anjali Teckchandani, Natalie Toida, Jake Goodchild, Christine Henderson, Julian Watts, Bernd Wollscheid, and Jonathan A. Cooper. Quantitative proteomics identifies a Dab2/integrin module regulating cell migration. Journal of Cell Biology, 186(1):99–??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/186/1/99.

Teis:2006:PMM

[TTK^{+}06] David Teis, Nicole Taub, Robert Kurzbauer, Diana Hilber, Mariana E. de Araujo, Miriam Erlacher, Martin Offterdinger, Andreas Villunger, Stephan Geley, Georg Bohn, Christoph Klein, Michael W. Hess, and Lukas A. Huber. p14–MP1-MEK1 signaling regulates endosomal traffic and cellular proliferation during tissue homeostasis. Journal of Cell Biology, 175(6):861–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
175/6/861.

Timney:2006:SKR

[TTNA^{+}06] Benjamin L. Timney, Jaclyn Tetenbaum-Novatt, Diana S. Agate, Rosemary Williams, Wenzhu Zhang, Brian T. Chait, and Michael P. Rout. Simple kinetic relationships and nonspecific competition govern nuclear import rates in vivo. Journal of Cell Biology, 175(4):579–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/175/4/579.

Tahara:2008:ISG

[TTO^{+}08] Kiyoshi Tahara, Masatoshi Takagi, Miho Ohsugi, Takefumi Sone, Fumiko Nishiumi, Kazuhiro Maeshima, Yasuomi Horiuchi, Noriko Tokai-Nishizumi, Fumio Imamoto, Tadashi Yamamoto, Shingo Kose, and Naoko Imamoto. Importin-β and the small guanosine triphosphatase Ran mediate chromosome loading of the human chromokinesin Kid. Journal of Cell Biology, 180(3):493–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/3/493.

Tallada:2009:PMR

[TTYH09] Victor A. Tallada, Kenji Tanaka, Mitsuhiro Yanagida, and Iain M. Hagan. The S. pombe mitotic regulator Cut12 promotes spindle pole body activation and integration into the nuclear envelope. Journal of Cell Biology, 185(5):875–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/185/5/
875.

Tuma:2005:AN

[Tum05a] Rabiya S. Tuma. Adaptor to nucleus. Journal of Cell Biology, 170(2):??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/2/162.3.

Tuma:2005:ADS

[Tum05b] Rabiya S. Tuma. Agrin drives synaptogenesis. Journal of Cell Biology, 169(3):??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/3/374.2.

Tuma:2005:AUS

[Tum05c] Rabiya S. Tuma. Anchoring under shear stress. Journal of Cell Biology, 171(6):??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/6/913.2.

Tuma:2005:AJC

[Tum05d] Rabiya S. Tuma. Another job for β-catenin. Journal of Cell Biology, 168(4):??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/4/521.1.

Tuma:2005:AC

[Tum05e] Rabiya S. Tuma. Asymmetry in cytokinesis. Journal of Cell Biology, 171(2):??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/2/191.1.

Tuma:2005:BBT

[Tum05f] Rabiya S. Tuma. Basal body timing. Journal of Cell Biology, 169(5):??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/5/700.1.

Tuma:2005:BUP

[Tum05g] Rabiya S. Tuma. Beating under pressure. Journal of Cell Biology, 168(6):??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/6/849.1.

Tuma:2005:BP

[Tum05h] Rabiya S. Tuma. Boundary proteomics. Journal of Cell Biology, 169(1):??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/1/15.2.

Tuma:2005:CDS

[Tum05i] Rabiya S. Tuma. Calcium, direct from the source. Journal of Cell Biology, 170(7):??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/170/7/1014.3.

Tuma:2005:CCb

[Tum05j] Rabiya S. Tuma. Clipping coreceptors. Journal of Cell Biology, 171(4):??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/4/581.2.

Tuma:2005:CBY

[Tum05k] Rabiya S. Tuma. Cohesins and breaks in yeast. Journal of Cell Biology, 170(4):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/4/513.3.

Tuma:2005:CCa

[Tum05l] Rabiya S. Tuma. Controlling cadherin. Journal of Cell Biology, 170(5):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/5/697.2.

Tuma:2005:CDP

[Tum05m] Rabiya S. Tuma. Controlling division planes. Journal of Cell Biology, 171(6):??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/6/912.1.

Tuma:2005:C

[Tum05n] Rabiya S. Tuma. Correction. Journal of Cell Biology, 170 (1):153–??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/1/153.

Tuma:2005:CS

[Tum05o] Rabiya S. Tuma. Counting spores. Journal of Cell Biology, 171 (4):??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/4/580.2.

Tuma:2005:DF

[Tum05p] Rabiya S. Tuma. Decondensation at the fork. Journal of Cell Biology, 168(6):??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/6/849.3.

Tuma:2005:DTS

[Tum05q] Rabiya S. Tuma. The degranulation two-step. Journal of Cell Biology, 170(1):??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/1/9.3.

Tuma:2005:DHS

[Tum05r] Rabiya S. Tuma. Dynein holds still. Journal of Cell Biology, 170(4):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/4/513.2.

Tuma:2005:ES

[Tum05s] Rabiya S. Tuma. Eating Staph. Journal of Cell Biology, 170(3):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/3/335.2.

Tuma:2005:ECF

[Tum05t] Rabiya S. Tuma. ECM controls flow-induced NF-κB. Journal of Cell Biology, 169(1):??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/1/14.2.

Tuma:2005:EAA

[Tum05u] Rabiya S. Tuma. Evidence for AMPAR association. Journal of Cell Biology, 169(3):??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/3/375.2.

Tuma:2005:ESS

[Tum05v] Rabiya S. Tuma. Exocytosis in small steps. Journal of Cell Biology, 168(6):??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/6/848.2.

Tuma:2005:EER

[Tum05w] Rabiya S. Tuma. Extra exocytosis rescues polarity. Journal of Cell Biology, 170(2):??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/2/162.2.

Tuma:2005:FD

[Tum05x] Rabiya S. Tuma. Fission and disease. Journal of Cell Biology, 170(7):??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/7/1014.2.

Tuma:2005:FWC

[Tum05y] Rabiya S. Tuma. Fragile when condensed. Journal of Cell Biology, 171(6):??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/6/915.2.

Tuma:2005:GPF

[Tum05z] Rabiya S. Tuma. G proteins find partners. Journal of Cell Biology, 171(6):??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/6/915.1.

Tuma:2005:GPR

[Tum05-27] Rabiya S. Tuma. A gas pedal for replication? Journal of Cell Biology, 171(6):??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/6/912.2.

Tuma:2005:GRG

[Tum05-28] Rabiya S. Tuma. GLUT4 ready to go. Journal of Cell Biology, 169(3):??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/3/375.1.

Tuma:2005:GAL

[Tum05-29] Rabiya S. Tuma. Glycolipids anchor laminin. Journal of Cell Biology, 169(1):??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/1/15.1.

Tuma:2005:HPN

[Tum05-30] Rabiya S. Tuma. Hiding proteins in the nucleolus. Journal of Cell Biology, 170(5):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/5/696.1.

Tuma:2005:HIA

[Tum05-31] Rabiya S. Tuma. History’s influence on actin. Journal of Cell Biology, 171(6):??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/6/915.3.

Tuma:2005:HP

[Tum05-32] Rabiya S. Tuma. Holding on to peroxisomes. Journal of Cell Biology, 169(5):??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/5/701.1.

Tuma:2005:ID

[Tum05-33] Rabiya S. Tuma. Immortal DNA. Journal of Cell Biology, 170(5):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/5/696.2.

Tuma:2005:ICM

[Tum05-34] Rabiya S. Tuma. Integrins control migration style. Journal of Cell Biology, 169(3):??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/3/374.1.

Tuma:2005:IDC

[Tum05-35] Rabiya S. Tuma. Interleukins drive the cell cycle. Journal of Cell Biology, 169(5):??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/5/700.2.

Tuma:2005:JTM

[Tum05-36] Rabiya S. Tuma. Junctions tighten myelin sheaths. Journal of Cell Biology, 169(3):??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/3/375.3.

Tuma:2005:JRS

[Tum05-37] Rabiya S. Tuma. Just-right spirals. Journal of Cell Biology, 170(7):??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/7/1014.1.

Tuma:2005:KBE

[Tum05-38] Rabiya S. Tuma. Kinesin at both ends. Journal of Cell Biology, 171(2):??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/2/191.3.

Tuma:2005:LLW

[Tum05-39] Rabiya S. Tuma. Lamellae lead the way. Journal of Cell Biology, 168(4):??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/4/520.1.

Tuma:2005:LPL

[Tum05-40] Rabiya S. Tuma. Less p53 for life. Journal of Cell Biology, 171 (6):??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/6/914.2.

Tuma:2005:LCA

[Tum05-41] Rabiya S. Tuma. Linking centrosomes and actin. Journal of Cell Biology, 170(4):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/4/512.2.

Tuma:2005:LPP

[Tum05-42] Rabiya S. Tuma. Linking pore and pole. Journal of Cell Biology, 170(2):??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/2/162.1.

Tuma:2005:LRG

[Tum05-43] Rabiya S. Tuma. A little Rac goes a long way. Journal of Cell Biology, 170(5):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/5/697.1.

Tuma:2005:MRA

[Tum05-44] Rabiya S. Tuma. Making ROS for apoptosis. Journal of Cell Biology, 170(4):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/4/513.1.

Tuma:2005:MC

[Tum05-45] Rabiya S. Tuma. Mapping the checkpoint. Journal of Cell Biology, 169(1):??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/1/14.1.

Tuma:2005:MD

[Tum05-46] Rabiya S. Tuma. Moving in on desmosomes. Journal of Cell Biology, 171(6):??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/6/913.1.

Tuma:2005:NR

[Tum05-47] Rabiya S. Tuma. NAD to the rescue. Journal of Cell Biology, 170(3):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/3/334.2.

Tuma:2005:NMM

[Tum05-48] Rabiya S. Tuma. A new mode of moving. Journal of Cell Biology, 170(1):??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/1/8.2.

Tuma:2005:NA

[Tum05-49] Rabiya S. Tuma. Nonubiquitous autophagy. Journal of Cell Biology, 171(4):??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/4/581.3.

Tuma:2005:NEP

[Tum05-50] Rabiya S. Tuma. A novel export pathway. Journal of Cell Biology, 171(2):??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/2/190.2.

Tuma:2005:ORL

[Tum05-51] Rabiya S. Tuma. One ROCK is not like another. Journal of Cell Biology, 170(3):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/3/334.1.

Tuma:2005:RRM

[Tum05-52] Rabiya S. Tuma. Raf-1 regulates migration. Journal of Cell Biology, 168(6):??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/6/849.2.

Tuma:2005:RYD

[Tum05-53] Rabiya S. Tuma. Regulated yeast death — in colonies. Journal of Cell Biology, 169(5):??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/5/700.3.

Tuma:2005:RBB

[Tum05-54] Rabiya S. Tuma. Regulating basal body replication. Journal of Cell Biology, 171(6):??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/171/6/913.3.

Tuma:2005:RP

[Tum05-55] Rabiya S. Tuma. Regulating the prostate. Journal of Cell Biology, 170(1):??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/1/9.2.

Tuma:2005:RO

[Tum05-56] Rabiya S. Tuma. Repair orders. Journal of Cell Biology, 170(2): ??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
170/2/163.1.

Tuma:2005:RTa

[Tum05-57] Rabiya S. Tuma. Retrograde tRNAs. Journal of Cell Biology, 170(4):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/4/512.1.

Tuma:2005:RTb

[Tum05-58] Rabiya S. Tuma. A reversal of Tat. Journal of Cell Biology, 171(2):??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/2/190.1.

Tuma:2005:RCF

[Tum05-59] Rabiya S. Tuma. RhoA controls furrow formation. Journal of Cell Biology, 170(1):??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/1/8.1.

Tuma:2005:RC

[Tum05-60] Rabiya S. Tuma. Ribosome cycling. Journal of Cell Biology, 170(3):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/3/335.1.

Tuma:2005:SMEb

[Tum05-61] Rabiya S. Tuma. Seeing muscle evolution. Journal of Cell Biology, 171(4):??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/4/581.1.

Tuma:2005:SSM

[Tum05-62] Rabiya S. Tuma. Severing saves microtubules. Journal of Cell Biology, 168(4):??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/4/521.3.

Tuma:2005:SAD

[Tum05-63] Rabiya S. Tuma. Signal for ER-associated degradation. Journal of Cell Biology, 169(1):??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/1/15.3.

Tuma:2005:SMC

[Tum05-64] Rabiya S. Tuma. Signaling memory from cell shape. Journal of Cell Biology, 170(7):??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/170/7/1015.1.

Tuma:2005:SQC

[Tum05-65] Rabiya S. Tuma. Sperm’s quick calcium response. Journal of Cell Biology, 169(5):??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/5/701.2.

Tuma:2005:SMEa

[Tum05-66] Rabiya S. Tuma. Stabilizing the minus ends? Journal of Cell Biology, 170(7):??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/7/1015.2.

Tuma:2005:SFS

[Tum05-67] Rabiya S. Tuma. A stalled fork stalls recombination. Journal of Cell Biology, 168(4):??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/168/4/520.2.

Tuma:2005:SNE

[Tum05-68] Rabiya S. Tuma. Sticking to the nuclear envelope. Journal of Cell Biology, 168(6):??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/6/848.1.

Tuma:2005:SI

[Tum05-69] Rabiya S. Tuma. Sticky interactions. Journal of Cell Biology, 170(5):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/5/697.3.

Tuma:2005:ST

[Tum05-70] Rabiya S. Tuma. Stretching and twisting. Journal of Cell Biology, 171(2):??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/2/191.2.

Tuma:2005:SSH

[Tum05-71] Rabiya S. Tuma. Sun and survival with Hif1α. Journal of Cell Biology, 170(1):??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/1/9.1.

Tuma:2005:TAP

[Tum05-72] Rabiya S. Tuma. TCRs alive on the periphery. Journal of Cell Biology, 171(6):??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/6/914.1.

Tuma:2005:TSAb

[Tum05-73] Rabiya S. Tuma. There’s something about the midline. Journal of Cell Biology, 171(4):??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/171/4/580.1.

Tuma:2005:TSAa

[Tum05-74] Rabiya S. Tuma. Two-step adhesion. Journal of Cell Biology, 170(3):??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/3/335.3.

Tuma:2005:VCW

[Tum05-75] Rabiya S. Tuma. Viruses catch a wave. Journal of Cell Biology, 170(2):??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/2/163.2.

Tuma:2005:WLM

[Tum05-76] Rabiya S. Tuma. Working up to the last minute. Journal of Cell Biology, 168(4):??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/4/521.2.

Tuma:2005:AMT

[Tum05-77] Rbiya S. Tuma. Actin-myosin tears it up. Journal of Cell Biology, 168(2):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/2/173.1.

Tuma:2005:GT

[Tum05-78] Rbiya S. Tuma. Growth and trafficking. Journal of Cell Biology, 168(2):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/2/172.3.

Tuma:2005:GSN

[Tum05-79] Rbiya S. Tuma. A GTP signal to the nucleolus. Journal of Cell Biology, 168(2):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/2/172.2.

Tuma:2005:MDC

[Tum05-80] Rbiya S. Tuma. Modeling distinct compartments. Journal of Cell Biology, 168(2):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/2/173.2.

Tuma:2005:MGC

[Tum05-81] Rbiya S. Tuma. Myo6 gets choosy. Journal of Cell Biology, 168(2):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/2/172.1.

Tuma:2006:ASS

[Tum06a] Rabiya S. Tuma. Actin senses self-incompatibility. Journal of Cell Biology, 174(2):165–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/174/2/165.

Tuma:2006:AR

[Tum06b] Rabiya S. Tuma. Activating Rac. Journal of Cell Biology, 174(5):??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/5/609a.

Tuma:2006:AML

[Tum06c] Rabiya S. Tuma. Aged muscle lacks sparks. Journal of Cell Biology, 174(5):609–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/5/609.

Tuma:2006:ADE

[Tum06d] Rabiya S. Tuma. Aneuploidy does not equal aging. Journal of Cell Biology, 172(4):??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/172/4/481b.

Tuma:2006:AP

[Tum06e] Rabiya S. Tuma. Assembling pores. Journal of Cell Biology, 173(3):??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/3/315a.

Tuma:2006:AW

[Tum06f] Rabiya S. Tuma. An axonal winch. Journal of Cell Biology, 173(3):??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/3/314a.

Tuma:2006:CFS

[Tum06g] Rabiya S. Tuma. Cell front supports the back. Journal of Cell Biology, 174(3):??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/3/318a.

Tuma:2006:CZ

[Tum06h] Rabiya S. Tuma. Chromatin zigzags. Journal of Cell Biology, 174(1):2–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/1/2.

Tuma:2006:CRF

[Tum06i] Rabiya S. Tuma. Contractile ring formation. Journal of Cell Biology, 174(3):??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/3/319b.

Tuma:2006:CM

[Tum06j] Rabiya S. Tuma. Cooperative motors. Journal of Cell Biology, 174(7):906–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/7/906.

Tuma:2006:DC

[Tum06k] Rabiya S. Tuma. Death by Cdc6. Journal of Cell Biology, 174(1):??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/1/3a.

Tuma:2006:DSP

[Tum06l] Rabiya S. Tuma. DNA stays put. Journal of Cell Biology, 172 (6):784–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/6/784.

Tuma:2006:ET

[Tum06m] Rabiya S. Tuma. Early transcription. Journal of Cell Biology, 174(3):??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/3/319a.

Tuma:2006:ESb

[Tum06n] Rabiya S. Tuma. Endosome signaling. Journal of Cell Biology, 173(1):??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/1/3a.

Tuma:2006:ESa

[Tum06o] Rabiya S. Tuma. Exit without separase. Journal of Cell Biology, 172(6):??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/6/785b.

Tuma:2006:EPM

[Tum06p] Rabiya S. Tuma. Exosomes at the plasma membrane. Journal of Cell Biology, 172(6):785–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/172/6/785.

Tuma:2006:ECO

[Tum06q] Rabiya S. Tuma. An eye on cone outer segments. Journal of Cell Biology, 173(1):??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/1/2a.

Tuma:2006:FL

[Tum06r] Rabiya S. Tuma. Formin localization. Journal of Cell Biology, 174(5):??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/5/608a.

Tuma:2006:FFP

[Tum06s] Rabiya S. Tuma. Funky FAK proliferation. Journal of Cell Biology, 174(2):??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/2/165a.

Tuma:2006:HT

[Tum06t] Rabiya S. Tuma. Homing therapy. Journal of Cell Biology, 174(2):164–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/2/164.

Tuma:2006:S

[Tum06u] Rabiya S. Tuma. In-and-out signaling. Journal of Cell Biology, 173(3):315–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/3/315.

Tuma:2006:ITT

[Tum06v] Rabiya S. Tuma. It takes two to regulate. Journal of Cell Biology, 173(5):??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/5/641b.

Tuma:2006:KMO

[Tum06w] Rabiya S. Tuma. Knocking out malaria in one punch. Journal of Cell Biology, 172(6):??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/6/784a.

Tuma:2006:LD

[Tum06x] Rabiya S. Tuma. Lamins for differentiation. Journal of Cell Biology, 173(1):2–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/1/2.

Tuma:2006:LHM

[Tum06y] Rabiya S. Tuma. Linking Huntingtin to motility. Journal of Cell Biology, 172(4):??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/4/481a.

Tuma:2006:LSC

[Tum06z] Rabiya S. Tuma. Linking secretion and cell cycle. Journal of Cell Biology, 174(2):??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/2/165b.

Tuma:2006:MTR

[Tum06-27] Rabiya S. Tuma. Making γ-tubulin ring. Journal of Cell Biology, 172(4):??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/4/480a.

Tuma:2006:MRR

[Tum06-28] Rabiya S. Tuma. Mannose receptor regulates motility. Journal of Cell Biology, 174(3):319–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/174/3/319.

Tuma:2006:MS

[Tum06-29] Rabiya S. Tuma. Microclustered signaling. Journal of Cell Biology, 174(1):??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/1/3b.

Tuma:2006:MHC

[Tum06-30] Rabiya S. Tuma. miRNAs halt cell cycle. Journal of Cell Biology, 174(5):608–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/5/608.

Tuma:2006:MG

[Tum06-31] Rabiya S. Tuma. Mobile genes. Journal of Cell Biology, 172 (2):??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/2/165b.

Tuma:2006:NCT

[Tum06-32] Rabiya S. Tuma. Neighbors control trafficking. Journal of Cell Biology, 172(6):??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/6/785a.

Tuma:2006:NIM

[Tum06-33] Rabiya S. Tuma. NO induces myoblast fusion. Journal of Cell Biology, 172(2):??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/2/165a.

Tuma:2006:OCT

[Tum06-34] Rabiya S. Tuma. One cell, two chemotaxis systems. Journal of Cell Biology, 173(3):??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/3/315b.

Tuma:2006:OAa

[Tum06-35] Rabiya S. Tuma. Opposites attract. Journal of Cell Biology, 174(1):??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/1/2a.

Tuma:2006:OAb

[Tum06-36] Rabiya S. Tuma. Ordered adhesions. Journal of Cell Biology, 174(5):??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/5/609b.

Tuma:2006:RPM

[Tum06-37] Rabiya S. Tuma. Rab proteins move integrins. Journal of Cell Biology, 173(5):??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/5/641a.

Tuma:2006:RM

[Tum06-38] Rabiya S. Tuma. Reinforced microtubules. Journal of Cell Biology, 173(5):640–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/5/640.

Tuma:2006:RNP

[Tum06-39] Rabiya S. Tuma. Repair at the nuclear periphery. Journal of Cell Biology, 172(2):164–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/172/2/164.

Tuma:2006:SSV

[Tum06-40] Rabiya S. Tuma. Scattering synaptic vesicles. Journal of Cell Biology, 174(2):??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/2/164a.

Tuma:2006:STM

[Tum06-41] Rabiya S. Tuma. Sitting in — not through — a membrane. Journal of Cell Biology, 174(3):318–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/174/3/318.

Tuma:2006:SD

[Tum06-42] Rabiya S. Tuma. A slow death. Journal of Cell Biology, 172 (4):481–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/4/481.

Tuma:2006:SS

[Tum06-43] Rabiya S. Tuma. Soaking up stress. Journal of Cell Biology, 174(1):3–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/1/3.

Tuma:2006:SSD

[Tum06-44] Rabiya S. Tuma. Sometimes size does matter. Journal of Cell Biology, 172(2):165–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/2/165.

Tuma:2006:SJR

[Tum06-45] Rabiya S. Tuma. Spaced just right. Journal of Cell Biology, 174(7):907–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/7/907.

Tuma:2006:SNI

[Tum06-46] Rabiya S. Tuma. Speeding nuclear import. Journal of Cell Biology, 174(7):??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/7/907a.

Tuma:2006:SSM

[Tum06-47] Rabiya S. Tuma. Stat3 stabilizes microtubules. Journal of Cell Biology, 172(2):??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/2/164a.

Tuma:2006:SFT

[Tum06-48] Rabiya S. Tuma. Stress fibers from two sources. Journal of Cell Biology, 173(3):314–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/3/314.

Tuma:2006:TC

[Tum06-49] Rabiya S. Tuma. There’s a catch. Journal of Cell Biology, 174 (7):??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/7/906a.

Tuma:2006:TE

[Tum06-50] Rabiya S. Tuma. Trimming ERAD. Journal of Cell Biology, 174(7):??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/7/907b.

Tuma:2006:TTC

[Tum06-51] Rabiya S. Tuma. Tunneling through cells. Journal of Cell Biology, 173(5):641–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/5/641.

Tuma:2006:UTW

[Tum06-52] Rabiya S. Tuma. Ubiquitin tug–O–war. Journal of Cell Biology, 173(1):3–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/1/3.

Tuma:2006:US

[Tum06-53] Rabiya S. Tuma. Unscrambling secretion. Journal of Cell Biology, 173(1):??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/1/3b.

Tuma:2006:UR

[Tum06-54] Rabiya S. Tuma. Unspliced regulator. Journal of Cell Biology, 172(4):480–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/4/480.

Tuma:2006:ULF

[Tum06-55] Rabiya S. Tuma. An unused license to fire. Journal of Cell Biology, 173(5):??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/5/640a.

Thuault:2006:TGF

[TVP^{+}06] Sylvie Thuault, Ulrich Valcourt, Maj Petersen, Guidalberto Manfioletti, Carl-Henrik Heldin, and Aristidis Moustakas. Transforming growth factor-β employs HMGA2 to elicit epithelial–mesenchymal transition. Journal of Cell Biology, 174(2):175–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/2/175.

Terry:2007:NME

[TW07] Laura J. Terry and Susan R. Wente. Nuclear mRNA export requires specific FG nucleoporins for translocation through the nuclear pore complex. Journal of Cell Biology, 178(7): 1121–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/7/1121.

Tsang:2007:SNC

[TWC^{+}07] William Y. Tsang, Leyu Wang, Zhihong Chen, Irma Sánchez, and Brian David Dynlacht. SCAPER, a novel cyclin A– interacting protein that regulates cell cycle progression. Journal of Cell Biology, 178(4):621–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/178/4/621.

Toso:2009:KGP

[TWG^{+}09] Alberto Toso, Jennifer R. Winter, Ainslie J. Garrod, Ana C. Amaro, Patrick Meraldi, and Andrew D. McAinsh. Kinetochoregenerated pushing forces separate centrosomes during bipolar spindle assembly. Journal of Cell Biology, 184(3):365–??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/184/3/
365.

Tam:2007:CRK

[TWL07] Lai-Wa Tam, Nedra F. Wilson, and Paul A. Lefebvre. A CDKrelated kinase regulates the length and assembly of flagella in Chlamydomonas. Journal of Cell Biology, 176(6):819–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
176/6/819.

Uren:2007:MPR

[UDC^{+}07] Rachel T. Uren, Grant Dewson, Lin Chen, Stephanie C. Coyne, David C. S. Huang, Jerry M. Adams, and Ruth M. Kluck. Mitochondrial permeabilization relies on BH3 ligands engaging multiple prosurvival Bcl-2 relatives, not Bak. Journal of Cell Biology, 177(2):277–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/2/277.

Ueda:2007:IPN

[UFY^{+}07] Hiroshi Ueda, Ryousuke Fujita, Akira Yoshida, Hayato Matsunaga, and Mutsumi Ueda. Identification of prothymosin-α1, the necrosis–apoptosis switch molecule in cortical neuronal cultures. Journal of Cell Biology, 176(6):853–??, March 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/176/6/853.

Uteng:2008:PTE

[UHM^{+}08] Marianne Uteng, Christian Hentrich, Kota Miura, Peter Bieling, and Thomas Surrey. Poleward transport of Eg5 by dynein– dynactin in Xenopus laevis egg extract spindles. Journal of Cell Biology, 182(4):715–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/4/715.

Uetake:2007:CCP

[ULN^{+}07] Yumi Uetake, Jadranka Lončarek, Joshua J. Nordberg, Christopher N. English, Sabrina La Terra, Alexey Khodjakov, and Greenfield Sluder. Cell cycle progression and de novo centriole assembly after centrosomal removal in untransformed human cells. Journal of Cell Biology, 176(2):173–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/176/2/173.

Ulbert:2006:DMP

[UPBM06] Sebastian Ulbert, Melpomeni Platani, Stephanie Boue, and Iain W. Mattaj. Direct membrane protein–DNA interactions required early in nuclear envelope assembly. Journal of Cell Biology, 173(4):469–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/4/469.

Uchida:2009:KSI

[UTK^{+}09] Kazuhiko S. K. Uchida, Kentaro Takagaki, Kazuki Kumada, Youko Hirayama, Tetsuo Noda, and Toru Hirota. Kinetochore stretching inactivates the spindle assembly checkpoint. Journal of Cell Biology, 184(3):383–??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/184/3/383.

Uematsu:2007:ADP

[UWiY^{+}07] Naoya Uematsu, Eric Weterings, Ken ichi Yano, Keiko Morotomi-Yano, Burkhard Jakob, Gisela Taucher-Scholz, Pierre-Olivier Mari, Dik C. van Gent, Benjamin P. C. Chen, and David J. Chen. Autophosphorylation of DNA–PK_{CS} regulates its dynamics at DNA double-strand breaks. Journal of Cell Biology, 177(2):219–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/177/2/219.

Uniacke:2008:SIA

[UZ08] James Uniacke and William Zerges. Stress induces the assembly of RNA granules in the chloroplast of Chlamydomonas reinhardtii. Journal of Cell Biology, 182(4):641–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/182/4/641.

Vaughan:2005:TMT

[Vau05] Kevin T. Vaughan. TIP maker and TIP marker; EB1 as a master controller of microtubule plus ends. Journal of Cell Biology, 171(2):197–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/2/197.

vanBuul:2007:RRE

[vBAS^{+}07] Jaap D. van Buul, Michael J. Allingham, Thomas Samson, Julia Meller, Etienne Boulter, Rafael García-Mata, and Keith Burridge. RhoG regulates endothelial apical cup assembly downstream from ICAM1 engagement and is involved in leukocyte trans-endothelial migration. Journal of Cell Biology, 178(7): 1279–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/7/1279.

vonBlume:2009:ARA

[vBDF^{+}09] Julia von Blume, Juan M. Duran, Elena Forlanelli, Anne-Marie Alleaume, Mikhail Egorov, Roman Polishchuk, Henrik Molina, and Vivek Malhotra. Actin remodeling by ADF/cofilin is required for cargo sorting at the trans-Golgi network. Journal of Cell Biology, 187(7):1055–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/187/7/1055.

VandenAbeele:2006:FIC

[VBG^{+}06] Fabien Vanden Abeele, Gabriel Bidaux, Dmitri Gordienko, Benjamin Beck, Yuri V. Panchin, Ancha V. Baranova, Dmitry V. Ivanov, Roman Skryma, and Natalia Prevarskaya. Functional implications of calcium permeability of the channel formed by pannexin 1. Journal of Cell Biology, 174(4):535–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
174/4/535.

Valiyaveettil:2008:NRM

[VBG^{+}08] Manojkumar Valiyaveettil, Amber A. Bentley, Priya Gursahaney, Rajaa Hussien, Ritu Chakravarti, Nina Kureishy, Soren Prag, and Josephine C. Adams. Novel role of the muskelin– RanBP9 complex as a nucleocytoplasmic mediator of cell morphology regulation. Journal of Cell Biology, 182(4):727–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
182/4/727.

Vitali:2005:AME

[VBL^{+}05] Patrice Vitali, Eugenia Basyuk, Elodie Le Meur, Edouard Bertrand, Françoise Muscatelli, Jérôme Cavaillé, and Alexander Huttenhofer. ADAR2-mediated editing of RNA substrates in the nucleolus is inhibited by C/D small nucleolar RNAs. Journal of Cell Biology, 169(5):745–??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/169/5/745.

Vogel:2006:DSM

[VBNR06] Frank Vogel, Carsten Bornhövd, Walter Neupert, and Andreas S. Reichert. Dynamic subcompartmentalization of the mitochondrial inner membrane. Journal of Cell Biology, 175 (2):237–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/2/237.

VanKomen:2006:ISC

[VBSM06] Jeffrey S. Van Komen, Xiaoyang Bai, Brenton L. Scott, and James A. McNew. An intramolecular t–SNARE complex functions in vivo without the syntaxin NH_{2}-terminal regulatory domain. Journal of Cell Biology, 172(2):295–??, January 2006.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/172/2/295.

Vince:2008:TFS

[VCC^{+}08] James E. Vince, Diep Chau, Bernard Callus, W. Wei-Lynn Wong, Christine J. Hawkins, Pascal Schneider, Mark McKinlay, Christopher A. Benetatos, Stephen M. Condon, Srinivas K. Chunduru, George Yeoh, Robert Brink, David L. Vaux, and John Silke. TWEAK-FN14 signaling induces lysosomal degradation of a cIAP1–TRAF2 complex to sensitize tumor cells to TNFα. Journal of Cell Biology, 182(1):171–??, July 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/182/1/171.

Verollet:2006:DMT

[VCD^{+}06] Christel Vérollet, Nathalie Colombié, Thomas Daubon, Henri-Marc Bourbon, Michel Wright, and Brigitte Raynaud-Messina. Drosophila melanogaster γ-TuRC is dispensable for targeting γ-tubulin to the centrosome and microtubule nucleation. Journal of Cell Biology, 172(4):517–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/172/4/517.

Vincent:2005:SND

[VCN^{+}05] Patrick Vincent, Michael Chua, Fabien Nogue, Ashley Fairbrother, Hal Mekeel, Yue Xu, Nina Allen, Tatiana N. Bibikova, Simon Gilroy, and Vytas A. Bankaitis. A Sec14p-nodulin domain phosphatidylinositol transfer protein polarizes membrane growth of Arabidopsis thaliana root hairs. Journal of Cell Biology, 168(5):801–??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/5/801.

Venkatesh:2007:MDM

[VCSG07] Karthik Venkatesh, Onanong Chivatakarn, Shey-Shing Sheu, and Roman J. Giger. Molecular dissection of the myelinassociated glycoprotein receptor complex reveals cell type– specific mechanisms for neurite outgrowth inhibition. Journal of Cell Biology, 177(3):393–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/177/3/393.

Vecchi:2005:HH

[VD05] Manuela Vecchi and Pier Paolo Di Fiore. It’s HIP to be a hub. Journal of Cell Biology, 170(2):169–??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/170/2/169.

Vale:2009:BSI

[VD09] Ronald D. Vale and Karen Dell. The biological sciences in India. Journal of Cell Biology, 184(3):342–??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/184/3/342.

Vadaie:2008:CSM

[VDA^{+}08] Nadia Vadaie, Heather Dionne, Darowan S. Akajagbor, Seth R. Nickerson, Damian J. Krysan, and Paul J. Cullen. Cleavage of the signaling mucin Msb2 by the aspartyl protease Yps1 is required for MAPK activation in yeast. Journal of Cell Biology, 181(7):1073–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/7/1073.

vandenBoom:2007:UCA

[vdBKB^{+}07] Vincent van den Boom, Susanne M. Kooistra, Marije Boesjes, Bart Geverts, Adriaan B. Houtsmuller, Koshiro Monzen, Issei Komuro, Jeroen Essers, Loes J. Drenth-Diephuis, and Bart J. L. Eggen. UTF1 is a chromatin-associated protein involved in ES cell differentiation. Journal of Cell Biology, 178(6): 913–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/6/913.

Varma:2006:RKC

[VDSV06] Dileep Varma, Denis L. Dujardin, Stephanie A. Stehman, and Richard B. Vallee. Role of the kinetochore/cell cycle checkpoint protein ZW10 in interphase cytoplasmic dynein function. Journal of Cell Biology, 172(5):655–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/172/5/655.

vonDassow:2009:ADD

[vDVM^{+}09] George von Dassow, Koen J. C. Verbrugghe, Ann L. Miller, Jenny R. Sider, and William M. Bement. Action at a distance during cytokinesis. Journal of Cell Biology, 187(6):831–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
187/6/831.

Vercammen:2007:MC

[VDVV07] Dominique Vercammen, Wim Declercq, Peter Vandenabeele, and Frank Van Breusegem. Are metacaspases caspases? Journal of Cell Biology, 179(3):375–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/3/375.

Villadsen:2007:ESC

[VFRJ^{+}07] René Villadsen, Agla J. Fridriksdottir, Lone Rønnov-Jessen, Thorarinn Gudjonsson, Fritz Rank, Mark A. LaBarge, Mina J. Bissell, and Ole W. Petersen. Evidence for a stem cell hierarchy in the adult human breast. Journal of Cell Biology, 177 (1):87–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/1/87.

Vignjevic:2006:RFF

[ViKA^{+}06] Danijela Vignjevic, Shin ichiro Kojima, Yvonne Aratyn, Oana Danciu, Tatyana Svitkina, and Gary G. Borisy. Role of fascin in filopodial protrusion. Journal of Cell Biology, 174(6):863– ??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/6/863.

vanHaastert:2007:ERP

[vKGK07] Peter J. M. van Haastert, Ineke Keizer-Gunnink, and Arjan Kortholt. Essential role of PI3-kinase and phospholipase A2 in Dictyostelium discoideum chemotaxis. Journal of Cell Biology, 177(5):809–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/5/809.

Veltman:2008:FKS

[VKGV08] Douwe M. Veltman, Ineke Keizer-Gunnik, and Peter J. M. Van Haastert. Four key signaling pathways mediating chemotaxis in Dictyostelium discoideum. Journal of Cell Biology, 180(4): 747–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/4/747.

Vijayaraj:2009:KRP

[VKR^{+}09] Preethi Vijayaraj, Cornelia Kröger, Ursula Reuter, Reinhard Windoffer, Rudolf E. Leube, and Thomas M. Magin. Keratins regulate protein biosynthesis through localization of GLUT1 and -3 upstream of AMP kinase and Raptor. Journal of Cell Biology, 187(2):175–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/2/175.

vonKnethen:2007:PAC

[vKST^{+}07] Andreas von Knethen, Mathias Soller, Nico Tzieply, Andreas Weigert, Axel M. Johann, Carla Jennewein, Roman Köhl, and Bernhard Brüne. PPARγ1 attenuates cytosol to membrane translocation of PKCα to desensitize monocytes/macrophages. Journal of Cell Biology, 176(5):681–??, February 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/176/5/681.

Vascotto:2007:ABM

[VLFA^{+}07] Fulvia Vascotto, Danielle Lankar, Gabrielle Faure-André, Pablo Vargas, Jheimmy Diaz, Delphine Le Roux, Maria-Isabel Yuseff, Jean-Baptiste Sibarita, Marianne Boes, Graça Raposo, Evelyne Mougneau, Nicolas Glaichenhaus, Christian Bonnerot, Bénédicte Manoury, and Ana-Maria Lennon-Duménil. The actin-based motor protein myosin II regulates MHC class II trafficking and BCR-driven antigen presentation. Journal of Cell Biology, 176(7):1007–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/176/7/1007.

Vaccari:2008:EER

[VLK^{+}08] Thomas Vaccari, Han Lu, Ritu Kanwar, Mark E. Fortini, and David Bilder. Endosomal entry regulates Notch receptor activation in Drosophila melanogaster. Journal of Cell Biology, 180 (4):755–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/4/755.

Vicente-Manzanares:2008:SAM

[VMKW^{+}08] Miguel Vicente-Manzanares, Margaret A. Koach, Leanna Whitmore, Marcelo L. Lamers, and Alan F. Horwitz. Segregation and activation of myosin IIB creates a rear in migrating cells. Journal of Cell Biology, 183(3):543–??, November 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/183/3/543.

Vader:2006:CPC

[VML06] Gerben Vader, René H. Medema, and Susanne M. A. Lens. The chromosomal passenger complex: guiding Aurora–B through mitosis. Journal of Cell Biology, 173(6):833–??, June 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/173/6/833.

Varma:2008:DRD

[VMSV08] Dileep Varma, Pascale Monzo, Stephanie A. Stehman, and Richard B. Vallee. Direct role of dynein motor in stable kinetochore-microtubule attachment, orientation, and alignment. Journal of Cell Biology, 182(6):1045–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/182/6/
1045.

VanKeymeulen:2009:EPG

[VMY^{+}09] Alexandra Van Keymeulen, Guilhem Mascre, Khalil Kass Youseff, Itamar Harel, Cindy Michaux, Natalie De Geest, Caroline Szpalski, Younes Achouri, Wilhelm Bloch, Bassem A. Hassan, and Cédric Blanpain. Epidermal progenitors give rise to Merkel cells during embryonic development and adult homeostasis. Journal of Cell Biology, 187(1):91–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/187/1/91.

Vicente-Manzanares:2007:RPAa

[VMZW^{+}07a] Miguel Vicente-Manzanares, Jessica Zareno, Leanna Whitmore, Colin K. Choi, and Alan F. Horwitz. Regulation of protrusion, adhesion dynamics, and polarity by myosins IIA and IIB in migrating cells. Journal of Cell Biology, 176(5):573–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/176/5/

Vicente-Manzanares:2007:RPAb

[VMZW^{+}07b] Miguel Vicente-Manzanares, Jessica Zareno, Leanna Whitmore, Colin K. Choi, and Alan F. Horwitz. Regulation of protrusion, adhesion dynamics, and polarity by myosins IIA and IIB in migrating cells. Journal of Cell Biology, 176(7):1073–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/176/7/
1073.

Vachova:2005:PRY

[VP05] Libuše Váchová and Zdena Palková. Physiological regulation of yeast cell death in multicellular colonies is triggered by ammonia. Journal of Cell Biology, 169(5):711–??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/169/5/711.

vanRoyen:2007:CAR

[vRCB^{+}07] Martin E. van Royen, Sónia M. Cunha, Maartje C. Brink, Karin A. Mattern, Alex L. Nigg, Hendrikus J. Dubbink, Pernette J. Verschure, Jan Trapman, and Adriaan B. Houtsmuller. Compartmentalization of androgen receptor protein–protein interactions in living cells. Journal of Cell Biology, 177(1):63–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
177/1/63.

vonRoretz:2008:DMD

[vRG08] Christopher von Roretz and Imed-Eddine Gallouzi. Decoding ARE-mediated decay: is microRNA part of the equation? Journal of Cell Biology, 181(2):189–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/181/2/189.

vanRheenen:2007:EIP

[vRSvR^{+}07] Jacco van Rheenen, Xiaoyan Song, Wies van Roosmalen, Michael Cammer, Xiaoming Chen, Vera DesMarais, Shu-Chin Yip, Jonathan M. Backer, Robert J. Eddy, and John S. Condeelis. EGF-induced PIP_{2} hydrolysis releases and activates cofilin locally in carcinoma cells. Journal of Cell Biology, 179(6): 1247–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/6/1247.

Vladar:2007:MCC

[VS07] Eszter K. Vladar and Tim Stearns. Molecular characterization of centriole assembly in ciliated epithelial cells. Journal of Cell Biology, 178(1):31–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/1/31.

Vale:2009:DMM

[VSG09] Ronald D. Vale, James A. Spudich, and Eric R. Griffis. Dynamics of myosin, microtubules, and Kinesin-6 at the cortex during cytokinesis in Drosophila S2 cells. Journal of Cell Biology, 186(5):727–??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/186/5/727.

Veracini:2008:CBP

[VSR^{+}08] Laurence Veracini, Valérie Simon, Véronique Richard, Burkhart Schraven, Vaclav Horejsi, Serge Roche, and Christine Benistant. The Csk-binding protein PAG regulates PDGF-induced Src mitogenic signaling via GM1. Journal of Cell Biology, 182(3): 603–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/3/603.

Varnai:2006:RIC

[VTRB06] Peter Varnai, Baskaran Thyagarajan, Tibor Rohacs, and Tamas Balla. Rapidly inducible changes in phosphatidylinositol 4,5bisphosphate levels influence multiple regulatory functions of the lipid in intact living cells. Journal of Cell Biology, 175(3): 377–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/3/377.

Vacaru:2009:SSR

[VTT^{+}09a] Ana M. Vacaru, Fikadu G. Tafesse, Philipp Ternes, Vangelis Kondylis, Martin Hermansson, Jos F. H. M. Brouwers, Pentti Somerharju, Catherine Rabouille, and Joost C. M. Holthuis. Sphingomyelin synthase-related protein SMSr controls ceramide homeostasis in the ER. Journal of Cell Biology, 185(6):1013–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
185/6/1013.

Voronina:2009:ICA

[VTT^{+}09b] Vera A. Voronina, Ken-Ichi Takemaru, Piper Treuting, Damon Love, Barbara R. Grubb, Adeline M. Hajjar, Allison Adams, Feng-Qian Li, and Randall T. Moon. Inactivation of Chibby affects function of motile airway cilia. Journal of Cell Biology, 185(2):225–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/2/225.

VanDamme:2008:PFN

[VVL^{+}08] Philip Van Damme, Annelies Van Hoecke, Diether Lambrechts, Peter Vanacker, Elke Bogaert, John van Swieten, Peter Carmeliet, Ludo Van Den Bosch, and Wim Robberecht. Progranulin functions as a neurotrophic factor to regulate neurite outgrowth and enhance neuronal survival. Journal of Cell Biology, 181(1):37–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/1/37.

Vieira:2005:FIT

[VVMS05] Otilia V. Vieira, Paul Verkade, Aki Manninen, and Kai Simons. FAPP2 is involved in the transport of apical cargo in polarized MDCK cells. Journal of Cell Biology, 170(4):521–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/170/4/
521.

Valiathan:2008:PAM

[VW08] Rajeshwari R. Valiathan and Lois S. Weisman. Pushing for answers: is myosin V directly involved in moving mitochondria? Journal of Cell Biology, 181(1):15–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/181/1/15.

VanKeymeulen:2006:SNP

[VWK^{+}06] Alexandra Van Keymeulen, Kit Wong, Zachary A. Knight, Cedric Govaerts, Klaus M. Hahn, Kevan M. Shokat, and Henry R. Bourne. To stabilize neutrophil polarity, PIP3 and Cdc42 augment RhoA activity at the back as well as signals at the front. Journal of Cell Biology, 174(3):437–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/174/3/437.

Vanderluit:2007:RFM

[VWM^{+}07] Jacqueline L. Vanderluit, Crystal A. Wylie, Kelly A. McClellan, Noel Ghanem, Andre Fortin, Steve Callaghan, Jason G. MacLaurin, David S. Park, and Ruth S. Slack. The Retinoblastoma family member p107 regulates the rate of progenitor commitment to a neuronal fate. Journal of Cell Biology, 178(1):129– ??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
178/1/129.

vanZeijl:2007:RCG

[vZPG^{+}07] Leonie van Zeijl, Bas Ponsioen, Ben N. G. Giepmans, Aafke Ariaens, Friso R. Postma, Péter Várnai, Tamas Balla, Nullin Divecha, Kees Jalink, and Wouter H. Moolenaar. Regulation of connexin43 gap junctional communication by phosphatidylinositol 4,5-bisphosphate. Journal of Cell Biology, 177(5):881–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
177/5/881.

Wang:2008:RGB

[WAKS08] Fei Wang, Birgit Agne, Felix Kessler, and Danny J. Schnell. The role of GTP binding and hydrolysis at the atToc159 preprotein receptor during protein import into chloroplasts. Journal of Cell Biology, 183(1):87–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/183/1/87.

Waters:2009:APQ

[Wat09] Jennifer C. Waters. Accuracy and precision in quantitative fluorescence microscopy. Journal of Cell Biology, 185(7):1135–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
185/7/1135.

Winter:2008:PIM

[WAW08] Lilli Winter, Christina Abrahamsberg, and Gerhard Wiche. Plectin isoform 1b mediates mitochondrion–intermediate filament network linkage and controls organelle shape. Journal of Cell Biology, 181(6):903–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/181/6/903.

Wang:2006:ARE

[WBC^{+}06] Xuan Wang, Diana Bolotin, David H. Chu, Lisa Polak, Trevor Williams, and Elaine Fuchs. AP-2α: a regulator of EGF receptor signaling and proliferation in skin epidermis. Journal of Cell Biology, 172(3):409–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/172/3/409.

Weimer:2008:DRA

[WBE^{+}08] Carolin Weimer, Rainer Beck, Priska Eckert, Ingeborg Reckmann, Jörg Moelleken, Britta Brügger, and Felix Wieland. Differential roles of ArfGAP1, ArfGAP2, and ArfGAP3 in COPI trafficking. Journal of Cell Biology, 183(4):725–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/183/4/
725.

Wheeler:2008:CLE

[WBG^{+}08] Damian G. Wheeler, Curtis F. Barrett, Rachel D. Groth, Parsa Safa, and Richard W. Tsien. CaMKII locally encodes Ltype channel activity to signal to nuclear CREB in excitation– transcription coupling. Journal of Cell Biology, 183(5):849– ??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/5/849.

Weisleder:2006:MAA

[WBK^{+}06] Noah Weisleder, Marco Brotto, Shinji Komazaki, Zui Pan, Xiaoli Zhao, Thomas Nosek, Jerome Parness, Hiroshi Takeshima, and Jianjie Ma. Muscle aging is associated with compromised Ca^{2+} spark signaling and segregated intracellular Ca^{2+} release. Journal of Cell Biology, 174(5):639–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/174/5/639.

Weinert:2006:MLD

[WBL^{+}06] Stefanie Weinert, Nora Bergmann, Xiuju Luo, Bettina Erdmann, and Michael Gotthardt. M line–deficient titin causes cardiac lethality through impaired maturation of the sarcomere. Journal of Cell Biology, 173(4):559–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/173/4/559.

Wu:2006:CSD

[WBLL06] Minnie M. Wu, JoAnn Buchanan, Riina M. Luik, and Richard S. Lewis. Ca^{2+} store depletion causes STIM1 to accumulate in ER regions closely associated with the plasma membrane. Journal of Cell Biology, 174(6):803–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/174/6/803.

Watanabe:2008:MCA

[WBP^{+}08] Naohide Watanabe, Laurent Bodin, Manjula Pandey, Matthias Krause, Shaun Coughlin, Vassiliki A. Boussiotis, Mark H. Ginsberg, and Sanford J. Shattil. Mechanisms and consequences of agonist-induced talin recruitment to platelet integrin αIIbβ3. Journal of Cell Biology, 181(7):1211–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/181/7/1211.

Whyte:2008:PRT

[WBT^{+}08] Jacqueline Whyte, Jason R. Bader, Sinji B. F. Tauhata, Maurice Raycroft, Jessica Hornick, K. Kevin Pfister, William S. Lane, Gordon K. Chan, Edward H. Hinchcliffe, Patricia S. Vaughan, and Kevin T. Vaughan. Phosphorylation regulates targeting of cytoplasmic dynein to kinetochores during mitosis. Journal of Cell Biology, 183(5):819–??, December 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/183/5/819.

Wang:2007:STG

[WC07a] Claire Q. F. Wang and C. Yan Cheng. A seamless trespass: germ cell migration across the seminiferous epithelium during spermatogenesis. Journal of Cell Biology, 178(4):549–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
178/4/549.

Weterings:2007:DDP

[WC07b] Eric Weterings and David J. Chen. DNA-dependent protein kinase in nonhomologous end joining: a lock with multiple keys? Journal of Cell Biology, 179(2):183–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/2/183.

Weaver:2009:RAP

[WC09] Beth A. A. Weaver and Don W. Cleveland. The role of aneuploidy in promoting and suppressing tumors. Journal of Cell Biology, 185(6):935–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/6/935.

Wang:2008:ILK

[WCB^{+}08] Hao-Ven Wang, Ling-Wei Chang, Klara Brixius, Sara A. Wickström, Eloi Montanez, Ingo Thievessen, Martin Schwander, Ulrich Müller, Wilhelm Bloch, Ulrike Mayer, and Reinhard Fässler. Integrin-linked kinase stabilizes myotendinous junctions and protects muscle from stress-induced damage. Journal of Cell Biology, 180(5):1037–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/5/1037.

White:2007:IRP

[WCN07] Dominic P. White, Patrick T. Caswell, and Jim C. Norman. α vβ3 and α5β1 integrin recycling pathways dictate downstream Rho kinase signaling to regulate persistent cell migration. Journal of Cell Biology, 177(3):515–??, May 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/177/3/515.

Wei:2005:RIC

[WCR^{+}05] Ying Wei, Ralf-Peter Czekay, Liliane Robillard, Matthias C. Kugler, Feng Zhang, Kevin K. Kim, Jian ping Xiong, Martin J. Humphries, and Harold A. Chapman. Regulation of α5β1 integrin conformation and function by urokinase receptor binding. Journal of Cell Biology, 168(3):501–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/168/3/501.

Wu:2008:CDL

[WCR^{+}08] Jun Wu, Hyekyung P. Cho, David B. Rhee, Dabney K. Johnson, John Dunlap, Yie Liu, and Yisong Wang. Cdc14B depletion leads to centriole amplification, and its overexpression prevents unscheduled centriole duplication. Journal of Cell Biology, 181(3):475–??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/3/475.

Wagenbach:2008:KMC

[WDWC08] Michael Wagenbach, Sarah Domnitz, Linda Wordeman, and Jeremy Cooper. A kinesin-13 mutant catalytically depolymerizes microtubules in ADP. Journal of Cell Biology, 183(4): 617–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/4/617.

Wells:2005:AW

[Wel05a] William A. Wells. Antiviral walls. Journal of Cell Biology, 171(1):??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/1/11.1.

Wells:2005:AU

[Wel05b] William A. Wells. Autophagy unveiled. Journal of Cell Biology, 168(5):??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/5/676.2.

Wells:2005:BKO

[Wel05c] William A. Wells. Bacteria keep oxygen flowing. Journal of Cell Biology, 168(5):??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/5/675.3.

Wells:2005:BGD

[Wel05d] William A. Wells. Bugs get decrepit too. Journal of Cell Biology, 168(5):??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/5/674.1.

Wells:2005:CDR

[Wel05e] William A. Wells. Catalytic dry run. Journal of Cell Biology, 171(5):??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/5/757.3.

Wells:2005:CSL

[Wel05f] William A. Wells. Catching sight of lysosomes. Journal of Cell Biology, 168(2):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/2/174.1.

Wells:2005:CLU

[Wel05g] William A. Wells. A cell line that is under control. Journal of Cell Biology, 168(7):988–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/168/7/988.

Wells:2005:CLS

[Wel05h] William A. Wells. Chemotaxis by local steering. Journal of Cell Biology, 168(5):??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/5/674.2.

Wells:2005:CGA

[Wel05i] William A. Wells. Cilia get arms for bending. Journal of Cell Biology, 168(3):350–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/3/350.

Wells:2005:CER

[Wel05j] William A. Wells. ClpX eats randomly. Journal of Cell Biology, 171(3):??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/3/406.1.

Wells:2005:DGJ

[Wel05k] William A. Wells. Defining gap junctions. Journal of Cell Biology, 169(3):379–??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/3/379.

Wells:2005:DJC

[Wel05l] William A. Wells. Defining junctional complexes. Journal of Cell Biology, 168(7):989–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/168/7/989.

Wells:2005:DE

[Wel05m] William A. Wells. Development is easy. Journal of Cell Biology, 168(3):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/3/349.3.

Wells:2005:DAD

[Wel05n] William A. Wells. Diet affects DNA. Journal of Cell Biology, 168(5):??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/5/675.2.

Wells:2005:DDD

[Wel05o] William A. Wells. Difference drives division. Journal of Cell Biology, 171(5):??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/5/756.2.

Wells:2005:DSV

[Wel05p] William A. Wells. The discovery of synaptic vesicles. Journal of Cell Biology, 168(1):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/168/1/12.1.

Wells:2005:DT

[Wel05q] William A. Wells. The discovery of tubulin. Journal of Cell Biology, 169(4):552–??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/4/552.

Wells:2005:DIN

[Wel05r] William A. Wells. Divided identities in the niche. Journal of Cell Biology, 171(3):??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/171/3/406.2.

Wells:2005:DMM

[Wel05s] William A. Wells. Division as modified migration. Journal of Cell Biology, 169(2):??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/2/215.1.

Wells:2005:ES

[Wel05t] William A. Wells. Exclusion is spreading. Journal of Cell Biology, 168(1):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/1/11.1.

Wells:2005:FR

[Wel05u] William A. Wells. A fat ride. Journal of Cell Biology, 169(4):??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
169/4/550.1.

Wells:2005:FT

[Wel05v] William A. Wells. Fate trafficking. Journal of Cell Biology, 171(1):??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/1/10.2.

Wells:2005:FA

[Wel05w] William A. Wells. Freedom not aggregation. Journal of Cell Biology, 170(6):??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/6/863.1.

Wells:2005:ET

[Wel05x] William A. Wells. From error to tetraploid. Journal of Cell Biology, 171(3):??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/3/407.1.

Wells:2005:GAS

[Wel05y] William A. Wells. Getting ahead by staying still. Journal of Cell Biology, 169(4):??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/4/551.2.

Wells:2005:GNM

[Wel05z] William A. Wells. Getting nuclei moving. Journal of Cell Biology, 170(6):??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/6/863.3.

Wells:2005:HL

[Wel05-27] William A. Wells. Heterochromatin is late. Journal of Cell Biology, 168(4):??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/4/524.1.

Wells:2005:HGS

[Wel05-28] William A. Wells. Highly goes slowly. Journal of Cell Biology, 171(1):??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/1/10.1.

Wells:2005:HSS

[Wel05-29] William A. Wells. How to spot a satellite cell. Journal of Cell Biology, 168(5):??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/5/676.1.

Wells:2005:HVB

[Wel05-30] William A. Wells. How vessels become leaky. Journal of Cell Biology, 168(5):677–??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/5/677.

Wells:2005:IFF

[Wel05-31] William A. Wells. The invention of freeze fracture EM and the determination of membrane structure. Journal of Cell Biology, 168(2):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/2/174.2.

Wells:2005:IDP

[Wel05-32] William A. Wells. Is it DNA or protein? Journal of Cell Biology, 171(3):??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/3/407.3.

Wells:2005:KRD

[Wel05-33] William A. Wells. Keeping RNA away from DNA. Journal of Cell Biology, 170(6):??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/170/6/862.1.

Wells:2005:KAB

[Wel05-34] William A. Wells. Kept alive by a bacterial pull. Journal of Cell Biology, 169(2):??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/2/214.3.

Wells:2005:KHR

[Wel05-35] William A. Wells. Kinetochores hold on with a ring. Journal of Cell Biology, 168(3):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/168/3/349.2.

Wells:2005:LLG

[Wel05-36] William A. Wells. Lengthening without letting go. Journal of Cell Biology, 168(1):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/168/1/10.1.

Wells:2005:LDS

[Wel05-37] William A. Wells. Lipid-driven sorting and fission. Journal of Cell Biology, 169(2):??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/2/214.1.

Wells:2005:MYO

[Wel05-38] William A. Wells. Make your own network. Journal of Cell Biology, 169(4):??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/4/550.2.

Wells:2005:MNF

[Wel05-39] William A. Wells. Mice with no fat carriers are skinny. Journal of Cell Biology, 168(7):??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/7/987.2.

Wells:2005:MV

[Wel05-40] William A. Wells. Microsomes are the in vitro ER. Journal of Cell Biology, 168(1):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/168/1/13.2.

Wells:2005:MGN

[Wel05-41] William A. Wells. Microtubules get a name. Journal of Cell Biology, 168(6):852–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/6/852.

Wells:2005:MI

[Wel05-42] William A. Wells. Mito immunity. Journal of Cell Biology, 171(1):??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/1/11.3.

Wells:2005:NTO

[Wel05-43] William A. Wells. A new take on the old. Journal of Cell Biology, 168(1):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/1/13.1.

Wells:2005:NEC

[Wel05-44] William A. Wells. No exit for Ca^{2+}. Journal of Cell Biology, 168(5):??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/5/675.1.

Wells:2005:NLD

[Wel05-45] William A. Wells. NO link to destruction. Journal of Cell Biology, 171(3):??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/3/407.2.

Wells:2005:NOR

[Wel05-46] William A. Wells. The nucleolar origin of rRNA. Journal of Cell Biology, 168(4):??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/4/524.2.

Wells:2005:OW

[Wel05-47] William A. Wells. The oenocytes went 3 by 3. Journal of Cell Biology, 168(1):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/1/10.2.

Wells:2005:OMC

[Wel05-48] William A. Wells. One more C means no hearing. Journal of Cell Biology, 168(1):??, January 2005. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/168/1/11.3.

Wells:2005:OB

[Wel05-49] William A. Wells. An optical brain. Journal of Cell Biology, 170(6):??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/6/862.2.

Wells:2005:PS

[Wel05-50] William A. Wells. A pathway for secretion. Journal of Cell Biology, 168(4):525–??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/4/525.

Wells:2005:PGC

[Wel05-51] William A. Wells. PML goes to the centrosome. Journal of Cell Biology, 168(7):??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/7/986.3.

Wells:2005:PBS

[Wel05-52] William A. Wells. Pore before seal. Journal of Cell Biology, 168(3):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/3/349.1.

Wells:2005:RSW

[Wel05-53] William A. Wells. Ready, set, wait. Journal of Cell Biology, 168(7):??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/7/987.1.

Wells:2005:ROS

[Wel05-54] William A. Wells. Really outside signaling. Journal of Cell Biology, 169(2):??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/2/215.2.

Wells:2005:RD

[Wel05-55] William A. Wells. Repairing for death. Journal of Cell Biology, 171(1):??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/1/11.2.

Wells:2005:RPP

[Wel05-56] William A. Wells. Ribosomes, or the particles of Palade. Journal of Cell Biology, 168(1):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/168/1/12.2.

Wells:2005:RM

[Wel05-57] William A. Wells. RNA in the machine. Journal of Cell Biology, 169(4):??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/4/551.3.

Wells:2005:SVS

[Wel05-58] William A. Wells. Saving varicella from suicide. Journal of Cell Biology, 168(3):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/3/348.1.

Wells:2005:SM

[Wel05-59] William A. Wells. Split motifs. Journal of Cell Biology, 168(7):??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/7/986.2.

Wells:2005:SDN

[Wel05-60] William A. Wells. Squeezing dendritic necks. Journal of Cell Biology, 171(5):??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/5/757.1.

Wells:2005:SSA

[Wel05-61] William A. Wells. Staying smart with age. Journal of Cell Biology, 171(5):??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/5/757.2.

Wells:2005:SDF

[Wel05-62] William A. Wells. Stretch but don’t follow. Journal of Cell Biology, 168(1):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/1/11.2.

Wells:2005:TBI

[Wel05-63] William A. Wells. TB bug inhibits any which way. Journal of Cell Biology, 168(7):??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/7/986.1.

Wells:2005:TDO

[Wel05-64] William A. Wells. There’s DNA in those organelles. Journal of Cell Biology, 168(6):853–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/168/6/853.

Wells:2005:TTa

[Wel05-65] William A. Wells. Thinner tubes. Journal of Cell Biology, 168(3):??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/168/3/348.2.

Wells:2005:TEE

[Wel05-66] William A. Wells. Time encoding by ERKs. Journal of Cell Biology, 169(2):??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/2/214.2.

Wells:2005:TTb

[Wel05-67] William A. Wells. Turning on by turning. Journal of Cell Biology, 170(6):??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/170/6/863.2.

Wells:2005:WW

[Wel05-68] William A. Wells. Walk this (and that) way. Journal of Cell Biology, 169(4):??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/4/551.1.

Wells:2005:WCS

[Wel05-69] William A. Wells. Wiping the cellular slate clean. Journal of Cell Biology, 171(5):??, December 2005. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/171/5/756.1.

Wells:2006:AB

[Wel06a] William A. Wells. Analogue brains. Journal of Cell Biology, 172(7):958–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/7/958.

Wells:2006:ACT

[Wel06b] William A. Wells. Automatic cell topology. Journal of Cell Biology, 174(5):??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/5/611b.

Wells:2006:BSF

[Wel06c] William A. Wells. Biased spindle formation. Journal of Cell Biology, 173(6):??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/6/825b.

Wells:2006:CS

[Wel06d] William A. Wells. Charged signaling. Journal of Cell Biology, 174(4):??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/4/476a.

Wells:2006:CD

[Wel06e] William A. Wells. Computing direction. Journal of Cell Biology, 175(1):4–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/1/4.

Wells:2006:CP

[Wel06f] William A. Wells. Correction 174, p. 477. Journal of Cell Biology, 174(6):901–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/174/6/901.

Wells:2006:CTS

[Wel06g] William A. Wells. Curing TB with sunlight. Journal of Cell Biology, 172(7):??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/7/958a.

Wells:2006:DU

[Wel06h] William A. Wells. Dynamic uncapping. Journal of Cell Biology, 175(6):841–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/6/841.

Wells:2006:EH

[Wel06i] William A. Wells. Electric healing. Journal of Cell Biology, 174 (4):477–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/4/477.

Wells:2006:EA

[Wel06j] William A. Wells. Essential adaptor. Journal of Cell Biology, 175(6):840–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/6/840.

Wells:2006:FBL

[Wel06k] William A. Wells. Fever brings in the lymphocytes. Journal of Cell Biology, 175(5):??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/175/5/678a.

Wells:2006:FH

[Wel06l] William A. Wells. Fixing the hole. Journal of Cell Biology, 175(1):??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/1/5a.

Wells:2006:FTM

[Wel06m] William A. Wells. Flexible tumor metabolism. Journal of Cell Biology, 174(2):167–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/2/167.

Wells:2006:FMV

[Wel06n] William A. Wells. Flow makes vein. Journal of Cell Biology, 173(4):??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/4/457a.

Wells:2006:FSF

[Wel06o] William A. Wells. Full of sound and fury, but signifying something: XVI International AIDS Conference, Toronto, Canada, August 13–18, 2006. Journal of Cell Biology, 175(1):6–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/175/1/
6.

Wells:2006:GT

[Wel06p] William A. Wells. Gellular transport. Journal of Cell Biology, 175(5):678–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/5/678.

Wells:2006:GDD

[Wel06q] William A. Wells. Golgi directs dendritic traffic. Journal of Cell Biology, 172(1):??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/1/5.3.

Wells:2006:GEP

[Wel06r] William A. Wells. Group effort from poliovirus. Journal of Cell Biology, 172(1):??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/1/5.1.

Wells:2006:HCS

[Wel06s] William A. Wells. How chronic stress exacerbates cancer. Journal of Cell Biology, 174(4):476–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/174/4/476.

Wells:2006:HHL

[Wel06t] William A. Wells. How HIV-1 lost control. Journal of Cell Biology, 174(2):166–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/2/166.

Wells:2006:IWL

[Wel06u] William A. Wells. Immune to weight loss. Journal of Cell Biology, 173(2):149–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/2/149.

Wells:2006:IT

[Wel06v] William A. Wells. Innate tumorigenesis. Journal of Cell Biology, 175(2):197–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/2/197.

Wells:2006:KVC

[Wel06w] William A. Wells. Keeping valves clear. Journal of Cell Biology, 175(1):??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/1/5b.

Wells:2006:KAP

[Wel06x] William A. Wells. Killers aid a pregnancy. Journal of Cell Biology, 174(5):??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/5/611a.

Wells:2006:KO

[Wel06y] William A. Wells. Kinetochore to order. Journal of Cell Biology, 172(7):??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/7/959b.

Wells:2006:LYE

[Wel06z] William A. Wells. Lamins as youth elixir. Journal of Cell Biology, 173(4):??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/4/457b.

Wells:2006:LM

[Wel06-27] William A. Wells. Lipid microscopy. Journal of Cell Biology, 175(2):??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/2/196a.

Wells:2006:LCB

[Wel06-28] William A. Wells. The liver that caveolin built. Journal of Cell Biology, 175(1):??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/1/4a.

Wells:2006:LT

[Wel06-29] William A. Wells. Lymphocyte tracks. Journal of Cell Biology, 175(5):??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/5/679a.

Wells:2006:LL

[Wel06-30] William A. Wells. Lysosomes for lunch. Journal of Cell Biology, 173(4):456–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/4/456.

Wells:2006:MSC

[Wel06-31] William A. Wells. Making a single centrosome. Journal of Cell Biology, 174(4):??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/4/477b.

Wells:2006:MA

[Wel06-32] William A. Wells. Making an autophagosome. Journal of Cell Biology, 175(6):??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/6/841a.

Wells:2006:MU

[Wel06-33] William A. Wells. Malaria undercover. Journal of Cell Biology, 174(5):610–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/5/610.

Wells:2006:MWS

[Wel06-34] William A. Wells. Metabolic waste smoothes consumption. Journal of Cell Biology, 172(5):??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/172/5/641b.

Wells:2006:MGD

[Wel06-35] William A. Wells. Microtubules get dynamic. Journal of Cell Biology, 172(5):643–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/172/5/643.

Wells:2006:MKF

[Wel06-36] William A. Wells. Mixed kinetochore fiber. Journal of Cell Biology, 174(2):??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/2/167b.

Wells:2006:MD

[Wel06-37] William A. Wells. Mixed-up DNA. Journal of Cell Biology, 173(2):??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/2/149a.

Wells:2006:MC

[Wel06-38] William A. Wells. Mopping up chemokines. Journal of Cell Biology, 175(2):??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/2/197a.

Wells:2006:MCL

[Wel06-39] William A. Wells. Moving cells link membrane and actin. Journal of Cell Biology, 175(5):??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/175/5/679b.

Wells:2006:MMS

[Wel06-40] William A. Wells. Muscular measuring stick. Journal of Cell Biology, 173(6):825–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/6/825.

Wells:2006:NAF

[Wel06-41] William A. Wells. Nuclear actin filaments. Journal of Cell Biology, 172(7):??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/7/959a.

Wells:2006:ODM

[Wel06-42] William A. Wells. Outdoor decision-making. Journal of Cell Biology, 173(2):??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/2/148a.

Wells:2006:PC

[Wel06-43] William A. Wells. The physics of chaperones. Journal of Cell Biology, 173(2):148–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/2/148.

Wells:2006:PLS

[Wel06-44] William A. Wells. Pilings in a lipid sea. Journal of Cell Biology, 173(6):824–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/6/824.

Wells:2006:RCN

[Wel06-45] William A. Wells. Rac controls nuclear entry. Journal of Cell Biology, 175(6):??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/6/841b.

Wells:2006:RAD

[Wel06-46] William A. Wells. Relaxing after damage. Journal of Cell Biology, 174(4):??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/4/477a.

Wells:2006:RR

[Wel06-47] William A. Wells. Remodeling replication. Journal of Cell Biology, 172(1):??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/1/5.2.

Wells:2006:RE

[Wel06-48] William A. Wells. Rosettes for elongation. Journal of Cell Biology, 175(2):196–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/2/196.

Wells:2006:STF

[Wel06-49] William A. Wells. Seeing tube formation. Journal of Cell Biology, 174(2):??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/2/167a.

Wells:2006:SE

[Wel06-50] William A. Wells. Segregation by entropy. Journal of Cell Biology, 174(5):611–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/5/611.

Wells:2006:SAP

[Wel06-51] William A. Wells. Sending the advance party for metastases. Journal of Cell Biology, 172(1):??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/172/1/4.2.

Wells:2006:SEM

[Wel06-52] William A. Wells. Slow escape from mitosis. Journal of Cell Biology, 174(2):??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/2/166a.

Wells:2006:SGM

[Wel06-53] William A. Wells. Smell genes must wander. Journal of Cell Biology, 174(5):??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/5/610a.

Wells:2006:SBB

[Wel06-54] William A. Wells. SPB building blocks. Journal of Cell Biology, 173(6):??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/6/824a.

Wells:2006:SS

[Wel06-55] William A. Wells. Spinning septins. Journal of Cell Biology, 175(2):??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/2/197b.

Wells:2006:SRA

[Wel06-56] William A. Wells. Stepwise receptor activation. Journal of Cell Biology, 172(7):959–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/7/959.

Wells:2006:SB

[Wel06-57] William A. Wells. Stressful betrayal. Journal of Cell Biology, 173(2):??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/2/149b.

Wells:2006:SLE

[Wel06-58] William A. Wells. Surviving on low energy. Journal of Cell Biology, 175(6):??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/6/840a.

Wells:2006:SSC

[Wel06-59] William A. Wells. Sympathy for stem cells. Journal of Cell Biology, 172(5):640–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/5/640.

Wells:2006:SVS

[Wel06-60] William A. Wells. Synaptic vesicle structure. Journal of Cell Biology, 175(5):679–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/5/679.

Wells:2006:TBS

[Wel06-61] William A. Wells. Tether before synapse. Journal of Cell Biology, 175(1):5–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/1/5.

Wells:2006:TPB

[Wel06-62] William A. Wells. Tortoise puts brake on hare. Journal of Cell Biology, 172(5):??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/5/640a.

Wells:2006:TC

[Wel06-63] William A. Wells. Tracks for cellulose. Journal of Cell Biology, 173(4):457–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/4/457.

Wells:2006:TG

[Wel06-64] William A. Wells. Transcriptional glue. Journal of Cell Biology, 173(6):??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/6/825a.

Wells:2006:TP

[Wel06-65] William A. Wells. Trip to the pore. Journal of Cell Biology, 172(5):??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/5/641a.

Wells:2006:TE

[Wel06-66] William A. Wells. Twisting endocytosis. Journal of Cell Biology, 173(4):??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/4/456a.

Wells:2006:USH

[Wel06-67] William A. Wells. Unpleasant surprises: how the Introduction has wandered into the Discussion. Journal of Cell Biology, 174(6):741–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/6/741.

Wells:2006:VAJ

[Wel06-68] William A. Wells. Virtual adherens junctions. Journal of Cell Biology, 172(1):??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/1/4.1.

Wells:2006:VSA

[Wel06-69] William A. Wells. VWF springs into action. Journal of Cell Biology, 172(5):641–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/5/641.

Wells:2007:BMB

[Wel07a] William A. Wells. Big mammals have big (or slow) cells. Journal of Cell Biology, 176(7):??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/7/893b.

Wells:2007:CB

[Wel07b] William A. Wells. Chemokine blockade. Journal of Cell Biology, 176(5):551–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/5/551.

Wells:2007:CDA

[Wel07c] William A. Wells. Clonal diversity by asymmetry. Journal of Cell Biology, 176(7):892–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/176/7/892.

Wells:2007:DM

[Wel07d] William A. Wells. Destroying the messenger. Journal of Cell Biology, 176(7):893–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/7/893.

Wells:2007:EF

[Wel07e] William A. Wells. Evolving folding. Journal of Cell Biology, 176(5):??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/5/550a.

Wells:2007:GMI

[Wel07f] William A. Wells. Gero Miesenböck: instructing the nervous system. Journal of Cell Biology, 177(3):374–??, May 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/177/3/374.

Wells:2007:HML

[Wel07g] William A. Wells. How to make less webby webs. Journal of Cell Biology, 176(5):550–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/176/5/550.

Wells:2007:IME

[Wel07h] William A. Wells. Ira Mellman: from endosomes to industry. Journal of Cell Biology, 177(4):570–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/177/4/570.

Wells:2007:MRB

[Wel07i] William A. Wells. Michael Rape: the biology of destruction. Journal of Cell Biology, 177(2):188–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/177/2/188.

Wells:2007:MTA

[Wel07j] William A. Wells. Migrating toward adhesion. Journal of Cell Biology, 176(7):??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/7/893a.

Wells:2007:MSC

[Wel07k] William A. Wells. Motoring to a signaling check-up. Journal of Cell Biology, 176(5):??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/176/5/551b.

Wells:2007:PC

[Wel07l] William A. Wells. Pass the cycloheximide. Journal of Cell Biology, 176(5):??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/5/551a.

Wells:2007:RT

[Wel07m] William A. Wells. The returning tide. Journal of Cell Biology, 176(4):376–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/4/376.

White:2007:RCT

[WEL^{+}07n] Susan Roehl White, Katia J. Evans, Jeffrey Lary, James L. Cole, and Brett Lauring. Recognition of C-terminal amino acids in tubulin by pore loops in Spastin is important for microtubule severing. Journal of Cell Biology, 176(7):995–??, March 2007.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/176/7/995.

Wong:2005:PKR

[WF05] Oi Kwan Wong and Guowei Fang. Plx1 is the 3F3/2 kinase responsible for targeting spindle checkpoint proteins to kinetochores. Journal of Cell Biology, 170(5):709–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/170/5/709.

Wong:2006:HCS

[WF06] Jim Wong and Guowei Fang. HURP controls spindle dynamics to promote proper interkinetochore tension and efficient kinetochore capture. Journal of Cell Biology, 173(6):879–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/173/6/
879.

Wong:2007:CPB

[WF07] Oi Kwan Wong and Guowei Fang. Cdk1 phosphorylation of BubR1 controls spindle checkpoint arrest and Plk1-mediated formation of the 3F3/2 epitope. Journal of Cell Biology, 179 (4):611–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/4/611.

Wood:2006:DMR

[WFJ06] Will Wood, Celia Faria, and Antonio Jacinto. Distinct mechanisms regulate hemocyte chemotaxis during development and wound healing in Drosophila melanogaster. Journal of Cell Biology, 173(3):405–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/3/405.

White:2008:AOF

[WGKA^{+}08] Robin White, Constantin Gonsior, Eva-Maria Krämer-Albers, Nadine Stöhr, Stefan Hüttelmaier, and Jacqueline Trotter. Activation of oligodendroglial Fyn kinase enhances translation of mRNAs transported in hnRNP A2-dependent RNA granules. Journal of Cell Biology, 181(4):579–??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/181/4/579.

Woodward:2006:EML

[WGL^{+}06] Anna M. Woodward, Thomas Göhler, M. Gloria Luciani, Maren Oehlmann, Xinquan Ge, Anton Gartner, Dean A. Jackson, and J. Julian Blow. Excess Mcm2–7 license dormant origins of replication that can be used under conditions of replicative stress. Journal of Cell Biology, 173(5):673–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/173/5/673.

Wang:2006:SDM

[WH06] Hong Wang and Christer Höög. Structural damage to meiotic chromosomes impairs DNA recombination and checkpoint control in mammalian oocytes. Journal of Cell Biology, 173 (4):485–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/173/4/485.

Wen:2007:BGS

[WHB^{+}07] Zhexing Wen, Liang Han, James R. Bamburg, Sangwoo Shim, Guo li Ming, and James Q. Zheng. BMP gradients steer nerve growth cones by a balancing act of LIM kinase and Slingshot phosphatase on ADF/cofilin. Journal of Cell Biology, 178 (1):107–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/1/107.

Wang:2007:UST

[WHC^{+}07] Xiaoli Wang, Roger A. Herr, Wei-Jen Chua, Lonnie Lybarger, Emmanuel J. H. J. Wiertz, and Ted H. Hansen. Ubiquitination of serine, threonine, or lysine residues on the cytoplasmic tail can induce ERAD of MHC–I by viral E3 ligase mK3. Journal of Cell Biology, 177(4):613–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/177/4/613.

Wixler:2007:DLO

[WHM^{+}07] Viktor Wixler, Stephanie Hirner, Judith M. Müller, Lucia Gullotti, Carola Will, Jutta Kirfel, Thomas Günther, Holm Schneider, Anja Bosserhoff, Hubert Schorle, Jung Park, Roland Schüle, and Reinhard Buettner. Deficiency in the LIM-only protein Fhl2 impairs skin wound healing. Journal of Cell Biology, 177(1):163–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/1/163.

Wang:2006:ECC

[WHOS06] Chao-Wen Wang, Susan Hamamoto, Lelio Orci, and Randy Schekman. Exomer: a coat complex for transport of select membrane proteins from the trans-Golgi network to the plasma membrane in yeast. Journal of Cell Biology, 174(7): 973–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/7/973.

Wang:2009:UUH

[WHR^{+}09] Xiaoli Wang, Roger A. Herr, Martijn Rabelink, Rob C. Hoeben, Emmanuel J. H. J. Wiertz, and Ted H. Hansen. Ube2j2 ubiquitinates hydroxylated amino acids on ER-associated degradation substrates. Journal of Cell Biology, 187(5):655–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/187/5/
655.

Wang:2005:DBR

[WIL^{+}05] Huidong Wang, Anna Iacoangeli, Daisy Lin, Keith Williams, Robert B. Denman, Christopher U. T. Hellen, and Henri Tiedge. Dendritic BC1 RNA in translational control mechanisms. Journal of Cell Biology, 171(5):811–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/171/5/811.

Wang:2006:C

[WIL^{+}06a] Huidong Wang, Anna Iacoangeli, Daisy Lin, Keith Williams, Robert B. Denman, Christopher U. T. Hellen, and Henri Tiedge. Correction. Journal of Cell Biology, 172(4):635– ??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/4/635.

Wilde:2006:HWK

[Wil06b] Andrew Wilde. “HURP on” we’re off to the kinetochore! Journal of Cell Biology, 173(6):829–??, June 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/173/6/829.

Williams:2006:CDD

[Wil06c] Ruth Williams. Centrosomes deliver the death blow. Journal of Cell Biology, 175(3):366–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/175/3/366.

Williams:2006:CP

[Wil06d] Ruth Williams. Correction 175, p. 367a. Journal of Cell Biology, 175(4):673–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/4/673.

Williams:2006:CC

[Wil06e] Ruth Williams. Cosmc converter. Journal of Cell Biology, 175 (3):??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/3/366a.

Williams:2006:FF

[Wil06f] Ruth Williams. The fragility factor. Journal of Cell Biology, 174(3):??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/3/321b.

Williams:2006:HTH

[Wil06g] Ruth Williams. Handy tool for high-resolution microscopy. Journal of Cell Biology, 174(6):738–??, September 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/174/6/738.

Williams:2006:LLC

[Wil06h] Ruth Williams. Live long and cancer-free. Journal of Cell Biology, 174(6):??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/6/739b.

Williams:2006:MM

[Wil06i] Ruth Williams. Microtubule munching. Journal of Cell Biology, 174(6):??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/6/739a.

Williams:2006:NSC

[Wil06j] Ruth Williams. Nucleotides: saviors of the cell. Journal of Cell Biology, 174(3):321–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/3/321.

Williams:2006:OCC

[Wil06k] Ruth Williams. An open or closed case for HMGA. Journal of Cell Biology, 174(6):739–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/174/6/739.

Williams:2006:PNR

[Wil06l] Ruth Williams. Point of no return for senescence. Journal of Cell Biology, 175(3):??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/175/3/367b.

Williams:2006:PP

[Wil06m] Ruth Williams. Positioning proteasomes. Journal of Cell Biology, 174(3):??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/3/321a.

Williams:2006:SSP

[Wil06n] Ruth Williams. Salmonella sneaks past security. Journal of Cell Biology, 174(3):320–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/3/320.

Williams:2006:TAT

[Wil06o] Ruth Williams. Transforming antibiotic treatment. Journal of Cell Biology, 174(3):??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/3/320a.

Williams:2006:TTN

[Wil06p] Ruth Williams. tRNA transport to the nucleus. Journal of Cell Biology, 175(3):367–??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/175/3/367.

Williams:2006:UDD

[Wil06q] Ruth Williams. Ubp6 delays destruction. Journal of Cell Biology, 175(3):??, November 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/3/367a.

Williams:2007:AGO

[Wil07a] Ruth Williams. Actin, the great orchestrator. Journal of Cell Biology, 178(2):182–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/2/182.

Williams:2007:ASC

[Wil07b] Ruth Williams. Actin’s stable cable. Journal of Cell Biology, 177(5):746–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/5/746.

Williams:2007:AHU

[Wil07c] Ruth Williams. Anthony Hyman: From unlikely scientist to Royal Society Fellow. Journal of Cell Biology, 179(7):1330– ??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/7/1330.

Williams:2007:BMF

[Wil07d] Ruth Williams. Bending membrane to form filopodia. Journal of Cell Biology, 176(7):??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/7/890a.

Williams:2007:CCSb

[Wil07e] Ruth Williams. Calcium channels SMA. Journal of Cell Biology, 179(1):??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/1/3.2.

Williams:2007:CCD

[Wil07f] Ruth Williams. Cellular career decisions start early. Journal of Cell Biology, 178(4):??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/4/542.2.

Williams:2007:CSP

[Wil07g] Ruth Williams. Cellular suicide pact. Journal of Cell Biology, 178(4):??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/4/543.1.

Williams:2007:CS

[Wil07h] Ruth Williams. Cilia signaling. Journal of Cell Biology, 177 (5):747–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/5/747.

Williams:2007:CGC

[Wil07i] Ruth Williams. Clean up or go crazy. Journal of Cell Biology, 179(3):??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/3/360.1.

Williams:2007:CCSa

[Wil07j] Ruth Williams. Clustering creates a spindle. Journal of Cell Biology, 176(3):??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/3/247b.

Williams:2007:CPB

[Wil07k] Ruth Williams. Constructing P-bodies. Journal of Cell Biology, 179(3):??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/3/361.2.

Williams:2007:CRN

[Wil07l] Ruth Williams. Could rejuvenating neurons protect against Parkinson’s? Journal of Cell Biology, 178(1):??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/178/1/4a.

Williams:2007:CCB

[Wil07m] Ruth Williams. The critical chaperone balance. Journal of Cell Biology, 177(2):??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/2/185b.

Williams:2007:CA

[Wil07n] Ruth Williams. Curves are attractive. Journal of Cell Biology, 177(6):947–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/6/947.

Williams:2007:DDT

[Wil07o] Ruth Williams. Death-defying tactics. Journal of Cell Biology, 178(1):5–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/1/5.

Williams:2007:DSC

[Wil07p] Ruth Williams. Dendritic spines cut loose to mature. Journal of Cell Biology, 178(4):??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/4/543.3.

Williams:2007:DD

[Wil07q] Ruth Williams. DIX dynamics. Journal of Cell Biology, 178(1): ??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
178/1/5a.

Williams:2007:DDD

[Wil07r] Ruth Williams. Dynactin drives dynein without microtubule binding. Journal of Cell Biology, 176(5):549–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/176/5/549.

Williams:2007:ELR

[Wil07s] Ruth Williams. Early and late recombination roadblocks. Journal of Cell Biology, 176(5):??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/176/5/548a.

Williams:2007:EC

[Wil07t] Ruth Williams. Eat to compete. Journal of Cell Biology, 178(2): ??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
178/2/183a.

Williams:2007:EE

[Wil07u] Ruth Williams. The exosome exchange. Journal of Cell Biology, 177(5):??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/5/747a.

Williams:2007:FYS

[Wil07v] Ruth Williams. Fibulin-5 for youthful skin. Journal of Cell Biology, 176(7):891–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/7/891.

Williams:2007:GWR

[Wil07w] Ruth Williams. Gilbert Weidinger: Regeneration researcher. Journal of Cell Biology, 179(6):1088–??, December 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/179/6/1088.

Williams:2007:GP

[Wil07x] Ruth Williams. Glycosylation PERKs. Journal of Cell Biology, 176(5):??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/5/549b.

Williams:2007:GLN

[Wil07y] Ruth Williams. Golgi lands new role. Journal of Cell Biology, 178(1):??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/1/5b.

Williams:2007:HHR

[Wil07z] Ruth Williams. Heart hormone release. Journal of Cell Biology, 179(3):??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/3/361.3.

Williams:2007:HISa

[Wil07-27] Ruth Williams. Hey, I’m still replicating! Journal of Cell Biology, 177(1):4–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/1/4.

Williams:2007:HEP

[Wil07-28] Ruth Williams. How ERK5 prompts proliferation. Journal of Cell Biology, 177(2):??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/2/185a.

Williams:2007:HISb

[Wil07-29] Ruth Williams. How Ire1 senses stress. Journal of Cell Biology, 179(1):??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/1/3.3.

Williams:2007:IIT

[Wil07-30] Ruth Williams. Imaging individual transcription factors. Journal of Cell Biology, 177(6):??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/177/6/946a.

Williams:2007:JLP

[Wil07-31] Ruth Williams. Jidong Liu: Probing P-bodies for the secrets of silencing. Journal of Cell Biology, 178(6):894–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/178/6/
894.

Williams:2007:KHN

[Wil07-32] Ruth Williams. Konrad Hochedlinger: The new kid of nuclear reprogramming. Journal of Cell Biology, 178(7):1098– ??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/7/1098.

Williams:2007:LCL

[Wil07-33] Ruth Williams. Langerhans cells limit HIV invasion. Journal of Cell Biology, 177(1):??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/1/5a.

Williams:2007:MHH

[Wil07-34] Ruth Williams. The macrophage home of HIV. Journal of Cell Biology, 177(2):184–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/2/184.

Williams:2007:MBH

[Wil07-35] Ruth Williams. Macrophages boost hair growth. Journal of Cell Biology, 176(7):??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/7/891a.

Williams:2007:MM

[Wil07-36] Ruth Williams. Methylate to memorize. Journal of Cell Biology, 177(1):??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/1/4a.

Williams:2007:MDG

[Wil07-37] Ruth Williams. Microbots deliver the goods. Journal of Cell Biology, 178(2):183–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/2/183.

Williams:2007:MNK

[Wil07-38] Ruth Williams. Migrating neurons keep their limbs tucked in. Journal of Cell Biology, 178(2):??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/178/2/183b.

Williams:2007:MPO

[Wil07-39] Ruth Williams. Mother puts organelles in their place. Journal of Cell Biology, 177(6):946–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/177/6/946.

Williams:2007:MSS

[Wil07-40] Ruth Williams. mRNAs sit out the stress in EGP bodies. Journal of Cell Biology, 179(1):??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/179/1/2.2.

Williams:2007:MB

[Wil07-41] Ruth Williams. Multifunctional BAF-1. Journal of Cell Biology, 178(4):??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/4/542.1.

Williams:2007:MMM

[Wil07-42] Ruth Williams. Myosin mobilizes MHC. Journal of Cell Biology, 176(7):890–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/7/890.

Williams:2007:NNF

[Wil07-43] Ruth Williams. New neurons’ fleeting flexibility. Journal of Cell Biology, 177(6):??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/6/947b.

Williams:2007:N

[Wil07-44] Ruth Williams. Nucleator or not? Journal of Cell Biology, 179(1):??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/1/3.1.

Williams:2007:ORT

[Wil07-45] Ruth Williams. Oliver Rando: Taking chromatin analysis to the genomic scale. Journal of Cell Biology, 177(6):948–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/177/6/
948.

Williams:2007:PPP

[Wil07-46] Ruth Williams. Peptides prompt prions. Journal of Cell Biology, 177(5):??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/5/747b.

Williams:2007:PE

[Wil07-47] Ruth Williams. PIP_{2} in endocytosis. Journal of Cell Biology, 177(2):185–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/2/185.

Williams:2007:PPT

[Wil07-48] Ruth Williams. Podosomes push through. Journal of Cell Biology, 178(2):??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/2/182a.

Williams:2007:PPL

[Wil07-49] Ruth Williams. A polarity/proton loop. Journal of Cell Biology, 179(3):??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/3/361.1.

Williams:2007:PPS

[Wil07-50] Ruth Williams. Pom1p prevents a spreading middle. Journal of Cell Biology, 176(3):??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/176/3/247a.

Williams:2007:PXI

[Wil07-51] Ruth Williams. Positioning X for inactivity. Journal of Cell Biology, 177(5):??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/5/746a.

Williams:2007:PCP

[Wil07-52] Ruth Williams. Proteasomes on ES cell patrol. Journal of Cell Biology, 176(3):246–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/3/246.

Williams:2007:PP

[Wil07-53] Ruth Williams. Putting ER in its place. Journal of Cell Biology, 179(3):??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/3/360.2.

Williams:2007:RNR

[Wil07-54] Ruth Williams. Recruitment but no repair. Journal of Cell Biology, 177(2):??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/2/184a.

Williams:2007:RTB

[Wil07-55] Ruth Williams. Roger Tsien: Bringing color to cell biology. Journal of Cell Biology, 179(1):6–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/179/1/6.

Williams:2007:SSS

[Wil07-56] Ruth Williams. SCAPER speeds and slows the cell cycle. Journal of Cell Biology, 178(4):??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/178/4/543.2.

Williams:2007:SFL

[Wil07-57] Ruth Williams. Scott Fraser: live action embryology. Journal of Cell Biology, 177(5):748–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/177/5/748.

Williams:2007:SEE

[Wil07-58] Ruth Williams. Self-eating embryos. Journal of Cell Biology, 177(1):5–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/1/5.

Williams:2007:SPT

[Wil07-59] Ruth Williams. Spastin pulls tubulin out of shape. Journal of Cell Biology, 176(7):??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/7/891b.

Williams:2007:SMS

[Wil07-60] Ruth Williams. Stabilizing the microtubule scaffold. Journal of Cell Biology, 176(3):??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/176/3/246a.

Williams:2007:SSA

[Wil07-61] Ruth Williams. Stalling secretase assembly. Journal of Cell Biology, 176(5):??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/5/549a.

Williams:2007:SS

[Wil07-62] Ruth Williams. Steps to STIMulation. Journal of Cell Biology, 177(6):??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/6/947a.

Williams:2007:TP

[Wil07-63] Ruth Williams. Tanning with p53. Journal of Cell Biology, 177(1):??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/1/5b.

Williams:2007:TTA

[Wil07-64] Ruth Williams. Tension testing in adherent cells. Journal of Cell Biology, 179(1):??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/1/2.1.

Williams:2007:TSM

[Wil07-65] Ruth Williams. Transient silencing made permanent. Journal of Cell Biology, 176(3):247–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/176/3/247.

Williams:2007:TCG

[Wil07-66] Ruth Williams. Tumor cells go with the flow. Journal of Cell Biology, 178(1):4–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/1/4.

Williams:2007:VPF

[Wil07-67] Ruth Williams. Vascular permeability: the flow factor. Journal of Cell Biology, 176(5):548–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/176/5/548.

Williams:2007:WZC

[Wil07-68] Ruth Williams. W. Zacheus Cande: Evolutionary biologist in cell biologist’s clothing. Journal of Cell Biology, 178(5):718–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
178/5/718.

Williams:2008:AFP

[Wil08a] Ruth Williams. The actin flow paradox. Journal of Cell Biology, 183(6):??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/6/965.1.

Williams:2008:ALG

[Wil08b] Ruth Williams. Andrea Ladd: Getting to the heart of alternative splicing control. Journal of Cell Biology, 183(6):966–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
183/6/966.

Williams:2008:AAA

[Wil08c] Ruth Williams. APC ain’t always necessary for axons. Journal of Cell Biology, 183(2):??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/183/2/172.3.

Williams:2008:AHV

[Wil08d] Ruth Williams. Ari Helenius: viruses under surveillance. Journal of Cell Biology, 182(3):414–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/182/3/414.

Williams:2008:ACJ

[Wil08e] Ruth Williams. Astrocyte’s clean up job. Journal of Cell Biology, 183(2):??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/2/173.1.

Williams:2008:BLD

[Wil08f] Ruth Williams. Brett Lindenbach: Deconstructing and reconstructing hepatitis C. Journal of Cell Biology, 183(5): 754–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/5/754.

Williams:2008:BCM

[Wil08g] Ruth Williams. Brian Chait: Master of mass spectrometry. Journal of Cell Biology, 180(4):652–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/180/4/652.

Williams:2008:BAE

[Wil08h] Ruth Williams. Bruce Alberts: Education, education, education. Journal of Cell Biology, 182(5):820–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/182/5/820.

Williams:2008:CCC

[Wil08i] Ruth Williams. CK2: channel controller. Journal of Cell Biology, 183(6):??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/6/964.3.

Williams:2008:EBB

[Wil08j] Ruth Williams. Elizabeth Blackburn: Because science is worth it. Journal of Cell Biology, 180(6):1056–??, March 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/180/6/1056.

Williams:2008:FRA

[Wil08k] Ruth Williams. Flavivirus reveals its access code. Journal of Cell Biology, 183(2):??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/2/172.1.

Williams:2008:IB

[Wil08l] Ruth Williams. IP blacklist. Journal of Cell Biology, 183(2):??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
183/2/172.2.

Williams:2008:JZD

[Wil08m] Ruth Williams. Jennifer Zallen: Decoding the developmental dance. Journal of Cell Biology, 180(5):850–??, March 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/180/5/850.

Williams:2008:JUR

[Wil08n] Ruth Williams. Jernej Ule: an RNA runaway success. Journal of Cell Biology, 180(3):448–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/180/3/448.

Williams:2008:JPG

[Wil08o] Ruth Williams. Junction protein goes on the road. Journal of Cell Biology, 183(6):??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/183/6/964.2.

Williams:2008:KDF

[Wil08p] Ruth Williams. Kris DeMali: Fixed on cell adhesion. Journal of Cell Biology, 182(6):1032–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/182/6/1032.

Williams:2008:MSB

[Wil08q] Ruth Williams. Microdomain switching is a bad move for APP. Journal of Cell Biology, 183(2):??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/183/2/173.2.

Williams:2008:MAA

[Wil08r] Ruth Williams. Multiple axons and actions with PSD-95. Journal of Cell Biology, 183(6):??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/183/6/964.1.

Williams:2008:PMA

[Wil08s] Ruth Williams. Paul Mischel: All about brains. Journal of Cell Biology, 181(7):1044–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/7/1044.

Williams:2008:RGN

[Wil08t] Ruth Williams. Richard Gardner: Nuclear garbologist. Journal of Cell Biology, 181(1):6–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/181/1/6.

Williams:2008:SJG

[Wil08u] Ruth Williams. Sir John Gurdon: Godfather of cloning. Journal of Cell Biology, 181(2):178–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/181/2/178.

Williams:2008:THKb

[Wil08v] Ruth Williams. Tae Hoon Kim: Knows his boundaries. Journal of Cell Biology, 183(1):4–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/183/1/4.

Williams:2008:THKa

[Wil08w] Ruth Williams. Tony Hunter: Kinase king. Journal of Cell Biology, 181(4):572–??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/4/572.

Williams:2008:YSS

[Wil08x] Ruth Williams. Yasushi Saka: Stirring a melting pot of math and morphogens. Journal of Cell Biology, 181(5):716–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/181/5/
716.

Williams:2009:LI

[Wil09a] Ruth Williams. 1D is like 3D, but 2D isn’t. Journal of Cell Biology, 184(4):??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/4/464.2.

Williams:2009:AGT

[Wil09b] Ruth Williams. Antonio Giraldez: At the tip of the microRNA iceberg. Journal of Cell Biology, 185(7):1132–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/185/7/
1132.

Williams:2009:BST

[Wil09c] Ruth Williams. Blood, stress, and tiRNAs. Journal of Cell Biology, 185(1):3–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/1/3.

Williams:2009:BB

[Wil09d] Ruth Williams. Brain building. Journal of Cell Biology, 185(1):??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/1/2.1.

Williams:2009:CS

[Wil09e] Ruth Williams. Centromere stabilizer. Journal of Cell Biology, 185(3):??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/3/372.1.

Williams:2009:DD

[Wil09f] Ruth Williams. Drainage developers. Journal of Cell Biology, 185(3):373–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/3/373.

Williams:2009:DSP

[Wil09g] Ruth Williams. Dynamics of staying put. Journal of Cell Biology, 185(5):??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/5/756.3.

Williams:2009:FFT

[Wil09h] Ruth Williams. Following the fat trail to death. Journal of Cell Biology, 185(5):757–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/5/757.

Williams:2009:HSM

[Wil09i] Ruth Williams. Heart saves muscle. Journal of Cell Biology, 185(5):??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/5/756.2.

Williams:2009:HCS

[Wil09j] Ruth Williams. How T cells shape their signals. Journal of Cell Biology, 185(3):??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/3/372.3.

Williams:2009:HWN

[Wil09k] Ruth Williams. How Wld^{S} neuroprotects. Journal of Cell Biology, 184(4):??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/4/464.1.

Williams:2009:HDR

[Wil09l] Ruth Williams. Hyung Don Ryoo: a healthy career in cellular death. Journal of Cell Biology, 185(5):758–??, June 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/185/5/758.

Williams:2009:MFF

[Wil09m] Ruth Williams. Marc Freeman: Fishing for the function of fly glia. Journal of Cell Biology, 186(3):302–??, August 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/186/3/302.

Williams:2009:PP

[Wil09n] Ruth Williams. Portioning up PIE-1. Journal of Cell Biology, 184(4):??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/184/4/464.3.

Williams:2009:POC

[Wil09o] Ruth Williams. Pygo2 opens chromatin and cycles cells. Journal of Cell Biology, 185(5):??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/185/5/756.1.

Williams:2009:SVS

[Wil09p] Ruth Williams. Speedy versus sluggish cells. Journal of Cell Biology, 185(1):??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/1/2.2.

Williams:2009:TNB

[Wil09q] Ruth Williams. Tankyrase: a new breed of telomere security guard. Journal of Cell Biology, 184(4):??, February 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/184/4/465.1.

Williams:2009:UTS

[Wil09r] Ruth Williams. Ubiquitin traffic signals. Journal of Cell Biology, 185(3):??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/3/372.2.

Williams:2009:WDK

[Wil09s] Ruth Williams. What do kidneys and embryonic fish skin have in common? Journal of Cell Biology, 185(1):??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/185/1/2.3.

Wittmann:2008:ECC

[Wit08] Torsten Wittmann. EBs clip CLIPs to growing microtubule ends. Journal of Cell Biology, 183(7):1183–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/183/7/
1183.

Wang:2005:DIC

[WJM05] Tao Wang, Yuchen Jiao, and Craig Montell. Dissecting independent channel and scaffolding roles of the Drosophila transient receptor potential channel. Journal of Cell Biology, 171(4): 685–??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/4/685.

Wang:2007:DPR

[WJM07] Tao Wang, Yuchen Jiao, and Craig Montell. Dissection of the pathway required for generation of vitamin A and for Drosophila phototransduction. Journal of Cell Biology, 177(2):305–??, April 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
177/2/305.

Winderlich:2009:VPC

[WKC^{+}09] Mark Winderlich, Linda Keller, Giuseppe Cagna, Andre Broermann, Olena Kamenyeva, Friedemann Kiefer, Urban Deutsch, Astrid F. Nottebaum, and Dietmar Vestweber. VE-PTP controls blood vessel development by balancing Tie-2 activity. Journal of Cell Biology, 185(4):657–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/185/4/657.

Wilson-Kubalek:2008:OSN

[WKCY^{+}08] Elizabeth M. Wilson-Kubalek, Iain M. Cheeseman, Craig Yoshioka, Arshad Desai, and Ronald A. Milligan. Orientation and structure of the Ndc80 complex on the microtubule lattice. Journal of Cell Biology, 182(6):1055–??, September 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/182/6/1055.

Wang:2006:PFN

[WKH^{+}06] Xueying Wang, Hui-Yi Kua, Yuanyu Hu, Ke Guo, Qi Zeng, Qiang Wu, Huck-Hui Ng, Gerard Karsenty, Benoit de Crombrugghe, James Yeh, and Baojie Li. p53 functions as a negative regulator of osteoblastogenesis, osteoblast-dependent osteoclastogenesis, and bone remodeling. Journal of Cell Biology, 172(1):115–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/1/115.

Winograd-Katz:2009:MAF

[WKIKG09] Sabina E. Winograd-Katz, Shalev Itzkovitz, Zvi Kam, and Benjamin Geiger. Multiparametric analysis of focal adhesion formation by RNAi-mediated gene knockdown. Journal of Cell Biology, 186(3):423–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/186/3/423.

Wilczynski:2008:IRM

[WKW^{+}08] Grzegorz M. Wilczynski, Filip A. Konopacki, Ewa Wilczek, Zofia Lasiecka, Adam Gorlewicz, Piotr Michaluk, Marcin Wawrzyniak, Monika Malinowska, Pawel Okulski, Lukasz R. Kolodziej, Witold Konopka, Kamila Duniec, Barbara Mioduszewska, Evgeni Nikolaev, Agnieszka Walczak, Dorota Owczarek, Dariusz C. Gorecki, Werner Zuschratter, Ole Petter Ottersen, and Leszek Kaczmarek. Important role of matrix metalloproteinase 9 in epileptogenesis. Journal of Cell Biology, 180(5):1021–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/180/5/1021.

Windoffer:2006:FAH

[WKWL06] Reinhard Windoffer, Anne Kölsch, Stefan Wöll, and Rudolf E. Leube. Focal adhesions are hotspots for keratin filament precursor formation. Journal of Cell Biology, 173(3):341–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/173/3/
341.

Wells:2007:CSA

[WL07] William A. Wells and Nicole LeBrasseur. Cells in the Sun: The American Society for Cell Biology, San Diego, CA, December 9–13, 2006. Journal of Cell Biology, 176(5):552–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/176/5/

Warner:2009:CAR

[WL09a] Stephen J. Warner and Gregory D. Longmore. Cdc42 antagonizes Rho1 activity at adherens junctions to limit epithelial cell apical tension. Journal of Cell Biology, 187(1):119–??, October 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/187/1/
119.

Warner:2009:DFR

[WL09b] Stephen J. Warner and Gregory D. Longmore. Distinct functions for Rho1 in maintaining adherens junctions and apical tension in remodeling epithelia. Journal of Cell Biology, 185 (6):1111–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/185/6/1111.

Wang:2008:JSP

[WLC08] Liwei Wang, Zhouhua Li, and Yu Cai. The JAK/STAT pathway positively regulates DPP signaling in the Drosophila germline stem cell niche. Journal of Cell Biology, 180(4):721–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/180/4/
721.

Wang:2009:TMA

[WLE09] Tao Wang, Uyen Lao, and Bruce A. Edgar. TOR-mediated autophagy regulates cell death in Drosophila neurodegenerative disease. Journal of Cell Biology, 186(5):703–??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/186/5/
703.

Wilhelmsen:2005:NNO

[WLK^{+}05] Kevin Wilhelmsen, Sandy H. M. Litjens, Ingrid Kuikman, Ntambua Tshimbalanga, Hans Janssen, Iman van den Bout, Karine Raymond, and Arnoud Sonnenberg. Nesprin-3, a novel outer nuclear membrane protein, associates with the cytoskeletal linker protein plectin. Journal of Cell Biology, 171(5): 799–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/5/799.

Weis:2008:CRP

[WLLF^{+}08] Sara M. Weis, Ssang-Taek Lim, Kimberly M. Lutu-Fuga, Leo A. Barnes, Xiao Lei Chen, Joachim R. Göthert, Tang-Long Shen, Jun-Lin Guan, David D. Schlaepfer, and David A. Cheresh. Compensatory role for Pyk2 during angiogenesis in adult mice lacking endothelial cell FAK. Journal of Cell Biology, 181 (1):43–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/1/43.

Wu:2009:CSS

[WLPB^{+}09] Zhuoru Wu, Katherine Luby-Phelps, Abhijit Bugde, Laura A. Molyneux, Bray Denard, Wen-Hong Li, Gürol M. Süel, and David L. Garbers. Capacity for stochastic self-renewal and differentiation in mammalian spermatogonial stem cells. Journal of Cell Biology, 187(4):513–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/187/4/513.

Wang:2009:HHM

[WLS^{+}09] Yanming Wang, Ming Li, Sonja Stadler, Sarah Correll, Pingxin Li, Danchen Wang, Ryo Hayama, Lauriebeth Leonelli, Hyunsil Han, Sergei A. Grigoryev, C. David Allis, and Scott A. Coonrod. Histone hypercitrullination mediates chromatin decondensation and neutrophil extracellular trap formation. Journal of Cell Biology, 184(2):205–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/184/2/205.

Wang:2006:RRP

[WLY06] Qiuyan Wang, Lianyun Li, and Yihong Ye. Regulation of retrotranslocation by p97-associated deubiquitinating enzyme ataxin-3. Journal of Cell Biology, 174(7):963–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/174/7/
963.

Wellbrock:2005:EEM

[WM05] Claudia Wellbrock and Richard Marais. Elevated expression of MITF counteracts B–RAF–stimulated melanocyte and melanoma cell proliferation. Journal of Cell Biology, 170(5): 703–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/5/703.

Wall:2009:PCP

[WMM^{+}09] Dana S. Wall, Alan J. Mears, Brian McNeill, Chantal Mazerolle, Sherry Thurig, Yaping Wang, Ryoichiro Kageyama, and Valerie A. Wallace. Progenitor cell proliferation in the retina is dependent on Notch-independent Sonic hedgehog/Hes1 activity. Journal of Cell Biology, 184(1):101–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/184/1/101.

Wang:2006:ASC

[WMS^{+}06] Weigang Wang, Ghassan Mouneimne, Mazen Sidani, Jeffrey Wyckoff, Xiaoming Chen, Anastasia Makris, Sumanta Goswami, Anne R. Bresnick, and John S. Condeelis. The activity status of cofilin is directly related to invasion, intravasation, and metastasis of mammary tumors. Journal of Cell Biology, 173(3):395–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/173/3/395.

Willaime-Morawek:2006:ECN

[WMSB^{+}06] Sandrine Willaime-Morawek, Raewyn M. Seaberg, Claudia Batista, Etienne Labbé, Liliana Attisano, Jessica A. Gorski, Kevin R. Jones, Angela Kam, Cindi M. Morshead, and Derek van der Kooy. Embryonic cortical neural stem cells migrate ventrally and persist as postnatal striatal stem cells. Journal of Cell Biology, 175(1):159–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/175/1/159.

Wolfe:2007:UDR

[WMT07] Breann L. Wolfe, Adriano Marchese, and JoAnn Trejo. Ubiquitination differentially regulates clathrin-dependent internalization of protease-activated receptor-1. Journal of Cell Biology, 177(5):905–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/177/5/905.

Witte:2008:MSS

[WNB08] Harald Witte, Dorothee Neukirchen, and Frank Bradke. Microtubule stabilization specifies initial neuronal polarization.

Journal of Cell Biology, 180(3):619–??, February 2008. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/180/3/619.

Wood:2005:C

[WNF^{+}05a] Christopher D. Wood, Takuya Nishigaki, Toshiaki Furuta, Shoji A. Baba, and Alberto Darszon. Correction. Journal of Cell Biology, 170(7):1171–??, September 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/170/7/1171.

Wood:2005:RTA

[WNF^{+}05b] Christopher D. Wood, Takuya Nishigaki, Toshiaki Furuta, Shoji A. Baba, and Alberto Darszon. Real-time analysis of the role of Ca^{2+} in flagellar movement and motility in single sea urchin sperm. Journal of Cell Biology, 169(5):725–??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/169/5/
725.

Watanabe:2009:ENR

[WNL^{+}09] Kazuhide Watanabe, Tadahiro Nagaoka, Joseph M. Lee, Caterina Bianco, Monica Gonzales, Nadia P. Castro, Maria Cristina Rangel, Kei Sakamoto, Youping Sun, Robert Callahan, and David S. Salomon. Enhancement of Notch receptor maturation and signaling sensitivity by Cripto-1. Journal of Cell Biology, 187(3):343–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/187/3/343.

Wu:2006:CIE

[WOP^{+}06] Boqian Wu, Kim Ottow, Peter Poulsen, Richard F. Gaber, Eva Albers, and Morten C. Kielland-Brandt. Competitive intraand extracellular nutrient sensing by the transporter homologue Ssy1p. Journal of Cell Biology, 173(3):327–??, May 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/173/3/327.

Woolner:2008:MAF

[WOWB08] Sarah Woolner, Lori L. O’Brien, Christiane Wiese, and William M. Bement. Myosin-10 and actin filaments are essential for mitotic spindle function. Journal of Cell Biology, 182(1):77–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/1/77.

Williams:2008:MRS

[WP08] Dumaine Williams and Jeffrey E. Pessin. Mapping of R– SNARE function at distinct intracellular GLUT4 trafficking steps in adipocytes. Journal of Cell Biology, 180(2):375–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
180/2/375.

West:2008:TLI

[WPC^{+}08] Michele A. West, Alan R. Prescott, Kui Ming Chan, Zhongjun Zhou, Stefan Rose-John, Jürgen Scheller, and Colin Watts. TLR ligand–induced podosome disassembly in dendritic cells is ADAM17 dependent. Journal of Cell Biology, 182(5):993– ??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/5/993.

Wong:2006:NCD

[WPDZ^{+}06] Christine E. Wong, Christian Paratore, María T. Dours-Zimmermann, Ariane Rochat, Thomas Pietri, Ueli Suter, Dieter R. Zimmermann, Sylvie Dufour, Jean Paul Thiery, Dies Meijer, Friedrich Beermann, Yann Barrandon, and Lukas Sommer. Neural crest–derived cells with stem cell features can be traced back to multiple lineages in the adult skin. Journal of Cell Biology, 175(6):1005–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/175/6/1005.

Weber:2007:BIA

[WPH^{+}07] Arnim Weber, Stefan A. Paschen, Klaus Heger, Florian Wilfling, Tobias Frankenberg, Heike Bauerschmitt, Barbara M. Seiffert, Susanne Kirschnek, Hermann Wagner, and Georg Häcker. Bim_{S}-induced apoptosis requires mitochondrial localization but not interaction with anti-apoptotic Bcl-2 proteins. Journal of Cell Biology, 177(4):625–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/177/4/625.

Wiedemann:2006:TDC

[WPL^{+}06] Agnès Wiedemann, Jayesh C. Patel, Jenson Lim, Andy Tsun, Yvette van Kooyk, and Emmanuelle Caron. Two distinct cytoplasmic regions of the β_{2} integrin chain regulate RhoA function during phagocytosis. Journal of Cell Biology, 172(7):1069–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
172/7/1069.

Wang:2008:AFA

[WPWF08] Xuan Wang, H. Amalia Pasolli, Trevor Williams, and Elaine Fuchs. AP-2 factors act in concert with Notch to orchestrate terminal differentiation in skin epidermis. Journal of Cell Biology, 183(1):37–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/1/37.

Wang:2009:MDA

[WRKC^{+}09] Xin Wang, Raghavendra Pralhada Rao, Teresa Kosakowska-Cholody, M. Athar Masood, Eileen Southon, Helin Zhang, Cyril Berthet, Kunio Nagashim, Timothy K. Veenstra, Lino Tessarollo, Usha Acharya, and Jairaj K. Acharya. Mitochondrial degeneration and not apoptosis is the primary cause of embryonic lethality in ceramide transfer protein mutant mice. Journal of Cell Biology, 184(1):143–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/184/1/143.

Wipff:2007:MCA

[WRMH07] Pierre-Jean Wipff, Daniel B. Rifkin, Jean-Jacques Meister, and Boris Hinz. Myofibroblast contraction activates latent TGF-β1 from the extracellular matrix. Journal of Cell Biology, 179(6): 1311–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/6/1311.

Wasmeier:2006:RRC

[WRP^{+}06] Christina Wasmeier, Maryse Romao, Lynn Plowright, Dorothy C. Bennett, Graça Raposo, and Miguel C. Seabra. Rab38 and Rab32 control post-Golgi trafficking of melanogenic enzymes. Journal of Cell Biology, 175(2):271–??, October 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/175/2/271.

Wei:2009:MSM

[WS09] Jen-Hsuan Wei and Joachim Seemann. The mitotic spindle mediates inheritance of the Golgi ribbon structure. Journal of Cell Biology, 184(3):391–??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/184/3/391.

Waschke:2006:IRA

[WSB^{+}06] Jens Waschke, Volker Spindler, Paola Bruggeman, Detlef Zillikens, Gudula Schmidt, and Detlev Drenckhahn. Inhibition of Rho A activity causes pemphigus skin blistering. Journal of Cell Biology, 175(5):721–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/175/5/721.

Wood:2009:PPR

[WSB^{+}09] Christopher S. Wood, Karl R. Schmitz, Nicholas J. Bessman, Thanuja Gangi Setty, Kathryn M. Ferguson, and Christopher G. Burd. PtdIns4 P recognition by Vps74/GOLPH3 links PtdIns 4-kinase signaling to retrograde Golgi trafficking. Journal of Cell Biology, 187(7):967–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/187/7/967.

Wright:2007:CMA

[WSFD07] Kevin M. Wright, Michelle I. Smith, Lila Farrag, and Mohanish Deshmukh. Chromatin modification of Apaf-1 restricts the apoptotic pathway in mature neurons. Journal of Cell Biology, 179(5):825–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/5/825.

Wu:2006:UGF

[WSG06] Rong Wu, Prim B. Singh, and David M. Gilbert. Uncoupling global and fine-tuning replication timing determinants for mouse pericentric heterochromatin. Journal of Cell Biology, 174(2):185–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/2/185.

Wu:2006:ACC

[WSK^{+}06] Jian-Qiu Wu, Vladimir Sirotkin, David R. Kovar, Matthew Lord, Christopher C. Beltzner, Jeffrey R. Kuhn, and Thomas D. Pollard. Assembly of the cytokinetic contractile ring from a broad band of nodes in fission yeast. Journal of Cell Biology, 174(3):391–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/3/391.

Waterhouse:2006:CLI

[WSS^{+}06] Nigel J. Waterhouse, Vivien R. Sutton, Karin A. Sedelies, Annette Ciccone, Misty Jenkins, Stephen J. Turner, Phillip I. Bird, and Joseph A. Trapani. Cytotoxic T lymphocyte–induced killing in the absence of granzymes A and B is unique and distinct from both apoptosis and perforin-dependent lysis. Journal of Cell Biology, 173(1):133–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/173/1/133.

Wang:2008:MID

[WTE^{+}08] Haibo Wang, Ambika Tewari, Steven Einheber, James L. Salzer, and Carmen V. Melendez-Vasquez. Myosin II has distinct functions in PNS and CNS myelin sheath formation. Journal of Cell Biology, 182(6):1171–??, September 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/182/6/1171.

Wu:2005:MMV

[WTH05] Xufeng S. Wu, Grace L. Tsan, and John A. Hammer. Melanophilin and myosin Va track the microtubule plus end on EB1. Journal of Cell Biology, 171(2):201–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/171/2/201.

Wu:2006:C

[WTH06] Xufeng S. Wu, Grace L. Tsan, and John A. Hammer. Correction. Journal of Cell Biology, 172(4):633–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/172/4/633.

Wong:2007:PMI

[WVB07] Kit Wong, Alexandra Van Keymeulen, and Henry R. Bourne. PDZRhoGEF and myosin II localize RhoA activity to the back of polarizing neutrophil-like cells. Journal of Cell Biology, 179(6):1141–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/179/6/1141.

Wiedemann:2007:SSM

[WvdLH^{+}07] Nils Wiedemann, Martin van der Laan, Dana P. Hutu, Peter Rehling, and Nikolaus Pfanner. Sorting switch of mitochondrial presequence translocase involves coupling of motor module to respiratory chain. Journal of Cell Biology, 179(6):1115–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
179/6/1115.

Willis:2007:ESS

[WvNS^{+}07] Dianna E. Willis, Erna A. van Niekerk, Yukio Sasaki, Mariano Mesngon, Tanuja T. Merianda, Gervan G. Williams, Marvin Kendall, Deanna S. Smith, Gary J. Bassell, and Jeffery L. Twiss. Extracellular stimuli specifically regulate localized levels of individual neuronal mRNAs. Journal of Cell Biology, 178(6): 965–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/6/965.

Wyatt:2007:PTP

[WWC^{+}07] Leila Wyatt, Carol Wadham, Lesley A. Crocker, Michael Lardelli, and Yeesim Khew-Goodall. The protein tyrosine phosphatase Pez regulates TGFβ, epithelial–mesenchymal transition, and organ development. Journal of Cell Biology, 178(7): 1223–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/7/1223.

Wirth:2006:SUT

[WWK^{+}06] Karin G. Wirth, Gordana Wutz, Nobuaki R. Kudo, Chantal Desdouets, Anders Zetterberg, Shahryar Taghybeeglu, Janina Seznec, Germain M. Ducos, Romeo Ricci, Nicole Firnberg, Jan-Michael Peters, and Kim Nasmyth. Separase: a universal trigger for sister chromatid disjunction but not chromosome cycle progression. Journal of Cell Biology, 172(6):847–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/172/6/
847.

Woll:2007:PMD

[WWL07] Stefan Wöll, Reinhard Windoffer, and Rudolf E. Leube. p38 MAPK-dependent shaping of the keratin cytoskeleton in cultured cells. Journal of Cell Biology, 177(5):795–??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/177/5/795.

Wang:2008:IUP

[WWO^{+}08] Jianjun Wang, Chuan-En Wang, Adam Orr, Suzanne Tydlacka, Shi-Hua Li, and Xiao-Jiang Li. Impaired ubiquitin–proteasome system activity in the synapses of Huntington’s disease mice. Journal of Cell Biology, 180(6):1177–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/180/6/1177.

Wittmann:2005:C

[WWS05a] Torsten Wittmann and Clare M. Waterman-Storer. Correction. Journal of Cell Biology, 171(2):393–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/171/2/393.

Wittmann:2005:SRC

[WWS05b] Torsten Wittmann and Clare M. Waterman-Storer. Spatial regulation of CLASP affinity for microtubules by Rac1 and GSK3β in migrating epithelial cells. Journal of Cell Biology, 169(6):929–??, June 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/6/929.

Wozniak:2006:RBC

[WWS^{+}06] Ann L. Wozniak, Xinmin Wang, Emily S. Stieren, Shelby G. Scarbrough, Cornelis J. Elferink, and Darren Boehning. Requirement of biphasic calcium release from the endoplasmic reticulum for Fas-mediated apoptosis. Journal of Cell Biology, 175(5):709–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/175/5/709.

Wordeman:2007:MFC

[WWvD07] Linda Wordeman, Michael Wagenbach, and George von Dassow. MCAK facilitates chromosome movement by promoting kinetochore microtubule turnover. Journal of Cell Biology, 179 (5):869–??, December 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/179/5/869.

Wang:2006:REP

[WWX^{+}06] Jiahu Wang, Ningning Wang, Jinling Xie, Staci C. Walton, Robert L. McKown, Ronald W. Raab, Peisong Ma, Shannon L. Beck, George L. Coffman, Isa M. Hussaini, and Gordon W. Laurie. Restricted epithelial proliferation by lacritin via PKCαdependent NFAT and mTOR pathways. Journal of Cell Biology, 174(5):689–??, August 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/5/689.

Wu:2005:STO

[WXM^{+}05] Ru Feng Wu, You Cheng Xu, Zhenyi Ma, Fiemu E. Nwariaku, George A. Sarosi, and Lance S. Terada. Subcellular targeting of oxidants during endothelial cell migration. Journal of Cell Biology, 171(5):893–??, December 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/171/5/893.

Wang:2005:LMM

[WZC^{+}05] Jing Wang, Qiwei Zhai, Ying Chen, Estelle Lin, Wei Gu, Michael W. McBurney, and Zhigang He. A local mechanism mediates NAD-dependent protection of axon degeneration. Journal of Cell Biology, 170(3):349–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/170/3/349.

Wang:2008:CEL

[WZC^{+}08] Xuelin Wang, Wei Zhang, Thomas Cheever, Valentin Schwarz, Karla Opperman, Harald Hutter, Deanna Koepp, and Lihsia Chen. The C. elegans L1CAM homologue LAD-2 functions as a coreceptor in MAB-20/Sema2-mediated axon guidance. Journal of Cell Biology, 180(1):233–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/180/1/233.

Witzel:2006:WCC

[WZCB^{+}06] Sabine Witzel, Vitaly Zimyanin, Filipa Carreira-Barbosa, Masazumi Tada, and Carl-Philipp Heisenberg. Wnt11 controls cell contact persistence by local accumulation of Frizzled 7 at the plasma membrane. Journal of Cell Biology, 175(5): 791–??, December 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/175/5/791.

Wasiak:2007:BBP

[WZM07] Sylwia Wasiak, Rodolfo Zunino, and Heidi M. McBride. Bax/ Bak promote sumoylation of DRP1 and its stable association with mitochondria during apoptotic cell death. Journal of Cell Biology, 177(3):439–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/177/3/439.

Wang:2008:SNA

[WZM^{+}08] Chuan-En Wang, Hui Zhou, John R. McGuire, Vincenzo Cerullo, Brendan Lee, Shi-Hua Li, and Xiao-Jiang Li. Suppression of neuropil aggregates and neurological symptoms by an intracellular antibody implicates the cytoplasmic toxicity of mutant huntingtin. Journal of Cell Biology, 181(5):803–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
181/5/803.

Walker:2005:MMM

[WZO^{+}05] William Walker, Zi-Qiang Zhou, Sara Ota, Anthony Wynshaw-Boris, and Peter J. Hurlin. Mnt–max to Myc–Max complex switching regulates cell cycle entry. Journal of Cell Biology, 169(3):405–??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/169/3/405.

Wakabayashi:2009:DRG

[WZW^{+}09] Junko Wakabayashi, Zhongyan Zhang, Nobunao Wakabayashi, Yasushi Tamura, Masahiro Fukaya, Thomas W. Kensler, Miho Iijima, and Hiromi Sesaki. The dynamin-related GTPase Drp1 is required for embryonic and brain development in mice. Journal of Cell Biology, 186(6):805–??, September 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/186/6/805.

Xu:2008:SAG

[XC08] Meng Xu and Peter R. Cook. Similar active genes cluster in specialized transcription factories. Journal of Cell Biology, 181 (4):615–??, May 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/4/615.

Xu:2009:RHH

[XGX^{+}09] Zhuojin Xu, Qiang Gong, Bin Xia, Benjamin Groves, Marc Zimmermann, Chris Mugler, Dezhi Mu, Brian Matsumoto, Matthew Seaman, and Dzwokai Ma. A role of histone H3 lysine 4 methyltransferase components in endosomal trafficking. Journal of Cell Biology, 186(3):343–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/186/3/343.

Xiang:2006:TST

[Xia06] Xin Xiang. A +TIP for a smooth trip. Journal of Cell Biology, 172(5):651–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/5/651.

Xie:2005:IND

[XJF05] Liang Xie, Randall S. Johnson, and Robert S. Freeman. Inhibition of NGF deprivation–induced death by low oxygen involves suppression of BIM_{EL} and activation of HIF-1. Journal of Cell Biology, 168(6):911–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/6/911.

Xiong:2007:NRI

[XKYW07] Lei Xiong, Fei Kou, Ying Yang, and Jiarui Wu. A novel role for IGF-1R in p53-mediated apoptosis through translational modulation of the p53-Mdm2 feedback loop. Journal of Cell Biology, 178(6):995–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/6/995.

Xie:2007:FVD

[XLH^{+}07] Mian Xie, Xiang Li, Jing Han, Daniel L. Vogt, Silke Wittemann, Melanie D. Mark, and Stefan Herlitze. Facilitation versus depression in cultured hippocampal neurons determined by targeting of Ca^{2+} channel Ca_{v} β_{4} versus Ca_{v} β_{2} subunits to synaptic terminals. Journal of Cell Biology, 178(3):489–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/178/3/
489.

Xiong:2009:CSC

[XMA^{+}09] Jian-Ping Xiong, Bhuvaneshwari Mahalingham, Jose Luis Alonso, Laura Ann Borrelli, Xianliang Rui, Saurabh Anand, Bradley T. Hyman, Thomas Rysiok, Dirk Müller-Pompalla, Simon L. Goodman, and M. Amin Arnaout. Crystal structure of the complete integrin αVβ3 ectodomain plus an α /β transmembrane fragment. Journal of Cell Biology, 186(4):589–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
186/4/589.

Xu:2007:LCI

[XMSYJ07] Xuehua Xu, Martin Meier-Schellersheim, Jianshe Yan, and Tian Jin. Locally controlled inhibitory mechanisms are involved in eukaryotic GPCR-mediated chemosensing. Journal of Cell Biology, 178(1):141–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/1/141.

Xu:2009:SAS

[XNM^{+}09] Ren Xu, Celeste M. Nelson, John L. Muschler, Mandana Veiseh, Barbara K. Vonderhaar, and Mina J. Bissell. Sustained activation of STAT5 is essential for chromatin remodeling and maintenance of mammary-specific function. Journal of Cell Biology, 184(1):57–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/184/1/57.

Xu:2009:IAB

[XOV^{+}09] Zhenjie Xu, Hiromi Ogawa, Paola Vagnarelli, Jan H. Bergmann, Damien F. Hudson, Sandrine Ruchaud, Tatsuo Fukagawa, William C. Earnshaw, and Kumiko Samejima. INCENP-aurora B interactions modulate kinase activity and chromosome passenger complex localization. Journal of Cell Biology, 187(5): 637–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/187/5/637.

Xian:2005:PEF

[XSVGR05] Wa Xian, Kathryn L. Schwertfeger, Tracy Vargo-Gogola, and Jeffrey M. Rosen. Pleiotropic effects of FGFR1 on cell proliferation, survival, and migration in a 3D mammary epithelial cell model. Journal of Cell Biology, 171(4):663–??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/171/4/
663.

Xu:2007:MRN

[XYC^{+}07] Lan Xu, Xiaohao Yao, Xiaochu Chen, Peiyuan Lu, Biliang Zhang, and Y. Tony Ip. Msk is required for nuclear import of TGF-β /BMP-activated Smads. Journal of Cell Biology, 178(6):981–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/178/6/981.

Yanagisawa:2006:PCE

[YA06] Masahiro Yanagisawa and Panos Z. Anastasiadis. p120 catenin is essential for mesenchymal cadherin-mediated regulation of cell motility and invasiveness. Journal of Cell Biology, 174(7): 1087–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/7/1087.

Yu:2005:FRA

[YAV^{+}05] Vionnie W. C. Yu, Gourgen Ambartsoumian, Lieve Verlinden, Janet M. Moir, Josée Prud’homme, Claude Gauthier, Peter J. Roughley, and René St-Arnaud. FIAT represses ATF4mediated transcription to regulate bone mass in transgenic mice. Journal of Cell Biology, 169(4):591–??, May 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/169/4/591.

Yin:2006:ENF

[YBK^{+}06] Xinghua Yin, Rena C. Baek, Daniel A. Kirschner, Alan Peterson, Yasuhisa Fujii, Klaus-Armin Nave, Wendy B. Macklin, and Bruce D. Trapp. Evolution of a neuroprotective function of central nervous system myelin. Journal of Cell Biology, 172 (3):469–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/3/469.

Yang:2005:CCA

[YBT^{+}05] Dongren Yang, Jesse Bierman, Yukie S. Tarumi, Yong-Ping Zhong, Reshma Rangwala, Thomas M. Proctor, Yuko Miyagoe-Suzuki, Shin’ichi Takeda, Jeffrey H. Miner, Larry S. Sherman, Bruce G. Gold, and Bruce L. Patton. Coordinate control of axon defasciculation and myelination by laminin-2 and -8. Journal of Cell Biology, 168(4):655–??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/168/4/655.

Yamamoto:2009:ABS

[YBT^{+}09] Masamichi Yamamoto, Hideyuki Beppu, Katsuyoshi Takaoka, Chikara Meno, En Li, Kohei Miyazono, and Hiroshi Hamada. Antagonism between Smad1 and Smad2 signaling determines the site of distal visceral endoderm formation in the mouse embryo. Journal of Cell Biology, 184(2):323–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/184/2/323.

Yu:2005:MNA

[YCK^{+}05] W. Haung Yu, Ana Maria Cuervo, Asok Kumar, Corrinne M. Peterhoff, Stephen D. Schmidt, Ju-Hyun Lee, Panaiyur S. Mohan, Marc Mercken, Mark R. Farmery, Lars O. Tjernberg, Ying Jiang, Karen Duff, Yasuo Uchiyama, Jan Näslund, Paul M. Mathews, Anne M. Cataldo, and Ralph A. Nixon. Macroautophagy — a novel β-amyloid peptide-generating pathway activated in Alzheimer’s disease. Journal of Cell Biology, 171 (1):87–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/1/87.

Yang:2008:RVM

[YCM^{+}08] Ge Yang, Lisa A. Cameron, Paul S. Maddox, Edward D. Salmon, and Gaudenz Danuser. Regional variation of microtubule flux reveals microtubule organization in the metaphase meiotic spindle. Journal of Cell Biology, 182(4):631–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/182/4/
631.

Yamamoto:2006:AMC

[YCR06] Ai Yamamoto, M. Laura Cremona, and James E. Rothman. Autophagy-mediated clearance of huntingtin aggregates triggered by the insulin-signaling pathway. Journal of Cell Biology, 172(5):719–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/5/719.

Yue:2008:DSC

[YCX^{+}08] Zuojun Yue, Ana Carvalho, Zhenjie Xu, Xuemei Yuan, Stefano Cardinale, Susana Ribeiro, Fan Lai, Hiromi Ogawa, Elisabet Gudmundsdottir, Reto Gassmann, Ciaran G. Morrison, Sandrine Ruchaud, and William C. Earnshaw. Deconstructing Survivin: comprehensive genetic analysis of Survivin function by conditional knockout in a vertebrate cell line. Journal of Cell Biology, 183(2):279–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/183/2/279.

Yuen:2008:WC

[YD08] Karen W. Yuen and Arshad Desai. The wages of CIN. Journal of Cell Biology, 180(4):661–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/180/4/661.

Yang:2006:MSR

[YDL^{+}06] Fang Yang, Rabindranath De La Fuente, N. Adrian Leu, Claudia Baumann, K. John McLaughlin, and P. Jeremy Wang. Mouse SYCP2 is required for synaptonemal complex assembly and chromosomal synapsis during male meiosis. Journal of Cell Biology, 173(4):497–??, May 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/173/4/497.

Yang:2008:MTE

[YEL^{+}08] Fang Yang, Sigrid Eckardt, N. Adrian Leu, K. John McLaughlin, and Peijing Jeremy Wang. Mouse TEX15 is essential for DNA double-strand break repair and chromosomal synapsis during male meiosis. Journal of Cell Biology, 180(4): 673–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/4/673.

Yamashita:2008:ACB

[YF08] Yukiko M. Yamashita and Margaret T. Fuller. Asymmetric centrosome behavior and the mechanisms of stem cell division.

Journal of Cell Biology, 180(2):261–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/180/2/261.

Yokoyama:2008:CRD

[YGR^{+}08] Hideki Yokoyama, Oliver J. Gruss, Sofia Rybina, Maïwen Caudron, Malgorzata Schelder, Matthias Wilm, Iain W. Mattaj, and Eric Karsenti. Cdk11 is a RanGTP-dependent microtubule stabilization factor that regulates spindle assembly rate. Journal of Cell Biology, 180(5):867–??, March 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/180/5/867.

Yan:2006:RDA

[YGW06] Dong Yan, Li Guo, and Yizheng Wang. Requirement of dendritic Akt degradation by the ubiquitin–proteasome system for neuronal polarity. Journal of Cell Biology, 174(3):415–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/174/3/
415.

Yamaguchi:2005:PCI

[YGY^{+}05] Tomoya Yamaguchi, Hidemasa Goto, Tomoya Yokoyama, Herman Silljé, Anja Hanisch, Andreas Uldschmid, Yasushi Takai, Takashi Oguri, Erich A. Nigg, and Masaki Inagaki. Phosphorylation by Cdk1 induces Plk1-mediated vimentin phosphorylation during mitosis. Journal of Cell Biology, 171(3): 431–??, November 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/3/431.

Yim:2006:CCC

[YHC^{+}06] Hyungshin Yim, In Sun Hwang, Joon-Seok Choi, Kwang-Hoon Chun, Ying Hua Jin, Young-Mi Ham, Kwang Youl Lee, and Seung Ki Lee. Cleavage of Cdc6 by caspase-3 promotes ATM/ ATR kinase-mediated apoptosis of HeLa cells. Journal of Cell Biology, 174(1):77–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/1/77.

Yeung:2009:CPM

[YHD^{+}09] Tony Yeung, Bryan Heit, Jean-Francois Dubuisson, Gregory D. Fairn, Basil Chiu, Robert Inman, Andras Kapus, Michele Swanson, and Sergio Grinstein. Contribution of phosphatidylserine to membrane surface charge and protein targeting during phagosome maturation. Journal of Cell Biology, 185(5):917–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
185/5/917.

Yoshigi:2005:MFM

[YHJ^{+}05] Masaaki Yoshigi, Laura M. Hoffman, Christopher C. Jensen, H. Joseph Yost, and Mary C. Beckerle. Mechanical force mobilizes zyxin from focal adhesions to actin filaments and regulates cytoskeletal reinforcement. Journal of Cell Biology, 171 (2):209–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/2/209.

Yamasaki:2009:ACT

[YIfHA09] Satoshi Yamasaki, Pavel Ivanov, Guo fu Hu, and Paul Anderson. Angiogenin cleaves tRNA and promotes stress-induced translational repression. Journal of Cell Biology, 185(1):35–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
185/1/35.

Yang:2009:PBI

[YJJ^{+}09] Feng Yang, Hyun-Soo Je, Yuanyuan Ji, Guhan Nagappan, Barbara Hempstead, and Bai Lu. Pro-BDNF-induced synaptic depression and retraction at developing neuromuscular synapses. Journal of Cell Biology, 185(4):727–??, May 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/185/4/727.

Yu:2007:AKI

[YK07] Hong-Guo Yu and Douglas Koshland. The Aurora kinase Ipl1 maintains the centromeric localization of PP2A to protect cohesin during meiosis. Journal of Cell Biology, 176(7):911–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
176/7/911.

Yang:2009:CSM

[YKBR09] Zhenye Yang, Alison E. Kenny, Daniela A. Brito, and Conly L. Rieder. Cells satisfy the mitotic checkpoint in Taxol, and do so faster in concentrations that stabilize syntelic attachments. Journal of Cell Biology, 186(5):675–??, September 2009. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/186/5/675.

Yamamoto:2008:SCA

[YKH^{+}08] Ayumu Yamamoto, Kenji Kitamura, Daisuke Hihara, Yukinobu Hirose, Satoshi Katsuyama, and Yasushi Hiraoka. Spindle checkpoint activation at meiosis I advances anaphase II onset via meiosis-specific APC/C regulation. Journal of Cell Biology, 182(2):277–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/182/2/277.

Yin:2008:CPC

[YKP^{+}08] Chunyue Yin, Maria Kiskowski, Philippe-Alexandre Pouille, Emmanuel Farge, and Lilianna Solnica-Krezel. Cooperation of polarized cell intercalations drives convergence and extension of presomitic mesoderm during zebrafish gastrulation. Journal of Cell Biology, 180(1):221–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/1/221.

Yamaguchi:2005:MMI

[YLK^{+}05] Hideki Yamaguchi, Mike Lorenz, Stephan Kempiak, Corina Sarmiento, Salvatore Coniglio, Marc Symons, Jeffrey Segall, Robert Eddy, Hiroaki Miki, Tadaomi Takenawa, and John Condeelis. Molecular mechanisms of invadopodium formation. Journal of Cell Biology, 168(3):441–??, January 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/168/3/441.

Yan:2006:DRX

[YLM06a] Shan Yan, Howard D. Lindsay, and W. Matthew Michael. Direct requirement for Xmus101 in ATR-mediated phosphorylation of Claspin bound Chk1 during checkpoint signaling. Journal of Cell Biology, 173(2):181–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/173/2/181.

Yuan:2006:CEI

[YLM^{+}06b] Weiping Yuan, Tina M. Leisner, Andrew W. McFadden, Zhengyan Wang, Mark K. Larson, Shantres Clark, Christel Boudignon-Proudhon, Stephen C.-T. Lam, and Leslie V. Parise. CIB1 is an endogenous inhibitor of agonist-induced integrin αIIbβ3 activation. Journal of Cell Biology, 172(2):169–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
172/2/169.

Yang:2006:NIT

[YM06] Weidong Yang and Siegfried M. Musser. Nuclear import time and transport efficiency depend on importin β concentration. Journal of Cell Biology, 174(7):951–??, September 2006. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/174/7/951.

Yan:2009:TDP

[YM09] Shan Yan and W. Matthew Michael. TopBP1 and DNA polymerase-α directly recruit the 9-1-1 complex to stalled DNA replication forks. Journal of Cell Biology, 184(6):793–??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/184/6/
793.

Yoneda:2005:RKI

[YMC05] Atsuko Yoneda, Hinke A. B. Multhaupt, and John R. Couchman. The Rho kinases I and II regulate different aspects of myosin II activity. Journal of Cell Biology, 170(3):443–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
170/3/443.

Yamauchi:2008:EDA

[YMCT08] Junji Yamauchi, Yuki Miyamoto, Jonah R. Chan, and Akito Tanoue. ErbB2 directly activates the exchange factor Dock7 to promote Schwann cell migration. Journal of Cell Biology, 181 (2):351–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/181/2/351.

Yang:2007:AIM

[YMD^{+}07] Zhiwei Yang, Zhenyu Mu, Branka Dabovic, Vladimir Jurukovski, Dawen Yu, Joanne Sung, Xiaozhong Xiong, and John S. Munger. Absence of integrin-mediated TGFβ1 activation in vivo recapitulates the phenotype of TGFβ1-null mice.

Journal of Cell Biology, 176(6):787–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/176/6/787.

Yuan:2009:CMS

[YMDL09] Jian Yuan, Katherine Minter-Dykhouse, and Zhenkun Lou. A c–Myc–SIRT1 feedback loop regulates cell growth and transformation. Journal of Cell Biology, 185(2):203–??, April 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/185/2/203.

Yan:2005:AXW

[YMTC05] Hong Yan, Jill McCane, Thomas Toczylowski, and Chinyi Chen. Analysis of the Xenopus Werner syndrome protein in DNA double-strand break repair. Journal of Cell Biology, 171 (2):217–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/171/2/217.

Yamada:2007:LZR

[YN07] Soichiro Yamada and W. James Nelson. Localized zones of Rho and Rac activities drive initiation and expansion of epithelial cell–cell adhesion. Journal of Cell Biology, 178(3):517–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/178/3/
517.

Yang:2008:BIC

[YNL^{+}08] Hui-Ju Yang, Hideki Nakanishi, Song Liu, James A. McNew, and Aaron M. Neiman. Binding interactions control SNARE specificity in vivo. Journal of Cell Biology, 183(6):1089–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
183/6/1089.

Yang:2007:SRA

[YOHR07] Yang Yang, Yasuhiro Ogawa, Kristian L. Hedstrom, and Matthew N. Rasband. βiv spectrin is recruited to axon initial segments and nodes of Ranvier by ankyrinG. Journal of Cell Biology, 176(4):509–??, February 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/176/4/509.

Yamada:2005:APS

[YOS^{+}05] Eijiro Yamada, Shuichi Okada, Tsugumichi Saito, Kihachi Ohshima, Minoru Sato, Takafumi Tsuchiya, Yutaka Uehara, Hiroyuki Shimizu, and Masatomo Mori. Akt2 phosphorylates Synip to regulate docking and fusion of GLUT4-containing vesicles. Journal of Cell Biology, 168(6):921–??, March 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/168/6/921.

Yoshida:2006:PUE

[YOSM06] Hiderou Yoshida, Masaya Oku, Mie Suzuki, and Kazutoshi Mori. pXBP1(U) encoded in XBP1 pre-mRNA negatively regulates unfolded protein response activator pXBP1(S) in mammalian ER stress response. Journal of Cell Biology, 172(4): 565–??, February 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/172/4/565.

Yang:2008:DHS

[YOY08] Chun Yang, Heather A. Owen, and Pinfen Yang. Dimeric heat shock protein 40 binds radial spokes for generating coupled power strokes and recovery strokes of 9 + 2 flagella. Journal of Cell Biology, 180(2):403–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/2/403.

Yuce:2005:ECC

[YPG05] Özlem Yüce, Alisa Piekny, and Michael Glotzer. An ECT2– centralspindlin complex regulates the localization and function of RhoA. Journal of Cell Biology, 170(4):571–??, August 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/170/4/571.

Ydenberg:2009:ARF

[YR09] Casey A. Ydenberg and Mark D. Rose. Antagonistic regulation of Fus2p nuclear localization by pheromone signaling and the cell cycle. Journal of Cell Biology, 184(3):409–??, February 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/184/3/

Yokoyama:2009:IRD

[YRSM^{+}09] Hideki Yokoyama, Sofia Rybina, Rachel Santarella-Mellwig, Iain W. Mattaj, and Eric Karsenti. ISWI is a RanGTPdependent MAP required for chromosome segregation. Journal of Cell Biology, 187(6):813–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/187/6/813.

Yi:2007:GWR

[YSB^{+}07] Caroline H. Yi, Dodzie K. Sogah, Michael Boyce, Alexei Degterev, Dana E. Christofferson, and Junying Yuan. A genome-wide RNAi screen reveals multiple regulators of caspase activation. Journal of Cell Biology, 179(4):619–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/179/4/
619.

Yu:2006:MRH

[YSP^{+}06] Sidney Yu, Ayano Satoh, Marc Pypaert, Karl Mullen, Jesse C. Hay, and Susan Ferro-Novick. mBet3p is required for homotypic COPII vesicle tethering in mammalian cells. Journal of Cell Biology, 174(3):359–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/3/359.

Yamasaki:2007:ZNI

[YSSH^{+}07] Satoru Yamasaki, Kumiko Sakata-Sogawa, Aiko Hasegawa, Tomoyuki Suzuki, Koki Kabu, Emi Sato, Tomohiro Kurosaki, Susumu Yamashita, Makio Tokunaga, Keigo Nishida, and Toshio Hirano. Zinc is a novel intracellular second messenger. Journal of Cell Biology, 177(4):637–??, May 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/177/4/637.

Yin:2007:RIA

[YTB07] Viravuth P. Yin, Carl S. Thummel, and Arash Bashirullah. Down-regulation of inhibitor of apoptosis levels provides competence for steroid-triggered cell death. Journal of Cell Biology, 178(1):85–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/1/85.

Yamakoshi:2009:RTV

[YTH^{+}09] Kimi Yamakoshi, Akiko Takahashi, Fumiko Hirota, Rika Nakayama, Naozumi Ishimaru, Yoshiaki Kubo, David J. Mann, Masako Ohmura, Atsushi Hirao, Hideyuki Saya, Seiji Arase, Yoshio Hayashi, Kazuki Nakao, Mitsuru Matsumoto, Naoko Ohtani, and Eiji Hara. Real-time in vivo imaging of p16^{Ink4a} reveals cross talk with p53. Journal of Cell Biology, 186(3): 393–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/186/3/393.

Yokoyama:2007:RCI

[YTY^{+}07] Takashi Yokoyama, Kazunori Takano, Akira Yoshida, Fumiko Katada, Peng Sun, Tadaomi Takenawa, Toshiwo Andoh, and Takeshi Endo. DA-raf1, a competent intrinsic dominantnegative antagonist of the Ras–ERK pathway, is required for myogenic differentiation. Journal of Cell Biology, 177(5):781– ??, June 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
177/5/781.

Yu:2006:SAM

[Yu06] Hongtao Yu. Structural activation of Mad2 in the mitotic spindle checkpoint: the two-state Mad2 model versus the Mad2 template model. Journal of Cell Biology, 173(2):153–??, April 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/173/2/
153.

Yamamoto:2008:SFK

[YWEK08] Takaharu G. Yamamoto, Sonoko Watanabe, Anthony Essex, and Risa Kitagawa. SPDL-1 functions as a kinetochore receptor for MDF-1 in Caenorhabditis elegans. Journal of Cell Biology, 183(2):187–??, October 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/2/187.

Yam:2007:AMN

[YWJ^{+}07] Patricia T. Yam, Cyrus A. Wilson, Lin Ji, Benedict Hebert, Erin L. Barnhart, Natalie A. Dye, Paul W. Wiseman, Gaudenz Danuser, and Julie A. Theriot. Actin–myosin network reorganization breaks symmetry at the cell rear to spontaneously initiate polarized cell motility. Journal of Cell Biology, 178(7): 1207–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/7/1207.

Yap:2008:SER

[YWK^{+}08] Chan Choo Yap, Dolora Wisco, Pekka Kujala, Zofia M. Lasiecka, Johanna T. Cannon, Michael C. Chang, Harald Hirling, Judith Klumperman, and Bettina Winckler. The somatodendritic endosomal regulator NEEP21 facilitates axonal targeting of L1/NgCAM. Journal of Cell Biology, 180(4): 827–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/4/827.

Yun:2008:NPB

[YWM^{+}08] Chawon Yun, Yonggang Wang, Debaditya Mukhopadhyay, Peter Backlund, Nagamalleswari Kolli, Alfred Yergey, Keith D. Wilkinson, and Mary Dasso. Nucleolar protein B23/ nucleophosmin regulates the vertebrate SUMO pathway through SENP3 and SENP5 proteases. Journal of Cell Biology, 183(4):589–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/183/4/589.

Yung:2007:SAL

[YWRA07] Yuval Yung, Janice L. Walker, James M. Roberts, and Richard K. Assoian. A Skp2 autoinduction loop and restriction point control. Journal of Cell Biology, 178(5):741–??, August 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/178/5/
741.

Yago:2005:C

[YWW^{+}05] Tadayuki Yago, Jianhua Wu, C. Diana Wey, Arkadiusz G. Klopocki, Cheng Zhu, and Rodger P. McEver. Correction. Journal of Cell Biology, 168(6):975–??, March 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/168/6/975.

Yu:2008:CDM

[YWW^{+}08] Zuoren Yu, Chenguang Wang, Min Wang, Zhiping Li, Mathew C. Casimiro, Manran Liu, Kongming Wu, James Whittle, Xiaoming Ju, Terry Hyslop, Peter McCue, and Richard G. Pestell. A cyclin D1/microRNA 17/20 regulatory feedback loop in control of breast cancer cell proliferation. Journal of Cell Biology, 182(3):509–??, August 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/182/3/509.

Yang:2005:KMT

[yYMM05] Hsin ya Yang, Paul E. Mains, and Francis J. McNally. Kinesin-1 mediates translocation of the meiotic spindle to the oocyte cortex through KCA-1, a novel cargo adapter. Journal of Cell Biology, 169(3):447–??, May 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/169/3/447.

Yuneva:2007:DGG

[YZO^{+}07] Mariia Yuneva, Nicola Zamboni, Peter Oefner, Ravi Sachidanandam, and Yuri Lazebnik. Deficiency in glutamine but not glucose induces MYC-dependent apoptosis in human cells. Journal of Cell Biology, 178(1):93–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/178/1/93.

Zhang:2007:BAEa

[ZAM07a] Jiayin Zhang, Sana Ahmad, and Yinghui Mao. BubR1 and APC/EB1 cooperate to maintain metaphase chromosome alignment. Journal of Cell Biology, 178(5):773–??, August 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/178/5/773.

Zhang:2007:BAEb

[ZAM07b] Jiayin Zhang, Sana Ahmad, and Yinghui Mao. BubR1 and APC/EB1 cooperate to maintain metaphase chromosome alignment. Journal of Cell Biology, 179(2):357–??, October 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/179/2/357.

Zaidel-Bar:2009:ECI

[ZB09] Ronen Zaidel-Bar. Evolution of complexity in the integrin adhesome. Journal of Cell Biology, 186(3):317–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/186/3/317.

Zheng:2007:FBR

[ZBKJ07] Lihua Zheng, Hwa-Jin Baek, Gerard Karsenty, and Monica J. Justice. Filamin B represses chondrocyte hypertrophy in a Runx2/Smad3-dependent manner. Journal of Cell Biology, 178 (1):121–??, July 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/178/1/121.

Zhu:2008:FPM

[ZCJ^{+}08] Hui Zhu, Judith A. Coppinger, Chang-Young Jang, John R. Yates, and Guowei Fang. FAM29A promotes microtubule amplification via recruitment of the NEDD1-γ-tubulin complex to the mitotic spindle. Journal of Cell Biology, 183(5):835–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
183/5/835.

Zappulla:2006:MME

[ZDD^{+}06] Jacques P. Zappulla, Patrice Dubreuil, Sabine Desbois, Sébastien Létard, Nadine Ben Hamouda, Marc Daëron, Georges Delsol, Michel Arock, and Roland S. Liblau. Mastocytosis in mice expressing human Kit receptor with the activating Asp816Val mutation. Journal of Cell Biology, 172(1):??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
172/1/i1.

Zhong:2006:BDR

[ZDMD06] Fei Zhong, Michael C. Davis, Karen S. McColl, and Clark W. Distelhorst. Bcl-2 differentially regulates Ca^{2+} signals according to the strength of T cell receptor activation. Journal of Cell Biology, 172(1):127–??, January 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/172/1/127.

Zheng:2008:DPP

[ZEC^{+}08] Dinghai Zheng, Nader Ezzeddine, Chyi-Ying A. Chen, Wenmiao Zhu, Xiangwei He, and Ann-Bin Shyu. Deadenylation is prerequisite for P-body formation and mRNA decay in mammalian cells. Journal of Cell Biology, 182(1):89–??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/182/1/89.

Zaoui:2008:MRM

[ZHI^{+}08] Kossay Zaoui, Stéphane Honoré, Daniel Isnardon, Diane Braguer, and Ali Badache. Memo-RhoA-mDia1 signaling controls microtubules, the actin network, and adhesion site formation in migrating cells. Journal of Cell Biology, 183(3):401–??, November 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/183/3/
401.

Zambetti:2006:SPT

[ZHS06] Gerard P. Zambetti, Edwin M. Horwitz, and Ernestina Schipani. Skeletons in the p53 tumor suppressor closet: genetic evidence that p53 blocks bone differentiation and development. Journal of Cell Biology, 172(6):795–??, March 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/172/6/795.

Zou:2008:MDK

[ZHYE08] Jianwei Zou, Mark A. Hallen, Christine D. Yankel, and Sharyn A. Endow. A microtubule-destabilizing kinesin motor regulates spindle length and anchoring in oocytes. Journal of Cell Biology, 180(3):459–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/180/3/459.

Zeidan:2008:RCC

[ZJH08] Youssef H. Zeidan, Russell W. Jenkins, and Yusuf A. Hannun. Remodeling of cellular cytoskeleton by the acid sphingomyelinase/ceramide pathway. Journal of Cell Biology, 181(2):335–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
181/2/335.

Zhao:2008:MCH

[ZKG^{+}08] Rongmin Zhao, Yoshito Kakihara, Anna Gribun, Jennifer Huen, Guocheng Yang, May Khanna, Michael Costanzo, Renée L. Brost, Charles Boone, Timothy R. Hughes, Christopher M. Yip, and Walid A. Houry. Molecular chaperone Hsp90 stabilizes Pih1/Nop17 to maintain R2TP complex activity that regulates snoRNA accumulation. Journal of Cell Biology, 180 (3):563–??, February 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/180/3/563.

Zhang:2006:EMT

[ZKN^{+}06] Gang Zhang, Rohini Kashimshetty, Kwee Eng Ng, Heng Buck Tan, and Foong May Yeong. Exit from mitosis triggers Chs2p transport from the endoplasmic reticulum to mother–daughter neck via the secretory pathway in budding yeast. Journal of Cell Biology, 174(2):207–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/174/2/207.

Zou:2007:SCS

[ZKR^{+}07] Wei Zou, Hideki Kitaura, Jennifer Reeve, Fanxin Long, Victor L. J. Tybulewicz, Sanford J. Shattil, Mark H. Ginsberg, F. Patrick Ross, and Steven L. Teitelbaum. Syk, c–Src, the α vβ3 integrin, and ITAM immunoreceptors, in concert, regulate osteoclastic bone resorption. Journal of Cell Biology, 176 (6):877–??, March 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/176/6/877.

Zhang:2008:HMC

[ZKS^{+}08] Nenggang Zhang, Sergey G. Kuznetsov, Shyam K. Sharan, Kaiyi Li, Pulivarthi H. Rao, and Debananda Pati. A handcuff model for the cohesin complex. Journal of Cell Biology, 183(6): 1019–??, December 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/183/6/1019.

Zolov:2005:CDB

[ZL05] Sergey N. Zolov and Vladimir V. Lupashin. Cog3p depletion blocks vesicle-mediated Golgi retrograde trafficking in HeLa cells. Journal of Cell Biology, 168(5):747–??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/168/5/747.

Zou:2005:C

[ZLB^{+}05] Chaozhong Zou, Jun Li, Yujie Bai, William T. Gunning, David E. Wazer, Vimla Band, and Qingshen Gao. Centrobin. Journal of Cell Biology, 171(3):437–??, November 2005. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/171/3/437.

Zisman:2007:PMC

[ZMA^{+}07] Sophie Zisman, Karen Marom, Oshri Avraham, Lilah Rinsky-Halivni, Uri Gai, Gilit Kligun, Vered Tzarfaty-Majar, Tatsuo Suzuki, and Avihu Klar. Proteolysis and membrane capture of F-spondin generates combinatorial guidance cues from a single molecule. Journal of Cell Biology, 178(7):1237–??, September 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/178/7/
1237.

Zaitseva:2005:CIF

[ZMGC05] Elena Zaitseva, Aditya Mittal, Diane E. Griffin, and Leonid V. Chernomordik. Class II fusion protein of alphaviruses drives membrane fusion through the same pathway as class I proteins. Journal of Cell Biology, 169(1):167–??, April 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/169/1/167.

Zhu:2009:GST

[ZMH^{+}09] Qubo Zhu, Lingjun Meng, Joseph K. Hsu, Tao Lin, Jun Teishima, and Robert Y. L. Tsai. GNL3L stabilizes the TRF1 complex and promotes mitotic transition. Journal of Cell Biology, 185(5):827–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/5/827.

Zoeller:2008:CFP

[ZMW^{+}08] Jason J. Zoeller, Angela McQuillan, John Whitelock, Shiu-Ying Ho, and Renato V. Iozzo. A central function for perlecan in skeletal muscle and cardiovascular development. Journal of Cell Biology, 181(2):381–??, April 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/181/2/381.

Zhang:2009:RNG

[ZNP^{+}09] Xin Zhang, Irina Neganova, Stefan Przyborski, Chunbo Yang, Michael Cooke, Stuart P. Atkinson, George Anyfantis, Stefan Fenyk, W. Nicol Keith, Stacey F. Hoare, Owen Hughes, Tom Strachan, Miodrag Stojkovic, Philip W. Hinds, Lyle Armstrong, and Majlinda Lako. A role for NANOG in G1 to S transition in human embryonic stem cells through direct binding of CDK6 and CDC25A. Journal of Cell Biology, 184(1):67–??, January 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/184/1/
67.

Zhang:2008:MAF

[ZOH^{+}08] Xiaoyu Zhang, Kelly Orlando, Bing He, Fengong Xi, Jian Zhang, Allison Zajac, and Wei Guo. Membrane association and functional regulation of Sec3 by phospholipids and Cdc42. Journal of Cell Biology, 180(1):145–??, January 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/180/1/145.

Zagorska:2007:RAL

[ZPGB^{+}07] Anna Zagórska, Eulalia Pozo-Guisado, Jérôme Boudeau, Alberto C. Vitari, Fatema H. Rafiqi, Jacob Thastrup, Maria Deak, David G. Campbell, Nick A. Morrice, Alan R. Prescott, and Dario R. Alessi. Regulation of activity and localization of the WNK1 protein kinase by hyperosmotic stress. Journal of Cell Biology, 176(1):89–??, January 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/176/1/89.

Zhang:2007:TMS

[ZRBS07] Dong Zhang, Gregory C. Rogers, Daniel W. Buster, and David J. Sharp. Three microtubule severing enzymes contribute to the “Pacman-flux” machinery that moves chromosomes. Journal of Cell Biology, 177(2):231–??, April 2007. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/177/2/231.

Zhou:2009:ZDI

[ZRH^{+}09] Kang Zhou, Melissa M. Rolls, David H. Hall, Christian J. Malone, and Wendy Hanna-Rose. A ZYG-12–dynein interaction at the nuclear envelope defines cytoskeletal architecture in the C. elegans gonad. Journal of Cell Biology, 186(2):229–??, July 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/186/2/
229.

Zhang:2008:SRL

[ZS08] Yu-Qian Zhang and Kevin D. Sarge. Sumoylation regulates lamin A function and is lost in lamin A mutants associated with familial cardiomyopathies. Journal of Cell Biology, 182(1):35– ??, July 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
182/1/35.

Zimniak:2009:PBY

[ZSMW09] Tomasz Zimniak, Katharina Stengl, Karl Mechtler, and Stefan Westermann. Phosphoregulation of the budding yeast EB1 homologue Bim1p by Aurora/Ipl1p. Journal of Cell Biology, 186 (3):379–??, August 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/186/3/379.

Zhang:2005:CRT

[ZTKK05] Jennifer Y. Zhang, Shiying Tao, Robin Kimmel, and Paul A. Khavari. CDK4 regulation by TNFR1 and JNK is required for NF-κB-mediated epidermal growth control. Journal of Cell Biology, 168(4):561–??, February 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/168/4/561.

Zonta:2008:GNI

[ZTM^{+}08] Barbara Zonta, Steven Tait, Shona Melrose, Heather Anderson, Sheila Harroch, Jennifer Higginson, Diane L. Sherman, and Peter J. Brophy. Glial and neuronal isoforms of Neurofascin have distinct roles in the assembly of nodes of Ranvier in the central nervous system. Journal of Cell Biology, 181(7):1169–??, June 2008. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/181/7/
1169.

Zhang:2009:RIL

[ZTZ^{+}09] Yongjun Zhang, Yizeng Tu, Jianping Zhao, Ka Chen, and Chuanyue Wu. Reversion-induced LIM interaction with Src reveals a novel Src inactivation cycle. Journal of Cell Biology, 184(6):785–??, March 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/184/6/785.

Zhang:2005:LGL

[ZWG^{+}05] Xiaoyu Zhang, Puyue Wang, Akanksha Gangar, Jian Zhang, Patrick Brennwald, Daniel TerBush, and Wei Guo. Lethal giant larvae proteins interact with the exocyst complex and are involved in polarized exocytosis. Journal of Cell Biology, 170 (2):273–??, July 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/170/2/273.

Zemp:2009:DCM

[ZWO^{+}09] Ivo Zemp, Thomas Wild, Marie-Françoise O’Donohue, Franziska Wandrey, Barbara Widmann, Pierre-Emmanuel Gleizes, and Ulrike Kutay. Distinct cytoplasmic maturation steps of 40S ribosomal subunit precursors require hRio2. Journal of Cell Biology, 185(7):1167–??, June 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/185/7/1167.

Zeng:2006:HCS

[ZXM^{+}06] Lingfang Zeng, Qingzhong Xiao, Andriana Margariti, Zhongyi Zhang, Anna Zampetaki, Seema Patel, Maurizio C. Capogrossi, Yanhua Hu, and Qingbo Xu. HDAC3 is crucial in shearand VEGF-induced stem cell differentiation toward endothelial cells. Journal of Cell Biology, 174(7):1059–??, September 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/174/7/